A Complete Bibliography of IEEE/ACM Transactions on Audio, Speech, and Language Processing

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 April 2024 Version 1.17

Title word cross-reference

1 [BSR^{+}23, DG23, GCJ^{+}14]. 1000 [RGK24]. 2 [MASC23, ZHCB23]. 2.5 [WSFS19, WAS16]. 3 [BWH18, BSR^{+}23, CT14, DGMB21, ECH23, GDM18, HB17, SMH14, YLLC19, ZHCB23, vMM14]. ^{2} [WSJ^{+}23]. ^{3} [LCRY24]. α [VZC14]. β [FSO17]. ℓ_{1} [FWF17]. ℓ_{p} [HCBY18]. F 0 [ZWW23a]. F_{0} [LTE^{+}21, WTY^{+}20b]. I [CL14a, KWGW17, MPC16, MZM15, SLK17, BHS^{+}16, CPL14, GH17, KMAB18, PS17, RSR^{+}19, XYH14, ZSB17]. M [YHC^{+}20]. N [MHBT15]. O(n^{3}) [LZZP22]. Q [ZYW^{+}24]. R^{2} [LQW^{+}24]. t [CS19, LBR18]. U [JLKK21]. X [CZN^{+}24, ZLX^{+}24].

-D [CT14, HB17, WAS16]. -Divergence-Based [FSO17]. -Estimate [YHC^{+}20]. -Gram [MHBT15]. -Hour [RGK24]. -integration [VZC14]. -Net [JLKK21]. -Network [ZYW^{+}24]. -Noise-Robust [LTE^{+}21]. -Norm [GCJ^{+}14, HCBY18]. -Optimal [FWF17]. -Source [CS19, LBR18]. -Vector [BHS^{+}16, GH17, KMAB18, XYH14, ZSB17, CPL14, MPC16, CL14a]. -Vector-Based [SLK17]. -Vectors [PS17, KWGW17, MZM15].

/r/ [SGB22].

14 [FWM^{+}14a]. 19 [ZTM^{+}23].

2 [MZC^{+}24]. 2.0 [LLC^{+}23b]. 2.0-Based [FSN24]. 2.5D [FFSS17]. 2017 [MDE^{+}19].

2019 [PMA^{+}21]. 2021 [Ano21a, LWS^{+}23]. 29 [Ano21a].

3D [OGZ18, ZHCB21].

4 [CRH^{+}14].

A2SA [LCZ23]. AAC [CRH^{+}14]. Ability [HWW23b]. Abstractive [CSS23, LPM^{+}23, ZZZ16]. Abuse [KAA^{+}24]. Accelerating [LKM23]. Acceleration [LYH^{+}22, MYC^{+}23]. Accelerometer [MVH16]. Accent [BHS^{+}16, CAFFR23, SGY^{+}24, ZGO19, ZDGO21, GH23a]. Accent/ [GH23a]. Accented [ZZZ^{+}24]. Accents [ZWTL23]. Access [Ano14-82, Ano14-83, Ano14-84, Ano14-85, Ano14-86, Ano14-87, Ano14-88, Ano14-89, Ano15-71, Ano15-72, Ano15-73, Ano15-74, Ano17-70, Ano17-71, Ano17-72, Ano17-73, Ano17-74, Ano17-75, Ano17-76, Ano17-77, Ano18-43, Ano18-67, Ano18-68, KYL24, Ano14-90, Ano14-91]. Accompanied [CWJ16]. Accompaniment [RLS^{+}18, WC16b]. Accordance [WC16b]. Accumulated [SIUW21]. Accuracy [FSN24, GBF21, HB17, OHN16, WTQ^{+}19]. Accurate [MBM15]. ACE [EGMN16]. ACM [Ano14-45, Ano14-77, Ano14-54, Ano14-66, Ano14-78, Ano14-55, Ano14-67, Ano14-79, Ano14-56, Ano14-68, Ano14-80, Ano14-57, Ano14-69, Ano14-46, Ano14-58, Ano14-70, Ano14-47, Ano14-59, Ano14-71, Ano14-48, Ano14-60, Ano14-72, Ano14-49, Ano14-61, Ano14-73, Ano14-50, Ano14-62, Ano14-74, Ano14-51, Ano14-63, Ano14-75, Ano14-52, Ano14-64, Ano14-76, Ano14-53, Ano14-65, Ano14a, Ano15-40, Ano15-50, Ano15-41, Ano15-51, Ano15-42, Ano15-52, Ano15-43, Ano15-53, Ano15-44, Ano15-54, Ano15-45, Ano15-55, Ano15-46, Ano15-56, Ano15-47, Ano15-57, Ano15-48, Ano15-58, Ano15-49, Ano15-59, Ano16k, Ano16l, Ano16-32, Ano16-38, Ano16-33, Ano16-39, Ano16-29, Ano16-30, Ano16-31, Ano16-34, Ano16-36, Ano16-40, Ano16-35, Ano16-37, Ano16-41, Ano17-59, Ano17-60, Ano21a, Ano16i, Ano16j, Ano17r, Ano17z, Ano17s, Ano17-27, Ano17t, Ano17u, Ano17v]. ACM [Ano17w, Ano17x, Ano17y, Ano18v, Ano18w, Ano18x, Ano18y, Ano18z, Ano18-27, Ano18-28, Ano18-29, Ano18-30, Ano18-31, Ano18a, Ano19a]. Acoustic [AB18, ATS18, AA17, AM18a, AHH^{+}20, ADM^{+}19, BBM17, BSR17, BH19, BSER17, BSR^{+}23, BMVG17, CATS20, CTSC14, CH16, CG17, CWJ15, Chi15b, CWJ16, CG21, CC15, CTLL17, CL18a, CPSFC^{+}14, CSAR^{+}14, ÇH21, CGK15, CZK^{+}21, DMHZ22, DLY^{+}19, DCAS22, DAMvW15, DKO^{+}18, DH20, DLMM^{+}18, DG15, DMP^{+}18, DKDN20, EGMN16, Emu18, EV23, EN18, FZZ^{+}15, FZM24, FdDPG21, GTS^{+}19, GDC19, GWLL22, GR14, GCvWMJ14, GHAP17, GPF17, GYdL^{+}21, Hac14, HH14, HBK24, HHZ^{+}21, HF15, HKM^{+}24, HCY21a, HZBW23, HSLB23, HGAPB15, HHC16, IO17, IO19, JNP^{+}21, JJB16, JGSC^{+}22, KJG16, KMG21, KLK17, KWGW17, KKGK18, KD16, KD17, KNH^{+}17, KSHH18, KBT19, LLR^{+}16, yLsL14, LFO^{+}24, LAK16, LWZG21, LH23, LSH^{+}23, LZAY23, LK16, LYB^{+}23, LHJ^{+}23, MXH21, MKA^{+}14, MAB^{+}16, MVH16, MHB^{+}18, MZM15, NW15, NDV18, OCS21, OBP^{+}24, PC19, PMMM15, RJCW17]. Acoustic [RLPK15, RvWM22, SV18, SDF18, SS16, SHAGPRZ17, SD16b, SD16a, SK16a, SGA15, SMA^{+}17, SLGA15, SSSS22, SHDG16, SMSV17, SBG17, SSS18, SR16b, SLR16, TD16, TQY16, TALP24, TK14, VH19, VAH19, VHBS21, VRN18, WLL^{+}15a, WCY19, WTW20, WDW^{+}23, WST^{+}21, WSY14, WRK18, XGKM18, YK17, YZCW21, YC22, YsL15, ZZH16, ZHH18, ZW19, ZHH19, ZLL^{+}19, ZDD21, ZTDD23a, ZLLZ23, ZYG^{+}14, ZGZ23, ZH15, ZLD21, ZZX^{+}23b, vMM14]. Acoustic- [MXH21].

Acoustic-Articulatory [GHAP17, HGAPB15]. Acoustic-Phonetic [CWJ15, CWJ16]. Acoustic-to-Articulatory [JJB16, SSSS22]. Acoustic-to-Word [DLY^{+}19]. Acoustical [LLLL19, YK21b]. Acoustics [AGV24, AGD^{+}22, BHBS16, BS14, CMM^{+}23, DHCS15, HB17, KE23, SRH15, SLH17]. Acquisition [PCW15, RV17, TH16]. Acronyms [MK15a]. Across [CTSC14, TALP24, YLD15]. Act [MHBT15]. Acted [OR16a]. Action [WBSG18]. Activation [CAFFR23, JLKK21, ZYG^{+}14]. Active [AA17, ATR^{+}21, AFdDG23, AKN^{+}16, BSB14, BJ24, BBJ18, CC14a, CE14, CD17, FdDPG21, GdC14, KBI^{+}21, LLZ^{+}23a, LV22, LC16b, LFdDG14, LP22, LSGH24, MX17, MZA20, MMS^{+}20, NNL15, PCG17, SGL^{+}21, SGLS23, SGL^{+}23, VE24, XCZ^{+}23, YPL21, YH17, ZAZ^{+}18, ZTQB19, ZZZ^{+}21, ZPW23, ZLX^{+}24, ZLSC23]. Activities [SBG17]. Activity [BE17, CE15, CC14b, DWX^{+}21, DTC15, DTC16, DTC17, DB17, DSH18, GG18, KM23, MLMT24, McL14, TON^{+}22, YLY15, ZW16a]. ACTUAL [ZYD^{+}23]. Actuator [RvWM22]. ACVAE [KKTH19]. ACVAE-VC [KKTH19]. Ad [TJCB16, WC18b]. Ad-hoc [WC18b]. Adaptable [NNC21]. Adaptation [ARSLN16, BDV^{+}14, BE17, CWYC17, CQ22, CLL^{+}21b, Chi18a, DHD^{+}19, DKO^{+}18, DXW^{+}23, DBM15, FCG24, GR14, HBK24, HMO21, KWGW17, LCWC16, LMC18, LM23a, MK15a, MH18, MSD14, NXCL16, PKR^{+}24, RLPK15, SML^{+}24, SS16, SDK17, SHJT21, SSAS21, SZL19, SR16b, SLR16, WUF^{+}18, WH22, WWG^{+}23, WSZN22, XLLW21, XYH14, XSC^{+}23, XAHJ^{+}14]. Adapted [DAGS^{+}18, SW20]. Adapters [GZX^{+}21, HZW^{+}22]. Adaption [GHJ^{+}23, ZGL^{+}23]. Adaptive [AM16, BBC^{+}16, BvWWM17, BvWM18, CE15, CBS23, CD17, CWR^{+}22, DKO^{+}18, DH20, DWX^{+}16, DGCY23, EV23, FWM^{+}14a, FWM^{+}14b, FNA15, FdDARF16, GCvWMJ14, GZX^{+}21, GYdL^{+}21, HCBY18, HDLL23, HBCC22, HQTX23, HLZ^{+}23, HGAPB15, KKJN17, KWZ24, LCLM18, LRG21, LWLL24, LC16b, LG16, LFdDG14, MLL^{+}22, MMCF18, MSB18, MZM15, MDM15, MZL21a, NTM16, NW15, PPSG17, QCN^{+}21, RG15b, RG17, RLK14, SGL^{+}21, SP15, TM16, TZA14, TQY16, TQH^{+}18, TNS^{+}18, VE24, WBLC22, WYHL20, WXL^{+}21, YK14, YEY17, Yan22, YPL21, YHC^{+}20, ZYL^{+}21, ZGZ23, ZDZ^{+}21, ZEM18]. aDCF [MMR^{+}22]. Additive [HAGA21, HCJT15]. Addressable [QVZY24]. Addressing [LLW21]. Adequacy [BDL15]. Adjustable [WW15]. Adjusted [SAB16]. Adjustment [CWH^{+}16]. ADL [ZXY^{+}21]. ADL-MVDR [ZXY^{+}21]. Admittance [MSS17]. Advanced [GWL^{+}24]. Advancements [LH14]. Advances [RPK23, XXWY24]. Advancing [DLY^{+}19, Ano14-115, Ano14-116]. Adversarial [AB18, BCCZ19, BGD^{+}18, CWZ^{+}23, CHL^{+}23, DLL^{+}21, HRL^{+}21, JGD^{+}24, LY21, LFQL19, LLCT22, rLcHcC^{+}22, LSG^{+}23b, LCTA19, MV23, NLYK24, PKP^{+}21, STS18, SWH22, TZL^{+}23, TL19, WWG^{+}23, WYZ^{+}23, WLL^{+}22, XHRC24, YTWB19, YTF^{+}23, ZMZ24]. Adverse [LC24, VMOL16]. AdvExpander [SWH22]. Aerial [YLH23]. AF [CBS23]. Affective [HHZ^{+}21, HSWC21, WYLZ16, WC16b]. Affects [WTQ^{+}19]. Affine [CZN^{+}24, FdDPG21, NZL23, ZN14]. African [HK24]. African-American [HK24]. Against [CH18, DLL^{+}21, TK14]. AGC [YLD15]. Agent [LWCC21, YYZY24]. AgentGraph [CCT^{+}19]. Agglutinative [LHZ^{+}24]. Aggregating [TM22]. Aggregation [CMY^{+}21, GCG21, GGZ23, LWJ19, YHL^{+}23]. Aging [KH16]. Agnostic [LM23a, SKT21, ZTL21]. Agreement [PCB16b]. AHP [AATKS21]. Aid [FPTJ17, MWL22, SSFP18, SP15, WW15, YLJ16, KA16]. Aided [MFG15, QMS16, hShWsL15]. Aids [AHH^{+}19, ABKP19, FWM^{+}14a, FWM^{+}14b, GvdP23, IC19, KNBC19, KPH15, MHD15, MHGD15, NNL15, PPSG17, SGM^{+}15, SD16b, SD16a, SNTD19, TNS^{+}18, YKPMP18, ZEM18]. Air [LWY^{+}23, SFK15, WCZR23]. Air- [SFK15]. Aircraft [VSL21]. Airflow [CWCP17, CG21, MYC^{+}23]. Alaryngeal [DTN^{+}14]. Algorithm [AAA19, CBA^{+}15, CC14a, CC15, CZN^{+}24, DC16, DG15, DPHG18, FdDPG21, GTS^{+}19, GLQ16, GB14, GJ22, HKI23, HLZ^{+}23, JT16, KBGAP^{+}16, LK19, LLZ^{+}23a, LKM23, LWLS24, NZL23, PC21, PCG17, RCR^{+}21, SOF15, SGH15a, SGH16, SGLS23, SRH15, Sug22, VDC^{+}17, WYHL20, YBC^{+}14, YK14, YND14, YHC^{+}20, ZZN^{+}16, ZZH16, ZLX^{+}24, ZLSC23]. Algorithmic [WWWR23]. Algorithms [BvWM18, BY15, CMG^{+}17, CRH^{+}14, EIPB^{+}19, FdDARF16, GaKGB21, HZF18, KBT19, LRG21, LFdDG14, MLMT24, MHGD15, NIK^{+}22, PACPJF^{+}23, PZZZ22, SG14, SGH17a, SMVW14, TR14a, ZP15, ZN14, ZGZ23]. Algorithms-Based [ZGZ23]. Aliasing [AR16, RKN14, WSFS19]. Aligned [APWB23, BHAMR21, ZB14]. Aligned-Parallel-Corpora [ZB14]. Aligning [BCA15]. Alignment [BHAMR19, CWJ16, DHD^{+}19, FCCA23, HS15, IK23, LSZ23, LCZ23, LZD^{+}23, NNS16, SFDRB21, TZS^{+}21, TXH^{+}23, TS16, WED16, ZLS^{+}21, ZKKC22]. Alignment-Based [FCCA23]. Alignment-Guided [ZKKC22]. Alignments [YCM^{+}21]. All-Frame-Level [AL23]. All-Neural [GZZY23]. All-Pole [SKH^{+}15a]. Alleviating [DCYY22, LLZ23b, ZLW^{+}22]. Alliance [XXS^{+}23]. Allocation [RV14]. Allophones [SGB22]. Allpass [VM15]. Also [WTQ^{+}19]. Alternating [BH18, ES16, LH23, WRKG17]. Alternative [HWW16, MHD^{+}17, ZGC^{+}23]. AM [VS14]. Amalgamation [RPL^{+}23]. Ambient [HTG14, HGT15a, HGT15b, LLS24]. Ambiguity [SSS18]. Ambiguous [ZLG^{+}22]. AmbiSep [HCH23]. Ambisonic [CM18, HCH23, HPPM24, KD24, SNK21]. Ambisonic-to-Ambisonic [HCH23]. Ambisonics [BHAMR21, BBHAR24, GLWQ22]. Amen [DM16]. American [HK24]. Amplitude [AKUS23, AL23, DAR16, FWF17, KTJ19, KBG16b, MAT15, MAK15, XWWW18]. Analog [BBFS21, VE24, YLD15]. Analysis [ATS18, AP21, ARSA14, AM16, ALH^{+}21, AZGN14, BDV^{+}14, BSB14, BBHAR24, CZG^{+}15, CH15, CCT21, CQ22, CHGS23, CZN^{+}24, CPL14, DV14a, DST^{+}19, DBLČ20, DMP^{+}18, DG23, ECJF21, GJYW23, GMD21, GLL^{+}21, Hac14, HMDG15, HDG16, HH14, HNT^{+}23, HZF18, HLZ^{+}23, HM14, IIY16, IO17, IO19, JMD^{+}17, JNP^{+}21, JYC^{+}24, JQL24, KH21, KOS^{+}16, KD18, KPA^{+}24, LHD16, LTE^{+}21, LFL^{+}23, LCRY24, LZD^{+}23, LCB^{+}23, MAZ21, MMT^{+}19, MLL^{+}22, MHGD15, MT18, MAL18, MTK^{+}20, MDM15, MGT15, MSEC21, NKS21, NIK^{+}22, NZL23, NCB23, NK22, OWA15, OBP^{+}24, PDS^{+}16, PCB15, PEO^{+}20, PM19, RG17, RVB^{+}17, SAP14, SIKN23, SGH17b, SB18, VM15, WLLZ17, WCY19, WYLZ20, WC16a, WK24b, YEY17, Yan22, YK21a, YWLZ18, YHC^{+}20, ZZH23, ZA14, ZGL^{+}20, ZDLK18, Ano16-52, Ano16-53, Ano16-54, Ano16-55, Ano16-56, Ano16-57]. Analysis-By-Synthesis [AP21]. Analytic [VS14, WAS16]. Analytical [LMGG21]. Analyzed [YHO^{+}17]. Analyzing [ARW^{+}16, MXH21, ZH19].

Anaphora [BZ21]. ANC [GLQ16, LP22]. Anchor [Lib21]. Anchor-Based [Lib21]. Anechoic [ZWW15]. Angle [PCB14]. Angular [WK24a]. Animations [CWCP17]. Anisotropic [TOKS14]. Annotation [MSW^{+}22, MHBT15]. Annotations [LCZ^{+}17, ZWCT15]. Annotator [MMM23]. Anomalous [KSU^{+}19, WK24a]. Anonymization [MWC^{+}23]. ANSD [HDLL23]. ANSD-MA-MSE [HDLL23]. ANSI [YLJ16]. Answer [CLC^{+}22, LCL^{+}21, TWZ^{+}18]. Answering [ACK21, KCL21, LWJ19, WQ24a, ZLZ^{+}24, ZLW^{+}24, ZXH^{+}16, ZBW^{+}24]. Anthropometric [GMFG16]. Anti [HWW23b, KHKY22, WDD^{+}16, ZLL^{+}23, ZXN^{+}21]. Anti-Collusion [ZXN^{+}21]. Anti-Interference [HWW23b]. Anti-Spoofing [KHKY22, WDD^{+}16, ZLL^{+}23]. Any [LCW^{+}21, XSC^{+}23]. Any-to-Any [XSC^{+}23]. Any-to-Many [LCW^{+}21]. APA [ZYG^{+}14]. Aperiodic [CT14]. APNet [AL23]. Apple [ZYZU22]. Application [AFG^{+}23, BATS21, CHC^{+}14, CZN^{+}24, GKC16, LLR^{+}16, LLdCF23, MAL18, OOK17, PDS^{+}16, PT21, PM19, RKL14, SAB16, SHD14, SMVW14, TR14a, TKH21, WZS14, YK21a, ZCDS15, ZZH16, ZGZ23]. Applications [Bor22, CD17, CC15, CTLL17, CL18a, FPTJ17, GH23a, KSG24, MS23, NEWN23, RG17, SSFP18, SRH15, SP15, TGGE21, WXL^{+}21, dSCB24]. Applied [AS17b, Ano17a, BSER17, Bre15, DM16, SB18, YHO^{+}18]. Approach [ATR^{+}21, ABA17, APH14, AY15, AKN^{+}16, BC17, BC18, BCA15, BvWWM17, BM15, Bor23, BMV^{+}23, BZ21, CPCBFSF16, CH18, CBN^{+}24, CH16, CZ17, CZY^{+}18, CK19, DWX^{+}21, DTDL16, DZF^{+}23, FGQ^{+}15, GMFG16, IISN21, ICB24, JT15, LGTG18, LS14a, LD16, LSDM19, LLZ^{+}21a, LLW23, LYY24, LCL^{+}21, LWY^{+}23, LW19, LPI^{+}21, MH16, MSD14, OR16b, OBP^{+}24, SD16a, SGF^{+}21, TXH^{+}23, TLW^{+}17, TMH^{+}14, WHRC16, WDDL17, WDDL18, WDW^{+}23, XWZ16, XYH14, XDDL15, ZA14, ZAZ^{+}18, ZHB^{+}18, ZJD^{+}15, ZSIA16]. Approaches [BHH22, BD16, FK19, FLMS16, RH18, RHvdP^{+}16, SCY22]. Approximate [SOF15]. Approximation [CY15, FK19, HCJT15, KE23, KHHJ19a, MBM15, SMVW14, VDC^{+}17]. Approximations [WAS14]. Arabic [AJK24, AAA19, BZ21, RARR15, ZB14]. Arbitrary [ABB^{+}23, KSHH18, NNS16, OIY^{+}14]. Architecture [CRVB^{+}15, CMY^{+}22, DBP^{+}24, HXC^{+}22, LCQ23, LFM^{+}23, PACPJF^{+}23, RARR15, WW22, YLD15, YLJ16, ZWW23a, ZLW^{+}24]. Architectures [DMSM21, SMA^{+}17, WDDL18, YH21]. Area [AP21]. Areas [SAP14]. Argument [GLT23, LZ16, LPL^{+}22, ZZZ16]. ARoBERT [WDW^{+}22]. Arousal [WYLZ16]. Arpeggios [BTBT18]. Arrange [ZMX^{+}21]. Array [ABB^{+}23, BJ24, BBJ18, CMM^{+}23, FSA18, GEC15, GS21, HSK17, HL14, HGYN23, HAGA21, IO17, JSR16, JEP14, LWZ^{+}14, LH23, LMH23, LAP23, MLP^{+}24, MFH21, NR14a, NEWN23, NHK16, OFF^{+}17, OBP^{+}24, OISO14, PCB15, PC19, SR14, SvWRK15, SPM17a, SPM17b, TP15, TCY15, TR14b, TR15, WYO23, WC16a, YZCW21, ZZL^{+}23]. Array-Based [BJ24]. Arrays [ABB^{+}23, AR16, BBAS19, BCB18, BABC19, CM18, CCT21, CT14, FFSS18, FSEI21, GvdP23, Hac14, HF15, HNF20, HCY21b, HBC17, HZF17, HCB18, HCB20, JHW^{+}21, LHJ^{+}23, MAZ21, MMS^{+}20, MXNB19, PCB16a, PCBS19, PC24, SSTS18, SCB21, SLC^{+}15, Son21, SKDN14, TH14, TH16, TJCB16, TC19, TC22, WD16, WCCB19, WBC^{+}21, WRKG17, ZA14, ZBC14, ZHCB21, ZHCB23]. Arrival [BKTR24, DGMB23, ED18, GWLL22, HMN17, HMO21, HAS21, KK14, NV14, SLC^{+}15, TP15, TTH14, TR15]. Arrivals [KB14]. Art [MSRND14]. Article [Ano19a, GHS^{+}21, Ano14-90, Ano14-91]. Articulation [DB17, LPI^{+}21, LRP^{+}23, MLC^{+}23]. Articulatory [CHWL14, CMG^{+}17, GBL24, GHAP17, GCG^{+}17, HGAPB15, JJB16, KCMW17, NRČR14, SSSS22, SLH17, SGB22, TKT17, TGGE21, WSY14, YMM19, YLLC19]. Artificial [AF18, AF19, AHF21, KBG^{+}23]. Artificially [AKG^{+}17]. ARX [LTE^{+}21]. ARX-LF [LTE^{+}21]. Ascending [BTBT18]. Ascending/Descending [BTBT18]. Asian [TGZ22]. Asking [KG19]. Aspect [CZG^{+}15, CQ22, JBWZ24, JYC^{+}24, LYL21, LFL^{+}23, LZL^{+}23, LZD^{+}23, LLL^{+}19, PFLX24, WFR^{+}21, YJX19, ZLC^{+}24]. Aspect-Based [CZG^{+}15, JBWZ24, JYC^{+}24, LFL^{+}23, LZD^{+}23, ZLC^{+}24]. Aspects [FBE19]. Aspiratory [CWCP17]. ASR [CMY^{+}22, DAGS^{+}18, DCYY22, GWG^{+}14, HSPK18, HIA^{+}17, HCZC24, LC24, LAK16, LLLL19, SV18, SZW23, TZS^{+}21, TVJ15, WDW^{+}22, WQ24b, WSY14, WSZN22, YCM^{+}21, YCL22, dlCSVAGAPM15]. Assessing [CDLL21, GAM23]. Assessment [BNvdB^{+}20, FBE19, KKK15, KNH^{+}18, LLLL19, LPI^{+}21, LRP^{+}23, PFLX24, WZC23, ZFC^{+}23, ZDZ^{+}23]. Assignment [DCZ19]. Assisted [BAS20, CSY^{+}21, WsL15]. Assistive [HdHTJ22, ZZ22]. Association [AM18a, DCZ19, SSS18]. Associative [OOK17]. AST [JSG^{+}22]. ASVspoof [LWS^{+}23]. Asymmetric [AHH^{+}19, BCB18]. Asymmetrical [YXX^{+}22]. Asynchronous [CBA^{+}15, CZK^{+}21, HZBW23]. ATF [ZSC^{+}23b]. Atomic [WSM19]. Attack [LLCT22, MKMDM17, NLYK24, WYZ^{+}23, YDZ19]. Attacks [TZL^{+}23, XNG^{+}14, YLQZ24, ZZX^{+}21, ZZX^{+}23a]. Attending [LLT^{+}21a, PW22]. Attention [AGB^{+}24, AGC23, BLZ21, BB21, CHWQ24, CMY^{+}22, CLHK21, CLZ22, DLY^{+}19, FLM23, GQXZ19, HFL^{+}23, IK23, JHLK18, KSK21, LKC^{+}21, LC24, LLT^{+}21a, LWL^{+}21, LPM^{+}23, LSWN23, LLHW24, LLC^{+}24, LLWQ24, LYL21, LSG^{+}23b, LHYJ23, LFM^{+}23, MYT^{+}19, MYL^{+}21, NW24, ODNA23, PW21, PLQZ23, QWW^{+}23, RGK24, SWC20, SRC^{+}23, TM22, WPY^{+}21, WZG^{+}23, WZH22, WLL^{+}23, XH23, XLL^{+}19, XZL^{+}21, XLZ^{+}23, XZL^{+}23, XRWL21, YCM^{+}21, YJX20, ZZLZ19, ZQN^{+}23, ZLZ^{+}24, ZDZ^{+}21, ZZX^{+}23b]. Attention-Based [CHWQ24, CMY^{+}22, IK23, LPM^{+}23, LLHW24, NW24, SWC20, XLL^{+}19, YCM^{+}21]. Attention-Copy [AGB^{+}24]. Attentional [SZX^{+}18]. Attentions [QZLH21]. Attentive [DZWC20, GLT23, LWL^{+}19, LdLvWWS21, LCQ23, PW23, ZPW23, ZM23]. Attractor [LCM18]. Attractors [HFW^{+}22, HWG^{+}23]. Attribute [HCC^{+}23, LWCL19]. Attributes [ASX22, BHS^{+}16]. Audibility [LMV21]. Audible [McL14]. Audio [AZK15, AGM23, AKB^{+}16, AM16, Ano14-45, Ano14-77, Ano14-54, Ano14-66, Ano14-78, Ano14-55, Ano14-67, Ano14-79, Ano14-56, Ano14-68, Ano14-80, Ano14-57, Ano14-69, Ano14-46, Ano14-58, Ano14-70, Ano14-47, Ano14-59, Ano14-71, Ano14-48, Ano14-60, Ano14-72, Ano14-49, Ano14-61, Ano14-73, Ano14-50, Ano14-62, Ano14-74, Ano14-51, Ano14-63, Ano14-75, Ano14-52, Ano14-64, Ano14-76, Ano14-53, Ano14-65, Ano14a, Ano15-40, Ano15-50, Ano15-41, Ano15-51, Ano15-42, Ano15-52, Ano15-43, Ano15-53, Ano15-44, Ano15-54, Ano15-45, Ano15-55, Ano15-46, Ano15-56, Ano15-47, Ano15-57, Ano15-48, Ano15-58, Ano15-49, Ano15-59, Ano16k, Ano16l, Ano16-32, Ano16-38, Ano16-33, Ano16-39, Ano16-29, Ano16-30, Ano16-31, Ano16-34, Ano16-36, Ano16-40, Ano16-35, Ano16-37, Ano16-41, Ano16i, Ano16j, Ano17-59, Ano17-60, Ano17r, Ano17z, Ano17s, Ano17-27]. Audio [Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17-57, Ano18v, Ano18w, Ano18x, Ano18y, Ano18z, Ano18-27, Ano18-28, Ano18-29, Ano18-30, Ano18-31, Ano18a, Ano19a, Ano19c, Ano19t, Ano21a, AB14, AB15, BF18, BV16, BCA15, Bay15, BBFS21, BMV^{+}23, BS14, Bre15, CWL22, CHJ14, CHL^{+}19, CLW24, CWJ16, CWT22, DHSG15, DTC15, DVLH17, DSH18, DDNN19, EDO17, FSO17, FM15, FLB16, FBE19, GEC15, GaKGB21, GdBBS21, GCG21, GBF21, GRSP17, HF21, HRL^{+}21, HTG14, HGT15a, HGT15b, HPPM24, HWRC15, HGT15c, IBN^{+}18, IISN21, JVR16, JYB^{+}15, KYL24, KA16, KSH17, KML18, KLCL24, KBGAP^{+}16, KEzW21, KGG19, KCL21, LG19, LM17, LBR16, LBR18, LLZ^{+}21c, LDG^{+}23, LSL24, LLWQ24, LCZ23, LB16, rLcHcC^{+}22, LDZ19, LLS24, MBD18, MV19, MMT^{+}19, MPHM19, MMCF18, MP16, MSRND14, MF17, MCF21]. Audio [MT18, MMMDL^{+}14, MTZ^{+}21, MPC^{+}24, MBTB14, MK16, NNS16, NKS21, NY21, NR14b, NXH^{+}17, NM21, NCC15, NK22, NTO^{+}23, NLV16, OVN^{+}22, ODG16, PKA14, PEO^{+}20, PPM^{+}24, PGS^{+}22, PCW15, PHMB18, PHM^{+}16, PHM^{+}17, QCW21, QWW^{+}23, RG15a, Rak17, ROL^{+}15, SUK24, SSGL21, SW16, SLOVB15, SB18, SLWY20, Sug22, SKH^{+}15a, SKH^{+}15b, SKM19, SRT23, TZZ^{+}15a, TKH21, VML14, WPY^{+}21, WGLZ21, Wid18, Wid19, WL19, XNG^{+}14, XNRG15, XNP^{+}18, XV21, XO16, XHW^{+}17, XXWY24, YZW^{+}21, ZM21, ZLO^{+}22, ZSH20, ZYD^{+}23, ZAL24, ZDZ^{+}21, ZLL^{+}24a, ZCQ^{+}19, ZXN^{+}21, ZTC23, vHK14]. Audio-Aware [KCL21]. Audio-Based [ROL^{+}15, WPY^{+}21]. Audio-Lyrics [ZLL^{+}24a]. Audio-to-Score [LCZ23, NNS16]. Audio-Visual [AZK15, CWT22, DTC15, LLZ^{+}21c, LDG^{+}23, LLWQ24, MTZ^{+}21, OVN^{+}22, PEO^{+}20, QCW21, QWW^{+}23, YZW^{+}21, ZAL24, ZDZ^{+}21]. Audiobook [CBG15]. AudioLM [BMV^{+}23]. Audioradio [OWW^{+}23]. Audiovisual [Abd18, TB18]. Audition [MAZS22, TR14b, YLH23]. Auditory [CE15, CHWL14, DSMD15, GA20, HHZ^{+}21, JZJØ19, KGH^{+}16, MAK15, SP16, SMvdP22, TMKO17, WWS^{+}18, XGKM18, XRWL21]. Auditory-Inspired [XGKM18]. Auditory-Receptive-Field [WWS^{+}18]. Aug [FWM^{+}14a]. Augmentation [CHGS23, CGK15, DZZ23, HMO21, JCYwH21, JGD^{+}24, LGL^{+}22, LLZ23b, ÖM24, WDW^{+}23]. Augmented [CWDA17, HMN17, LLZ^{+}23a, LLWJ16, NNC19, RG15b, XWZ16, ZBW^{+}24]. Auralization [GS21]. Authenticator [MUK22]. Authors [Ano14-66, Ano14-67, Ano14-68, Ano14-57, Ano14-58, Ano14-59, Ano14-60, Ano14-61, Ano14-62, Ano14-63, Ano14-64, Ano14-65, Ano15-60, Ano15-61, Ano15-62, Ano15-63, Ano15-64, Ano15-65, Ano15-66, Ano15-67, Ano15-68, Ano15-69, Ano16-44, Ano16-45, Ano16-36, Ano16-37, Ano16-42, Ano16-43, Ano16g, Ano16h, Ano17z, Ano17-27, Ano17-57, Ano17-58, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u]. Autistic [CPGL24]. Auto [HBB17, SZVZ16, XLZ^{+}23]. Auto-Encoder [XLZ^{+}23]. Auto-Regressive [HBB17]. Autocorrelation [AS17b, JFG^{+}17, PT14]. Autoencoder [BCCZ19, DZWC20, GJ17, KKTH19, KLCL24, LKM23, WTY^{+}20b, WLT24, XHRC24]. Autoencoder-Based [LKM23]. Autoencoders [CWBv19, DXZ^{+}18, HL18, JMD^{+}17, LMC18, MDCS20, PMB^{+}20]. Autoencoding [CHL^{+}19]. AutoGuitarTab [MFG15]. AutoMashUpper [DHYG14]. Automated [AATKS21]. Automatic [ACRS14, AZGN14, BVGV15, BK24b, BHS^{+}16, BD16, CJ15, CLW^{+}16, CSZZ24, CLSyL21, DHYG14, ESVK17, FCCA23, GLW20, HK24, HHZ^{+}21, HYI^{+}15, HXWS15, KSH17, KKK15, KG16, LKHU14, LDGHU14, LCZ^{+}19, LLC^{+}24, LS14b, LK16, LPI^{+}21, LRP^{+}23, MSFK15, MSRND14, MAK15, MKA16, NRČR14, OS21, O’S24, OR16a, RGR^{+}22, RARR15, RWF16, SWW^{+}17, SHJY21, SW20, SBM^{+}19, SSKP16, SBG17, SLW15, SKT21, TZL^{+}23, WW16, WTY^{+}23, WDS^{+}18, WL19, WZH22, ZH19, ZDZ^{+}23, ZGC^{+}23]. Automation [MTS^{+}19]. Autoregressive [BYT^{+}21a, CLZ^{+}24, FCCA23, GMH14, cHrLLyL23, LLC22, LSDM19, MIYO14, NCC15, PDKG24, WTY18, WTY^{+}20b, WHO^{+}21, WHT^{+}21, YCL22]. Auxiliary [AA17, KKTH19, SHJT21, ZCY16]. Availability [QCL23]. Averaging [CW18, WFN^{+}18]. Awards [Ano19a]. Aware [AGB^{+}24, ADLX23, BM21, Bor23, CMY^{+}21, CLW24, DZWC20, DFY^{+}24, DB18, DZZ23, HCQ24, HDLL23, KCL21, LWC^{+}18, LLZ^{+}21c, LSWN23, LHL^{+}22a, LHZ^{+}24, LCG^{+}24, LLZ^{+}23d, MZZ23, MV19, MXH21, MWT^{+}23, MTKS21, MZL21b, NDS^{+}23, QTY16, QCN^{+}21, STS21, SF18, SWC^{+}21, TWZ^{+}18, VDH24, WXS24, XZL^{+}23, XSL^{+}19, YLH23, ZXSD17, ZXSL19, ZL21a, ZZZ22a, ZLZ^{+}24, ZM17, ZZ19, ZTC23]. Awareness [GYXY23, LSLK24]. Away [ZYZU22]. Axial [HFL^{+}23]. Axial-Attention-Based [HFL^{+}23]. Azimuth [SCB21, XBZ^{+}24].

Back [ACRS14, LH14, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano15n, Ano15o]. Back-end [LH14]. Back-off [ACRS14]. Backend [GH17]. Background [CZ17, RDP14]. Backward [LFQL19, ZTWY19]. Bad [MIR^{+}21]. Bag [GBF21, GPF17, LCYW21]. Bag-of-Audio-Words [GBF21]. Bag-of-Features [GPF17]. Bag-of-Words [LCYW21]. BaNa [YBC^{+}14]. Band [AA17, KKJN17, LSWN23, LY23, WCC^{+}23]. Band-Dependent [KKJN17]. Band-Split [LY23]. Banded [MUT^{+}20]. Bandwidth [AKG^{+}17, AF18, AF19, BNvdB^{+}20, GDC19, HJZB23, LAGD18, LB16, MV23, Ziv15, ZPP15]. Bandwidth-Extended [AKG^{+}17]. Bank [AM16, MWL22, MAK15, SGM^{+}15, WW15, WXL^{+}21, YLJ16]. Banking [LWYC23]. Barriers [LLZ^{+}21b]. BART [CXvG^{+}23]. Base [FWF17, GJB15, LWJ19, WQ24a]. Based [AZK15, Abd18, AFG^{+}23, AF19, AGW15, AB21, AJLW14, AMK^{+}20, ADGE19, BWH18, BVGV15, BV17, BIK^{+}18, BA19, BV23, BCA15, Bay15, BGQO^{+}17, BHAMR19, BHAMR21, BKTR24, BJ24, BvWWM17, BG21, BH19, BM15, BSTM21, BSL15, BSL16, BH18, BD16, BG23, CWL22, CvdPHH21, CCTO21, CSJD16, CDLL19, CDLL21, CW18, CBN^{+}24, CZG^{+}15, CWW14, CHWL14, CJ15, CH15, CZY^{+}18, CQ22, CHWQ24, CMY^{+}22, CCC21, CWJ16, CMG^{+}17, CG21, CMD^{+}23, CY15, CC14b, CC15, CPSFC^{+}14, CSAR^{+}14, CG16, CL17b, DMHZ22, DMSM21, DvWDM14, DKO^{+}18, DW17, DMVP16, DXW^{+}23, DDMvW20b, DBLČ20, DFY^{+}24, DTN^{+}14, DRK18, DG15, DY22, DSH18, DDNN19, DG23, FSO17, FM15, FCCA23, FWM^{+}14a, FWM^{+}14b, FBWG23, FLM23, FLMS16, FNM^{+}19, FD21, FLB16, FSN24, FSEI21, FdDARF16, GDC19, GWQ23, GBL24, GY14, GaKGB21, GTV19, GG18, GR14, GCJ^{+}14]. Based [GvdP23, GAM23, GLL20, GGZ23, GXW^{+}23, HSPK18, HF21, HW14, HSM19, HSK17, HBK24, HTG14, HGT15b, HBF^{+}16, HBB17, HFL^{+}23, HS15, HIA^{+}17, HdHTJ22, HKHP23, HHO^{+}21, HFW^{+}22, HCKW23, HCY21a, HAS21, HCY21b, HZBW23, HGT15c, HZF17, HBCC22, HQTX23, HHMK18, IIY16, IK23, ISZA15, IBN^{+}18, IISN21, JP14, JFG^{+}17, JT14, JBWZ24, JYC^{+}24, JvWGD15, KGP14, KAA^{+}24, KLK17, KVMV^{+}24, KNBC19, KHK17, KDY14, KPH15, KK14, KKGK18, KSK21, KD21, KD16, KD18, KNH^{+}18, KSU^{+}19, KMG15, KGH^{+}16, KBI^{+}21, KBG16b, KGG19, LGTG16, LGTG18, LG20, LC16a, LTWT17, LK19, LNNL22, LRWG23, LGHG17, LGGH18, LCSL19, LGGH19a, LCZ^{+}19, LWL^{+}21, LWCC21, LTE^{+}21, LPL^{+}22, LPM^{+}23, LH23, LLZ^{+}23a, LHC^{+}23, LKM23, LJWC23, LLHW24, LHW^{+}21, Lib21, LMC18, Lin19, LFL^{+}23, LCRY24, LB16, LK16, LWCL19, rLcHcC^{+}22, LGL^{+}22, LWYC23, LSG^{+}23b, LFWZ23, LZD^{+}23, LSYN19, LSGH24, MSW^{+}22, MIYO14]. Based [MBD18, MK15a, MCL22, MHD15, MUK22, MP20, MLC^{+}23, May18, MTZ^{+}21, MC17, MPC^{+}24, MTKS21, MTK^{+}20, MMZ23, MTBB15, MBD^{+}15, MTS^{+}19, MK16, MVN22, NYS17, NKS21, NXH^{+}17, NW24, NLL^{+}18, NNL15, NCB23, NV14, NDS^{+}23, NAK^{+}17, ODNA23, OHN16, OCS21, OOK17, OVN^{+}22, OIY^{+}14, OD17, ODG16, ODSM16, OWW^{+}23, PS19, PBC18, PACPJF^{+}23, PW19, PLZL17, PCB16b, PLQZ23, PT14, PM15, PZP19, PHM^{+}17, PBQ15, QDSL19, QYCT23, QTY16, QLZ^{+}21, RV14, RDP14, RSR^{+}19, RPK23, RH18, RA23, RG18, RMBK14, RWL^{+}23, ROL^{+}15, SML^{+}24, SV17, SSFP18, SS16, SAC17, SDK17, STT^{+}19, SDDCC15, SSI21, SNTD19, SWH^{+}17, SHKB17, SGH17a, SKSD14, SMVW14, SHCV23, SK16b, SAS17, SLC^{+}15, SSGL21, SGL^{+}21, SGLS23, SGL^{+}23, SZS22, SBM^{+}19, SLK17, SSK17, SCW21, SW16, SKDN14, SBP^{+}24, SWC20, SHZ^{+}23]. Based [SYG^{+}24, SKT21, SMvdP22, TOKS14, TOS14, TM14, TM16, TQY16, TMM18, TZW21, TW21, TZS^{+}21, TH18, TON^{+}22, TTH14, TLW^{+}17, TQM^{+}23, TKT17, TDL19, UKS19, VML14, WFN^{+}18, WS14, WYH^{+}15, WYLZ16, WD16, WLL^{+}16, WDDL17, WYH17, WB18, WC18a, WW19, WCY19, WZW19, WPY^{+}21, WCC^{+}21, WYY22, WDW^{+}23, WTY^{+}23, WLT24, WWG^{+}24, WYY24, WSM19, WVCL14, WHLG14, WLL15b, WK16, WRKG17, WZ21, WCL^{+}21a, XH23, XNG^{+}14, XNRG15, XHRC24, XWWW18, XCZ^{+}23, XLL^{+}19, XLZ^{+}23, XZW15, XBZ^{+}24, XDDL15, XO16, XHW^{+}17, XYY^{+}23, XAHJ^{+}14, XMBN21, YK14, YZL^{+}18, YLPL19, YCM^{+}21, YC22, YK21a, YWT23, YLY15, YHO^{+}17, YHO^{+}18, YH21, YH17, YMM19, YJX19, YHC^{+}20, ZP15, ZCF14, ZSB17, ZWCT15, ZFC^{+}23, ZKA15, ZW16a, ZW17, ZH18, ZCHH18, ZHH18, ZKH18, ZGL^{+}20, ZLZW21, ZL21a, ZZ22, ZSC^{+}23b, ZQZ^{+}23, ZGO19, ZZX^{+}21, ZGZ23, ZZX^{+}23a, ZWS^{+}23, ZZ19, ZH15, ZJD^{+}15, ZJL21, ZLC^{+}24]. Based [ZCQ^{+}19, ZLSC23, ZB14, ZXN^{+}21, ZSIA16, dSCB24, Ano16-50, Ano16-51, TMH^{+}14, MYNO23, RG21, SSS18, ZM21]. Bases [BK24b]. Basis [KHTL18, PSL17, VDC^{+}17]. Bass [AS17a, MGT15]. Bayes [CY15]. Bayesian [AV16, BIK^{+}18, Bor22, CNFG23, CH18, CS19, CY16, Chi18b, DBLČ20, ET18, HAGSZ16, HXL^{+}21, IBN^{+}18, LG19, LHD16, LSGH24, MIYO14, OISO14, OIY^{+}14, SEM^{+}14, SCW21, VMOL16, WAS14, WSH15, XLLW21, YSCZ15, ZDC14, ZM23]. Beam [Son21]. Beamformed [LWZ^{+}14]. Beamformer [AVM19, ABKP19, CBHC21, CMM^{+}23, GMD21, HDG16, HIA^{+}17, JTHN14, JSR16, KSHH18, LYN14, PCB14, SNTD19, SGH17c, ZH14, ZCHH18, ZZ22, ZEM18]. Beamformer-Based [SNTD19]. Beamformers [BC17, BC18, BCB15, DMVP16, HMDG15, IBC21, LHJ^{+}23, Non14, YOM21, YBHC21, ZDD21, ZHCB21]. Beamforming [ABB^{+}23, AR16, ATP15, BCB15, BABC19, CP21, CBC19, CT14, GZZY23, HL14, HCB18, JSR16, JBCH23, KHHJ19a, MGGK17, MYK^{+}24, NIK^{+}22, ODNA23, PCB15, PCB16a, PC24, SDF18, SR14, SBM^{+}19, WCCB19, WBC^{+}21, WST^{+}21, YG24, ZPØ^{+}21, ZBCH21]. Beampattern [ZBCH21]. Beampatterns [HBC17, PCBS19]. Beamwidth [PC24]. Beamwidths [SCB21]. Beat [CMD^{+}23, DLH23, ZDG14]. Bee [TON^{+}22]. Beginners [WPY^{+}21]. BEHM [MV23]. BEHM-GAN [MV23]. Belief [SHD14, WWS^{+}18]. Benefit [CHL^{+}24]. Bernoulli [Lin19]. BERT [BYT^{+}21a, CCL^{+}21, GZX^{+}21, MWT^{+}23, MVN22, QLZ^{+}21, WCZ21]. Better [AAA19, LLZ^{+}21a, LLJ^{+}23, YKL^{+}23]. Between [AJB^{+}18, BTBT18, KTJ19, LHC^{+}23, MVH16, PBA^{+}23, WTW20, XLG^{+}23, YLLC19, ZQZ^{+}23]. Beyond [sLGyLaC15, LLWQ24, LCB^{+}23, WTQ^{+}19, XXWY24]. BFGAN [LFQL19]. Bi [BV23, DZWC20, LWL^{+}21]. Bi-LSTM [DZWC20, LWL^{+}21]. Bi-Phase [BV23]. Bias [FPTJ18, FLM23, LLZ23b, WSZN21]. Bias-Compensated [FPTJ18]. Biasing [SZW23]. Bidirectional [CH16, DAGS^{+}18, JGSC^{+}22, LLC^{+}24, LLL^{+}19, XZL^{+}21]. Bilateral [BHAMR21, BBHAR24, GH23b, KM21]. Bilevel [YHL^{+}23]. Bilinear [ZDZ^{+}21]. Bilingual [BCCZ19, CLZ^{+}23, LFU^{+}17, WZL^{+}15, YsL15]. BiLSTM [LPM^{+}23, LKR23]. Bin [YOM21]. Binary [ATS18, BWZC22, DAFS15, DVLH17, HL18, RBBG15, WWS^{+}18, YK17]. Binaural [AJLW14, AHH^{+}19, AdHTJ16, ABKP19, BHAMR21, BOL^{+}18, BATS21, DHSG15, FBE19, GN23, GvdP23, GMD21, GS21, HMDG15, HDG16, HF21, HAGA21, IC19, JSR16, JWLF14, KNBC19, KHHJ19a, KGBPM24, MMB17, MGB18, MHD15, MHGD15, MD18, May18, PLZL17, RJW19, SR14, SvWRK15, SKM17, SBVM16, TE23, TMKO17, VDH24, VS14, WSM19, YLL21, YND14, ZW17, ZL21a, ZZ22, ZEM18, ZM20]. Binaural-Projection [TE23]. Binaurally [YKPMP18]. Bins [GWLL22]. Bio [ARW^{+}16, JGM23]. Bio-Inspired [ARW^{+}16, JGM23]. Biologically [dlCSVAGAPM15]. Biomedical [SUK24, ZQF^{+}24]. Biomimetic [YLD15]. BioPRO [ZQF^{+}24]. Biosignal [Ano16-50, Ano16-51, SWH^{+}17, SHKB17]. Biosignal-Based [SWH^{+}17, SHKB17, Ano16-50, Ano16-51]. Bipartite [BWZC22]. Bird [JK19, KMG17, SBG17]. Bit [CLAG16, LJK17, SKH^{+}15b, XYH^{+}21]. Bitrate [BF18]. Bivariate [ZDLK18]. Black [TZL^{+}23]. Black-Box [TZL^{+}23]. Blank [Ano14d, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano16c, Ano16d, Ano16e, Ano16f, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14b, Ano14c, Ano14e, Ano14f, Ano14g, Ano14h, Ano15n, Ano15o, Ano15a, Ano16a, Ano16b, Ano19b]. Blind [BBM17, BV16, CONVVC18, CG17, CG20, CTE21, CK19, DB19, FK19, HCBY18, ISP19, IISN21, KOS^{+}16, KD24, MTKS21, MTK^{+}20, NIK^{+}22, NV14, OOK17, OVN21, SH14a, SOF15, SIKN23, SH14b, SGH17a, TH17, TH18, UNI^{+}24, WC18b, WW19, WGLZ21, WYY22, WRDP17, YG24]. Block [BvWWM17, FdDARF16, LG16, Yan22, YH21]. Block-Based [FdDARF16, YH21]. Block-Sparse [LG16]. Blocking [WGD15]. BLSTM [LCSL19, WL19]. BLSTM-Based [LCSL19]. BLSTM-CRF [WL19]. Board [MAZS22]. Body [HJZB23, MSS17]. Body-Conducted [HJZB23]. Boltzmann [NTA15, NTM16, NTY19, NY21]. Bone [LYY24, SFK15, WCZR23, ZH19]. Bone-Conducted [LYY24]. Bone-Conduction [SFK15]. Boosting [GWCT23, ZW16a, ZCW^{+}24]. Bootstrap [CW18]. Both [EMTF18, LSL24]. Bottleneck [CCC^{+}24, DMSM21, JMD^{+}17, LNNL22, STT^{+}19, WK16]. Bound [CCC21, SGH17b, WCY19]. Boundaries [ZXZZ23]. Boundary [BHBS16, SSTS18, YTF^{+}23]. Bounded [RLPK15]. Bounds [TC22]. Box [TZL^{+}23]. BPTT [CH16]. Branch [CC21, SYG^{+}24, TYZ^{+}22, ZZX^{+}23b]. Branching [WTQ^{+}19]. Break [DM16]. Bridge [ACK21, MSS17]. Bridging [LLZ^{+}21c, WTW20]. Broad [LCY^{+}23]. Broadband [BC17, BC18, BCB15, CT14, MZA20, Non14, RMBK14]. Broadcast [CLC^{+}15, DHD^{+}19, PACPJF^{+}23, PBQ15, WHB^{+}21, XO16]. BSS [CY15, OG14, YK21a]. Build [ARSLN16, PCB16b]. Building [DOL^{+}23, LGL^{+}24, PM23, Zhe23]. BYOL [NTO^{+}23].

C50 [PSL^{+}16]. Cabin [AC24]. Calculation [MV14]. Calibration [Bor23, Cum21, CS23, FNM^{+}19, HCY21a, HCY21b, KH16, KH21, MR20, TMM18, WCY19, WSG21]. Calls [AMK^{+}20]. Can [CHL^{+}24, Ano15z, Ano15-27]. Canceler [YEY17, ZGL^{+}23]. Cancelers [Emu18]. Cancellation [AA17, AR16, BvWWM17, BvWM18, CS19, CSAR^{+}14, DST^{+}19, DDMvW20a, DLMM^{+}18, FWM^{+}14a, FWM^{+}14b, GCvWMJ14, GYdL^{+}21, HBK24, KJ21, LV22, LP22, MV14, MMK21, PC19, RvWM22, SNTD19, SK16a, SLGA15, SP15, ZZH16, ZLLZ23, ZYG^{+}14, ZGZ23]. Canceller [ABvM19, PPSG17]. Cancer [AFG^{+}23]. Candidate [CWH^{+}16, ZLJ22, ZHZ19]. Capabilities [CDA^{+}14]. Capacity [XNRG15]. Capsule [LFM^{+}23, WCL^{+}21b, ZWHD23]. Caption [ZYD^{+}23]. Captioning [XXWY24, ZYD^{+}23]. Capture [HJZB23]. Car [AC24]. Care [KAA^{+}24]. CASA [BA19, LW19]. CASA-Based [BA19]. Cascade [WW22, ZWW23a]. Cascaded [GHAP17, HGAPB15, LSG^{+}23b, NR14b]. Cascading [sLGyLaC15]. Case [BC17, GSA18, LP22, MBDV23, PKR^{+}24, WC16a]. Case-Enhanced [LP22]. Cases [GHS^{+}21]. Casting [OR16a]. Categories [Ano16k, Ano16l]. Causal [GS21, MYK^{+}24, WZZ^{+}24, Wid18, Wid19, ZLC^{+}24]. CausalABSC [ZLC^{+}24]. Causality [KJ21]. Causality-Constrained [KJ21]. Cause [ADLX23, CJY^{+}21, FYG^{+}21, HZL23, SCCM23, SHZ^{+}23]. CCi [GH23b]. CCi-MOBILE [GH23b]. CEGM [CDLL21]. Cellular [CZ17]. Center [AMK^{+}20, Ano16-48, Ano16-49]. Centered [AS17a]. Centric [ICB24]. Cepstral [DP16, EMTF17, EMTF18, JT15, KHTL18, KS16, SK18]. Cepstrum [IO17, YHO^{+}18]. CET2 [XZFN23]. CFDRN [LWYD23]. CGMM [ZSC^{+}23b]. CGMM-Based [ZSC^{+}23b]. Challenge [EGMN16, KSG^{+}21, KGG^{+}21, MDE^{+}19, YCC^{+}24, MHB^{+}18]. Challenges [SYKL21, TD16, XXWY24]. Chance [BC18]. Chance-Constrained [BC18]. Change [ALH^{+}21, PCB16b, SHJT21, YUK21]. Changes [MASC23]. Channel [AM18b, BXY14, BV23, CBN^{+}24, CDLX18, CS19, CH23, DST^{+}19, DB18, DBN^{+}17, DTDL16, FD21, GDC19, GRSP17, JvWGD15, KD16, KD17, KG14, LG19, LK19, LDG^{+}23, MXH21, MLMT24, MHD15, MHGD15, MF17, MR20, MK15b, MK15c, MBK17, PSL^{+}16, PZP19, PT21, QCL23, SGM^{+}15, SLOVB15, SKAK16, TZA14, TG23, TG24, WFN^{+}18, WWD14, WDDL17, WLT24, WYSD15, XGKM19, XBZ^{+}24, YC22, YND14, YZW^{+}21, ZTQB19, ZXY^{+}21, ZWZ^{+}21, ZLLZ23, ZTDD23b, ZZY^{+}24, ZPP15]. Channel-Aware [MXH21]. Channel-Combination [MLMT24]. Channels [YLD15]. Character [NNC20, WZ21, ZLJ22]. Character-Level [WZ21]. Characteristics [JG19]. Characterization [CPGL24, PHK16, TGGE21]. Characterizing [CTSC14]. Chasing [Sug22]. Chat [LSYN19]. Chat-Based [LSYN19]. Chatbots [GLL20, LXS^{+}18]. Check [NNC21]. Checking [ZZGD21]. Children [PLT^{+}18]. Chinese [Chi18a, CCL^{+}21, LZC^{+}14, LZ16, LXK^{+}21, NNC21, QLL^{+}21, QVZY24, WSY14, WZ21, XHS^{+}18, YLGG22, YCL22, ZLZW21, ZZH21, ZCY^{+}22]. Choice [KCL21, XPC^{+}22, ZZD23]. Chord [CB14, DDZS17, MSRND14, OS21, WL19]. Chroma [ZM21]. Chunk [CH16]. CI [SF14]. Circuits [BBFS21, GdBBS21]. Circulant [SGLS23]. Circular [BCB18, FLL^{+}23, HF15, HNF20, HBC17, HCB18, MAZ21, OFF^{+}17, OBP^{+}24, SCB21, SCW21, WAS16, ZA14, ZHCB21]. CircularE [FLL^{+}23]. Citation [DZWC20]. Cities [TALP24]. CL [LFL^{+}23]. CL-XABSA [LFL^{+}23]. Clarification [KG19]. Clarity [LCC^{+}15]. Class [CAFFR23, CSJD16, CJ15, DVO^{+}23, LCY^{+}23, SKT21, YBHC21]. Class-Based [CSJD16]. Class-Normalized [CJ15]. Classes [LWZZ16]. Classical [CMD^{+}23]. Classification [ATS18, AHF21, BLZ21, BSER17, BMVG17, CAFFR23, CWW14, CHJ14, CWR^{+}22, CSZZ24, DCYY22, DTL15, DR22, FLM23, GWCT23, GBF21, GPF17, HS15, HKM^{+}24, JWLF14, JBWZ24, KEzW21, KM23, LWL^{+}19, LLW23, LSG^{+}23a, LJX^{+}16, LYL21, LLWJ16, LFWZ23, LHG^{+}24, MYT^{+}19, MYL^{+}21, MMCF18, MZX^{+}15, MHB^{+}18, MSEC21, MZL21b, NM21, OHN16, OGZ18, PKA14, PPM^{+}24, PMMM15, PHM^{+}17, RG15a, Rak17, SK17, SHAGPRZ17, STBC17, SUK24, SLY14, SKT21, TALP24, TQW^{+}15, TON^{+}22, TP23, WAS14, WYH17, WJ23, XXS^{+}23, XLL^{+}19, XV21, YLC^{+}24, YK17, YJX20, ZW19, ZSH20, ZZY^{+}24, ZM17, ZWHD23, ZLC^{+}24, VZC14]. Classification-Based [CWW14, WYH17]. Classifier [KKTH19, SMA^{+}17, ZLG^{+}22]. Classifiers [PCB16b]. Classify [CJY^{+}22]. Clean [BPG23, HCZC24, MC17]. Clear [NHK16]. Cleft [MP20, VAP19]. Clinical [LXK^{+}21]. Clip [LSL24]. Clip-Level [LSL24]. CLOpinionMiner [ZWX15]. Closed [ARSA14, dCBC23]. Closed-Form [dCBC23]. Closure [DP16, KDY14, KKGH16, MG18]. Closure/Opening [KKGH16]. Clues [DVO^{+}23]. Cluster [HCQ24, LJWC23, TQY16]. Cluster-Aware-DINO [HCQ24]. Clustering [AB14, AB15, CCY^{+}24, DSH18, HPB15, KHTL18, KD21, KSK^{+}15, KSG24, LDZ19, ODG16, PKP^{+}21, SG21, SKDN14, WLL^{+}15a, XYH14, XO16, YLPL19, YH17, ZG21, ZJD^{+}15]. Clustering-Based [ODG16, SKDN14]. CNN [BSTM21, DG23, LWL^{+}21, MASC23, PW19, PW21, SGL^{+}23, VHBS21, WYLZ20, YH21, YDW^{+}21, ZQZH19]. CNN-Based [PW19, SGL^{+}23]. CNN-LSTM [WYLZ20]. CNNs [DGMB23]. Co [CBCL15, DHSG15, HL15, SLC^{+}15, WZH22, YJX19]. Co-Attention [WZH22]. Co-Extraction [YJX19]. Co-Factorization [HL15]. Co-Localization [DHSG15]. Co-occurrence [CBCL15]. Co-Prime [SLC^{+}15]. Cochannel [ZWW15]. Cochlear [CCC21, CFMM14, GH23b, KM21, NWM16, SF14, SMVW14, YLD15, dSCB24]. Cochleograms [dlCSVAGAPM15]. Cocktail [PWS^{+}23]. Code [AVK^{+}15, HLL15, JSG^{+}22, LWL^{+}21, WLLZ17, WSY14, WSH15, YsL15]. Code-Mixed [HLL15]. Code-Switched [YsL15]. Code-Switching [AVK^{+}15, WLLZ17, WSY14, WSH15].

Codebook [GXW^{+}23, HBB17, HCZC24, VML14, WSM19]. Codebook-Based [HBB17, VML14, WSM19]. Codec [LHM21, ZLO^{+}22]. Coded [XZF23, ZLF19]. Coder [SKH^{+}15a, SKH^{+}15b]. Codes [XAHJ^{+}14]. Coding [AHH^{+}19, BF18, CLAG16, CDW16, GH23b, HPPM24, JSYW16, JG19, JPX^{+}23, KM21, MK16, PZW^{+}21, SKH^{+}15a, SKM19, ZSL^{+}22]. Coefficient [ZZX^{+}21]. Coefficients [AS17b, DP16, GI15, ISZA15, KS16, KBG16b, MKA^{+}14, ZGL^{+}23]. Coefficients-Switched [ZGL^{+}23]. Cognition [LWYD23]. Cognition-Inspired [LWYD23]. Cognitive [VSL21]. Coherence [BBM17, FSA20, MHD15, MD18, RJW19, SGH16, XZW15]. Coherent [JC17, MLP^{+}24, SK15, XZFN23, ZYW^{+}24, ZCHW24]. Coherent-to-Diffuse [SK15]. Collaborative [CHGS23, MLL^{+}22, TLWV17]. Collabratec [Ano16-46, Ano16-47, Ano17-61, Ano17-62, Ano17-63, Ano17-64, Ano17-65, Ano17-66, Ano18-64, Ano18-65]. Collecting [KNH^{+}17]. Collective [HZC^{+}22]. Collisions [PBA^{+}23]. Collusion [ZXN^{+}21]. Colored [SK18]. Colorful [ZYZU22]. Colouration [WWH^{+}18]. Comb [RJW19]. Combination [CPCBFSF16, CJ15, FdDARF16, IIY16, MLMT24, YOM21]. Combinatorial [LLCT22]. Combined [BCB15, MGGK17, MHGD15, PZZZ22, RvWM22]. Combined-Sample [PZZZ22]. Combining [CMY16, CLL^{+}21b, KSK^{+}15, LCC^{+}15, MGB18, RDP14, SKM17, SY15, WJ15a, WW19, YK17]. Command [PCG17]. Commands [VSL21]. Comments [CE14, LLS24]. Common [GSP23, PP19, SD16b, SD16a]. Common-Slope [GSP23]. Commonsense [LPL^{+}21, ZLW^{+}22]. Communication [Ano16-50, Ano16-51, KBV17, LWCC21, LHM21, MHD^{+}17, RCR^{+}21, SWH^{+}17, SHKB17]. Community [WYLZ16, ZXH^{+}16]. Community-Based [WYLZ16]. Compact [KYL24, MBD^{+}15, OS21, SSI21, SRH15, WDDL18]. Compaction [NEWN23]. Companding [Ziv15]. Comparative [BSR^{+}23, DST^{+}19]. Comparing [Abd18, TM21]. Comparison [AJB^{+}18, AHF21, BKS^{+}18, KBV16, TON^{+}22, WDD^{+}16, vHK14]. Comparisons [RGK24]. Compensated [FPTJ18]. Compensation [BGQO^{+}17, BBHAR24, DV23, DBP^{+}24, DH14, HSK17, LC16b, MGF18, MCF21, MJJ^{+}14, Nos16, TZA14, WVCL14]. Competence [MMM23]. Complementary [DZF^{+}23]. Complete [CE14]. Completely [HSM19]. Completion [LHL^{+}23b, WWG^{+}24]. Complex [Emu18, FLL^{+}23, GTSM24, HLL15, HIA^{+}17, HKI23, HHC16, KBG16b, LK19, LLZ^{+}21a, MV19, MF18, NTY19, TW20, WWW21, WWX^{+}24, WWW16, WW17, ZMC^{+}24, ZLZ^{+}24]. Complex-Domain [HKI23]. Complex-Valued [Emu18, HHC16, LK19, NTY19]. Complexity [CBA^{+}15, DDZS17, FWM^{+}14a, FWM^{+}14b, FMD14, HL14, LZZP22, NWM16, SDF18, SLC^{+}15, SLW15, VH19, WW15]. Component [BV23, CWYC17, GJYW23, IIY16, ZGL^{+}20]. Components [CB14]. Composers [TP22]. Composers-Performers [TP22]. Composite [LLW23]. Composition [CLAG16, MFG15]. Compositional [ZM17]. Compositionality [DWX^{+}16]. Comprehension [LLW^{+}19, LYCX22, MSW^{+}22, MZZ23, ZZL^{+}19, ZLZ21, ZZZ22a, ZZZ22b, ZZD23, ZZZL22]. Comprehensive [KPA^{+}24]. Compressed [KMM^{+}18, LWL^{+}21, LWLS24, MCD14]. Compression [CZG^{+}15, HPPM24, HDW^{+}23, KGH^{+}16, LYH^{+}22, MF17, NR14b, TW21, WLT^{+}18]. Compressive [WLL^{+}16, WZS14].

Computation [AB21, FB18, LPI^{+}21, LLL24]. Computational [SSKP16, SMO^{+}21]. Computationally [KAJC16, MDM15]. Computer [MFG15, QMS16, hShWsL15, WsL15]. Computer-Aided [MFG15, QMS16, hShWsL15]. Computer-Assisted [WsL15]. Concatenative [ZLD21]. Concealment [LC16a]. Concentric [HCB18, PC24, SCB21, ZHCB21]. Concept [AFG^{+}23, ZCHW24]. Concertos [ÖM24]. Concurrent [MASC23, SCEH19]. Condition [Bor23, XT23]. Condition-Aware [Bor23]. Conditional [LZAY23, May18, RLPK15, TL19, WLT24, YLL15]. Conditioned [BSW^{+}24, DVO^{+}23, DYY^{+}23, SHG21]. Conditioning [NNC21]. Conditions [BHBS16, CW18, DW17, GLQ21, KK23, RKN14, SLGZ15, ZWW14, ZWW15]. Conducted [HJZB23, LYY24, WCZR23]. Conduction [SFK15]. Confidence [DAGS^{+}18, DXW^{+}23, RGR^{+}22]. Configurable [HJZB23]. Configuration [TR14b]. Conformer [DXW^{+}23, WDW^{+}23]. Confused [RARR15]. Confusion [LHD16]. Conjugate [SLZ^{+}21, SBP^{+}24, XCZ^{+}23]. Connectionist [OOK17, WJ23]. Connectives [MHPB15]. Conquer [LW19]. Considering [HSWC21, WYY24, WZS14]. Consistency [CCH^{+}21, HCC^{+}23, LZ16, LMW^{+}24]. Consistent [JT15, KWZ24]. Consolidated [GVMGO17]. Consonant [MLC^{+}23]. Consonant-Vowel [MLC^{+}23]. Constant [PC24]. Constituent [LZZP22, ZDC14]. Constrained [AHH^{+}19, AHH^{+}20, ABKP19, BC18, CXvG^{+}23, FD21, GQXZ19, KJ21, KSHH18, LFQL19, MHGD15, RMBK14, SGL^{+}21, SGLS23, SZW23, Wid19, YSCZ15, YH17, ZW19, ZDD21]. Constraint [BV17, HCBY18, LHL^{+}23b, LWYC23]. Constraint-Based [LWYC23]. Constraints [AVM19, HMDG15, SBV17, SLZ^{+}21, Wid18, ZSC^{+}23b]. Constructing [Chi18a, XPC^{+}22]. Construction [BK24b, DJY^{+}19, HZC^{+}22, SMB18, VS14, WSY14]. Contact [AMK^{+}20]. Containing [NNS16]. Contemporary [XXWY24]. Content [GWL^{+}24, yLsL14, sLGyLaC15, LCW^{+}18, LLW^{+}19, LXS^{+}18, XSC^{+}23]. Content-Oriented [LXS^{+}18]. Contents [Ano14-92, Ano14-93, Ano14-96, Ano14-97, Ano14-100, Ano14-101, Ano14-104, Ano14-105, Ano14-108, Ano14-109, Ano14-112, Ano14-113, Ano15-79, Ano15-80, Ano15-81, Ano15-82, Ano15-83, Ano15-84, Ano15-85, Ano15-86, Ano15-87, Ano15-88, Ano15-89, Ano15-90, Ano15-91, Ano15-92, Ano15-93, Ano16-58, Ano16-59, Ano16-62, Ano16-63, Ano16-64, Ano16-65, Ano16-66, Ano16-67, Ano16-68, Ano16-69, Ano16-70, Ano16-71, Ano16-72, Ano16-73, Ano16-74, Ano16-75, Ano17-99, Ano17-78, Ano17-79, Ano17-80, Ano17-81, Ano17-82, Ano17-83, Ano17-84, Ano17-85, Ano17-86, Ano17-87, Ano17-88, Ano17-89, Ano17-90, Ano17-91, Ano17-92, Ano17-93, Ano17-94, Ano17-95, Ano17-96, Ano17-97, Ano17-98, Ano18-80, Ano18-81, Ano18-82, Ano18-83, Ano18-84, Ano18-85, Ano18-86, Ano18-87, Ano18-88, Ano18-89, Ano18-90, Ano18-70, Ano18-71, Ano18-72, Ano18-73, Ano18-74, Ano18-75, Ano18-76, Ano18-77, Ano18-78, Ano18-79, Ano19-31]. Contents [Ano19v, Ano19w, Ano19x, Ano19y, Ano19z, Ano19-28, Ano19-29, Ano19-30, Ano21c, Ano21d, Ano14-94, Ano14-95, Ano14-98, Ano14-99, Ano14-102, Ano14-103, Ano14-106, Ano14-107, Ano14-110, Ano14-111, Ano14-114, Ano15-75, Ano15-76, Ano15-77, Ano15-78, Ano16-60, Ano16-61, Ano18-69, Ano19-32, Ano19u, Ano19-27, Ano19-33, Ano19-34, Ano19-35, Ano19-36]. Context [AF23, BSR17, BSTM21, CGL^{+}15, CH16, CZY^{+}18, CWU^{+}19, CLW24, CBCL15, CDW16, DZWC20, DKO^{+}18, DFY^{+}24, HQTX23, JCYwH21, LKC^{+}21, LZC^{+}23, LHM21, MPHM19, MZL21b, QCN^{+}21, RGR^{+}22, SF18, SFIL17, SgL20, TWZ^{+}18, WS14, WZZ^{+}24, WZ24, WHLG14, XH23, XSL^{+}19, ZZJ^{+}23, ZDC14, ZXSD17, ZQF^{+}24]. Context-Aware [DZWC20, DFY^{+}24, MZL21b, QCN^{+}21, SF18, TWZ^{+}18, XSL^{+}19, ZXSD17]. Context-Based [HQTX23]. Context-Dependent [BSR17, CDW16, WS14, WHLG14]. Context-Infused [ZQF^{+}24]. Context-Sensitive-Chunk [CH16]. Contexts [AJK24, CLW^{+}19, NAK^{+}17, SBG17]. Contextual [LWJ19, LLZ^{+}21b, LLZ23b, LCY^{+}23, QCN^{+}21, SZW23, ZW16a]. Contextualized [BZ23, GLL^{+}21, LCG^{+}24, QVZY24, WZ24]. Continuity [PSL17]. Continuity/Discontinuity [PSL17]. Continuous [Abd18, BDL15, DVO^{+}23, DB15, FGQ^{+}15, HQTX23, LFH^{+}15, LCQ22, LLLL19, LCTA19, MAL18, RF19, SFIL17, WZL^{+}15, WWW21, vNKB^{+}23]. Continuous-Space [WZL^{+}15]. Contours [KYI^{+}15]. Contraction [WHLG14]. Contrast [SDDCC15, Wid19]. Contrastive [LYY^{+}23, LHL^{+}23a, LLHW24, LFL^{+}23, LWY^{+}23, LLZ^{+}23c, STT^{+}19, SHL^{+}21, ZYD^{+}21, ZSC^{+}23a, ZWW^{+}23b, ZZZW23]. Contrasts [SMvdP22]. Contribution [KKGK18]. Contributions [SLR16]. Control [AKUS23, AA17, ATR^{+}21, AC24, AFdDG23, AKN^{+}16, BSB14, BBJ18, CC14a, CE14, CD17, CDA^{+}14, FdDPG21, GdC14, HBK24, KCD20, KBI^{+}21, LYWX22, LLZ^{+}23a, LWY^{+}23, LFdDG14, LSGH24, MX17, MZA20, MMS^{+}20, MMZ23, MNFGO19, SGL^{+}21, SGLS23, SGL^{+}23, SLW15, SKM19, TM16, TNS^{+}18, VH19, VE24, XCZ^{+}23, YEY17, YPL21, ZAZ^{+}18, ZHB^{+}18, ZTQB19, ZZZ^{+}21, ZQZ^{+}23, ZPW23, ZLX^{+}24, ZLSC23]. Controllable [BRC24, DY22, HCC^{+}23, LNNL22, MOT^{+}23, TKT17, WFY^{+}22, YWT23]. Controlled [HWT^{+}17, KGH^{+}16, LCWC16, MJ18, VR19, YZW^{+}18, CHC^{+}14]. Controller [VE24]. Controllers [BSB14]. Controlling [SCB21]. Controls [KKJK23, PCW15]. Conv [LM19]. Conv-TasNet [LM19]. Conventional [Zhe23]. Convergence [YEY17]. Conversation [JMD^{+}17, WWLL14, YZW^{+}18]. Conversational [ESVK17, LLT21b, MSFK15, SWC20, WTHJ14, WXS24, XDH^{+}17, XWS^{+}21, YB14]. Conversations [ADLX23, HPB15, HSWC21, LWLL24, PMB^{+}20, SFEB23, SWZ16, XZFN23, YLGG22, YB14]. Conversion [ATA16, CLLD14, DZLGO20, DLL^{+}21, DTN^{+}14, HSM19, HHW^{+}21, JRA17, KKTH19, LNNL22, Lib21, LCW^{+}21, LCTA19, LLL^{+}23, MZC^{+}24, NTA15, NTM16, PS17, PM15, SHCV23, SZL19, SYKL21, TLW^{+}17, WWX^{+}23, WVCL14, XT23, XSC^{+}23, ZLL^{+}19, ZLD20, ZZZL21, ZKDS18, ZGO19, ZWTL23]. Converting [ACRS14, ZDGO21]. Convex [JBTL15, KHHJ19a, MMK21, PC21]. Convolution [GZG23, MXH21, MZL21a, WFR^{+}21, WHO^{+}21, WHT^{+}21, XLZ^{+}23]. Convolutional [AHMJ^{+}14, AdTJ18, CC21, CP21, CDW16, DGMB21, FWT^{+}18, GBL24, GTSM24, GAPGG19, JGM23, JSYW16, JG19, KH24, KEzW21, LLC22, LdLvWWS21, LHG^{+}24, MPC^{+}24, MZL21b, NIK^{+}22, OCS21, OGZ18, PHM^{+}17, QBTY16, QCN^{+}21, SP16, SGLS23, TQH^{+}18, TW20, VDH24, WPY^{+}21, YHL^{+}23, ZKH18, ZPD19, ZWZ^{+}21, ZZY^{+}24, ZLF19, rPH^{+}17]. Convolutions [TCW19]. Convolutive [BP14, CY15, FK19, KBGAP^{+}16, LGGH18, LGGH19b, MD16, SOF15, WWL21, WYY22].

Convolved [SDDCC15]. Cooperative [HKM^{+}24, ZCDS15]. Coprime [MLP^{+}24]. Copy [AGB^{+}24, GBL24]. CorefDPR [YLGG22]. Coreference [YLGG22]. Corpora [LFU^{+}17, ZB14]. Corpus [CWH^{+}16, CMY^{+}21, LLWC21, MMBT21, PW22, WCZR23, ZS20]. Corpus-Aware [CMY^{+}21]. Corrected [APWB23]. Correction [AAA19, BHAMR19, FWM^{+}14a, KDJJ17, ZWHD23, ZLJ22]. Corrections [SPM17a]. Correlated [PSL^{+}16]. Correlates [HHZ^{+}21]. Correlation [ATP16, CTE21, FLL^{+}23, MC17, PACPJF^{+}23, SAC17, WD16, PT14]. Correlation-Based [PACPJF^{+}23]. Correlations [Lin18, SM14]. Correlator [AM20]. Correntropy [ZGZ23]. Correspondence [GHS^{+}21]. Corrupted [DBN^{+}17]. Corruption [MIR^{+}21]. Cortical [SKM17]. Cosine [SKSD14]. Cost [CPSFC^{+}14, DCW^{+}21, KSHH18, SSKP16, SLK17, ZN14]. Counterfactual [WZZ^{+}24]. Counterfactually [SHJY21]. Countermeasures [WDD^{+}16, ZWC^{+}23]. Counting [CH23, NK22, WHRC16, YLPL19]. CountNet [SCEH19]. Coupled [AZK15, AGD^{+}22, BVGV15, LCZ^{+}17]. Coupled-HMM-Based [AZK15]. Coupling [DSD23, LSTK21]. Course [yLSY^{+}14]. Covariance [CONVVC18, EJ16, MUT^{+}20, MXNB19, NV14, PC19, SIKN23, YLH23, ZF14, ZDLK18]. Cover [Ano14u, Ano14v, Ano14-29, Ano14-30, Ano14-31, Ano14-32, Ano15r, Ano15s, Ano15v, Ano15w, Ano15x, Ano15y, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano17-28, Ano17-29, Ano17-30, Ano17-31, Ano17-32, Ano17-33, Ano17-34, Ano17-35, Ano17-36, Ano17-37, Ano18-33, Ano18-34, Ano18-35, Ano18-36, Ano18-37, Ano18-38, Ano18-39, Ano18-40, Ano18-41, Ano18-42, FDK15, HC19, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14w, Ano14x, Ano14y, Ano14z, Ano14-27, Ano14-28, Ano15n, Ano15o, Ano15p, Ano15q, Ano15t, Ano15u, Ano16m, Ano16n]. Covert [STBC17]. COVID [ZTM^{+}23]. COVID-19 [ZTM^{+}23]. Cramér [SGH17b, WCY19]. CRBs [LLGG20, LG20]. Creation [DHYG14]. CRF [WL19]. Cries [SK16b]. Criteria [JBTL15, WJ15b, vNKB^{+}23]. Criterion [HXWS15, SH19, WSH15]. Cross [AF23, ASC^{+}18, BYT^{+}21a, BWZC22, CDLL21, CHWL14, Chi18a, CTE21, CLHK21, DAFS15, DKDN20, FL19, FLC^{+}23, GDC19, GHAP17, GHJ^{+}23, HZW^{+}22, JHLK18, JBWZ24, LCLM18, LYZ21, LWX^{+}22, LHC^{+}23, LSZ23, LLC^{+}24, LLWQ24, LJX^{+}16, Lin18, LFL^{+}23, LYCX22, LGR14, PW22, PT14, PM15, QWW^{+}23, SV18, SDK17, SPM17a, SPM17b, SLWY20, TXH^{+}23, WJ23, WQ24b, XLZ^{+}23, XHS^{+}18, YsL15, ZFC^{+}23, ZZZ16, ZS20, ZSC^{+}23a, ZQZ^{+}23, ZZS^{+}23, ZWX15, ZTL21, ZWTL23, ZLL^{+}24a, ZCW^{+}24, ZLL^{+}24b]. Cross-Attention [QWW^{+}23, XLZ^{+}23]. Cross-Channel [GDC19]. Cross-Corpus [PW22, ZS20]. Cross-correlation [PT14]. Cross-Dialect [Chi18a]. Cross-Domain [JBWZ24, LSZ23, LYCX22, SLWY20, XHS^{+}18, ZFC^{+}23, ZCW^{+}24]. Cross-Entropy [BWZC22, SV18, WJ23]. Cross-Entropy-Guided [CDLL21]. Cross-Language [JHLK18, YsL15, ZZZ16, ZWX15]. Cross-Lingual [ASC^{+}18, CHWL14, DKDN20, FL19, FLC^{+}23, GHJ^{+}23, HZW^{+}22, LYZ21, LHC^{+}23, LLC^{+}24, LJX^{+}16, LFL^{+}23, LGR14, PM15, SDK17, TXH^{+}23, WQ24b, ZTL21, ZWTL23, ZLL^{+}24b]. Cross-Modal [AF23, BYT^{+}21a, LLWQ24, ZSC^{+}23a, ZLL^{+}24a]. Cross-Modality [CLHK21]. Cross-Recording [LCLM18]. Cross-Show [DAFS15]. Cross-Speaker [GHAP17, LWX^{+}22, ZQZ^{+}23, ZLL^{+}24b]. Cross-Spectra [SPM17a, SPM17b]. Cross-Utterance [ZZS^{+}23]. Crosslingual [SV17, YGS19]. Crosstalk [KJ21, MV14, MMK21]. Crowdsourced [LB19, MMM23]. Crowdsourcing [GMH17]. CSS [SHL^{+}21]. CSS-LM [SHL^{+}21]. CTC [DLY^{+}19, FCCA23, YKL^{+}23, ZM21]. CTC-Based [ZM21]. CTNet [LLT21b]. Cube [WBC^{+}21]. Cue [HMDG15, MHGD15, SBVM16, TE23, XCHX24]. Cue-Preserving [TE23]. Cued [LLL24]. Cues [ABKP19, May18]. CUHK [LGH^{+}21]. Cumulative [CPSFC^{+}14]. Cumulative-Sum-Based [CPSFC^{+}14]. CUPVC [LWYC23]. Current [SLGA15]. Curriculum [BWZC22, CZWC23, GHJ^{+}23, LB19, RH18, ZMX^{+}21, ZWL^{+}23a]. Curriculum-Style [GHJ^{+}23]. Curve [MTBB15]. Custom [ISZA15]. Customer [AMK^{+}20, MSW^{+}22]. Customized [LYY^{+}23]. Cylindrical [KFH^{+}14, MFH21, PMB^{+}19].

D [BWH18, BSR^{+}23, CT14, DGMB21, DG23, ECH23, GDM18, HB17, MASC23, SMH14, WSJ^{+}23, WAS16, WSFS19, YLLC19, ZLHL21, ZHCB23, vMM14]. D-CNN [MASC23]. D-Score [ZLHL21]. Damped [HCY21b]. Dance [GKC16]. Data [ARSLN16, AGV24, AGB^{+}24, AM18a, BYT^{+}21b, Bre15, CBG15, CMY16, CGK15, DCZ19, DDZS17, DWX^{+}21, DZZ23, FMB16, GCG^{+}17, HRL^{+}21, HSM19, HMO21, JQL24, JWH^{+}22, JGD^{+}24, KXS^{+}19, LGTG18, LAK16, LMW^{+}24, LGL^{+}22, LFWZ23, MSD14, NSNN20, PJTJ24, QZ22, QCL23, RG21, SDK17, SSAS21, WDW^{+}23, WQ24b, WK24b, XYH14, XYY^{+}23, YLFL22, ZZZL21, ZZZ^{+}24]. Data-Association [AM18a]. Data-Driven [DDZS17, DWX^{+}21, LGTG18, PJTJ24]. Data-Efficient [SSAS21]. Data-to-Text [AGB^{+}24]. Database [KG19, SLH17, WDD^{+}16, ZWC^{+}23]. Dataset [FFP^{+}22, LWY^{+}23, LLC^{+}23b]. Datastore [DZD^{+}24]. DBSA [ZZX^{+}23b]. DBSA-Net [ZZX^{+}23b]. DCASE [MHB^{+}18, MDE^{+}19, PMA^{+}21, SMA^{+}17]. DCT [SB18, XH23]. De-synchronization [XNG^{+}14]. Deal [KMAB18]. Dealing [WZMH22]. Debiasing [WZZ^{+}24, ZLC^{+}24]. Decay [BY15, SH14b]. Decaying [BV17]. Decentralized [CZK^{+}21, ZTQB19]. Decide [LCL^{+}21]. Decimation [HKHP23]. Decision [SHAGPRZ17, SWC20, ZXZZ23]. Declipping [GaKGB21]. Decoder [CHWQ24, DOL^{+}23, HFW^{+}22, JSG^{+}22, LKC^{+}21, LPM^{+}23, LWCL19, TZS^{+}21]. Decoding [ACRS14, CXvG^{+}23, GLQ21, KLK17, MHBT15, MF17, WL19, ZTWY19]. Decomposed [ZYW^{+}24]. Decomposition [ADM^{+}19, BIK^{+}18, BG21, CSS23, DGCA23, HWW23a, JP14, KD18, LWYD23, MMS^{+}20, MK15b, NK21, NEN21, PBC18, PSL17, SBP^{+}24, TZA14, TZZ^{+}15a, WCY15, WCC^{+}21]. Decomposition-Based [KD18, SBP^{+}24]. Decompositions [WBC^{+}21]. Deconvolution [DVLH17]. Decorrelation [SFV23, ZZH16]. Decoupled [HZF17]. Decoupling [CBCL15, DCYY22, LLL^{+}23, SGY^{+}24]. Deep [AFG^{+}23, AF18, BHP^{+}23, BSR17, CLA^{+}21, CLAG16, CDLL19, CLLD14, CRVB^{+}15, CM15, CH16, CCT^{+}19, CLL^{+}21b, CTLL17, CGK15, CG16, DAGS^{+}18, DQY18, DTDL16, FLMS16, GN23, GTV19, GTSM24, GH17, GvdP23, GRSP17, GJ17, GLL^{+}21, HBK24, HHO^{+}21, HL18, HKHJS15, ICB24, JMD^{+}17, JWLF14, JGSC^{+}22, KK19a, KKGK18, KSU^{+}19, KYTJ17, KK19b, KG19, KHE21, KEzW21, LC16a, LCW^{+}18, LK19, LLC22, LS14a, LCSL19, LWLS24, LYL21, LM22, LYH^{+}22, LM23b, LW19, LR17, LDZ19, LCM18, LSGH24, MMB17, MGB18, MLC^{+}23, MZX^{+}15, MZM15, MTZ^{+}21, MMR^{+}22, MPC^{+}24, MKA16, NW15, NR22, NLL^{+}18, NLV16, OCS21, PS19, PDS^{+}16, PGS^{+}22, QDSL19, QYCT23, QBTY16, QCDW17, QCW21, QVZY24, RARR15, SWW^{+}17, STS21, SS16, SHD14, STT^{+}19, SBN^{+}19, SSSS22, SAS17, SYKL21, SLWY20, SHZ^{+}23, SWEV23, SZVZ16, SWCN19, TD23, TQY16, TQH^{+}18, TZW21, TW21, TG23, TG24, TWIN23]. Deep [TDL19, WS14, WDDL17, WWS^{+}18, WC18a, WW19, WZW19, WBSG18, WYSD15, WGR^{+}18, WL19, WZ21, XHRC24, XLLW21, XDDL15, XHW^{+}17, XAHJ^{+}14, YLL21, YYZY24, YMM19, YK21b, YWW^{+}23, ZFC^{+}23, ZW16a, ZW16b, ZZN^{+}16, ZW17, ZH18, ZLW^{+}22, ZM17, ZQZH19, ZWW19, ZWS^{+}23, ZYW^{+}24, ZZ19, ZJD^{+}15, ZJL21]. Deep-Learning-Based [MTZ^{+}21]. Deep-Scale [ZWS^{+}23]. Deep-Sparse-Representation-Based [SAS17]. Deepfake [LWS^{+}23]. Deeply [HNT^{+}23, MMT^{+}19]. Deficient [LLGG20, Yan22]. Deficient-Length [Yan22]. Defined [MTBB15]. Deformable [KH24, MYL^{+}21]. Deformed [SLOVB15]. Degraded [AY17]. Delay [APH14, AM16, AGD^{+}22, BRD15, BGQO^{+}17, CB21, CRH^{+}14, DV14b, DSD23, DHCS15, FB18, FB19, HL14, HXL^{+}21, HXC^{+}22, LMV21, SH17a, SFV23, SKH^{+}15a, SKH^{+}15b, ZH14]. Delay-Free [DV14b, FB18, FB19]. Delay-Lines [HL14]. Delayed [FK24]. Delayless [CD17, KKJN17, LSGH24, PPSG17]. Delta [WSH15]. Delta-Bayesian [WSH15]. Demixing [FLB16]. Demodulation [AS15, VS14]. Denoised [JMD^{+}17, TMMZ18]. Denoising [BV17, Bay15, DZWC20, DB19, GJ17, HWW^{+}15, MDCS20, QL24, VRN18, WW17, ZLJ22, ZZX^{+}23b]. Dense [PW21]. Density [BKS^{+}18, ÇOH18, KD18, NCB23, SH17a, WGD15]. Dependence [CZY^{+}18, TKH21, WK24b, YZL^{+}18, ZLJ^{+}17]. Dependencies [BLZ21]. Dependency [CZZ15, GHJ^{+}23, JLZD24, LZC^{+}14, LZZP22, LSG^{+}23a, LLWJ16, LLL^{+}19, ODSM16, SKK14, WZZ^{+}18, WZ21, ZCY^{+}22]. Dependency-to-Dependency [WZZ^{+}18]. Dependent [BSR17, BSTM24, CDW16, CS23, DSD23, DMVP16, DDZL19, DSH18, KKJN17, KD17, LM23a, MASC23, MMR^{+}22, NTY19, STT^{+}19, SKAK16, SAK16, WS14, WHLG14, WHO^{+}21, WHT^{+}21, ZSB17]. Depth [LCQ23]. Depth-First [LCQ23]. Dereverberation [ADM^{+}19, BV17, BH18, DST^{+}19, DDMvW20a, DB19, FCG24, HWW^{+}15, HBCC22, ISP19, JvWGD15, KD16, KH24, LRWG23, LGGH18, LGGH19a, LDG^{+}23, LLW23, LZHN14, LLdCF23, MIYO14, May18, MD16, OIY^{+}14, PDKG24, QL24, RKL14, RWL^{+}23, SF18, SIKN23, SEM^{+}14, SGH15b, SGH15a, SGH16, SGH17b, SK15, TK14, UNI^{+}24, WWL21, WYY24, WW17, XBZ^{+}24, YG24, YK14, YC22]. Derived [LRG21, MVH16, SLY14]. Descending [BTBT18]. Descent [QYCT23]. Description [RCL18]. Descriptions [WQZ^{+}21, WvdHZ^{+}21]. Descriptor [FMD14]. Design [BC17, BC18, BSB14, BBHAR24, BBJ18, BCB18, CZ17, CMM^{+}23, CFMM14, CT14, DV23, HL14, HBC17, HCB20, IBC21, JEP14, JK15, KJ21, LYN14, LHJ^{+}23, Non14, PBC16, PCBS19, Son21, WW15, ZBC14, ZHCB21, ZHCB23, ZSL^{+}22]. Desired [YG24]. Desynchronization [ZZX^{+}21, ZZX^{+}23a]. Detect [CWZ^{+}23, KTH^{+}20, SK16b, XZSX23]. Detecting [BHH22, KAA^{+}24, KB14, LLZ^{+}21b]. Detection [AC24, ALH^{+}21, AAA19, BE17, BJ24, BPG23, CE15, CC21, CC14b, CTLL17, CL18a, DH20, DQY18, DWY21, DWX^{+}21, DTC15, DTC16, DTC17, DKDN20, DSH18, FZZ^{+}15, GMD24, GG18, GAPGG19, GJ22, GPF17, GdC14, HWT^{+}17, HHZ^{+}21, HFL^{+}23, HXWS15, JGM23, JLP^{+}21, JWBR21, JYC^{+}24, JGK^{+}22, KSH17, KDY14, KPH15, KSU^{+}19, KMG17, KXS^{+}19, KHE21, KKGH16, KM23, LLR^{+}16, LSW18, LCSL19, LWL^{+}21, LPL^{+}21, LJWC23, LWS^{+}23, LHL^{+}22b, LLZ^{+}23c, MAP14, MLMT24, MSB18, McL14, MCZ21, MHB^{+}18, MDE^{+}19, MAT15, MKMDM17, NSA21, PMMM15, PMA^{+}21, PLT^{+}18, QMS16, QCDW17, MG18, RAB18, RA23, RPL^{+}23, STH^{+}18, SHJT21, SGA15, SH17b, SLK17, SM18, TH18, Ter16, TMKO17, TSKO20, VAP19, WAS14, WsL15, WDDL17, WSL^{+}20, WWG^{+}23, WDW^{+}23, WK24a, WSH15, WSG21, XYY^{+}23, XLG^{+}23, YBC^{+}14, YDZ19, YWDQ21, YB14, YLY15, YH21, YUK21, ZW16a]. Detection [ZWW23a, ZXZZ23, ZWC^{+}23, ZLJB23, ZH15, ZTM^{+}23, ZB14, rPH^{+}17]. Detection-Estimation [SM18]. Detector [AFG^{+}23]. Detectors [DLL^{+}21]. Determinantal [XO16]. Determination [ZKA15]. Determine [BM15]. Determined [KOS^{+}16, LKM23, MMT^{+}19, MTK^{+}20, OG14, WRC16, WC18b, YK21a]. Deterministic [WFY^{+}22, WC16a]. DetTrans [XZSX23]. Development [MTS^{+}19]. Developmental [BHH22]. Device [HWRC15, MHD^{+}17, TYZ^{+}22, XYY^{+}23]. Devices [HdHTJ22, LV22, TALP24, ZL21a]. DFT [LSNC21, SH19, YEY17]. DFT-Domain [LSNC21, SH19, YEY17]. Diacritization [RARR15]. Diagnosis [AATKS21, DKDN20, QMS16]. Diagonal [SDF18]. Diagonalization [IISN21, SOF15]. Dialect [AJK24, BDV^{+}14, Chi18a, GH23a, GJB15, HYI^{+}15, ZH18]. Dialects [CTSC14]. Dialog [BYY^{+}23, CLW24, JHLK18, KSG^{+}21, KGG^{+}21, KSK21, LLZ^{+}21c, LCL^{+}23b, SWC20, TDDN24, VSK^{+}24, WLL15b, YCC^{+}24]. DialogMCF [CLW24]. Dialogue [BM21, CCT^{+}19, CZWC23, CLHK21, GY14, GLL^{+}21, HCC^{+}23, JGHN23, KG17, KG19, LXK^{+}21, LPL^{+}22, LWLL24, LCL^{+}21, LZX^{+}23, LGL^{+}24, LSYN19, MSW^{+}22, MZZ23, MMBT21, MHBT15, QXW^{+}23, RLK14, RC23, SSAS21, hShWsL15, SWC^{+}21, WSJ^{+}23, WZZ^{+}24, WBSG18, WXS24, YSCZ15, ZYL^{+}21, ZLHL21, ZLZ21, ZDZ^{+}23, ZYW^{+}24, ZCHW24, Zhe23, ZJL21, ZTC23]. Dialogues [ZZH21, ZJL21]. Diarization [BSW^{+}24, CHWQ24, CCY^{+}24, DMSM21, DAFS15, DBLČ20, FMB16, HDLL23, HFW^{+}22, HWG^{+}23, LCLM18, NDS^{+}23, PKP^{+}21, SKSD14, SG21, YB14, YH17]. DiCLET [LHC^{+}23]. DiCLET-TTS [LHC^{+}23]. Dictionaries [BVGV15, BK24b, CWDA17, LKPR18]. Dictionary [BXY14, DCAS22, DVLH17, GFG^{+}14, LLC23a, RBR16, VRN18, WED17]. Dictionary-Free [LLC23a]. Diffeomorphic [SHCV23]. Difference [BS14, CvdPHH21, CSY^{+}21, CC14b, GYXY23, LHW^{+}21, MPKS^{+}23, PS19, PZW^{+}21, SvWRK15, TWIN23, dCBC23]. Differences [CONVVC18, CTSC14, IC19, MYNO23]. Different [GLQ21, QCL23]. Differentiable [MS23, SFRK^{+}23, SR16b]. Differential [ABB^{+}23, BBAS19, BCB18, CCT21, CMM^{+}23, CBC19, HCB20, IBC21, JHW^{+}21, JBCH23, PCB15, PCBS19, TC19, TC22, WC16a, YBHC21, ZBC14, ZBCH21, ZHCB23]. Diffraction [KE23]. Diffsound [YYW^{+}23]. Diffuse [SK15]. Diffuseness [EJ16]. Diffusion [CZN^{+}24, DTC15, LRWG23, LLZ^{+}23a, LHC^{+}23, RWL^{+}23, YYW^{+}23]. Diffusion-Based [LRWG23, RWL^{+}23]. Digit [SAK16]. Digital [Ano15-28, Ano15-29, BBFS21, GdBBS21, GDM18, LC16a, MSS17, PPSG17, PCLP14, SAA15, SRH15, SMH14, VE24]. Dilated [TCW19, WHO^{+}21, WHT^{+}21, ZWZ^{+}21]. Dimension [LFWZ23, MJJ^{+}14]. Dimensional [AG17, ADGE19, DCZ19, KBV17, SMB18, UKS19, WCY19, WYLZ20, XDC^{+}17, ZA14]. DINO [HCQ24]. DIP [DSH18]. DIP-Based [DSH18]. Diphthong [GDM18]. Diphthongs [AG17, ADGE19]. Diplophonic [AHSS^{+}18, AP21]. Direct [AL23, CASZ17, DGCA23, FWT^{+}18, GCG^{+}17, JD17, LGHG17, MTNR21, NR14a, RJCW17, SAC17, SLK17, YLL21, ZM20]. Direct-Path [LGHG17, NR14a, YLL21]. Direct-to-Reverberant [CASZ17, SAC17, ZM20]. Directed [ZLZ^{+}24]. Direction [BKTR24, DGMB23, ED18, GWLL22, HMN17, HMO21, HAS21, KK14, NV14, PCB16b, SLC^{+}15, TP15, TTH14, TR15]. Direction-of-Arrival [BKTR24, HMO21, KK14, SLC^{+}15, TTH14]. Directional [BH19, HPPM24, KSK^{+}15, PAB19, TS14, WLT24]. Directions [ES16]. Directive [WCCB19]. Directivity [AGV24, AB21, LHJ^{+}23, PBC16, RHvdP^{+}16, WC16a]. Directly [TM21]. Dirichlet [Chi15a, HSN18, KGG19, TP16]. Disambiguating [MHPB15]. Discontinuity [PSL17]. Discontinuous [LJL^{+}23]. Discourse [BSL15, HNY^{+}15, HW23, JQL24, MHPB15, MIR^{+}21, MZL21b, WYH^{+}15, WSJ^{+}23]. Discovered [Kam23]. Discovery [CL18a, KJG16, SHDG16, WsL15, WSZN21, XO16]. Discrepancy [CLL^{+}21b, LMC18]. Discrete [SKM19, YYW^{+}23]. Discriminant [CPL14, LSZ23, WCC^{+}15, XAHJ^{+}14, ZS20]. Discriminating [AY15]. Discrimination [BTBT18]. Discriminative [BXY14, HXWS15, LAK16, LSG^{+}23b, LSTK21, TR14a, WHQY19, WJ15b, ZSH20, ZXZZ23]. Discriminatively [HSPK18, ZF14]. Discriminator [CLL^{+}21b]. Disease [AATKS21, LRP^{+}23, NSA21, NRČR14, RA23]. Disentangled [AJK24, HCC^{+}23, LZAY23, WWG^{+}23, ZLD20, ZMZ24]. Disentanglement [BK24a, CHGS23, LCW24, TALP24, ZQZ^{+}23]. Disentangling [ASX22, LWX^{+}22, QLW^{+}24]. Disfluencies [KHE21]. Disorder [LLLL19]. Disorders [AFG^{+}23, NRČR14]. Dissonance [MBTB14]. Distance [AB21, CBCL15, GSA18, KHTL18, KGBPM24, SKSD14, YXX^{+}22, ZM20]. Distance-Based [SKSD14]. Distant [MLMT24, MH16, NHK16, WHRC16]. Distantly [MSW^{+}22]. Distill [TP23]. Distillation [CLCW18, DZD^{+}24, HDW^{+}23, IK23, JQL24, JKU18, KWZ24, LWL^{+}22, LLZ23b, TP23, YLK^{+}21, YKL^{+}23, ZXSL19]. Distinct [Lin18]. Distortion [AM18b, DH14, FNA15, HSK17, MJJ^{+}14, PMB^{+}19, SWEV23, WFN^{+}18, WTW20, ZTDD23b]. Distortion-Independent [WTW20]. Distortionless [CP21, MYK^{+}24, YOM21]. Distractor [XPC^{+}22]. Distributed [AFdDG23, BF18, CBA^{+}15, CZZ15, CZK^{+}21, DPHG18, FSEI21, HGYN23, HL18, HCY21a, HCY21b, HZBW23, HSLB23, IO17, KSHH18, LGTG17, LLZ^{+}23a, MMS^{+}20, MH16, RvWM22, SAP14, SKDN14, SKH^{+}15a, TH14, TH16, TCY15, WD16, WRC16, WLT^{+}18, ZH14, ZCY16, ZHH18]. Distribution [ATP15, Cum21, FYG^{+}21, LB23, WXZ^{+}19, YXX^{+}22, ZDLK18]. Distributions [ATP16, CDLL19, CS23, SKM19, WAS16, WK24b, CPL14]. Disturbances [NCC15]. Divergence [FSO17, SLGZ15]. Divergences [ISP19, SDDCC15]. Diverse [DY22, RG21, TLD^{+}23, TZZ15b, XPC^{+}22]. Diversity [AZGN14, FL19]. Divide [LW19, TP23]. Diving [WWS^{+}18]. Division [FFSS18]. DNN [Abd18, BSTM24, CDLL21, EMTF18, FSEI21, GDC19, GLWQ22, HSM19, KNH^{+}18, LMC17, MYNO23, RG21, SV18, TMM18, WSY14, WK16, XZF23]. DNN-Based [MYNO23, RG21, Abd18, CDLL21, FSEI21, GDC19, KNH^{+}18, TMM18, WK16].

DNN-Driven [LMC17]. DNN-Extracted [WSY14]. DNN-Supported [EMTF18]. DNNs [SNK21, SMA^{+}17]. Do [GSA18, WK24a]. DOA [AM18a, BSTM21, CH18, ÇH21, FSA20, HNF20, HZF17, JTHN14, JCBJ15, JWBR21, KAJC16, LJWC23, MLP^{+}24, RKN14, VDH24, WCY19, WAJL15, WRK18, XBZ^{+}24]. DOA-Based [WCY19]. Doctor [ZYZU22]. Document [AGC23, DDZL19, HPB15, HNY^{+}15, yLsL14, LLJ^{+}23, LHG^{+}24, TDDN24, YW14, YCLS17, ZZJ^{+}23, ZYW^{+}20]. Document-Grounded [TDDN24]. Document-Level [ZZJ^{+}23]. Document-to-Document [LLJ^{+}23]. Document/Query [yLsL14]. Documents [GMH17, MBD^{+}15]. Does [YLC^{+}24]. Dolphin [KD21]. Domain [AB18, ARSLN16, AGW15, BP14, BM21, BvWWM17, BHBS16, BAS20, BCB18, CQ22, CS23, DLG18, DMVP16, DB18, DB19, DBM15, Emu20, FCG24, GEC15, GH23a, GCvWMJ14, GZZY23, GWSC23, HMN17, HCBY18, HKI23, HHC16, JTHN14, JPX^{+}23, JBWZ24, JGHN23, JGK^{+}22, KJ21, KYL24, KVMV^{+}24, KFH^{+}14, LPHCB14, LSNC21, LKHU14, LSZ23, LWLL24, LMC18, LYCX22, LZ23, LGL^{+}24, MMS^{+}20, MHZ^{+}24, MAP14, MCZ21, MPKS^{+}23, MH18, MUT^{+}20, MTKS21, MBK17, NKS21, NNC21, NV21, OG14, PW19, PW21, PKR^{+}24, PDKG24, QCL23, RC23, SK16a, SH19, SSAS21, SGLS23, SCW21, SB18, SLWY20, SKM19, SKT21, TM16, Ter16, TKH21, TM22, VH19, VDM15, WB18, WUF^{+}18, WCC^{+}21, WH22, WW22, WWG^{+}23, WW17, XWWW18, XHS^{+}18, XMBN18, XMBN21, YEY17, Yan22, YPC^{+}22, ZYL^{+}21, ZFC^{+}23, ZAZ^{+}18, ZHB^{+}18, ZWL^{+}23b, ZCQ^{+}19, ZCW^{+}24, WZS14]. Domain [WWWR23]. Domain-Agnostic [SKT21]. Domain-Aware [BM21]. Domain-Invariant [ZWL^{+}23b].

Domain-Sensitive [MHZ^{+}24]. Domain-Shift [NNC21]. Domains [RKL14, YXX^{+}22]. Dominance [NR14a]. Dominant [YZLY17]. Double [CTE21, FLYJ23, GCvWMJ14, ICB24, SWC20]. Double-Cross-Correlation [CTE21]. Double-Talk [ICB24]. Double-Talk-Robust [GCvWMJ14]. Doubly [BGD^{+}18, TP16]. DP [CJ15]. DP-Based [CJ15]. DPGMM [WSZN21, WSZN22]. DPGMM-RNN [WSZN21, WSZN22]. Dramatic [MSEC21]. Drift [BBM17]. Drilling [ZH19]. Driven [BHP^{+}23, BE17, DDZS17, DWX^{+}21, DLMM^{+}18, HDW^{+}23, LGTG18, LMC17, PJTJ24, WYL^{+}23, ZTDD23a]. Driving [FFSS17]. Drone [MAZS22]. DropAttack [NLYK24]. Dropped [NLYK24, YLGG22]. Drug [KAA^{+}24]. Drum [DM16, LKPR18, WDS^{+}18]. DSTC10 [TDDN24, YCC^{+}24]. DSTC8 [JCYwH21, KGG^{+}21]. DSTC9 [TDDN24]. Dual [CLZ^{+}23, CZ17, CLZ^{+}24, CC14b, JEP14, KK14, LCQ22, LKR23, LLZ^{+}23c, LCTA19, PLQZ23, SYG^{+}24, TZW21, WLT24, ZWZ^{+}21, ZZY^{+}24, ZQZH19, ZZX^{+}23b]. Dual-Branch [SYG^{+}24]. Dual-Channel [WLT24, ZZY^{+}24]. Dual-Microphone [CC14b, KK14, TZW21]. Dual-Path [LCQ22, ZWZ^{+}21]. Dual-Query [CLZ^{+}24]. Dual-Radius [JEP14]. Dual-TBNet [LKR23]. Dual-Transformer-BiLSTM [LKR23]. Dubbing [LLC^{+}24]. Due [AFG^{+}23, AGV24, TGGE21, ACB17]. DUMA [ZZZL22]. Duration [DWY21, HWT^{+}17, LNNL22, WHQY19]. Duration-Controlled [HWT^{+}17]. During [RF19]. Dyadic [SWZ16]. Dynamic [AZK15, BGQO^{+}17, DBP^{+}24, GZG23, GDM18, Kam23, KMM^{+}18, KB14, KGH^{+}16, KGG19, MAP14, MV14, NDS^{+}23, TYZ^{+}22, TP19, WLL15b, XSL^{+}19, YLFL22, ZCYZ18]. Dynamical [BLLP17]. Dynamics [JNP^{+}21, PK15]. Dysarthric [JGD^{+}24, KKK15, LWYD23, LPI^{+}21, LGH^{+}21]. Dysphonic [ZLJB23].

E-Model [RCR^{+}21]. Ear [BHAMR19, BHAMR21]. Ear-Aligned [BHAMR21]. Earbuds [LP22]. Early [DDBW23, DDMvW20b, SGH15b, XGKM19]. Early-To-Late [XGKM19]. East [TGZ22]. EBEN [HJZB23]. EC [LP22]. EC-ANC [LP22]. Echo [CSAR^{+}14, DDBW20, DDBW23, DLMM^{+}18, Emu18, Emu20, GCvWMJ14, GYdL^{+}21, HBK24, HGT15c, ICB24, LWY^{+}23, LB16, PC19, PC21, RvWM22, SH17a, SK16a, SLGA15, TK14, VH19, YEY17, ZZH16, ZGL^{+}23, ZLLZ23, ZYG^{+}14, ZGZ23, ZXN^{+}21]. Echo-Based [HGT15c]. Edge [KE23, LP22, WFR^{+}21]. Edge-Enhanced [WFR^{+}21]. EDICS [Ano19-32, Ano16-32, Ano16-33, Ano16-34, Ano16-35, Ano16i, Ano16j, Ano17-59, Ano17-60, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17-99, Ano18v, Ano18w, Ano18x, Ano18y, Ano18z, Ano18-27, Ano18-28, Ano18-29, Ano18-30, Ano18-31, Ano18-80, Ano18-81, Ano18-82, Ano18-83, Ano18-84, Ano18-85, Ano18-86, Ano18-87, Ano18-88, Ano18-89, Ano18-90, Ano19-31, Ano19-33, Ano19-34, Ano19-35, Ano19-36, Ano14-45, Ano14-54, Ano14-55, Ano14-56, Ano14-46, Ano14-47, Ano14-48, Ano14-49, Ano14-50, Ano14-51, Ano14-52, Ano14-53, Ano15-40, Ano15-41, Ano15-42, Ano15-43, Ano15-44, Ano15-45, Ano15-46, Ano15-47, Ano15-48, Ano15-49, Ano16k, Ano16l]. Edit [CQ22]. Editing [BBC^{+}16, SWC^{+}21]. Editorial [Ano18-32, Ano19t, DRFO14, KSG^{+}21, Li15, Den14]. EEG [STBC17]. Effect [KH21, LC24, LD16, LV22, MTNR21, RJW19]. Effective [CBCL15, CL14a, HC19, SSK17, YND14, ZYF18, ZZL^{+}19]. Effects [BRC24, DB17, HAGA21, NWM16, SIUW21, SWEV23]. Efficiency [FSN24]. Efficient [ACRS14, BHAMR19, BZ21, CLW^{+}16, CXX^{+}23, DHCS15, GJ22, GZX^{+}21, HAGSZ16, IISN21, KAJC16, LHW^{+}21, LCW^{+}16, LLCT22, LCQ23, LLL24, MSS17, MBM15, MYKO24, MDM15, Nos16, PKR^{+}24, QWL^{+}19, SAPB15, SSAS21, TZA14, TZL^{+}23, TS16, WED16, WBSG18, WXL^{+}21, YLJ16, YSCZ15, ZTDD23a, ZSL^{+}22]. EfficientTTS [MZC^{+}24]. Effort [SAB16, WHB^{+}21]. EGG [CMY16]. Egocentric [JNP^{+}21]. Eigenspace [HSM19]. Eigenvalue [JP14, KD18, NEN21]. Eigenvalues [ZTQB19]. Eigenvector [PZP19]. Eigenvector-Based [PZP19]. Eigenvoice [DBLČ20, DTN^{+}14, MSD14]. Eigenvoices [SDK17]. Eighth [KSG^{+}21, KGG^{+}21]. ELD [CRH^{+}14]. Elderly [JGD^{+}24]. Electroglottographic [BMVG17]. Electrolaryngeal [XWWW19]. Electrolarynx [XWWW18]. Electromyographic [JD17]. Electronics [Ano16u, Ano16v]. Element [AG17, ADGE19, LWL^{+}21]. Elements [JK19]. Elevation [PC24, SCB21]. Elevation-Beamwidth [PC24]. Elicitation [LSYN19]. Elimination [NCC15]. Elitist [HHMK18]. Ellipsis [BZ21]. Embedded [DvWDM14, ZWC^{+}23]. Embedded-Optimization-Based [DvWDM14]. Embedding [BS21, CLZ^{+}23, CHWQ24, CLSyL21, FLL^{+}23, GMD24, GZG23, HDLL23, HWW23a, LWL^{+}21, LYZ21, LCQ22, LM22, LM23a, LJL^{+}23, LHYJ23, LSLK24, PDS^{+}16, QVZY24, SCY22, SGF^{+}21, TXH^{+}23, TM22, WHQY19, XNRG15, XNP^{+}18, XZL^{+}23, ZLZW21, ZLJ22, ZTL21, ZTC23]. Embedding-Aware [ZTC23]. Embeddings [AGM23, BSW^{+}24, DR22, KJG16, KMG21, LCZ^{+}19, MUK22, NNC20, NNC21, PHM^{+}17, SSI21, WCZ21, WZ24, XV21, YM17, YTF^{+}23, YWLZ18, YUK21, ZKH18, ZLJB23, ZM23, ŞUY^{+}18]. Embracing [Li15]. EMD [ZCF14]. EMG [JD17]. EMG-to-Speech [JD17]. EmoInt [SFEB23]. EmoInt-Trans [SFEB23]. Emotion [AB18, ADLX23, CWYC17, CCH^{+}21, CJY^{+}21, DXZ^{+}18, DG23, FYG^{+}21, GFG^{+}14, GWSC23, HZC^{+}22, HSWC21, HZL23, LFO^{+}24, LYWX22, LWX^{+}22, LHC^{+}23, LSZ23, LSH^{+}23, LLT21b, LB23, LSG^{+}23b, LKR23, LB19, LSYN19, MCL22, PCB16b, SCCM23, SHCV23, SHZ^{+}23, TD16, TP22, WCC^{+}15, WLLZ17, WC16b, WCL^{+}21b, XLL^{+}19, XDC^{+}17, YHC19, YHL^{+}23, ZCDS15, ZS20, ZQZ^{+}23, ZWS^{+}23, ZDZ^{+}21, ZLG^{+}22, ZLL^{+}24b]. Emotion-Cause [CJY^{+}21, FYG^{+}21, HZL23, SCCM23]. Emotion-Semantic-Enhanced [YHC19]. Emotional [KKJK23, LYWX22, LCTA19, ZLL^{+}24b]. Emotions [LLL^{+}23, SFEB23]. Emotive [WCL^{+}21a]. Empathetic [HCKW23]. Empathy [HCKW23]. Emphasis [DTN^{+}17, DSN18]. Empirical [ARSLN16, CSJD16, WSZN22, ZKA15]. En-HACN [LFM^{+}23]. Encoder [AM20, CHWQ24, DOL^{+}23, HFW^{+}22, KM21, LKC^{+}21, LPM^{+}23, LWLS24, LLyL21, LHYJ23, MPHM19, MUK22, RCL18, SZVZ16, TZS^{+}21, TLD^{+}23, WLZH21, XLZ^{+}23, ZXSD17]. Encoder-Decoder [CHWQ24, DOL^{+}23, HFW^{+}22, LKC^{+}21, TZS^{+}21]. Encoding [GLWQ22, JYB^{+}15, KCA21a, KCA21b, Lin19, YPC^{+}22, ZZXH22]. End [BYT^{+}21a, BYT^{+}21b, BWZC22, BNA24, BB21, CHWQ24, CMY^{+}22, CLHK21, DQY18, FYT^{+}21, FCCA23, FWT^{+}18, FJT^{+}23, GH23a, HBK24, HCJT15, HFW^{+}22, KLK17, KHE21, LWM^{+}20, LY21, LWX^{+}22, LDG^{+}23, LZL^{+}23, LFM^{+}23, MZC^{+}24, MMR^{+}22, NW14, PC19, PHS^{+}24, QXW^{+}23, RGK24, SZS22, SBD16, TZS^{+}21, TYW^{+}23, WCZR23, YCM^{+}21, YLK^{+}21, YČS23, ZG21, ZLO^{+}22, ZTWY19, LH14]. End-to-End [BYT^{+}21a, BYT^{+}21b, BWZC22, BB21, CHWQ24, CMY^{+}22, CLHK21, DQY18, FYT^{+}21, FCCA23, FWT^{+}18, GH23a, HBK24, HFW^{+}22, KLK17, KHE21, LWM^{+}20, LWX^{+}22, LDG^{+}23, LZL^{+}23, MZC^{+}24, MMR^{+}22, PHS^{+}24, QXW^{+}23, RGK24, SZS22, SBD16, TZS^{+}21, TYW^{+}23, WCZR23, YCM^{+}21, YLK^{+}21, YČS23, ZG21, ZLO^{+}22, ZTWY19, LFM^{+}23]. End2End [CQ22]. Ending [LHL^{+}22a, ZHZ19]. Endings [ZHZ19]. Energetic [BKTR24]. Energies [CSS^{+}19]. Energy [CASZ17, ISZA15, LFWZ23, SAC17, SH17b, XWY^{+}23, ZTDD23a, ZM20, ZSIA16]. Energy-Efficient [ZTDD23a]. Energy-Preserving [XWY^{+}23]. Engaging [XZFN23]. Engineering [DM16, QCDW17]. English [CTSC14, CHL^{+}24, HK24, LHC^{+}23, WSY14]. Enhance [HJZB23, ZLF19]. Enhanced [BOSR23, CvdPHH21, CSY^{+}21, GWG^{+}14, HC19, JSG^{+}22, JGHN23, LJL^{+}23, LP22, MZZ23, NR14b, PK15, WFR^{+}21, YHC19, ZZZ16, ZWZ^{+}21, ZMC^{+}24, ZZGD21]. Enhancement [AVM19, BVGV15, BIK^{+}18, BA19, BHP^{+}23, BV23, BB21, CDLL19, CDLL21, CBN^{+}24, CH15, CLL^{+}21b, CWT22, DP16, DBK15, DB18, DTN^{+}14, DBN^{+}17, EMTF18, FBWG23, FD21, FCG24, FWT^{+}18, FJT^{+}23, FSEI21, GN23, GVMGO17, GWQ23, GTV19, GTSM24, GAM23, HBB17, HdHTJ22, HSLB23, ISZA15, JVR16, JLKK21, JY23, JRA17, KGP14, KNBC19, KHK17, KK14, KNH^{+}18, KTJ19, KMG17, KBG16a, KBG18, KG14, KDJJ16, KDJJ17, LG19, LK19, LFO^{+}24, LRWG23, LGGH19b, LY21, LLZ^{+}21a, LYZ^{+}23, LSWN23, LdLvWWS21, LYH^{+}22, LCY^{+}23, LSH14, MWT16, MPAEBLB15, MCZ21, MTZ^{+}21, MC17, MK15b, MK15c, MBK17, NW24, NEN21, NV21, NJC16a, OWW^{+}23, PS19, PW19, PW21, PW22, PW23, PZP19, QDSL19, RG18, RG21, RWL^{+}23, SBN^{+}19, SSSS22, SBM^{+}19, SM18, SLGZ15, SZVZ16, SYG^{+}24, SK18, TOS14, TM14, TD23, TCW19, TW20, TZW21]. Enhancement [TW21, TJCB16, TG21, TG23, TE23, TP19, TWSS16, TDL19, TE16, VBY18, WFN^{+}18, WLL^{+}16, WB18, WDDL18, WTW20, WW22, WZG^{+}23, WWWR23, WRDP17, WSM19, WZS14, XHRC24, XDDL15, XMBN18, XMBN21, YOM21, ZCF14, ZH14, ZWZ^{+}21, ZCDH21, ZTDD23a, ZQN^{+}23, ZWW19, ZZ19, ZAL24, ZZZD23]. Enhancer [CHWQ24, TM16]. Enhancing [CCC^{+}24, EZS14, HHC16, LSG^{+}23a, LMW^{+}24, LFM^{+}23, MCL22, XFH^{+}23]. Enjoyable [MSFK15]. Enroll [LWK^{+}22]. Enroll-Test [LWK^{+}22]. Enrollment [DVO^{+}23, LH14]. Ensemble [AZGN14, GBF21, KWZ24, LM23b, LB16, SV18, TMH^{+}14, ZW16b]. Ensembling [WDDL18]. Entire [MYKO24]. Entity [CXvG^{+}23, CCC^{+}24, FLC^{+}23, FLYJ23, HSZ^{+}18, JLZD24, LWCC21, QZLH21, VSK^{+}24, WZMH22, XHS^{+}18, YJX20, ZCHW24, ZWW^{+}23b, ZCY^{+}22, ZQF^{+}24]. Entity-Level [YJX20]. Entity-Relation [JLZD24]. Entity-Sensitive [YJX20]. Entropy [BWZC22, CDLL21, MK16, NAK^{+}17, SV18, WJ23, WLL15b]. Entropy-Based [WLL15b]. Envelope [AF18, CCC21, DSMD15, DAR16, EMTF18, dSCB24]. Envelope-Based [CCC21, dSCB24]. Envelopes [AY17]. Environment [AC24, CWR^{+}22, CZN^{+}24, FSA18, FNA15, GWQ23, MMBT21, TMH^{+}14, WRK18]. Environmental [SK17, SSKP16, TP23, TMKO17, XHW^{+}17, YND14]. Environments [BHP^{+}23, DBK15, DDMvW20b, DPHG18, FSO17, FZZ^{+}15, GAM23, JLKK21, JWBR21, LSH14, MMB17, MGGK17, NV21, SGH17a, SK16b, SWCN19, VMOL16, XCHX24, ZW17]. Epoch [VP17, VAP19]. Epoch-Synchronous [VAP19]. Equalization [FdDARF16, KD16, KD17, LGGH18, LMV21, PMB^{+}19, PGS^{+}22, PT21, PAB19, RKL14]. Equalizer [LC16b, RVB14, VR19]. Equivalent [DCAS22]. Equivariance [SNK21]. Era [Ano19t]. Ergodic [TP16]. Erratum [Ano19c]. Error [CDLL19, DKDN20, GCJ^{+}14, GCvWMJ14, HBCC22, JT15, KTJ19, LCB^{+}23, OHN16, TNS^{+}18, WsL15, WK16, WTQ^{+}19, YC22, ZLX^{+}24]. Error-Based [HBCC22, YC22]. Errors [AAA19, CH18, MBDV23, NNS16, Non14, SSFL22, SZW23]. ESPRIT [HZF18, WAJL15]. ESPRIT-Type [HZF18]. Essay [MIR^{+}21]. Estimate [Emu18, LHW^{+}21, SGL^{+}21, YHC^{+}20]. Estimated [AVM19, BSW^{+}24, WCW^{+}21]. Estimates [APH14, AM18a, JTHN14, KGH^{+}16, TTH14]. Estimating [HHMK18, KGP14, OHN16, RGR^{+}22, SAC17, SBV17, SCEH19]. Estimation [AF18, AM18a, AMK^{+}20, AB15, AM20, BBM17, BV23, BKTR24, BSTM21, Bor22, CATS20, CH18, CH15, CASZ17, CSY^{+}21, CS19, CTE21, CP21, ÇH21, CL14a, CL17a, DAR16, DDBW20, DDBW23, DGMB23, DDMvW20b, EGMN16, EMTF17, EMTF18, ESJ17, FSA18, FSA20, FWM^{+}14a, FWM^{+}14b, FSEI21, GTS^{+}19, GWLL22, GKC16, GB14, GMD21, HMN17, HAGSZ16, HJC19, HTG14, HGT15a, HMO21, HNF20, HKI23, HZBW23, HZF17, IIY16, JP14, JEvW23, JFG^{+}17, JCBJ15, JT15, JLP^{+}21, JWBR21, JBTL15, KAJC16, KK19a, KK14, KD18, KKGH16, KHHJ19b, KGBPM24, KBG16b, KDJJ16, KDJJ17, LLGG20, LG20, LWZG21, LH23, LMH23, LJWC23, Lin19, LFU^{+}17, LLS24, MLP^{+}24, MPAEBLB15, MKA^{+}14, MD18, MMM23, MYNO23, MSRND14, MCS21, MAT15, MXNB19, MYC^{+}23, MR20, MK15b, MK15c, NJC16a, NJC16b, PS17, PC21, PT14, PZP19]. Estimation [PCLP14, RKN14, RG17, SIUW21, SM14, SD16b, SD16a, SH14b, SGH16, SK15, SLC^{+}15, SFK15, SMO^{+}21, SM18, SY15, SKH^{+}15a, SLGZ15, SZVZ16, SLJ18, TD23, TH17, TP15, Ter16, TP19, TG16, TR15, TVJ15, TMH^{+}14, TE16, VDH24, VSL21, WGD15, WD16, WYH17, WXZ^{+}19, WAJL15, WRK18, XGKM18, XGKM19, XBZ^{+}24, XLLY18, YG24, YLH23, YK21b, ZG16, ZZN^{+}16, ZHH19, ZWW23a, ZM20]. Estimations [HAS21]. Estimator [PSL^{+}16]. Estimators [BKS^{+}18, NCB23, SH14b, SGH17b]. Estonian [ESVK17]. ETEH [CMY^{+}22]. Evaluated [PM15, PBA^{+}23]. Evaluating [BDL15, SCY22]. Evaluation [BBC^{+}16, BvWM18, BKS^{+}18, Bre15, CLW^{+}16, CMG^{+}17, FWT^{+}18, GaKGB21, GLQ21, GTV19, GLW20, JSR16, JEP14, LCZ^{+}19, MLC^{+}23, MVN22, NRČR14, PMA^{+}21, SW20, SMB18, TKH21, TL19, VKH18, WPY^{+}21, WK24b, YLD15, ZLHL21, ZWL^{+}23a, vNKB^{+}23]. Evanescent [MTKS21]. Evanescent-Region-Aware [MTKS21]. EVD [NEWN23]. Event [AGM23, BLLP17, CC21, DWY21, FZZ^{+}15, GMD24, GPF17, HWT^{+}17, HKM^{+}24, JP14, KXS^{+}19, LLR^{+}16, LZ16, LPL^{+}22, LFWZ23, MMCF18, MSB18, MZX^{+}15, MDE^{+}19, PMMM15, PMA^{+}21, RVB^{+}17, SGA15, TSKO20, WWS^{+}18, WDW^{+}23, WSH15, WSG21, rPH^{+}17, Ano16-52, Ano16-53, Ano16-54, Ano16-55, Ano16-56, Ano16-57]. Event-Based [JP14]. Events [CPSFC^{+}14, DH20, FFP^{+}22, JGM23, MMM23, MHB^{+}18, PHM^{+}16, SMA^{+}17]. Evidence [ZZD23]. Evolution [LB16, MTS^{+}19, SWZ16, WLL15b]. Evolution-Strategy-Based [MTS^{+}19]. Evolutionary [HHMK18]. Evolving [YDW^{+}21]. Examining [MDCS20].

Example [MAP14, RAB18]. Examples [CWZ^{+}23, JWH^{+}22, SWH22]. Excitation [EMTF17, EMTF18, JBTA19, SSK17, WFY^{+}22]. Exemplar [BVGV15, BV17, KGP14, LCW24, RA23, TLW^{+}17, WWD14, WVCL14, WCL^{+}21a, YGV14]. Exemplar-Based [BVGV15, BV17, KGP14, RA23, TLW^{+}17, WVCL14, WCL^{+}21a]. Exemplar-Guided [LCW24]. Exemplars [YGS19]. Expanding [DRFO14, GLWQ22, SWH22]. Expansion [CWH^{+}16, GDC19, GH23a, LPHCB14, yLsL14, PC21, YC22]. Expectation [BG23, DG15, DPHG18, LSH14, SGH16]. Expectation-Maximization [BG23, DG15, DPHG18, LSH14, SGH16]. Experts [LZ23, ZDZ^{+}23, Zhe23]. Explain [CLC^{+}22]. Explainable [CLL^{+}21a]. Explicit [DV14b, FB18, SRH15, ZLS^{+}21, vMM14]. Exploit [LPL^{+}22, ZQL22]. Exploiting [BSTM21, CLW^{+}19, DDNN19, FL19, HZW^{+}22, JWH^{+}22, LLWC21, LSB^{+}21, MMB17, MLP^{+}24, ODSM16, PSL17, QYCT23, SWZ16, WWLL14]. Exploits [May18]. Exploration [CHL^{+}23]. Exploring [JYC^{+}24, LLL^{+}22, NTO^{+}23, PM23, SRC^{+}23, TZZ15b, XGKM18, ZZH23, ZWW^{+}23b]. Exponent [LYH^{+}22]. Exponential [CPCBFSF16, FOBP15, HOF15]. Exposure [LLZ23b]. Expression [CBG15, WZH22]. Expressions [WCL^{+}23]. Expressive [CMD^{+}23, LZC^{+}23, LSGL21, MSFK15, QDSL19]. Expressiveness [CBG15]. Extended [AKG^{+}17, GMD21, NAK^{+}17]. Extending [ABvM19, GHAP17]. Extensible [ES16]. Extension [AF18, AF19, HJZB23, LAGD18, LB16, MV23, NIK^{+}22, ZPP15]. Extensively [QL24]. External [ABvM19, BYT^{+}21b, HLB^{+}24, YKPMP18, ZL21a]. Extra [WED17]. Extract [WSZN22]. Extracted [WSY14]. Extracting [HL18, ZL21b].

Extraction [AV16, ADLX23, AY17, BSTM24, CXvG^{+}23, CJY^{+}21, CWJ15, CCC^{+}24, CL14a, DVO^{+}23, FYG^{+}21, GMH14, HPB15, HTG14, HGT15a, HGT15b, HZL23, JLZD24, LZ16, LPL^{+}22, LLHW24, LCRY24, LLW21, LZL^{+}23, LLS24, LLL^{+}19, MAB^{+}16, MYK^{+}24, MAK15, NM21, RMBK14, SAB16, SCCM23, SHZ^{+}23, TH14, TH17, TON^{+}22, VP17, WLT24, WFR^{+}21, XLZ^{+}23, YDZ19, YJX19, ZZJ^{+}23, ZCYZ18, ZW19, ZYD^{+}21, ZQL22, ZMC^{+}24, ZWX15, ZJL21]. Extractive [CLC^{+}15, CLCW18, LCC^{+}15, MSW^{+}22, YW14, ZYW^{+}20]. Extractor [GGL23, ZSB17]. Extreme [HJZB23].

F [LK19]. F0 [AB15, CWJ15, IIY16, KK19a, LCTA19, TGGE21, WTY18]. F1 [HXWS15]. F1-Score [HXWS15]. Face [EMXD20]. Facial [JD17]. Fact [GHS^{+}21, ZZGD21]. Fact-Article [GHS^{+}21]. Factor [BDV^{+}14, CC15, HH14, MAL18, NK21, QTY16, SGL^{+}21, SGLS23, WLLZ17, YHO^{+}17]. Factored [AVK^{+}15]. Factorization [ATA16, BV16, BOL^{+}18, BSER17, CONVVC18, CBG15, CY16, ED18, FSO17, HL15, JLKK21, KHTL18, KSK^{+}15, KOS^{+}16, LKPR18, MH16, MUT^{+}20, MTKS21, PPM^{+}24, WGLZ21, WYY22, ZBCH21, ZXH^{+}16]. Factorized [CL14a, GWL^{+}24, SS16, ZDZ^{+}21]. Factors [HK24, KKJN17, SKAK16, WWLL14, WC16a, ZYG^{+}14]. Facts [ACK21, BYY^{+}23]. Fail [FNM^{+}19]. Fail-Safe [FNM^{+}19]. Fair [SHJY21]. Fake [ZWC^{+}23]. Families [PM23]. Family [HBLC14, TR14a]. Far [ATP15, QCL23, ZSIA16]. Far-Field [QCL23, ATP15]. Farewell [Ano18-32, Den14]. Fast [BF18, BYT^{+}21a, CRH^{+}14, CL14a, Cum15, DAFS15, KWZ24, KKGH16, KBT19, LSNC21, LKM23, LM23a, LFM^{+}23, MOT^{+}23, ODG16, XAHJ^{+}14, YPH23]. FastMVAE2 [LKM23]. Fate [PP19]. FC [NR22]. FC-U2-Net [NR22]. FDAF [GCvWMJ14]. FDTD [HB17, SRH15, vMM14]. Feature [AV16, AATKS21, BY15, BSER17, BSTM24, CWW14, CHJ14, CZZ15, CWR^{+}22, DH14, DGCY23, GMH14, GGZ23, HC19, IO17, KGP14, KKGK18, KNH^{+}17, LFO^{+}24, LWZ^{+}14, LJH^{+}24, LWYD23, LCRY24, LCQ23, LSG^{+}23b, LFWZ23, MAK15, NM21, NXCL16, QCDW17, SAB16, SKDN14, SLWY20, SMvdP22, TALP24, TON^{+}22, TQM^{+}23, VML14, WLT^{+}18, WL19, XHW^{+}17, ZDG14, ZW19, ZLZW21, ZQL22, ZYD^{+}23, ZLJB23]. Features [AFG^{+}23, AVK^{+}15, BPG23, BMVG17, CSJD16, CHWL14, CTLL17, DHSG15, DW17, DB15, FOBP15, GKC16, GMFG16, GPF17, HLL15, HL18, JMD^{+}17, JT15, KTH^{+}20, LGHG17, LWZG21, LLLL19, LSB^{+}21, LPI^{+}21, LKR23, LCTA19, PKA14, RG21, STT^{+}19, SGA15, SMA^{+}17, SAS17, SW20, SKK14, SKM17, TD16, TZZ15b, VAH19, VDH24, VAP19, WYH17, WW19, WSY14, WK16, WSZN22, XGKM18, XYY^{+}23, YK17, YB14, ZM21, ZFC^{+}23, ZTM^{+}23, dlCSVAGAPM15, VZC14]. Feed [ATR^{+}21]. Feed-Forward [ATR^{+}21]. Feedback [AA17, BRD15, BE17, BvWWM17, BvWM18, DSD23, FWM^{+}14a, FWM^{+}14b, PPSG17, SD16b, SD16a, SNTD19, SH17a, SFV23, SLGA15, SP15, TNS^{+}18, XYH14, ZLL^{+}24a]. Feedback-Driven [BE17]. Feedforward [BSB14, CC14a, MMS^{+}20, SNS15, VE24]. Few [CWR^{+}22, DYY^{+}23, ED18]. Few-Shot [CWR^{+}22, DYY^{+}23]. Fewer [HRL^{+}21]. Fibonacci [FM15]. Fidelity [WWX^{+}23, YPH23, YWT23]. Field [AKUS23, ATP16, ADM^{+}17, BGQO^{+}17, BATS21, EJ16, FFSS17, FFSS18, GFS22, KMM^{+}18, KBV16, KBV17, KEzW21, KFH^{+}14, KCD20, KBI^{+}21, LLGG20, LG20, LMGG21, MAZ21, MFH21, MAB^{+}16, MF18, OBP^{+}24, QCL23, UKS19, VAH19, WYO23, WWS^{+}18, WAS16, WWH^{+}18, WSFS19, ZA14, ZHB^{+}18, ZSIA16, ATP15]. Fields [CE15, HNF20, MZA20, NYD17, YLL15]. FifthNet [OS21]. Figure [KTH^{+}20]. Figure-of-Merit [KTH^{+}20]. Filling [LYCX22, MDY^{+}15]. Filter [AM16, AM20, BvWWM17, DV14a, DV14b, DDMvW20a, EV23, FZZ^{+}15, FB18, FB19, GCvWMJ14, HAGSZ16, HHMK18, IISN21, KJ21, KKJN17, KE23, KMG17, Lin19, LLL^{+}23, LSGH24, MWL22, May18, MAK15, MKA16, MJJ^{+}14, MG18, RJW19, SGM^{+}15, SD16b, SH19, SGH15a, SMVW14, SGL^{+}23, SRH15, SBP^{+}24, TE23, TTH14, TCY15, TE16, VP17, WTY20a, WW15, WRKG17, WXL^{+}21, XCZ^{+}23, YLJ16, YPH23, YPL21, YWT23, YHC^{+}20, ZCY16, ZCYZ18, ZTDD23b, ZGZ23, ZDLK18, dCBC23]. Filter-Bank [SGM^{+}15]. Filterbank [CFMM14, SP16, SGA15, YLD15]. Filterbanks [BRC24, CRH^{+}14]. Filtered [CC14a, CZN^{+}24, FdDARF16, ZLX^{+}24, dlCSVAGAPM15]. Filtered- [CZN^{+}24, ZLX^{+}24]. Filtered-x [FdDARF16, CC14a]. Filtering [ARSA14, ABA17, AY15, CH18, CD17, CMG^{+}17, DMVP16, DB18, DB19, ET18, GMH14, GJ17, GS21, GYdL^{+}21, HBCC22, JCBJ15, KFH^{+}14, LRG21, LGGH19a, LG16, MHD15, MHGD15, NNC21, NCC15, OOK17, PZZZ22, MG19, RG15b, SR16a, TD23, TWW22, TH14, TH16, TG21, TGGE21, WXL^{+}21, XBZ^{+}24, XMBN18, XMBN21, ZHH18, ZH15]. Filtering-Based [MHD15]. Filters [AM18b, BGQO^{+}17, BG21, BH18, CCR16, CBS23, DRK18, FD21, FNA15, GEC15, Hel14, HBLC14, JGM23, JBC16, KHCW15, LSNC21, MV14, MZL21b, PGS^{+}22, PT21, SSTS18, TG23, TG24, WCC^{+}21, WST^{+}21, Wid18, Wid19, Yan22, ZG16]. Financial [LPM^{+}23]. Finding [HLL15, ZHZ19]. Fine [FSN24, GGL23, GHS^{+}21, GHJ^{+}23, LJL^{+}23, MLL^{+}22, SHL^{+}21, WZMH22, WY23, YM17, YPC^{+}22]. Fine-Grained [GHS^{+}21, GHJ^{+}23, LJL^{+}23, WZMH22, WY23, YPC^{+}22]. Fine-Tuned [FSN24]. Fine-Tuning [GGL23, MLL^{+}22, SHL^{+}21]. Fingerprint [XO16]. Fingerprinting [HWRC15, ODG16, SW16]. Finite [AB21, AG17, ADGE19, BHBS16, BS14, HNF20, MPKS^{+}23, OVN21, SvWRK15, ZH15]. Finite-Difference [MPKS^{+}23]. Finite-Distance [AB21]. Finnish [ESVK17]. First [BBAS19, BBJ18, CASZ17, LCQ23, MFH21, TC19, TC22, WC16a, Ano15z, Ano15-27]. First- [TC19]. First-Order [BBAS19, BBJ18, CASZ17, TC22, WC16a]. Fisher [ATP15]. Fixed [FB18, LSGH24, SGL^{+}23]. Fixed-Filter [LSGH24, SGL^{+}23]. Fixed-Point [FB18]. Flamenco [KG16]. Flat [HSPK18]. Flat-Start [HSPK18]. Flattening [HHZ^{+}21]. Flesh [KCMW17]. Flesh-Point [KCMW17]. Flexible [ARW^{+}16, KBG16a, PACPJF^{+}23, YLK^{+}21]. Flexibly [ACK21]. Floating [LYH^{+}22]. Floating-Points [LYH^{+}22]. Flow [CLW24, GI15, SHCV23, WC16b, ZCHW24]. Flow-Based [SHCV23]. Fluency [BDL15]. Fluent [ZCHW24]. FluentNet [KHE21]. FluGCF [ZCHW24]. Fly [EDO17]. FM [VS14]. FMLLR [JKU18]. Focusing [ACK21]. Following [LD16]. Footprint [LR17]. Foreground [DP16]. Foreign [BHS^{+}16, MK15a, ZDGO21, ZWTL23]. Forensic [WH22]. Foresight [LLT^{+}21a]. Forests [PMMM15]. Forgetting [CC15]. Fork [PWS^{+}23]. Form [dCBC23]. Formalism [KDY14]. Formant [FWM^{+}14a, FWM^{+}14b, JFG^{+}17, YLY15]. Formant-Based [YLY15]. Formants [SIUW21]. Formulation [Bay15]. Forward [ATR^{+}21, LFQL19, ZTWY19]. Four [WDW^{+}23]. Four-Stage [WDW^{+}23]. Fourier [WCC^{+}21, LYB^{+}23, VH19, XCZ^{+}23].

Fractional [BGQO^{+}17, Hel14, HCB20, MMZ23]. Fractional-Order [Hel14]. Fragment [HLL15]. Fragmentation [WLF23]. Frame [AL23, CHWL14, DAR16, HQTX23, LWZG21, LSL24, LSGL21, SIKN23, TD23, TM22, YCM^{+}21, YsL15, YH21, ZXY^{+}21, ZGO19]. Frame-Level [HQTX23, LWZG21, LSL24, TM22, YsL15, YH21]. Frame-Synchronous [YCM^{+}21]. Frames [OD17]. Framework [BDL15, BZ23, BAS20, CE15, CLCW18, CXX^{+}23, Chi18a, DV23, DDZL19, EDO17, ES16, FYT^{+}21, GCvWMJ14, GHAP17, GJB15, GZG23, JQL24, JKU18, KMM^{+}18, KBG16a, KTH^{+}20, LQW^{+}24, LRG21, LHL^{+}23a, LJX^{+}16, LHYJ23, LSTK21, LHL^{+}22b, LSH14, MV14, MØ20, NNC19, PW19, PW23, PFLX24, QMS16, QWL^{+}19, RARR15, SHCV23, SM18, SHL^{+}21, TQW^{+}15, TJCB16, TGZ22, TR14b, WW16, WTY^{+}23, WZM21, WLL15b, XDC^{+}17, XZSX23, YsL15, ZHCB23, ZYW^{+}20, ZZZD23]. Free [DV14b, FB18, FB19, HSM19, LKC^{+}21, LLC23a, SSFP18, TYW^{+}23]. Frequencies [AHSS^{+}18, AY17, BSR^{+}23, HJC19]. Frequency [AM16, AY15, AY17, BP14, BOSR23, BvWWM17, BCB18, BABC19, CvdPHH21, DSD23, DSH18, EV23, GLWQ22, GJYW23, GWLL22, GCvWMJ14, GZZY23, HSK17, HCBY18, HEMN21, HBC17, HCB18, JP14, JFG^{+}17, JCBJ15, JT15, JWBR21, KJ21, KYI^{+}15, KXS^{+}19, KGH^{+}16, KBG16a, LSWN23, LFWZ23, LM19, MTNR21, MAP14, MOT^{+}23, MGF18, MCD14, NV21, NJC16b, OG14, PDKG24, RBR16, RG15a, RKL14, SUK24, SGLS23, Son21, SKH^{+}15a, SKM19, SLJ18, TM16, TWIN23, TLW^{+}17, TZZ^{+}15a, WCC^{+}15, WZW19, WW17, YOM21, YBC^{+}14, Yan22, YK21a, ZKA15, ZG16, ZTQB19, ZQN^{+}23, ZZX^{+}21, ZSIA16, ZH15]. Frequency-Band [LSWN23].

Frequency-Dependent [DSD23, DSH18]. Frequency-Domain [BvWWM17, BCB18, GCvWMJ14, HCBY18, KJ21, SGLS23, SKM19, TM16, Yan22]. Frequency-Energy [LFWZ23]. Frequency-Invariant [BABC19, HBC17, HCB18, Son21]. Frequency-Resolution [EV23]. Fricatives [XWWW19]. Friendly [EDO17]. Front [Ano14u, Ano14v, Ano14-29, Ano14-30, Ano14-31, Ano14-32, Ano15r, Ano15s, Ano15v, Ano15w, Ano15x, Ano15y, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano17-28, Ano17-29, Ano17-30, Ano17-31, Ano17-32, Ano17-33, Ano17-34, Ano17-35, Ano17-36, Ano17-37, Ano18-33, Ano18-34, Ano18-35, Ano18-36, Ano18-37, Ano18-38, Ano18-39, Ano18-40, Ano18-41, Ano18-42, NW14, Ano14w, Ano14x, Ano14y, Ano14z, Ano14-27, Ano14-28, Ano15p, Ano15q, Ano15t, Ano15u, Ano16m, Ano16n]. Front-End [NW14]. Frontends [SRT23]. Frozen [CWDA17]. FSD50K [FFP^{+}22]. FSPRM [ZLZW21]. Full [BNvdB^{+}20, Cum15, GLC^{+}14, JLZD24, MC17, SIKN23, WCC^{+}23, WY23]. Full- [WCC^{+}23]. Full-Rank [SIKN23]. Full-Sentence [MC17]. Full-Song [WY23]. Fully [BD16, FWT^{+}18, LLC22]. Function [BKK15, CSY^{+}21, CG20, HMDG15, ISZA15, KMG15, LGGH18, LGGH19b, LETJ21, MVH16, MMR^{+}22, MD16, MNFGO19, MR20, PCB14, PAB19, RF19, SAPB15, SDDCC15, SMH14, WYY22, YLL21, ZHH19, ZDD21]. Function-Based [WYY22]. Functional [BSL16, CSAR^{+}14, WSZN21, ZA14, ZKA15]. Functions [APWB23, BHAMR19, FPTJ17, FPTJ18, FFSS17, GSA18, KD17, MBM15, MPKS^{+}23, MDCS20, MAT15, SGH15b, VDC^{+}17, ZLLS16]. Fundamental [AHSS^{+}18, AY17, HJC19, HEMN21, JP14, JCBJ15, KYI^{+}15, KBG16a, MOT^{+}23, NJC16b, SLJ18, YBC^{+}14, ZG16]. Fusing [YYY^{+}23, ZZZ16]. Fusion [Abd18, CXvG^{+}23, CWR^{+}22, FYT^{+}21, GN23, HQTX23, JVR16, LLHW24, LJH^{+}24, LCQ23, LLL24, PS17, QCN^{+}21, SHAGPRZ17, SPM17a, SPM17b, TQM^{+}23, XLZ^{+}23, YMM19, YJX20, ZWS^{+}23, ZDZ^{+}21, ZTM^{+}23, VZC14]. Future [CLW^{+}19, DCW^{+}21, LLZ^{+}23d, ZXSL19]. Future-Aware [LLZ^{+}23d, ZXSL19]. Fuzzy [AATKS21]. FxLMS [LLZ^{+}23a, PCG17].

Gabor [OD17, SGA15]. Gain [AA17, AC24, MAT15, NZL23, Non14]. Gains [HBB17]. Game [hShWsL15]. Gamma [NY21]. Gammatone [JGM23]. GAN [BK24a, CHL^{+}23, KVMV^{+}24, MV23]. Gap [GAM23, WTW20]. Gate [MWT^{+}23]. Gated [FYT^{+}21, GAPGG19, LKC^{+}21, TWZ^{+}18, TCW19, TW20, ZQZH19]. Gating [TB18]. Gaussian [ATP16, BDV^{+}14, BSB14, CDLL19, CY15, DDNN19, GY14, GR14, GHAP17, HIA^{+}17, HGAPB15, JRA17, KK19b, LGR14, MTK^{+}20, NNL15, NCB23, RG18, RBBG15, SKH^{+}15a, SKM19, TZZ^{+}15a, VZC14, YUK21]. Gaussian-Impulse [TZZ^{+}15a]. Gaussianization [Ano19c, MMMDL^{+}14]. GCC [WRDP17]. GCC-NMF [WRDP17]. GCN [YHL^{+}23]. Gender [WDDL17]. General [BZ23, BHBS16, FFSS18, JQL24, LYZ^{+}23, NTO^{+}23]. General-Purpose [NTO^{+}23]. Generality [ZJL21]. Generalization [GAM23, HWW23b, PW22, ZWL^{+}23b]. Generalized [ABvM19, CDLL19, CWR^{+}22, DV14a, DV14b, DTL15, DST^{+}19, JSR16, KPH15, Lin19, MH16, PLZL17, RPL^{+}23, SDDCC15, SR14, SLW15, SKH^{+}15a, SKM19, VS14, XCZ^{+}23, ZDZ^{+}23]. Generalized-Gaussian-Distributed [SKH^{+}15a]. Generalizing [LSLK24, XLLY18]. Generate [CLL^{+}21a]. Generated [EZS14, WWD14]. Generating [LM17, SWH22, WQZ^{+}21, WCL^{+}23, ZZH21].

Generation [AGB^{+}24, ASC^{+}18, BYY^{+}23, BGD^{+}18, BTD^{+}19, BMV^{+}23, CLL^{+}21a, CZWC23, CSZZ24, CLZ^{+}24, CHL^{+}23, DYY^{+}23, EMXD20, HRL^{+}21, HCKW23, cHrLLyL23, HCC^{+}23, HW23, HWW16, HSZ^{+}18, JD17, JSG^{+}22, JWH^{+}22, LSNC21, LXK^{+}21, LCW24, LAGD18, LFQL19, LLW21, LCYW21, LHL^{+}22a, MSFK15, OFF^{+}17, RLK14, SLZ^{+}21, SWC20, SLWY20, SWC^{+}21, WFY^{+}22, WSJ^{+}23, WQ24b, WC16b, WK16, WTQ^{+}19, WZH22, WLF23, XPC^{+}22, XFH^{+}23, XSL^{+}19, YLFL22, YYW^{+}23, YYZY24, YZW^{+}18, ZZZ22a, ZSC^{+}23b, ZCHW24, ZTL21]. Generative [Bor23, CLLD14, CRVB^{+}15, CHL^{+}23, Cum21, CS23, GBH16, HRL^{+}21, KYI^{+}15, LY21, LFQL19, rLcHcC^{+}22, LSTK21, LSGH24, MV23, PKP^{+}21, RWL^{+}23, STS18, STS21, SSAS21, TL19, WHO^{+}21, WHT^{+}21, YLC^{+}24, ZYD^{+}21]. Generative-Retrieval [SSAS21]. Generator [SZW23]. Generic [DDNN19, RRdM16]. Genre [DHD^{+}19, PKA14, RV14, ROL^{+}15]. Genre-Based [RV14]. Geometric [APH14, JBCH23, WSFS19]. Geometrical [KE23]. Geometries [DAMvW15]. Geometry [BWH18, HCY21a, HCY21b, PC21, WCY19, ZCQ^{+}19]. German [SGB22]. Gestures [HM14]. Get [Ano17-68, Ano17-69, Ano15z, Ano15-27]. Get-A-Member [Ano17-68, Ano17-69]. Given [ZZH21]. Glimpses [SMvdP22]. Global [AJK24, ADLX23, BSL15, BBJ18, ET18, FWRJ21, HWG^{+}23, MYT^{+}19, Nos16, XH23, YJX19]. Global-Local [MYT^{+}19]. Global-View [ADLX23]. GlotNet [JBTA19]. Glottal [AP21, ARSA14, ABA17, AJB^{+}18, CMG^{+}17, CG21, DP16, JBTA19, KDY14, KKGH16, LTE^{+}21, MYC^{+}23, MG18, MG19, SR16a]. GMM [DY22, WWD14]. GMM-Based [DY22]. Golay [KMG17]. Gossip [ZH14]. GPT [LCL^{+}23b]. GPU [BGQO^{+}17, LFdDG14, ODG16, SRH15]. GPU-Based [BGQO^{+}17]. Grad [CAFFR23]. Grad-Transfer [CAFFR23]. Gradformer [PFLX24]. Gradient [CAFFR23, HF15, QYCT23, SLZ^{+}21, SBP^{+}24, XCZ^{+}23, YPL21]. Gradient-Weighted [CAFFR23]. Gradients [KWZ24, RG15a]. Gradual [WZLL24]. Grained [GHS^{+}21, GHJ^{+}23, HCC^{+}23, LJL^{+}23, WZMH22, WY23, YPC^{+}22, ZSC^{+}23a, ZZD23]. Gram [MHBT15, WCZ21]. Grammar [ZDC14]. Granularity [LXK^{+}21, LHG^{+}24, PFLX24, TZS^{+}21, WYL^{+}23]. Graph [BLZ21, CBN^{+}24, CMY^{+}21, FLL^{+}23, JBWZ24, KD21, LPM^{+}23, LHL^{+}23b, LK16, LHG^{+}24, QCN^{+}21, SGF^{+}21, WYLZ16, WLLZ17, WYL^{+}23, WWG^{+}24, WZM21, WZ21, WFR^{+}21, XZL^{+}21, XLZ^{+}23, XZL^{+}23, YHL^{+}23, ZWCT15, ZZXH22, ZZY^{+}24, ZLZ^{+}24, ZCY^{+}22]. Graph-Based [KD21, LK16, WZ21, ZWCT15]. Graph-to-Sequence [WZM21]. Graphic [RVB14, VR19]. Graphone [SML^{+}24]. Graphone-Based [SML^{+}24]. Graphs [yLsL14, PHMB18]. Greek [PKR^{+}24]. Grid [ÇOH18, HKHP23, LJL^{+}23, SLJ18]. Grid-Based [HKHP23]. Grid-Tagging [LJL^{+}23]. GridNet [WCC^{+}23]. Grounded [BYY^{+}23, KSL19, TDDN24, XZFN23]. Group [Chi15b, DCAS22, LMV21, LHM21, ONKP16, SZL19]. Group-Delay [LMV21]. Grouped [DSD23]. Grouping [LSDM19, YYY^{+}23]. Growing [WZL^{+}15]. Growth [Li15]. GRTr [SSAS21]. GTrans [YYY^{+}23]. Guaranteeing [SNK21]. Guest [HLL15]. Guidance [RC23]. Guided [AGB^{+}24, BNA24, CDLL21, FZZ^{+}15, HRL^{+}21, KBG^{+}23, LPL^{+}22, LCW24, LCZ23, NIK^{+}22, WZW19, XCHX24, ZZJ^{+}23, ZA14, ZKKC22]. Guitar [AS17a, MFG15, SWEV23].

H [PMB^{+}20]. H-Space [PMB^{+}20].

HAAQI [KA16]. HACN [LFM^{+}23]. Hallucination [SSFL22]. Hammerstein [DvWDM14]. Handling [HBF^{+}16]. Handwritten [GMH17]. Hankel [JP14]. Hard [ZH19]. Harmonic [AKB^{+}16, AB21, BPG23, BAS20, DMVP16, DHP14, ESJ17, EJ16, FNA15, GJ22, HMN17, JTHN14, JC17, KFH^{+}14, LKPR18, MMS^{+}20, MOT^{+}23, MPC^{+}24, MTKS21, MBTB14, MDM15, MK15b, NK21, PSL17, PHK16, SCW21, TOKS14, TOS14, TM16, Ter16, VAH19, WYH17, WZG^{+}23, YK21a, Ziv15]. Harmonic-Net [MOT^{+}23]. Harmonic-Percussive [BPG23, PSL17]. Harmonic-Temporal [NK21]. Harmonic/Percussive [LKPR18, TOKS14, TOS14]. Harmonics [HZF17, HZF18, SAC17, ZKA15, Ziv15]. Harmonics-Based [SAC17, ZKA15]. Hate [LLZ^{+}23c]. HDRS [MMBT21]. Head [AFG^{+}23, APWB23, BKK15, BHAMR19, BBHAR24, GSA18, MMB17, MBM15, MGF18, MPKS^{+}23, PAB19, RF19]. Head-Related [APWB23, BKK15, BHAMR19, GSA18, MBM15, MPKS^{+}23, PAB19, RF19]. Head-Tracked [BBHAR24]. Headline [ASC^{+}18]. Headphones [GdC14, LV22, RG15b]. Heads [LLZ^{+}21a]. Health [KAA^{+}24, ZLW^{+}24]. Hearing [AHH^{+}19, ABKP19, DV23, DBP^{+}24, FWM^{+}14a, FWM^{+}14b, FPTJ17, GvdP23, HdHTJ22, IC19, KNBC19, KPH15, LV22, MWL22, MHD15, MHGD15, NWM16, NNL15, PPSG17, SSFP18, SGM^{+}15, SD16b, SD16a, SNTD19, SP15, TNS^{+}18, WW15, YLJ16, YKPMP18, ZZ22, ZEM18, KA16]. Hearing-Aid [YLJ16, KA16]. Hearing-Loss [DV23]. Height [KFH^{+}14]. Heterogeneous [CL18b, FLC^{+}23, LCZ^{+}17, ZZXH22]. HGEN [ZZXH22]. HHT [MCZ21]. Hidden [DBK15, HBT^{+}21, HL18, KSG24, PCW15, RV17, SS16, SLR16, TP16]. Hierarchical [ARW^{+}16, AMK^{+}20, CXX^{+}23, Chi15a, Chi18b, ÇOH18, DJY^{+}19, FZM24, JHLK18, JLP^{+}21, KM23, LG19, LZC^{+}23, LAGD18, LWCL19, LZL^{+}23, LHG^{+}24, LCG^{+}24, NNC20, NNC21, PLQZ23, RV14, RC23, TWZ^{+}18, TP16, WCY15, WWS^{+}18, WSL^{+}20, WZMH22, WLL^{+}23, WLF23, XZL^{+}21, YLC^{+}24, ZZXH22, ZWHD23, ZLW^{+}24]. Hierarchical-Diving [WWS^{+}18]. Hierarchies [FGQ^{+}15]. Hierarchy [SZX^{+}18, YLC^{+}24, ZQL22]. Hierarchy-to-Sequence [SZX^{+}18]. Hiformer [WLL^{+}23]. High [AY17, BP14, DTDL16, FNA15, GB14, HB17, HCQ24, KLS^{+}21, KKGH16, MCS21, MTS^{+}19, NR14a, RVB14, RF19, WCY15, WDDL18, WCCB19, WWX^{+}23, WFR^{+}21, XNRG15, XPC^{+}22, YPH23, YPC^{+}22, YWT23, YH21]. High-Fidelity [WWX^{+}23, YPH23, YWT23]. High-Order [HB17, MCS21, WCY15, WFR^{+}21, YPC^{+}22]. High-Performance [HCQ24, MTS^{+}19, WDDL18]. High-Precision [FNA15, RVB14]. High-Quality [XPC^{+}22]. High-Resolution [BP14, DTDL16, KLS^{+}21, KKGH16, RF19]. Higher [AS17b, BSR^{+}23, GLWQ22, HPPM24, HF23, SAP14, ZZZ^{+}21, vMM14]. Higher-Lag [AS17b]. Higher-Order [HPPM24, HF23, vMM14]. Highlight [HSZ^{+}18]. Highly [YsL15]. Highway [LR17]. HIM [LZL^{+}23]. Hindi [MMBT21]. Hints [ZHZ19]. Histogram [RG15a]. Histograms [MJ18]. Historical [MV23]. History [CBCL15, KSK21, LCYW21, LLZ^{+}23d]. HLTPR [TDDN24]. HMM [AZK15, BD16, DBLČ20, HSPK18, NYS17, PM15, SDK17, WWD14, YHO^{+}17, ZSB17]. HMM-Based [BD16, HSPK18, PM15, SDK17, ZSB17]. HMMs [SLW15]. HMRFs [AB15]. Hoc [TJCB16, WC18b]. Holistic [ZLHL21].

Holography [DCAS22]. Homogeneous [LG20]. Homomorphic [JY23]. Honey [TON^{+}22]. Hop [ACK21]. Horizon [MØ20]. Hospital [BHP^{+}23]. Hough [DTL15]. Hour [RGK24]. Householder [MWC^{+}23]. Howling [AC24, KPH15]. HPM [LCWC16]. HPSG [LZZP22]. HPSG-Inspired [LZZP22]. HRIR [MAL18]. HRTF [RHvdP^{+}16, TJV24, ZKA15]. HRTFs [BHAMR21, GMFG16, ZGL^{+}20]. HuBERT [HBT^{+}21]. Hull [PC21]. Human [BBC^{+}16, FFP^{+}22, GA20, JZJØ19, NW24, WDD^{+}16, XDH^{+}17]. Human-Labeled [FFP^{+}22]. Humanoid [TR14b, TR15]. Humans [TM21]. Humming [WJ15a]. Hurst [ZCF14]. Hurst-Based [ZCF14]. Hybrid [ALH^{+}21, EV23, LKPR18, LGTG18, LLdCF23, LFM^{+}23, RGK24, WSZN21, WSZN22, XWWW19, ZSL^{+}22, ZLSC23]. Hybrid- [EV23]. Hyperbolic [XCZ^{+}23]. Hypotheses [LAK16]. Hypothesis [HEMN21]. Hysteresis [MTBB15].

I-Vector [CL17b, XLLY18, JKU18].

I-Vectors [LMC18, TMMZ18]. Icosahedral [DGMB23]. idea [Ano15z, Ano15-27]. Ideal [LM19]. Identification [AJK24, Ano19c, AB14, BZ21, CG20, EIPB^{+}19, EV23, HCBY18, HL15, HC19, JSMD18, KD24, LGGH18, LG16, MMMDL^{+}14, PBC18, PMB^{+}20, PBQ15, ROL^{+}15, SH14a, TWC^{+}18, VTN15, WYZ^{+}23, WYHL20, YsL15, ZH19, ZWW14, ZWW15]. Identifying [FDK15, SFEB23, WED17]. IDLMA [HNT^{+}23]. IEEE [Ano14-45, Ano14-77, Ano14-54, Ano14-66, Ano14-78, Ano14-55, Ano14-67, Ano14-79, Ano14-56, Ano14-68, Ano14-80, Ano14-57, Ano14-69, Ano14-46, Ano14-58, Ano14-70, Ano14-47, Ano14-59, Ano14-71, Ano14-48, Ano14-60, Ano14-72, Ano14-49, Ano14-61, Ano14-73, Ano14-50, Ano14-62, Ano14-74, Ano14-51, Ano14-63, Ano14-75, Ano14-52, Ano14-64, Ano14-76, Ano14-53, Ano14-65, Ano14a, Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano14-39, Ano14-40, Ano14-41, Ano14-42, Ano14-43, Ano14-44, Ano14-90, Ano14-91, Ano15-40, Ano15-50, Ano15-41, Ano15-51, Ano15-42, Ano15-52, Ano15-43, Ano15-53, Ano15-44, Ano15-54, Ano15-45, Ano15-55, Ano15-46, Ano15-56, Ano15-47, Ano15-57, Ano15-48, Ano15-58, Ano15-49, Ano15-59, Ano15-28, Ano15-29, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano15-35, Ano15-36, Ano15-37, Ano15-38, Ano15-39, Ano16k]. IEEE [Ano16l, Ano16-32, Ano16-38, Ano16-33, Ano16-39, Ano16-29, Ano16-30, Ano16-31, Ano16-34, Ano16-36, Ano16-40, Ano16-35, Ano16-37, Ano16-41, Ano16-46, Ano16-47, Ano16-48, Ano16-49, Ano16u, Ano16v, Ano16-27, Ano16-28, Ano16w, Ano16x, Ano16y, Ano16z, Ano17-59, Ano17-60, Ano17-61, Ano17-62, Ano17-63, Ano17-64, Ano17-65, Ano17-66, Ano17-38, Ano17-39, Ano17-40, Ano17-41, Ano17-42, Ano17-43, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano17-51, Ano17-52, Ano17-53, Ano17-54, Ano17-55, Ano17-56, Ano17-57, Ano17-58, Ano18-64, Ano18-65, Ano18-43, Ano18-44, Ano18-45, Ano18-46, Ano18-47, Ano18-48, Ano18-49, Ano18-50, Ano18-51, Ano18-52, Ano18-53, Ano18-54, Ano18-55, Ano18-56, Ano18-57, Ano18-58, Ano18-59, Ano18-60, Ano18-61, Ano18-62, Ano18-63, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i]. IEEE [Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano21a, Ano21b, Ano22, Ano16i, Ano16j, Ano16g, Ano16h, Ano17r, Ano17z, Ano17s, Ano17-27, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17a, Ano18v, Ano18w, Ano18x, Ano18y, Ano18z, Ano18-27, Ano18-28, Ano18-29, Ano18-30, Ano18-31, Ano18a, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano19a]. IEEE/ [Ano17-60]. IEEE/ACM [Ano14-45, Ano14-77, Ano14-54, Ano14-66, Ano14-78, Ano14-55, Ano14-67, Ano14-79, Ano14-56, Ano14-68, Ano14-80, Ano14-57, Ano14-69, Ano14-46, Ano14-58, Ano14-70, Ano14-47, Ano14-59, Ano14-71, Ano14-48, Ano14-60, Ano14-72, Ano14-49, Ano14-61, Ano14-73, Ano14-50, Ano14-62, Ano14-74, Ano14-51, Ano14-63, Ano14-75, Ano14-52, Ano14-64, Ano14-76, Ano14-53, Ano14-65, Ano14a, Ano15-40, Ano15-50, Ano15-41, Ano15-51, Ano15-42, Ano15-52, Ano15-43, Ano15-53, Ano15-44, Ano15-54, Ano15-45, Ano15-55, Ano15-46, Ano15-56, Ano15-47, Ano15-57, Ano15-48, Ano15-58, Ano15-49, Ano15-59, Ano16k, Ano16l, Ano16-32, Ano16-38, Ano16-33, Ano16-39, Ano16-29, Ano16-30, Ano16-31, Ano16-34, Ano16-36, Ano16-40, Ano16-35, Ano16-37, Ano16-41, Ano17-59, Ano21a, Ano16i, Ano16j, Ano17r, Ano17z, Ano17s, Ano17-27, Ano17t, Ano17u, Ano17v, Ano17w]. IEEE/ACM [Ano17x, Ano17y, Ano18v, Ano18w, Ano18x, Ano18y, Ano18z, Ano18-27, Ano18-28, Ano18-29, Ano18-30, Ano18-31, Ano18a, Ano19a]. iEmoTTS [ZQZ^{+}23]. II [DV14b]. IIR [PGS^{+}22, Wid18]. Illustration [QDSL19]. Illustrations [QYCT23]. Image [RJCW17]. Images [DHSG15, WQZ^{+}21, WvdHZ^{+}21, ZAL24]. Imaginary [LYB^{+}23]. Imaging [MAST15, MF18, OBP^{+}24]. Imbalanced [YsL15]. Imitation [ZPD19]. IMM [TCY15]. IMM-Unscented [TCY15]. Immersive [HGT15a, KBV17, LLS24]. Impact [ZLL^{+}23]. Impaired [ACB17]. Impedance [BHBS16, SRH15]. Imperceptibility [HGT15c]. Imperfection [CCT21]. Imperfections [TC22]. Imperfectly [LYZ21]. Implant [CFMM14, GH23b, SF14, YLD15, dSCB24]. Implants [CCC21, KM21, SMVW14].

Implementation [DV14b, GdBBS21, HKHP23, LFdDG14, Nos16, SRH15, ZN14]. Implications [SMA^{+}17]. Implicit [MZL21b]. Importance [BKTR24, CB14, RG18]. Improve [CDLL19, CHL^{+}23, HBF^{+}16, KGP14, LSB^{+}21, MLL^{+}22, PW22, RRdM16]. Improved [ATR^{+}21, CWT22, DH14, FFSS17, KMG21, KG14, yLsL14, LLC^{+}23b, PCB15, PHM^{+}17, PPSG17, SF14, SGM^{+}15, TG23, TDL19, VHBS21, YsL15, YND14, ZZS^{+}23]. Improvement [BNA24, CAFFR23, HCJT15, HWW23b, KNBC19, SZS22]. Improvements [BD16, YHC^{+}20, ZSIA16]. Improves [GBF21]. Improving [CLSyL21, DC16, FSN24, GCG21, HWW16, LWZZ16, LCSL19, LKM23, LWYC23, LKR23, LCY^{+}23, LLL^{+}19, MS23, NW15, SRT23, TL19, WJ15a, WCZ21, WK16, WGR^{+}18, WLL^{+}22, YCLS17, ZSC^{+}23a, ZCY^{+}22]. Impulse [ADM^{+}17, BWH18, CCR16, CCTO21, CNFG23, DGCA23, GS21, JEvW23, KB14, KD24, KBT19, MMK21, MCD14, TP15, Ter16, TZZ^{+}15a, VDC^{+}17, YK21b]. Impulsive [MCZ21, NCC15, RBBG15, ZLSC23]. Imputation [RLPK15]. In-Domain [JGK^{+}22, QCL23]. Inactive [JWH^{+}22]. Inaugural [Ano19t, Li15]. Incidence [PCB14]. Incoherent [BABC19]. Incoming [SNTD19]. Incomplete [KMG15]. Inconsistency [ZLW^{+}22]. Incorporate [GZX^{+}21]. Incorporated [LGL^{+}24]. Incorporating [AL23, BYY^{+}23, CWL22, HCJT15, MWT^{+}23, MIR^{+}21, RCR^{+}21, STS18, WTZ18, ZL21a, ZAL24]. Incorrect [LC24]. Increase [KNH^{+}18]. Increased [DVO^{+}23, SK16a]. Increasing [RGR^{+}22]. Incremental [CGL^{+}15, MP16, RLK14]. Independent [CMY16, EZS14, GZG23, HNT^{+}23, JJB16, KCMW17, KOS^{+}16, LWLS24, LW19, LPI^{+}21, LCM18, LLL^{+}23, MMT^{+}19, MTK^{+}20, NIK^{+}22, WHQY19, WTW20, WH22, WDD^{+}16, XH23, YM17, ZSB17, ZKH18, ZM23]. Index [Ano14a, Ano17a, Ano18a, Ano21a, KA16]. Indian [MVN22, PM19]. Indic [PM23]. Indirect [PCW15]. Individual [CB14, GSA18, JK19, KKJN17, ZGL^{+}20, ZYG^{+}14]. Individual-Weighting-Factors [KKJN17]. Individualised [DV23]. Indoor [DCZ19, LYN14, SCW21]. Induced [GA20]. Induction [BCCZ19, CLZ^{+}23, ZDC14]. Inductive [SML^{+}24]. Infant [AHF21]. Inference [AV16, BWH18, Bre15, KWZ24, LZ16, LHL^{+}23a, SEM^{+}14, TGZ22, WZZ^{+}24, YJX19, ZZD23, ZLC^{+}24]. Inferring [KHCW15]. Infinite [MIYO14]. Infinite-Order [MIYO14]. Influence [HK24, MNFGO19, RF19]. Influences [CHC^{+}14]. Information [ARSLN16, Ano14-66, Ano14-67, Ano14-68, Ano14-57, Ano14-58, Ano14-59, Ano14-60, Ano14-61, Ano14-62, Ano14-63, Ano14-64, Ano14-65, Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano14-39, Ano14-40, Ano14-41, Ano14-42, Ano14-43, Ano14-44, Ano15-60, Ano15-61, Ano15-62, Ano15-63, Ano15-64, Ano15-65, Ano15-66, Ano15-67, Ano15-68, Ano15-69, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano15-35, Ano15-36, Ano15-37, Ano15-38, Ano15-39, Ano16-42, Ano16-43, Ano16-44, Ano16-45, Ano16u, Ano16v, Ano16-27, Ano16-28, ACB17, BKTR24, BSL15, CWL22, CLCW18, CBCL15, CCC^{+}24, DMSM21, FDK15, GA20, HWW23a, JZJØ19, JT14, KHK17, KML18, KNH^{+}17, KBG^{+}23, LNNL22, LZC^{+}23, LLL^{+}22, LCY^{+}23, MK15b, NSA21, NM21, PDS^{+}16, QL24, RRdM16, RMBK14, RLK14, SP17, TM14, TM16, TM22, VML14, WJ15a, WLT24, WSH15, YDZ19, YWDQ21]. Information [YKL^{+}23, YMM19, ZW16a, ZYD^{+}21, ZMC^{+}24, ZMZ24, ZDZ^{+}21, Ano14-77, Ano14-78, Ano14-79, Ano14-80, Ano14-69, Ano14-70, Ano14-71, Ano14-72, Ano14-73, Ano14-74, Ano14-75, Ano14-76, Ano15-50, Ano15-51, Ano15-52, Ano15-53, Ano15-54, Ano15-55, Ano15-56, Ano15-57, Ano15-58, Ano15-59, Ano16-38, Ano16-39, Ano16-36, Ano16-40, Ano16-37, Ano16-41, Ano16g, Ano16h, Ano17-57, Ano17-58, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u]. Information-Based [RMBK14]. Information-Theoretic [FDK15, ZMZ24]. Informative [KNH^{+}17]. Informed [AF23, CH18, DM16, FPTJ17, FPTJ18, KKGK18, KBG16a, SFDRB21, SBM^{+}19, TH14, TTH14, WWD14, WED17]. Infused [ZQF^{+}24]. Initial [WDD^{+}16]. Innovation [HNF20]. Innovations [Ano19t, Den14]. Inpainting [MPHM19, MPC^{+}24, PHMB18]. Input [Ano19c, CD17, CHJ14, EV23, HM14, MMMDL^{+}14]. Input-Output [EV23]. Inputs [XZSX23]. Insights [HCB18, SMO^{+}21, TG23]. Inspired [ARW^{+}16, JGM23, LZZP22, LSH^{+}23, LWYD23, MAK15, XGKM18, dlCSVAGAPM15]. Instance [BY15, LZD^{+}23]. Instant [KKGH16, MG18]. Instantaneous [EMTF17, JP14, KGH^{+}16, NJC16b, TTH14]. Instants [DP16, KDY14]. Instrument [AS17a, KM23, MSS17, SCY22, SHG21]. Instrument-Centered [AS17a]. Instrumental [AKG^{+}17, PCW15, VKH18]. Instruments [AGV24]. Integral [RPL^{+}23, SG21]. Integrated [DDMvW20a, GGL23, IBN^{+}18, ZZY^{+}24, ZEM18]. Integrating [BYT^{+}21b, CWUS22, FBWG23, TYW^{+}23, WCC^{+}23]. Integration [AVM19, GJB15, JY23, MKA16, SKDN14, WLZH21, YZW^{+}21, ZZJ^{+}23, VZC14]. Intelligence [WRKG17]. Intelligent [KAA^{+}24]. Intelligibility [AdHTJ16, AdTJ18, BNA24, ECJF21, EZS14, GTSM24, HCJT15, JT14, JT16, JRA17, KNBC19, KHK17, KKK15, KTJ19, LY21, MJZ15, PJTJ24, PK15, SF14, SW20, SZS22, TM14, VKH18, ZSIA16]. Intelligible [LM17]. Intensity [Hac14, HMN17, KKJK23, YWLZ18]. Intent [AGB^{+}24, CJY^{+}22, ZXZZ23]. Intention [TP22]. Intents [SFEB23]. Inter [BV23, CHGS23, LYL21, PM19, TWIN23, WZC23, ZZJ^{+}23]. Inter- [CHGS23]. Inter-Aspect [LYL21]. Inter-Component [BV23]. Inter-Frequency [TWIN23]. Inter-Pair [ZZJ^{+}23]. Inter-Pausal [PM19]. Inter-Rater [WZC23]. Interacting [SGY^{+}24]. Interaction [LZL^{+}23, LHG^{+}24, WWG^{+}24, WWLL14, ZLL^{+}24a]. Interactions [JGHN23, LLL^{+}22]. Interactive [GLL20, LCW^{+}18, ZWW^{+}23b, ZLG^{+}22]. Interaural [CvdPHH21, IC19, MHD15, MD18, PZW^{+}21, RJW19, dCBC23]. Interference [HMDG15, HWW23b, TK14, ZLSC23]. Interfering [DDMvW20a, HdHTJ22]. Interframe [AM18b]. Interleaved [VP21]. Intermittently [IO19]. Internet [LGL^{+}24]. Interpolated [JGSC^{+}22]. Interpolation [ADM^{+}17, APWB23, KBI^{+}21, MCD14, RBBG15, TJV24, WYO23]. Interpretability [PPM^{+}24, WGR^{+}18, ŞUY^{+}18]. Interpretable [YLQZ24, ZZY19]. Interpreting [AFG^{+}23]. Interrelate [CCY^{+}24]. Interval [AATKS21]. Intonation [TG16]. Intra [CHGS23, JYB^{+}15]. Intra-Object [JYB^{+}15]. Introducing [Ano16-46, Ano16-47, Ano16-48, Ano16-49, Ano17-61, Ano17-62, Ano17-63, Ano17-64, Ano17-65, Ano17-66, Ano18-64, Ano18-65]. Introduction [LFH^{+}15, RVB^{+}17, SHKB17]. Intrusive [CSS^{+}19, PSL^{+}16, PJTJ24, VKH18, XZF23, ZFC^{+}23]. Intuitive [CDA^{+}14, SGLS23]. Invariant [BABC19, HBC17, HCB18, KYTJ17, LSW18, LM15, Son21, WYO23, ZWL^{+}23b]. Invasive [BSWP18]. Inventories [WCW^{+}21]. Inverse [ARSA14, ABA17, AKN^{+}16, CMG^{+}17, GJ17, LGGH19a, PT21, MG19, SR16a, SH19, SGL^{+}21, ZF14]. Inversion [DSMD15, HGAPB15, JJB16, PT21, SSSS22, TKT17]. Invertible [HKHP23]. Investigating [BLZ21, DQY18]. Investigation [CSJD16, CD17, CCC21, GA20, LH14, NW14, PCLP14]. Investigations [CWCP17]. Irregularities [GdC14]. ISNMF [MV19]. Isomorphic [LYZ21, TXH^{+}23]. Issue [Ano16-50, Ano16-51, KSG^{+}21, SHKB17, Ano16-52, Ano16-53, Ano16-54, Ano16-55, Ano16-56, Ano16-57]. Italian [MBDV23]. Iterative [CTLL17, JLP^{+}21, LRG21, MCL22, PC21, SOF15, SH19, WHRC16, WCC^{+}21]. IVN [DH14].

Jitter [MCS21]. JMS [ZLW^{+}24]. JMS-QA [ZLW^{+}24]. JoinER [CXvG^{+}23]. JoinER-BART [CXvG^{+}23]. Joining [TSKO20]. Joint [ABA17, AJLW14, ADM^{+}19, BV17, BSW^{+}24, BB21, CXvG^{+}23, CDLX18, CL17a, DDBW23, DDMvW20a, DZF^{+}23, FYT^{+}21, GDC19, HCH23, HKHJS15, IISN21, JCBJ15, JLP^{+}21, KMAB18, KD16, KHHJ19b, LK19, LZC^{+}14, LZ16, LZZP22, LMH23, LLC^{+}24, LLdCF23, MSS17, MYKO24, MAT15, MK16, NW15, PKA14, QTY16, QZLH21, QL24, SOF15, SHJT21, SSSS22, SGY^{+}24, TQW^{+}15, TLWV17, UNI^{+}24, WW16, WED16, WLLZ17, WQ24a, WAJL15, XGKM19, XBZ^{+}24, YG24, YZL^{+}18, YLGG22, ZYF18, ZMC^{+}24, ZLW^{+}24, ZYW^{+}20, ZZZD23]. Jointing [JWBR21]. Jointly [ACK21]. Judgment [LQW^{+}24].

Kaiser [KPH15, SH17b]. Kalman [BvWWM17, BH18, DDMvW20a, DB18, DB19, EV23, MYC^{+}23, NCC15, SGH15a, TCY15, XMBN18, XMBN21]. Karhunen [LPHCB14]. Keeping [Den14]. Keeps [ZYZU22]. Kernel [CLL^{+}21b, DTC16, DTC17, DSH18, GI15, KBI^{+}21, XDC^{+}17]. Kernels [ÇH21, WTHJ14]. Key [DAFS15, YZW^{+}18]. Keyphrase [LLW21, YYZY24]. Keyword [HPB15, HBF^{+}16, LYY^{+}23, LWL^{+}19, LETJ21, WJ15b, YCM^{+}21, YGS19, YČS23]. Keyword-Attentive [LWL^{+}19]. KGAgent [YYZY24]. Kinematic [JJB16]. Kinematic-Independent [JJB16]. KINet [BYY^{+}23]. KLT [RKL14]. Knowing [QCN^{+}21]. Knowledge [BYT^{+}21a, BYT^{+}21b, BYY^{+}23, BK24b, CWUS22, DYY^{+}23, FLL^{+}23, GJB15, HLB^{+}24, HDW^{+}23, IK23, JQL24, LWJ19, yLSY^{+}14, LWC^{+}18, LPL^{+}21, LHL^{+}23b, LZD^{+}23, QVZY24, RPL^{+}23, SGF^{+}21, TP23, WTZ18, WWG^{+}24, WQ24a, XFH^{+}23, XZL^{+}23, XZFN23, YLK^{+}21, YKL^{+}23, ZXSL19, ZLZ21, ZLW^{+}22, ZLZ^{+}24, ZZGD21]. Knowledge-Aware [LWC^{+}18]. Knowledge-Conditioned [DYY^{+}23]. Knowledge-Grounded [BYY^{+}23, XZFN23]. Knowledge-Language [ZLW^{+}22]. Knowledge-Regularized [QVZY24]. Knowledgeable [ZZZ22b]. Kronecker [BG21, CBC19, HBCC22, IBC21, PBC18, SBP^{+}24, WBC^{+}21, WCC^{+}21]. Kullback [SLGZ15]. KWS [CMY^{+}22].

L1 [MBDV23]. Label [CLZ22, JCYwH21, KTH^{+}20, PHM^{+}17, WZLL24, WZMH22, XXS^{+}23, YLC^{+}24, ZWHD23]. Label-Correction [ZWHD23]. Label-Tree [PHM^{+}17]. Labeled [FFP^{+}22, Lin19, MSB18, OVN21]. Labeling [CMY^{+}21, CLZ^{+}24, CJY^{+}22, DBM15, FWRJ21, GCG21, MMM23, MZL21a, PC21, SgL20, WXS24, XWS^{+}21, ZQZH19, ZGC^{+}23]. Labelled [HRL^{+}21, KXS^{+}19]. Labels [Bor22, DVO^{+}23, FCG24, JGK^{+}22, LC24, LCSL19, LB19, MMM23, SKT21, YCLS17]. Ladder [DV14a, DV14b]. Lag [AS17b]. Landmark [EMXD20]. Language [AVK^{+}15, AF23, Ano14-45, Ano14-77, Ano14-54, Ano14-66, Ano14-78, Ano14-55, Ano14-67, Ano14-79, Ano14-56, Ano14-68, Ano14-80, Ano14-57, Ano14-69, Ano14-46, Ano14-58, Ano14-70, Ano14-47, Ano14-59, Ano14-71, Ano14-48, Ano14-60, Ano14-72, Ano14-49, Ano14-61, Ano14-73, Ano14-50, Ano14-62, Ano14-74, Ano14-51, Ano14-63, Ano14-75, Ano14-52, Ano14-64, Ano14-76, Ano14-53, Ano14-65, Ano14a, Ano15-40, Ano15-50, Ano15-41, Ano15-51, Ano15-42, Ano15-52, Ano15-43, Ano15-53, Ano15-44, Ano15-54, Ano15-45, Ano15-55, Ano15-46, Ano15-56, Ano15-47, Ano15-57, Ano15-48, Ano15-58, Ano15-49, Ano15-59, Ano16k, Ano16l, Ano16-32, Ano16-38, Ano16-33, Ano16-39, Ano16-29, Ano16-30, Ano16-31, Ano16-34, Ano16-36, Ano16-40, Ano16-35, Ano16-37, Ano16-41, Ano16i, Ano16j, Ano17-59, Ano17-60, Ano17r, Ano17z, Ano17s, Ano17-27, Ano17t, Ano17u]. Language [Ano17v, Ano17w, Ano17x, Ano17y, Ano17-57, Ano18v, Ano18w, Ano18x, Ano18y, Ano18z, Ano18-27, Ano18-28, Ano18-29, Ano18-30, Ano18-31, Ano18a, Ano19a, Ano19t, Ano21a, ACRS14, BDV^{+}14, BMV^{+}23, CPCBFSF16, CSJD16, CLC^{+}15, CLW^{+}16, CLW^{+}19, CHL^{+}24, Chi15a, CBCL15, CHL^{+}23, DHD^{+}19, DBM15, FOBP15, FLMS16, HLL15, HYI^{+}15, HDW^{+}23, HOF15, JHLK18, JWH^{+}22, JSMD18, JGSC^{+}22, KAA^{+}24, KWZ24, KG17, KTH^{+}20, KCL21, LTWT17, LFH^{+}15, LHL^{+}23a, LCC^{+}15, LCW^{+}16, LLW21, LPI^{+}21, LLC^{+}23b, LHZ^{+}24, MWT^{+}23, MDY^{+}15, MBDV23, MVN22, NLYK24, PM23, QCN^{+}21, RV14, RV17, RLK14, SML^{+}24, SHD14, SWH22, SHDG16, SLK17, SHL^{+}21, SNS15, TWC^{+}18, VTN15, WZL^{+}15, WsL15, WDW^{+}22, WZLL24, WWX^{+}24, WZC23, WSH15, WTQ^{+}19, WCL^{+}23, XWY^{+}23, XYH^{+}21, YsL15, YTWB19, YCL22, ZZZ16, ZH18, ZZY19, ZMX^{+}21, ZLW^{+}22, Zhe23, ZWX15, ZTL21]. Language-Adversarial [YTWB19]. Language-Independent [LPI^{+}21]. Language/Dialect [ZH18]. Languages [DCLHJ18, FLC^{+}23, KMG21, NNC20, PM19, TGZ22, ZP15]. Laplace [Chi15b]. Large [CL14b, DB15, ESVK17, GaKGB21, JHLK18, SAP14, SFIL17, TB18, WBSG18, ZCY^{+}22]. Large-Scale [CL14b, GaKGB21, ZCY^{+}22]. Laryngeal [DB17]. Laryngectomy [MHD^{+}17]. Lasso [RBR16]. Lasso-LS [RBR16]. Late [BRD15, BKS^{+}18, CS19, DDBW20, DDBW23, GSP23, HCJT15, KD18, VP21, XGKM19]. Late-Reverberation [VP21]. Latency [CK19, EV23, WWWR23, ZPW23]. Latent [GN23, HM14, JPX^{+}23, KGG19, LWCL19, LCL^{+}23b, PKP^{+}21, RV14, RCL18, SGL^{+}23, WTHJ14, WSH15, YPC^{+}22, ZWCT15]. Latent-Domain [JPX^{+}23]. Latent-State [LCL^{+}23b]. Lattice [CM18, LCW^{+}16, TYW^{+}23]. Lattice-Free [TYW^{+}23]. Law [SAB16]. Layer [CLLD14, GYXY23, QCN^{+}21, SS16, XSC^{+}23]. Layer-Level [GYXY23]. Layer-Wise [CLLD14]. Layers [GN23, TYZ^{+}22, YYY^{+}23]. Layouts [CT14]. LCMV [HDG16, MGGK17, SGH17c]. LCMV-TRINICON [MGGK17]. Lead [MFG15, RLS^{+}18]. Leaderboard [GLW20]. Leak [SGL^{+}21]. Learn [YLC^{+}24]. Learnable [CWZ^{+}23]. Learned [HNT^{+}23, LLC23a, MMT^{+}19, ZW19]. Learner [MBDV23, TJV24]. Learners [LCSL19]. Learning [AZK15, AGC23, BXY14, BHP^{+}23, BHH22, BSR17, BPG23, BSER17, BNA24, BZ21, BG23, CWDA17, CBS23, CDLL19, CJ15, CM15, CCT^{+}19, CHL^{+}19, CLL^{+}21a, CZWC23, CJY^{+}22, CY16, Chi18b, CWBv19, CTLL17, DVO^{+}23, DCYY22, DZLGO20, DCLHJ18, DVLH17, DDZL19, DKDN20, FOBP15, FZM24, FGQ^{+}15, GFG^{+}14, GWQ23, GHS^{+}21, GH17, GvdP23, GCG^{+}17, GAM23, GMFG16, GZG23, GGZ23, GWSC23, HK24, HWW^{+}15, HCQ24, HRL^{+}21, HBK24, HLB^{+}24, HWW23a, HBT^{+}21, HXL^{+}21, ISP19, JG19, JKC^{+}24, KKGK18, KLCL24, KSU^{+}19, KG19, KHE21, KSG24, KBG^{+}23, yLSY^{+}14, LCW^{+}18, LK19, LYY^{+}23, LYZ21, LPL^{+}22, LHL^{+}23a, LLHW24, LCW24, LM22, LFL^{+}23, LM23a, LWY^{+}23, LCRY24, LK16, LLyL21, rLcHcC^{+}22, LB19, LSTK21, LLZ^{+}23c, LSGH24, MHZ^{+}24, MCL22, MP16, MLC^{+}23, MTZ^{+}21, MMR^{+}22, MZL21b, NNC20, NLL^{+}18, NNL15, OCS21, PKP^{+}21, PEO^{+}20, PLQZ23, PHM^{+}16, QVZY24, QL24, Rak17, RH18, RARR15, RG18]. Learning [RV17, RC23, SP16, STS21, SSFP18, STT^{+}19, SSFL22, SGY^{+}24, SK16b, SCY22, SgL20, SG21, SYKL21, SCEH19, SWC20, SHZ^{+}23, SLR16, TMM18, TW20, TZW21, TW21, TR14a, TDL19, VRN18, WsL15, WED17, WDDL18, WC18a, WW19, WHQY19, WZW19, WYL^{+}23, WZM21, WBSG18, WGR^{+}18, WZH22, WLL^{+}22, XHRC24, XLLW21, XDC^{+}17, XHW^{+}17, XCHX24, YLL21, YYZY24, YTWB19, YTF^{+}23, YH17, YWW^{+}23, ZM21, ZYL^{+}21, ZFC^{+}23, ZCDS15, ZW16b, ZW17, ZH18, ZS20, ZMX^{+}21, ZZZL21, ZZXH22, ZSC^{+}23a, ZXZZ23, ZWL^{+}23a, ZMC^{+}24, ZWW19, ZYW^{+}24, ZMZ24, ZXH^{+}16, ZJL21, ZLG^{+}22, ZZZW23, ZZZD23, ZQF^{+}24, ZB14, ZPP15, ZTC23]. Learning-Based [CDLL19, GvdP23, GAM23, HBK24, KKGK18, LK19, LPL^{+}22, SSFP18, SK16b, WZW19, XHRC24, ZFC^{+}23]. Least [BBM17, EIPB^{+}19, KJ21, LH23, LZHN14, SOF15, SD16b, SH19, ZZH16, ZGL^{+}23, ZLX^{+}24]. Least-Mean- [ZGL^{+}23]. Least-Mean-Square [ZZH16]. Least-Squares [BBM17, EIPB^{+}19, KJ21, SOF15, SD16b, SH19]. Least-Squares-Based [LH23]. Lecture [yLSY^{+}14]. Lectures [YsL15]. Legal [GHS^{+}21, LQW^{+}24]. Legendre [CCR16, ÇH21]. LegoNN [DOL^{+}23]. LEGOs [ED18]. Leibler [SLGZ15]. Lemma [KSU^{+}19]. Length [CC14a, SIUW21, Yan22]. Less [vNKB^{+}23]. Level [AL23, CvdPHH21, CONVVC18, CHJ14, CWU^{+}19, CZWC23, CC14b, DTN^{+}17, DY22, FNA15, GWLL22, GYXY23, HQTX23, KYTJ17, LWZG21, LHL^{+}23a, LSL24, LYL21, LLCT22, MASC23, MHZ^{+}24, OGZ18, QZLH21, SFDRB21, SGH17b, SgL20, Sug22, TQW^{+}15, TM22, WWG^{+}24, WZ21, YsL15, YH21, YJX20, ZZJ^{+}23, ZL21b, ZDZ^{+}21, dCBC23]. Level-Dependent [MASC23]. Levels [GB14]. Leverage [QCN^{+}21]. Leveraging [BJ24, CLC^{+}15, DSH18, LB23, LCY^{+}23, XYH14, YW14, YKL^{+}23]. Lexical [QLL^{+}21, QLZ^{+}21]. Lexically [LFQL19]. Lexicon [BCCZ19, CLZ^{+}23, DJY^{+}19, KJG16, ZWCT15]. LF [HXL^{+}21, HXC^{+}22, LTE^{+}21]. LF-MMI [HXL^{+}21, HXC^{+}22]. Library [Ano15-28, Ano15-29]. LID [JSMD18]. LID-Senones [JSMD18]. Lightweight [LJH^{+}24, LHM21, XZSX23, YPH23]. Likelihood [CWJ15, CWJ16, CP21, CG16, HH14, KDJJ16, KDJJ17, LGHG17, LMH23, MH18, SHAGPRZ17, TWIN23]. Limited [HLL15, HOF15, NNC19, SDK17, ZZZ^{+}24]. Limits [KM21, MK15c]. Limits-Potential [MK15c]. Line [TM16]. Linear [ARSA14, AS17b, BSB14, BLLP17, BG21, BSL15, BBAS19, BH18, CCT21, CMM^{+}23, CS19, CPL14, DV14a, DHSG15, DST^{+}19, DDMvW20a, FFSS18, GR14, GCJ^{+}14, HBK24, HTG14, HBCC22, JBC16, JHW^{+}21, JvWGD15, KKK15, LHJ^{+}23, MAP14, MH18, MMZ23, NXCL16, OOK17, PBC18, PCB16a, MG19, SAB16, TG23, TG24, WYY24, XBZ^{+}24, XYY^{+}23, YOM21, YK14, ZTM^{+}23, ZXN^{+}21]. Linear-in-The-Parameters [BG21]. Linearity [ATR^{+}21]. Linearly [ABKP19, KSHH18, MHGD15, RMBK14, ZDD21]. Linearly-Constrained [ZDD21]. Lines [HL14]. Lingual [ASC^{+}18, CHWL14, DKDN20, FL19, FLC^{+}23, GHJ^{+}23, HZW^{+}22, JKC^{+}24, LYZ21, LHC^{+}23, LLC^{+}24, LJX^{+}16, LFL^{+}23, LGR14, PM15, SDK17, TXH^{+}23, WQ24b, ZTL21, ZWTL23, ZLL^{+}24b]. Linguistic [DFY^{+}24, DBM15, LWC^{+}18, NNC19, WSL^{+}20, ZLD20]. Linguistics [OL15, ZWTL23]. Link [CSAR^{+}14, FLL^{+}23]. Linked [YKPMP18]. Linking [FMB16, ZQF^{+}24]. Links [ACK21]. Lip [AF23, MP20]. List [Ano14-81, Ano15-70, Ano17-67, Ano18-66]. Listener [KGBPM24]. Listening [BA16, FJT^{+}23, LLW^{+}19, LSDM19, NV21, RG15b, TM21, TP22, WHB^{+}21]. Listwise [ZLLS16]. Lite [CWT22]. Live [JGSC^{+}22, Sug22]. Liveness [STH^{+}18]. LM [SHL^{+}21]. LMD [CWZ^{+}23]. LMS [CC14a]. Load [VSL21, WSZN21]. Local [ATS18, AJK24, BSL15, CMD^{+}23, CC15, GMD24, HWG^{+}23, LFdDG14, MYT^{+}19, SMvdP22, WAS16, WWH^{+}18, YK17]. Local-Global [AJK24]. Localisation [VHBS21]. Locality [GQXZ19]. Localization [AKB^{+}16, APH14, ADM^{+}19, BHP^{+}23, BOSR23, BOL^{+}18, BJ24, BAS20, CvdPHH21, CBA^{+}15, CSY^{+}21, CPSFC^{+}14, ÇOH18, DCZ19, DMHZ22, DHSG15, DG15, DPHG18, FPTJ17, FPTJ18, GJYW23, GSA18, GvdP23, HGYN23, HWRC15, JGM23, KK23, KGBPM24, LGTG16, LGTG17, LGHG17, Lin18, MMB17, MGB18, MAZS22, MBM15, NR14a, PS19, PZW^{+}21, PLZL17, PHK16, PMA^{+}21, RJCW17, SDF18, SCW21, SPM17a, SPM17b, SSS18, TWSS16, TSKO20, VAH19, WRC16, WHRC16, WZW19, WDW^{+}23, WRKG17, YLPL19, YLL21, YZCW21, YZLY17, ZEM18]. Localization-Driven [BHP^{+}23]. Localizing [JC17]. Locally [DHSG15]. Locally-Linear [DHSG15]. Location [AM18a, BSTM24, LCW^{+}21, SKDN14]. Location-Relative [LCW^{+}21]. Loève [LPHCB14]. Log [DLG18]. Logarithmic [LKHU14]. LOGEN [DYY^{+}23]. Logic [BM15, WCL^{+}23, Zhe23]. Logic-Based [BM15]. Logical [DYY^{+}23, KYL24, LLC^{+}23b]. LogiQA [LLC^{+}23b]. Logographic [NNC20]. Lombard [JRA17, KH21]. Long [CBCL15, DCYY22, EV23, GWL^{+}24, LCYW21, LHG^{+}24, MKMDM17, NR14b, PDS^{+}16, PHMB18, XXS^{+}23, YB14, ZLJ^{+}17]. Long-Content [GWL^{+}24]. Long-Tailed [DCYY22, XXS^{+}23]. Long-Term [MKMDM17, YB14, ZLJ^{+}17]. Lookup [HCZC24]. Loop [DV14b, FB18, FB19, XYH14]. Loss [ACB17, DV23, DBP^{+}24, JZJØ19, LC16a, LWL^{+}19, LSGL21, LCYW21, LETJ21, MMR^{+}22, NWM16, WW22, WJ23, XWZ16]. Loss-Augmented [XWZ16]. Losses [WK24a, ZWTL23]. Lost [WZ24]. Loudness [JLP^{+}21, LSW18, Sug22]. Loudness-Invariant [LSW18]. Loudness-Level-Chasing [Sug22]. Loudspeaker [ATR^{+}21, BA16, BBJ18, DvWDM14, FNA15, GEC15, GFS22, HF21, HSK17, KBV16, MFH21, OFF^{+}17, SSTS18, VDM15, ZA14]. Loudspeaker-Based [HF21]. Loudspeaker-Room [BA16]. Loudspeaker-Signal-Domain [VDM15]. Loudspeakers [LCL23a, MMZ23, MJJ^{+}14]. Low [AM16, BF18, BIK^{+}18, CvdPHH21, CBA^{+}15, CLAG16, CWW14, CM15, CSS23, CHL^{+}24, CRH^{+}14, CK19, CPSFC^{+}14, DZF^{+}23, EIPB^{+}19, EV23, FWM^{+}14a, FWM^{+}14b, FLYJ23, GQXZ19, Hel14, HZW^{+}22, HOF15, JEvW23, JKC^{+}24, KAA^{+}24, KSHH18, KSG24, LJK17, LHL^{+}23a, LMW^{+}24, LHZ^{+}24, LGR14, MGF18, MCD14, MTK^{+}20, MYC^{+}23, PKA14, QYCT23, QZ22, SML^{+}24, SV18, SDF18, SMVW14, SLC^{+}15, SKH^{+}15a, SKH^{+}15b, SLGZ15, TXH^{+}23, VH19, WWWR23, WYY24, WQ24b, WW15, WSZN22, XYH^{+}21, YLD15, YTWB19, YGS19, ZN14, ZPW23]. Low-Bit [XYH^{+}21]. Low-Bit-Rate [LJK17, SKH^{+}15b]. Low-Bitrate [BF18]. Low-Complexity [CBA^{+}15, SDF18, SLC^{+}15, VH19]. Low-Cost [CPSFC^{+}14, KSHH18, ZN14]. Low-Delay [CRH^{+}14, SKH^{+}15a, SKH^{+}15b]. Low-Frequency [MGF18]. Low-Latency [ZPW23]. Low-Order [MYC^{+}23]. Low-Pass [Hel14]. Low-Rank [BIK^{+}18, EIPB^{+}19, GQXZ19, HOF15, JEvW23, MTK^{+}20, PKA14, QYCT23, SLGZ15, WYY24, SMVW14]. Low-Resource [CM15, CSS23, CHL^{+}24, DZF^{+}23, FLYJ23, HZW^{+}22, JKC^{+}24, LHL^{+}23a, LMW^{+}24, LHZ^{+}24, LGR14, QZ22, SML^{+}24, SV18, WSZN22, YTWB19]. Lowband [AF19]. Lower [TC22]. LPC [WCY15]. LS [RBR16]. LSBert [QLZ^{+}21]. LSTM [CSS^{+}19, CH16, DZWC20, DG23, HWT^{+}17, JGSC^{+}22, KCMW17, LWL^{+}21, MXH21, SgL20, SSK17, SNS15, WYLZ20, XLL^{+}19]. LSTM-Network [CSS^{+}19]. LVCSR [CH16]. Lyric [MSEC21, WJ15a]. Lyrics [CWJ16, GGL23, SFDRB21, ZLL^{+}24a].

M [LCRY24]. M3S [WYL^{+}23]. MA [HDLL23]. Macaron [CC21]. Machine [BHH22, BK24b, BBC^{+}16, BY15, CZY^{+}18, CWU^{+}19, CWUS22, DZD^{+}24, DFY^{+}24, DCW^{+}21, DZZ23, GCG^{+}17, GZX^{+}21, GYXY23, HK24, HSZ^{+}18, LLW^{+}19, LWC^{+}18, LCZ^{+}19, LLT^{+}21a, LLZ^{+}21b, LLJ^{+}23, LWL^{+}22, LWCL19, LYCX22, LLZ23b, LLL^{+}22, LZ23, LLZ^{+}23d, MSW^{+}22, MHZ^{+}24, MHPB15, MVN22, NTM16, NTY19, NY21, NNC19, NSNN20, QVZY24, RG18, SZX^{+}18, TYZ^{+}22, TZZ15b, WZL^{+}15, WTZ18, WUF^{+}18, WXZ^{+}19, WZZ^{+}18, XWZ16, XZW15, YYY^{+}23, ZLLS16, ZXSD17, ZZLZ19, ZXSL19, ZLS^{+}21, ZWL^{+}23a, ZZZ22b, ZKKC22]. Machine-Learning [HK24, RG18]. Machines [CLC^{+}22, CL14b, NTA15, TM21, TP22]. Magnitude [APWB23, BIK^{+}18, LGGH19a, LM19, PBS17, SLY14, YWDQ21]. Magnitude-Corrected [APWB23]. Magnitude-Phase [YWDQ21]. Mainlobe [TC19]. Management [CCT^{+}19, WLL15b]. Manager [GY14]. Mandarin [CHC^{+}14, CWCP17, HWW16, LCSL19, LHC^{+}23, LCWC16, SLWY20, XWWW19, YTF^{+}23]. Manifold [BG23, CJY^{+}22, GMFG16, LGTG16, SV17, TR14a]. Manifold-Learning [BG23]. Manifolds [LGTG17]. Manipulation [EMTF17]. Many [ATA16, DTN^{+}14, HSM19, KHT^{+}21, LCW^{+}21]. Many-to-Many [ATA16, HSM19, KHT^{+}21]. MAP [MSD14, TMH^{+}14]. MAP-based [TMH^{+}14]. Mapping [BE17, CAFFR23, GHAP17, HWW^{+}15, KG19, LWZ^{+}14, LLZ^{+}21a, MDCS20, PS17, PLZL17, SSSS22, TW20, TZS^{+}21, WWW21, ZSH20]. Mappings [LYZ21, TKT17]. Maps [DTC15, MAB^{+}16, WCC^{+}15]. Margin [WK24a]. Marginal [WXZ^{+}19]. Marginalization [ET18]. Marginalized [ZCY16]. market [Ano15z, Ano15-27]. Markov [DBK15, NYD17, PCW15, RV17, TP16, YSCZ15]. Mashups [DHYG14]. Mask [BA19, CWZ^{+}23, FSEI21, ODNA23, PZP19, RBBG15]. Mask-Based [ODNA23]. Masked [HBT^{+}21, JT16, KLCL24]. Maskers [JT16]. Masking [BB21, CCL^{+}21, DW17, DRK18, FZM24, JY23, KML18, LS14a, LM19, MTNR21, NDS^{+}23, WZW19, WWW16, WW17, YK21a, ZH15]. Masking-Based [DW17]. Masks [HKHJS15, LK19]. Matching [AKUS23, GLQ21, GLL20, LWJ19, MKA^{+}14, MS23, OFF^{+}17, SSTS18, UKS19, WZS14, YGV14, YXX^{+}22, ZSC^{+}23b]. Matching-Aggregation [LWJ19]. Math [LGL^{+}22, WZH22, ZZXH22]. Matrices [LHD16, MUT^{+}20, ZF14, ZDLK18]. Matrix [ATA16, BSER17, CONVVC18, DVLH17, ED18, HNT^{+}23, HL15, JP14, JLKK21, KHTL18, KSK^{+}15, KOS^{+}16, LKPR18, MMT^{+}19, MUT^{+}20, MTK^{+}20, MXNB19, NEN21, NZL23, PPM^{+}24, PT21, SGH16, TH17, WGD15, WGLZ21, WYY22, YLH23, ZTQB19, ZXH^{+}16]. Matter [YLC^{+}24]. max [SD16a]. Maximal [KTH^{+}20]. Maximization [BG23, DG15, DPHG18, LGHG17, LSH14, SGH16, WD16]. Maximization-Based [WD16]. Maximize [KNH^{+}17]. Maximizing [LCL23a]. Maximum [CLL^{+}21b, CP21, CG16, HH14, HBLC14, HXWS15, KLK17, KDJJ16, KDJJ17, LMH23, LMC18, LHJ^{+}23, MH18, PBC16, SHAGPRZ17, TWIN23]. Maximum-a-Posteriori-Based [KLK17]. Maximum-Likelihood [CP21, MH18]. MaxNSR [WGD15]. MDN [DY22]. Mean [CLL^{+}21b, JT15, KTJ19, LMC18, RLPK15, SKSD14, ZZH16, ZGL^{+}23, ZLX^{+}24]. Mean-Square [JT15, KTJ19]. Meaningful [SLZ^{+}21]. Means [DDMvW20b, EZS14, SIUW21]. Measure [AKG^{+}17, CDLL21, CG21, LCC^{+}15, MJZ15]. Measurement [CCR16, CCTO21, FNA15, MAL18, SP17, XZF23]. Measurements [BKK15, Hac14, KMM^{+}18, KMG15, RF19]. Measures [CPGL24, DAGS^{+}18, FDK15, MVH16, SSFP18, TM14, TKH21, ZKDS18]. Mechanism [HCKW23, LWL^{+}19, LLT^{+}21a, LLC^{+}24, LHYJ23, OGZ18, PLQZ23, WPY^{+}21, WZH22, ZXN^{+}21]. Mechanisms [LWL^{+}21, SRC^{+}23, WLL^{+}23]. Media [KAA^{+}24, XHS^{+}18, ZZH23]. Medical [LXK^{+}21]. Meeting [BSL16, FMB16, TS16, YB14]. Meetings [BSL15, DMSM21, vNKB^{+}23]. Mel [DP16, JT15, LKHU14, YHO^{+}18]. Mel-Cepstrum-Based [YHO^{+}18]. Mel-Frequency [JT15]. Melody [CWJ15, RDP14, WJ15a, ZCYZ18]. Member [Ano15-28, Ano15-29, Ano17-68, Ano17-69, Ano17-68, Ano17-69]. Memes [LGL^{+}24]. Memory [CL14a, GWG^{+}14, GGZ23, HDLL23, HKPR21, LCQ22, LYL21, OOK17, PDS^{+}16, QVZY24, SWC^{+}21, XSL^{+}19, ZYG^{+}14]. Memory-Aware [HDLL23]. Memory-Enhanced [GWG^{+}14]. Mental [ZLW^{+}24]. Mention [ZZJ^{+}23, ZB14]. Merged [NYS17]. Merged-Output [NYS17]. Merit [KTH^{+}20]. Mesh [GDM18, SMH14]. Meshes [SMO^{+}21]. Meta [BHP^{+}23, CBS23, CHL^{+}23, LM23a, PKP^{+}21, TJV24, ZYL^{+}21, ZWL^{+}23b, CBS23]. Meta-AF [CBS23]. Meta-Generalization [ZWL^{+}23b]. Meta-Learner [TJV24]. Meta-Learning [CBS23, LM23a, PKP^{+}21, ZYL^{+}21]. MetaEx [CHL^{+}23]. MetaEx-GAN [CHL^{+}23]. Metaphor [SGF^{+}21]. Meter [GKC16]. Method [AGW15, BBFS21, CCTO21, CMM^{+}23, CCC21, DV14b, DCAS22, DTC16, DTC17, EMXD20, FFSS18, GFS22, GWQ23, GLQ21, GCvWMJ14, HF21, HGYN23, HZL23, JP14, JFG^{+}17, KJ21, KKGH16, LCLM18, LYZ^{+}23, LSG^{+}23b, MF18, MGF18, MCF21, MPKS^{+}23, MGT15, OFF^{+}17, PT21, PCLP14, PBS17, SR16a, SAA15, SK16b, SLZ^{+}21, SBP^{+}24, TG16, TNS^{+}18, UKS19, WCY15, WAJL15, XNG^{+}14, XNRG15, XHRC24, ZW16b, ZZX^{+}21, ZZX^{+}23a, ZH15]. Methodology [CATS20, PM15]. Methods [ABvM19, ALH^{+}21, GaKGB21, GPF17, HB17, JVR16, JCBJ15, JK15, KBV16, KML18, LFH^{+}15, LWL^{+}21, LCW^{+}16, O’S24, RDP14, RJCW17, TON^{+}22, WLL^{+}15a, WAS16, WXS24, YZLY17]. Metric [GSA18, LCZ^{+}19, LY21, MMR^{+}22, MVN22, SF14]. Metrical [KHCW15]. Metrics [BDL15, FWT^{+}18, GTSM24, VKH18]. METTS [ZLL^{+}24b]. MFMVDR [FD21]. MGM [Ano17-68, Ano17-69]. Micro [DB15]. Micro-Modulation [DB15]. Microcanonical [KDY14]. Microphone [ABB^{+}23, AR16, AFdDG23, BC18, BCB18, CBA^{+}15, CM18, CZ17, CASZ17, CCT21, CC14b, DDMvW20a, ED18, ECH23, FSA18, FSEI21, GvdP23, GS21, Hac14, HGYN23, HAGA21, HCY21b, HBC17, HCB18, HCB20, IO17, JEP14, JHW^{+}21, KK14, LWZ^{+}14, LH23, LMH23, LAP23, MVH16, MXNB19, NR14a, NHK16, Non14, OBP^{+}24, OISO14, PCB15, PCB16a, PCBS19, PC19, SSTS18, SNTD19, SGH15b, SCB21, SKDN14, SPM17a, SPM17b, SBVM16, TD23, TZW21, TH14, TJCB16, TP15, TCY15, TR14b, TR15, TNS^{+}18, TC19, TC22, TE16, VH19, WD16, WC18b, WCCB19, WBC^{+}21, WC16a, WRKG17, YZCW21, YKPMP18, YLL15, ZCY16, ZCHH18, ZCDH21, ZLLZ23, ZBC14, ZHCB21, ZHCB23, WWW21]. Microphones [ABvM19, BBAS19, LGTG17, MAZS22, MH16, STH^{+}18, SAP14, SFK15, WRC16, WHRC16]. Microwave [BSWP18]. Mid [MMCF18]. Mid-Term [MMCF18]. MIDI [WL19]. MIDI-Trained [WL19]. Min [SD16a]. Min-max [SD16a]. Minimal [MSD14]. Minimax [Non14]. Minimising [SZW23]. Minimization [CY15, GCJ^{+}14]. Minimum [ABKP19, FJT^{+}23, JT15, RMBK14, WGLZ21, WST^{+}21, WK16, ZPØ^{+}21]. Minimum-Volume [WGLZ21]. Mining [GLT23, LLWC21, XLG^{+}23]. Misarticulated [MP20]. Misequalization [LC16b]. Mises [ATP15]. Mismatch [CS23, LWK^{+}22, SH14a, SAB16]. Mismatched [DCLHJ18]. Mismatches [BC18, WC16a]. Mispronunciation [HXWS15, LCSL19, PLT^{+}18, QMS16]. Missed [GA20]. Missing [IO19, NSNN20, WED17]. Mitigating [LC24, NWM16]. Mixed [BATS21, CM18, CHJ14, DLY^{+}19, GDC19, HLL15, HYI^{+}15, LSW18, TZZ^{+}15a, XYH^{+}21]. Mixed-Bandwidth [GDC19]. Mixed-Order [CM18]. Mixed-to-Panned [CHJ14]. Mixed-Units [DLY^{+}19]. Mixing [AJLW14, HYI^{+}15, LBR18, SH17a]. Mixture [BDV^{+}14, CWYC17, CY15, DH14, GR14, GHAP17, HSN18, HIA^{+}17, HGAPB15, LMC17, LM22, LGR14, LZ23, MPC16, WDDL17, ZGZ23]. Mixtures [BP14, BV16, CH23, DHP14, HJC19, HSN18, KBGAP^{+}16, LKM23, SDDCC15, SMvdP22]. Mixup [CJY^{+}22]. ML [LLGG20, LG20]. MMI [HXL^{+}21, HXC^{+}22, TYW^{+}23]. MMSE [ET18, KBG16b]. MMSE-Based [KBG16b]. mmWave [OWW^{+}23]. mmWave-Based [OWW^{+}23]. Mobile [TZW21, GH23b]. Modal [AF23, BYT^{+}21a, BMVG17, CHGS23, FSA20, GYXY23, LLWQ24, LLL24, TZA14, TLD^{+}23, WCZR23, WYL^{+}23, YTF^{+}23, ZSC^{+}23a, ZLL^{+}24a, ZCQ^{+}19]. Modality [CLHK21, GYXY23]. Mode [UKS19, WWG^{+}23, ZCF14]. Mode-Matching [UKS19]. Model [ARW^{+}16, AJLW14, AMK^{+}20, BDV^{+}14, BDL15, BCA15, CONVVC18, CW18, CH16, CWYC17, CWCP17, CZY^{+}18, CJY^{+}21, CWJ15, Chi15a, CDA^{+}14, CS23, DV14b, DLY^{+}19, DHD^{+}19, DvWDM14, DLG18, DJY^{+}19, DDZS17, DFY^{+}24, DBM15, DDNN19, FLL^{+}23, GTS^{+}19, GLWQ22, GWQ23, GR14, HSN18, HIA^{+}17, HDW^{+}23, HOF15, IO19, JFG^{+}17, JBTA19, KSL19, KNBC19, KMAB18, KKK15, KGH^{+}16, KHHJ19b, KGG19, LLR^{+}16, LWJ19, LG19, LTWT17, LRWG23, LLWC21, LKHU14, LAK16, LWCC21, LTE^{+}21, LCQ22, LHC^{+}23, LHL^{+}23b, LCW24, LWL^{+}22, LMW^{+}24, LHW^{+}21, LJX^{+}16, LM23a, LS14b, LHL^{+}22a, LZL^{+}23, LLdCF23, LJL^{+}23, LHZ^{+}24, LSTK21, MWT16, MYT^{+}19, MBD18, MV19, MPKS^{+}23, MBDV23, MTK^{+}20, MMZ23, MDM15, NCB23, NV14, OOK17, PBA^{+}23, QZLH21, MG18, RLPK15, RCR^{+}21, SAC17, SGH17a, SMO^{+}21, SMB18, SZX^{+}18, SYG^{+}24, SKT21, SR16b, SLR16, TQY16, TW21, TLWV17, TWC^{+}18, TQM^{+}23, TZZ^{+}15a]. Model [TP16, TG16, TWSS16, TZZ15b, VMOL16, WZL^{+}15, WYLZ16, WLLZ17, WTY18, WB18, WYLZ20, WTY^{+}20b, WDW^{+}22, WFY^{+}22, WJ23, WZLL24, WQ24a, WSFS19, WSH15, WL19, WHO^{+}21, WHT^{+}21, WSZN21, WSZN22, WLF23, XZL^{+}23, XHS^{+}18, YLGG22, YYW^{+}23, ZFC^{+}23, ZDC14, ZZZ16, ZYF18, ZZY19, ZLZW21, ZCHW24, ZJL21, ZLD21]. Model-Agnostic [LM23a]. Model-Based [BCA15, CW18, GWQ23, KNBC19, KGH^{+}16, MBD18, OOK17, SGH17a, WB18]. Modeling [AF23, AS17b, AKN^{+}16, BP14, BHS^{+}16, BBFS21, BMV^{+}23, CDLL19, CHC^{+}14, CLC^{+}15, CDLX18, CWJ16, CBCL15, CC15, CGK15, DAMvW15, DAFS15, DKDN20, DCW^{+}21, DDNN19, FOBP15, FL19, GLQ16, GDC19, GBH16, GCJ^{+}14, GdBBS21, GSP23, GQXZ19, HSN18, ISZA15, JK19, JWBR21, JBWZ24, KYI^{+}15, KVMV^{+}24, KKGK18, LZC^{+}23, LM15, LWZZ16, LCQ22, LZAY23, LYL21, LB23, LAGD18, LCC^{+}15, LK16, LLWJ16, LXS^{+}18, LCW^{+}21, MX17, MSS17, MD16, MJJ^{+}14, NY21, NW24, NCC15, OL15, PDKG24, QWL^{+}19, RV14, RAB18, RJW19, RCL18, STS21, SS16, SLL^{+}16, SGL^{+}21, SYKL21, SSK17, SMH14, SMSV17, SLY14, SYG^{+}24, SNS15, TMH^{+}14, VM15, WS14, WLL^{+}15a, WCC^{+}15, WTW20, WDW^{+}23, WWX^{+}23, WCC^{+}23, WSZN22, WLL^{+}23, WXS24, XH23, XZW15, YsL15, YCL22, ZKA15, ZHB^{+}18, ZLL^{+}19, ZGL^{+}20, ZJD^{+}15, VZC14]. Modeling-Based [CC15, DDNN19]. Modelling [DY22, HKM^{+}24, LYB^{+}23, ONKP16, SFIL17, WZC23, WK16, XZFN23, ZZS^{+}23]. Models [Abd18, AVK^{+}15, ACRS14, AGBC14, BYT^{+}21a, BSR17, BS14, CAFFR23, CPCBFSF16, CLW^{+}16, CLW^{+}19, CHL^{+}24, Chi18a, Chi15b, CZK^{+}21, Cum15, CL18b, Cum21, DOL^{+}23, DKO^{+}18, DBK15, DSN18, DWX^{+}16, DH14, GMH14, HSPK18, HK24, HH14, HYI^{+}15, IBN^{+}18, JGSC^{+}22, KLK17, KWGW17, KK19a, KWZ24, KPA^{+}24, KCL21, LGTG18, LBR18, LCSL19, LCW^{+}16, LLW21, LGR14, MGB18, MASC23, MLC^{+}23, MZM15, NW15, OCS21, PCW15, RV17, RWL^{+}23, SV18, SFRK^{+}23, SSSS22, SHL^{+}21, SWEV23, WTY20a, XHW^{+}17, XYH^{+}21, YKL^{+}23, YCL22, ZM17]. Modern [PKR^{+}24]. Modification [DAR16, MCF21, SZS22, TKT17, ZZX^{+}21]. Modifications [EZS14]. Modified [PCG17, YWDQ21]. Modify [TTB^{+}16]. Modular [DOL^{+}23]. Modularized [QXW^{+}23, SCCM23]. Modulated [AP21, JT16]. Modulation [BV16, CSS^{+}19, DB15, DB18, DB19, ECJF21, GMH14, HSK17, HHC16, MAK15, MKA16, MBK17, TTB^{+}16, WB18, XWWW18, XMBN18, XMBN21, ZTM^{+}23]. Modulation-Based [HSK17]. Modulation-Domain [DB18, DB19, MBK17, XMBN18, XMBN21]. Module [LJH^{+}24, ZQN^{+}23]. Modules [ACK21, YPH23]. momentum [Den14]. Monaural [BS21, CH15, CY16, DW17, DB19, FBE19, GvdP23, HKHJS15, JSYW16, KH24, LSDM19, LLZ^{+}21a, LW19, MV19, NK21, SWCN19, TOS14, TCW19, TW20, WTW20, WWL21, WZG^{+}23, WWW16, XGKM18, YDW^{+}21, ZW16b]. Mongolian [LSB^{+}21]. Monolingual [PM15]. Montage [OR16b]. Moog [DV14a, DV14b]. Morph [MK15a]. Morph-Based [MK15a]. MorPhEMe [QVZY24]. Morphological [FOBP15, HBF^{+}16, LLR^{+}16, LSB^{+}21]. Morphologically [ZP15, dlCSVAGAPM15]. Morphology [LHZ^{+}24]. Morphology-Aware [LHZ^{+}24]. Motif [XO16]. Motifs [XO16]. Motion [AF23, SLH17]. Motivated [LYZ^{+}23, ÖM24, VDC^{+}17, VS14]. Movements [KCMW17, MMB17, TGGE21, YLLC19]. Moving [KGBPM24, MØ20, NDV18, ODNA23, OVN21, TH18, TR15]. MPEG [CRH^{+}14]. MPEG-4 [CRH^{+}14]. MRI [ADGE19, KS21]. MRI-Based [ADGE19]. MSE [HDLL23, Wid19]. MSE-Optimal [Wid19]. MsEmoTTS [LYWX22]. MSM [WWX^{+}23]. MSM-VC [WWX^{+}23]. MSMC [GXW^{+}23]. MSMC-TTS [GXW^{+}23]. MSStyleTTS [LZC^{+}23]. MT [BDL15]. Multi [AF23, ACK21, AMK^{+}20, AGC23, BM21, BHP^{+}23, BPG23, BSL15, CC21, CBN^{+}24, CD17, CZWC23, CLL^{+}21b, CS19, DCZ19, DHYG14, DHD^{+}19, DAR16, DST^{+}19, DDMvW20a, DDMvW20b, DH20, DWX^{+}16, DHP14, FSA20, FYG^{+}21, FOBP15, GTS^{+}19, GLT23, GJYW23, GWLL22, GKC16, GJ22, GLL20, GYXY23, GXW^{+}23, HMO21, HDLL23, HLB^{+}24, HWRC15, HZC^{+}22, HCC^{+}23, JKC^{+}24, JWBR21, JQL24, JGHN23, JvWGD15, KKJK23, KD16, KD17, KTH^{+}20, LYWX22, LZC^{+}23, LWM^{+}20, LWCC21, LY21, LXK^{+}21, LDG^{+}23, LHL^{+}23a, LSZ23, LWL^{+}22, LWLL24, Lin19, LdLvWWS21, LCRY24, LH14, LLL24, LHG^{+}24, LYB^{+}23, LLL^{+}22, LZ23, MZZ23, MHZ^{+}24, MHD15, MHGD15, May18, MPC^{+}24, NSNN20, OVN^{+}22, PZW^{+}21, PC19, PFLX24, PZP19, PP19, PT21, QTY16, QCL23, RPK23, RC23, STS21, SIKN23, SNTD19, SGH15b, SHCV23, SGY^{+}24, SCW21, SLOVB15, TZA14, TD23, TYZ^{+}22, TZS^{+}21, TLD^{+}23, TG23, TG24, VH19, WLZH21]. Multi [WWW21, WH22, WSJ^{+}23, WCZR23, WYL^{+}23, WWX^{+}23, WZZ^{+}24, WWG^{+}24, WYSD15, XXS^{+}23, XBZ^{+}24, XRWL21, XSC^{+}23, XCHX24, YW14, YLL15, YTF^{+}23, YND14, YJX19, YZW^{+}21, YDW^{+}21, ZDG14, ZFC^{+}23, ZXY^{+}21, ZLZ21, ZCDH21, ZZZ22a, ZSC^{+}23a, ZLLZ23, ZZD23, ZDZ^{+}21, ZLG^{+}22, VZC14]. Multi-Agent [LWCC21]. Multi-Aspect [PFLX24]. Multi-Bernoulli [Lin19]. Multi-Branch [CC21, TYZ^{+}22]. Multi-Channel [CBN^{+}24, CS19, DST^{+}19, JvWGD15, KD16, KD17, LDG^{+}23, MHD15, MHGD15, PZP19, PT21, QCL23, SLOVB15, TZA14, TG23, TG24, XBZ^{+}24, YND14, YZW^{+}21, ZXY^{+}21, ZLLZ23]. Multi-Choice [ZZD23]. Multi-Classifier [ZLG^{+}22]. Multi-Codebook [GXW^{+}23]. Multi-Compositionality [DWX^{+}16]. Multi-Conditional [May18]. Multi-Cue [XCHX24]. Multi-Data [JQL24]. Multi-Device [HWRC15]. Multi-Dimensional [DCZ19]. Multi-Document [AGC23, YW14]. Multi-Domain [BM21, JGHN23, LWLL24, LZ23, MHZ^{+}24, RC23]. Multi-Factor [QTY16]. Multi-Feature [ZDG14]. Multi-Frame [DAR16, SIKN23, TD23, ZXY^{+}21]. Multi-Genre [DHD^{+}19]. Multi-Grained [HCC^{+}23, ZSC^{+}23a, ZZD23]. Multi-Granularity [LXK^{+}21, LHG^{+}24, PFLX24, TZS^{+}21, WYL^{+}23]. Multi-Hop [ACK21]. Multi-Input [CD17]. Multi-Kernel [CLL^{+}21b]. Multi-Label [KTH^{+}20, XXS^{+}23]. Multi-Layer [XSC^{+}23]. Multi-Level [CZWC23, GWLL22, LHL^{+}23a, WWG^{+}24, ZDZ^{+}21]. Multi-Lingual [JKC^{+}24]. Multi-Metric [LY21]. Multi-Microphone [DDMvW20a, PC19, SNTD19, SGH15b, VH19, ZCDH21, ZLLZ23, WWW21]. Multi-Modal [GYXY23, LLL24, TLD^{+}23, WCZR23, WYL^{+}23, YTF^{+}23]. Multi-Motion-Informed [AF23]. Multi-Objective [ZFC^{+}23]. Multi-Output [CD17]. Multi-Party [BSL15, MZZ23, WSJ^{+}23]. Multi-Pitch [GJ22, DHP14]. Multi-Purpose [GKC16]. Multi-Resolution [MPC^{+}24, PP19, YDW^{+}21]. Multi-Scale [DH20, HLB^{+}24, LYWX22, LZC^{+}23, LCRY24, WWX^{+}23]. Multi-Session [LH14]. Multi-Singer [KKJK23]. Multi-Song [DHYG14]. Multi-Source [DDMvW20b, FSA20, GJYW23, JWBR21, LSZ23, LLL^{+}22, NSNN20, OVN^{+}22, WH22]. Multi-Spans [ZZZ22a]. Multi-Speaker [BHP^{+}23, HMO21, HDLL23, JKC^{+}24, STS21, SHCV23, SCW21]. Multi-Stage [GTS^{+}19, GXW^{+}23, LdLvWWS21, LLL^{+}22]. Multi-Stream [LWM^{+}20, LYB^{+}23]. multi-style [VZC14]. Multi-Talker [RPK23, WYSD15, XRWL21, YLL15]. Multi-Task [AMK^{+}20, BPG23, FYG^{+}21, FOBP15, GLT23, JQL24, LCRY24, SGY^{+}24, WYL^{+}23, YTF^{+}23, YJX19]. Multi-Teacher [LWL^{+}22]. Multi-Tone [PZW^{+}21]. Multi-Turn [CZWC23, GLL20, MZZ23, WLZH21, WZZ^{+}24, ZLZ21]. Multi-View [HZC^{+}22]. Multiband [PZZZ22, PPSG17, WFY^{+}22, XWWW18]. Multiband-Structured [PZZZ22, PPSG17]. Multichannel [BP14, BIK^{+}18, Bay15, CONVVC18, DDNN19, Emu18, HNT^{+}23, HCBY18, HdHTJ22, HBCC22, IBN^{+}18, IISN21, KSK^{+}15, LPHCB14, LBR16, LBR18, LGGH18, LGGH19a, LGGH19b, LZHN14, Lin18, LFdDG14, MYKO24, MUT^{+}20, MTKS21, NDV18, NLV16, OIY^{+}14, QL24, RKL14, SWW^{+}17, SEM^{+}14, SK16a, SBN^{+}19, SMVW14, SBM^{+}19, TH17, TG21, TE23, TS14, WYO23, WCC^{+}21, WGLZ21, WYY22, WYY24, XMBN18, XMBN21, YK14, ZDLK18, dSCB24, dCBC23]. Multiformat [Sug22]. Multilayer [NXH^{+}17]. Multilingual [KMG21, LYY^{+}23, LLC23a, SV17, YČS23, ZXH^{+}16, ZLL^{+}24b]. Multiloudspeaker [FWF17]. Multimedia [Ano17-58, Ano16g, Ano16h, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u]. Multimicrophone [GVMGO17, KHHJ19b, SGH16]. Multimodal [CCH^{+}21, CHGS23, CLW24, CCC^{+}24, DTC17, GMH17, HQTX23, HM14, KS21, LLZ^{+}21c, LCRY24, MXH21, NK22, SFEB23, SLH17, VSK^{+}24, YJX20, ZKKC22, ZWS^{+}23]. Multiobjective [WDDL18]. Multipitch [SY15]. Multiple [AHSS^{+}18, ATA16, AM18a, AB15, BA16, BCA15, BOL^{+}18, BBFS21, BG23, ÇOH18, DCZ19, ESJ17, HEMN21, HAS21, HZL23, JYB^{+}15, KCL21, LGTG17, LWZ^{+}14, LAK16, LGHG17, LWL^{+}21, LJWC23, MMB17, MAZ21, MGGK17, NR14a, NYS17, OVN21, PHK16, SLOVB15, SSS18, TOS14, TS16, WED16, WLZH21, XNP^{+}18, XPC^{+}22, XDC^{+}17, YLPL19, YZCW21, YZLY17, YHO^{+}17, ZA14, ZH15, ZJD^{+}15, KMG17]. Multiple-Choice [KCL21]. Multiple-Integration [WLZH21]. Multiple-Speaker [LGHG17]. Multiple-Speech-Source [LJWC23]. Multiplicative [FK19, HBB17]. Multiplicity [PBC16]. Multirate [YLJ16]. Multiresolution [DG23, NKS21]. Multiscale [KDY14, LLC^{+}24]. Multisession [GH17]. Multisource [BSTM21, LMC18, WRK18]. Multispeaker [SGH17c]. Multistage [ZHCB23]. Multitalker [KYTJ17, SMvdP22]. Multitask [BSR17, CM15, DCLHJ18, TMM18, TLWV17]. Multiview [DR22, GFG^{+}14, MAST15, MBD^{+}15]. Multizone [AKUS23, CH14, DRK18, JK15, RBR16]. MuSE [KKJK23]. MuSE-SVS [KKJK23].

MuseMorphose [WY23]. Music [AS17a, AB15, BPG23, CWYC17, CMD^{+}23, CDW16, DHYG14, GGL23, HJC19, HNT^{+}23, JSYW16, JLP^{+}21, KM23, KHCW15, KG16, LSW18, LSH^{+}23, LCZ23, LY23, MV23, NWM16, NNS16, NYS17, PKA14, RV14, RLS^{+}18, RRdM16, SBV17, SFRK^{+}23, SBD16, SY15, TOS14, TP22, VML14, WY23, WLF23, YBC^{+}14, ZM21, ZCYZ18, BAS20, HAS21]. Musical [AGV24, ARW^{+}16, AB14, FMD14, MP16, SCY22, SHG21, WED16]. Musically [ÖM24]. Musicians [TP22]. MusicYOLO [WTY^{+}23]. Mutual [IIY16, JT14, KHK17, KNH^{+}17, MYT^{+}19, RMBK14, TM14, TM16, TK14]. Mutuality [SCCM23]. MVDR [AVM19, GMD21, HMDG15, HIA^{+}17, PCB14, SM14, TD23, ZCHH18, ZXY^{+}21]. MWF [ZL21a]. MWSG [KMG17].

Named [FLC^{+}23, FLYJ23, LWCC21, QZLH21, XHS^{+}18, ZWW^{+}23b, ZCY^{+}22]. Names [MK15a, PBQ15]. Narrow [AA17]. Narrow-Band [AA17]. Narrowband [AS15, CE14, FK19, HFL^{+}23, HKI23, JTHN14, LC16b, MX17, ZLSC23]. Nasalized [VAP19]. Native [DKDN20, LCSL19, Lib21]. Native-Nonnative [Lib21]. Natural [CSJD16, CWH^{+}16, CHL^{+}23, JWH^{+}22, LFH^{+}15, LHL^{+}23a, LLC^{+}23b, NLYK24, RG15b, RLK14, SHD14, SWH22, WCL^{+}23, XWY^{+}23, ZMX^{+}21]. Naturalistic [DSH18, KSH17]. Near [ATP16, BNA24, FJT^{+}23, HCJT15, LMGG21, LY21, MF18, PC19, VAH19, ZSIA16]. Near-End [BNA24, FJT^{+}23, HCJT15, LY21, PC19]. Near-Field [ATP16, LMGG21, MF18, VAH19]. Nearby [YKPMP18]. Nearest [BG21, DZD^{+}24, MSD14, SDK17]. Nearest-Neighbors [MSD14, SDK17]. Nearfield [DCAS22]. Neck [AFG^{+}23, MVH16, MYC^{+}23]. Neck-Surface [MVH16]. Need [GSA18]. Negative [ATA16, BDV^{+}14, BOL^{+}18, DVLH17, ED18, HLB^{+}24, MUT^{+}20, MTKS21, MD16, ONKP16, ZXH^{+}16, PPM^{+}24]. Neighbor [DZD^{+}24, XZL^{+}21]. Neighborhood [LHL^{+}23b]. Neighbors [MSD14, SDK17]. NER [LJL^{+}23, WYL^{+}23]. Nested [LWCC21, ZWW^{+}23b]. Net [JLKK21, MOT^{+}23, CC21, LYH^{+}22, HFL^{+}23, JLKK21, NR22, ZZX^{+}23b]. Nets [BGD^{+}18]. Network [AHF21, ACRS14, AGD^{+}22, BLZ21, BSR17, BKTR24, BB21, CSS^{+}19, CLC^{+}15, CLW^{+}16, CLW^{+}19, CMY^{+}21, CB21, CHGS23, CWZ^{+}23, CLL^{+}21b, CGK15, CLZ22, DHD^{+}19, DKO^{+}18, DWX^{+}16, DV23, DBP^{+}24, FBWG23, FLMS16, GTV19, GAPGG19, GLL20, GGZ23, HW14, HRL^{+}21, HJZB23, HMO21, HFL^{+}23, HHO^{+}21, HCY21a, HZBW23, HQTX23, JLP^{+}21, KHT^{+}21, KCMW17, KSK21, LTWT17, LWLS24, LLT21b, LYL21, LWYD23, LCW^{+}16, rLcHcC^{+}22, LSG^{+}23b, LHG^{+}24, LCM18, LFM^{+}23, MXH21, May18, MZM15, MWC^{+}23, NW15, NR22, OGZ18, PKP^{+}21, QDSL19, QTY16, QCW21, QWW^{+}23, QCN^{+}21, SS16, SCCM23, SGY^{+}24, SSGL21, SBD16, SWEV23, TQY16, TQH^{+}18, TB18, WDDL18, WWS^{+}18, WPY^{+}21, WHO^{+}21, WHT^{+}21, XXS^{+}23, XLLW21, XZL^{+}21, XYH^{+}21, XAHJ^{+}14, YHC19, YPC^{+}22, YHO^{+}18, YJX20, YHL^{+}23, ZW16a, ZZN^{+}16, ZKH18, ZWZ^{+}21, ZPW23, ZZY^{+}24, ZLZ^{+}24, ZZZ22b, ZWHD23, ZWS^{+}23, ZYW^{+}24, ZZX^{+}23b]. Network-Based [BKTR24, FBWG23, GTV19, HHO^{+}21, May18]. Networks [AHMJ^{+}14, AF18, AM18a, AHH^{+}20, AdTJ18, AFdDG23, BBM17, BRD15, CBA^{+}15, CLAG16, CBN^{+}24, CLLD14, CM15, CG17, CHL^{+}23, CPSFC^{+}14, CG16, DMHZ22, DSD23, DHCS15, DAGS^{+}18, DGMB21, DQY18, DTDL16, DG23, FdDPG21, FB18, FB19, FWT^{+}18, GLT23, GBL24, GWG^{+}14, GTSM24, GRSP17, HXL^{+}21, HCY21a, HXC^{+}22, HZBW23, HSLB23, HKHJS15, JGM23, JG19, JWLF14, KWGW17, KK19a, KYTJ17, KH24, KEzW21, KSHH18, LC16a, LS14a, LY21, LWZG21, LdLvWWS21, LYH^{+}22, LAGD18, LB16, LFQL19, LR17, LCTA19, LLL^{+}23, MMB17, MGB18, MZX^{+}15, MDY^{+}15, MPC^{+}24, MV23, MKA16, NKS21, NNC20, NLV16, OS21, OCS21, PS19, PDS^{+}16, PZW^{+}21, PPM^{+}24, PHM^{+}17, QYCT23, QBTY16, RvWM22, SWW^{+}17, STS18, SHAGPRZ17, SF18, SHJT21, SHD14, SH17a, SFV23, STBC17, SWCN19, SNS15, TWZ^{+}18, TCW19, TW20, TWIN23, TCY15, TL19, VDH24, VMOL16, WAS14, WS14, WDDL17]. Networks [WC18b, WCY19, WYSD15, WHB^{+}21, WCL^{+}21b, WLL^{+}23, XWY^{+}23, XDDL15, YZL^{+}18, YZCW21, YC22, YWT23, YMM19, YJX19, YK21b, ZCY16, ZHH18, ZHH19, ZPD19, ZTDD23a, ZLF19, ZZ19, ZJD^{+}15, ZDZ^{+}21, rPH^{+}17]. Neural [AHMJ^{+}14, AF18, AL23, AHF21, AdTJ18, ACRS14, ASC^{+}18, BLZ21, BRC24, BSR17, BKTR24, CLAG16, CBN^{+}24, CLLD14, CLC^{+}15, CM15, CLW^{+}16, CZY^{+}18, CLW^{+}19, CWU^{+}19, CLL^{+}21a, CB21, CWUS22, CHWQ24, CGK15, CG16, DHD^{+}19, DAGS^{+}18, DKO^{+}18, DGMB21, DFY^{+}24, DQY18, DBM15, DWX^{+}16, DV23, DBP^{+}24, DTDL16, DCW^{+}21, DZZ23, FBWG23, FLMS16, FWT^{+}18, GBL24, GWG^{+}14, GTV19, GAPGG19, GWL^{+}24, GRSP17, GZZY23, GZX^{+}21, GYXY23, GXW^{+}23, HW14, HRL^{+}21, HMO21, HDLL23, HHO^{+}21, HFW^{+}22, HWG^{+}23, HXL^{+}21, HXC^{+}22, HKHJS15, HSZ^{+}18, JGM23, JHLK18, JG19, JWLF14, JPX^{+}23, JY23, KWGW17, KK19a, KCMW17, KYTJ17, KEzW21, LC16a, LTWT17, LLC22, LS14a, LWC^{+}18, LLT^{+}21a, LLZ^{+}21b, LWZG21, LZZP22, LLJ^{+}23, LWLS24, LWL^{+}22, LYH^{+}22, LAGD18, LCW^{+}16, LWCL19, LLZ23b, LCQ23, LR17, LLL^{+}22, LLZ^{+}23d, MMB17, MGB18, MHZ^{+}24, MOT^{+}23, May18, MZX^{+}15, MDY^{+}15, MZM15, MWC^{+}23, MBDV23, MPC^{+}24, MKA16, MVN22]. Neural [NKS21, NW15, NNC20, NR22, NSNN20, NLV16, OS21, ODNA23, OCS21, PS19, PZW^{+}21, PHM^{+}17, QDSL19, QYCT23, QTY16, QBTY16, QCW21, QVZY24, RCL18, SWW^{+}17, SS16, SF18, SSI21, STBC17, SSGL21, SZS22, SgL20, SBD16, SZX^{+}18, SWEV23, SWCN19, SNS15, TQY16, TWZ^{+}18, TYZ^{+}22, TWW22, TWC^{+}18, TB18, TWIN23, TJV24, VDH24, WS14, WDDL17, WTY18, WTZ18, WDDL18, WUF^{+}18, WHQY19, WXZ^{+}19, WSL^{+}20, WTY20a, WTY^{+}20b, WPY^{+}21, WW22, WFY^{+}22, WWWR23, WYSD15, WHB^{+}21, WZZ^{+}18, WHO^{+}21, WHT^{+}21, XWY^{+}23, XLLW21, XDDL15, XYH^{+}21, XAHJ^{+}14, YZL^{+}18, YHC19, YC22, YPH23, YYY^{+}23, YWT23, YHO^{+}18, YMM19, YJX19, YK21b, YČS23, ZLO^{+}22, ZW16a, ZXSD17, ZLJ^{+}17, ZYF18, ZKH18, ZZLZ19, ZXSL19, ZPD19, ZZY19, ZLS^{+}21, ZLLZ23, ZQN^{+}23, ZLF19, ZKKC22, ZSL^{+}22, ZZ19, ZJD^{+}15, ZYW^{+}20, rPH^{+}17]. Neural- [BKTR24]. Neural-Network [DV23]. Neural-Network-Based [YHO^{+}18]. NeuralDPS [WFY^{+}22]. Neurally [VR19]. Neuro [AFG^{+}23]. Neuro-Based [AFG^{+}23]. Neurogram [MJZ15]. News [CLC^{+}15, DDZL19, LPM^{+}23, PACPJF^{+}23]. Newton [BBFS21, FB19, HCY21a, HCY21b, HZBW23]. Neyman [KSU^{+}19]. NLP [CHL^{+}24, LMW^{+}24]. NMF [BP14, GWG^{+}14, NLL^{+}18, NDV18, SBM^{+}19, SLGZ15, WRDP17]. NMF-Informed [SBM^{+}19]. NMF-Style [NLL^{+}18]. Node [WCY19]. Noise [AA17, ATR^{+}21, AS17b, AHH^{+}19, AHH^{+}20, AM16, AM18b, AFdDG23, AKN^{+}16, BSB14, BB21, BH18, BBJ18, CC14a, CE14, CD17, CWW14, CZ17, CSZZ24, CCC21, DBK15, DST^{+}19, DDMvW20a, DBN^{+}17, EZS14, EMXD20, FdDPG21, GTV19, GI15, GB14, GMD21, GdC14, HH14, HCBY18, HAGA21, HCJT15, HCH23, HIA^{+}17, HCZC24, HBLC14, JFG^{+}17, JTHN14, JT16, JBC16, JLKK21, KGP14, KAJC16, KK19a, KD16, KBI^{+}21, KDJJ16, KDJJ17, LPHCB14, LLGG20, LG20, LDGHU14, LS14a, LGGH18, LTE^{+}21, LWZG21, LLZ^{+}23a, LV22, LC16b, LETJ21, LFdDG14, LSH14, LP22, LSGH24, MX17, MZA20, MMS^{+}20, MPAEBLB15, MPC16, MHD15, MHGD15, MD18, MCZ21, MNFGO19, MXNB19, NW14, NCB23, NJC16a, PK15, QBTY16, QCDW17, MG18, RH18, RG17, RBBG15, RvWM22, SK17, SM14, SGM^{+}15, SEM^{+}14, SGH15b, SGH16, SGH17b, SGH17c, SMVW14, SGL^{+}21, SGLS23]. Noise [SGL^{+}23, SZS22, SBM^{+}19, SLW15, SLGZ15, SZVZ16, SK18, SBVM16, TM16, TQH^{+}18, TMM18, TH17, TP19, TK14, TZZ^{+}15a, TVJ15, TDL19, UNI^{+}24, VP21, VE24, WGD15, WCC^{+}21, WFY^{+}22, WZS14, WZMH22, XWWW18, XWWW19, XCZ^{+}23, YBC^{+}14, YKPMP18, YLH23, YGV14, YPL21, YHO^{+}18, ZAZ^{+}18, ZCHH18, ZHH18, ZHB^{+}18, ZTQB19, ZL21a, ZZZ^{+}21, ZZ22, ZPW23, ZTDD23b, ZLX^{+}24, ZZZD23, ZLSC23, ZSIA16, dSCB24, VZC14]. Noise-Aware [YLH23]. Noise-Controllable [WFY^{+}22]. Noise-Field [ZHB^{+}18]. Noise-Independent [EZS14]. Noise-Reduced [GTV19]. Noise-Resilient [EMXD20]. Noise-Robust [AS17b, GI15, HCZC24, LDGHU14, LS14a, LETJ21, SBM^{+}19, TDL19, ZZZD23]. NoiseBandNet [BRC24]. Noises [GA20]. Noisy [AC24, AdHTJ16, BHP^{+}23, Bor22, CH23, DH14, ET18, FSA18, FZZ^{+}15, FNA15, GAM23, HW14, LFO^{+}24, LKHU14, LH23, LH14, MGGK17, NEN21, NV21, SGH17a, SZS22, VBY18, WWL21, XT23, XZSX23, ZWW23a, ZWW14, ZWW19]. Noisy-Reverberant [WWL21, ZWW19]. Noisy-to-Noisy [XT23]. Non [ATR^{+}21, ATA16, BP14, BDV^{+}14, BYT^{+}21a, BOL^{+}18, BSWP18, CSS^{+}19, CLZ^{+}24, CHL^{+}24, DBK15, DVLH17, DKDN20, ED18, FCCA23, KKTH19, LLC22, LCSL19, LB23, LHYJ23, LSH14, MPAEBLB15, MAP14, MUT^{+}20, MTKS21, MD16, NTM16, NJC16a, ONKP16, PSL^{+}16, PPM^{+}24, PJTJ24, SOF15, TG23, TG24, TP16, TP22, WWX^{+}23, WJ15b, WHO^{+}21, XZF23, YCL22, ZFC^{+}23, ZLD20, ZZZL21, ZXH^{+}16, ZXN^{+}21]. Non-Autoregressive [BYT^{+}21a, CLZ^{+}24, FCCA23, LLC22, WHO^{+}21, YCL22]. Non-English [CHL^{+}24]. Non-Ergodic [TP16]. Non-Intrusive [CSS^{+}19, PSL^{+}16, PJTJ24, XZF23, ZFC^{+}23]. Non-Invasive [BSWP18]. Non-Linear [TG23, TG24]. Non-Linear-Echo [ZXN^{+}21]. Non-Linearity [ATR^{+}21]. Non-Musicians [TP22]. Non-Native [DKDN20, LCSL19]. Non-Negative [ATA16, BDV^{+}14, BOL^{+}18, DVLH17, ED18, MUT^{+}20, MTKS21, MD16, ONKP16, ZXH^{+}16, PPM^{+}24]. Non-Orthogonal [SOF15]. Non-Parallel [KKTH19, LHYJ23, NTM16, WWX^{+}23, ZLD20, ZZZL21]. Non-Segmental [MAP14]. Non-Stationary [BP14, DBK15, LSH14, MPAEBLB15, NJC16a]. Non-Uniform [LB23, WJ15b]. Noncontinuous [MAT15]. Nonintrusive [AdTJ18]. Noniterative [PBS17]. Nonlinear [Ano19c, BG21, CCR16, CSAR^{+}14, CG16, CL17a, CL17b, CL18b, DV14b, FB18, FB19, GdBBS21, HSK17, HHMK18, MMMDL^{+}14, PC19, PMB^{+}19, SLGA15, TG21, XCZ^{+}23, YPL21]. Nonlinearities [BBFS21, CCTO21]. Nonlinearly [AdHTJ16]. Nonmodal [BMVG17]. Nonnative [Lib21]. Nonnegative [BV16, CONVVC18, FSO17, HL15, JLKK21, KHTL18, KSK^{+}15, KOS^{+}16, LKPR18, LGGH18, MH16, WGLZ21, WYY22]. Nonparametric [Chi18b, LM15, MIYO14, TVJ15]. Nonparametrics [OISO14, OIY^{+}14, SCW21]. Nonrecurrent [ZLJ^{+}17]. Nonspeech [AY15, PHM^{+}16]. Nonstationary [MT18, NJC16b, OD17, TH17, WAS14]. Nonuniform [MWL22]. Nonuniqueness [SK16a]. Nonverbal [HSWC21]. NOPM [MJZ15]. Norm [ABA17, BV17, DC16, GCJ^{+}14, HCBY18]. Norm-1 [ABA17]. Norm-2 [ABA17]. Normal [JRA17, ZDLK18]. Normal-to-Lombard [JRA17]. Normalization [CLA^{+}21, JKU18, SIUW21]. Normalized [CJ15, DGCY23, GYdL^{+}21, KS16, YHC^{+}20, ZZH16, ZGL^{+}23, dlCSVAGAPM15]. Notch [GSA18, WXL^{+}21]. Notches [SMO^{+}21]. Note [DLH23, NYD17, ONKP16]. Notes [WED17]. Novel [ATR^{+}21, BKK15, CBN^{+}24, CMM^{+}23, DV14b, DMSM21, EDO17, LQW^{+}24, LLT^{+}21a, LETJ21, LSH14, NR22, PCLP14, RJCW17, SR16a, SP16, TXH^{+}23, TG16, WCY15, WQ24a]. NTF [FK24]. Nuisance [KMAB18]. Null [PBC16, SNTD19]. Null-Steering [SNTD19]. Nullors [GdBBS21]. Number [OIY^{+}14, SCEH19]. Numbers [FM15, HWG^{+}23]. Numerical [ZBW^{+}24]. Nutrition [KG17].

Object [JYB^{+}15, PBA^{+}23]. Objective [BNvdB^{+}20, BvWM18, FBE19, GTSM24, KNH^{+}18, KTJ19, MLC^{+}23, MGT15, PCLP14, TM14, TKH21, ZFC^{+}23]. Objects [JYB^{+}15, MF17, MCF21]. Oblique [CG20, SK18]. Obscurity [ZWL^{+}23a]. Obscurity-Quantified [ZWL^{+}23a]. Observation [LKHU14, NHK16, RLPK15]. Observations [IO19]. Obstacle [PBA^{+}23]. Occlusion [LV22]. occurrence [CBCL15]. Octave [YDZ19]. Off [LSNC21, SLJ18, ACRS14]. Off-Grid [SLJ18]. Offset [BBM17, CTE21, HZBW23, KLS^{+}21, WD16]. Offsets [MYKO24]. On-Bird [SBG17]. On-Board [MAZS22]. On-Device [TYZ^{+}22]. On-the-Fly [EDO17]. One [BBFS21, DTN^{+}14, ED18, JQL24, LLZ^{+}21a, MJJ^{+}14, SMB18, WY23]. One-Dimension [MJJ^{+}14]. One-Dimensional [SMB18]. One-Port [BBFS21]. One-to-Many [DTN^{+}14]. Online [AA17, AKN^{+}16, BH18, CH15, CCY^{+}24, DDBW20, DDBW23, DBN^{+}17, ESJ17, EV23, HIA^{+}17, HWG^{+}23, IK23, LKC^{+}21, LSW18, LSDM19, LGGH19a, MX17, MP16, MYNO23, MYK^{+}24, MR20, OVN^{+}22, SGH15a, TMH^{+}14, UNI^{+}24, YK14, ZHB^{+}18]. Only [BYT^{+}21b, LYH^{+}22, LFU^{+}17]. Onset [KLS^{+}21, MP20]. Onsets [Lin18]. Ontologies [JHLK18]. OOV [GWCT23, HBF^{+}16, SFIL17]. Open [Ano14-82, Ano14-83, Ano14-84, Ano14-85, Ano14-86, Ano14-87, Ano14-88, Ano14-89, Ano15-71, Ano15-72, Ano15-73, Ano15-74, Ano17-70, Ano17-71, Ano17-72, Ano17-73, Ano17-74, Ano17-75, Ano17-76, Ano17-77, Ano18-43, Ano18-67, Ano18-68, CWR^{+}22, CJY^{+}22, FFP^{+}22, GH23b, Hac14, LGL^{+}24, SSI21, XLZ^{+}23, YČS23, ZXZZ23]. Open-Domain [LGL^{+}24]. Open-Source [GH23b]. Open-Vocabulary [SSI21]. Opening [KKGH16]. Opera [MSEC21]. Operated [ISZA15]. Operation [LGL^{+}22, ZBW^{+}24]. Operation-Augmented [ZBW^{+}24]. Operator [KPH15]. Operators [MSB18, MCS21, SH17b]. Opinion [MSFK15, WFR^{+}21, XLG^{+}23, YJX19, ZWX15]. Opinionated [WZM21]. Opportunities [Li15]. Optimal [AR16, CCC21, FWF17, HCJT15, JBC16, NHK16, NJC16b, SGL^{+}21, SKH^{+}15a, Wid19, YEY17]. Optimization [ABA17, BC17, CFMM14, DvWDM14, FWT^{+}18, GEC15, GDC19, GY14, GG18, HCY21a, HCY21b, HZBW23, HKHJS15, JEP14, KHHJ19a, LZC^{+}14, LY21, LLCT22, MYKO24, MMK21, PGS^{+}22, RBR16, SH19, SSSS22, SLK17, TK14, TR14b, UNI^{+}24, YHO^{+}17, ZWTL23, ZCQ^{+}19]. Optimized [GJYW23, HGT15c, MKA16, MK16]. Optimizing [BA16, BY15, CDLL21, KHKY22, KCD20, QZ22]. Optimum [GA20]. Option [FNM^{+}19]. Oracle [YKL^{+}23]. Oracles [ZYZU22]. Orchestral [KM23, LCZ23]. Order [BBAS19, BBJ18, CM18, CASZ17, FWRJ21, GLWQ22, HB17, Hel14, HPPM24, HF23, JGHN23, MFH21, MIYO14, MCS21, MMZ23, MYC^{+}23, PCB16a, SAP14, TC19, TC22, WCY15, WC16a, WFR^{+}21, YLC^{+}24, YPC^{+}22, ZZZ^{+}21, vMM14, NCC15]. Orders [HCB20]. Organization [yLSY^{+}14]. Orientation [GI15, TC19]. Oriented [CJY^{+}21, HZL23, LZX^{+}23, LXS^{+}18, LCL^{+}23b, MMBT21, QXW^{+}23, TDDN24, WYH^{+}15, ZYW^{+}24]. Origin [NNC20]. Orthogonal [CCTO21, DMVP16, DDMvW20b, HF21, MJZ15, MWC^{+}23, SOF15, XNP^{+}18]. Orthonormal [VDC^{+}17]. Orthotist [XXS^{+}23]. Other [DSMD15, ZP15]. our [DRFO14]. Out-of-Domain [JGK^{+}22]. Outcome [MHB^{+}18]. Outdoor [SMSV17]. Outgoing [MZA20]. Outlier [CWDA17]. Outlooks [TP23]. Output [CD17, EV23, NYS17, SGL^{+}21, SGLS23]. Output-Constrained [SGL^{+}21, SGLS23]. Over-Determined [OG14, WRC16]. Overdetermined [WYY22]. Overlapped [BPG23, CDLX18, MLMT24, SH17b, YZW^{+}21]. Overlapping [HEMN21, LWZ^{+}14, YB14, YH21, Ziv15]. Overview [GaKGB21, KGG^{+}21, KCD20, LDGHU14, MTZ^{+}21, PMA^{+}21, RLS^{+}18, SYKL21, WC18a, YCC^{+}24]. Own [HdHTJ22].

P [LLJ^{+}23]. P-Transformer [LLJ^{+}23]. Packet [ISZA15, LC16a]. Packets [FZZ^{+}15]. Page [Ano17c, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15n, Ano15o, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17b, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano19b]. Pair [CJY^{+}21, CBCL15, FYG^{+}21, HZL23, LYY^{+}23, SCCM23, WFR^{+}21, ZZJ^{+}23]. Pair-Wise [WFR^{+}21]. Pairing [ZGO19]. Pairs [TLD^{+}23]. Pairwise [CL14b, ZZN^{+}16, ZHZ19]. Palate [MP20, VAP19]. Panel [ZDZ^{+}23]. Panned [CHJ14]. Panning [FWF17, MF18, MGF18]. Paralinguistic [CXX^{+}23, GBF21]. Parallel [CE14, HSM19, cHrLLyL23, JJB16, KKTH19, LLC22, LLWC21, LHYJ23, NTM16, RVB14, SWC20, VE24, WWX^{+}23, WHO^{+}21, XYY^{+}23, ZLD20, ZZZL21, ZB14]. Parallel-Data-Free [HSM19]. Parameter [AM20, LH23, LLL24, SLW15, TD23, XGKM18, YG24, YK21b]. Parameterization [DV14a, NTY19, SAPB15]. Parameters [BG21, DDBW20, EGMN16, GTS^{+}19, HHMK18, KHHJ19b, LMH23, PCLP14, RCR^{+}21, WCY15]. Parametric [AJB^{+}18, CRVB^{+}15, DLG18, HSK17, HL18, JBTA19, KK19b, LS14b, MJJ^{+}14, Nos16, NAK^{+}17, PLZL17, PGS^{+}22, STS18, SFRK^{+}23, SMB18, TTB^{+}16, TTH14, WTY18, WTY20a, WTY^{+}20b, XMBN21, ZZ22, ZZL^{+}23]. Parametrization [VDM15]. Paraphrase [LCW24, XFH^{+}23]. Parity [XDH^{+}17]. Parkinson [AATKS21, NSA21, NRČR14, RA23].

Parkinson&#x0027 [LRP^{+}23]. Parse [SKK14]. Parsing [CZZ15, HW23, JLZD24, LZC^{+}14, LZZP22, ODSM16, WYH^{+}15, WLZH21, WSJ^{+}23, WZ21, YZL^{+}18]. Part [SD16a, DV14a, DV14b]. Partial [GMD21, HL15, MTNR21, MD18, WYHL20]. Partially [GLQ21]. PartialSpoof [ZWC^{+}23]. Particle [HHMK18, KHCW15, WRKG17, ZCY16, ZG16, ZCYZ18, ZH15]. Partition [LJH^{+}24]. Partitioned [BvWWM17, Yan22]. Partitioned-Block [BvWWM17, Yan22]. Partitioning [LC24]. Parts [LYB^{+}23, MFG15]. Party [BSL15, MZZ23, WSJ^{+}23]. Pass [Hel14, QMS16]. Passive [HCY21b]. Patchwork [NXH^{+}17, XNG^{+}14]. Patchwork-Based [NXH^{+}17, XNG^{+}14]. Path [ATR^{+}21, AKN^{+}16, GLQ16, GdC14, LGHG17, LCQ22, LSG^{+}23a, MX17, NR14a, SG21, YLL21, ZHB^{+}18, ZWZ^{+}21, ZGL^{+}23, ZQZH19]. Pathology [GA20]. Pathology-Induced [GA20]. Paths [LLWJ16, SD16b, SD16a]. Pattern [ATS18, CATS20, DTL15, FSA20, Son21, WWS^{+}18, YK17, ZZH21]. Patterns [PBC16, PHM^{+}16, RHvdP^{+}16, WsL15]. Pausal [PM19]. PCFG [ZWCT15]. Pearson [KSU^{+}19]. Pedals [KLS^{+}21]. PEFAC [GB14]. PEM [GCvWMJ14, BvWWM17]. PEM-Based [BvWWM17]. Penalty [KD17, SGLS23]. Perception [HHZ^{+}21, MTNR21, NW24, NNL15, SWZ16, TP22, WSZN21]. Perception-Based [NNL15]. Perceptual [ACB17, BATS21, ECH23, ISZA15, JSR16, LCL23a, MGT15, STS21, SMA^{+}17, TKH21, VDM15, WSZN22]. Perceptual-Similarity-Aware [STS21]. Perceptually [IC19, MJ18, MR20, PBA^{+}23]. Perceptually-Transparent [MR20]. Percussive [BPG23, LKPR18, PSL17, TOKS14, TOS14]. Perfect [CCR16]. Performance [ATR^{+}21, BC17, CDLL21, CCC21, CZN^{+}24, DC16, DVO^{+}23, GTSM24, GMD21, HDG16, HCQ24, HNF20, HZF18, HLZ^{+}23, KPA^{+}24, LCZ23, LCY^{+}23, MTS^{+}19, NZL23, PCB14, PSL^{+}16, SH14b, SSKP16, TC22, WDDL18, WPY^{+}21, WDD^{+}16, YH21, YHC^{+}20]. Performances [LD16, NNS16, SHG21, WED16]. Performers [TP22]. Period [HAGSZ16]. Periodic [CCTO21, WHO^{+}21, WHT^{+}21]. Periodically [NZL23]. Periodically-Updated [NZL23]. Periodicity [CH15, CG21, CMD^{+}23, VDH24]. Periodicity-Based [CMD^{+}23]. Periphony [CM18]. Periphony-Lattice [CM18]. Permutation [KYTJ17, SDDCC15]. Perpendicular [SPM17a, SPM17b]. Person [QCW21]. Personal [GEC15, SLZ^{+}21, Wid19, ZCQ^{+}19]. Personality [SWZ16]. Personalization [CWYC17, NNL15]. Personalized [DZWC20, HWC^{+}17, JGD^{+}24, yLSY^{+}14, LXS^{+}18, hShWsL15, WW15, ZTL21]. Personalizing [GMFG16, LTWT17]. Persons [MHD^{+}17]. Perspective [GVMGO17, JBCH23, LMGG21, LYCX22, ZBCH21, ZMZ24]. Perturb [ZZZW23]. Perturbation [FLM23, HCKW23, TGGE21, ZSC^{+}23b]. Perturbation-Based [FLM23]. Perturbations [LMW^{+}24]. Perturbed [LHYJ23]. PES [Ano16-48, Ano16-49]. PESQ [XZF23]. PESQ-DNN [XZF23]. Phase [AL23, ARSA14, AM20, BV23, BHAMR19, CSY^{+}21, DB18, KBG16b, KG14, MBD18, MV19, MYNO23, MK15b, MK15c, Non14, OD17, PS19, PZW^{+}21, PCG17, PBS17, SLY14, TWIN23, VM15, WFN^{+}18, YWDQ21, ZWZ^{+}21, ZZ19]. Phase-Aware [DB18, MV19, ZZ19]. Phase-Correction [BHAMR19]. Phase-Derived [SLY14]. Phase-Enhanced [ZWZ^{+}21]. Phase-Scheduled-Command [PCG17]. Phase-Uncertainty [KBG16b]. PhaseDCN [ZWZ^{+}21]. PHAT [DGMB21]. Phonation [LRP^{+}23, WWG^{+}23]. Phone [DCLHJ18, DY22, Kam23, rLcHcC^{+}22, LCB^{+}23, WSY14]. Phone-Level [DY22]. Phoneme [BSWP18, CMY16, CB21, KBG^{+}23, LNNL22, LHD16, SFDRB21, WSZN21, YCM^{+}21]. Phoneme-Based [LNNL22]. Phoneme-Unit-Specific [CB21]. Phones [CZ17, TZW21]. Phonetic [AFG^{+}23, CGL^{+}15, CTSC14, CWJ15, CWJ16, FL19, HWW23a, KTH^{+}20, LLC23a, LCB^{+}23, LCY^{+}23, NAK^{+}17, TWC^{+}18, ZP15, ZGO19]. Phonological [LSB^{+}21, NNC20]. Phonologically [KKK15]. Phonologically-Structured [KKK15]. Phonology [NNC19]. Phonology-Augmented [NNC19]. Phrase [SKT21, ZSB17, ZLS^{+}21]. Phrase-Based [SKT21]. Phrase-Independent [ZSB17]. Phrasing [LSB^{+}21]. Physical [GTS^{+}19, GWQ23, SLZ^{+}21]. Physically [VDC^{+}17]. Physiologically [CMG^{+}17]. PI [MMZ23]. Piano [CMD^{+}23, CDW16, ES16, KLS^{+}21, LD16, NYD17, NYS17, ÖM24, SBD16, WED17, WPY^{+}21, WY23]. Pieces [SBV17]. Pilot [CGL^{+}24]. Pinna [SMO^{+}21]. Pitch [BTBT18, FWM^{+}14a, FWM^{+}14b, GB14, GJ22, HAGSZ16, HW14, HKI23, KAJC16, Lin19, MTBB15, RDP14, SM18, WYH17, WAJL15, WHO^{+}21, WHT^{+}21, YWT23, ZZN^{+}16, DHP14]. Pitch-Based [RDP14]. Pitch-Controllable [YWT23]. Pitch-Dependent [WHO^{+}21, WHT^{+}21]. Pitch-Synchronous [SM18]. Pitch-Time [MTBB15]. Pitched [Ziv15]. Pitman [Chi15a, RV14]. Pitman-Yor [RV14]. Pivot [ZLZ21]. Placement [GFS22, KBV16, KCD20]. Planar [ABB^{+}23, BBJ18, CT14, HCB20, SPM17a, SPM17b]. Plane [ADM^{+}19]. Planning [RLK14, YLFL22]. Platform [GH23b]. Platforms [KAA^{+}24]. Play [HCKW23].

Playback [GLQ21]. Playing

[SLY14, WBLC22]. PLCA [AB15]. PLDA [Bor22, Cum15, CL17a, CL17b, CL18b, LM15, LMC17, MPC16]. PLDA-Based [CL17b]. Plots [CLL^{+}21a]. Plug [HCKW23]. Plug-and-Play [HCKW23]. Plus [HOF15, WFY^{+}22]. PN [XNP^{+}18]. PNCC [KS16]. PNLMS [DC16]. PoE [ZDZ^{+}23]. Point [FB18, KCMW17, MP20, XO16]. Pointer [SZW23]. Points [JWBR21, LYH^{+}22]. Polarities [XNP^{+}18]. Pole [SD16b, SKH^{+}15a, VP17]. Pole-Zero [SD16b]. Policy [RC23, SWC20, ZYW^{+}24, ZTC23]. Polyglot [CHWL14]. Polynomial [CC15, MJZ15, NEN21, NEWN23, YSCZ15]. Polyphonic [AGM23, AB14, AB15, BLLP17, GGL23, HWT^{+}17, KG16, MBTB14, NYS17, NR14b, ONKP16, SBD16, SY15, ZCYZ18, rPH^{+}17]. PoLyScriber [GGL23]. POMDP [GY14]. POMDP-Based [GY14]. Pooling [MSB18, OGZ18, SR16b, YDW^{+}21, ZSH20, ZDZ^{+}21]. PoP [HNT^{+}23]. PoP-IDLMA [HNT^{+}23]. Popular [SBV17]. Port [BBFS21]. POS [LZC^{+}14, LCZ^{+}17, YZL^{+}18, ZYF18]. Position [Non14, TH16]. Position-Selective [TH16]. Positions [BA16]. Positive [LSYN19, TLD^{+}23]. Post [BBC^{+}16, DLL^{+}21, May18]. Post-Editing [BBC^{+}16]. Post-Filter [May18]. Post-Processing [DLL^{+}21]. Posterior [CY15, Cum15, LLLL19, CPL14]. Posteriorgram [ZGO19]. Posteriori [KLK17]. Postfilter [SGH17c, YKPMP18]. Postfiltering [CRVB^{+}15, MD18]. Postfilters [TTB^{+}16]. Potential [MK15c]. Power [Ano16u, Ano16v, BKS^{+}18, CC14b, ÇOH18, GTSM24, GI15, KS16, KD18, LKHU14, MIYO14, MPAEBLB15, NCB23, QDSL19, SAB16, SK15, TWSS16, WGD15, WLF23, YLD15, dlCSVAGAPM15]. Power-Law [SAB16]. Power-Normalized [KS16, dlCSVAGAPM15]. Powered [SKH^{+}15a]. PQG [LCZ23]. PQG-A2SA [LCZ23]. Practical [MS23]. Pre [CCL^{+}21, KCL21, LLW21, LHL^{+}22a, LHZ^{+}24, MLL^{+}22, MIR^{+}21, NTA15, NTO^{+}23, QXW^{+}23, SHL^{+}21, WDW^{+}22, WZLL24, WWX^{+}24, YCL22, ZZS^{+}23]. Pre-Trained [KCL21, LLW21, LHL^{+}22a, NTA15, NTO^{+}23, SHL^{+}21, WDW^{+}22, YCL22, ZZS^{+}23]. Pre-Training [CCL^{+}21, LHZ^{+}24, MLL^{+}22, MIR^{+}21, QXW^{+}23, WZLL24, WWX^{+}24]. Precision [FNA15, RVB14, XYH^{+}21]. Precompensation [DvWDM14]. Precompensator [Wid18]. Predicate [ZZZ16]. Predict [LCL^{+}21]. Predict-Then-Decide [LCL^{+}21]. Predicting [AdHTJ16, JT16, MASC23, YLLC19, ZL21b]. Prediction [AL23, ARSA14, AdTJ18, BH18, CSS^{+}19, CPCBFSF16, CLSyL21, CLZ22, DST^{+}19, DDMvW20a, ECJF21, FLL^{+}23, GCvWMJ14, HBT^{+}21, HZL23, HBCC22, HQTX23, JNP^{+}21, JT14, JvWGD15, LQW^{+}24, LYWX22, LLW21, MJZ15, MAP14, MP16, NR14b, PJTJ24, MG19, SAB16, SF14, TNS^{+}18, WYLZ16, WWL21, WYY24, WSFS19, XZL^{+}21, XBZ^{+}24, YK14, YC22, YTF^{+}23, ZTM^{+}23]. Prediction-Based [BH18, JvWGD15, WYY24]. Predictive [AS17b, GTSM24, JG19, JPX^{+}23, LCL^{+}21]. Predictors [CS19, GCJ^{+}14]. Preference [CSS23]. Preliminary [BS21]. Preordering [KCA21a, KCA21b]. Presence [Bor22, DTC16, KGH^{+}16, PJTJ24, TDL19, WSM19, YND14]. Presentation [MKMDM17, RLK14]. Preservation [ABKP19, HMDG15, IC19, MHD15, MHGD15, MD18, SNTD19, SBVM16, dCBC23]. Preserving [DTN^{+}17, NM21, RSR^{+}19, TE23, XWY^{+}23]. Pressure [HF15, HAS21, OFF^{+}17, ZSC^{+}23b].

Pretrained [CHL^{+}24, KWZ24].

Pretraining [GCG21, HHW^{+}21, LZD^{+}23]. Prewhitening [BvWWM17]. Primary [HTG14, HGT15a, HGT15b, LLS24]. Primary-Ambient [HTG14, HGT15a, HGT15b, LLS24]. Prime [SLC^{+}15]. Principal [IIY16, ZGL^{+}20]. Principle [LWK^{+}22]. Principled [RGK24]. Prior [BV23, CWUS22, CS19, HNT^{+}23, HIA^{+}17, KGG19, MPC^{+}24, PS17, SBN^{+}19, XFH^{+}23, YG24]. Prior-Based [MPC^{+}24]. Priori [AVM19, CH15, EMTF17, SFK15]. Priors [DBLČ20, ET18, FK24, JvWGD15, LBR16, RG18, SLL^{+}16, WYY24]. Privacy [DRK18, NM21, RSR^{+}19]. Privacy- [DRK18]. Privacy-Preserving [NM21, RSR^{+}19]. Private [OFF^{+}17]. Probabilistic [BCA15, CPL14, LBR16, MG18, MG19, TWSS16, WLL15b]. Probability [KGH^{+}16, MWT^{+}23, PJTJ24, TDL19, WSM19, ZLZW21]. Probability-Aware [MWT^{+}23]. Problem [AM18a, AKN^{+}16, DCYY22, DDMvW20b, GWCT23, HGYN23, KHHJ19a, PWS^{+}23, SDDCC15, WZH22, ZZXH22]. Problems [LGL^{+}22]. Procedures [TD23]. Process [GLWQ22, HSN18, JRA17, RV14, TJV24, TP16, XO16]. Processed [AdHTJ16]. Processes [GY14, KK19b, MIYO14, NNL15]. Processing [Ano14-45, Ano14-54, Ano14-55, Ano14-56, Ano14-47, Ano14-48, Ano14-49, Ano14-50, Ano14-51, Ano14-52, Ano14-53, Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano14-39, Ano14-40, Ano14-41, Ano14-42, Ano14-43, Ano14-44, Ano15-40, Ano15-41, Ano15-42, Ano15-43, Ano15-44, Ano15-45, Ano15-46, Ano15-47, Ano15-48, Ano15-49, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano15-35, Ano15-36, Ano15-37, Ano15-38, Ano15-39, Ano16-27, Ano16-28, Ano16w, Ano16x, Ano16y, Ano16z, Ano17-60, Ano17-38, Ano17-39, Ano17-40, Ano17-41, Ano17-42, Ano17-43, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano17-51, Ano17-52, Ano17-53, Ano17-54, Ano17-55, Ano17-56, Ano17-57, Ano18-44, Ano18-45, Ano18-46, Ano18-47, Ano18-48, Ano18-49, Ano18-50, Ano18-51, Ano18-52, Ano18-53, Ano18-54, Ano18-55, Ano18-56, Ano18-57]. Processing [Ano18-58, Ano18-59, Ano18-60, Ano18-61, Ano18-62, Ano18-63, Ano19t, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano21a, Ano21b, Ano22, CXX^{+}23, CTE21, CT14, DLL^{+}21, DBP^{+}24, FWM^{+}14a, FWM^{+}14b, FJT^{+}23, GH23b, KS21, KGG19, LFH^{+}15, MAB^{+}16, NEWN23, OISO14, QYCT23, SWW^{+}17, SNK21, SM14, SvWRK15, SGF^{+}21, TM14, TR15, VS14, WYO23, YND14, ZPØ^{+}21, Ano14-77, Ano14-66, Ano14-78, Ano14-67, Ano14-79, Ano14-68, Ano14-80, Ano14-57, Ano14-69, Ano14-46, Ano14-58, Ano14-70, Ano14-59, Ano14-71, Ano14-60, Ano14-72, Ano14-61, Ano14-73, Ano14-62, Ano14-74, Ano14-63, Ano14-75, Ano14-64, Ano14-76, Ano14-65, Ano14a, Ano15-50, Ano15-51, Ano15-52, Ano15-53, Ano15-54, Ano15-55, Ano15-56, Ano15-57, Ano15-58]. Processing [Ano15-59, Ano16k, Ano16l, Ano16-32, Ano16-38, Ano16-33, Ano16-39, Ano16-29, Ano16-30, Ano16-31, Ano16-34, Ano16-36, Ano16-40, Ano16-35, Ano16-37, Ano16-41, Ano17-59, Ano16i, Ano16j, Ano17r, Ano17z, Ano17s, Ano17-27, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano18v, Ano18w, Ano18x, Ano18y, Ano18z, Ano18-27, Ano18-28, Ano18-29, Ano18-30, Ano18-31, Ano18a, Ano19a]. Processor [YLD15]. Procrustes [DDMvW20b]. Product [BG21, CBC19, HNT^{+}23, HBCC22, IBC21, PBC18, SBP^{+}24, WBC^{+}21, WCC^{+}21]. Product-of-Prior [HNT^{+}23]. Production [ACB17, Sug22, TKT17]. Professional [MSEC21]. Profiling [PCG17]. Program [Ano17-68, Ano17-69]. Programming [ABA17, BC17, BC18, Kam23, SD16a, ZCYZ18, Zhe23]. Progress [LGH^{+}21]. Progressive [CDLX18, GYXY23]. Projection [CG20, CZN^{+}24, FdDPG21, FLB16, GLQ21, NZL23, SK18, TE23, ZN14]. Projection-Based [FLB16]. Projective [LKPR18, MAST15]. Promoting [NZL23]. Prompt [ZQF^{+}24]. Pronominal [BZ21]. Pronoun [YLGG22]. Pronunciation [CJ15, DKDN20, HBF^{+}16, LLC23a, PFLX24, QMS16, hShWsL15, WsL15]. Pronunciation-Based [HBF^{+}16]. Propagation [PMB^{+}19, TVJ15, WTQ^{+}19, XZL^{+}21]. Properties [BS14]. Proportionate [LRG21, LG16, PPSG17, ZN14, ZYG^{+}14]. Prosodic [Chi18a, LSB^{+}21, NAK^{+}17, YTF^{+}23]. Prosody [CHC^{+}14, CWH^{+}16, DY22, HWW16, LWYC23, OL15, QLW^{+}24, SZL19, ZQZ^{+}23, ZZS^{+}23]. Prototype [yLSY^{+}14, XXS^{+}23, SWC^{+}21]. PROTOTYPE-TO-STYLE [SWC^{+}21]. Proximal [GYdL^{+}21]. Pruning [SWEV23, TZZ15b]. PSD [KDJJ17, DDBW23, DDMvW20b, ET18, FSA18, KBG18, KDJJ16, RG17, TH17]. PSDs [LLGG20, LG20]. Pseudo [FCG24, GBH16, JKU18, LC24, WC18b, ZGC^{+}23]. Pseudo-Determined [WC18b]. Pseudo-FMLLR [JKU18]. Pseudo-Labeling [ZGC^{+}23]. Pseudo-Labels [LC24]. Pseudo-Whisper [GBH16]. Pseudospectral [BS14]. Pseudospectrum [MKA^{+}14]. PSFM [MG18]. PSG [WSJ^{+}23]. PSLA [GCG21]. Psychological [WWLL14]. Psychophysical [YLD15]. publication [Ano14-77, Ano14-78, Ano14-79, Ano14-80, Ano14-69, Ano14-70, Ano14-71, Ano14-72, Ano14-73, Ano14-74, Ano14-75, Ano14-76, Ano15-50, Ano15-51, Ano15-52, Ano15-53, Ano15-54, Ano15-55, Ano15-56, Ano15-57, Ano15-58, Ano15-59, Ano16-38, Ano16-39, Ano16-40, Ano16-41]. Publish [Ano14-90, Ano14-91]. Publishing [Ano18-43]. Pulse [DLG18]. Pulses [PT14]. Purpose [GKC16, NTO^{+}23]. Pursuit [WZS14]. Pyramid [OGZ18]. Pyramidal [ZSH20].

Q [SWC20, YPL21]. Q-Gradient [YPL21]. Q-Learning [SWC20]. QA [ZLW^{+}24]. QDM [NDS^{+}23]. QDM-SSD [NDS^{+}23]. QIM [LJK17]. Quad [SW16]. Quad-Based [SW16]. Quadratic [ABA17, BSB14, Wid18]. Quality [AKG^{+}17, BNvdB^{+}20, BvWM18, BMVG17, Bre15, CSS^{+}19, DRK18, FBE19, GLW20, KA16, KNH^{+}18, LYY24, LFU^{+}17, MGT15, NW24, NDS^{+}23, SSFP18, TKH21, VP17, XPC^{+}22, XZF23, ZLJB23]. Quality-Aware [NDS^{+}23]. Quality-Based [DRK18]. Quantification [GA20, LCZ23]. Quantified [ZWL^{+}23a]. Quantile [TP19]. Quantization [CJ15, MK16, WCY15, XYH^{+}21, YHO^{+}18, ZL21a]. Quantization-Aware [ZL21a]. Quantized [DGCY23, WTY^{+}20b]. Quasi [ARSA14, WHO^{+}21, WHT^{+}21]. Quasi-Periodic [WHO^{+}21, WHT^{+}21]. Quaternion [GWSC23]. Query [CLZ^{+}24, yLsL14, MAP14, RAB18, TZL^{+}23, WJ15a]. Query-by-Example [MAP14]. Query-by-Singing [WJ15a]. Query-by-Singing/Humming [WJ15a]. Query-Efficient [TZL^{+}23]. Question [ACK21, BGD^{+}18, KCL21, LWJ19, SKT21, WQ24a, XFH^{+}23, ZMC^{+}24, ZLZ^{+}24, ZLW^{+}24, ZXH^{+}16, ZBW^{+}24]. Question-Directed [ZLZ^{+}24]. Question-Specific [LWJ19]. Questions [KG19, XPC^{+}22]. Quo [XXWY24].

r [SGB22]. Radial [ATP16]. Radiance [BRD15]. Radiated [XWWW18].

Radiation [CATS20]. Radiativity [MSS17]. Radio [LWY^{+}23]. Radio-Echo [LWY^{+}23]. RadioSES [OWW^{+}23]. Radius [JEP14]. Rake [WST^{+}21]. Random [yLsL14, NYD17, NLYK24, OVN21, PMMM15, SAK16, XCZ^{+}23, YLC^{+}24, YLL15, ZH15, ZZL^{+}23]. Randomization [BF18]. Randomized [ZH14]. Randomly [KSK21]. Rank [BIK^{+}18, CJ15, EIPB^{+}19, GQXZ19, HOF15, JEvW23, LLGG20, MTK^{+}20, PKA14, PCB16b, QYCT23, SIKN23, SLGZ15, WYY24, SMVW14]. Rank-Based [PCB16b]. Rank-Deficient [LLGG20]. Ranking [BWZC22, LQW^{+}24, LHL^{+}23b, LXS^{+}18, ZLLS16]. RANSAC [TS14]. Rao [SGH17b, WCY19]. Raphson [BBFS21, FB19]. Rapid [DKO^{+}18, KWGW17, VTN15, XLLY18]. RARS [SSGL21]. Rate [AHH^{+}19, AHH^{+}20, BBM17, CLAG16, CHC^{+}14, CTE21, HNF20, HZBW23, JBTL15, LJK17, LCWC16, LCB^{+}23, MYKO24, MOT^{+}23, SH14b, SKH^{+}15b, WD16, ZG16, ZHH18]. Rate-Constrained [AHH^{+}19, AHH^{+}20]. Rate-Controlled [LCWC16, CHC^{+}14]. Rate-Distributed [ZHH18]. Rater [WZC23]. Ratio [BA19, CASZ17, CC14b, LWZG21, SAC17, SK15, WWW16, XGKM19, ZM20]. Rational [WTHJ14]. Ratios [CWW14]. Raw [DQY18, JBTA19, LYB^{+}23, WHO^{+}21, WHT^{+}21]. Ray [MAST15, MAB^{+}16]. RBA [YHL^{+}23]. RBA-GCN [YHL^{+}23]. RBM [SP16]. Re [LHL^{+}23b, PS17]. Re-Estimation [PS17]. Re-Ranking [LHL^{+}23b]. Reach [DRFO14]. Read [CLC^{+}22]. Readable [BK24b]. Reading [AF23, LYCX22, MSW^{+}22, PLT^{+}18, ZLZ21, ZZZ22a, ZZZ22b, ZZD23, ZZZL22]. Real [AAA19, HZF17, KS21, LYB^{+}23, MAZ21, NNS16, PMB^{+}20, PWS^{+}23, SK17, TZW21, XCHX24]. Real-Time [KS21, MAZ21, NNS16, SK17, TZW21]. Real-Word [AAA19]. Real-World [PWS^{+}23]. Realistic [SP15]. Reality [RG15b]. Reallocation [ZSIA16]. Reasoning [LSG^{+}23a, LLC^{+}23b, WZZ^{+}24, WWX^{+}24, ZZJ^{+}23, ZLZ^{+}24, ZBW^{+}24]. Recall [LQW^{+}24]. Receptive [CE15, KEzW21, WWS^{+}18]. Recognition [AHMJ^{+}14, AZK15, Abd18, AB18, AS17b, ACRS14, AZGN14, BVGV15, BDV^{+}14, BYT^{+}21a, BYT^{+}21b, BHS^{+}16, BSWP18, CWL22, CDLL21, CM15, CLW^{+}16, CWYC17, CDLX18, CLW^{+}19, CCH^{+}21, CB14, CP21, CLSyL21, CL14b, CL17b, CL18c, DHD^{+}19, DXZ^{+}18, DXW^{+}23, DB15, DCLHJ18, DH14, DZF^{+}23, DG23, ESVK17, FSA20, FYT^{+}21, FCCA23, FZM24, FLMS16, FNM^{+}19, FLC^{+}23, FLYJ23, GFG^{+}14, GDC19, GI15, GBH16, GH17, GH23a, GWL^{+}24, GJ17, HK24, HLL15, HYI^{+}15, HZW^{+}22, HZC^{+}22, HSWC21, HXL^{+}21, HHC16, IK23, JK19, JGD^{+}24, JGSC^{+}22, KGP14, KH21, KS16, KCMW17, KBG^{+}23, KTH^{+}20, sLGyLaC15, LKC^{+}21, LFO^{+}24, LKHU14, LWZ^{+}14, LDGHU14, LS14a, LWZZ16, LWM^{+}20, LWCC21, LWK^{+}22, LDG^{+}23, LSZ23, LSH^{+}23, LLWQ24, LLT21b, LWYD23, LWY^{+}23, LH14, LS14b, LK16, LGH^{+}21, rLcHcC^{+}22, LSG^{+}23b, LKR23, LLL24, LB19, LGR14, LR17, LFM^{+}23, MWT16, MK15a, MCL22, MHD^{+}17, MC17, MH16, MAK15, MKA16]. Recognition [MTS^{+}19, NYD17, NW14, NXCL16, OS21, OHN16, PSL^{+}16, PKR^{+}24, PHS^{+}24, QTY16, QBTY16, QZLH21, QZ22, RGR^{+}22, RPK23, RH18, RWF16, RGK24, SAB16, SV17, SP16, SWW^{+}17, SHJY21, SSFL22, SGY^{+}24, SAS17, SFIL17, SSGL21, SCY22, SBM^{+}19, SHDG16, SSKP16, SLK17, SKM17, SKAK16, SAK16, SBG17, SLW15, SLWY20, TD16, TQH^{+}18, TLWV17, TB18, TYW^{+}23, TR14a, TZL^{+}23, TDL19, TP22, WS14, WW16, WWS^{+}18, WLT^{+}18, WTW20, WBLC22, WH22, WCZR23, WWLL14, WYSD15, WL19, WCL^{+}21b, WK24b, XH23, XDH^{+}17, XHS^{+}18, XLLY18, XYH^{+}21, XAHJ^{+}14, YLQZ24, YTWB19, YLK^{+}21, YZW^{+}21, YCL22, YHL^{+}23, ZF14, ZCDS15, ZH18, ZS20, ZWS^{+}23, ZWW^{+}23b, ZJD^{+}15, ZDZ^{+}21, ZLG^{+}22, ZCY^{+}22, ZGC^{+}23, ZZZD23, ZCW^{+}24]. Recognizing [XDC^{+}17, YsL15]. Recombination [LWYD23]. Recommendation [DZWC20, HPB15]. Reconfigurable [MWL22]. Reconstructing [MLP^{+}24]. Reconstruction [CNFG23, ECH23, GCG^{+}17, HKPR21, KFH^{+}14, KG14, LSGL21, MYNO23, PBS17, QLW^{+}24, SSI21, SB18, TWIN23, WFN^{+}18, XWWW19]. Reconstructions [NLL^{+}18]. Recording [KFH^{+}14, LCLM18, NHK16, SSGL21]. Recordings [AS17a, DM16, KM23, KG16, NCC15, SBG17, TS16, WED17]. Recovery [HLL15, MBD18, PK15, SAA15, YLGG22]. Rectangular [DAMvW15]. Recurrent [CLC^{+}15, CLW^{+}16, CLHK21, DHD^{+}19, DAGS^{+}18, FYT^{+}21, GBL24, GTSM24, GAPGG19, HKHJS15, KK19a, KYTJ17, LTWT17, LKC^{+}21, LWZG21, LAGD18, LCW^{+}16, MDY^{+}15, NTA15, SF18, SNS15, TWZ^{+}18, TW20, TLWV17, WHB^{+}21, YMM19, ZXSD17, ZPW23, rPH^{+}17]. Recurrent-Neural-Network-Based [LTWT17]. Recursive [BG23, DDMvW20b, DWX^{+}16, DG15, EIPB^{+}19, NNC20, RG17, SG14, hShWsL15]. Reduce [ZWTL23]. Reduced [GTV19, HL14, KD24, PCB16a]. Reduced-Order [PCB16a]. Reduction [ATR^{+}21, AHH^{+}19, AHH^{+}20, AM16, AM18b, BH18, CZ17, CSZZ24, CCC21, DST^{+}19, DDMvW20a, Emu18, Emu20, HCH23, HBLC14, JTHN14, JBC16, KD16, LPHCB14, LGGH18, LFWZ23, MHD15, MHGD15, MD18, MBTB14, NWM16, RBBG15, RvWM22, SM14, SGM^{+}15, SEM^{+}14, SGH15b, SGH16, SGH17b, SGH17c, SMVW14, SK18, SBVM16, TM16, TK14, UNI^{+}24, WCC^{+}21, WHB^{+}21, YKPMP18, ZCHH18, ZHH18, ZL21a, ZZ22, ZTDD23b, dSCB24]. Reference [GLW20, JJB16, MMS^{+}20, SSFP18, ZLHL21, ZCDH21, ZDGO21, ZZZ22b]. Reference-Free [SSFP18]. References [SLOVB15]. Referencing [FFSS17]. Refinement [ÇOH18, FYG^{+}21, FWRJ21]. Refining [LLZ^{+}23d, YPC^{+}22, YWLZ18]. Reflecting [SMSV17]. Reflection [BAS20, MKA^{+}14]. Reflections [MFH21, TP15, Ter16]. Reflector [RJCW17]. Regeneration [LRWG23]. Region [MTKS21, ZKKC22]. Regional [WYLZ20]. Regressing [KLS^{+}21]. Regression [CNFG23, CDLL19, DHSG15, DTDL16, GR14, GHAP17, HGAPB15, JRA17, KK19a, LLW23, LS14b, MAL18, PMMM15, QDSL19, WS14, XDDL15]. Regression-Based [WS14]. Regressive [HBB17]. Regularity [SBV17]. Regularization [ADM^{+}19, DMHZ22, DC16, KHTL18, KEzW21, LGTG16, LGHG17, RHvdP^{+}16, VRN18, WGR^{+}18, ZWCT15, ZZLZ19, ZYD^{+}23]. Regularized [BCB15, QVZY24, STBC17, SKT21, UNI^{+}24, YXX^{+}22]. Regularizing [ZTWY19]. Regulated [JLP^{+}21]. Reinforced [YYZY24]. Reinforcement [AC24, AGC23, CCT^{+}19, CH14, LCW^{+}18, LPL^{+}22, RC23, WBSG18, WZH22, ZLL^{+}24a]. Reject [FNM^{+}19]. Rejection [HF21]. Rejuvenation [JWH^{+}22]. Related [APWB23, BKK15, BHAMR19, GSA18, LFH^{+}15, MBM15, MPKS^{+}23, PAB19, RF19]. Relation [CXvG^{+}23, CWR^{+}22, CCC^{+}24, FFSS18, FWF17, HZC^{+}22, JLZD24, JQL24, LWL^{+}19, LSG^{+}23a, LHL^{+}23b, LLWJ16, MZL21b, XZL^{+}21, XLZ^{+}23, ZZJ^{+}23, ZQL22, ZLZ^{+}24, ZJL21]. Relation-Aware [ZLZ^{+}24]. Relational [FLL^{+}23, XZL^{+}23, YHL^{+}23]. Relations [LWJ19, LPL^{+}22, LJL^{+}23, XLG^{+}23, ZZH23, ZWW^{+}23b]. Relationship [KTJ19, YLLC19]. Relationships [MVH16]. Relative [CG20, CB14, FPTJ17, FPTJ18, HAS21, JLP^{+}21, KMG15, LCW^{+}21, SGH15b, YLL21, ZHH19, ZDD21]. Relativistic [CLL^{+}21b]. Relaxation [YLJ16]. Relaxed [KHHJ19a, LZHN14, MYK^{+}24]. Relaxed-Distortionless [MYK^{+}24]. Relevance [LZ16, LCC^{+}15]. Relevant [BYY^{+}23, IC19]. Reliability [CSY^{+}21]. Reminiscent [HKPR21]. Remote [AFdDG23, SBVM16]. Rendering [HAGA21]. Reorganization [HWW23a, MBTB14]. Repeated [JFG^{+}17]. Repeats [NNS16]. Replacement [WZLL24]. Replay [XYY^{+}23]. Representation [AGW15, ADM^{+}17, BA19, BHAMR19, CXvG^{+}23, CHL^{+}19, CHGS23, CWBv19, DCYY22, DDZL19, GJYW23, GBF21, GGZ23, HRL^{+}21, HWW23b, HBT^{+}21, KYL24, KLCL24, KSG24, LSH^{+}23, LHW^{+}21, Lib21, LM22, LLyL21, LSG^{+}23b, LFWZ23, LLL^{+}19, MHZ^{+}24, MBD^{+}15, PKA14, PEO^{+}20, Rak17, STS21, SAS17, SUK24, SG21, SZL19, SKH^{+}15b, VML14, WVCL14, WCL^{+}23, XHRC24, YXX^{+}22, YWW^{+}23, ZKA15, ZLZW21, ZMZ24, ZZZD23, ZPP15]. Representation-Based [AGW15]. Representations [AFG^{+}23, AB21, AJK24, ADGE19, BV17, CZZ15, CSAR^{+}14, ÇH21, DSMD15, DZLGO20, GLL^{+}21, GWSC23, HCZC24, KS21, LRP^{+}23, LLC23a, MMCF18, NNC20, NTO^{+}23, PHM^{+}16, QLW^{+}24, RG15a, RA23, SCY22, SY15, WSL^{+}20, YCLS17, YGV14, YČS23, ZLD20, ZYZU22, ZXZZ23, ZZS^{+}23, ZXH^{+}16, ZZZW23, vHK14]. Representing [WLL15b]. Reproduction [BHAMR21, DRK18, GFS22, HF21, HGT15a, HGT15b, JSR16, JK15, KBV16, KBV17, KFH^{+}14, LLS24, MFH21, SR14, SvWRK15, TZA14, UKS19, VDM15, ZA14]. Reranking [SKK14]. Rescoring [LCW^{+}16]. Research [GH23b]. Reserved [WYZ^{+}23]. Reshaping [MJ18, MMK21]. Residual [DGCA23, DDBW20, DDBW23, Emu18, Emu20, GLT23, ICB24, Lib21, RBBG15, SSGL21, TQH^{+}18, TCW19, WVCL14]. Residual-Echo [ICB24]. Resilient [EMXD20, YBC^{+}14, ZZX^{+}21, ZZX^{+}23a]. ResNet [WDW^{+}23]. ResNet-Conformer [WDW^{+}23]. Resolution [BP14, BZ21, DTDL16, EV23, KVMV^{+}24, KLCL24, KLS^{+}21, KKGH16, MPC^{+}24, PP19, RARR15, RF19, SKH^{+}15b, TOS14, VSK^{+}24, WW21, YLGG22, YDW^{+}21]. Resource [Ano16-48, Ano16-49, CM15, CSS23, CHL^{+}24, DZF^{+}23, FLYJ23, HLL15, HZW^{+}22, HOF15, JKC^{+}24, KMG21, KAA^{+}24, KSG24, LHL^{+}23a, LMW^{+}24, LHZ^{+}24, LGR14, QZ22, SML^{+}24, SV18, WQ24b, WSZN22, YTWB19, YGS19]. Resource-Limited [HLL15]. Resources [DBM15, NNC19]. Response [ADM^{+}17, BWH18, CCR16, CCTO21, CNFG23, CP21, ÇOH18, GLL20, GLL^{+}21, HCKW23, JEvW23, JCYwH21, KSK21, KD24, KBT19, LXS^{+}18, MYK^{+}24, MMK21, SWC20, TWSS16, XSL^{+}19]. Responses [BA16, DGCA23, DMP^{+}18, EV23, GS21, KB14, MCD14, TP15, Ter16, VDC^{+}17, YK21b, ZZH21]. Restaurant [MMBT21]. Restoration [DM16, KSK^{+}15, LYY24]. Restore [HCZC24]. Restricted [NTA15, NTM16, NTY19]. Results [MBK17]. RETF [DDMvW20b]. Retrieval [ARSLN16, GLL20, KSL19, yLsL14, sLGyLaC15, LCW^{+}18, PDS^{+}16, RRdM16, SSAS21, SWC^{+}21, VML14, ZM21, ZXH^{+}16, ZLL^{+}24a]. Retrieval-Based [GLL20]. Retrieve [CQ22]. Retrieve-and-Edit [CQ22]. Reuse [CXvG^{+}23, LCYW21]. Reused [ZLX^{+}24]. Reverberant [AGBC14, BKTR24, CW18, CASZ17, DW17, DDMvW20b, DPHG18, FSA18, FSA20, FSO17, FK19, FK24, GWQ23, GAM23, JWBR21, JWLF14, LKHU14, LH23, LMH23, MMB17, MTNR21, MGGK17, MCF21, NEN21, SAC17, SWCN19, VBY18, WWL21, WRK18, ZW17, ZWW14, ZWW15, ZWW19, ZM20]. Reverberation [BRD15, BB21, BKS^{+}18, CS19, DDBW20, DBN^{+}17, GSP23, HCJT15, KD18, KDJJ16, KDJJ17, LLGG20, LG20, LBR16, LYN14, Lin18, LLdCF23, NR14a, PLZL17, SH14a, SH14b, SGH17b, VP21, XGKM19]. Reverberation-Robust [Lin18]. Reverberator [PCLP14]. Reverse [DM16, LGL^{+}22]. Reversion [RJCW17]. Review [BHH22, MLL^{+}22, MSRND14, O’S24, WDS^{+}18, ZMX^{+}21]. Reviewed [TKH21]. Reviewers [Ano14-81, Ano15-70, Ano17-67, Ano18-66]. Reviews [XLG^{+}23]. Reward [HCC^{+}23]. Reweighting [LRG21]. Rewriting [BZ23]. Rhythm [GKC16, HS15, MFG15, NYS17, RDP14]. Rhythm-Based [RDP14]. Rich [SKK14, YZW^{+}18, ZP15]. Riemannian [QYCT23]. Riesz [AS15]. Ring [PC24]. Risk [CY15]. RLS [CC15, ZN14]. RNN [GG18, LY23, NTA15, PW22, SSK17, WSZN21, WSZN22]. RNN-Based [GG18, SSK17]. RoBERTa [KSK21]. Robot [TR14b]. Robotic [QWW^{+}23]. Robots [TR15]. Robust [AGM23, AV16, AS17b, AM18b, ATP15, ABKP19, AM20, AATKS21, BC17, BC18, BE17, BCB15, CBA^{+}15, CATS20, CLZ^{+}23, CCR16, CCTO21, CZ17, CCT21, CP21, DMHZ22, DGMB21, DWY21, DSH18, FYT^{+}21, FD21, GMH14, GWG^{+}14, GI15, GBH16, GCvWMJ14, GAPGG19, GB14, GJB15, HAGSZ16, HCQ24, HH14, HCBY18, HIA^{+}17, HCZC24, HHC16, IIY16, JFG^{+}17, KGP14, KAJC16, KK19a, KS16, KLCL24, KD17, KMG17, KSHH18, KHHJ19b, LWZ^{+}14, LDGHU14, LS14a, LM15, LMC17, LTE^{+}21, LLWQ24, LZHN14, Lin18, Lin19, LM23b, LETJ21, LHL^{+}22b, MMB17, MGB18, MPAEBLB15, MPC16, MLMT24, MMCF18, May18, MZX^{+}15, MAK15, NW14, NXCL16, NNC21, Non14, PCB16a, QTY16, QBTY16, QCDW17, QCW21, QCL23, MG18, RH18, RG21, SH14a, STH^{+}18, SDF18, SBM^{+}19, SW16, SKM17, SLW15, TD16, TQH^{+}18, TMM18, TMMZ18, TZZ^{+}15a, TWSS16, TVJ15, TMKO17]. Robust [TS16, TDL19, WW16, WED16, WYH17, WCCB19, WZW19, WW21, WDW^{+}22, XNG^{+}14, YGV14, YPL21, YLY15, ZZH16, ZSC^{+}23b, ZQZ^{+}23, ZLJB23, ZBC14, ZWW14, ZGZ23, ZCQ^{+}19, ZZZD23, dlCSVAGAPM15, VZC14]. Robustness [CH18, CBHC21, DLL^{+}21, HGT15c, LKR23, MBDV23, NW15, SK16a, WST^{+}21, WLL^{+}22, ZMZ24]. RODA [LGL^{+}22]. Role [DBM15, FWRJ21, MZL21a, PM23, WXS24, XWS^{+}21]. Roles [YW14]. Rolling [CDA^{+}14]. Room [AGV24, ADM^{+}17, BWH18, BA16, BGQO^{+}17, BHBS16, BS14, CCR16, CCTO21, CNFG23, DHCS15, DAMvW15, DGCA23, EGMN16, FdDARF16, GS21, HB17, JEvW23, KB14, KD24, MCF21, MMK21, MCD14, MNFGO19, MR20, PC21, PT21, SAPB15, SRH15, SWCN19, TZA14, TP15, Ter16, VDC^{+}17, XGKM18, YB14, YK21b, vMM14]. Rooms [MZA20, ZZZ^{+}21]. Root [DDMvW20b]. Root-Based [DDMvW20b]. Rotating [MXNB19]. Rotation [MGF18, RF19, SNK21, WYO23]. Rotation-Invariant [WYO23]. Rotor [YLH23]. RST [HW23]. Rule [ZP15]. Rumor [LHL^{+}22b]. RWTH [TDDN24].

S [MUK22, LRP^{+}23]. S-Vectors [MUK22]. S1.11 [YLJ16]. SA [LCRY24]. Safe [FNM^{+}19]. Samba [DLH23]. Sample [Bre15, PKR^{+}24, PZZZ22, WBSG18, ZDLK18]. Sample-Efficient [PKR^{+}24]. Sampling [BBM17, BKK15, BHAMR19, CTE21, GCG21, HLB^{+}24, HZBW23, MYKO24, WD16]. Sampling-Rate [CTE21]. SANet [LWLS24]. Sarcasm [LPL^{+}21]. Satisfaction [AMK^{+}20]. Saturation [ATR^{+}21]. Savitzky [KMG17].

SBSim [MVN22]. Scalable [MK16, VDC^{+}17, XO16, ZSL^{+}22]. Scale [CL14b, DH20, GaKGB21, HLB^{+}24, LYWX22, LZC^{+}23, LCRY24, RGK24, SNK21, WCC^{+}15, WWX^{+}23, ZWS^{+}23, ZCY^{+}22]. Scale-Frequency [WCC^{+}15]. Scattered [WCCB19]. Scattergrams [SUK24]. Scattering [AGD^{+}22, DHCS15, WBLC22]. Scenario [LH23, LMH23, ZWX15]. Scenarios [CHL^{+}24, HOF15, LH14, QCL23]. Scene [ATS18, BSER17, CLW24, HKM^{+}24, IO17, IO19, LLZ^{+}21c, MF17, NK22, PEO^{+}20, PHM^{+}17, RG15a, Rak17, RVB^{+}17, TALP24, WYL^{+}23, YK17, ZW19, Ano16-52, Ano16-53, Ano16-54, Ano16-55, Ano16-56, Ano16-57]. Scene-Aware [CLW24, LLZ^{+}21c]. Scene-Event [HKM^{+}24]. Scenes [JNP^{+}21, LLR^{+}16, MHB^{+}18, SMA^{+}17, SMSV17, TMKO17]. Scheduled [PCG17]. Scheduling [AA17]. Scheme [BKK15, CM18]. Schemes [FFSS17, FdDARF16, vMM14]. Schizophrenia [HHZ^{+}21, HFL^{+}23]. Scope [JYC^{+}24]. Score [Bor23, CJ15, DXW^{+}23, DM16, HXWS15, KH16, KNH^{+}18, LD16, LCZ23, NNS16, TMM18, WED17, WK24b, ZM21, ZLHL21]. Score-Aging [KH16]. Score-Informed [DM16, WED17]. Scores [Cum21, CS23, MASC23, RGR^{+}22, VMOL16, YWLZ18]. Scoring [CJ15, Cum15, CL18b, Kam23, LG20, MIR^{+}21, ZYW^{+}20]. Scoring-Based [LG20]. Screams [SK16b]. Screening [MP20]. SDP [HGYN23]. SDP-SOCP [HGYN23]. SDW [ZTDD23b]. SDW-SWF [ZTDD23b]. SEAMNET [BB21]. Search [HBF^{+}16, HXC^{+}22, MMBT21, OR16a, YCM^{+}21, YGS19, YČS23, ZP15, ZPD19]. Second [FWRJ21, JGHN23, TC19]. Second-Order [FWRJ21, JGHN23, TC19]. Secondary [ATR^{+}21, AKN^{+}16, GLQ16, GdC14, MX17, WAS16, ZHB^{+}18]. Secondary-Path [GdC14, MX17]. Section [LFH^{+}15, RVB^{+}17]. Segment [MCL22, SgL20, WLL^{+}15a, vNKB^{+}23]. Segment-Based [MCL22]. Segment-Less [vNKB^{+}23]. Segment-Level [SgL20]. Segmental [MAP14, ZGO19]. Segmentation [BSL15, BSL16, BD16, CMY16, CHL^{+}19, DP16, FSN24, HEMN21, KJG16, Kam23, KXS^{+}19, NJC16b, PACPJF^{+}23, SBV17, SMvdP22, TQW^{+}15, TGZ22, WLF23, ZYF18, ZZL^{+}19]. Segments [PHMB18, ZWC^{+}23]. Segregation [JWLF14, TSKO20]. Select [HLB^{+}24]. Selected [GaKGB21]. Selecting [WLT^{+}18]. Selection [AATKS21, BY15, CHWL14, GWLL22, GLL20, GLL^{+}21, HSLB23, JCYwH21, KSK21, KNH^{+}17, LAK16, NAK^{+}17, TWZ^{+}18, WUF^{+}18, XYH14, ZCF14, ZCHH18, ZCDH21, ZDD21, ZHZ19, ZYW^{+}20]. Selective [AGC23, BSTM24, LFO^{+}24, LYL21, SGL^{+}23, TH16, TG24, XRWL21]. Selectors [FLYJ23]. Self [CWL22, DYY^{+}23, DLH23, FLM23, FZM24, FLYJ23, GWQ23, GQXZ19, HCQ24, HWRC15, HBT^{+}21, Kam23, KKGK18, KWZ24, LC24, LSL24, LdLvWWS21, LLyL21, LLZ23b, MYL^{+}21, MCL22, ODNA23, PW21, PW22, SAA15, SG21, SRC^{+}23, TLD^{+}23, TM22, WLL^{+}22, ZZX^{+}23b, ZM23, ZZZD23, ZCW^{+}24]. Self-Attending [PW22]. Self-Attention [GQXZ19, MYL^{+}21, PW21, SRC^{+}23, TM22, ZZX^{+}23b]. Self-Attention-Based [ODNA23]. Self-Attentive [LdLvWWS21, ZM23]. Self-Distillation [KWZ24]. Self-Informed [KKGK18]. Self-Learning [MCL22]. Self-Localization [HWRC15]. Self-Recovery [SAA15]. Self-Supervised [CWL22, DLH23, FLM23, FZM24, GWQ23, HCQ24, HBT^{+}21, Kam23, LC24, LSL24, LLyL21, SG21, TLD^{+}23, WLL^{+}22, ZZZD23]. Self-Supervision [ZCW^{+}24]. Self-Training [DYY^{+}23, FLYJ23]. Semantic [AVK^{+}15, BM15, CCH^{+}21, DBM15, FWRJ21, FGQ^{+}15, HM14, JLZD24, KSL19, yLsL14, yLSY^{+}14, LSG^{+}23a, MZL21a, NNC20, PACPJF^{+}23, SFIL17, WLZH21, WTHJ14, WXS24, XV21, XWS^{+}21, YW14, YHC19, YZW^{+}18, YUK21, ZZH23, ZZY^{+}24, ŞUY^{+}18]. Semantic-Key-Controlled [YZW^{+}18]. Semantics [LHW^{+}21]. Sememe [BK24b]. Semi [DZF^{+}23, GMD24, KK23, LGTG16, LGTG17, LAK16, LS14b, LCL^{+}23b, MS23, SBN^{+}19, SHL^{+}21, WXZ^{+}19, WK24a, XCHX24, ZGC^{+}23, ZB14]. Semi-Parametric [LS14b]. Semi-Supervised [DZF^{+}23, GMD24, KK23, LGTG16, LGTG17, LAK16, LCL^{+}23b, MS23, SBN^{+}19, SHL^{+}21, WXZ^{+}19, WK24a, XCHX24, ZGC^{+}23, ZB14]. Semidefinite [BC17, SD16a]. Semispontaneous [HWC^{+}17]. Semisupervised [DXZ^{+}18, LK16, XHS^{+}18]. Senone [FLMS16, TMMZ18]. Senone-Based [FLMS16]. Senones [JSMD18]. Sense [WZ24]. Sense-Wise [WZ24]. Sensing [Chi15b, KMM^{+}18, MCD14, WLL^{+}16, WZS14]. Sensing-Based [WLL^{+}16]. Sensitive [CH16, MHZ^{+}24, YJX20]. Sensitivity [MBDV23, WCY19]. Sensor [AM18a, AHH^{+}20, BBM17, CCT21, CMM^{+}23, CG17, CPSFC^{+}14, DMHZ22, FdDPG21, GWLL22, GCG^{+}17, HZBW23, HSLB23, KSHH18, KCD20, LHJ^{+}23, MMS^{+}20, RvWM22, SHAGPRZ17, TC22, WCY19, WRK18, ZHH18, ZHH19, ZDD21, ZTDD23a]. Sensors [HF15, MAZ21]. Sensory [BE17]. Sentence [CZG^{+}15, CWU^{+}19, LWL^{+}19, LHW^{+}21, LFQL19, MSFK15, MC17, NAK^{+}17, OGZ18, PDS^{+}16, QZLH21, QWL^{+}19, TWZ^{+}18, TQW^{+}15, WUF^{+}18, ZM17, ZYW^{+}20, ZZZW23, MVN22]. Sentence-BERT [MVN22]. Sentence-Level [CWU^{+}19, OGZ18, QZLH21]. Sentences [HS15, PLT^{+}18, RCL18]. Sentiment [BLZ21, CZG^{+}15, CQ22, CHGS23, DJY^{+}19, HQTX23, JBWZ24, JYC^{+}24, KSH17, LJX^{+}16, LYL21, LFL^{+}23, LCRY24, LHL^{+}22a, LZL^{+}23, LHYJ23, LZD^{+}23, MXH21, TQW^{+}15, WYLZ20, WSG21, YWLZ18, YJX20, ZZH23, ZZY^{+}24, ZLC^{+}24]. SEOFP [LYH^{+}22]. SEOFP-NET [LYH^{+}22]. Sep [SPM17a, BSW^{+}24]. Separable [SZVZ16]. Separately [LLW21]. Separating [MGGK17]. Separation [AV16, AJLW14, AGBC14, BXY14, BV16, BSW^{+}24, BSTM24, BS21, CONVVC18, CW18, CWW14, CH23, CY16, CK19, ÇH21, DW17, DM16, DTDL16, DDNN19, EDO17, FSA18, FSO17, FK19, FK24, GVMGO17, GRSP17, GZZY23, HNT^{+}23, HCH23, HL15, HKHJS15, IIY16, IBN^{+}18, IISN21, JSYW16, JVR16, KML18, KSK^{+}15, KOS^{+}16, KYTJ17, KBGAP^{+}16, LKPR18, LBR16, LBR18, LSDM19, LGGH19b, LCQ22, LDG^{+}23, LKM23, LW19, LLdCF23, LDZ19, LCM18, LM19, LHM21, LY23, MBD18, MV19, MMT^{+}19, MTZ^{+}21, MDCS20, MTKS21, MTK^{+}20, NK21, NKS21, NW14, NW15, NR22, NLL^{+}18, NV14, NDV18, NDS^{+}23, NLV16, OOK17, OVN21, OVN^{+}22, OIY^{+}14, ÖM24, OWW^{+}23, PSL17, PWS^{+}23, QL24, RDP14, RLS^{+}18, RJW19, SOF15, SIKN23, SFDRB21, SFRK^{+}23, SGH17c, SGH17a, SHG21, SKDN14, SLOVB15, SRC^{+}23, SWCN19, TOKS14, TOS14, TH18]. Separation [TG24, TGZ22, TS14, UNI^{+}24, WWD14, WNW14, WRC16, WDDL17, WC18b, WC18a, WW19, WWL21, WWW21, WGLZ21, WCW^{+}21, WYY22, WCC^{+}23, WWW16, WRDP17, XCHX24, YPC^{+}22, YLL15, YDW^{+}21, YWW^{+}23, ZG21, ZW16b, ZZN^{+}16, ZW17, ZXY^{+}21, Ziv15, ZPP15, vNKB^{+}23]. Separation-Based [NDS^{+}23]. Seq2Seq [SRT23]. Sequence [CHL^{+}19, CMY^{+}21, CLZ^{+}24, CJY^{+}21, DDZS17, DSN18, FYG^{+}21, GQXZ19, HHW^{+}21, LCW^{+}21, SSFL22, SgL20, SZX^{+}18, TZS^{+}21, TL19, WSJ^{+}23, WZM21, WL19, WTQ^{+}19, WLL^{+}23, YZW^{+}18, YXX^{+}22, ZLL^{+}19, ZLD20, ZQZH19, ZTWY19, ZLD21]. Sequence-to-Sequence [CHL^{+}19, CJY^{+}21, DSN18, HHW^{+}21, LCW^{+}21, ZLL^{+}19, ZLD20, ZLD21]. Sequences [BCA15, CCR16, CCTO21, rLcHcC^{+}22, VP21, XNP^{+}18]. Sequential [FMD14, GLC^{+}14, LZAY23, LB23, RCL18, WCL^{+}21b]. Series [WSG21]. Service [MSW^{+}22]. Services [LWYC23]. Session [GZG23, LH14]. Session-Independent [GZG23]. Set [CLZ^{+}24, FD21, RARR15, TD16, WSY14, ZH15]. Sets [MKA16, OVN21, PAB19]. Severe [SAB16]. Shallow [JLZD24]. Shape [SKM19, YG24]. Shaped [LSWN23]. Shaping [KBT19, YHO^{+}18, ZTQB19]. Shared [JLKK21]. Shift [GJ22, NNC21, SKSD14]. Shifting [HGT15b]. Short [GWCT23, HCJT15, KTJ19, LWZZ16, PDS^{+}16, VH19, WHQY19, WCC^{+}21, YZLY17, YZW^{+}18, ZWC^{+}23]. Short-Duration [WHQY19]. Short-Term [PDS^{+}16, YZLY17]. Short-Time [HCJT15, KTJ19, VH19]. Short-Time-Fourier-Transform [WCC^{+}21]. Shortest [LSG^{+}23a]. Shot [ASC^{+}18, CWR^{+}22, DYY^{+}23, JKC^{+}24, XV21, ZMC^{+}24]. Show [DAFS15]. Siamese [ZPD19]. Sidelobe [ABvM19, DST^{+}19, DDMvW20a, ZZL^{+}23]. SIFTER [LHL^{+}22b]. Sign [KKJN17, LYH^{+}22]. Sign-Exponent-Only [LYH^{+}22]. Signal [Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano14-39, Ano14-40, Ano14-41, Ano14-42, Ano14-43, Ano14-44, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano15-35, Ano15-36, Ano15-37, Ano15-38, Ano15-39, Ano16-27, Ano16-28, Ano16w, Ano16x, Ano16y, Ano16z, Ano17-38, Ano17-39, Ano17-40, Ano17-41, Ano17-42, Ano17-43, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano17-51, Ano17-52, Ano17-53, Ano17-54, Ano17-55, Ano17-56, Ano18-44, Ano18-45, Ano18-46, Ano18-47, Ano18-48, Ano18-49, Ano18-50, Ano18-51, Ano18-52, Ano18-53, Ano18-54, Ano18-55, Ano18-56, Ano18-57, Ano18-58, Ano18-59, Ano18-60, Ano18-61, Ano18-62, Ano18-63, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano21b]. Signal [Ano22, BvWWM17, CWW14, DMVP16, EJ16, GWQ23, KSK^{+}15, KD17, KHHJ19b, LWZG21, MLP^{+}24, MYKO24, NEWN23, PCW15, PBA^{+}23, RMBK14, SWW^{+}17, SNK21, SNTD19, SUK24, SK18, SBVM16, VS14, VDM15, WYO23, YOM21, YZLY17, ZZX^{+}23b]. Signal-Dependent [DMVP16, KD17]. Signal-to-Noise [CWW14, LWZG21]. Signals [AKG^{+}17, BP14, BSWP18, CT14, ESJ17, HTG14, HAGA21, HPPM24, HKI23, JD17, LSW18, LJK17, LAP23, MMS^{+}20, MP16, NWM16, NR14b, SK17, SB18, TOS14, TZZ^{+}15a, VM15, WAS14, WHB^{+}21, XNRG15, XGKM19, ZG16, ZZX^{+}23a, Ziv15, ZXN^{+}21, vHK14]. Signature [AB21]. SII [HCJT15]. Silence [ZLL^{+}23]. Silent [BSWP18, KCMW17, MHD^{+}17]. Similar [LYY^{+}23, MWT16]. Similar-Pair [LYY^{+}23]. Similarities [PHM^{+}16]. Similarity [BA16, BM15, FMD14, FDK15, yLsL14, MVN22, OR16a, PHMB18, QWL^{+}19, STS21]. Similarity-Based [MVN22]. Simple [SMO^{+}21, ZYF18]. Simplification [QLL^{+}21, QLZ^{+}21]. Simplified [GLQ16]. Simulated [MPKS^{+}23, VSL21]. Simulating [LLR^{+}16]. Simulation [AGV24, BHBS16, BH19, CFMM14, HB17, Hel14, KBV17, vMM14]. Simulations [BSR^{+}23, DAMvW15, SvWRK15]. Simulator [AGD^{+}22]. Simultaneous [SM18, TK14, YHO^{+}17]. Simultaneously [XZSX23]. Sine [HLZ^{+}23]. Singer [HL15, KKJK23]. Singers [MSEC21]. Singing [CWJ16, DAR16, GLW20, HHO^{+}21, HL15, IIY16, KKJK23, KG16, MDCS20, MTBB15, NR22, SFDRB21, SW20, TOS14, WWG^{+}23, WTY^{+}23, WJ23, YDW^{+}21, YWW^{+}23]. Singing/Humming [WJ15a]. Single [AM18b, AY15, BXY14, BV23, CDLX18, CH23, DAFS15, DB18, DBN^{+}17, DTDL16, FD21, GWLL22, GH17, GRSP17, HSPK18, KG14, LK19, LH23, LMH23, LJWC23, LWL^{+}22, MK15b, MK15c, MBK17, PSL^{+}16, PT21, SGM^{+}15, TD23, Ter16, VP17, WFN^{+}18, WWD14, WDDL17, WYSD15, WRK18, XGKM19, XRWL21, YKPMP18, YLL15, ZTQB19, ZWZ^{+}21, ZTDD23b, ZPP15]. Single- [DAFS15, PT21]. Single-Channel [AM18b, BXY14, BV23, CDLX18, CH23, DB18, DBN^{+}17, DTDL16, FD21, GRSP17, KG14, LK19, MK15b, MK15c, MBK17, PSL^{+}16, SGM^{+}15, WFN^{+}18, WWD14, WDDL17, WYSD15, XGKM19, ZWZ^{+}21, ZTDD23b, ZPP15]. Single-Microphone [TD23, YKPMP18, YLL15]. Single-Source [LH23, LJWC23]. Single-Stage [HSPK18]. Singular [ZZX^{+}21]. Sinsy [HHO^{+}21]. Sinusoidal [AF19, AJB^{+}18]. Sinusoidal-Based [AF19]. SiPTH [MTBB15]. Sister [LV22]. Situated [VSK^{+}24]. Size [AA17, Bre15, CC14a, TM16, ZZH16]. Sized [HWC^{+}17]. Sizes [KKJN17]. Skip [WCZ21]. Skip-Gram [WCZ21]. Skips [NNS16]. SLAM [EN18]. Slavic [ZP15]. Sliding [ZN14]. Sliding-Window [ZN14]. Slope [GSP23]. Slot [JGHN23, LYCX22, MDY^{+}15]. SM [TQM^{+}23]. Small [Bre15, CWH^{+}16, HWC^{+}17, LR17, TD16, WCCB19]. Small-Footprint [LR17]. Small-Sized [HWC^{+}17]. Small-Spacing [WCCB19]. Smoothing [BOSR23, HWW16, JC17, MYC^{+}23, RG17].

Smoothness [TOKS14]. SMT [ARSLN16]. Snapshot [Ter16]. Snowman [MPKS^{+}23]. SNR [AY17, BV23, CH15, EMTF17, ET18, HBLC14, KK14, LM15, MK15b, RG21, SFK15, SLGZ15]. SNR-Based [RG21, BV23]. SNR-Invariant [LM15]. Social [KAA^{+}24, LTWT17, SFEB23, XHS^{+}18, ZZH23]. Society [Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano14-39, Ano14-40, Ano14-41, Ano14-42, Ano14-43, Ano14-44, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano15-35, Ano15-36, Ano15-37, Ano15-38, Ano15-39, Ano16u, Ano16v, Ano16-27, Ano16-28, Ano16w, Ano16x, Ano16y, Ano16z, Ano17-38, Ano17-39, Ano17-40, Ano17-41, Ano17-42, Ano17-43, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano17-51, Ano17-52, Ano17-53, Ano17-54, Ano17-55, Ano17-56, Ano18-44, Ano18-45, Ano18-46, Ano18-47, Ano18-48, Ano18-49, Ano18-50, Ano18-51, Ano18-52, Ano18-53, Ano18-54, Ano18-55, Ano18-56, Ano18-57, Ano18-58, Ano18-59, Ano18-60, Ano18-61, Ano18-62, Ano18-63, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r]. Society [Ano19s, Ano21b, Ano22]. SOCP [HGYN23]. Soft [CJY^{+}22, LCSL19, ZH19]. Soft-Target [LCSL19]. Softmax [LKC^{+}21]. Softmax-Free [LKC^{+}21]. Solo [AS17a]. Solution [FB19, LWK^{+}22, PCB15, dCBC23]. Solutions [NV21, WW15]. Solving [GWCT23, LGL^{+}22, SDDCC15, ZZXH22]. Some [ZP15]. Song [DHYG14, HC19, SW20, WY23]. Songs [FDK15]. Sonority [SP17]. Sound [AKUS23, AGM23, APH14, AM18a, ADM^{+}17, BRC24, BOL^{+}18, BLLP17, BvWM18, BAS20, BATS21, CONVVC18, CC21, CSY^{+}21, CPSFC^{+}14, ÇOH18, DCZ19, DMHZ22, DVO^{+}23, DGMB21, DGMB23, DWY21, DTC17, DHP14, EJ16, FPTJ17, FPTJ18, FFP^{+}22, GFS22, GMD24, HWT^{+}17, HHZ^{+}21, HNF20, HKPR21, HAS21, JGM23, JC17, KM21, KMM^{+}18, KBV16, KBV17, KK23, KNH^{+}18, KSU^{+}19, KMG17, KXS^{+}19, KCD20, KBI^{+}21, KGBPM24, LLR^{+}16, LGTG16, LSNC21, LCL23a, LFWZ23, MGB18, MAZ21, MTNR21, MFH21, MSS17, MJ18, MAZS22, MBM15, MAB^{+}16, MMM23, MS23, MSB18, MZX^{+}15, MDE^{+}19, MNFGO19, MØ20, MR20, NV14, NDV18, NHK16, OR16b, OFF^{+}17, OBP^{+}24, OIY^{+}14, PS19, PZW^{+}21, PLZL17, PSL17, PCG17, PMA^{+}21, RBR16, RVB^{+}17, SLZ^{+}21, SSKP16, SPM17a, SPM17b, SBG17, TOKS14, TOS14, TH14, TH16, TON^{+}22, TP23, TSKO20]. Sound [UKS19, VDM15, WYO23, WWS^{+}18, WDW^{+}23, WK24a, WAS16, WSFS19, YLPL19, YLL21, YYW^{+}23, ZA14, ZPD19, ZSC^{+}23b, rPH^{+}17, Ano16-52, Ano16-53, Ano16-54, Ano16-55, Ano16-56, Ano16-57]. Sound-Based [TON^{+}22]. Sound-Class [DVO^{+}23]. Sound-Field [KMM^{+}18]. Sound-Quality [BvWM18]. SoundBeam [DVO^{+}23]. Soundfield [DRK18, ECH23, FSA20, JK15, MAST15, TZA14]. Sounds [CDA^{+}14, NK21, SCY22, TMKO17, ZH19]. Soundscape [DR22]. SoundStream [ZLO^{+}22]. Soundtracks [PWS^{+}23]. Source [AV16, AB21, APH14, AM18a, AJLW14, ATP16, AB14, AB15, BAS20, BATS21, BG23, CBA^{+}15, CONVVC18, CW18, CZY^{+}18, CSY^{+}21, CS19, CY16, CK19, ÇOH18, ÇH21, DCZ19, DMHZ22, DCAS22, DGMB21, DDMvW20b, DDNN19, EDO17, FSA18, FSA20, FSO17, FPTJ17, FPTJ18, FK19, FK24, GVMGO17, GJYW23, GH23b, GvdP23, GRSP17, HNT^{+}23, HAS21, HKHJS15, IBN^{+}18, IISN21, JSYW16, JVR16, JWBR21, KK23, KOS^{+}16, KNH^{+}18, KMG15, KCD20, KGBPM24, LKPR18, LGTG16, LGTG17, LBR16, LBR18, LLZ^{+}21b, LTE^{+}21, LH23, LKM23, LMH23, LSZ23, LJWC23, LLL^{+}22, LHM21, LLL^{+}23, LY23, MGB18, MBD18, MV19, MMT^{+}19, MAZS22, MYK^{+}24, MTKS21, MTK^{+}20, NKS21, NSA21, NK22, NV14, NSNN20, NLV16, OOK17, OVN^{+}22, OIY^{+}14, ÖM24, PCB14, PZW^{+}21, MG18, RJCW17, RJW19, SOF15, SDF18, SIKN23, SFRK^{+}23, SGH17c, SGH17a, SP17]. Source [SSGL21, SLL^{+}16, SHG21, SLOVB15, SPM17a, SPM17b, SWCN19, SSS18, TH18, TS14, TWSS16, TE16, UNI^{+}24, VAH19, VHBS21, WWD14, WRC16, WHRC16, WC18b, WTY20a, WGLZ21, WYY22, WH22, WWX^{+}23, XWWW19, YG24, YLPL19, YLL21, YZCW21, YPH23, YZLY17, YWT23, ZH15, ZZL^{+}23, ZPP15]. Source-Filter [LLL^{+}23, WTY20a, YPH23, YWT23]. Source-Specific [SLL^{+}16]. Sources [AKB^{+}16, BOL^{+}18, BH19, CATS20, DHSG15, DGMB23, DG15, DTC17, JC17, KNH^{+}17, LWZ^{+}14, MMB17, MAB^{+}16, MGGK17, NDV18, NHK16, OVN21, OIY^{+}14, PHK16, TH18, WCCB19, WAS16, ZZZ^{+}21]. Space [BDL15, CPGL24, CL14a, FLL^{+}23, FGQ^{+}15, HB17, KMAB18, LFH^{+}15, LM23b, LSLK24, MAST15, MAB^{+}16, MFG15, MBD^{+}15, PKP^{+}21, PMB^{+}20, PC19, SYG^{+}24, WZL^{+}15, WSH15, ZAZ^{+}18, ZYD^{+}23]. Spaces [LYZ21, TXH^{+}23, WBSG18]. Spacing [WCCB19]. Spam [XLG^{+}23]. Span [AGC23, JBC16, LSNC21]. Spans [ZZZ22a]. Sparse [AKB^{+}16, AGW15, ADM^{+}17, ADM^{+}19, BV17, BIK^{+}18, BHAMR19, BABC19, Chi18b, CDW16, CSAR^{+}14, ÇH21, CT14, DMHZ22, DCAS22, DBK15, DZLGO20, FK19, FWF17, GLWQ22, HL14, HOF15, JSYW16, JK15, JvWGD15, KKK15, Lib21, LG16, LAP23, MMS^{+}20, PKA14, PAB19, RAB18, RA23, SAS17, SZL19, Son21, SB18, SLY14, SLGZ15, VDC^{+}17, WAS14, WYY24, WYHL20, WVCL14, YGV14, ZF14, ZS20, ZGL^{+}23, ZKDS18, ZZL^{+}23]. Sparse-Echo-Path [ZGL^{+}23]. Sparsely [SMSV17]. Sparsity [AGBC14, Bay15, GJYW23, GaKGB21, JYB^{+}15, LGHG17, NZL23, ONKP16, TH18, XBZ^{+}24, ZZLZ19, ZTDD23a]. Sparsity-Based [GaKGB21, TH18]. Sparsity-Driven [ZTDD23a]. Sparsity-Promoting [NZL23]. Spatial [ATP16, ABKP19, BKTR24, CONVVC18, CH18, DMVP16, DGCA23, FLB16, FK24, HGT15a, HGT15b, HIA^{+}17, IO17, IBN^{+}18, KMG15, KBI^{+}21, LCL23a, LGHG17, LLS24, LAP23, MF17, MUT^{+}20, NV14, OGZ18, PACPJF^{+}23, PAB19, QL24, RKN14, SAC17, SIKN23, TZA14, TH14, TH16, TG21, TTH14, TQM^{+}23, TSKO20, VDH24, VDM15, WW19, WSFS19, YLPL19, ZHH18, ZHB^{+}18, ZGL^{+}20, ZZZ^{+}21]. Spatially [ATP15, BSTM24, FSEI21, LG20, NIK^{+}22, TG24, UNI^{+}24]. Spatially-Regularized [UNI^{+}24]. SpatialNet [QL24]. Spatio [ADM^{+}17, CNFG23, JCBJ15]. Spatio-Temporal [ADM^{+}17, CNFG23, JCBJ15]. Speaker [AS17b, ASX22, ADLX23, ALH^{+}21, BWZC22, BHP^{+}23, BOSR23, BJ24, BSW^{+}24, BSTM24, Bor22, Bor23, BS21, CLA^{+}21, CWL22, CBG15, CB21, CWZ^{+}23, CCY^{+}24, CH23, CL14b, CL17a, CL17b, CL18b, CL18c, Cum21, CS23, DMSM21, DAGS^{+}18, DAFS15, DXW^{+}23, DBLČ20, DQY18, DPHG18, DGCY23, FL19, FMB16, FNM^{+}19, GH17, GHAP17, GZG23, GGZ23, HCQ24, HH14, HMO21, HDLL23, HEMN21, HWW23a, HWW23b, HGAPB15, JG19, JKC^{+}24, JJB16, JKU18, KHKY22, KVMV^{+}24, KH16, KH21, KML18, KCMW17, LGTG18, LCLM18, LM15, LWZZ16, LMC17, LGHG17, LWX^{+}22, LWK^{+}22, LLHW24, LJH^{+}24, LCWC16, LM22, LM23a, LM23b, LH14, LW19, LCQ23, LSLK24, LSTK21, LCM18, LLL^{+}23, MZZ23, MPC16, MUK22, MZM15, MWC^{+}23, MMR^{+}22, MH18, MSD14, NTY19, NDS^{+}23, O’S24, PKP^{+}21, PBQ15, QWW^{+}23, QCL23, RSR^{+}19, RH18, SH14a, SAB16, STH^{+}18, STS21, SIUW21, SDK17, SHJT21, STT^{+}19]. Speaker [SG14, SHCV23, SCY22, SG21, SCW21, SKAK16, SAK16, TMM18, TMMZ18, TLWV17, TLD^{+}23, TCY15, TMH^{+}14, TM22, VMOL16, WW19, WHQY19, WZW19, WWL21, WCW^{+}21, WH22, WYZ^{+}23, WLT24, WDD^{+}16, WLL^{+}22, WK24b, XH23, XLLY18, XRWL21, XSC^{+}23, XCHX24, YLQZ24, YB14, YH17, ZG21, ZSB17, ZCY16, ZKH18, ZLD20, ZQZ^{+}23, ZWL^{+}23b, ZWW14, ZWW15, ZTL21, ZM23, ZLL^{+}24b, ZEM18, ZM20, VZC14]. Speaker-Adapted [DAGS^{+}18]. Speaker-Adaptive [HGAPB15]. Speaker-Aware [ADLX23, MZZ23]. Speaker-Dependent [NTY19]. Speaker-Independent [KCMW17, LCM18, LLL^{+}23]. Speaker-Normalization [SIUW21]. Speakers [HWG^{+}23, Lin18, Lin19, NR14a, ODNA23, SCEH19]. Speaking [CHC^{+}14, JBTL15, LLC^{+}24, LCWC16, TMH^{+}14]. SPEC [CSS23]. Special [Ano16-50, Ano16-51, Ano16-52, Ano16-53, Ano16-54, Ano16-55, Ano16-56, Ano16-57, KSG^{+}21, LFH^{+}15, RVB^{+}17, SHKB17]. Specialized [ACK21]. Specialty [ZJL21]. Species [JK19]. Specific [CB21, GvdP23, LWJ19, SW20, SLL^{+}16]. Specification [HCC^{+}23, YLJ16]. Spectra [AL23, NTY19, SPM17a, SPM17b, SLY14, SKH^{+}15a, YDZ19]. Spectral [AF18, Bay15, BS14, BKS^{+}18, DDNN19, FFSS18, GSA18, GKC16, HWW^{+}15, HL18, ISP19, JFG^{+}17, JY23, JRA17, KD18, KTJ19, KBG16b, LKHU14, LS14a, LGGH19a, LLZ^{+}21a, MGB18, MD18, MT18, MAT15, MKMDM17, NWM16, NCB23, PS17, PK15, SH14a, SSK17, SMO^{+}21, SY15, SKH^{+}15b, TW20, TM22, VRN18, WLL^{+}15a, WGD15, WW19, WWW21, ZKDS18]. Spectral-Amplitude [KTJ19]. Spectro [ECJF21, MD16, NXCL16, SGA15, vHK14]. Spectro-Temporal [ECJF21, MD16, NXCL16, SGA15, vHK14]. Spectrogram [KSK^{+}15, KMG17, PSL17]. Spectrograms [AS15, BIK^{+}18, BV16, CAFFR23, DSMD15, HFL^{+}23, MIYO14, TOKS14, TOS14]. Spectrospatial [TWW22]. Spectrotemporal [ATS18]. Spectrum [HGT15a, HHC16, LLS24, MPAEBLB15, SZL19, SKH^{+}15a, TTB^{+}16, VM15, XNRG15, XNP^{+}18, YLPL19, YWDQ21, YLQZ24, YMM19, ZTM^{+}23]. Spectrum-Based [XNRG15]. Speech [AHMJ^{+}14, AZK15, Abd18, AFG^{+}23, AKG^{+}17, AF18, AF19, AGW15, AJB^{+}18, AVM19, AC24, AS17b, AdHTJ16, AdTJ18, AM18b, AY15, AY17, ALH^{+}21, Ano14-45, Ano14-77, Ano14-54, Ano14-66, Ano14-78, Ano14-55, Ano14-67, Ano14-79, Ano14-56, Ano14-68, Ano14-80, Ano14-57, Ano14-69, Ano14-46, Ano14-58, Ano14-70, Ano14-47, Ano14-59, Ano14-71, Ano14-48, Ano14-60, Ano14-72, Ano14-49, Ano14-61, Ano14-73, Ano14-50, Ano14-62, Ano14-74, Ano14-51, Ano14-63, Ano14-75, Ano14-52, Ano14-64, Ano14-76, Ano14-53, Ano14-65, Ano14a, Ano15-40, Ano15-50, Ano15-41, Ano15-51, Ano15-42, Ano15-52, Ano15-43, Ano15-53, Ano15-44, Ano15-54, Ano15-45, Ano15-55, Ano15-46, Ano15-56, Ano15-47, Ano15-57, Ano15-48, Ano15-58, Ano15-49, Ano15-59, Ano16k, Ano16l, Ano16-32, Ano16-38, Ano16-33, Ano16-39, Ano16-29, Ano16-30, Ano16-31, Ano16-34, Ano16-36]. Speech [Ano16-40, Ano16-35, Ano16-37, Ano16-41, Ano16i, Ano16j, Ano17-59, Ano17-60, Ano17r, Ano17z, Ano17s, Ano17-27, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17-57, Ano18v, Ano18w, Ano18x, Ano18y, Ano18z, Ano18-27, Ano18-28, Ano18-29, Ano18-30, Ano18-31, Ano18a, Ano19a, Ano19t, Ano21a, AS15, ACRS14, AGBC14, ACB17, AZGN14, AM20, BK24a, BVGV15, BF18, BYT^{+}21a, BYT^{+}21b, BIK^{+}18, BXY14, BA19, BHP^{+}23, BV23, BNvdB^{+}20, BHS^{+}16, BE17, BKTR24, BPG23, BNA24, BSR^{+}23, BD16, BS21, CE15, CSS^{+}19, CLAG16, CDLL19, CDLL21, CBN^{+}24, CWW14, CHC^{+}14, CHWL14, CRVB^{+}15, CM15, CH15, CWH^{+}16, CLW^{+}16, CMY16, CDLX18, CLW^{+}19, CXX^{+}23, CLL^{+}21b, Chi18a, CMG^{+}17, CP21, CWBv19, CLSyL21, CWT22, CTLL17, CH14, DHD^{+}19, DP16, DLG18, DW17, DBK15, DXZ^{+}18, DXW^{+}23, DTL15, DST^{+}19, DDMvW20a, DB15]. Speech [DLL^{+}21, DB18, DB19, DTN^{+}17, DSN18, DTN^{+}14, DBN^{+}17, DRK18, DH14, DTDL16, DY22, DGCY23, DZF^{+}23, DSH18, DG23, ECJF21, EMTF18, ESVK17, ET18, EZS14, EMXD20, FYT^{+}21, FCCA23, FBWG23, FD21, FCG24, FSN24, FSEI21, GN23, GFG^{+}14, GVMGO17, GDC19, GBL24, GTV19, GTSM24, GG18, GI15, GBH16, GH23a, GCJ^{+}14, GvdP23, GWL^{+}24, GCG^{+}17, GAM23, GJB15, GJ17, GZZY23, GWSC23, HW14, HWW^{+}15, HJZB23, HBB17, HLL15, HCJT15, HS15, HCH23, HYI^{+}15, HdHTJ22, HKPR21, HKI23, HZW^{+}22, HZC^{+}22, HBT^{+}21, HSWC21, cHrLLyL23, HL18, HXL^{+}21, HSLB23, HCZC24, HWC^{+}17, HBCC22, HM14, HHC16, ISP19, IK23, ISZA15, JP14, JFG^{+}17, JD17, JVR16, JT14, JT16, JLKK21, JKC^{+}24, JWLF14, JPX^{+}23, JY23, JGD^{+}24, JRA17, JGSC^{+}22, JvWGD15, JBTA19, KGP14, KSL19, KVMV^{+}24, KNBC19, KML18, KK14, KKK15, KS16, KCMW17, KLCL24, KYTJ17, KTJ19, KK19b]. Speech [KGH^{+}16, KS21, KH24, KPA^{+}24, KHE21, KKGH16, KBG16a, KBG16b, KBG18, KG14, KSG24, KDJJ16, KDJJ17, LG19, LLGG20, LG20, LM17, sLGyLaC15, LC16a, LK19, LKC^{+}21, LLC22, LFO^{+}24, LYWX22, LZC^{+}23, LRWG23, LKHU14, LWZ^{+}14, LDGHU14, LS14a, LWZZ16, LJK17, LSDM19, LGGH19b, LWM^{+}20, LWL^{+}21, LY21, LLZ^{+}21a, LWX^{+}22, LCQ22, LDG^{+}23, LLW23, LHC^{+}23, LYZ^{+}23, LSZ23, LJWC23, LSH^{+}23, LSWN23, LWLS24, LYY24, LLWQ24, Lin18, LdLvWWS21, LYH^{+}22, LWYD23, LB23, LWY^{+}23, LAGD18, LS14b, LCC^{+}15, LK16, LLLL19, LSB^{+}21, LPI^{+}21, LGH^{+}21, LLyL21, LRP^{+}23, LWS^{+}23, LSG^{+}23b, LKR23, LLdCF23, LLC23a, LLL24, LHZ^{+}24, LB19, LGR14, LR17, LDZ19, LLZ^{+}23c, LCY^{+}23, LSH14, LCM18, LM19, LFM^{+}23, MWT16, MTNR21, MPAEBLB15, MJZ15, MK15a, MCL22, MLMT24, MGGK17, MP20, MOT^{+}23, May18, MCZ21, MHD^{+}17, MZC^{+}24, MTZ^{+}21, MC17, MH16, MSD14, MD16, MAT15, MDM15, MAK15, MKA16]. Speech [MTS^{+}19, MK15b, MK15c, MBK17, NTY19, NW14, NW15, NSA21, NW24, NEN21, NXCL16, NLL^{+}18, NJC16a, NJC16b, Nos16, NAK^{+}17, OHN16, OWA15, OWW^{+}23, PW19, PW21, PW22, PW23, PSL^{+}16, PKR^{+}24, PC19, PJTJ24, PK15, PZP19, PHM^{+}16, PHS^{+}24, PM19, PM23, PDKG24, QDSL19, QYCT23, QTY16, QBTY16, QZ22, QLW^{+}24, QL24, RGR^{+}22, RPK23, RKL14, RWF16, RKN14, RA23, RG18, RG21, RWL^{+}23, RGK24, SML^{+}24, SV17, SP16, SWW^{+}17, STS18, SSFP18, SF18, SHJY21, SDDCC15, SAA15, SM14, SGH15b, SGH15a, SGH16, SBN^{+}19, SKSD14, SSFL22, STBC17, SR14, SSSS22, SGY^{+}24, SAS17, SFIL17, SZS22, SBM^{+}19, SH17b, SSK17, SKDN14, SKM17, SM18, SLW15, SLWY20, SRC^{+}23, SLGZ15, SZVZ16, SRT23, SYG^{+}24, SK18, TM14, TTB^{+}16, TD23, TQH^{+}18, TCW19, TW20, TZW21, TW21, TLWV17, TB18, TH17]. Speech [TJCB16, TG21, TG23, TG24, TE23, TYW^{+}23, TQM^{+}23, TP19, TKT17, TK14, TR14a, TZL^{+}23, TDL19, TE16, VBY18, VDH24, VM15, VP17, VAP19, VSL21, WFN^{+}18, WWD14, WS14, WNW14, WCC^{+}15, WLL^{+}16, WW16, WDDL17, WTY18, WB18, WDDL18, WC18a, WLT^{+}18, WTW20, WTY20a, WTY^{+}20b, WWL21, WWW21, WCW^{+}21, WW21, WW22, WWG^{+}23, WCZR23, WZG^{+}23, WWWR23, WCC^{+}23, WYY24, WTHJ14, WYSD15, WJ15b, WWW16, WW17, WRDP17, WSM19, WHLG14, WZS14, WK16, WCL^{+}21a, WCL^{+}21b, XHRC24, XWWW18, XLL^{+}19, XDH^{+}17, XGKM18, XGKM19, XDDL15, XDC^{+}17, XYH^{+}21, XRWL21, XZF23, XYY^{+}23, XAHJ^{+}14, XMBN18, XMBN21, YBC^{+}14, YK14, YC22, YPC^{+}22, YB14, YLL15, YTWB19, YGV14, YZLY17, YLK^{+}21, YHO^{+}17, YHO^{+}18, YH21, YLLC19, YZW^{+}21, YCL22, ZCF14, ZG21, ZH14, ZFC^{+}23, ZF14, ZCDS15, ZW16b, ZG16, ZZN^{+}16, ZW17, ZS20, ZL21b, ZXY^{+}21, ZWZ^{+}21, ZCDH21, ZZZL21, ZTDD23a, ZWW23a]. Speech [ZLL^{+}23, ZSC^{+}23a, ZQZ^{+}23, ZWC^{+}23, ZZS^{+}23, ZTDD23b, ZQN^{+}23, ZLF19, ZWW19, ZDGO21, ZSL^{+}22, ZZ19, ZAL24, ZJD^{+}15, ZTL21, ZLD21, ZLG^{+}22, ZZZ^{+}24, ZGC^{+}23, ZTM^{+}23, ZZZD23, ZCW^{+}24, ZLL^{+}24b, ZSIA16, vNKB^{+}23, VZC14, BB21]. Speech-Aware [VDH24]. Speech-Distortion [AM18b]. Speech-in-Noise [SZS22, ZSIA16]. Speech-Music [BPG23]. Speech-to-Speech [DTN^{+}17]. SpeechFormer [CXX^{+}23]. Spell [NNC21]. Sphere [BKK15]. Spherical [AB21, AR16, ATP16, CNFG23, CM18, DMVP16, EJ16, FSA18, FD21, Hac14, HMN17, HAGA21, HZF17, HZF18, JTHN14, JSR16, JEP14, JC17, LAP23, MMS^{+}20, MTKS21, NR14a, NEWN23, SSTS18, SAC17, SR14, SvWRK15, TP15, Ter16, TJV24, VAH19, ZKA15]. Spherical-Harmonic-Domain [MMS^{+}20]. Spiking [CLAG16, PZW^{+}21, XWY^{+}23]. Split [LY23]. Spoken [Ano16-50, Ano16-51, CTLL17, CL18a, FLMS16, HS15, KG17, KTH^{+}20, KCL21, yLsL14, yLSY^{+}14, sLGyLaC15, LCW^{+}18, LLW^{+}19, LWLL24, MAP14, MWT^{+}23, MDY^{+}15, QCN^{+}21, RAB18, RV17, SWH^{+}17, SHKB17, SLK17, WQZ^{+}21, WvdHZ^{+}21, WDW^{+}22, WWLL14, WZC23]. Spontaneous [HWC^{+}17, WJ15b, WHLG14]. Spoof [LSLK24]. Spoofed [KYL24, LWS^{+}23]. Spoofing [DLL^{+}21, DQY18, GAPGG19, KHKY22, QCDW17, RPL^{+}23, WDD^{+}16, YDZ19, YWDQ21, ZLL^{+}23]. Spot} [LCL23a]. Spotforming [TH16]. Spotting [LYY^{+}23, LETJ21, WTHJ14, WJ15b]. Spread [HGT15c, XNRG15, XNP^{+}18]. SQL [WLZH21]. SQuAD [LLW^{+}19]. Square [DDMvW20b, JT15, KTJ19, ZZH16, ZLX^{+}24]. Squared [HLZ^{+}23]. Squares [BBM17, EIPB^{+}19, KJ21, LH23, LZHN14, SOF15, SD16b, SH19, ZGL^{+}23]. SR [LCWC16]. SR-HPM [LCWC16]. SRMR [SF14]. SRP [DGMB21]. SRP-PHAT [DGMB21]. SRRank [YW14]. SSAP [LHL^{+}22a]. SSD [NDS^{+}23]. SSVD [ZZX^{+}23a]. SSVS [ZZX^{+}23a]. SSVS-SSVD [ZZX^{+}23a]. Stable [GCJ^{+}14]. Stably [HKHP23]. Stacked [DZWC20, WK16]. Stacking [ZZN^{+}16]. Stacks [BWH18]. Stage [ABB^{+}23, GTS^{+}19, GRSP17, GXW^{+}23, HSPK18, HZF17, LLZ^{+}21a, LYY24, LdLvWWS21, LFWZ23, LLL^{+}22, SGM^{+}15, SK16b, SWCN19, WDW^{+}23, XHRC24, ZWW19, TOS14]. Stance [WSL^{+}20]. Standard [GLW20]. Start [HSPK18]. Startle [DMP^{+}18]. State [BM21, JHLK18, JGHN23, LWLL24, LB16, LCL^{+}23b, MMBT21, MSRND14, PC19, SYG^{+}24, WHLG14, WLL15b, YSCZ15, ZYL^{+}21, ZJD^{+}15]. State-Based [WHLG14]. State-Clustering [ZJD^{+}15]. State-Space [PC19, SYG^{+}24]. Stationary [BP14, DBK15, LSH14, MPAEBLB15, NJC16a]. Statistical [AJB^{+}18, AKN^{+}16, BY15, Bre15, CRVB^{+}15, CZY^{+}18, GJB15, HL18, HSZ^{+}18, JBTA19, KK19a, KK19b, KBG16a, MHPB15, MSD14, NNC19, NAK^{+}17, RLK14, STS18, SYKL21, TTB^{+}16, TKT17, TZZ15b, WZL^{+}15, WTY18, WTZ18, WTY20a, WTY^{+}20b, WK24b, XZW15, YEY17, ZLLS16, ZDLK18]. Statistically [BNA24, EZS14]. Statistics [BV23, JSMD18, MKMDM17, NJC16a, OL15, TS14]. Status [LXK^{+}21, XXWY24]. STD [LCZ^{+}19]. Steerable [BBAS19, CMM^{+}23, WC16a, ZHCB21, ZZL^{+}23]. Steered [ÇOH18, TWSS16, ZDD21]. Steering [CSY^{+}21, CP21, JHW^{+}21, SNTD19]. Steganalysis [LJK17, LWLS24, PLQZ23, TQM^{+}23]. Steganography [AGW15, DFY^{+}24, LJK17, LWL^{+}21, LWLS24]. Step [AA17, CC14a, CGL^{+}24, CC14b, HZF18, KKJN17, TM16, ZZH16]. Step-by-Step [CGL^{+}24]. Step-Size [AA17, CC14a, TM16, ZZH16]. Step-Sizes [KKJN17]. Stepwise [CGL^{+}24, TL19, YPC^{+}22]. Stepwise-Refining [YPC^{+}22]. Stereo [AJLW14, CHJ14, HTG14, LP22, NCC15, ZZX^{+}23a]. Stereo-Input [CHJ14]. Stereophonic [HJC19]. Stereophony [HF23]. STFF [TQM^{+}23]. STFF-SM [TQM^{+}23]. STFT [KG14, MBD18, PBS17, WWWR23]. STFT-Domain [WWWR23]. Stimulation [LWL^{+}19]. Stochastic [LRWG23, NCB23, WFY^{+}22, WLL15b]. STOI [SW20]. Stops [MP20, VAP19]. Storable [GGZ23]. StoRM [LRWG23]. Story [CLL^{+}21a, LHL^{+}22a, PACPJF^{+}23, ZHZ19]. Storylines [LHL^{+}22a]. STRAIGHT [AJB^{+}18]. Strain [CG21]. Strategies [GH23b, HHMK18, TM21]. Strategy [KM21, LSG^{+}23b, LETJ21, MX17, MLP^{+}24, MLL^{+}22, MTS^{+}19]. Stream [AZK15, LWM^{+}20, LYB^{+}23]. Streaming [DHP14, IK23, JGSC^{+}22]. Streamlined [PT14]. Streams [DSH18, TQM^{+}23, XO16]. Strengths [XNP^{+}18]. String [MSS17]. Strings [SAK16]. Strong [MMM23, OWA15]. Structural [FWRJ21, SBV17, WCL^{+}23, WLF23, ZKDS18]. Structure [JSG^{+}22, KK23, KHCW15, MIR^{+}21, TP16, WWX^{+}24, WXL^{+}21, WXS24, ZZZ16, ZLJ^{+}17, ŞUY^{+}18]. Structure-Aware [WXS24]. Structured [AGBC14, CCT^{+}19, CL18a, CLZ22, DZLGO20, KWGW17, KKK15, OS21, PZZZ22, PPSG17, RCL18, WYLZ20, ZKDS18]. Structuring [yLSY^{+}14]. Student [DWX^{+}21, ISZA15, LSL24, TDL19, CS19, LBR18]. Studio [ES16]. Study [BSR^{+}23, CWW14, CGL^{+}24, CFMM14, FLMS16, GSA18, HF15, JHW^{+}21, KBV17, LHC^{+}23, MBDV23, PCB14, PKR^{+}24, SKSD14, ZCDH21, ZLW^{+}22]. Stuttered [KHE21]. Stuttering [BHH22]. Style [ASX22, GKC16, GHJ^{+}23, HCC^{+}23, LZC^{+}23, LLC^{+}24, LSGL21, LHYJ23, NLL^{+}18, SWC^{+}21, WWX^{+}23, WWLL14, WY23, ZPD19, ZL21b, VZC14, SWC^{+}21]. Style-Aware [SWC^{+}21]. Sub [CL14a, MTK^{+}20, RARR15, WCC^{+}23, ZLSC23]. Sub-Algorithm [ZLSC23]. Sub-Band [WCC^{+}23]. Sub-Gaussian [MTK^{+}20]. Sub-Set [RARR15]. Sub-Space [CL14a]. Subband [CD17, FZZ^{+}15, GLQ16, GYdL^{+}21, KKJN17, KGH^{+}16, Lin19, PZZZ22, PPSG17, SM14, YPL21, YHC^{+}20, ZGZ23]. Subbands [MBM15]. Subglottal [MVH16]. Subgroup [KPA^{+}24]. Subharmonic [IIY16]. Subject [Ano17a, RF19, Wid18, Wid19]. Subjected [TM14]. Subjective [BNvdB^{+}20, BvWM18, Bre15, FBE19, GTV19, GTSM24, PCLP14, WHB^{+}21]. Subsequence [ZLZW21]. Subset [ZCHH18]. Subsources [FK24]. Subspace [DGCA23, Emu20, GR14, LM15, LSZ23, LGR14, MLP^{+}24, ONKP16, RAB18, SK18, XDC^{+}17, YZLY17, ZS20]. Subtraction [KMG15, SK18]. Subtree [SKK14]. Subword [FL19, HBF^{+}16, HSN18, SSI21, ZZL^{+}19]. Subword-Based [SSI21]. Successive [CG20]. Suitable [FCG24]. Sum [CPSFC^{+}14, ZH14]. Summarization [AGC23, BZ23, CLC^{+}15, CLCW18, CSS23, CGL^{+}24, GLC^{+}14, LPM^{+}23, LZX^{+}23, LCC^{+}15, MSW^{+}22, RRdM16, WYH^{+}15, YW14, ZZZ16, ZZLZ19, ZYW^{+}20]. Summarizing [BSL16, HNY^{+}15, HC19, WZM21]. Summary [CSS23, CGL^{+}24, LCYW21]. Summation [IIY16]. Super [CY15, KVMV^{+}24, KLCL24, McL14, RG18, WW21]. Super-Audible [McL14]. Super-Gaussian [CY15, RG18]. Super-Resolution [KVMV^{+}24, KLCL24, WW21]. Superconductivity [Ano17a]. Superdirective [BCB15, CBHC21, CT14, PCB16a]. Superdirectivity [LMGG21]. Superfast [PT21]. Supergaussian [RBBG15]. Superpositional [TG16]. Supertags [ODSM16]. Supervised [CWL22, DLH23, DWY21, DZF^{+}23, FLM23, FZM24, GFG^{+}14, GWQ23, GMD24, HCQ24, HBT^{+}21, Kam23, KK23, LGTG16, LGTG17, LC24, LAK16, LHL^{+}23a, LSL24, LLyL21, LCL^{+}23b, LCTA19, MSW^{+}22, MS23, PEO^{+}20, PLQZ23, Rak17, RV17, SBN^{+}19, SK16b, SG21, SCEH19, SHL^{+}21, TLD^{+}23, WNW14, WsL15, WC18a, WXZ^{+}19, WK24a, WLL^{+}22, XCHX24, YLL15, ZGC^{+}23, ZZZD23, ZB14]. Supervision [BJ24, DJY^{+}19, XSC^{+}23, ZCW^{+}24]. Supervisions [HZL23]. Support [CL14b]. Supported [EMTF18]. Supporting [ACK21]. Suppression [BB21, DDBW20, DDBW23, HdHTJ22, ICB24, TZZ^{+}15a, WLT^{+}18, XWWW18, ZZL^{+}23]. Suprasegmental [SP17]. Surface [MVH16, MYC^{+}23]. Surnames [ZP15].

Surpassing [LM19]. SURT [RPK23]. Survey [PHS^{+}24, SWH^{+}17, XXWY24]. Sustained [LD16]. SVD [ZCQ^{+}19]. SVD-Based [ZCQ^{+}19]. SVS [KKJK23]. Swarm [WRKG17]. Sweet [LCL23a]. SWF [ZTDD23b]. Switched [YsL15, ZGL^{+}23]. Switching [AVK^{+}15, ISP19, NIK^{+}22, WLLZ17, WSY14, WSH15]. Syllable [CGL^{+}15, WHLG14]. Syllable-Context [CGL^{+}15]. Symbolic [OL15, WLF23]. Synchronization [CG17, XNG^{+}14]. Synchronized [KS21]. Synchronous [SM18, VAP19, YCM^{+}21]. Synchrosqueezing [MCS21]. Syntactic [AVK^{+}15, BLZ21, WZLL24, WFR^{+}21, XWZ16, ZZY^{+}24, ZZH21, ZCY^{+}22]. Syntax [DZZ23, ZZZ22a]. Syntax-Aware [DZZ23, ZZZ22a]. Synthesis [AF19, AP21, AJB^{+}18, AG17, ADGE19, BK24a, BRD15, BRC24, BGQO^{+}17, BABC19, CHWL14, CRVB^{+}15, CWH^{+}16, CDA^{+}14, DHCS15, DLG18, DY22, ECH23, FFSS17, FFSS18, GBL24, GDM18, HHO^{+}21, cHrLLyL23, HL18, HWC^{+}17, JY23, JBTA19, KK19b, LLC22, LYWX22, LZC^{+}23, LWX^{+}22, LSB^{+}21, LLC23a, MSS17, MZC^{+}24, MSD14, MDM15, Nos16, NAK^{+}17, OR16b, PM19, SML^{+}24, STS18, SSK17, SLH17, SGB22, TTB^{+}16, VP21, WTY18, WTY20a, WTY^{+}20b, WAS16, WWH^{+}18, WSFS19, WHLG14, WK16, WCL^{+}21a, YHO^{+}17, YHO^{+}18, ZL21b, ZZZL21, ZQZ^{+}23, ZZS^{+}23, ZLD21, ZZZ^{+}24]. Synthesisers [PM23]. Synthesized [YGS19]. Synthesizer [CMG^{+}17, GBL24, KKJK23, MS23]. Synthesizing [RHvdP^{+}16, WvdHZ^{+}21, YLLC19]. Synthetic [ARSLN16, EZS14, LWL^{+}19]. System [BB21, DG23, GA20, HHO^{+}21, JZJØ19, KAA^{+}24, KBV17, KSG^{+}21, KGG^{+}21, KG17, yLSY^{+}14, LAK16, LG16, LGH^{+}21, LGL^{+}24, MTS^{+}19, MGT15, NK22, ODG16, OWW^{+}23, PS17, PBC18, RBR16, SGM^{+}15, SP17, SWC20, SBVM16, WYHL20, YCC^{+}24, ZLSC23]. Systematic [BHH22, YLJ16]. Systems [AA17, ATR^{+}21, Ano19c, AZGN14, BM21, BLLP17, CE14, CHL^{+}24, Chi18a, CB14, DXW^{+}23, EIPB^{+}19, GAM23, HHMK18, KHKY22, KG19, LK19, LCL^{+}21, LC16b, LCL^{+}23b, LSYN19, MWL22, MSFK15, MMMDL^{+}14, MMR^{+}22, MSD14, MVN22, PKR^{+}24, RLK14, SSK17, TDDN24, WBSG18, WLL15b, XWZ16, YLJ16, YLQZ24].

T [LK19]. T60 [LLW23]. Tablature [MFG15]. Table [Ano14-92, Ano14-93, Ano14-94, Ano14-95, Ano14-96, Ano14-97, Ano14-98, Ano14-99, Ano14-100, Ano14-101, Ano14-102, Ano14-103, Ano14-104, Ano14-105, Ano14-106, Ano14-107, Ano14-108, Ano14-109, Ano14-110, Ano14-111, Ano14-112, Ano14-113, Ano14-114, Ano15-75, Ano15-76, Ano15-77, Ano15-78, Ano15-79, Ano15-80, Ano15-81, Ano15-82, Ano15-83, Ano15-84, Ano15-85, Ano15-86, Ano15-87, Ano15-88, Ano15-89, Ano15-90, Ano15-91, Ano15-92, Ano15-93, Ano16-58, Ano16-59, Ano16-60, Ano16-61, Ano16-62, Ano16-63, Ano16-64, Ano16-65, Ano16-66, Ano16-67, Ano16-68, Ano16-69, Ano16-70, Ano16-71, Ano16-72, Ano16-73, Ano16-74, Ano16-75, Ano17-99, Ano17-78, Ano17-79, Ano17-80, Ano17-81, Ano17-82, Ano17-83, Ano17-84, Ano17-85, Ano17-86, Ano17-87, Ano17-88, Ano17-89, Ano17-90, Ano17-91, Ano17-92, Ano17-93, Ano17-94, Ano17-95, Ano17-96, Ano17-97, Ano17-98, Ano18-80, Ano18-81, Ano18-82, Ano18-83, Ano18-84]. Table [Ano18-85, Ano18-86, Ano18-87, Ano18-88, Ano18-89, Ano18-90, Ano18-69, Ano18-70, Ano18-71, Ano18-72, Ano18-73, Ano18-74, Ano18-75, Ano18-76, Ano18-77, Ano18-78, Ano18-79, Ano19-31, Ano19-32, Ano19u, Ano19v, Ano19-33, Ano19w, Ano19-34, Ano19x, Ano19-35, Ano19y, Ano19z, Ano19-27, Ano19-28, Ano19-36, Ano19-29, Ano19-30, Ano21c, Ano21d]. Tables [BTD^{+}19]. Tackling [PPM^{+}24, PWS^{+}23, WSZN21]. Tag [FYG^{+}21, LJL^{+}23]. Tag/ [LJL^{+}23]. Tagging [AGM23, FYG^{+}21, GCG21, KEzW21, LZC^{+}14, LCZ^{+}17, LJL^{+}23, XHW^{+}17, YZL^{+}18, ZYF18]. Tags [LJL^{+}23]. Tailed [DCYY22, XXS^{+}23]. Tailoring [ZZY19]. Taking [SWZ16]. Talk [GCvWMJ14, ICB24]. Talker [LW19, RPK23, WYSD15, WRKG17, XRWL21, YLL15]. Talker-Independent [LW19]. Tandem [KHKY22, OOK17]. Tangent [XCZ^{+}23]. Tangential [HF15]. Tap [CC14a]. Tap-Length [CC14a]. Tapped [HL14]. TARGAT [XZL^{+}23]. Target [ABKP19, CJY^{+}21, DVO^{+}23, DDZL19, FZZ^{+}15, FZM24, GTV19, KNH^{+}17, LCSL19, LLHW24, MGB18, MYK^{+}24, PCBS19, WLT24, XRWL21, XCHX24, YKL^{+}23, ZXY^{+}21, ZWX15]. Target-Dependent [DDZL19]. Target-Oriented [CJY^{+}21]. Targeted [BLZ21, ZZY^{+}24]. Targets [BSR17, WNW14]. Task [AMK^{+}20, BPG23, CWR^{+}22, FYG^{+}21, FOBP15, GLT23, JLP^{+}21, JQL24, LCL^{+}21, LCRY24, LCL^{+}23b, MMBT21, QXW^{+}23, SGY^{+}24, TDDN24, WYH^{+}15, WYL^{+}23, YM17, YTF^{+}23, YJX19, ZYW^{+}24]. Task-Adaptive [CWR^{+}22]. Task-Oriented [LCL^{+}23b, MMBT21, QXW^{+}23, TDDN24, WYH^{+}15, ZYW^{+}24]. Tasks [DB15, LSL24, MLL^{+}22, PM15, RRdM16, WWX^{+}24]. TasNet [LM19]. TAU [JLKK21]. TAU-Net [JLKK21]. Taylor [LYZ^{+}23]. TBNet [LKR23]. TDAC [CRH^{+}14]. TDFNet [ZWS^{+}23]. TDNNs [SMA^{+}17]. TDOA [DMHZ22, LGTG18, SSS18]. TDOA-Based [SSS18, DMHZ22]. Teacher [DWX^{+}21, JQL24, LSL24, LWL^{+}22, TDL19, YKL^{+}23]. Teacher-Student [DWX^{+}21, LSL24].

Teaching [CLC^{+}22]. Teager [ISZA15, KPH15, SH17b]. Technical [DRFO14]. Technique [AFdDG23, CC14b, ODG16, SLY14, WBLC22, ZZL^{+}23]. Techniques [CLC^{+}15, HWW16, HHW^{+}21, KEzW21, ÖM24, RG15b, SSK17]. Technology [KSG^{+}21, KGG^{+}21, YCC^{+}24, Ano14-115, Ano14-116]. Telephone [JMD^{+}17, JRA17, LWYC23, PMB^{+}20, SKSD14, VP17]. Telephony [KVMV^{+}24]. Tempo [GKC16, PT14]. Temporal [AY17, ADM^{+}17, BKTR24, BSTM21, CNFG23, CCH^{+}21, ECJF21, GMD24, HS15, JGM23, JNP^{+}21, JCBJ15, JLKK21, KM21, LdLvWWS21, LB16, MXH21, MD16, NK21, NTA15, NXCL16, PACPJF^{+}23, SGA15, SY15, SWZ16, TWC^{+}18, TQM^{+}23, VRN18, WJ23, XZL^{+}23, YK17, ZSH20, vHK14]. Temporally [LS14b]. TEND [DDZL19]. Tensor [BV16, BOL^{+}18, MH16, MTKS21, QYCT23]. Tensor-Train [QYCT23]. Tenth [YCC^{+}24]. TERA [LLyL21]. Term [CTLL17, CL18a, LXK^{+}21, LLL^{+}19, MAP14, MMCF18, MKMDM17, NR14b, PDS^{+}16, RAB18, YB14, YZLY17, ZLJ^{+}17]. Terms [WFR^{+}21, YJX19]. TESA [MUK22]. Test [LLW^{+}19, LWK^{+}22, NR14a]. Tests [Bre15]. Text [AGB^{+}24, ALH^{+}21, BYT^{+}21b, BTD^{+}19, BZ23, CMY16, CSZZ24, Chi18a, DYY^{+}23, DGCY23, FLM23, GWCT23, HW23, JKC^{+}24, sLGyLaC15, LLZ^{+}21c, LHC^{+}23, LM23a, LHYJ23, LHZ^{+}24, MYT^{+}19, MYL^{+}21, MZC^{+}24, MMR^{+}22, PLQZ23, PM19, SML^{+}24, STT^{+}19, SFDRB21, SWH22, SKAK16, SAK16, SLH17, VBY18, WLLZ17, WHQY19, WLZH21, WH22, WDD^{+}16, XH23, XXS^{+}23, YLC^{+}24, YLFL22, YYW^{+}23, YZW^{+}18, YLLC19, YXX^{+}22, ZP15, ZSB17, ZKH18, ZZL^{+}19, ZWHD23, ZMZ24, ZJL21, ZZZ^{+}24, ZM23, ZLL^{+}24b]. Text-Dependent [LM23a, MMR^{+}22, STT^{+}19, SKAK16, SAK16, ZSB17].

Text-Independent [CMY16, WHQY19, WH22, WDD^{+}16, XH23, ZKH18, ZM23]. Text-Inductive [SML^{+}24]. Text-Informed [SFDRB21]. Text-Only [BYT^{+}21b]. Text-to-Sound [YYW^{+}23]. Text-to-Speech [Chi18a, DGCY23, JKC^{+}24, LHC^{+}23, LHZ^{+}24, MZC^{+}24, PM19, VBY18, ZZZ^{+}24, ZLL^{+}24b]. Text-to-SQL [WLZH21]. Text-to-Text [HW23]. Texts [AAA19, BZ21, WZM21]. Textual [BM15, JGK^{+}22, LLCT22, SHZ^{+}23]. Texture [OR16b]. TF [WCC^{+}23, YLPL19]. TF-GridNet [WCC^{+}23]. TF-Wise [YLPL19]. Their [JSMD18, PHM^{+}16, SH14b, TR14a, TKH21]. Theme [PMB^{+}20]. Theorem [LYZ^{+}23]. Theoretic [FDK15, ZMZ24]. Theoretical [AZGN14, Hac14, HMDG15, HF15, MHGD15, PCB15, TR14b, TC22]. Theoretically [JT15]. Theory [JSR16, JK15, LSH^{+}23, QDSL19, WZS14]. Theory-Inspired [LSH^{+}23]. Thinking [ZZZL22]. Three [ADGE19, KBV17, UKS19, WCY19, WK24b, ZA14]. Three-Dimensional [ADGE19, KBV17, UKS19, WCY19]. Thresholding [ISZA15]. Throat [STH^{+}18, TE16]. Throat-Microphone [TE16]. Tied [CL18b]. Tilt [JRA17]. Timbre [ARW^{+}16, DGCY23, WYZ^{+}23, ZQZ^{+}23]. Timbre-Normalized [DGCY23]. Timbre-Reserved [WYZ^{+}23]. Time [APH14, AM16, APWB23, BP14, BRC24, BHBS16, CONVVC18, CB21, CTE21, FSN24, GEC15, GJYW23, GWLL22, GZZY23, HB17, HGT15b, HCJT15, HXL^{+}21, HXC^{+}22, HGT15c, IC19, JWBR21, KYL24, KVMV^{+}24, KB14, KTJ19, KXS^{+}19, KS21, KBGAP^{+}16, LG19, LZZP22, LM19, MAZ21, MTNR21, MAP14, MPKS^{+}23, MTK^{+}20, MTBB15, NNS16, NKS21, NV21, PZW^{+}21, PW19, PW21, RG15a, SK17, STT^{+}19, SNK21, SH17a, SH14b, SUK24, TZW21, TZZ^{+}15a, VH19, WZW19, WCC^{+}21, WW17, WZS14, WSG21, XWWW18, XZL^{+}23, XGKM19, YOM21, YK21a, ZQN^{+}23, ZSIA16, ZH15]. Time-Aligned [APWB23]. Time-Aware [XZL^{+}23]. Time-Contrastive [STT^{+}19]. Time-Delay [APH14, CB21]. Time-Domain [KYL24, KVMV^{+}24, MPKS^{+}23, NKS21, PW21, XWWW18, WZS14]. Time-Frequency [AM16, BP14, GJYW23, GWLL22, JWBR21, MTNR21, SUK24, TZZ^{+}15a, WZW19, WW17, YK21a, ZQN^{+}23]. Time-Frequency-Bin-Wise [YOM21]. Time-Offset [CTE21]. Time-Shifting [HGT15b]. Time-Spread [HGT15c]. Time-Variant [MTK^{+}20]. Time-Varying [BRC24, KBGAP^{+}16, LG19]. Times [KLS^{+}21]. Tissues [ZH19]. TOE [LJL^{+}23]. TOEFL [LLW^{+}19]. Toeplitz [PT21]. Together [Ano14-115, Ano14-116]. Tokens [CTLL17, CL18a, FZM24]. Tolerance [CCT21]. Tonal [DDZS17]. Tone [LCSL19, PZW^{+}21]. Tones [LCSL19]. Tongue [SLH17, ZAL24]. Topic [CWU^{+}19, DJY^{+}19, IO19, LJX^{+}16, LZX^{+}23, LCG^{+}24, SKT21, WTHJ14, WZH22, XZW15, XZFN23, ZM17]. Topic-Aware [LCG^{+}24, ZM17]. Topic-Based [XZW15]. Topic-Expression [WZH22]. Topic-Oriented [LZX^{+}23]. Topics [Chi18b, GLC^{+}14]. Topology [KSHH18]. Total [MBD^{+}15, SLL^{+}16]. Tracked [BBHAR24]. Tracker [BM21, JHLK18, JGHN23]. Tracking [AHSS^{+}18, BLLP17, BG23, CMD^{+}23, DLH23, DGMB21, Emu20, HAGSZ16, HW14, HEMN21, LGTG18, LWLL24, Lin19, MMBT21, NDV18, ODNA23, QWW^{+}23, SG14, TH18, TCY15, TP19, TS14, WRKG17, YSCZ15, ZDG14, ZYL^{+}21, ZCY16, ZH15, ZLSC23]. Tract [ADGE19, LTE^{+}21, SIUW21, SMH14, SMB18]. Tracts [AG17]. Trade [LSNC21]. Trade-Off [LSNC21]. Tradeoff [JTHN14, JSR16]. Traditional [DSMD15, NCB23]. Traffic [LWY^{+}23]. Train [QYCT23]. Trained [HSPK18, HXL^{+}21, HXC^{+}22, KCL21, LLW21, LHL^{+}22a, NTA15, NTO^{+}23, SHL^{+}21, WDW^{+}22, WL19, YCL22, ZF14, ZZS^{+}23]. Training [ACK21, Bor22, BG23, CLLD14, CLW^{+}16, CH16, CCY^{+}24, CZK^{+}21, CCL^{+}21, CL14b, DYY^{+}23, DWX^{+}21, DH14, DZF^{+}23, EMXD20, FYT^{+}21, FLYJ23, GTV19, HXWS15, KYTJ17, LAK16, LMW^{+}24, LSGL21, LHZ^{+}24, LETJ21, MLL^{+}22, May18, MZM15, MIR^{+}21, NTM16, NW15, NLYK24, PW23, QTY16, QMS16, QXW^{+}23, RG21, SV18, hShWsL15, TQY16, TLWV17, TLD^{+}23, VHBS21, WNW14, WW16, WJ23, WZLL24, WQ24a, WWX^{+}24, WJ15b, WK16, XHRC24, XWZ16, XYY^{+}23, ZZZL21, vNKB^{+}23, VZC14]. Training-Based [BG23]. Trait [SWZ16]. Traits [CPGL24]. Trajectory [LS14b, WK16]. Trans [SFEB23]. Transactions [Ano14-45, Ano14-77, Ano14-54, Ano14-66, Ano14-78, Ano14-55, Ano14-67, Ano14-79, Ano14-56, Ano14-68, Ano14-80, Ano14-57, Ano14-69, Ano14-46, Ano14-58, Ano14-70, Ano14-47, Ano14-59, Ano14-71, Ano14-48, Ano14-60, Ano14-72, Ano14-49, Ano14-61, Ano14-73, Ano14-50, Ano14-62, Ano14-74, Ano14-51, Ano14-63, Ano14-75, Ano14-52, Ano14-64, Ano14-76, Ano14-53, Ano14-65, Ano14a, Ano15-40, Ano15-50, Ano15-41, Ano15-51, Ano15-42, Ano15-52, Ano15-43, Ano15-53, Ano15-44, Ano15-54, Ano15-45, Ano15-55, Ano15-46, Ano15-56, Ano15-47, Ano15-57, Ano15-48, Ano15-58, Ano15-49, Ano15-59, Ano16k, Ano16l, Ano16-32, Ano16-38, Ano16-33, Ano16-39, Ano16-29, Ano16-30, Ano16-31, Ano16-34, Ano16-36, Ano16-40, Ano16-35, Ano16-37, Ano16-41, Ano16i, Ano16j, Ano16g, Ano16h, Ano17-59, Ano17-60, Ano17r, Ano17z, Ano17s, Ano17-27, Ano17t, Ano17u]. Transactions [Ano17v, Ano17w, Ano17x, Ano17y, Ano17-57, Ano17-58, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17a, Ano18v, Ano18w, Ano18x, Ano18y, Ano18z, Ano18-27, Ano18-28, Ano18-29, Ano18-30, Ano18-31, Ano18a, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano19a, Ano21a, DRFO14]. Transceiver [HCY21a]. Transcribed [SRT23]. Transcriber [GGL23]. Transcribing [GMH17]. Transcription [AS17a, CDW16, DCLHJ18, ES16, KLS^{+}21, KG16, MTBB15, NYD17, NYS17, ONKP16, PWS^{+}23, SBD16, WTY^{+}23, WJ23, WDS^{+}18]. Transcripts [BSL16]. Transducer [GWL^{+}24, RPK23]. Transducer-Based [RPK23]. Transducers [MHBT15]. Transfer [ASX22, APWB23, BRD15, BKK15, BHAMR19, BNA24, CAFFR23, CG20, DKDN20, FPTJ17, FPTJ18, GSA18, GHJ^{+}23, HMDG15, JKC^{+}24, KMG21, KMG15, LYWX22, LGGH18, LGGH19b, LWX^{+}22, LHC^{+}23, LLC^{+}24, LWYC23, LHYJ23, MBM15, MPKS^{+}23, MD16, MNFGO19, MR20, PAB19, RF19, SAPB15, SGH15b, SMH14, WYY22, WWX^{+}23, WY23, YLL21, YTWB19, ZHH19, ZS20, ZDD21, ZZZL21, ZQZ^{+}23, ZMC^{+}24, ZLL^{+}24b]. Transferable [SGL^{+}23]. Transferring [BYT^{+}21a]. Transform [AS15, DTL15, LYB^{+}23, LCTA19, NXCL16, OWA15, PP19, VH19, WCC^{+}21]. Transformation [BNA24, LJH^{+}24, MH18, WHLG14, XYY^{+}23, YLQZ24]. Transformations [CTSC14, CG16, CL17a, CL17b, CL18b]. Transformer [FCCA23, GYXY23, KHT^{+}21, KCA21a, KCA21b, LLZ^{+}21c, LXK^{+}21, LPM^{+}23, LLJ^{+}23, LSWN23, LSL24, LLT21b, LLyL21, LKR23, LLL24, MUK22, SFEB23, WY23, WLF23, YYY^{+}23, ZYD^{+}21, ZWS^{+}23]. Transformer-Based [ZWS^{+}23]. Transformer-BiLSTM [LPM^{+}23]. Transformers [LCG^{+}24, SSAS21, VSK^{+}24]. Transforms [HKHP23, WBLC22, ZKDS18]. Transient [WAS14]. Transients [DTC16]. Transition [MLC^{+}23, ODSM16, YZL^{+}18]. Transition-Based [ODSM16, YZL^{+}18]. Transitions [XZFN23]. Translate [XZSX23]. Translation [BBC^{+}16, BY15, BATS21, CZY^{+}18, CWU^{+}19, CWUS22, CLSyL21, DZD^{+}24, DFY^{+}24, DTN^{+}17, DSN18, DCW^{+}21, DZZ23, FSN24, GZX^{+}21, GYXY23, HSZ^{+}18, KCA21a, KCA21b, LLWC21, LWC^{+}18, LCZ^{+}19, LLT^{+}21a, LLZ^{+}21b, LLJ^{+}23, LWL^{+}22, LFU^{+}17, LWCL19, LLZ23b, LLL^{+}22, LZ23, LLZ^{+}23d, MHZ^{+}24, MHPB15, MVN22, NSNN20, SNK21, SZX^{+}18, TYZ^{+}22, TZZ15b, WZL^{+}15, WTZ18, WUF^{+}18, WXZ^{+}19, WZZ^{+}18, XWZ16, XZW15, YYY^{+}23, ZZZ16, ZLLS16, ZXSD17, ZZLZ19, ZXSL19, ZLS^{+}21, ZSC^{+}23a, ZWL^{+}23a, ZKKC22, ZXH^{+}16]. Transliteration [NNC19]. Transmission [LC16a]. Transparent [MR20]. Transplantation [CBG15]. Transposition [ZZZL22]. Tree [DG15, HNY^{+}15, LLL^{+}19, PHM^{+}17, SZW23, WYLZ20, YHO^{+}17, ZZY^{+}24]. Tree-Based [DG15, YHO^{+}17]. Tree-Constrained [SZW23]. Tree-Structured [WYLZ20]. Trellis [MK16]. Trellis-Based [MK16]. Trending [GLC^{+}14]. Trends [SHZ^{+}23]. Trial [CS23, FNM^{+}19]. Trial-Based [FNM^{+}19]. Trial-Dependent [CS23]. TrICy [AGB^{+}24]. Trigger [AGB^{+}24]. Trigger-Guided [AGB^{+}24]. Trimming [HNY^{+}15]. Trimodal [CHGS23]. TRINICON [MGGK17]. Tripartite [JBWZ24]. Triple [WW22, XXS^{+}23]. Triple-Domain [WW22]. Triplet [LZL^{+}23, ZKH18]. True [LP22]. Trump [YLLC19]. TS [BSW^{+}24]. TS-SEP [BSW^{+}24]. TTS [ASX22, CHC^{+}14, GXW^{+}23, LHC^{+}23, LZAY23, LCWC16, LSGL21, SRT23, ZTWY19]. Tuned [AG17, FSN24]. Tuning [GGL23, MLL^{+}22, SHL^{+}21, YM17]. Tunisian [GJB15]. Turbo [RWF16]. Turn [CZWC23, GLL20, MZZ23, SWZ16, WLZH21, WZZ^{+}24, ZLZ21]. Turn-Taking [SWZ16]. Turns [ZLZ21]. TutorNet [YLK^{+}21]. TV [PBQ15, WHB^{+}21]. Twitter [GLC^{+}14]. Two [ABB^{+}23, AG17, CC14b, GRSP17, HZF17, HZF18, LLZ^{+}21a, LYY24, LCW^{+}16, LFWZ23, MR20, QMS16, SGM^{+}15, SGH17a, SK16b, SWCN19, TOS14, TNS^{+}18, WHRC16, XHRC24, XDC^{+}17, ZTQB19, ZWW19]. Two-Channel [MR20, ZTQB19]. Two-Dimensional [AG17, XDC^{+}17]. Two-Microphone [TNS^{+}18]. Two-Pass [QMS16]. Two-Stage [ABB^{+}23, GRSP17, HZF17, LLZ^{+}21a, LYY24, LFWZ23, SGM^{+}15, SK16b, SWCN19, XHRC24, ZWW19, TOS14]. Two-Step [CC14b, HZF18]. Tying [CWYC17]. Type [AATKS21, HZF18, MSEC21, OHN16, QZLH21]. Type-2 [AATKS21]. Typed [BLZ21]. Types [AHF21, WZMH22]. Typing [WZMH22].

U [LSWN23]. U-Shaped [LSWN23]. U2 [NR22]. Ubiquitous [Ano19t]. Ultra [YLD15]. Ultra-Low [YLD15]. Ultrasound [ZAL24]. Unambiguous [RKN14]. Uncalibrated [CS23]. Uncertainties [AGV24, RLPK15]. Uncertainty [FBWG23, FD21, HDW^{+}23, KGP14, KBG16b, KBG18, RLK14, TVJ15, WZC23]. Uncertainty-Driven [HDW^{+}23]. Unconditional [BK24a]. Unconstrained [FSEI21, ZZ22]. Uncorrelating [OWA15]. Underdetermined [CY15, FK19, IISN21, SIKN23].

Underresourced [DCLHJ18]. Understanding [CSJD16, GJB15, KG17, LLZ^{+}21b, LLC^{+}23b, MWT^{+}23, MDY^{+}15, NLYK24, QCN^{+}21, SHD14, WDW^{+}22, XWY^{+}23, ZMX^{+}21]. Underwater [FZM24, ZZX^{+}23b]. Unfolded [YWW^{+}23]. Unfolding [LYZ^{+}23]. Unfrequent [MWT16]. Unified [CJY^{+}21, CMY^{+}22, GZZY23, KMAB18, LLWQ24, LZD^{+}23, LZ23, TGZ22, XHS^{+}18, YWT23]. Uniform [ATP16, BBAS19, HBC17, LB23, PCB16a, WJ15b, ZHCB21]. Unifying [KOS^{+}16, WWX^{+}24]. Unimodal [VSK^{+}24]. Unique [PBC16]. Unit [CB21, KSG24, LWZZ16, SHDG16, SLR16]. UnitNet [ZLD21]. Units [ABB^{+}23, DLY^{+}19, HBF^{+}16, HBT^{+}21, HL18, Kam23, PM19]. Universal [CCT^{+}19, KE23, LLZ^{+}21c, WQ24b]. Unlimited [HWG^{+}23]. Unloading [SDF18]. Unmanned [YLH23]. Unpaired [rLcHcC^{+}22]. Unscented [TCY15]. Unseen [JLKK21, SZVZ16]. Unsegmented [HWC^{+}17, MHBT15]. Unsupervised [BCCZ19, CLZ^{+}23, CDLX18, CHL^{+}19, CWBv19, CTLL17, CL18a, Cum21, DMSM21, FL19, GHJ^{+}23, HSN18, JK19, JG19, KJG16, KSU^{+}19, LCW24, LZAY23, LJX^{+}16, LWYC23, LHYJ23, MP16, MH18, NK21, PKR^{+}24, PBQ15, QLW^{+}24, SP16, SK17, SFRK^{+}23, SBM^{+}19, SR16b, SLR16, TXH^{+}23, WsL15, WDDL17, WSZN21, WSZN22, XYH14, XHW^{+}17, YWW^{+}23, ZH18, ZLJ22, ZZZW23]. Unsynchronized [TS16]. Untranscribed [KSL19]. Unvoiced [KKGH16]. Update [DDMvW20b, HBB17, WYHL20]. Updated [NZL23]. Updating [SF14]. Upper [CCC21]. Upsampling [LAP23, PAB19]. Urban [SK16b]. Usage [QZ22]. Use [ARSLN16, DB15, SCY22, SHDG16, CPL14]. Used [BS14, GEC15]. User [EDO17, HM14, ICB24, LXS^{+}18, WWD14]. User-Centric [ICB24]. User-Friendly [EDO17]. User-Generated [WWD14]. Users [SF14]. Using [AF18, ATS18, AM18a, ACK21, AdTJ18, ATP15, AY17, ADM^{+}17, AS15, AG17, AB15, AGC23, AM20, AATKS21, BV17, BLZ21, BC17, BRC24, BHP^{+}23, BV23, BGQO^{+}17, BHAMR19, BLLP17, BvWWM17, BPG23, BSWP18, BSTM24, BMVG17, BH18, BG23, BS21, CONVVC18, CCR16, CE15, CSS^{+}19, CDLL19, CHJ14, CHWL14, CLLD14, CWH^{+}16, CASZ17, CWCP17, CMM^{+}23, CPGL24, CG20, CS19, CMG^{+}17, CWBv19, CL18b, CL18c, DC16, DP16, DvWDM14, DAGS^{+}18, DHSG15, DAFS15, DGMB21, DGMB23, DWX^{+}21, DBM15, DCLHJ18, DTC15, DG23, ECJF21, Emu20, ES16, FSA18, FSO17, FWM^{+}14a, FWM^{+}14b, FOBP15, FPTJ17, FPTJ18, FZZ^{+}15, FDK15, FK24, FCG24, FNA15, GMH14, GJYW23, GWLL22, GvdP23, GRSP17, GJ17, GDM18, GS21, HAGSZ16, HRL^{+}21, HGT15a, HBF^{+}16, HDLL23, HFL^{+}23, HEMN21, HPPM24, HWG^{+}23, HL15, HL18, HAS21, HWW16, HWC^{+}17]. Using [HGAPB15, HHMK18, IO17, IISN21, JGM23, JK19, JHLK18, JSYW16, JTHN14, JG19, JYB^{+}15, JWLF14, JK15, JRA17, JGSC^{+}22, KHTL18, KJG16, KMG21, KK19a, KML18, KKK15, KCMW17, KMG17, KK19b, KH24, KHCW15, LSNC21, LRG21, LYY^{+}23, LWZ^{+}14, LS14a, LGGH19b, LY21, LJH^{+}24, LMC18, Lin18, LdLvWWS21, LYH^{+}22, LM23b, LAGD18, LCW^{+}16, LFU^{+}17, LLLL19, LRP^{+}23, LLC23a, LYB^{+}23, LLS24, LCTA19, MFH21, MMS^{+}20, MASC23, MJZ15, MAZS22, MD18, MMM23, MYKO24, MZX^{+}15, MMK21, MDY^{+}15, MPKS^{+}23, MZM15, MWC^{+}23, MCD14, MPC^{+}24, MCS21, MUT^{+}20, MSD14, MD16, MBTB14, MV23, MAT15, MYC^{+}23, MK16, MK15b, MJJ^{+}14, NYD17, NTA15, NTM16, NSA21, NW24, NM21, NEN21, NEWN23, NNC19, NXCL16, NNC20, NNC21, NCC15, NNL15, NDV18, NAK^{+}17, OCS21, ODG16, ÖM24, PS19, PS17, PDS^{+}16, PACPJF^{+}23, PC19, PCB16b, PJTJ24, MG19, RG15b]. Using [RRdM16, SAB16, STH^{+}18, SP16, SOF15, SIUW21, SDF18, SAP14, SAC17, SDK17, SH19, SFRK^{+}23, SG14, SGH15b, SGH15a, STBC17, SLGA15, SP17, SW20, SvWRK15, SLL^{+}16, SFK15, SgL20, SLOVB15, SMH14, SBP^{+}24, SLH17, SMSV17, SCEH19, SWC20, SLGZ15, SZVZ16, SWCN19, TZA14, TH14, TH18, TP15, TG24, TWIN23, TJV24, TP19, TR15, TNS^{+}18, TDL19, VP21, VAH19, VRN18, VM15, VP17, VAP19, VDM15, WAS14, WWD14, WGD15, WCC^{+}15, WRC16, WHRC16, WED17, WWS^{+}18, WCZ21, WCW^{+}21, WJ23, WJ15b, WAS16, WSY14, WHLG14, WSH15, WAJL15, WK16, WRKG17, WRK18, WCL^{+}21b, WSZN21, XNP^{+}18, XHRC24, XLL^{+}19, XYH14, YCM^{+}21, YB14, YGV14, YZW^{+}18, YWLZ18, YK21b, YCL22, ZP15, ZZN^{+}16, ZCYZ18, ZZZ^{+}21, ZSC^{+}23b, ZPW23, ZZY^{+}24, ZGO19, ZDLK18, ZZL^{+}23, ZDZ^{+}21, DH14, LZHN14, MAP14, NR14a, ZA14]. Utilizing [VDH24, YKPMP18]. Utterance [FWT^{+}18, GLL20, GLL^{+}21, KYTJ17, LWZZ16, WWW21, ZWC^{+}23, ZZS^{+}23]. Utterance-Level [KYTJ17]. Utterance-to-Utterance [GLL20]. Utterance-wise [WWW21]. Utterances [HYI^{+}15].

VACE [YC22]. VACE-WPE [YC22]. VAE [WTY^{+}20b, GXW^{+}23, LZAY23, WY23]. Valence [WYLZ16]. Valence-Arousal [WYLZ16]. Validation [BATS21]. Value [NYD17, ZZX^{+}21]. Valued [Emu18, HHC16, LK19, NTY19]. Variabilities [KMAB18]. Variability [MBD^{+}15, SLL^{+}16, VMOL16]. Variable [AA17, CC14a, CC15, JBC16, KKJN17, KGG19, LSNC21, NCC15, RBR16, SLW15, XNP^{+}18, ZG16, ZZH16]. Variable-order [NCC15]. Variance [ABKP19, Nos16, WCC^{+}15, WZ24, WST^{+}21]. Variant [MTK^{+}20]. Variants [ATS18]. Variational [AV16, CCC^{+}24, KKTH19, KBGAP^{+}16, LKM23, LCL^{+}23b, MZC^{+}24, NM21, OCS21, SEM^{+}14, SHCV23, WTY^{+}20b, WLT24, XHRC24]. Variations [XT23, ZL21b]. Varying [BRC24, KBGAP^{+}16, LG19, LS14b]. VC [KKTH19, WWX^{+}23]. Vector [AJLW14, BHS^{+}16, CMM^{+}23, CSY^{+}21, CP21, CL14b, CL17b, DGCY23, FWF17, FGQ^{+}15, GWLL22, GH17, HF21, KMAB18, KOS^{+}16, LHJ^{+}23, MAZ21, MPC16, NIK^{+}22, NCC15, QDSL19, SLK17, WCY15, WTY^{+}20b, WRK18, XYH14, XLLY18, YK21a, ZSB17, CL14a, CPL14, JKU18, RSR^{+}19]. Vector-Based [RSR^{+}19]. Vector-Quantized [DGCY23]. Vector-to-Vector [QDSL19]. Vectors [CLA^{+}21, CL17a, CL18b, CL18c, HMN17, KWGW17, LMC18, MUK22, MZM15, PS17, PSL17, TMMZ18]. Vehicle [YLH23]. Velvet [VP21]. Velvet-Noise [VP21]. Verification [BWZC22, Bor22, Bor23, CB21, CWZ^{+}23, Cum21, CS23, HCQ24, HH14, HWW23b, KHKY22, KVMV^{+}24, KH16, LM15, LMC17, LJH^{+}24, LM23b, LCQ23, LSLK24, LSTK21, MPC16, MPKS^{+}23, MMR^{+}22, MH18, O’S24, QCW21, QCL23, RSR^{+}19, STH^{+}18, STT^{+}19, TMM18, TMMZ18, VMOL16, WCC^{+}15, WHQY19, WDD^{+}16, WLL^{+}22, XRWL21, ZSB17, ZKH18, ZWL^{+}23b, ZZGD21, ZM23]. Vernacular [HK24]. Versus [SSKP16, TP22]. Vertical [GSA18]. Very [CLAG16, ESVK17, HW14, QBTY16, TQH^{+}18, WWWR23]. Via [BYT^{+}21a, DTDL16, FYG^{+}21, HKPR21, HZC^{+}22, LLWJ16, MXH21, NNC21, SLGZ15, WBC^{+}21, XV21, AFG^{+}23, BWZC22, BB21, CWDA17, CJY^{+}22, CLSyL21, CHL^{+}23, CT14, DCZ19, DHCS15, ECH23, GDC19, HGYN23, HS15, HCZC24, HHC16, LNNL22, LWL^{+}19, LPM^{+}23, LLZ23b, LKR23, LLZ^{+}23c, LLL^{+}23, MYNO23, MIR^{+}21, NW15, NLL^{+}18, PKA14, RPL^{+}23, SM14, TALP24, WXZ^{+}19, YPC^{+}22, ZH14, ZZZ16, ZXH^{+}16, ZLL^{+}24a, ZTM^{+}23]. Vibrating [PBA^{+}23]. Vibrato [MSEC21]. Video [CLHK21, KS21, LLZ^{+}21c, ROL^{+}15]. View [ADLX23, GLC^{+}14, HZC^{+}22]. Violin [PCW15]. Virtual [BBFS21, BATS21, ECH23, MGT15, YC22, ZLSC23]. Vision [WTY^{+}23]. Vision-Based [WTY^{+}23]. Visual [AZK15, BJ24, CWL22, CLW24, CWT22, DTC15, HKPR21, KML18, KBG^{+}23, LM17, LLZ^{+}21c, LDG^{+}23, LLWQ24, LDZ19, MXH21, MTZ^{+}21, OVN^{+}22, PEO^{+}20, QCW21, QWW^{+}23, SLWY20, YZW^{+}21, ZYZU22, ZAL24, ZDZ^{+}21]. Visual-LSTM [MXH21]. Visualizing [YLLC19]. Visually [KSL19]. Vocabularies [ESVK17]. Vocabulary [DB15, SSI21, SFIL17, TB18, YČS23]. Vocal [AG17, ADGE19, CWJ15, CG21, DMP^{+}18, IIY16, KBG^{+}23, LSW18, LTE^{+}21, MVH16, SAB16, SIUW21, SW20, SMH14, SMB18, ZPD19, ZLJB23]. Vocal-Specific [SW20]. Vocalization [AHF21, HSWC21, JK19]. Vocalized [SGB22]. VocalTractLab [GBL24]. Vocoder [AL23, MOT^{+}23, OD17, SZL19, YPH23, YWT23]. Vocoding [AJB^{+}18, CGL^{+}15]. Voice [ATA16, AM20, AATKS21, BNA24, BMVG17, CLLD14, CC14b, DAR16, DZLGO20, DLL^{+}21, DWX^{+}21, DTC15, DTC16, HSM19, HdHTJ22, HKPR21, HHO^{+}21, HL15, HHW^{+}21, IIY16, KYI^{+}15, KKTH19, KHT^{+}21, KKJK23, LNNL22, Lib21, LLLL19, LCW^{+}21, LCTA19, LLL^{+}23, MLMT24, McL14, MZC^{+}24, MDCS20, MCS21, MKMDM17, MSEC21, NTA15, NTM16, NSA21, NR22, OR16a, PS17, PM15, RPL^{+}23, STH^{+}18, SFDRB21, SYKL21, TOS14, TLW^{+}17, WWX^{+}23, WVCL14, WHLG14, XT23, XSC^{+}23, YLY15, YDW^{+}21, YWW^{+}23, ZW16a, ZLL^{+}19, ZLD20, ZZZL21, ZLJB23, ZKDS18, ZWTL23]. Voice-Based [HdHTJ22]. Voiced [JP14, KKGH16, KG14, NJC16a, NJC16b, VAP19]. Voiced/Unvoiced [KKGH16]. Voices [AHSS^{+}18, NYS17, OR16a]. Voicing [ZWW23a]. Vol [Ano14a, Ano17a, Ano18a, Ano21a]. Volatility [NCB23]. Voltage [SLGA15]. Volterra [MJJ^{+}14]. Volume [ATP16, AGD^{+}22, BHBS16, WGLZ21]. Vowel [CPGL24, CWJ16, LPI^{+}21, MASC23, MP20, MLC^{+}23]. VQ [WTY^{+}20b, GXW^{+}23]. VQ-VAE [WTY^{+}20b, GXW^{+}23]. vs [BPG23, RGK24]. VTS [DH14].

Wait [LCL^{+}21]. Walk [yLsL14]. Wall [MFH21]. Warped [SKH^{+}15b]. Warping [KB14, Lib21, MAP14, TLW^{+}17]. Watermark [DLMM^{+}18]. Watermark-Driven [DLMM^{+}18]. Watermarking [FM15, HGT15c, NXH^{+}17, SAA15, XNG^{+}14, XNRG15, XNP^{+}18, ZZX^{+}21, ZZX^{+}23a]. Wav2code [HCZC24]. wav2vec [FSN24]. Wave [ADM^{+}19, BGQO^{+}17, BBFS21, BH19, DQY18, Emu20, FFSS17, FFSS18, GdBBS21, KFH^{+}14, SK16a, WWH^{+}18, ZAZ^{+}18, ZHB^{+}18]. Wave-Based [BH19]. Wave-Domain [Emu20, ZHB^{+}18]. Wavefield [SAP14]. Waveform [FWT^{+}18, JBTA19, LAGD18, OCS21, WTY20a, WFY^{+}22, WHO^{+}21, WHT^{+}21, YHO^{+}18]. Waveform-Based [OCS21]. Waveforms [AP21]. WaveGAN [WHO^{+}21]. Waveguide [GDM18, SMH14, SMSV17]. Waveguides [MSS17, PMB^{+}19]. Wavelet [AGW15, DG23, FZZ^{+}15, HKHP23, ISZA15, LCTA19]. Wavelets [WLT^{+}18]. WaveNet [CWBv19, SZL19, WHT^{+}21]. Wavenumber [MUT^{+}20]. Waves [CNFG23, HHZ^{+}21]. Wavesplit [ZG21]. Way [GZX^{+}21, YLC^{+}24]. WDSRL [MHZ^{+}24].

Weak [MMM23, XLG^{+}23]. Weakly [DWY21, KXS^{+}19, MSB18, PEO^{+}20, RV17]. Weapon [SHAGPRZ17]. Web [SMSV17]. Weight [BDV^{+}14, LC24, LM23b, LS14b, NLYK24]. Weighted [ARSA14, AM18b, BBM17, BWZC22, BOSR23, CAFFR23, DMVP16, GFS22, HF21, HBCC22, OFF^{+}17, MG19, SH19, UKS19, WYLZ16, YC22, ZTDD23b]. Weighted-Orthogonal [DMVP16]. Weighting [JJB16, JBTL15, KKJN17, PLZL17, SH14a, WUF^{+}18]. Weights [AZK15]. Well [WK24a]. Where [QCN^{+}21]. Which [ZYZU22]. Whisper [GBH16, KH21]. Whispered [GBH16, GJ17]. Whistles [KD21]. Whole [CCL^{+}21]. Wideband [AF18, HL14, HFL^{+}23, SLC^{+}15]. Wiener [AM18b, GS21, IISN21, MHD15, MHGD15, SMVW14, SBP^{+}24, TE23, ZTDD23b, ZDLK18, dCBC23]. Wild [DWX^{+}21, LWS^{+}23]. Win [MLL^{+}22]. Win-Win [MLL^{+}22]. Window [KMG17, YZLY17, ZN14]. Window-Dominant [YZLY17]. Windy [CZN^{+}24]. Wired [KSK21]. Wireless [AM18a, AHH^{+}20, BBM17, CG17, CPSFC^{+}14, DMHZ22, KSHH18, LP22, RCR^{+}21, RvWM22, SHAGPRZ17, ZHH18, ZHH19, ZL21a, ZTDD23a]. Wise [CLLD14, WZ24, WFR^{+}21, YOM21, YLPL19, WWW21]. With-Height [KFH^{+}14]. Without [GLQ16, GLW20, JGK^{+}22, MSW^{+}22, SSS18, ZLHL21, ZDGO21]. WNSA [HFL^{+}23]. WNSA-Net [HFL^{+}23]. Word [AAA19, CLZ^{+}23, CPCBFSF16, CSJD16, CLW^{+}19, CBCL15, CLSyL21, CCL^{+}21, DLY^{+}19, DTN^{+}17, KJG16, KMG21, Kam23, LCZ^{+}19, LLCT22, LGL^{+}22, LJL^{+}23, MWT16, MHZ^{+}24, SSI21, SZW23, TXH^{+}23, TGZ22, WTZ18, WZ24, WZH22, YCLS17, YM17, YWLZ18, YUK21, ZYF18, ZL21b, ZLZW21, ZZXH22, ZYZU22, ZKKC22, ZLJ22, ŞUY^{+}18]. Word-Level [DTN^{+}17, LLCT22, MHZ^{+}24, ZL21b]. Word-Pair [CBCL15]. Word-Region [ZKKC22]. Word2vec [CHL^{+}19]. Words [GBF21, LCYW21, SFIL17, WYLZ16]. Work [WK24a]. Working [XSL^{+}19]. Workplace [JNP^{+}21]. World [CWR^{+}22, PWS^{+}23]. Worst [BC17, WC16a]. Worst-Case [BC17, WC16a]. WPE [YC22]. Write [CGL^{+}24]. Written [PBQ15].

x [FdDARF16, CC14a]. XABSA [LFL^{+}23].

Yor [Chi15a, RV14].

Zero [ASC^{+}18, DLH23, JKC^{+}24, KMG21, SD16b, XV21, ZMC^{+}24]. Zero-Note [DLH23]. Zero-Resource [KMG21]. Zero-Shot [ASC^{+}18, JKC^{+}24, XV21, ZMC^{+}24]. Zone [MNFGO19, ZSC^{+}23b]. Zones [LSNC21, MØ20, SLZ^{+}21].

References

Ahmed:2017:GSA

[AA17] Shakeel Ahmed and Muhammad Tahir Akhtar. Gain scheduling of auxiliary noise and variable stepsize for online acoustic feedback cancellation in narrow-band active noise control systems. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(2): 333–343, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Azmi:2019:RWE

[AAA19] A. M. Azmi, M. N. Almutery, and H. A. Aboalsamh. Real-word errors in Arabic texts: a better algorithm for detection and correction. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(8): 1308–1320, August 2019. CODEN ???? ISSN 2329-9290.

Azadi:2021:RVF

[AATKS21] Hamid Azadi, Mohammad-R. Akbarzadeh-T, Hamid-R. Kobravi, and Ali Shoeibi. Robust voice feature selection using interval type-2 fuzzy AHP for automated diagnosis of Parkinson’s disease. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2792–2802, 2021. CODEN ???? ISSN 2329-9290.

Arora:2014:MSC

[AB14] V. Arora and L. Behera. Musical source clustering and identification in polyphonic audio. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (6):1003–1012, June 2014. CODEN ???? ISSN 2329-9290.

Arora:2015:MFE

[AB15] V. Arora and L. Behera. Multiple F0 estimation and source clustering of polyphonic music audio using PLCA and HM-RFs. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(2):278–287, February 2015. CODEN ???? ISSN 2329-9290.

Abdelwahab:2018:DAA

[AB18] M. Abdelwahab and C. Busso. Domain adversarial for acoustic emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2423– 2435, December 2018. CODEN ???? ISSN 2329-9290.

Ahrens:2021:CSH

[AB21] Jens Ahrens and Stefan Bilbao. Computation of spherical harmonic representations of source directivity based on the finite-distance signature. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):83–92, January 2021.

URL https://dl.acm.
org/doi/10.1109/TASLP.
2020.3037471.

Airaksinen:2017:QPA

[ABA17] Manu Airaksinen, Tom Bäckström, and Paavo Alku. Quadratic programming approach to glottal inverse filtering by joint norm-1 and norm-2 optimization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(5): 929–939, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7636994/.

Albertini:2023:TSB

[ABB^{+}23] Davide Albertini, Alberto Bernardini, Federico Borra, Fabio Antonacci, and Augusto Sarti. Twostage beamforming with arbitrary planar arrays of differential microphone array units. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):590– 602, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3231719.

Abdelaziz:2018:CFM

[Abd18] A. H. Abdelaziz. Comparing fusion models for DNN-based audiovisual continuous speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(3):475– 484, March 2018. CODEN ???? ISSN 2329-9290.

Asad:2019:RTL

[ABKP19] H. As’ad, M. Bouchard, and H. Kamkar-Parsi. A robust target linearly constrained minimum variance beamformer with spatial cues preservation for binaural hearing aids. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (10):1549–1563, October 2019. CODEN ???? ISSN 2329-9290.

Ali:2019:MEG

[ABvM19] R. Ali, G. Bernardi, T. van Waterschoot, and M. Moonen. Methods of extending a generalized sidelobe canceller with external microphones. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(9):1349– 1364, September 2019. CODEN ???? ISSN 2329-9290.

Alkaher:2024:HDG

[AC24] Yehav Alkaher and Israel Cohen. Howling detection and gain control for speech reinforcement in a noisy car cabin environment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1494– 1505, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3364091.

Asaei:2017:PIL

[ACB17] Afsaneh Asaei, Milos Cernak, and Hervé Bourlard. Perceptual information loss due to impaired speech production. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12): 2433–2443, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8007273/.

Alkhaldi:2021:FFS

[ACK21] Tareq Alkhaldi, Chenhui Chu, and Sadao Kurohashi. Flexibly focusing on supporting facts, using bridge links, and jointly training specialized modules for multi-hop question answering. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 3216–3225, 2021. CODEN ???? ISSN 2329-9290.

Arisoy:2014:CNN

[ACRS14] E. Arisoy, S. F. Chen, B. Ramabhadran, and A. Sethy. Converting neural network language models into back-off language models for efficient decoding in automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):184–192, January 2014. CODEN ???? ISSN 2329-9290.

Arnela:2019:MBV

[ADGE19] M. Arnela, S. Dabbaghchian, O. Guasch, and O. Engwall. MRIbased vocal tract representations for the threedimensional finite element synthesis of diphthongs. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2173–2182, December 2019. CODEN ???? ISSN 2329-9290.

Andersen:2016:PIN

[AdHTJ16] Asger Heidemann Andersen, Jan Mark de Haan, Zheng-Hua Tan, and Jesper Jensen. Predicting the intelligibility of noisy and nonlinearly processed binaural speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(11): 1908–1920, ???? 2016. CODEN ???? ISSN 2329-9290.

An:2023:GVS

[ADLX23] Jiaming An, Zixiang Ding, Ke Li, and Rui Xia. Global-view and speaker-aware emotion cause extraction in conversations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3814– 3823, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3319990.

Antonello:2017:RIR

[ADM^{+}17] Niccolò Antonello, Enzo De Sena, Marc Moonen, Patrick A. Naylor, and Toon van Waterschoot. Room impulse response interpolation using a sparse spatiotemporal representation of the sound field. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(10): 1929–1941, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7987742/.

Antonello:2019:JAL

[ADM^{+}19] N. Antonello, E. De Sena, M. Moonen, P. A. Naylor, and T. van Waterschoot. Joint acoustic localization and dereverberation through plane wave decomposition and sparse regularization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27 (12):1893–1905, December 2019. CODEN ???? ISSN 2329-9290.

Andersen:2018:NSI

[AdTJ18] A. H. Andersen, J. M. de Haan, Z. Tan, and J. Jensen. Nonintrusive speech intelligibility prediction using convolutional neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):1925–1939, October 2018. CODEN ???? ISSN 2329-9290.

Abel:2018:ASB

[AF18] Johannes Abel and Tim Fingscheidt. Artificial speech bandwidth extension using deep neural networks for wideband spectral envelope estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(1):71–83, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8063328/.

Abel:2019:SBL

[AF19] J. Abel and T. Fingscheidt. Sinusoidalbased lowband synthesis for artificial speech bandwidth extension. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(4): 765–776, April 2019. CO-DEN ???? ISSN 2329-9290.

Ai:2023:CML

[AF23] Xi Ai and Bin Fang.

Cross-modal language modeling in multi-motion-informed context for lip reading. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2220– 2232, 2023. ISSN 2329-

9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2023.3282109.

Antonanzas:2023:RMT

[AFdDG23] Christian Antoñanzas, Miguel Ferrer, Maria de Diego, and Alberto Gonzalez. Remote microphone technique for active noise control over distributed networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1522–1535, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3264600.

Abderrazek:2023:IDR

[AFG^{+}23] Sondes Abderrazek, Corinne Fredouille, Alain Ghio, Muriel Lalain, Christine Meunier, and Virginie Woisard. Interpreting deep representations of phonetic features via neuro-based concept detector: Application to speech disorders due to head and neck cancer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):200– 214, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3221039.

Arnela:2017:FES

[AG17] Marc Arnela and Oriol Guasch. Finite element synthesis of diphthongs using tuned twodimensional vocal tracts. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (10):2013–2023, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8000370/.

Agarwal:2024:TTG

[AGB^{+}24] Vibhav Agarwal, Sourav Ghosh, Harichandana BSS, Himanshu Arora, and Barath Raj Kandur Raja. TrICy: Trigger-guided data-to-text generation with intent aware attentioncopy. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1173–1184, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3353574.

Asaei:2014:SSM

[AGBC14] A. Asaei, M. Golbabaee, H. Bourlard, and V. Cevher. Structured sparsity models for reverberant speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(3):620– 633, March 2014. CODEN ???? ISSN 2329-9290.

Atri:2023:MDS

[AGC23] Yash Kumar Atri, Vikram Goyal, and Tanmoy Chakraborty.

Multi-document summarization using selective attention span and reinforcement learning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 3457–3467, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3316459.

Atalay:2022:SDN

[AGD^{+}22] Timuçin Berk Atalay, Zühre Sü Gül, Enzo De Sena, Zoran Cvetković, and Hüseyin Hacıhabiboğlu. Scattering delay network simulator of coupled volume acoustics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 582–593, 2022. CODEN ???? ISSN 2329-9290.

Abesser:2023:HRA

[AGM23] Jakob Abeßer, Sascha Grollmisch, and Meinard Müller. How robust are audio embeddings for polyphonic sound event tagging? IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2658– 2667, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3293032.

Accolti:2024:URA

[AGV24] Ernesto Accolti, Javier Gimenez, and Michael Vorländer. Uncertainties of room acoustics simulation due to directivity data of musical instruments. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):300– 309, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3330117.

Ahani:2015:SRB

[AGW15] S. Ahani, S. Ghaemmaghami, and Z. J. Wang. A sparse representationbased wavelet domain speech steganography method. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (1):80–91, January 2015. CODEN ???? ISSN 2329-9290.

Anders:2021:CAN

[AHF21] Franz Anders, Mario Hlawitschka, and Mirco Fuchs. Comparison of artificial neural network types for infant vocalization classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):54–67, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3037414.

Amini:2019:ACR

[AHH^{+}19] J. Amini, R. C. Hendriks, R. Heusdens, M. Guo, and J. Jensen. Asymmetric coding for rateconstrained noise reduction in binaural hearing aids. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(1):154–167, January 2019. CODEN ???? ISSN 2329-9290.

Amini:2020:RCN

[AHH^{+}20] Jamal Amini, Richard Christian Hendriks, Richard Heusdens, Meng Guo, and Jesper Jensen. Rateconstrained noise reduction in wireless acoustic sensor networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 1–12, January 2020. CO-DEN ???? ISSN 2329-9290. URL https:/

                                        /dl.acm.org/doi/abs/
                                        10.1109/TASLP.2019.2947777.

Abdel-Hamid:2014:CNN

[AHMJ^{+}14] O. Abdel-Hamid, A.-R. Mohamed, Hui Jiang, Li Deng, G. Penn, and Dong Yu. Convolutional neural networks for speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(10):1533– 1545, October 2014. CO-DEN ???? ISSN 2329-9290.

Aichinger:2018:TMF

[AHSS^{+}18] Philipp Aichinger, Martin Hagmüller, Berit Schneider-Stickler, Jean Schoentgen, and Franz Pernkopf. Tracking of multiple fundamental frequencies in diplophonic voices. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(2): 330–341, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8062798/.

Airaksinen:2018:CBS

[AJB^{+}18] M. Airaksinen, L. Juvela, B. Bollepalli, J. Yamagishi, and P. Alku. A comparison between STRAIGHT, glottal, and sinusoidal vocoding in statistical parametric speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1658– 1670, September 2018. CODEN ???? ISSN 2329-9290.

Alhakeem:2024:DRL

[AJK24] Zainab Alhakeem, Se-In Jang, and Hong-Goo Kang. Disentangled representations in local-global contexts for Arabic dialect identification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 879–890, 2024. ISSN 2329-

9290. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2023.3341006.

Alinaghi:2014:JMV

[AJLW14] A. Alinaghi, P. J. Jackson, Qingju Liu, and Wenwu Wang. Joint mixing vector and binaural model based stereo source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (9):1434–1448, September 2014. CODEN ???? ISSN 2329-9290.

Adalbjornsson:2016:SLH

[AKB^{+}16] S. I. Adalbjornsson, T. Kronvall, S. Burgess, K. Astrom, and A. Jakobsson. Sparse localization of harmonic audio sources. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(1): 117–129, January 2016. CODEN ???? ISSN 2329-9290.

Abel:2017:IQM

[AKG^{+}17] Johannes Abel, Magdalena Kaniewska, Cyril Guillaumé, Wouter Tirry, and Tim Fingscheidt. An instrumental quality measure for artificially bandwidth-extended speech signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(2):384– 396, ???? 2017. CODEN ???? ISSN 2329-9290.

Ardekani:2016:SIP

[AKN^{+}16] I. T. Ardekani, J. P. Kaipio, A. Nasiri, H. Sharifzadeh, and W. H. Abdulla. A statistical inverse problem approach to online secondary path modeling in active noise control. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):54–64, January 2016. CODEN ???? ISSN 2329-9290.

Abe:2023:AMM

[AKUS23] Takumi Abe, Shoichi Koyama, Natsuki Ueno, and Hiroshi Saruwatari. Amplitude matching for multizone sound field control. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):656– 669, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3231715.

Ai:2023:AAF

[AL23] Yang Ai and Zhen-Hua Ling. APNet: an all-frame-level neural vocoder incorporating direct prediction of amplitude and phase spectra. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2145– 2157, 2023. ISSN 2329-

9290. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2023.3277276.

Anidjar:2021:HST

[ALH^{+}21] Or Haim Anidjar, Itshak Lapidot, Chen Hajaj, Amit Dvir, and Issachar Gilad. Hybrid speech and text analysis methods for speaker change detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2324–2338, 2021. CODEN ???? ISSN 2329-9290.

Andersen:2016:ATF

[AM16] Kristian Timm Andersen and Marc Moonen. Adaptive time-frequency analysis for noise reduction in an audio filter bank with low delay. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 784–795, April 2016. CO-DEN ???? ISSN 2329-9290.

Alexandridis:2018:MSS

[AM18a] Anastasios Alexandridis and Athanasios Mouchtaris. Multiple sound source location estimation in wireless acoustic sensor networks using DOA estimates: The dataassociation problem. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(2): 342–356, ???? 2018.

CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8105842/.

Andersen:2018:RSD

[AM18b] Kristian Timm Andersen and Marc Moonen. Robust speechdistortion weighted interframe Wiener filters for single-channel noise reduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(1):97–107, ???? 2018. CODEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/8063896/.

Azad:2020:RSF

[AM20] Abul Azad and Lamine Mili. Robust speech filter and voice encoder parameter estimation using the phase–phase correlator. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):592–604, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2961556.

Ando:2020:CSE

[AMK^{+}20] Atsushi Ando, Ryo Masumura, Hosana Kamiyama, Satoshi Kobashikawa, Yushi Aono, and Tomoki Toda.

Customer satisfaction estimation in contact center calls based on a hierarchical multi-task model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):715– 728, January 2020. CO-DEN ???? ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.2966857.

Anonymous:2014:IIA

[Ano14a] Anonymous. 2014 index IEEE/ACM Transactions on Audio, Speech, and Language Processing vol.

22. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2264–2288, December 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:BPa

[Ano14b] Anonymous. [Blank page]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (1):B287–B288, January 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:BPc

[Ano14c] Anonymous. [Blank page]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (3):B740, March 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:BPd

[Ano14d]  Anonymous.  Blank

page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(5):B997–B998, May 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:BPe

[Ano14e] Anonymous. [Blank page].

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (7):B1219–B1220, July 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:BPf

[Ano14f]  Anonymous.  [Blank

page]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(9):B1449–B1450, September 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:BPg

[Ano14g]  Anonymous.  [Blank

page]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(11):B1681–B1682, November 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:BPh

[Ano14h] Anonymous. [Blank page].

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):B1686, December 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:BPBa

[Ano14i] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): C4, January 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:BPBb

[Ano14j] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): C4, February 2014. CO-DEN ???? ISSN 2329-9290. Anonymous:2014:BPBc

[Ano14k] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(3): C4, March 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:BPBd

[Ano14l] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(4): C4, April 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:BPBe

[Ano14m] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(5):

C4, May 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:BPBf

[Ano14n] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(6): C4, June 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:BPBg

[Ano14o] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(7): C4, July 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:BPBh

[Ano14p] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(8): C4, August 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:BPBi

[Ano14q] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(9): C4, September 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:BPBj

[Ano14r] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(10): C4, October 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:BPBk

[Ano14s] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(11): C4, November 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:BPBl

[Ano14t] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(12): C4, December 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:FCa

[Ano14u] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (1):C1, January 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:FCb

[Ano14v] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (2):C1, February 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:FCc

[Ano14w]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(3): C1, March 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:FCd

[Ano14x]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(4): C1, April 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:FCe

[Ano14y]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(5):C1, May 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:FCf

[Ano14z]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(6): C1, June 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:FCg

[Ano14-27]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(7):C1, July 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:FCh

[Ano14-28]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):C1, August 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:FCi

[Ano14-29] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (9):C1, September 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:FCj

[Ano14-30] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (10):C1, October 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:FCk

[Ano14-31] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (11):C1, November 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:FCl

[Ano14-32] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):C1, December 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:ISPa

[Ano14-33] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):C3, January 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:ISPb

[Ano14-34] Anonymous. IEEE Signal Processing Society information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): C3, February 2014. CO-DEN ???? ISSN 2329-9290. Anonymous:2014:ISPc

[Ano14-35] Anonymous. IEEE Signal Processing Society information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(3): C3, March 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:ISPd

[Ano14-36] Anonymous. IEEE Signal Processing Society information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(4): C3, April 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:ISPe

[Ano14-37] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(5):C3, May 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:ISPf

[Ano14-38] Anonymous. IEEE Signal Processing Society information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(6): C3, June 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:ISPg

[Ano14-39] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(7):C3, July 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:ISPh

[Ano14-40] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):C3, August 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:ISPi

[Ano14-41] Anonymous. IEEE Signal Processing Society information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(9):

C3, September 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:ISPj

[Ano14-42] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(10):C3, October 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:ISPk

[Ano14-43] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(11):C3, November 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:ISPl

[Ano14-44] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):C3, December 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATa

[Ano14-45] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing — EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (1):289–290, January 2014.

CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATd

[Ano14-46] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing — EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (2):576–577, February 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATg

[Ano14-47]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (3):741–742, March 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATj

[Ano14-48]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (4):863–864, April 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATm

[Ano14-49]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (5):999–1000, May 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATp

[Ano14-50]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (6):1109–1110, June 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATs

[Ano14-51]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (7):1221–1222, July 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATv

[Ano14-52]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (8):1336–1337, August 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATy

[Ano14-53]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (9):1451–1452, September 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATab

[Ano14-54]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (10):1571–1572, October 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATae

[Ano14-55]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (11):1683–1684, November 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATah

[Ano14-56]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2260–2261, December 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATb

[Ano14-57]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): 291–292, January 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATe

[Ano14-58]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): 578–579, February 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATh

[Ano14-59]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(3): 743–744, March 2014. CODEN ???? ISSN 2329-

Anonymous:2014:IATk

[Ano14-60]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(4): 865–866, April 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:IATn

[Ano14-61] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(5):1001– 1002, May 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATq

[Ano14-62] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(6):1111– 1112, June 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATt

[Ano14-63] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(7):1223– 1224, July 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATw

[Ano14-64] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):1338– 1339, August 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:IATz

[Ano14-65] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(9):1453– 1454, September 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATac

[Ano14-66] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(10):1573– 1574, October 2014. CO-DEN ???? ISSN 2329-

Anonymous:2014:IATaf

[Ano14-67] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(11):1685– 1686, November 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATai

[Ano14-68] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2262– 2263, December 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATc

[Ano14-69] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):C2, January 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATf

[Ano14-70]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): C2, February 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:IATi

[Ano14-71]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(3): C2, March 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATl

[Ano14-72]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(4): C2, April 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATo

[Ano14-73] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(5):C2, May 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATr

[Ano14-74]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(6): C2, June 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATu

[Ano14-75] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(7):C2, July 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATx

[Ano14-76] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):C2, August 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:IATaa

[Ano14-77]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(9): C2, September 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:IATad

[Ano14-78]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(10): C2, October 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:IATag

[Ano14-79]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(11): C2, November 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:IATaj

[Ano14-80]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(12): C2, December 2014. CO-DEN ???? ISSN 2329-

Anonymous:2014:LR

[Ano14-81] Anonymous. List of reviewers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2257– 2259, December 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:OAa

[Ano14-82] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):293, January 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:OAb

[Ano14-83] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):580, February 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:OAc

[Ano14-84] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (3):745, March 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:OAd

[Ano14-85] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (4):867, April 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:OAe

[Ano14-86] Anonymous. Open access.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (7):1225, July 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:OAf

[Ano14-87] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):1340, August 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:OAg

[Ano14-88] Anonymous. Open access.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (9):1455, September 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:OAh

[Ano14-89] Anonymous. Open access.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (10):1575, October 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:PYAa

[Ano14-90] Anonymous. Publish your article in IEEE Access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (3):746, March 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:PYAb

[Ano14-91] Anonymous. Publish your article in IEEE Access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (4):868, April 2014. CO-DEN ???? ISSN 2329-9290.

Anonymous:2014:TCa

[Ano14-92] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):1–2, January 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:TCb

[Ano14-93] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):3–4, January 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:TCc

[Ano14-94] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):289–290, February 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:TCd

[Ano14-95] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):291–292, February 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCe

[Ano14-96] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(3):581–582, March 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCf

[Ano14-97] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(3):583–584, March 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCg

[Ano14-98] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(4):741–742, April 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCh

[Ano14-99] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(4):743–744, April 2014. CODEN ???? ISSN 2329-

Anonymous:2014:TCi

[Ano14-100] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(5):869–870, May 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCj

[Ano14-101] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(5):871–872, May 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCk

[Ano14-102] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(6):999–1000, June 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCl

[Ano14-103] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(6):1001– 1002, June 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCm

[Ano14-104] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(7):1113–1114, July 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:TCn

[Ano14-105] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(7):1115–1116, July 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCo

[Ano14-106] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):1221–1222, August 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCp

[Ano14-107] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):1223–1224, August 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCq

[Ano14-108] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(9):1341–1342, September 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCr

[Ano14-109] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(9):1343–1344, September 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:TCs

[Ano14-110] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(10):1451–1452, October 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCt

[Ano14-111] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(10):1453–1454, October 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCu

[Ano14-112] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(11):1577–1578, November 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCv

[Ano14-113] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(11):1579–1580, November 2014. CODEN ???? ISSN 2329-9290. Anonymous:2014:TCw

[Ano14-114] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1683–1685, December 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:TWAa

[Ano14-115] Anonymous. Together, we are advancing technology. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (9):1456, September 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2014:TWAb

[Ano14-116] Anonymous. Together, we are advancing technology. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (10):1576, October 2014. CODEN ???? ISSN 2329-9290.

Anonymous:2015:BPa

[Ano15a] Anonymous. [Blank page].

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (1):222, January 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:BPc

[Ano15b] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (3):B600, March 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:BPd

[Ano15c] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (3):C4, March 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:BPe

[Ano15d] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (4):C4, April 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:BPf

[Ano15e]  Anonymous.  Blank

page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(5):B951–B952, May 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:BPg

[Ano15f] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (5):C4, May 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:BPh

[Ano15g] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (6):C4, June 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:BPi

[Ano15h] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (7):B1243–B1244, July 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:BPj

[Ano15i] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (7):C4, July 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:BPk

[Ano15j] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (8):C4, August 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:BPl

[Ano15k]  Anonymous.  Blank

page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(9):B1533–B1534, September 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:BPm

[Ano15l] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (9):C4, September 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:BPn

[Ano15m] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (10):C4, October 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:BPBa

[Ano15n] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(1): C4, January 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:BPBb

[Ano15o] Anonymous. [Blank page — back cover]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): C4, February 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:FCa

[Ano15p]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):C1, January 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:FCb

[Ano15q]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(2): C1, February 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:FCc

[Ano15r] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (3):C1, March 2015. CO-DEN ???? ISSN 2329-9290. Anonymous:2015:FCd

[Ano15s] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (4):C1, April 2015. CO-DEN ???? ISSN 2329-9290. Anonymous:2015:FCe

[Ano15t]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(5):C1, May 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:FCf

[Ano15u]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(6): C1, June 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:FCg

[Ano15v] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (7):C1, July 2015. CO-DEN ???? ISSN 2329-

Anonymous:2015:FCh

[Ano15w] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (8):C1, August 2015. CO-DEN ???? ISSN 2329-9290. Anonymous:2015:FCi

[Ano15x] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (9):C1, September 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:FCj

[Ano15y] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (10):C1, October 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:HCYa

[Ano15z] Anonymous. How can you get your idea to market first? IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(9):1539, September 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:HCYb

[Ano15-27] Anonymous. How can you get your idea to market first? IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1708, October 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IMDa

[Ano15-28] Anonymous. IEEE member digital library. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(3): 606, March 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IMDb

[Ano15-29] Anonymous. IEEE member digital library. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(4): 812, April 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:ISPa

[Ano15-30] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):C3, January 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:ISPb

[Ano15-31] Anonymous. IEEE Signal Processing Society information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): C3, February 2015. CO-DEN ???? ISSN 2329-9290. Anonymous:2015:ISPc

[Ano15-32] Anonymous. IEEE Signal Processing Society information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(3): C3, March 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:ISPd

[Ano15-33] Anonymous. IEEE Signal Processing Society information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(4): C3, April 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:ISPe

[Ano15-34] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(5):C3, May 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:ISPf

[Ano15-35] Anonymous. IEEE Signal Processing Society information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(6): C3, June 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:ISPg

[Ano15-36] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(7):C3, July 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:ISPh

[Ano15-37] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):C3, August 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:ISPi

[Ano15-38] Anonymous. IEEE Signal Processing Society information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(9): C3, September 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:ISPj

[Ano15-39] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):C3, October 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATa

[Ano15-40]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (1):223–224, January 2015.

CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATc

[Ano15-41]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (2):420–421, February 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATe

[Ano15-42]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (3):601–602, March 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATg

[Ano15-43]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (4):807–808, April 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATi

[Ano15-44]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (5):953–954, May 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATk

[Ano15-45]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (6):1096–1097, June 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATm

[Ano15-46]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (7):1245–1246, July 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATo

[Ano15-47]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (8):1381–1382, August 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATq

[Ano15-48]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (9):1535–1536, September 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATs

[Ano15-49]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing EDICS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (10):1704–1705, October 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATb

[Ano15-50] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):C2, January 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATd

[Ano15-51]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): C2, February 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:IATf

[Ano15-52]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(3): C2, March 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATh

[Ano15-53]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(4): C2, April 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATj

[Ano15-54] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(5):C2, May 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATl

[Ano15-55]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(6): C2, June 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATn

[Ano15-56] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(7):C2, July 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATp

[Ano15-57] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):C2, August 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IATr

[Ano15-58]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(9): C2, September 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:IATt

[Ano15-59]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(10): C2, October 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:IAa

[Ano15-60]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(1): 225–226, January 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IAb

[Ano15-61]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): 422–423, February 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IAc

[Ano15-62]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(3): 603–604, March 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IAd

[Ano15-63]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(4): 809–810, April 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:IAe

[Ano15-64]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(5): 955–956, May 2015. CO-DEN ???? ISSN 2329-9290.

Anonymous:2015:IAf

[Ano15-65]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(6): 1098–1099, June 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IAg

[Ano15-66] Anonymous. Information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(7):1247– 1248, July 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:IAh

[Ano15-67]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(8): 1383–1384, August 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:IAi

[Ano15-68] Anonymous. Information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(9):1537– 1538, September 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:IAj

[Ano15-69]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(10): 1706–1707, October 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:LR

[Ano15-70] Anonymous. List of reviewers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2474– 2476, December 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:OAa

[Ano15-71] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):227, January 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:OAb

[Ano15-72] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(2):424, February 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:OAc

[Ano15-73] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(5):957, May 2015. CO-DEN ???? ISSN 2329-9290. Anonymous:2015:OAd

[Ano15-74] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (6):1100, June 2015. CO-DEN ???? ISSN 2329-9290. Anonymous:2015:TCa

[Ano15-75] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):1–2, January 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCb

[Ano15-76] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):3–4, January 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCc

[Ano15-77] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(2):223–224, February 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCd

[Ano15-78] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(2):225–226, February 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCe

[Ano15-79] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(3):425–426, March 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCf

[Ano15-80] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(4):601–602, April 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCg

[Ano15-81] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(4):603–604, April 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCh

[Ano15-82] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(5):813–814, May 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCi

[Ano15-83] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(5):815–816, May 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCj

[Ano15-84] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(6):953–954, June 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCk

[Ano15-85] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(6):955–956, June 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCl

[Ano15-86] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(7):1101–1102, July 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCm

[Ano15-87] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(7):1103–1104, July 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCn

[Ano15-88] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):1245–1246, August 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCo

[Ano15-89] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):1247–1248, August 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCp

[Ano15-90] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(9):1385–1386, September 2015. CODEN ???? ISSN 2329-9290. Anonymous:2015:TCq

[Ano15-91] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(9):1387–1388, September 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:TCr

[Ano15-92] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1535–1536, October 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2015:TCs

[Ano15-93] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1537–1538, October 2015. CODEN ???? ISSN 2329-9290.

Anonymous:2016:BPa

[Ano16a] Anonymous. [Blank page]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (1):C4, January 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:BPb

[Ano16b] Anonymous. [Blank page]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (2):C4, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:BPc

[Ano16c] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (5):C4, May 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:BPd

[Ano16d] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (6):C4, June 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:BPe

[Ano16e] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (9):C4, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:BPf

[Ano16f] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (10):C4, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:ITMa

[Ano16g] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (9):1679–1680, ???? 2016. CODEN ???? ISSN 2329-

Anonymous:2016:ITMb

[Ano16h]  Anonymous.  IEEE

Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1878–1879, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATn

[Ano16i] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(9): 1677–1678, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATo

[Ano16j] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(10): 1876–1877, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:ECIa

[Ano16k] Anonymous. EDICS categories for IEEE/ACM Transactions on Audio, Speech, and Language Processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(3):608– 609, March 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:ECIb

[Ano16l] Anonymous. EDICS categories for IEEE/ACM Transactions on Audio, Speech, and Language Processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(4):823– 824, April 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:FCa

[Ano16m]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):C1, January 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:FCb

[Ano16n]  Anonymous.  [Front

cover]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(2):C1, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:FCc

[Ano16o] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (3):C1, March 2016. CO-DEN ???? ISSN 2329-

Anonymous:2016:FCd

[Ano16p] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (4):C1, April 2016. CO-DEN ???? ISSN 2329-9290. Anonymous:2016:FCe

[Ano16q] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (5):C1, May 2016. CO-DEN ???? ISSN 2329-9290. Anonymous:2016:FCf

[Ano16r] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (6):C1, June 2016. CO-DEN ???? ISSN 2329-9290. Anonymous:2016:FCg

[Ano16s] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (9):C1, ???? 2016. CO-DEN ???? ISSN 2329-9290. Anonymous:2016:FCh

[Ano16t] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (10):C1, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:IPEa

[Ano16u] Anonymous. IEEE Power Electronics Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(5):C3, May 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IPEb

[Ano16v]  Anonymous.  IEEE

Power Electronics Society information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(6): C3, June 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:ISPc

[Ano16w] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (9):C2, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:ISPd

[Ano16x] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (9):C3, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:ISPe

[Ano16y] Anonymous. IEEE Signal Processing Society.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (10):C2, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:ISPf

[Ano16z] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (10):C3, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:ISPa

[Ano16-27] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):C3, January 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:ISPb

[Ano16-28] Anonymous. IEEE Signal Processing Society information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(2):C3, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATe

[Ano16-29]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): C2, March 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATf

[Ano16-30]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): C3, March 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATg

[Ano16-31]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): C2, April 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATa

[Ano16-32] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(1): 211–212, January 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATc

[Ano16-33] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(2): 400–401, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:IATh

[Ano16-34] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(5): 990–991, May 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:IATk

[Ano16-35] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(6): 1155–1156, June 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATi

[Ano16-36]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(5): 992–993, May 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:IATl

[Ano16-37] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(6):1157– 1158, June 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATb

[Ano16-38] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):C2, January 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATd

[Ano16-39] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(2):C2, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATj

[Ano16-40] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(5):C2, May 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IATm

[Ano16-41]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing publication information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(6): C2, June 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IAa

[Ano16-42]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(1): 213–214, January 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:IAb

[Ano16-43]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(2): 402–403, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:IAc

[Ano16-44]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): 610–611, March 2016.

CODEN ???? ISSN 2329-9290. Anonymous:2016:IAd

[Ano16-45]  Anonymous.  Informa-

tion for authors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 825–826, April 2016. CO-DEN ???? ISSN 2329-9290. Anonymous:2016:IICa

[Ano16-46] Anonymous. Introducing IEEE Collabratec. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): 613, March 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:IICb

[Ano16-47] Anonymous. Introducing IEEE Collabratec. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 828, April 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:IIPa

[Ano16-48] Anonymous. Introducing the IEEE PES Resource Center. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(9): 1681, ???? 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:IIPb

[Ano16-49] Anonymous. Introducing the IEEE PES Resource Center. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(10): 1880, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:SIBa

[Ano16-50] Anonymous. Special issue on biosignal-based spoken communication. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(5): 995, May 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:SIBb

[Ano16-51] Anonymous. Special issue on biosignal-based spoken communication. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(6): 1160, June 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:SISa

[Ano16-52] Anonymous. Special issue on sound scene and event analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(1): 215, January 2016. CO-DEN ???? ISSN 2329-9290.

Anonymous:2016:SISb

[Ano16-53] Anonymous. Special issue on sound scene and event analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(2):

404, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:SISc

[Ano16-54] Anonymous. Special issue on sound scene and event analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): 612, March 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:SISd

[Ano16-55] Anonymous. Special issue on sound scene and event analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 827, April 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:SISe

[Ano16-56] Anonymous. Special issue on sound scene and event analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(5): 994, May 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:SISf

[Ano16-57] Anonymous. Special issue on sound scene and event analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(6): 1159, June 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:TCa

[Ano16-58] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):1–2, January 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCb

[Ano16-59] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):3–4, January 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCc

[Ano16-60] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(2):211–212, ???? 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCd

[Ano16-61] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(2):213–214, ???? 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCe

[Ano16-62] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(3):405–406, March 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:TCf

[Ano16-63] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(3):407–408, March 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCg

[Ano16-64] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(4):608–609, April 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCh

[Ano16-65] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(4):610–611, April 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCi

[Ano16-66] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(5):829–830, May 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCj

[Ano16-67] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(5):831–832, May 2016. CODEN ???? ISSN 2329-

Anonymous:2016:TCk

[Ano16-68] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(6):990–991, June 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCl

[Ano16-69] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(6):992–993, June 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCm

[Ano16-70] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(9):1477–1478, ???? 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCn

[Ano16-71] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(9):1479–1480, ???? 2016. CODEN ???? ISSN 2329-9290. Anonymous:2016:TCo

[Ano16-72] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1677–1678, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:TCp

[Ano16-73] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1679–1680, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:TCq

[Ano16-74] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):1881–1882, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2016:TCr

[Ano16-75] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):1883–1884, ???? 2016. CODEN ???? ISSN 2329-9290.

Anonymous:2017:SII

[Ano17a] Anonymous. 2017 subject index IEEE Transactions on Applied Superconductivity vol. 27. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12): 2463–2488, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:BPa

[Ano17b] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):C4, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:BPb

[Ano17c] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):B1165–B1166, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:BPc

[Ano17d] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):C4, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:BPd

[Ano17e] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (6):C4, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:BPe

[Ano17f] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (7):C4, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:BPf

[Ano17g] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):C4, ???? 2017. CO-DEN ???? ISSN 2329-9290. Anonymous:2017:BPg

[Ano17h] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (9):C4, ???? 2017. CO-DEN ???? ISSN 2329-9290. Anonymous:2017:BPh

[Ano17i] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (10):C4, ???? 2017. CO-DEN ???? ISSN 2329-9290. Anonymous:2017:BPi

[Ano17j] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (11):C4, ???? 2017. CO-DEN ???? ISSN 2329-9290. Anonymous:2017:BPj

[Ano17k] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (12):C4, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ITMb

[Ano17l] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (7):1564–1565, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:ITMc

[Ano17m] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):1733–1734, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:ITMd

[Ano17n] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (9):1861–1862, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:ITMe

[Ano17o]  Anonymous.  IEEE

Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(10):2038–2039, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:ITMf

[Ano17p]  Anonymous.  IEEE

Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2254–2255, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:ITMg

[Ano17q]  Anonymous.  IEEE

Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(12):2460–2461, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IATc

[Ano17r] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(5): 1167–1168, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IATe

[Ano17s] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(6): 1398–1399, ???? 2017. CODEN ???? ISSN 2329-

Anonymous:2017:IATg

[Ano17t] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(7): 1562–1563, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IATh

[Ano17u] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(8): 1731–1732, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IATi

[Ano17v] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(9): 1859–1860, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IATj

[Ano17w] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(10): 2036–2037, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IATk

[Ano17x] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(11): 2252–2253, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IATl

[Ano17y] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12): 2458–2459, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IATd

[Ano17z] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):1169–1170, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IATf

[Ano17-27] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (6):1400–1401, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:FCa

[Ano17-28] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (2):C1, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:FCb

[Ano17-29] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):C1, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:FCc

[Ano17-30] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):C1, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:FCd

[Ano17-31] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (6):C1, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:FCe

[Ano17-32] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (7):C1, ???? 2017. CO-DEN ???? ISSN 2329-9290. Anonymous:2017:FCf

[Ano17-33] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):C1, ???? 2017. CO-DEN ???? ISSN 2329-9290. Anonymous:2017:FCg

[Ano17-34] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (9):C1, ???? 2017. CO-DEN ???? ISSN 2329-9290. Anonymous:2017:FCh

[Ano17-35] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (10):C1, ???? 2017. CO-DEN ???? ISSN 2329-9290. Anonymous:2017:FCi

[Ano17-36] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (11):C1, ???? 2017. CO-DEN ???? ISSN 2329-

Anonymous:2017:FCj

[Ano17-37] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (12):C1, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPa

[Ano17-38] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (2):C2, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPb

[Ano17-39] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):C2, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPc

[Ano17-40] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):C3, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPd

[Ano17-41] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):C2, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPe

[Ano17-42] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):C3, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPf

[Ano17-43] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (6):C2, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPg

[Ano17-44] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (6):C3, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPh

[Ano17-45] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (7):C2, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPi

[Ano17-46] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (7):C3, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPj

[Ano17-47] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):C2, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPk

[Ano17-48] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):C3, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPl

[Ano17-49] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (9):C2, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPm

[Ano17-50] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (9):C3, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPn

[Ano17-51] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (10):C2, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPo

[Ano17-52] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (10):C3, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPp

[Ano17-53] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (11):C2, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPq

[Ano17-54] Anonymous. IEEE Signal Processing Society.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (11):C3, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPr

[Ano17-55] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (12):C2, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ISPs

[Ano17-56] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (12):C3, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:ITA

[Ano17-57] Anonymous. IEEE transactions on audio, speech, and language processing information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):922–923, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:ITMa

[Ano17-58] Anonymous. IEEE transactions on multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (2):450–451, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IATa

[Ano17-59] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(2): 448–449, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:IATb

[Ano17-60] Anonymous. IEEE/ ACM transactions on audio, speech, and language processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(4): 920–921, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Anonymous:2017:IICa

[Ano17-61] Anonymous. Introducing IEEE Collabratec. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(2): 452, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IICb

[Ano17-62] Anonymous. Introducing IEEE Collabratec. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(4): 924, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IICc

[Ano17-63] Anonymous. Introducing IEEE Collabratec. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(5): 1172, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IICd

[Ano17-64] Anonymous. Introducing IEEE Collabratec. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(6): 1403, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IICe

[Ano17-65] Anonymous. Introducing IEEE Collabratec. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(7): 1567, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:IICf

[Ano17-66] Anonymous. Introducing IEEE Collabratec. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(8): 1736, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:LR

[Ano17-67] Anonymous. List of reviewers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(12):2454– 2457, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:MGMa

[Ano17-68]  Anonymous.  Member

Get-A-Member (MGM) Program. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(5):1173, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:MGMb

[Ano17-69]  Anonymous.  Member

Get-A-Member (MGM) Program. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(6):1404, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:OAa

[Ano17-70] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(5):1171, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:OAb

[Ano17-71] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(6):1402, ???? 2017. CODEN ???? ISSN 2329-

Anonymous:2017:OAc

[Ano17-72] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(7):1566, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:OAd

[Ano17-73] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(8):1735, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:OAe

[Ano17-74] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(9):1863, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:OAf

[Ano17-75] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(10):2040, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:OAg

[Ano17-76] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2256, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:OAh

[Ano17-77] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(12):2462, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCa

[Ano17-78] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(2):222–223, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCb

[Ano17-79] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(2):224–225, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCc

[Ano17-80] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(3):3–4, ???? 2017. CO-DEN ???? ISSN 2329-9290. Anonymous:2017:TCd

[Ano17-81] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(3):3–4, ???? 2017. CO-DEN ???? ISSN 2329-

Anonymous:2017:TCe

[Ano17-82] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(3):3–4, ???? 2017. CO-DEN ???? ISSN 2329-9290. Anonymous:2017:TCf

[Ano17-83] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(3):3–4, ???? 2017. CO-DEN ???? ISSN 2329-9290. Anonymous:2017:TCg

[Ano17-84] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(4):688–689, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCh

[Ano17-85] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(4):690–691, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCi

[Ano17-86] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(5):925–926, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:TCj

[Ano17-87] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(5):927–928, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCk

[Ano17-88] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(6):1167–1168, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCl

[Ano17-89] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(7):1405–1406, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCm

[Ano17-90] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(8):1562–1563, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCn

[Ano17-91] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(8):1564–1565, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:TCo

[Ano17-92] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(9):1737–1738, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCp

[Ano17-93] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(9):1739–1740, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCq

[Ano17-94] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(10):1859–1860, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCr

[Ano17-95] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(10):1861–1862, ???? 2017. CODEN ???? ISSN 2329-9290. Anonymous:2017:TCs

[Ano17-96] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2041–2042, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:TCt

[Ano17-97] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2043–2044, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:TCu

[Ano17-98] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(12):2252–2253, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2017:TCEa

[Ano17-99] Anonymous. Table of contents edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(7): 1407–1408, ???? 2017. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IIA

[Ano18a] Anonymous. 2018 index IEEE/ACM Transactions on Audio, Speech, and Language Processing vol.

26. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2503–2528, December 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:BPa

[Ano18b] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):C4, March 2018. CO-DEN ???? ISSN 2329-9290. Anonymous:2018:BPb

[Ano18c] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (4):C4, April 2018. CO-DEN ???? ISSN 2329-9290. Anonymous:2018:BPc

[Ano18d] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (5):C4, May 2018. CO-DEN ???? ISSN 2329-9290. Anonymous:2018:BPd

[Ano18e] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (6):C4, June 2018. CO-DEN ???? ISSN 2329-9290. Anonymous:2018:BPe

[Ano18f] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (7):C4, July 2018. CO-DEN ???? ISSN 2329-9290. Anonymous:2018:BPf

[Ano18g] Anonymous. Blank page. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (8):C4, August 2018. CO-DEN ???? ISSN 2329-9290. Anonymous:2018:BPg

[Ano18h] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (9):C4, September 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:BPh

[Ano18i] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):C4, October 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:BPi

[Ano18j] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):C4, November 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:BPj

[Ano18k] Anonymous. Blank page.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):C4, December 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:ITMa

[Ano18l] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):698–699, March 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ITMb

[Ano18m] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (4):849–850, April 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ITMc

[Ano18n] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (5):1027–1028, May 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ITMd

[Ano18o] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (6):1171–1172, June 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ITMe

[Ano18p] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (7):1305–1306, July 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ITMf

[Ano18q] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (8):1451–1452, August 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ITMg

[Ano18r] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (9):1700–1702, September 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ITMh

[Ano18s] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):1942–1944, October 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ITMi

[Ano18t] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):2253–2255, November 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ITMj

[Ano18u] Anonymous. IEEE Transactions on Multimedia information for authors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2499–2501, December 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IATa

[Ano18v] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(3): 696–697, March 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IATb

[Ano18w] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(4): 847–848, April 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:IATc

[Ano18x] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(5): 1025–1026, May 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IATd

[Ano18y] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(6): 1169–1170, June 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IATe

[Ano18z] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(7): 1303–1304, July 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IATf

[Ano18-27] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(8): 1449–1450, August 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IATg

[Ano18-28]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (9):1698–1699, September 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IATh

[Ano18-29] Anonymous. IEEE/ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(10): 1940–1941, October 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IATi

[Ano18-30]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):2251–2252, November 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IATj

[Ano18-31]  Anonymous.  IEEE/

ACM Transactions on Audio, Speech, and Language Processing edics. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2497–2498, December 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:FE

[Ano18-32] Anonymous. Farewell editorial. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2489, December 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:FCa

[Ano18-33] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):C1, March 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:FCb

[Ano18-34] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (4):C1, April 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:FCc

[Ano18-35] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (5):C1, May 2018. CO-DEN ???? ISSN 2329-

Anonymous:2018:FCd

[Ano18-36] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (6):C1, June 2018. CO-DEN ???? ISSN 2329-9290. Anonymous:2018:FCe

[Ano18-37] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (7):C1, July 2018. CO-DEN ???? ISSN 2329-9290. Anonymous:2018:FCf

[Ano18-38] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (8):C1, August 2018. CO-DEN ???? ISSN 2329-9290. Anonymous:2018:FCg

[Ano18-39] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (9):C1, September 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:FCh

[Ano18-40] Anonymous. Front cover. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):C1, October 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:FCi

[Ano18-41] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):C1, November 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:FCj

[Ano18-42] Anonymous. Front cover.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):C1, December 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IOA

[Ano18-43] Anonymous. IEEE open access publishing. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(12): 2502, December 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ISPa

[Ano18-44] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):C2, March 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPb

[Ano18-45] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):C3, March 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPc

[Ano18-46] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (4):C2, April 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPd

[Ano18-47] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (4):C3, April 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPe

[Ano18-48] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (5):C2, May 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPf

[Ano18-49] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (5):C3, May 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPg

[Ano18-50] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (6):C2, June 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPh

[Ano18-51] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (6):C3, June 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPi

[Ano18-52] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (7):C2, July 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPj

[Ano18-53] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (7):C3, July 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPk

[Ano18-54] Anonymous. IEEE Signal Processing Society.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (8):C2, August 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPl

[Ano18-55] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (8):C3, August 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:ISPm

[Ano18-56] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (9):C2, September 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ISPn

[Ano18-57] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (9):C3, September 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ISPo

[Ano18-58] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):C2, October 2018.

CODEN ???? ISSN 2329-9290.

Anonymous:2018:ISPp

[Ano18-59] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):C3, October 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ISPq

[Ano18-60] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):C2, November 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ISPr

[Ano18-61] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):C3, November 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:ISPu

[Ano18-62] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):C2, December 2018. CODEN ???? ISSN 2329-

Anonymous:2018:ISPv

[Ano18-63] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):C3, December 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IICa

[Ano18-64] Anonymous. Introducing IEEE Collabratec. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(3): 701, March 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:IICb

[Ano18-65] Anonymous. Introducing IEEE Collabratec. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(4): 852, April 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:LR

[Ano18-66] Anonymous. List of reviewers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2490– 2496, December 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:OAa

[Ano18-67] Anonymous. Open access. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):700, March 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:OAb

[Ano18-68] Anonymous. Open access.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (4):851, April 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:TCa

[Ano18-69] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(1):1–2, ???? 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:TCb

[Ano18-70] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(3):457–458, March 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:TCc

[Ano18-71] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(4):696–697, April 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:TCd

[Ano18-72] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(5):853–854, May 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:TCe

[Ano18-73] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(6):1025– 1026, June 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:TCf

[Ano18-74] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(7):1173–1174, July 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:TCg

[Ano18-75] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(8):1303–1304, August 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:TCh

[Ano18-76] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1453–1454, September 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:TCi

[Ano18-77] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(10):1698–1699, October 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:TCj

[Ano18-78] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):1945–1946, November 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:TCk

[Ano18-79] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2251–2252, December 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:TCEa

[Ano18-80] Anonymous. Table of contents [Edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(1):3– 4, ???? 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:TCEb

[Ano18-81] Anonymous. Table of contents [edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(3): 459–460, March 2018. CODEN ???? ISSN 2329-9290. Anonymous:2018:TCEc

[Ano18-82] Anonymous. Table of contents [edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(4): 698–699, April 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:TCEd

[Ano18-83] Anonymous. Table of contents [edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(5): 855–856, May 2018. CO-DEN ???? ISSN 2329-9290.

Anonymous:2018:TCEe

[Ano18-84] Anonymous. Table of contents [edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(6): 1027–1028, June 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:TCEf

[Ano18-85] Anonymous. Table of contents [edics]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(7):1175– 1176, July 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:TCEg

[Ano18-86] Anonymous. Table of contents [edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(8): 1305–1306, August 2018.

CODEN ???? ISSN 2329-9290.

Anonymous:2018:TCEh

[Ano18-87] Anonymous. Table of contents [edics]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1455– 1456, September 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:TCEi

[Ano18-88] Anonymous. Table of contents [edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(10): 1700–1701, October 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:TCEj

[Ano18-89] Anonymous. Table of contents [edics]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):1947– 1948, November 2018. CODEN ???? ISSN 2329-9290.

Anonymous:2018:TCEk

[Ano18-90] Anonymous. Table of contents [edics]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2253– 2254, December 2018. CODEN ???? ISSN 2329-

Anonymous:2019:AAI

[Ano19a]  Anonymous.  Article

awards for the IEEE/ ACM Transactions on Audio, Speech, and Language Processing. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(5): 988, May 2019. CODEN ???? ISSN 2329-9290.

Anonymous:2019:BP

[Ano19b] Anonymous. [Blank page]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (1):B4, January 2019. CODEN ???? ISSN 2329-9290.

Anonymous:2019:ENA

[Ano19c] Anonymous. Erratum for nonlinear audio systems identification through audio input Gaussianization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(1):240, January 2019. CODEN ???? ISSN 2329-9290.

Anonymous:2019:ISPa

[Ano19d] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (1):C2, January 2019. CODEN ???? ISSN 2329-9290.

Anonymous:2019:ISPb

[Ano19e] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (1):C3, January 2019. CODEN ???? ISSN 2329-9290.

Anonymous:2019:ISPc

[Ano19f] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (2):C2, February 2019. CODEN ???? ISSN 2329-9290.

Anonymous:2019:ISPd

[Ano19g] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (2):C3, February 2019. CODEN ???? ISSN 2329-9290.

Anonymous:2019:ISPe

[Ano19h] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (3):C2, March 2019. CO-DEN ???? ISSN 2329-9290.

Anonymous:2019:ISPf

[Ano19i] Anonymous. IEEE Signal Processing Society.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (3):C3, March 2019. CO-DEN ???? ISSN 2329-9290.

Anonymous:2019:ISPg

[Ano19j] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (4):C2, April 2019. CO-DEN ???? ISSN 2329-9290.

Anonymous:2019:ISPh

[Ano19k] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (4):C3, April 2019. CO-DEN ???? ISSN 2329-9290.

Anonymous:2019:ISPi

[Ano19l] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (5):C2, May 2019. CO-DEN ???? ISSN 2329-9290.

Anonymous:2019:ISPj

[Ano19m] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (5):C3, May 2019. CO-DEN ???? ISSN 2329-9290.

Anonymous:2019:ISPk

[Ano19n] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (6):C2, June 2019. CO-DEN ???? ISSN 2329-9290.

Anonymous:2019:ISPl

[Ano19o] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (6):C3, June 2019. CO-DEN ???? ISSN 2329-9290.

Anonymous:2019:ISPm

[Ano19p] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (7):C2, July 2019. CO-DEN ???? ISSN 2329-9290.

Anonymous:2019:ISPn

[Ano19q] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (7):C3, July 2019. CO-DEN ???? ISSN 2329-

Anonymous:2019:ISPo

[Ano19r] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (8):C2, August 2019. CO-DEN ???? ISSN 2329-9290. Anonymous:2019:ISPp

[Ano19s] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (8):C3, August 2019. CO-DEN ???? ISSN 2329-9290. Anonymous:2019:IEI

[Ano19t]  Anonymous.  Inaugu-

ral editorial innovations in an era of ubiquitous audio, speech, and language processing. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1):5– 6, January 2019. CODEN ???? ISSN 2329-9290. Anonymous:2019:TCa

[Ano19u] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(1):C1–1, January 2019. CODEN ???? ISSN 2329-9290. Anonymous:2019:TCb

[Ano19v] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(2):C1–241, February 2019. CODEN ???? ISSN 2329-9290. Anonymous:2019:TCd

[Ano19w] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(3):C1–469, March 2019. CODEN ???? ISSN 2329-9290. Anonymous:2019:TCg

[Ano19x] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(4):C1–660, April 2019. CODEN ???? ISSN 2329-9290. Anonymous:2019:TCi

[Ano19y] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(5):C1–878, May 2019. CODEN ???? ISSN 2329-9290. Anonymous:2019:TCj

[Ano19z] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(5):879–880, May 2019. CODEN ???? ISSN 2329-9290. Anonymous:2019:TCk

[Ano19-27] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(6):C1–989, June 2019. CODEN ???? ISSN 2329-9290. Anonymous:2019:TCl

[Ano19-28] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(7):C1–1109, July 2019. CODEN ???? ISSN 2329-9290. Anonymous:2019:TCn

[Ano19-29] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(8):C1–1213, August 2019. CODEN ???? ISSN 2329-9290. Anonymous:2019:TCo

[Ano19-30] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(8):1214–1215, August 2019. CODEN ???? ISSN 2329-9290.

Anonymous:2019:TCEa

[Ano19-31] Anonymous. Table of contents [edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1):2– 3, January 2019. CODEN ???? ISSN 2329-9290. Anonymous:2019:TCEb

[Ano19-32] Anonymous. Table of contents (EDICS). IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(6): 990–991, June 2019. CO-DEN ???? ISSN 2329-9290.

Anonymous:2019:TCc

[Ano19-33] Anonymous. Table of Contents[Edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(2): 242–243, February 2019. CODEN ???? ISSN 2329-9290.

Anonymous:2019:TCf

[Ano19-34] Anonymous. Table of Contents[Edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(3): 470–471, March 2019. CODEN ???? ISSN 2329-9290.

Anonymous:2019:TCh

[Ano19-35] Anonymous. Table of Contents[Edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(4): 661–662, April 2019. CO-DEN ???? ISSN 2329-9290.

Anonymous:2019:TCm

[Ano19-36] Anonymous. Table of Contents[Edics]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(7): 1110–1111, July 2019.

CODEN ???? ISSN 2329-9290. Anonymous:2021:IIA

[Ano21a] Anonymous. 2021 Index IEEE/ACM Transactions on Audio, Speech, and Language Processing Vol.

29. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3718–3760, 2021. CODEN ???? ISSN 2329-9290. Anonymous:2021:ISP

[Ano21b] Anonymous. IEEE Signal Processing Society. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):c2, 2021. CODEN ???? ISSN 2329-9290. Anonymous:2021:TCa

[Ano21c] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):c1–ix, 2021. CO-DEN ???? ISSN 2329-9290. Anonymous:2021:TCb

[Ano21d] Anonymous. Table of contents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):x–xx, 2021. CO-DEN ???? ISSN 2329-9290. Anonymous:2022:ISP

[Ano22] Anonymous. IEEE Signal Processing Society.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):C2, 2022. CODEN ???? ISSN 2329-9290.

Aichinger:2021:SAS

[AP21] Philipp Aichinger and Franz Pernkopf. Synthesis and analysis-by-synthesis of modulated diplophonic glottal area waveforms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):914–926, 2021. CO-DEN ???? ISSN 2329-9290.

Alameda-Pineda:2014:GAS

[APH14] X. Alameda-Pineda and R. Horaud. A geometric approach to sound source localization from timedelay estimates. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(6): 1082–1095, June 2014. CODEN ???? ISSN 2329-9290.

Arend:2023:MCT

[APWB23] Johannes M. Arend, Christoph Pörschmann, Stefan Weinzierl, and Fabian Brinkmann. Magnitude-corrected and time-aligned interpolation of head-related transfer functions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3783– 3799, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3313908.

Alon:2016:BOA

[AR16] D. L. Alon and B. Rafaely. Beamforming with optimal aliasing cancellation in spherical microphone arrays. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(1): 196–210, January 2016. CODEN ???? ISSN 2329-9290.

Airaksinen:2014:QCP

[ARSA14] M. Airaksinen, T. Raitio, B. Story, and P. Alku. Quasi closed phase glottal inverse filtering analysis with weighted linear prediction. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(3): 596–607, March 2014. CODEN ???? ISSN 2329-9290.

Abdul-Rauf:2016:EUI

[ARSLN16] Sadaf Abdul-Rauf, Holger Schwenk, Patrik Lambert, and Mohammad Nawaz. Empirical use of information retrieval to build synthetic data for SMT domain adaptation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 745–754, April 2016. CO-DEN ???? ISSN 2329-9290.

Adeli:2016:FBI

[ARW^{+}16] Mohammad Adeli, Jean Rouat, Sean Wood, Stéphane Molotchnikoff, and Eric Plourde. A flexible bio-inspired hierarchical model for analyzing musical timbre. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(5): 875–889, May 2016. CO-DEN ???? ISSN 2329-9290.

Aragonda:2015:DNS

[AS15] H. Aragonda and C. S.

Seelamantula. Demodulation of narrowband speech spectrograms using the Riesz transform. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1824–1834, November 2015. CODEN ???? ISSN 2329-9290.

Abesser:2017:ICM

[AS17a] Jakob Abeßer and Gerald Schuller. Instrumentcentered music transcription of solo bass guitar recordings. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(9): 1741–1750, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7921692/.

Alku:2017:LPM

[AS17b] Paavo Alku and Rahim Saeidi. The linear predictive modeling of speech from higher-lag autocorrelation coefficients applied to noise-robust speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(8):1606– 1617, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7924356/.

Ayana:2018:ZSC

[ASC^{+}18] Ayana, S. Shen, Y. Chen, C. Yang, Z. Liu, and M. Sun. Zero-shot crosslingual neural headline generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2319– 2327, December 2018. CODEN ???? ISSN 2329-9290.

An:2022:DSS

[ASX22] Xiaochun An, Frank K. Soong, and Lei Xie. Disentangling style and speaker attributes for TTS style transfer. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 646–658, 2022. CODEN ???? ISSN 2329-9290.

Aihara:2016:MNN

[ATA16] Ryo Aihara, Tetsuya Takiguchi, and Yasuo Ariki. Multiple nonnegative matrix factorization for many-to-many voice conversion. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(7): 1175–1184, July 2016. CODEN ???? ISSN 2329-9290.

Anderson:2015:SRF

[ATP15] C. A. Anderson, P. D.

Teal, and M. A. Poletti. Spatially robust far-field beamforming using the von Mises(–Fisher) distribution. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2189– 2197, December 2015. CODEN ???? ISSN 2329-9290.

Anderson:2016:SCR

[ATP16] C. A. Anderson, P. D.

Teal, and M. A. Poletti. Spatial correlation of radial Gaussian and uniform spherical volume near-field source distributions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):143–150, January 2016. CODEN ???? ISSN 2329-9290.

Ahmed:2021:NAI

[ATR^{+}21] Shakeel Ahmed, Muhammad Tufail, Muhammad Rehan, Tanveer Abbas, and Amna Majid. A novel approach for improved noise reduction performance in feed-forward active noise control systems with (loudspeaker) saturation non-linearity in the secondary path. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):187–197, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3039607.

Abidin:2018:SAU

[ATS18] S. Abidin, R. Togneri, and F. Sohel. Spectrotemporal analysis using local binary pattern variants for acoustic scene classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):2112–2121, November 2018. CODEN ???? ISSN 2329-9290.

Adiloglu:2016:VBI

[AV16] Kamil Adiloğlu and Emmanuel Vincent. Variational Bayesian inference for source separation and robust feature extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1746–1758, ????

2016. CODEN ???? ISSN 2329-9290.

Adel:2015:SSF

[AVK^{+}15] H. Adel, Ngoc Thang Vu, K. Kirchhoff, D. Telaar, and T. Schultz. Syntactic and semantic features for code-switching factored language models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (3):431–440, March 2015. CODEN ???? ISSN 2329-9290.

Ali:2019:IPE

[AVM19] R. Ali, T. Van Waterschoot, and M. Moonen. Integration of a priori and estimated constraints into an MVDR beamformer for speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2288–2300, December 2019. CODEN ???? ISSN 2329-9290.

Aneeja:2015:SFF

[AY15] G. Aneeja and B. Yegnanarayana. Single frequency filtering approach for discriminating speech and nonspeech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(4): 705–717, April 2015. CO-DEN ???? ISSN 2329-

Aneeja:2017:EFF

[AY17] G. Aneeja and B. Yegnanarayana. Extraction of fundamental frequency from degraded speech using temporal envelopes at high SNR frequencies. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):829–838, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7847347/.

Audhkhasi:2014:TAD

[AZGN14] K. Audhkhasi, A. M. Zavou, P. G. Georgiou, and S. S. Narayanan. Theoretical analysis of diversity in an ensemble of automatic speech recognition systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(3):711–726, March 2014. CODEN ???? ISSN 2329-9290.

Abdelaziz:2015:LDS

[AZK15] A. H. Abdelaziz, S. Zeiler, and D. Kolossa. Learning dynamic stream weights for coupled-HMM-based audio-visual speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(5):863– 876, May 2015. CODEN ???? ISSN 2329-9290.

Bahne:2016:OSL

[BA16] A. Bahne and A. Ahlen.

Optimizing the similarity of loudspeaker-room responses in multiple listening positions. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(2): 340–353, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Bao:2019:NRM

[BA19] F. Bao and W. H. Abdulla. A new ratio mask representation for CASAbased speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(1):7–19, January 2019. CODEN ???? ISSN 2329-9290.

Buchris:2019:ISS

[BABC19] Y. Buchris, A. Amar, J. Benesty, and I. Cohen. Incoherent synthesis of sparse arrays for frequency-invariant beamforming. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(3):482– 495, March 2019. CODEN ???? ISSN 2329-9290.

Birnie:2020:RAS

[BAS20] Lachlan I. Birnie, Thushara D.

Abhayapala, and Prasanga N. Samarasinghe. Reflection assisted sound source localization through a harmonic domain MU-SIC framework. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 279–293, January 2020. CODEN ???? ISSN 2329-9290. URL https:

//dl.acm.org/doi/abs/
10.1109/TASLP.2019.2953000.

Birnie:2021:MSS

[BATS21] Lachlan Birnie, Thushara Abhayapala, Vladimir Tourbabin, and Prasanga Samarasinghe. Mixed source sound field translation for virtual binaural application with perceptual validation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1188–1203, 2021. CODEN ???? ISSN 2329-9290.

Bayram:2015:MAD

[Bay15] I. Bayram. A multichannel audio denoising formulation based on spectral sparsity. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2272– 2285, December 2015. CODEN ???? ISSN 2329-9290.

Borgstrom:2021:SEA

[BB21] Bengt J. Borgström and Michael S. Brandstein. Speech Enhancement via Attention Masking Network (SEAMNET): an end-to-end system for joint suppression of noise and reverberation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):515–526, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3043655.

Borra:2019:ULA

[BBAS19] F. Borra, A. Bernardini, F. Antonacci, and A. Sarti. Uniform linear arrays of first-order steerable differential microphones. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):1906– 1918, December 2019. CODEN ???? ISSN 2329-9290.

Bentivogli:2016:EAM

[BBC^{+}16] L. Bentivogli, N. Bertoldi, M. Cettolo, M. Federico, M. Negri, and M. Turchi. On the evaluation of adaptive machine translation for human post-editing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (2):388–399, ???? 2016. CODEN ???? ISSN 2329-9290.

Bernardini:2021:WDN

[BBFS21] Alberto Bernardini, Enrico Bozzo, Federico Fontana, and Augusto Sarti. A wave digital Newton–Raphson method for virtual analog modeling of audio circuits with multiple one-port nonlinearities. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2162– 2173, 2021. CODEN ???? ISSN 2329-9290.

Berebi:2024:ADH

[BBHAR24] Or Berebi, Zamir Ben-Hur, David Lou Alon, and Boaz Rafaely. Analysis and design of headtracked compensation for bilateral ambisonics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 959–972, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3345140.

Bu:2018:DPF

[BBJ18] B. Bu, C. Bao, and M. Jia. Design of a planar firstorder loudspeaker array for global active noise control. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):2240–2250, November 2018. CODEN ???? ISSN 2329-9290.

Bahari:2017:BSR

[BBM17] Mohamad Hasan Bahari, Alexander Bertrand, and Marc Moonen. Blind sampling rate offset estimation for wireless acoustic sensor networks through weighted least-squares coherence drift estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (3):674–686, ???? 2017. CODEN ???? ISSN 2329-9290.

Bao:2017:DRB

[BC17] Yu Bao and Huawei Chen.

Design of robust broadband beamformers using worst-case performance optimization: a semidefinite programming approach. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(4):895– 907, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7864329/.

Bao:2018:CCP

[BC18] Y. Bao and H. Chen. A chance-constrained programming approach to the design of robust broadband beamformers with microphone mismatches. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2475–2488, December 2018. CODEN ???? ISSN 2329-9290.

Basaran:2015:PMB

[BCA15] D. Basaran, A. T. Cemgil, and E. Anarim. A probabilistic model-based approach for aligning multiple audio sequences. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (7):1160–1171, July 2015. CODEN ???? ISSN 2329-9290.

Berkun:2015:CBR

[BCB15] R. Berkun, I. Cohen, and J. Benesty. Combined beamformers for robust broadband regularized superdirective beamforming. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(5):877–886, May 2015. CODEN ???? ISSN 2329-9290.

Buchris:2018:FDD

[BCB18] Y. Buchris, I. Cohen, and J. Benesty. Frequencydomain design of asymmetric circular differential microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (4):760–773, April 2018. CODEN ???? ISSN 2329-9290.

Bai:2019:BAA

[BCCZ19] X. Bai, H. Cao, K. Chen, and T. Zhao. A bilingual adversarial autoencoder for unsupervised bilingual lexicon induction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(10):1639–1648, October 2019. CODEN ???? ISSN 2329-9290.

Brognaux:2016:HBS

[BD16] S. Brognaux and T. Drugman. HMM-based speech segmentation: Improvements of fully automatic approaches. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):5–15, January 2016. CODEN ???? ISSN 2329-9290.

Banchs:2015:AFM

[BDL15] R. E. Banchs, L. F.

D’Haro, and Haizhou Li. Adequacy–fluency metrics: Evaluating MT in the continuous space model framework. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(3): 472–482, March 2015. CODEN ???? ISSN 2329-9290.

Bahari:2014:NNF

[BDV^{+}14] M. H. Bahari, N. Dehak, H. Van hamme, L. Burget, A. M. Ali, and J. Glass. Non-negative factor analysis of Gaussian mixture model weight adaptation for language and dialect recognition.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (7):1117–1129, July 2014. CODEN ???? ISSN 2329-9290.

Bellur:2017:FDS

[BE17] Ashwin Bellur and Mounya Elhilali. Feedback-driven sensory mapping adaptation for robust speech activity detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(3): 481–492, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Backstrom:2018:FRD

[BF18] Tom Bäckstrøm and Johannes Fischer. Fast randomization for distributed low-bitrate coding of speech and audio. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(1):19–30, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8052578/.

Bhattacharjee:2021:NKP

[BG21] Sankha Subhra Bhattacharjee and Nithin V. George. Nearest Kronecker product decomposition based linear-in-the-parameters nonlinear filters. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2111–2122, 2021. CODEN ???? ISSN 2329-9290.

Bross:2023:TBM

[BG23] Avital Bross and Sharon Gannot. Training-based multiple source tracking using manifold-learning and recursive expectationmaximization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1124–1140, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3245414.

Bao:2018:QGD

[BGD^{+}18] J. Bao, Y. Gong, N. Duan, M. Zhou, and T. Zhao. Question generation with doubly adversarial nets. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):2230–2239, November 2018. CODEN ???? ISSN 2329-9290.

Belloch:2017:GBD

[BGQO^{+}17] Jose A. Belloch, Alberto Gonzalez, Enrique S. Quintana-Ortí, Miguel Ferrer, and Vesa Välimäki. GPU-based dynamic wave field synthesis using fractional delay filters and room compensation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(2):435–447, ???? 2017. CODEN ???? ISSN 2329-9290.

Braun:2018:LPB

[BH18] S. Braun and E. A. P. Habets. Linear predictionbased online dereverberation and noise reduction using alternating Kalman filters. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(6): 1119–1129, June 2018. CODEN ???? ISSN 2329-9290.

Bilbao:2019:DSW

[BH19] S. Bilbao and B. Hamilton. Directional sources in wave-based acoustic simulation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(2):415– 428, February 2019. CO-DEN ???? ISSN 2329-9290.

Ben-Hur:2019:ERS

[BHAMR19] Z. Ben-Hur, D. L. Alon, R. Mehra, and B. Rafaely. Efficient representation and sparse sampling of head-related transfer functions using phase-correction based on ear alignment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2249–2262, December 2019. CODEN ???? ISSN 2329-9290.

Ben-Hur:2021:BRB

[BHAMR21] Zamir Ben-Hur, David Lou Alon, Ravish Mehra, and Boaz Rafaely. Binaural reproduction based on bilateral ambisonics and ear-aligned HRTFs. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):901–913, 2021. CO-DEN ???? ISSN 2329-9290.

Bilbao:2016:FVT

[BHBS16] S. Bilbao, B. Hamilton, J. Botts, and L. Savioja. Finite volume time domain room acoustics simulation under general impedance boundary conditions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):161– 173, January 2016. CO-DEN ???? ISSN 2329-9290.

Barrett:2022:SRM

[BHH22] Liam Barrett, Junchao Hu, and Peter Howell. Systematic review of machine learning approaches for detecting developmental stuttering. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1160–1172, 2022. CODEN ???? ISSN 2329-9290.

Barhoush:2023:LDS

[BHP^{+}23] Mahdi Barhoush, Ahmed Hallawa, Arne Peine, Lukas Martin, and Anke Schmeink. Localizationdriven speech enhancement in noisy multispeaker hospital environments using deep learning and meta learning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 670–683, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3231700.

Behravan:2016:VMS

[BHS^{+}16] H. Behravan, V. Hautamaki, S. M. Siniscalchi, T. Kinnunen, and Chin-Hui Lee. i-vector modeling of speech attributes for automatic foreign accent recognition. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(1): 29–41, January 2016. CO-DEN ???? ISSN 2329-9290.

Bando:2018:SEB

[BIK^{+}18] Yoshiaki Bando, Katsutoshi Itoyama, Masashi Konyo, Satoshi Tadokoro, Kazuhiro Nakadai, Kazuyoshi Yoshii, Tatsuya Kawahara, and Hiroshi G. Okuno. Speech enhancement based on Bayesian low-rank and sparse decomposition of multichannel magnitude spectrograms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(2):215– 230, ???? 2018. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/8103886/.

Berghi:2024:LVS

[BJ24] Davide Berghi and Philip J. B. Jackson. Leveraging visual supervision for array-based active speaker detection and localization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):984– 995, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3346643.

Baas:2024:DGU

[BK24a] Matthew Baas and Herman Kamper. Disentanglement in a GAN for unconditional speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1324– 1335, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3359352.

Battal:2024:ACS

[BK24b] Omer Musa Battal and Aykut Koç. Automatic construction of sememe knowledge bases from machine readable dictionaries. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1023– 1035, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3347927.

Bates:2015:NSS

[BKK15] A. P. Bates, Z. Khalid, and R. A. Kennedy. Novel sampling scheme on the sphere for head-related transfer function measurements. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(6):1068– 1081, June 2015. CODEN ???? ISSN 2329-9290.

Braun:2018:ECL

[BKS^{+}18] S. Braun, A. Kuklasi ski, O. Schwartz, O. Thiergart, E. A. P. Habets, S. Gannot, S. Doclo, and J. Jensen. Evaluation and comparison of late reverberation power spectral density estimators. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (6):1056–1071, June 2018. CODEN ???? ISSN 2329-9290.

BenZaken:2024:NNB

[BKTR24] Orel Ben Zaken, Anurag Kumar, Vladimir Tourbabin, and Boaz Rafaely. Neuralnetwork-based directionof-arrival estimation for reverberant speech — the importance of energetic, temporal, and spatial information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1298– 1309, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3357037.

Benetos:2017:PSE

[BLLP17] Emmanouil Benetos, Grégoire Lafay, Mathieu Lagrange, and Mark D. Plumbley. Polyphonic sound event tracking using linear dynamical systems. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(6): 1266–1277, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7933041/.

Bai:2021:ITS

[BLZ21] Xuefeng Bai, Pengbo Liu, and Yue Zhang. Investigating typed syntactic dependencies for targeted sentiment classification using graph attention neural network. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):503–514, January

2021. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2020.3042009.

Blanco:2015:SLB

[BM15] E. Blanco and D. Moldovan. A semantic logic-based approach to determine textual similarity. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(4): 683–693, April 2015. CO-DEN ???? ISSN 2329-9290.

Balaraman:2021:DAD

[BM21] Vevake Balaraman and Bernardo Magnini. Domainaware dialogue state tracker for multi-domain dialogue systems. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 866–873, 2021. CODEN ???? ISSN 2329-9290.

Borsos:2023:ALM

[BMV^{+}23] Zalán Borsos, Raphaël Marinier, Damien Vincent, Eugene Kharitonov, Olivier Pietquin, Matt Sharifi, Dominik Roblek, Olivier Teboul, David Grangier, Marco Tagliasacchi, and Neil Zeghidour. AudioLM: a language modeling approach to audio generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2523– 2533, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3288409.

Borsky:2017:MNV

[BMVG17] Michal Borsky, Daryush D.

Mehta, Jarrad H. Van Stan, and Jon Gudnason. Modal and nonmodal voice quality classification using acoustic and electroglottographic features. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(12):2281–2291, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8114356/.

Biswas:2024:SGN

[BNA24] Ritujoy Biswas, Karan Nathwani, and Vinayak Abrol. Statistically guided near-end speech intelligibility improvement through voice transformation and transfer learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1445– 1456, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3359893.

Beerends:2020:SOA

[BNvdB^{+}20] John G. Beerends, Niels M. P. Neumann, Egon L. van den Broek, Anna Llagostera Casanovas, Jovana Torres Menendez, Christian Schmidmer, and Jens Berger. Subjective and objective assessment of full bandwidth speech quality. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):440– 449, January 2020. CO-DEN ???? ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2019.2957871.

Benaroya:2018:BLM

[BOL^{+}18] E. L. Benaroya, N. Obin, M. Liuni, A. Roebel, W. Raumel, and S. Argentieri. Binaural localization of multiple sound sources by non-negative tensor factorization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(6):1072– 1082, June 2018. CODEN ???? ISSN 2329-9290.

Borgstrom:2022:BEP

[Bor22] Bengt J. Borgström. Bayesian estimation of PLDA in the presence of noisy training labels, with applications to speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):414– 428, 2022. CODEN ???? ISSN 2329-9290.

Borgstrom:2023:GAC

[Bor23] Bengt J. Borgström. A generative approach to condition-aware score calibration for speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):891– 901, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3240663.

Beit-On:2023:WFS

[BOSR23] Hanan Beit-On, Tom Shlomo, and Boaz Rafaely. Weighted frequency smoothing for enhanced speaker localization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2090– 2099, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3277270.

Badeau:2014:MHR

[BP14] R. Badeau and M. D.

Plumbley. Multichannel high-resolution NMF for modeling convolutive mixtures of nonstationary signals in the time-frequency domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (11):1670–1680, November 2014. CODEN ???? ISSN 2329-9290.

Bhattacharjee:2023:CVO

[BPG23] Mrinmoy Bhattacharjee, S. R. M. Prasanna, and Prithwijit Guha. Clean vs. overlapped speechmusic detection using harmonic-percussive features and multi-task learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1–10, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3164199.

Barahona-Rios:2024:NCT

[BRC24] Adrián Barahona-Ríos and Tom Collins. Noise-BandNet: Controllable time-varying neural synthesis of sound effects using filterbanks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 1573–1585, 2024. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2024.3364616.

Bai:2015:LRS

[BRD15] Hequn Bai, G. Richard, and L. Daudet. Late reverberation synthesis: From radiance transfer to feedback delay networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2260–2271, December 2015. CODEN ???? ISSN 2329-9290.

Breebaart:2015:ESI

[Bre15] J. Breebaart. Evaluation of statistical inference tests applied to subjective audio quality data with small sample size. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (5):887–897, May 2015. CODEN ???? ISSN 2329-9290.

Botts:2014:SPP

[BS14] J. Botts and L. Savioja.

Spectral and pseudospectral properties of finite difference models used in audio and room acoustics. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(9):1403–1412, September 2014. CODEN ???? ISSN 2329-9290.

Byun:2021:MSS

[BS21] Jaeuk Byun and Jong Won Shin. Monaural speech separation using speaker embedding from preliminary separation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2753–2763, 2021. CODEN ???? ISSN 2329-9290.

Barkefors:2014:DAL

[BSB14] A. Barkefors, M. Sternad, and L.-J. Brannmark. Design and analysis of linear quadratic Gaussian feedforward controllers for active noise control. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1777–1791, December 2014. CODEN ???? ISSN 2329-9290.

Bisot:2017:FLM

[BSER17] Victor Bisot, Romain Serizel, Slim Essid, and Gaël Richard. Feature learning with matrix factorization applied to acoustic scene classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (6):1216–1229, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7933045/.

Bokaei:2015:LDS

[BSL15] M. H. Bokaei, H. Sameti, and Yang Liu. Linear discourse segmentation of multi-party meetings based on local and global information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):1879– 1891, November 2015. CODEN ???? ISSN 2329-9290.

Bokaei:2016:SMT

[BSL16] Mohammad Hadi Bokaei, Hossein Sameti, and Yang Liu. Summarizing meeting transcripts based on functional segmentation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1831–1841, ???? 2016. CODEN ???? ISSN 2329-9290.

Bell:2017:MLC

[BSR17] Peter Bell, Pawel Swietojanski, and Steve Renals. Multitask learning of context-dependent targets in deep neural network acoustic models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (2):238–247, ???? 2017. CODEN ???? ISSN 2329-9290.

Blandin:2023:CSA

[BSR^{+}23] Rémi Blandin, Simon Stone, Angélique Remacle, Vincent Didone, and Peter Birkholz. A comparative study of 3D and 1D acoustic simulations of the higher frequencies of speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3837– 3847, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313423.

Bohlender:2021:ETC

[BSTM21] Alexander Bohlender, Ann Spriet, Wouter Tirry, and Nilesh Madhu. Exploiting temporal context in CNN based multisource DOA estimation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1594–1608, 2021. CODEN ???? ISSN 2329-9290.

Bohlender:2024:SSS

[BSTM24] Alexander Bohlender, Ann Spriet, Wouter Tirry, and Nilesh Madhu. Spatially selective speaker separation using a DNN with a location dependent feature extraction. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 930–945, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3343605.

Boeddeker:2024:TSJ

[BSW^{+}24]  Christoph  Boeddeker,

Aswin Shanmugam Subramanian, Gordon Wichern, Reinhold Haeb-Umbach, and Jonathan Le Roux. TS-SEP: Joint diarization and separation conditioned on estimated speaker embeddings. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1185– 1197, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3350887.

Birkholz:2018:NIS

[BSWP18] P. Birkholz, S. Stone, K. Wolf, and D. Plettemeier. Non-invasive silent phoneme recognition using microwave signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2404–2411, December 2018. CODEN ???? ISSN 2329-9290.

Barbancho:2018:DBA

[BTBT18] I. Barbancho, G. Tzanetakis, A. M. Barbancho, and L. J. Tardón. Discrimination between ascending/descending pitch arpeggios. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):2194– 2203, November 2018. CODEN ???? ISSN 2329-9290.

Bao:2019:TGT

[BTD^{+}19] J. Bao, D. Tang, N. Duan, Z. Yan, M. Zhou, and T. Zhao. Text generation from tables. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(2): 311–320, February 2019. CODEN ???? ISSN 2329-9290.

Barker:2016:BSA

[BV16] Tom Barker and Tuomas Virtanen. Blind separation of audio mixtures through nonnegative tensor factorization of modulation spectrograms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2377– 2389, ???? 2016. CODEN ???? ISSN 2329-9290.

Baby:2017:JDD

[BV17] Deepak Baby and Hugo Van hamme. Joint denoising and dereverberation using exemplarbased sparse representations and decaying norm constraint. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(10):2024– 2035, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

                                        ieeexplore.ieee.org/
                                        document/8016372/.

Barysenka:2023:SBI

[BV23] Siarhei Y. Barysenka and Vasili I. Vorobiov. SNRbased inter-component phase estimation using biphase prior statistics for single-channel speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2365– 2381, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3284514.

Baby:2015:CDE

[BVGV15] D. Baby, T. Virtanen, J. F. Gemmeke, and H. Van hamme. Coupled dictionaries for exemplar-based speech enhancement and automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1788–1799, November 2015. CODEN ???? ISSN 2329-9290.

Bernardi:2018:SOS

[BvWM18] G. Bernardi, T. van Waterschoot, J. Wouters, and M. Moonen. Subjective and objective soundquality evaluation of adaptive feedback cancellation algorithms. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(5): 1010–1024, May 2018. CODEN ???? ISSN 2329-9290.

Bernardi:2017:AFC

[BvWWM17] Giuliano Bernardi, Toon van Waterschoot, Jan Wouters, and Marc Moonen. Adaptive feedback cancellation using a partitioned-block frequency-domain Kalman filter approach with PEMbased signal prewhitening. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(9):1784–1798, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7949026/.

Baba:2018:RGI

[BWH18] Y. E. Baba, A. Walther, and E. A. P. Habets. 3D room geometry inference based on room impulse response stacks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (5):857–872, May 2018. CODEN ???? ISSN 2329-9290.

Bai:2022:EES

[BWZC22] Zhongxin Bai, Jianyu Wang, Xiao-Lei Zhang, and Jingdong Chen. Endto-end speaker verification via curriculum bipartite ranking weighted binary cross-entropy. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1330–1344, 2022. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2022.3161155.

Bao:2014:LDD

[BXY14] Guangzhao Bao, Yangfei Xu, and Zhongfu Ye. Learning a discriminative dictionary for singlechannel speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(7):1130–1138, July 2014. CODEN ???? ISSN 2329-9290.

Bicici:2015:OIS

[BY15] E. Bicici and D. Yuret.

Optimizing instance selection for statistical machine translation with feature decay algorithms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (2):339–350, February 2015. CODEN ???? ISSN 2329-9290.

Bai:2021:FEE

[BYT^{+}21a] Ye Bai, Jiangyan Yi, Jianhua Tao, Zhengkun Tian, Zhengqi Wen, and Shuai Zhang. Fast end-to-end speech recognition via non-autoregressive models and cross-modal knowledge transferring from BERT. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1897– 1911, 2021. CODEN ???? ISSN 2329-9290.

Bai:2021:IKE

[BYT^{+}21b] Ye Bai, Jiangyan Yi, Jianhua Tao, Zhengqi Wen, Zhengkun Tian, and Shuai Zhang. Integrating knowledge into endto-end speech recognition from external text-only data. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1340–1351, 2021. CODEN ???? ISSN 2329-

Bai:2023:KIR

[BYY^{+}23] Jiaqi Bai, Ze Yang, Jian Yang, Hongcheng Guo, and Zhoujun Li. KINet: Incorporating relevant facts into knowledgegrounded dialog generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1213– 1222, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3240654.

Bouzid:2021:ELA

[BZ21] Saoussen Mathlouthi Bouzid and Chiraz Ben Othmane Zribi. Efficient learning approach for pronominal anaphora and ellipsis identification and resolution in Arabic texts. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):3335–3348, 2021. CODEN ???? ISSN 2329-9290.

Bao:2023:GCR

[BZ23] Guangsheng Bao and Yue Zhang. A general contextualized rewriting framework for text summarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1624– 1635, 2023. ISSN 2329-

                                        9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2023.3268569.

Carofilis:2023:IAC

[CAFFR23] Andrés Carofilis, Enrique Alegre, Eduardo Fidalgo, and Laura Fernández-Robles. Improvement of accent classification models through grad-transfer from spectrograms and gradient-weighted class activation mapping. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 2859–2871, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3297961.

Chen:2017:DRE

[CASZ17] Hanchi Chen, Thushara Dheemantha Abhayapala, Prasanga N. Samarasinghe, and Wen Zhang. Direct-to-reverberant energy ratio estimation using a first-order microphone. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(2):226– 237, ???? 2017. CODEN ???? ISSN 2329-9290.

Canclini:2020:MRE

[CATS20] Antonio Canclini, Fabio Antonacci, Stefano Tubaro, and Augusto Sarti. A methodology for the robust estimation of the radiation pattern of acoustic sources. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):211– 224, January 2020. CO-DEN ???? ISSN 2329-9290. URL https:/

                                        /dl.acm.org/doi/abs/
                                        10.1109/TASLP.2019.2950767.

Cho:2014:RII

[CB14] Taemin Cho and J. P. Bello. On the relative importance of individual components of chord recognition systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):477–492, February 2014. CODEN ???? ISSN 2329-9290.

Chen:2021:PUS

[CB21] Xianhong Chen and Changchun Bao. Phoneme-unit-specific time-delay neural network for speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1243– 1255, 2021. CODEN ???? ISSN 2329-9290.

Canclini:2015:RLC

[CBA^{+}15] A. Canclini, P. Bestagini, F. Antonacci, M. Compagnoni, A. Sarti, and S. Tubaro. A robust and low-complexity source localization algorithm for asynchronous distributed microphone networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1563– 1575, October 2015. CO-DEN ???? ISSN 2329-9290.

Cohen:2019:DKP

[CBC19] I. Cohen, J. Benesty, and J. Chen. Differential Kronecker product beamforming. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(5): 892–902, May 2019. CO-DEN ???? ISSN 2329-9290.

Chong:2015:DWP

[CBCL15] Tze Yuang Chong, R. E.

Banchs, Eng Siong Chng, and Haizhou Li. Decoupling word-pair distance and co-occurrence information for effective long history context language modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(7):1221– 1232, July 2015. CODEN ???? ISSN 2329-9290.

Chen:2015:SEF

[CBG15] Langzhou Chen, N. Braunschweiler, and M. J. F. Gales. Speaker and expression factorization for audiobook data: Expressiveness and transplantation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(4):605–618, April 2015.

CODEN ???? ISSN 2329-

9290.

Chen:2021:RSB

[CBHC21] Xi Chen, Jacob Benesty, Gongping Huang, and Jingdong Chen. On the robustness of the superdirective beamformer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):838–849, 2021. CO-DEN ???? ISSN 2329-9290.

Chau:2024:NAM

[CBN^{+}24] Hoang Ngoc Chau, Tien Dat Bui, Huu Binh Nguyen, Thanh Thi Hien Duong, and Quoc Cuong Nguyen. A novel approach to multi-channel speech enhancement based on graph neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32 (??):1133–1144, 2024. ISSN 2329-9290. URL

                                        https://dl.acm.org/doi/
                                        10.1109/TASLP.2024.3352259.

Casebeer:2023:MAM

[CBS23] Jonah Casebeer, Nicholas J. Bryan, and Paris Smaragdis. Meta-AF: Meta-learning for adaptive filters. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 355–370, 2023. ISSN 2329-

                                        9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2022.3224288.

Chang:2014:FAN

[CC14a] Dah-Chung Chang and Fei-Tao Chu. Feedforward active noise control with a new variable tap-length and step-size filtered-X LMS algorithm. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): 542–555, February 2014. CODEN ???? ISSN 2329-9290.

Choi:2014:DMV

[CC14b] Jae-Hun Choi and Joon-Hyuk Chang. Dualmicrophone voice activity detection technique based on two-step power level difference ratio. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(6): 1069–1081, June 2014. CODEN ???? ISSN 2329-9290.

Chu:2015:NLP

[CC15] Y. J. Chu and S. C.

Chan. A new local polynomial modeling-based variable forgetting factor RLS algorithm and its acoustic applications. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):2059–2069, November 2015. CODEN ???? ISSN 2329-9290.

Chan:2021:MBC

[CC21] Teck Kai Chan and Cheng Siong Chin. Multibranch convolutional macaron net for sound event detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2972– 2985, 2021. CODEN ???? ISSN 2329-9290.

Chiea:2021:OEB

[CCC21] Rafael A. Chiea, Márcio H. Costa, and Júlio A. Cordioli. An optimal envelopebased noise reduction method for cochlear implants: an upper bound performance investigation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1729–1739, 2021. CODEN ???? ISSN 2329-9290.

Cui:2024:EME

[CCC^{+}24] Shiyao Cui, Jiangxia Cao, Xin Cong, Jiawei Sheng, Quangang Li, Tingwen Liu, and Jinqiao Shi. Enhancing multimodal entity and relation extraction with variational information bottleneck. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1274– 1285, 2024. ISSN 2329-

                                        9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2023.3345146.

Chen:2021:MER

[CCH^{+}21] Bingzhi Chen, Qi Cao, Mixiao Hou, Zheng Zhang, Guangming Lu, and David Zhang. Multimodal emotion recognition with temporal and semantic consistency. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3592– 3603, 2021. CODEN ???? ISSN 2329-9290.

Cui:2021:PTW

[CCL^{+}21] Yiming Cui, Wanxiang Che, Ting Liu, Bing Qin, and Ziqing Yang. Pre-training with whole word masking for Chinese BERT. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3504– 3514, 2021. CODEN ???? ISSN 2329-9290.

Carini:2016:RRI

[CCR16] Alberto Carini, Stefania Cecchi, and Laura Romoli. Robust room impulse response measurement using perfect sequences for Legendre nonlinear filters. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):1969– 1982, ???? 2016. CODEN ???? ISSN 2329-9290.

Chen:2019:ATU

[CCT^{+}19] L. Chen, Z. Chen, B. Tan, S. Long, M. Gašić, and K. Yu. AgentGraph: Toward universal dialogue management with structured deep reinforcement learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(9):1378– 1391, September 2019. CODEN ???? ISSN 2329-9290.

Chen:2021:SIT

[CCT21] Zuolong Chen, Huawei Chen, and Quansheng Tu. Sensor imperfection tolerance analysis of robust linear differential microphone arrays. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2915–2929, 2021. CODEN ???? ISSN 2329-9290.

Carini:2021:RIR

[CCTO21] Alberto Carini, Stefania Cecchi, Alessandro Terenzi, and Simone Orcioni. A room impulse response measurement method robust towards nonlinearities based on orthogonal periodic sequences. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3104– 3117, 2021. CODEN ???? ISSN 2329-9290.

Chen:2024:ITC

[CCY^{+}24] Yifan Chen, Gaofeng Cheng, Runyan Yang, Pengyuan Zhang, and Yonghong Yan. Interrelate training and clustering for online speaker diarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1352– 1364, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3357033.

Cheer:2017:IDS

[CD17] Jordan Cheer and Stephen Daley. An investigation of delayless subband adaptive filtering for multiinput multi-output active noise control applications. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (2):359–373, ???? 2017. CODEN ???? ISSN 2329-9290.

Conan:2014:SMI

[CDA^{+}14] S. Conan, O. Derrien, M. Aramaki, S. Ystad, and R. Kronland-Martinet. A synthesis model with intuitive control capabilities for rolling sounds. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(8): 1260–1273, August 2014.

CODEN ???? ISSN 2329-9290.

Chai:2019:UGG

[CDLL19] L. Chai, J. Du, Q. Liu, and C. Lee. Using generalized Gaussian distributions to improve regression error modeling for deep learningbased speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):1919–1931, December 2019. CODEN ???? ISSN 2329-9290.

Chai:2021:CEG

[CDLL21] Li Chai, Jun Du, Qing-Feng Liu, and Chin-Hui Lee. A crossentropy-guided measure (CEGM) for assessing speech recognition performance and optimizing DNN-based speech enhancement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):106–117, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3036783.

Chen:2018:PJM

[CDLX18] Zhehuai Chen, Jasha Droppo, Jinyu Li, and Wayne Xiong. Progressive joint modeling in unsupervised single-channel overlapped speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (1):184–196, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8080252/.

Cogliati:2016:CDP

[CDW16] Andrea Cogliati, Zhiyao Duan, and Brendt Wohlberg. Context-dependent piano music transcription with convolutional sparse coding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2218–2230, ???? 2016. CODEN ???? ISSN 2329-9290.

Cheer:2014:CCP

[CE14] J. Cheer and S. J. Elliott. Comments on “Complete Parallel Narrowband Active Noise Control Systems”. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(5): 995–996, May 2014. CO-DEN ???? ISSN 2329-9290.

Carlin:2015:FSA

[CE15] M. A. Carlin and M. Elhilali. A framework for speech activity detection using adaptive auditory receptive fields. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2422–2433, December 2015. CODEN ???? ISSN 2329-9290.

Cosentino:2014:CIF

[CFMM14] S. Cosentino, T. H. Falk, D. McAlpine, and T. Marquardt. Cochlear implant filterbank design and optimization: A simulation study. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):347–353, February 2014. CODEN ???? ISSN 2329-9290.

Cui:2016:MLN

[CG16] Xiaodong Cui and Vaibhava Goel. Maximum likelihood nonlinear transformations based on deep neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(11): 2023–2031, ???? 2016. CODEN ???? ISSN 2329-9290. See correction [?].

Cherkassky:2017:BSW

[CG17] Dani Cherkassky and Sharon Gannot. Blind synchronization in wireless acoustic sensor networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(3):651– 661, ???? 2017. CODEN ???? ISSN 2329-9290.

Cherkassky:2020:SRT

[CG20] Dani Cherkassky and Sharon Gannot. Successive relative transfer function identification using blind oblique projection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):474–486, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2957883.

Chien:2021:AMV

[CG21] Yu-Ren Chien and Jón Gunason. Acoustic measure of vocal strain based on glottal airflow periodicity. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):563–574, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3044168.

Cui:2015:DAD

[CGK15] Xiaodong Cui, V. Goel, and B. Kingsbury. Data augmentation for deep neural network acoustic modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(9):1469– 1477, September 2015. CODEN ???? ISSN 2329-

Cernak:2015:ISC

[CGL^{+}15] M. Cernak, P. N. Garner, A. Lazaridis, P. Motlicek, and Xingyu Na. Incremental syllable-context phonetic vocoding. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(6): 1019–1030, June 2015. CODEN ???? ISSN 2329-9290.

Chen:2024:WSS

[CGL^{+}24] Xiuying Chen, Shen Gao, Mingzhe Li, Qingqing Zhu, Xin Gao, and Xiangliang Zhang. Write summary step-by-step: a pilot study of stepwise summarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1406– 1415, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3357040.

Crespo:2014:MSR

[CH14] J. B. Crespo and R. C. Hendriks. Multizone speech reinforcement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): 54–66, January 2014. CO-DEN ???? ISSN 2329-9290.

Chen:2015:OMS

[CH15] Zhangli Chen and V. Hohmann. Online monaural speech enhancement based on periodicity analysis and a priori SNR estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1904–1916, November 2015. CODEN ???? ISSN 2329-9290.

Chen:2016:TDB

[CH16] Kai Chen and Qiang Huo. Training deep bidirectional LSTM acoustic model for LVCSR by a context-sensitive-chunk BPTT approach. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (7):1185–1193, July 2016. CODEN ???? ISSN 2329-9290.

Chakrabarty:2018:BAI

[CH18] Soumitro Chakrabarty and Emanuël A. P. Habets. A Bayesian approach to informed spatial filtering with robustness against DOA estimation errors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(1): 145–160, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8167179/.

Coteli:2021:SRL

[ÇH21] Mert Burkay Çöteli and Hüseyin Hacıhabiboğlu.

Sparse representations with Legendre kernels for DOA estimation and acoustic source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2296–2309, 2021. CODEN ???? ISSN 2329-9290.

Chetupalli:2023:SCS

[CH23] Srikanth Raj Chetupalli and Emanuël A. P. Habets. Speaker counting and separation from single-channel noisy mixtures. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1681–1692, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3268572.

Chen:2014:MSR

[CHC^{+}14] Sin-Horng Chen, Chiao-Hua Hsieh, Chen-Yu Chiang, Hsi-Chun Hsiao, Yih-Ru Wang, Yuan-Fu Liao, and Hsiu-Min Yu. Modeling of speaking rate influences on Mandarin speech prosody and its application to speaking ratecontrolled TTS. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(7): 1158–1171, July 2014. CODEN ???? ISSN 2329-9290.

Chen:2023:IIM

[CHGS23] Chen Chen, Hansheng Hong, Jie Guo, and Bin Song. Inter- intra modal representation augmentation with trimodal collaborative disentanglement network for multimodal sentiment analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1476– 1488, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3263801.

Chien:2015:HPY

[Chi15a] Jen-Tzung Chien. Hierarchical Pitman–Yor– Dirichlet language model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (8):1259–1272, August 2015. CODEN ???? ISSN 2329-9290.

Chien:2015:LGS

[Chi15b] Jen-Tzung Chien. Laplace group sensing for acoustic models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(5):909– 922, May 2015. CODEN ???? ISSN 2329-9290.

Chiang:2018:CDA

[Chi18a] Chen-Yu Chiang. Crossdialect adaptation framework for constructing prosodic models for Chinese dialect text-to-speech systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(1):108– 121, ???? 2018. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/8066318/.

Chien:2018:BNL

[Chi18b] Jen-Tzung Chien. Bayesian nonparametric learning for hierarchical and sparse topics. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(2):422– 435, ???? 2018. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/8141927/.

Chen:2014:MSA

[CHJ14] A. Chen and M. A. Hasegawa-Johnson. Mixed stereo audio classification using a stereo-input mixed-to-panned level feature. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2025–2033, December 2014. CODEN ???? ISSN 2329-9290.

Chen:2019:AWS

[CHL^{+}19] Y. Chen, S. Huang, H. Lee, Y. Wang, and C. Shen. Audio Word2vec:

Sequence-to-sequence autoencoding for unsupervised learning of audio segmentation and representation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(9):1481– 1493, September 2019. CODEN ???? ISSN 2329-9290.

Chuang:2023:MGM

[CHL^{+}23] Yun-Yen Chuang, Hung-Min Hsu, Kevin Lin, Ray-I. Chang, and Hung-Yi Lee. MetaEx-GAN: Meta exploration to improve natural language generation via generative adversarial networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3968– 3980, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3317571.

Chi:2024:CPE

[CHL^{+}24] Zewen Chi, Heyan Huang, Luyang Liu, Yu Bai, Xiaoyan Gao, and Xian-Ling Mao. Can pretrained English language models benefit non-English NLP systems in low-resource scenarios? IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1061– 1074, 2024. ISSN 2329-

9290. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2023.3267618.

Hsu:2023:PSA

[cHrLLyL23] Po chun Hsu, Da rong Liu, Andy T. Liu, and Hung yi Lee. Parallel synthesis for autoregressive speech generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3095– 3111, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3301212.

Chen:2014:PSS

[CHWL14] Chia-Ping Chen, Yi-Chin Huang, Chung-Hsien Wu, and Kuan-De Lee. Polyglot speech synthesis based on cross-lingual frame selection using auditory and articulatory features. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(10):1558– 1570, October 2014. CO-DEN ???? ISSN 2329-9290.

Chen:2024:ABE

[CHWQ24] Zhengyang Chen, Bing Han, Shuai Wang, and Yanmin Qian. Attentionbased encoder-decoder end-to-end neural diarization with embedding enhancer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1636– 1649, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3366756.

Chen:2015:APS

[CJ15] Liang-Yu Chen and J.- S. R. Jang. Automatic pronunciation scoring with score combination by learning to rank and class-normalized DP-based quantization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1737–1749, November 2015. CODEN ???? ISSN 2329-9290.

Cheng:2021:UTO

[CJY^{+}21] Zifeng Cheng, Zhiwei Jiang, Yafeng Yin, Na Li, and Qing Gu. A unified target-oriented sequenceto-sequence model for emotion-cause pair extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2779– 2791, 2021. CODEN ???? ISSN 2329-9290.

Cheng:2022:LCO

[CJY^{+}22] Zifeng Cheng, Zhiwei Jiang, Yafeng Yin, Cong Wang, and Qing Gu. Learning to classify open intent via soft labeling and manifold mixup. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):635–645, 2022. CO-DEN ???? ISSN 2329-9290.

Chua:2019:LLA

[CK19] J. Chua and W. B. Kleijn. A low latency approach for blind source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(8):1280–1294, August 2019. CODEN ???? ISSN 2329-9290.

Cumani:2014:FSS

[CL14a] S. Cumani and P. Laface. Factorized sub-space estimation for fast and memory effective I-vector extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):248– 259, January 2014. CO-DEN ???? ISSN 2329-9290.

Cumani:2014:LST

[CL14b] S. Cumani and P. Laface. Large-scale training of pairwise support vector machines for speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(11):1590– 1600, November 2014. CODEN ???? ISSN 2329-9290.

Cumani:2017:JEP

[CL17a] Sandro Cumani and Pietro Laface. Joint estimation of PLDA and nonlinear transformations of speaker vectors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(10): 1890–1900, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7971950/.

Cumani:2017:NVT

[CL17b] Sandro Cumani and Pietro Laface. Nonlinear Ivector transformations for PLDA-based speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(4):908– 919, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7864395/.

Chung:2018:UDS

[CL18a] Cheng-Tao Chung and Lin-Shan Lee. Unsupervised discovery of structured acoustic tokens with applications to spoken term detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(2): 394–405, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8125178/.

Cumani:2018:SHS

[CL18b] S. Cumani and P. Laface. Scoring heterogeneous speaker vectors using nonlinear transformations and tied PLDA models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (5):995–1009, May 2018. CODEN ???? ISSN 2329-9290.

Cumani:2018:SRU

[CL18c] S. Cumani and P. Laface. Speaker recognition using e vectors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(4): 736–748, April 2018. CO-DEN ???? ISSN 2329-9290.

Cai:2021:DNS

[CLA^{+}21] Yunqi Cai, Lantian Li, Andrew Abel, Xiaoyan Zhu, and Dong Wang. Deep normalization for speaker vectors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 733–744, 2021. CODEN ???? ISSN 2329-9290.

Cernak:2016:CDS

[CLAG16] Milos Cernak, Alexandros Lazaridis, Afsaneh Asaei, and Philip N. Garner. Composition of deep and spiking neural networks for very low bit rate speech coding. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(12): 2301–2312, ???? 2016. CODEN ???? ISSN 2329-9290.

Chen:2015:EBN

[CLC^{+}15] Kuan-Yu Chen, Shih-Hung Liu, B. Chen, Hsin-Min Wang, Ea-Ee Jan, Wen-Lian Hsu, and Hsin-Hsi Chen. Extractive broadcast news summarization leveraging recurrent neural network language modeling techniques. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):1322– 1334, August 2015. CO-DEN ???? ISSN 2329-9290.

Cui:2022:TMR

[CLC^{+}22] Yiming Cui, Ting Liu, Wanxiang Che, Zhigang Chen, and Shijin Wang. Teaching machines to read, answer and explain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1483– 1492, 2022. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3156789.

Chen:2018:IDF

[CLCW18] Kuan-Yu Chen, Shih-Hung Liu, Berlin Chen, and Hsin-Min Wang. An information distillation framework for extractive summarization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(1): 161–170, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8074745/.

Chu:2021:EER

[CLHK21] Yun-Wei Chu, Kuan-Yen Lin, Chao-Chun Hsu, and Lun-Wei Ku. End-to-end recurrent cross-modality attention for video dialogue. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2456–2464, 2021. CODEN ???? ISSN 2329-9290.

Chen:2021:LGE

[CLL^{+}21a] Gang Chen, Yang Liu, Huanbo Luan, Meng Zhang, Qun Liu, and Maosong Sun. Learning to generate explainable plots for neural story generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):585–593, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3039606.

Cheng:2021:DAN

[CLL^{+}21b] Jiaming Cheng, Ruiyu Liang, Zhenlin Liang, Li Zhao, Chengwei Huang, and Björn Schuller. A deep adaptation network for speech enhancement: Combining a relativistic discriminator with multi-kernel maximum mean discrepancy. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):41–53, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3036611.

Chen:2014:VCU

[CLLD14] Ling-Hui Chen, Zhen-Hua Ling, Li-Juan Liu, and Li-Rong Dai. Voice conversion using deep neural networks with layerwise generative training. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1859–1872, December 2014. CODEN ???? ISSN 2329-9290.

Chuang:2021:IAS

[CLSyL21] Shun-Po Chuang, Alexander H. Liu, Tzu-Wei Sung, and Hung yi Lee. Improving automatic speech recognition and speech translation via word embedding prediction. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 93–105, January 2021.

URL https://dl.acm.

                                        org/doi/10.1109/TASLP.
                                        2020.3037543.

Chen:2016:ETE

[CLW^{+}16] Xie Chen, Xunying Liu, Yongqiang Wang, Mark J. F. Gales, and Philip C. Woodland. Efficient training and evaluation of recurrent neural network language models for automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):2146–2157, ???? 2016. CODEN ???? ISSN 2329-9290.

Chen:2019:EFW

[CLW^{+}19] X. Chen, X. Liu, Y. Wang, A. Ragni, J. H. M. Wong, and M. J. F. Gales. Exploiting future word contexts in neural network language models for speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (9):1444–1454, September 2019. CODEN ???? ISSN 2329-9290.

Chen:2024:DMC

[CLW24] Zhe Chen, Hongcheng Liu, and Yu Wang. DialogMCF: Multimodal context flow for audio visual scene-aware dialog. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):753– 764, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3284511.

Cui:2022:LAN

[CLZ22] Leyang Cui, Yafu Li, and Yue Zhang. Label attention network for structured prediction. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1235–1248, 2022. CODEN ???? ISSN 2329-9290.

Cao:2023:DWE

[CLZ^{+}23] Hailong Cao, Liguo Li, Conghui Zhu, Muyun Yang, and Tiejun Zhao. Dual word embedding for robust unsupervised bilingual lexicon induction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2606–2615, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3290425.

Chen:2024:SLN

[CLZ^{+}24] Xiang Chen, Lei Li, Yuqi Zhu, Shumin Deng, Chuanqi Tan, Fei Huang, Luo Si, Ningyu Zhang, and Huajun Chen. Sequence labeling as nonautoregressive dual-query set generation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??):

1546–1558, 2024. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2024.3358053.

Chen:2015:MLD

[CM15] Dongpeng Chen and B. K.-W. Mak. Multitask learning of deep neural networks for lowresource speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(7):1172–1183, July 2015. CODEN ???? ISSN 2329-9290.

Chang:2018:PLM

[CM18] J. Chang and M. Marschall. Periphony-lattice mixedorder ambisonic scheme for spherical microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(5):924– 936, May 2018. CODEN ???? ISSN 2329-9290.

Chiu:2023:LPB

[CMD^{+}23] Ching-Yu Chiu, Meinard Müller, Matthew E. P. Davies, Alvin Wen-Yu Su, and Yi-Hsuan Yang. Local periodicity-based beat tracking for expressive classical piano music. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2824– 2835, 2023. ISSN 2329-

                                        9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2023.3297956.

Chien:2017:EGI

[CMG^{+}17] Yu-Ren Chien, Daryush D.

Mehta, Jón Gunason, Matías Zañartu, and Thomas F. Quatieri. Evaluation of glottal inverse filtering algorithms using a physiologically based articulatory speech synthesizer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(8):1718–1730, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7946161/.

Chen:2023:NMD

[CMM^{+}23] Feng Chen, Ke Ma, Yapeng Mao, Desen Yang, Yi Zhang, Jie Shi, Shiqi Mo, Gui Chenyang, and Song Li. A novel method to design steerable differential beamformer using linear acoustics vector sensor array. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2836– 2849, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3297960.

Chen:2016:TIP

[CMY16] Lijiang Chen, Xia Mao, and Hong Yan. Textindependent phoneme segmentation combining EGG and speech data. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (6):1029–1037, June 2016. CODEN ???? ISSN 2329-9290.

Chen:2021:CAG

[CMY^{+}21] Haibin Chen, Qianli Ma, Liuhong Yu, Zhenxi Lin, and Jiangyue Yan. Corpus-aware graph aggregation network for sequence labeling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2048–2057, 2021. CODEN ???? ISSN 2329-9290.

Cheng:2022:EUA

[CMY^{+}22] Gaofeng Cheng, Haoran Miao, Runyan Yang, Keqi Deng, and Yonghong Yan. ETEH: Unified attention-based end-to-end ASR and KWS architecture. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1360– 1373, 2022. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3161159.

Caviedes-Nozal:2023:STB

[CNFG23] Diego Caviedes-Nozal and Efren Fernandez-Grande. Spatio-temporal Bayesian regression for room impulse response reconstruction with spherical waves. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3263– 3277, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3306708.

Coteli:2018:MSS

[ÇOH18] M. B. Çöteli, O. Olgun, and H. Hacıhabiboğlu. Multiple sound source localization with steered response power density and hierarchical grid refinement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):2215–2229, November 2018. CODEN ???? ISSN 2329-9290.

Carabias-Orti:2018:MBS

[CONVVC18] J. J. Carabias-Orti, J. Nikunen, T. Virtanen, and P. Vera-Candeas. Multichannel blind sound source separation using spatial covariance model with level and time differences and nonnegative matrix factorization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1512–1527, September 2018. CODEN ???? ISSN 2329-9290.

Cho:2021:CML

[CP21] Byung Joon Cho and Hyung-Min Park. Convolutional maximum-likelihood distortionless response beamforming with steering vector estimation for robust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1352–1367, 2021. CODEN ???? ISSN 2329-9290.

Cavalieri:2016:CLM

[CPCBFSF16] Daniel C. Cavalieri, Sira E. Palazuelos-Cagigas, Teodiano F. Bastos-Filho, and Mário Sarcinelli-Filho. Combination of language models for word prediction: An exponential approach. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(9): 1481–1494, ???? 2016. CODEN ???? ISSN 2329-9290.

Chen:2024:UMV

[CPGL24] Chin-Po Chen, Ho-Hsien Pan, Susan Shur-Fen Gau, and Chi-Chun Lee. Using measures of vowel space for autistic traits characterization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 591–607, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3330605.

Cumani:2014:UVP

[CPL14] S. Cumani, O. Plchot, and P. Laface. On the use of i-vector posterior distributions in probabilistic linear discriminant analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(4):846–857, April 2014. CODEN ???? ISSN 2329-9290.

Cobos:2014:CSB

[CPSFC^{+}14] M. Cobos, J. J. Perez-Solano, S. Felici-Castell, J. Segura, and J. M. Navarro. Cumulativesum-based localization of sound events in low-cost wireless acoustic sensor networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1792– 1802, December 2014. CODEN ???? ISSN 2329-9290.

Chen:2022:RED

[CQ22] Zhuang Chen and Tieyun Qian. Retrieve-and-edit domain adaptation for End2End aspect based sentiment analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):659–672, 2022. CO-DEN ???? ISSN 2329-9290.

Chivukula:2014:FAL

[CRH^{+}14] R. K. Chivukula, Y. A.

Reznik, Yanyan Hu, V. Devarajan, and M. Jayendra-Lakshman. Fast algorithms for low-delay TDAC filterbanks in MPEG-4 AAC–ELD. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1701–1712, December 2014. CODEN ???? ISSN 2329-9290.

Chen:2015:DGA

[CRVB^{+}15] L.-H. Chen, T. Raitio, C. Valentini-Botinhao, Z.- H. Ling, and J. Yamagishi. A deep generative architecture for postfiltering in statistical parametric speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):2003–2014, November 2015. CODEN ???? ISSN 2329-9290.

Chetupalli:2019:LRC

[CS19] S. R. Chetupalli and T. V. Sreenivas. Late reverberation cancellation using Bayesian estimation of multi-channel linear predictors and Student’s tsource prior. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(6):1007– 1018, June 2019. CODEN ???? ISSN 2329-9290.

Cumani:2023:DUS

[CS23] Sandro Cumani and Salvatore Sarni. The distributions of uncalibrated speaker verification scores: a generative model for domain mismatch and trialdependent calibration. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):2204–2219, 2023. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2023.3282096.

Comminiello:2014:NAE

[CSAR^{+}14] D. Comminiello, M. Scarpiniti, L. A. Azpicueta-Ruiz, J. Arenas-Garcia, and A. Uncini. Nonlinear acoustic echo cancellation based on sparse functional link representations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(7):1172–1183, July 2014. CODEN ???? ISSN 2329-9290.

Celikyilmaz:2016:EIW

[CSJD16] Asli Celikyilmaz, Ruhi Sarikaya, Minwoo Jeong, and Anoop Deoras. An empirical investigation of word class-based features for natural language understanding. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(6): 994–1005, June 2016. CO-DEN ???? ISSN 2329-9290.

Cauchi:2019:NIS

[CSS^{+}19] B. Cauchi, K. Siedenburg, J. F. Santos, T. H. Falk, S. Doclo, and S. Goetze. Non-intrusive speech quality prediction using modulation energies and LSTM-network. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (7):1151–1163, July 2019. CODEN ???? ISSN 2329-9290.

Chen:2023:SSP

[CSS23] Yi-Syuan Chen, Yun-Zhu Song, and Hong-Han Shuai. SPEC: Summary preference decomposition for lowresource abstractive summarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):603– 618, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3230539.

Cheng:2021:ERF

[CSY^{+}21] Longbiao Cheng, Xingwei Sun, Dingding Yao, Junfeng Li, and Yonghong Yan. Estimation reliability function assisted sound source localization with enhanced steering vector phase difference. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):421–435, January

2021. URL https://dl.

acm.org/doi/10.1109/
TASLP.2020.3043107.

Chen:2024:ANG

[CSZZ24] Huiyao Chen, Yueheng Sun, Meishan Zhang, and Min Zhang. Automatic noise generation and reduction for text classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):139– 150, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3325135.

Crocco:2014:DSP

[CT14] M. Crocco and A. Trucco.

Design of superdirective planar arrays with sparse aperiodic layouts for processing broadband signals via 3-D beamforming. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(4):800–815, April 2014. CODEN ???? ISSN 2329-9290.

Chinaev:2021:DCC

[CTE21] Aleksej Chinaev, Philipp Thüne, and Gerald Enzner. Double-cross-correlation processing for blind sampling-rate and time-offset estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):1881–1896, 2021.

CODEN ???? ISSN 2329-9290.

Chung:2017:UID

[CTLL17] Cheng-Tao Chung, Cheng-Yu Tsai, Chia-Hsiang Liu, and Lin-Shan Lee. Unsupervised iterative deep learning of speech features and acoustic tokens with applications to spoken term detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(10): 1914–1928, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7984872/.

Chen:2014:CPT

[CTSC14] N. F. Chen, S. W. Tam, Wade Shen, and J. P. Campbell. Characterizing phonetic transformations and acoustic differences across English dialects. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): 110–124, January 2014. CODEN ???? ISSN 2329-9290.

Cumani:2015:FSF

[Cum15] S. Cumani. Fast scoring of full posterior PLDA models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):2036–2045, November 2015. CODEN ???? ISSN 2329-9290.

Cumani:2021:DSV

[Cum21] Sandro Cumani. On the distribution of speaker verification scores: Generative models for unsupervised calibration. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):547–562, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3040103.

Calis:2021:LBE

[CvdPHH21] Metin Calis, Steven van de Par, Richard Heusdens, and Richard Christian Hendriks. Localization based on enhanced low frequency interaural level difference. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3025– 3039, 2021. CODEN ???? ISSN 2329-9290.

Chandna:2018:BAM

[CW18] S. Chandna and W. Wang.

Bootstrap averaging for model-based source separation in reverberant conditions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(4):806– 819, April 2018. CODEN ???? ISSN 2329-9290.

Chorowski:2019:USR

[CWBv19] J. Chorowski, R. J. Weiss, S. Bengio, and A. van den Oord. Unsupervised speech representation learning using WaveNet autoencoders. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2041–2053, December 2019. CODEN ???? ISSN 2329-9290.

Chen:2017:IMA

[CWCP17] Fei Chen, Lan Wang, Hui Chen, and Gang Peng. Investigations on Mandarin aspiratory animations using an airflow model. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12): 2399–2409, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8114353/.

Carroll:2017:OLA

[CWDA17] Brandon T. Carroll, Bradley M. Whitaker, Wayne Dayley, and David V. Anderson. Outlier learning via augmented frozen dictionaries. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(6):1207–1215, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7933057/.

Chen:2016:CEP

[CWH^{+}16] Yan-You Chen, Chung-Hsien Wu, Yi-Chin Huang, Shih-Lun Lin, and Jhing-Fa Wang. Candidate expansion and prosody adjustment for natural speech synthesis using a small corpus. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(6): 1052–1065, June 2016. CODEN ???? ISSN 2329-9290.

Chien:2015:APM

[CWJ15] Yu-Ren Chien, Hsin-Min Wang, and Shyh-Kang Jeng. An acousticphonetic model of F0 likelihood for vocal melody extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(9):1457– 1468, September 2015. CODEN ???? ISSN 2329-9290.

Chien:2016:ALA

[CWJ16] Yu-Ren Chien, Hsin-Min Wang, and Shyh-Kang Jeng. Alignment of lyrics with accompanied singing audio based on acousticphonetic vowel likelihood modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):1998– 2008, ???? 2016. CODEN ???? ISSN 2329-9290.

Cai:2022:IVI

[CWL22] Danwei Cai, Weiqing Wang, and Ming Li. Incorporating visual information in audio based self-supervised speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1422– 1435, 2022. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3162078.

Chen:2022:TAF

[CWR^{+}22] Xiaofeng Chen, Guohua Wang, Haopeng Ren, Yi Cai, Ho fung Leung, and Tao Wang. Taskadaptive feature fusion for generalized few-shot relation classification in an open world environment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1003–1015, 2022. CODEN ???? ISSN 2329-9290.

Chuang:2022:ILA

[CWT22] Shang-Yi Chuang, Hsin-Min Wang, and Yu Tsao. Improved Lite audiovisual speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1345–1359, 2022. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3153265.

Chen:2019:NMT

[CWU^{+}19] K. Chen, R. Wang, M. Utiyama, E. Sumita, and T. Zhao. Neural machine translation with sentence-level topic context. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):1970–1984, December 2019. CODEN ???? ISSN 2329-9290.

Chen:2022:IPT

[CWUS22] Kehai Chen, Rui Wang, Masao Utiyama, and Eiichiro Sumita. Integrating prior translation knowledge into neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):330– 339, 2022. CODEN ???? ISSN 2329-9290.

Chen:2014:FSC

[CWW14] Jitong Chen, Yuxuan Wang, and Deliang Wang. A feature study for classification-based speech separation at low signalto-noise ratios. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1993–2002, December 2014. CODEN ???? ISSN 2329-9290.

Chen:2017:CTM

[CWYC17] Yu-An Chen, Ju-Chiang Wang, Yi-Hsuan Yang, and Homer H. Chen. Component tying for mixture model adaptation in personalization of music emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (7):1409–1420, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7898399/.

Chen:2023:LLM

[CWZ^{+}23] Xing Chen, Jie Wang, Xiao-Lei Zhang, Wei-Qiang Zhang, and Kunde Yang. LMD: a learnable mask network to detect adversarial examples for speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2476– 2490, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3288417.

Chang:2023:JBJ

[CXvG^{+}23] Hongyang Chang, Hongfei Xu, Josef van Genabith, Deyi Xiong, and Hongying Zan. JoinER-BART: Joint entity and relation extraction with constrained decoding, representation reuse and fusion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3603– 3616, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3310879.

Chen:2023:SHE

[CXX^{+}23] Weidong Chen, Xiaofen Xing, Xiangmin Xu, Jianxin Pang, and Lan Du. SpeechFormer++: a hierarchical efficient framework for paralinguistic speech processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):775–788, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3235194.

Cho:2015:UCB

[CY15] Janghoon Cho and C. D.

Yoo. Underdetermined convolutive BSS: Bayes risk minimization based on a mixture of super-Gaussian posterior approximation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(5): 828–839, May 2015. CO-DEN ???? ISSN 2329-9290.

Chien:2016:BFL

[CY16] Jen-Tzung Chien and Po-Kai Yang. Bayesian factorization and learning for monaural source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):185–195, January 2016. CODEN ???? ISSN 2329-9290.

Chen:2017:BNR

[CZ17] Yung-Yue Chen and Jia-Hao Zhang. Background noise reduction design for dual microphone cellular phones: Robust approach. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):852–862, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7478101/.

Che:2015:SCA

[CZG^{+}15] Wanxiang Che, Yanyan Zhao, Honglei Guo, Zhong Su, and Ting Liu. Sentence compression for aspect-based sentiment analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2111– 2124, December 2015. CODEN ???? ISSN 2329-9290.

Cui:2021:ADD

[CZK^{+}21] Xiaodong Cui, Wei Zhang, Abdullah Kayi, Mingrui Liu, Ulrich Finkler, Brian Kingsbury, George Saon, and David Kung. Asynchronous decentralized distributed training of acoustic models. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 3565–3576, 2021. CODEN ???? ISSN 2329-9290.

Chu:2024:NDF

[CZN^{+}24] Yijing Chu, Sipei Zhao, Feng Niu, Yongzheng Dong, and Yuezhe Zhao. A new diffusion filtered-X affine projection algorithm: Performance analysis and application in windy environment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1596– 1608, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3363409.

Chen:2023:MLC

[CZWC23] Guanhua Chen, Runzhe Zhan, Derek F. Wong, and Lidia S. Chao. Multilevel curriculum learning for multi-turn dialogue generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3958– 3967, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3322583.

Chen:2018:NAS

[CZY^{+}18] Kehai Chen, Tiejun Zhao, Muyun Yang, Lemao Liu, Akihiro Tamura, Rui Wang, Masao Utiyama, and Eiichiro Sumita. A neural approach to source dependence based context model for statistical machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (2):266–280, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8105847/.

Chen:2015:DFR

[CZZ15] Wenliang Chen, Min Zhang, and Yue Zhang. Distributed feature representations for dependency parsing. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(3): 451–460, March 2015. CODEN ???? ISSN 2329-9290.

Delgado:2015:FSC

[DAFS15] H. Delgado, X. Anguera, C. Fredouille, and J. Serrano. Fast single- and cross-show speaker diarization using binary key speaker modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2286–2297, December 2015. CODEN ???? ISSN 2329-9290.

Del-Agua:2018:SAC

[DAGS^{+}18] M. Á. Del-Agua, A. Giménez, A. Sanchis, J. Civera, and A. Juan. Speaker-adapted confidence measures for ASR using deep bidirectional recurrent neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(7):1198– 1206, July 2018. CODEN ???? ISSN 2329-9290.

DeSena:2015:MRG

[DAMvW15] E. De Sena, N. Antonello, M. Moonen, and T. van Waterschoot. On the modeling of rectangular geometries in room acoustic simulations. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(4): 774–786, April 2015. CO-DEN ???? ISSN 2329-9290.

Degottex:2016:MFA

[DAR16] Gilles Degottex, Luc Ardaillon, and Axel Roebel. Multi-frame amplitude envelope estimation for modification of singing voice. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(7):1242–1254, July 2016. CODEN ???? ISSN 2329-9290.

Dimitriadis:2015:UMM

[DB15] D. Dimitriadis and E. Bocchieri. Use of micromodulation features in large vocabulary continuous speech recognition tasks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):1348–1357, August 2015. CODEN ???? ISSN 2329-9290.

Dromey:2017:ELA

[DB17] Christopher Dromey and Katherine M. Black. Effects of laryngeal activity on articulation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12): 2272–2280, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8114361/.

Dionelis:2018:PAS

[DB18] N. Dionelis and M. Brookes. Phase-aware single-channel speech enhancement with modulation-domain Kalman filtering. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(5):937– 950, May 2018. CODEN ???? ISSN 2329-9290.

Dionelis:2019:MDK

[DB19] N. Dionelis and M. Brookes. Modulation-domain Kalman filtering for monaural blind speech denoising and dereverberation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(4): 799–814, April 2019. CO-DEN ???? ISSN 2329-9290.

Deng:2015:SHM

[DBK15] Feng Deng, Changchun Bao, and W. B. Kleijn. Sparse hidden Markov models for speech enhancement in non-stationary noise environments. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(11): 1973–1987, November 2015. CODEN ???? ISSN 2329-9290.

Diez:2020:ASD

[DBLČ20] Mireia Diez, Lukáš Burget, Federico Landini, and Jan Černocký. Analysis of speaker diarization based on Bayesian HMM with eigenvoice priors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):355–368, January 2020. CODEN ???? ISSN 2329-9290.

URL https://dl.acm.
org/doi/abs/10.1109/
TASLP.2019.2955293.

Do:2015:DAS

[DBM15] Quynh Thi Ngoc Do, S. Bethard, and M.-F. Moens. Domain adaptation in semantic role labeling using a neural language model and linguistic resources. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):1812– 1823, November 2015. CODEN ???? ISSN 2329-9290.

Doire:2017:SCO

[DBN^{+}17] Clement S. J. Doire, Mike Brookes, Patrick A. Naylor, Christopher M. Hicks, Dave Betts, Mohammad A. Dmour, and Søren Holdt Jensen. Singlechannel online enhancement of speech corrupted by reverberation and noise. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(3):572–587, ???? 2017. CODEN ???? ISSN 2329-9290.

Drgas:2024:DPN

[DBP^{+}24] Szymon Drgas, Lars Bramsløw, Archontis Politis, Gaurav Naithani, and Tuomas Virtanen. Dynamic processing neural network architecture for hearing loss compensation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):203– 214, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3328285.

Das:2016:IPP

[DC16] Rajib Lochan Das and Mrityunjoy Chakraborty. Improving the performance of the PNLMS algorithm using norm regularization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(7):1280– 1290, July 2016. CODEN ???? ISSN 2329-9290.

DeLucia:2022:GDE

[DCAS22] Riccardo R. De Lucia, Antonio Canclini, Fabio Antonacci, and Augusto Sarti. Group dictionary equivalent source method for sparse nearfield acoustic holography. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 744–757, 2022. CODEN ???? ISSN 2329-9290.

doCarmo:2023:CFS

[dCBC23] Diego Marques do Carmo, Ricardo A. Borsoi, and Márcio Holsbach Costa. Closed-form solution to the multichannel Wiener filter with interaural level difference preservation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2796– 2811, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3291530.

Do:2018:MLP

[DCLHJ18] V. H. Do, N. F. Chen, B. P. Lim, and M. A. Hasegawa-Johnson. Multitask learning for phone recognition of underresourced languages using mismatched transcription. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):501–514, March 2018. CODEN ???? ISSN 2329-9290.

Duan:2021:MFC

[DCW^{+}21] Chaoqun Duan, Kehai Chen, Rui Wang, Masao Utiyama, Eiichiro Sumita, Conghui Zhu, and Tiejun Zhao. Modeling future cost for neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):770– 781, 2021. CODEN ???? ISSN 2329-9290.

Deng:2022:AAL

[DCYY22] Keqi Deng, Gaofeng Cheng, Runyan Yang, and Yonghong Yan. Alleviating ASR long-tailed problem by decoupling the learning of representation and classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):340–354, 2022. CO-DEN ???? ISSN 2329-9290.

Dang:2019:IMS

[DCZ19] X. Dang, Q. Cheng, and H. Zhu. Indoor multiple sound source localization via multi-dimensional assignment data association. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):1944–1956, December 2019. CODEN ???? ISSN 2329-9290.

Desiraju:2020:OER

[DDBW20] Naveen Kumar Desiraju, Simon Doclo, Markus Buck, and Tobias Wolff. Online estimation of reverberation parameters for late residual echo suppression. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):77–91, January 2020. CODEN ???? ISSN 2329-9290.

URL https://dl.acm.
org/doi/abs/10.1109/
TASLP.2019.2948765.

Desiraju:2023:JOE

[DDBW23] Naveen Kumar Desiraju, Simon Doclo, Markus Buck, and Tobias Wolff. Joint online estimation of early and late residual echo PSD for residual echo suppression. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 333–344, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3224292.

Dietzen:2020:ISC

[DDMvW20a] Thomas Dietzen, Simon Doclo, Marc Moonen, and Toon van Waterschoot. Integrated sidelobe cancellation and linear prediction Kalman filter for joint multi-microphone speech dereverberation, interfering speech cancellation, and noise reduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):740– 754, January 2020. CO-DEN ???? ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.2966869.

Dietzen:2020:SRB

[DDMvW20b] Thomas Dietzen, Simon Doclo, Marc Moonen, and Toon van Waterschoot. Square root-based multisource early PSD estimation and recursive RETF update in reverberant environments by means of the orthogonal Procrustes problem. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):755– 769, January 2020. CO-DEN ???? ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.2966891.

Duong:2019:GMB

[DDNN19] T. T. H. Duong, N. Q. K.

Duong, P. C. Nguyen, and C. Q. Nguyen. Gaussian modeling-based multichannel audio source separation exploiting generic source spectral model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (1):32–43, January 2019. CODEN ???? ISSN 2329-9290.

Duan:2019:TTD

[DDZL19] J. Duan, X. Ding, Y. Zhang, and T. Liu. TEND: A target-dependent representation learning framework for news document. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2313–2325, December 2019. CODEN ???? ISSN 2329-9290.

DiGiorgi:2017:DDM

[DDZS17] Bruno Di Giorgi, Simon Dixon, Massimiliano Zanoni, and Augusto Sarti. A data-driven model of tonal chord sequence complexity. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(11): 2237–2250, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8049362/.

Deng:2014:FEK

[Den14] Deng. Farewell editorial: Keeping up the momentum of innovations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1687, December 2014. CODEN ???? ISSN 2329-9290.

Ding:2024:CAL

[DFY^{+}24] Changhao Ding, Zhangjie Fu, Zhongliang Yang, Qi Yu, Daqiu Li, and Yongfeng Huang. Contextaware linguistic steganography model based on neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):868– 878, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3340601.

Dorfan:2015:TBR

[DG15] Y. Dorfan and S. Gannot. Tree-based recursive expectation-maximization algorithm for localization of acoustic sources. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (10):1692–1703, October 2015. CODEN ???? ISSN 2329-9290.

Dutt:2023:WMA

[DG23] Aditya Dutt and Paul Gader. Wavelet multiresolution analysis based speech emotion recognition system using 1D CNN LSTM networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):2043–2054, 2023. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2023.3277291.

Deppisch:2023:DRS

[DGCA23] Thomas Deppisch, Sebastià V. Amengual Garí, Paul Calamia, and Jens Ahrens. Direct and residual subspace decomposition of spatial room impulse responses. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 927–942, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3240657.

Du:2023:SAT

[DGCY23] Chenpeng Du, Yiwei Guo, Xie Chen, and Kai Yu. Speaker adaptive text-to-speech with timbre-normalized vectorquantized feature. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 3446–3456, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3308374.

Diaz-Guerra:2021:RSS

[DGMB21] David Diaz-Guerra, Antonio Miguel, and Jose R.

Beltran. Robust sound source tracking using SRP-PHAT and 3D convolutional neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):300–311, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3040031.

Diaz-Guerra:2023:DAE

[DGMB23] David Diaz-Guerra, Antonio Miguel, and Jose R. Beltran. Direction of arrival estimation of sound sources using icosahedral CNNs. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):313– 321, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3224282.

Du:2014:IVF

[DH14] Jun Du and Qiang Huo. An improved VTS feature compensation using mixture models of distortion and IVN training for noisy speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (11):1601–1611, November 2014. CODEN ???? ISSN 2329-9290.

Ding:2020:AMS

[DH20] Wenhao Ding and Liang He. Adaptive multiscale detection of acoustic events. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):294– 306, January 2020. CO-DEN ???? ISSN 2329-9290. URL https:/

/dl.acm.org/doi/abs/
10.1109/TASLP.2019.2953350.

DeSena:2015:ESR

[DHCS15] E. De Sena, H. Hacıhabiboğlu, Z. Cvetković, and J. O. Smith. Efficient synthesis of room acoustics via scattering delay networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (9):1478–1492, September 2015. CODEN ???? ISSN 2329-9290.

Deena:2019:RNN

[DHD^{+}19] S. Deena, M. Hasan, M. Doulaty, O. Saz, and T. Hain. Recurrent neural network language model adaptation for multi-genre broadcast speech recognition and alignment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (3):572–582, March 2019. CODEN ???? ISSN 2329-9290.

Duan:2014:MPS

[DHP14] Zhiyao Duan, Jinyu Han, and B. Pardo. Multi-pitch streaming of harmonic sound mixtures. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): 138–150, January 2014. CODEN ???? ISSN 2329-9290.

Deleforge:2015:CLA

[DHSG15] A. Deleforge, R. Horaud, Y. Y. Schechner, and L. Girin. Co-localization of audio sources in images using binaural features and locally-linear regression. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(4):718–731, April 2015. CODEN ???? ISSN 2329-9290.

Davies:2014:AAC

[DHYG14] M. E. P. Davies, P. Hamel, K. Yoshii, and M. Goto. AutoMashUpper: Automatic creation of multisong music mashups. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1726–1737, December 2014. CODEN ???? ISSN 2329-9290.

Deng:2019:SLC

[DJY^{+}19] D. Deng, L. Jing, J. Yu, S. Sun, and M. K. Ng. Sentiment lexicon construction with hierarchical supervision topic model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (4):704–718, April 2019. CODEN ???? ISSN 2329-9290.

Duan:2020:CLT

[DKDN20] Richeng Duan, Tatsuya Kawahara, Masatake Dantsuji, and Hiroaki Nanjo. Crosslingual transfer learning of non-native acoustic modeling for pronunciation error detection and diagnosis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):391– 401, January 2020. CO-DEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2955858.

Delcroix:2018:CAN

[DKO^{+}18] M. Delcroix, K. Kinoshita, A. Ogawa, C. Huemmer, and T. Nakatani. Context adaptive neural network based acoustic models for rapid adaptation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(5):895–908, May 2018. CODEN ???? ISSN 2329-9290.

de-la-Calle-Silos:2015:MFP

[dlCSVAGAPM15] F. de-la Calle-Silos, F. J.

Valverde-Albacete, A. Gallardo-Antolin, and C. Pelaez-Moreno. Morphologically filtered power-normalized cochleograms as robust, biologically inspired features for ASR. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(11): 2070–2080, November 2015. CODEN ???? ISSN 2329-9290.

Degottex:2018:LDP

[DLG18] Gilles Degottex, Pierre Lanchantin, and Mark Gales. A log domain pulse model for parametric speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (1):57–70, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8063442/.

Desblancs:2023:ZNS

[DLH23] Dorian Desblancs, Vincent Lostanlen, and Romain Hennequin. Zeronote samba: Self-supervised beat tracking. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 2922–2934, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3297963.

Ding:2021:RSS

[DLL^{+}21] Yi-Yang Ding, Hao-Jian Lin, Li-Juan Liu, Zhen-Hua Ling, and Yu Hu. Robustness of speech spoofing detectors against adversarial postprocessing of voice conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3415–3426, 2021. CODEN ???? ISSN 2329-9290.

Djaziri-Larbi:2018:WDA

[DLMM^{+}18] Sonia Djaziri-Larbi, Gaël Mahé, Imen Mezghani, Monia Turki, and Mériem Jaïdane. Watermarkdriven acoustic echo cancellation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(2):367– 378, ???? 2018. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/8122007/.

Das:2019:AAW

[DLY^{+}19] A. Das, J. Li, G. Ye, R. Zhao, and Y. Gong. Advancing acoustic-to-word CTC model with attention and mixed-units. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):1880–1892, December 2019. CODEN ???? ISSN 2329-9290.

Dittmar:2016:REA

[DM16] Christian Dittmar and Meinard Müller. Reverse engineering the amen break — score-informed separation and restoration applied to drum recordings. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(9):1535–1547, ???? 2016. CODEN ???? ISSN 2329-9290.

Dang:2022:TBR

[DMHZ22] Xudong Dang, Wen Ma, Emanuël A. P. Habets, and Hongyan Zhu. TDOA-based robust sound source localization with sparse regularization in wireless acoustic sensor networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1108– 1123, 2022. CODEN ???? ISSN 2329-9290.

Dropuljic:2018:VAA

[DMP^{+}18] Branimir Dropuljić, Igor Mijić, Davor Petrinović, Tanja Jovanovic, and Krešimir Ćosić. Vocal analysis of acoustic startle responses. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(2): 318–329, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8115287/.

Dawalatabad:2021:NAU

[DMSM21] Nauman Dawalatabad, Srikanth Madikeri, C. Chandra Sekhar, and Hema A.

Murthy. Novel architectures for unsupervised information bottleneck based speaker diarization of meetings. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 14–27, January 2021.

URL https://dl.acm.
org/doi/10.1109/TASLP.
2020.3036231.

Delikaris-Manias:2016:SDS

[DMVP16] Symeon Delikaris-Manias, Juha Vilkamo, and Ville Pulkki. Signal-dependent spatial filtering based on weighted-orthogonal beamformers in the spherical harmonic domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (9):1511–1523, ???? 2016. CODEN ???? ISSN 2329-9290.

Dalmia:2023:LBM

[DOL^{+}23] Siddharth Dalmia, Dmytro Okhonko, Mike Lewis, Sergey Edunov, Shinji Watanabe, Florian Metze, Luke Zettlemoyer, and Abdelrahman Mohamed. LegoNN: Building modular encoder-decoder models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3112–3126, 2023. ISSN 2329-

9290. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2023.3296019.

Deepak:2016:FSS

[DP16] K. T. Deepak and S. R. Mahadeva Prasanna. Foreground speech segmentation and enhancement using glottal closure instants and mel cepstral coefficients. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(7):1204– 1218, July 2016. CODEN ???? ISSN 2329-9290.

Dorfan:2018:DEM

[DPHG18] Y. Dorfan, A. Plinge, G. Hazan, and S. Gannot. Distributed expectationmaximization algorithm for speaker localization in reverberant environments. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):682–695, March 2018. CODEN ???? ISSN 2329-9290.

Dinkel:2018:IRW

[DQY18] H. Dinkel, Y. Qian, and K. Yu. Investigating raw wave deep neural networks for end-to-end speaker spoofing detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):2002–2014, November 2018. CODEN ???? ISSN 2329-9290.

Devalraju:2022:MES

[DR22] Dhanunjaya Varma Devalraju and Padmanabhan Rajan. Multiview embeddings for soundscape classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1197– 1206, 2022. CODEN ???? ISSN 2329-9290.

Deng:2014:EET

[DRFO14] L. Deng, S. Renals, M. Federico, and M. Ostendorf. Editorial: Expanding the technical reach of our transactions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (1):5, January 2014. CO-DEN ???? ISSN 2329-9290.

Donley:2018:MSR

[DRK18] J. Donley, C. Ritz, and W. B. Kleijn. Multizone soundfield reproduction with privacy- and qualitybased speech masking filters. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(6):1041–1055, June 2018. CODEN ???? ISSN 2329-9290.

deSouza:2024:EBM

[dSCB24] Luciana M. X. de Souza, Márcio H. Costa, and Renata Coelho Borges. Envelope-based multichannel noise reduction for cochlear implant applications. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1873– 1884, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3372891.

Das:2023:GFD

[DSD23] Orchisama Das, Sebastian J. Schlecht, and Enzo De Sena. Grouped feedback delay networks with frequency-dependent coupling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2004– 2015, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3277368.

Dubey:2018:LFD

[DSH18] H. Dubey, A. Sangwan, and J. H. L. Hansen. Leveraging frequencydependent kernel and DIP-based clustering for robust speech activity detection in naturalistic audio streams. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):2056– 2071, November 2018. CODEN ???? ISSN 2329-9290.

Decorsiere:2015:IAS

[DSMD15] R. Decorsiere, P. L.

Søndergaard, E. N. Mac-Donald, and T. Dau. Inversion of auditory spectrograms, traditional spectrograms, and other envelope representations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (1):46–56, January 2015. CODEN ???? ISSN 2329-9290.

Do:2018:SSM

[DSN18] Q. T. Do, S. Sakti, and S. Nakamura. Sequenceto-sequence models for emphasis speech translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(10):1873–1883, October 2018. CODEN ???? ISSN 2329-9290.

Dietzen:2019:CAG

[DST^{+}19] T. Dietzen, A. Spriet, W. Tirry, S. Doclo, M. Moonen, and T. van Waterschoot. Comparative analysis of generalized sidelobe cancellation and multi-channel linear prediction for speech dereverberation and noise reduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(3):544– 558, March 2019. CODEN ???? ISSN 2329-9290.

Dov:2015:AVV

[DTC15] D. Dov, R. Talmon, and I. Cohen. Audiovisual voice activity detection using diffusion maps. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (4):732–745, April 2015. CODEN ???? ISSN 2329-9290.

Dov:2016:KMV

[DTC16] David Dov, Ronen Talmon, and Israel Cohen. Kernel method for voice activity detection in the presence of transients. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2313–2326, ???? 2016. CODEN ???? ISSN 2329-9290.

Dov:2017:MKM

[DTC17] David Dov, Ronen Talmon, and Israel Cohen. Multimodal kernel method for activity detection of sound sources. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (6):1322–1334, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7933049/.

Du:2016:RAS

[DTDL16] Jun Du, Yanhui Tu, Li-Rong Dai, and Chin-Hui Lee. A regression approach to singlechannel speech separation via high-resolution deep neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(8): 1424–1437, ???? 2016. CODEN ???? ISSN 2329-9290.

Dennis:2015:GHT

[DTL15] J. Dennis, H. D. Tran, and Haizhou Li. Generalized Hough transform for speech pattern classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):1963– 1972, November 2015. CODEN ???? ISSN 2329-9290.

Doi:2014:ASE

[DTN^{+}14] H. Doi, T. Toda, K. Nakamura, H. Saruwatari, and K. Shikano. Alaryngeal speech enhancement based on one-to-many eigenvoice conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):172–183, January 2014. CODEN ???? ISSN 2329-9290.

Do:2017:PWL

[DTN^{+}17] Quoc Truong Do, Tomoki Toda, Graham Neubig, Sakriani Sakti, and Satoshi Nakamura. Preserving word-level emphasis in speech-to-speech translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(3):544– 556, ???? 2017. CODEN ???? ISSN 2329-9290.

DAngelo:2014:GMLa

[DV14a] S. D’Angelo and V. Valimaki. Generalized Moog ladder filter: Part I — linear analysis and parameterization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1825– 1832, December 2014. CODEN ???? ISSN 2329-9290.

DAngelo:2014:GMLb

[DV14b] S. D’Angelo and V. Valimaki. Generalized Moog ladder filter: Part II — explicit nonlinear model through a novel delayfree loop implementation method. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1873– 1883, December 2014. CODEN ???? ISSN 2329-9290.

Drakopoulos:2023:NNF

[DV23] Fotios Drakopoulos and Sarah Verhulst. A neuralnetwork framework for the design of individualised hearing-loss compensation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2395– 2409, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3282093.

Drgas:2017:BNN

[DVLH17] Szymon Drgas, Tuomas Virtanen, Jörg Lücke, and Antti Hurmalainen. Binary non-negative matrix deconvolution for audio dictionary learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):1644–1656, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7935444/.

Delcroix:2023:STS

[DVO^{+}23] Marc Delcroix, Jorge Bennasar Vazquez, Tsubasa Ochiai, Keisuke Kinoshita, Yasunori Ohishi, and Shoko Araki. Sound-Beam: Target sound extraction conditioned on sound-class labels and enrollment clues for increased performance and continuous learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):121–136, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2022.3221000.

Defraene:2014:EOB

[DvWDM14] B. Defraene, T. van Waterschoot, M. Diehl, and M. Moonen. Embeddedoptimization-based loudspeaker precompensation using a Hammerstein loudspeaker model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (11):1648–1659, November 2014. CODEN ???? ISSN 2329-9290.

Delfarah:2017:FMB

[DW17] Masood Delfarah and DeLiang Wang. Features for masking-based monaural speech separation in reverberant conditions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(5):1085–1094, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7887742/.

Dong:2016:AMC

[DWX^{+}16] Li Dong, Furu Wei, Ke Xu, Shixia Liu, and Ming Zhou. Adaptive multi-compositionality for recursive neural network models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(3):422– 431, March 2016. CODEN ???? ISSN 2329-9290.

Dinkel:2021:VAD

[DWX^{+}21] Heinrich Dinkel, Shuai Wang, Xuenan Xu, Mengyue Wu, and Kai Yu. Voice activity detection in the wild: a data-driven approach using teacherstudent training. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1542–1555, 2021. CODEN ???? ISSN 2329-9290.

Dinkel:2021:TDR

[DWY21] Heinrich Dinkel, Mengyue Wu, and Kai Yu. Towards duration robust weakly supervised sound event detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 887–900, 2021. CODEN ???? ISSN 2329-9290.

Deng:2023:CSB

[DXW^{+}23] Jiajun Deng, Xurong Xie, Tianzi Wang, Mingyu Cui, Boyang Xue, Zengrui Jin, Guinan Li, Shujie Hu, and Xunying Liu. Confidence score based speaker adaptation of conformer speech recognition systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1175– 1190, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3250842.

Deng:2018:SAS

[DXZ^{+}18] Jun Deng, Xinzhou Xu, Zixing Zhang, Sascha Frühholz, and Björn Schuller. Semisupervised autoencoders for speech emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(1):31–43, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8059872/.

Du:2022:PLP

[DY22] Chenpeng Du and Kai Yu. Phone-level prosody modelling with GMM-based MDN for diverse and controllable speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):190–201, 2022. CO-DEN ???? ISSN 2329-9290.

Deng:2023:LFS

[DYY^{+}23] Shumin Deng, Jiacheng Yang, Hongbin Ye, Chuanqi Tan, Mosha Chen, Songfang Huang, Fei Huang, Huajun Chen, and Ningyu Zhang. LOGEN: Fewshot logical knowledgeconditioned text generation with self-training. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):2124–2133, 2023. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2023.3275028.

Dai:2024:DDN

[DZD^{+}24] Yuhan Dai, Zhirui Zhang, Yichao Du, Shengcai Liu, Lemao Liu, and Tong Xu. Datastore distillation for nearest neighbor machine translation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 807–817, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3337633.

Du:2023:SSC

[DZF^{+}23] Ye-Qian Du, Jie Zhang, Xin Fang, Ming-Hui Wu, and Zhou-Wang Yang. A semi-supervised complementary joint training approach for lowresource speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3908– 3921, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313434.

Ding:2020:LSS

[DZLGO20] Shaojin Ding, Guanlong Zhao, Christopher Liberatore, and Ricardo Gutierrez-Osuna. Learning structured sparse representations for voice conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):343– 354, January 2020. CO-DEN ???? ISSN 2329-9290. URL https:/

                                        /dl.acm.org/doi/abs/
                                        10.1109/TASLP.2019.2955289.

Dai:2020:ASD

[DZWC20] Tao Dai, Li Zhu, Yaxiong Wang, and Kathleen M. Carley. Attentive stacked denoising autoencoder with Bi-LSTM for personalized contextaware citation recommendation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):553– 568, January 2020. CO-DEN ???? ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2019.2949925.

Duan:2023:SAD

[DZZ23] Sufeng Duan, Hai Zhao, and Dongdong Zhang. Syntax-aware data augmentation for neural machine translation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 2988–2999, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3301214.

Erdem:2023:PSR

[ECH23] Ege Erdem, Zoran Cvetković, and Hüseyin Hacıhabiboğlu. 3D perceptual soundfield reconstruction via virtual microphone synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1305– 1317, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3260703.

Edraki:2021:SIP

[ECJF21] Amin Edraki, Wai-Yip Chan, Jesper Jensen, and Daniel Fogerty. Speech intelligibility prediction using spectro-temporal modulation analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):210–225, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3039929.

ElBadawy:2018:DAO

[ED18] D. El Badawy and I. Dokmanić. Direction of arrival with one microphone, a few LEGOs, and nonnegative matrix factorization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2436–2446, December 2018. CODEN ???? ISSN 2329-9290.

ElBadawy:2017:FAS

[EDO17] Dalia El Badawy, Ngoc Q. K. Duong, and Alexey Ozerov. On-the-fly audio source separation — a novel user-friendly framework. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(2):261–272, ???? 2017. CODEN ???? ISSN 2329-9290.

Eaton:2016:ERA

[EGMN16] James Eaton, Nikolay D. Gaubitch, Alastair H. Moore, and Patrick A. Naylor. Estimation of room acoustic parameters: The ACE challenge. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1681–1693, ???? 2016. CODEN ???? ISSN 2329-9290.

Elisei-Iliescu:2019:RLS

[EIPB^{+}19] C. Elisei-Iliescu, C. Paleologu, J. Benesty, C. Stanciu, C. Anghel, and S. Ciochină. Recursive least-squares algorithms for the identification of low-rank systems. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(5): 903–918, May 2019. CO-DEN ???? ISSN 2329-9290.

Epain:2016:SHS

[EJ16] Nicolas Epain and Craig T. Jin. Spherical harmonic signal covariance and sound field diffuseness. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1796–1807, ???? 2016. CODEN ???? ISSN 2329-9290.

Elshamy:2017:IPS

[EMTF17] Samy Elshamy, Nilesh Madhu, Wouter Tirry, and Tim Fingscheidt. Instantaneous a priori SNR estimation by cepstral excitation manipulation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):1592–1605, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7922539/.

Elshamy:2018:DSS

[EMTF18] S. Elshamy, N. Madhu, W. Tirry, and T. Fingscheidt. DNN-supported speech enhancement with cepstral estimation of both excitation and envelope. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2460–2474, December 2018. CODEN ???? ISSN 2329-9290.

Emura:2018:RER

[Emu18] S. Emura. Residual echo reduction for multichannel acoustic echo cancelers with a complex-valued residual echo estimate. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):485–500, March 2018.

CODEN ???? ISSN 2329-

9290.

Emura:2020:WDR

[Emu20] Satoru Emura. Wavedomain residual echo reduction using subspace tracking. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):144– 156, January 2020. CO-DEN ???? ISSN 2329-9290. URL https:/

                                        /dl.acm.org/doi/abs/
                                        10.1109/TASLP.2019.2949709.

Eskimez:2020:NRT

[EMXD20] Sefik Emre Eskimez, Ross K. Maddox, Chenliang Xu, and Zhiyao Duan. Noise-resilient training method for face landmark generation from speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):27–38, January 2020. CODEN ???? ISSN 2329-9290.

                                        URL https://dl.acm.
                                        org/doi/abs/10.1109/
                                        TASLP.2019.2947741.

Evers:2018:AS

[EN18] C. Evers and P. A. Naylor. Acoustic SLAM. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (9):1484–1498, September 2018. CODEN ???? ISSN 2329-9290.

Ewert:2016:PTS

[ES16] Sebastian Ewert and Mark Sandler. Piano transcription in the studio using an extensible alternating directions framework. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):1983–1997, ???? 2016. CODEN ???? ISSN 2329-9290.

Elvander:2017:OEM

[ESJ17] Filip Elvander, Johan Swärd, and Andreas Jakobsson. Online estimation of multiple harmonic signals. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(2): 273–284, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Enarvi:2017:ASR

[ESVK17] Seppo Enarvi, Peter Smit, Sami Virpioja, and Mikko Kurimo. Automatic speech recognition with very large conversational Finnish and Estonian vocabularies. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2085– 2097, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/8015196/.

Enzner:2018:BMF

[ET18] G. Enzner and P. Thüne. Bayesian MMSE filtering of noisy speech by SNR marginalization with global PSD priors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2289–2304, December 2018. CODEN ???? ISSN 2329-9290.

Enzner:2023:HFR

[EV23] Gerald Enzner and Svantje Voit. Hybrid- frequencyresolution adaptive Kalman filter for online identification of long acoustic responses with low inputoutput latency. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 3550–3563, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3313435.

Erro:2014:EIS

[EZS14] D. Erro, T.-C. Zorila, and Y. Stylianou. Enhancing the intelligibility of statistically generated synthetic speech by means of noise-independent modifications. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2101– 2111, December 2014. CODEN ???? ISSN 2329-9290.

Fontana:2018:EFP

[FB18] F. Fontana and E. Bozzo.

Explicit fixed-point computation of nonlinear delay-free loop filter networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(10):1884– 1896, October 2018. CO-DEN ???? ISSN 2329-9290.

Fontana:2019:NRS

[FB19] F. Fontana and E. Bozzo.

Newton–Raphson solution of nonlinear delayfree loop filter networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (10):1590–1600, October 2019. CODEN ???? ISSN 2329-9290.

Flessner:2019:SOA

[FBE19] J. Fleßner, T. Biberger, and S. D. Ewert. Subjective and objective assessment of monaural and binaural aspects of audio quality. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(7):1112– 1125, July 2019. CODEN ???? ISSN 2329-9290.

Fang:2023:IUN

[FBWG23] Huajian Fang, Dennis Becker, Stefan Wermter, and Timo Gerkmann. Integrating uncertainty into neural networkbased speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1587–1600, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3265202.

Fan:2023:CAB

[FCCA23] Ruchao Fan, Wei Chu, Peng Chang, and Abeer Alwan. A CTC alignmentbased non-autoregressive transformer for end-to-end automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1436– 1448, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3263789.

Frenkel:2024:DAU

[FCG24] Lior Frenkel, Shlomo E. Chazan, and Jacob Goldberger. Domain adaptation using suitable pseudo labels for speech enhancement and dereverberation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1226– 1236, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3358051.

Fischer:2021:RCM

[FD21] Dörte Fischer and Simon Doclo. Robust constrained MFMVDR filters for single-channel speech enhancement based on spherical uncertainty set. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):618–631, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3042013.

Fuster:2016:AFX

[FdDARF16] Laura Fuster, Maria de Diego, Luis A. Azpicueta-Ruiz, and Miguel Ferrer. Adaptive filteredx algorithms for room equalization based on block-based combination schemes. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1732– 1745, ???? 2016. CODEN ???? ISSN 2329-9290.

Ferrer:2021:APA

[FdDPG21] Miguel Ferrer, Maria de Diego, Gema Piñero, and Alberto Gonzalez. Affine projection algorithm over acoustic sensor networks for active noise control. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):448–461, January

2021. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2020.3042590.

Foster:2015:ICS

[FDK15] P. Foster, S. Dixon, and A. Klapuri. Identifying cover songs using information-theoretic measures of similarity. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (6):993–1005, June 2015. CODEN ???? ISSN 2329-9290.

Fonseca:2022:FOD

[FFP^{+}22] Eduardo Fonseca, Xavier Favory, Jordi Pons, Frederic Font, and Xavier Serra. FSD50K: an open dataset of human-labeled sound events. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 829–852, 2022. CODEN ???? ISSN 2329-9290.

Firtha:2017:IRS

[FFSS17] Gergely Firtha, Péter Fiala, Frank Schultz, and Sascha Spors. Improved referencing schemes for 2.5D wave field synthesis driving functions. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(5): 1117–1127, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7888944/.

Firtha:2018:GRW

[FFSS18] G. Firtha, P. Fiala, F. Schultz, and S. Spors. On the general relation of wave field synthesis and spectral division method for linear arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2393–2403, December 2018. CODEN ???? ISSN 2329-9290.

Fu:2015:LSH

[FGQ^{+}15] Ruiji Fu, Jiang Guo, Bing Qin, Wanxiang Che, Haifeng Wang, and Ting Liu. Learning semantic hierarchies: a continuous vector space approach. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (3):461–471, March 2015. CODEN ???? ISSN 2329-9290.

Fuglsig:2023:MPN

[FJT^{+}23] Andreas Jonas Fuglsig, Jesper Jensen, Zheng-Hua Tan, Lars Søndergaard Bertelsen, Jens Christian Lindof, and Jan Østergaard. Minimum processing near-end listening enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2233– 2245, 2023. ISSN 2329-

9290. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2023.3282094.

Feng:2019:URB

[FK19] F. Feng and M. Kowalski. Underdetermined reverberant blind source separation: Sparse approaches for multiplicative and convolutive narrowband approximation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(2): 442–456, February 2019. CODEN ???? ISSN 2329-9290.

Fras:2024:RSS

[FK24] Mieszko Fraś and Konrad Kowalczyk. Reverberant source separation using NTF with delayed subsources and spatial priors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1954– 1967, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3374065.

Feng:2019:ECL

[FL19] S. Feng and T. Lee. Exploiting cross-lingual speaker and phonetic diversity for unsupervised subword modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2000–2011, December 2019. CODEN ???? ISSN 2329-9290.

FitzGerald:2016:PBD

[FLB16] Derry FitzGerald, Antoine Liutkus, and Roland Badeau. Projection-based demixing of spatial audio. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (9):1560–1572, ???? 2016. CODEN ???? ISSN 2329-9290.

Fu:2023:CLN

[FLC^{+}23] Yingwen Fu, Nankai Lin, Boyu Chen, Ziyu Yang, and Shengyi Jiang. Cross-lingual named entity recognition for heterogeneous languages. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 371–382, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3212698.

Fang:2023:CCS

[FLL^{+}23] Yan Fang, Wei Lu, Xiaodong Liu, Witold Pedrycz, Qi Lang, and Jianhua Yang. CircularE: a complex space circular correlation relational model for link prediction in knowledge graph embedding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3162– 3175, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3297959.

Feng:2023:PBS

[FLM23] Huawen Feng, Zhenxi Lin, and Qianli Ma. Perturbation-based selfsupervised attention for attention bias in text classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3139– 3151, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3302230.

Ferrer:2016:SSB

[FLMS16] L. Ferrer, Yun Lei, M. McLaren, and N. Scheffer. Study of senone-based deep neural network approaches for spoken language recognition. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(1): 105–116, January 2016. CODEN ???? ISSN 2329-9290.

Fu:2023:STD

[FLYJ23] Yingwen Fu, Nankai Lin, Xiaohui Yu, and Shengyi Jiang. Self-training with double selectors for lowresource named entity recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1265– 1275, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3250828.

Fallahpour:2015:AWB

[FM15] M. Fallahpour and D. Megias.

Audio watermarking based on Fibonacci numbers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (8):1273–1282, August 2015. CODEN ???? ISSN 2329-9290.

Ferras:2016:SDL

[FMB16] Marc Ferràs, Srikanth Madikeri, and Hervé Bourlard. Speaker diarization and linking of meeting data. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(11): 1935–1945, ???? 2016. CODEN ???? ISSN 2329-9290.

Foster:2014:SCD

[FMD14] P. Foster, M. Mauch, and S. Dixon. Sequential complexity as a descriptor for musical similarity. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1965–1977, December 2014. CODEN ???? ISSN 2329-9290.

Fujioka:2015:HPH

[FNA15] T. Fujioka, Y. Nagata, and M. Abe. Highprecision harmonic distortion level measurement of a loudspeaker using adaptive filters in a noisy environment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1613– 1622, October 2015. CO-DEN ???? ISSN 2329-9290.

Ferrer:2019:TFS

[FNM^{+}19] L. Ferrer, M. K. Nandwana, M. McLaren, D. Castan, and A. Lawson. Toward fail-safe speaker recognition: Trial-based calibration with a reject option. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1): 140–153, January 2019. CODEN ???? ISSN 2329-9290.

Fang:2015:ELM

[FOBP15] Hao Fang, M. Ostendorf, P. Baumann, and J. Pierrehumbert. Exponential language modeling using morphological features and multi-task learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2410–2421, December 2015. CODEN ???? ISSN 2329-9290.

Farmani:2017:ISS

[FPTJ17] Mojtaba Farmani, Michael Syskind Pedersen, Zheng-Hua Tan, and Jesper Jensen. Informed sound source localization using relative transfer functions for hearing aid applications. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (3):611–623, ???? 2017. CODEN ???? ISSN 2329-9290.

Farmani:2018:BCI

[FPTJ18] M. Farmani, M. S. Pedersen, Z. Tan, and J. Jensen. Bias-compensated informed sound source localization using relative transfer functions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(7):1275– 1289, July 2018. CODEN ???? ISSN 2329-9290.

Fahim:2018:PES

[FSA18] A. Fahim, P. N. Samarasinghe, and T. D. Abhayapala. PSD estimation and source separation in a noisy reverberant environment using a spherical microphone array. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1594–1607, September 2018. CODEN ???? ISSN 2329-9290.

Fahim:2020:MSD

[FSA20] Abdullah Fahim, Prasanga N.

Samarasinghe, and Thushara D. Abhayapala. Multisource DOA estimation through pattern recognition of the modal coherence of a reverberant soundfield. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 605–618, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2960734.

Furnon:2021:DBM

[FSEI21] Nicolas Furnon, Romain Serizel, Slim Essid, and Irina Illina. DNN-based mask estimation for distributed speech enhancement in spatially unconstrained microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2310–2323, 2021. CODEN ???? ISSN 2329-9290.

Fukuda:2024:IST

[FSN24] Ryo Fukuda, Katsuhito Sudoh, and Satoshi Nakamura. Improving speech translation accuracy and time efficiency with finetuned wav2vec 2.0-based speech segmentation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 906–916, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3343614.

Fakhry:2017:ASS

[FSO17] Mahmoud Fakhry, Piergiorgio Svaizer, and Maurizio Omologo. Audio source separation in reverberant environments using β-divergence-based nonnegative factorization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (7):1462–1476, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7904611/.

Franck:2017:SOM

[FWF17] Andreas Franck, Wenwu Wang, and Filippo Maria Fazi. Sparse ℓ_{1}-optimal multiloudspeaker panning and its relation to vector base amplitude panning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):996–1010, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7864372/.

FanChiang:2014:CLC

[FWM^{+}14a] Y. FanChiang, C.-W.

Wei, Y.-L. Meng, Y.- W. Lin, S.-J. Jou, and T.-S. Chang. Correction to “Low Complexity Formant Estimation Adaptive Feedback Cancellation for Hearing Aids Using Pitch Based Processing” [Aug 14 1248– 1259]. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2256, December 2014. CODEN ???? ISSN 2329-9290. See [FWM^{+}14b].

FanChiang:2014:LCF

[FWM^{+}14b] Yi FanChiang, Cheng-Wen Wei, Yi-Le Meng, Yu-Wen Lin, Shyh-Jye Jou, and Tian-Sheuan Chang. Low complexity formant estimation adaptive feedback cancellation for hearing aids using pitch based processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):1248–1259, August 2014. CODEN ???? ISSN 2329-9290. See correction [FWM^{+}14a].

Fei:2021:SOS

[FWRJ21] Hao Fei, Shengqiong Wu, Yafeng Ren, and Donghong Ji. Secondorder semantic role labeling with global structural refinement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1966–1976, 2021. CODEN ???? ISSN 2329-9290.

Fu:2018:EEW

[FWT^{+}18] S. Fu, T. Wang, Y. Tsao, X. Lu, and H. Kawai.

End-to-end waveform utterance enhancement for direct evaluation metrics optimization by fully convolutional neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1570– 1584, September 2018. CODEN ???? ISSN 2329-9290.

Fan:2021:MTS

[FYG^{+}21] Chuang Fan, Chaofa Yuan, Lin Gui, Yue Zhang, and Ruifeng Xu. Multi-task sequence tagging for emotion-cause pair extraction via tag distribution refinement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):2339–2350, 2021. CODEN ???? ISSN 2329-9290.

Fan:2021:GRF

[FYT^{+}21] Cunhang Fan, Jiangyan Yi, Jianhua Tao, Zhengkun Tian, Bin Liu, and Zhengqi Wen. Gated recurrent fusion with joint training framework for robust endto-end speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):198–209, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3039600.

Feng:2024:MHT

[FZM24] Sheng Feng, Xiaoqian Zhu, and Shuqing Ma. Masking hierarchical tokens for underwater acoustic target recognition with self-supervised learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1365–1379, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3358719.

Feng:2015:TGS

[FZZ^{+}15] Zu-Ren Feng, Qing Zhou, Jun Zhang, Ping Jiang, and Xue-Wen Yang. A target guided subband filter for acoustic event detection in noisy environments using wavelet packets. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(2):361–372, February 2015. CODEN ???? ISSN 2329-9290.

Ghasemzadeh:2020:TOQ

[GA20] Hamzeh Ghasemzadeh and Meisam K. Arjmandi. Toward optimum quantification of pathologyinduced noises: an investigation of information missed by human auditory system. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??):

519–528, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2959222.

Gaultier:2021:SBA

[GaKGB21] Clément Gaultier, Sr an Kiti , Rémi Gribonval, and Nancy Bertin. Sparsitybased audio declipping methods: Selected overview, new algorithms, and large-scale evaluation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):1174–1187, 2021. CODEN ???? ISSN 2329-9290.

Gonzalez:2023:AGG

[GAM23] Philippe Gonzalez, Tommy Sonne Alstrøm, and Tobias May. Assessing the generalization gap of learningbased speech enhancement systems in noisy and reverberant environments. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3390– 3403, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3318965.

Gomez-Alanis:2019:GRC

[GAPGG19] A. Gomez-Alanis, A. M.

Peinado, J. A. Gonzalez, and A. M. Gomez. A gated recurrent convolutional neural network for robust spoofing detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):1985–1999, December 2019. CODEN ???? ISSN 2329-9290.

Gonzalez:2014:PPE

[GB14] S. Gonzalez and M. Brookes. PEFAC — a pitch estimation algorithm robust to high levels of noise. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):518–530, February 2014. CODEN ???? ISSN 2329-9290.

Gosztolya:2021:EBA

[GBF21] Gábor Gosztolya and Róbert Busa-Fekete. Ensemble bag-of-audio-words representation improves paralinguistic classification accuracy. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):477–488, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3044465.

Ghaffarzadegan:2016:GMP

[GBH16] Shabnam Ghaffarzadegan, Hynek Bořil, and John H. L. Hansen. Generative modeling of pseudo-whisper for robust whispered speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1705–1720, ???? 2016. CODEN ???? ISSN 2329-9290.

Gao:2024:ACS

[GBL24] Yingming Gao, Peter Birkholz, and Ya Li. Articulatory copy synthesis based on the speech synthesizer VocalTractLab and convolutional recurrent neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1845– 1858, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3372874.

Gonzalez:2017:DSR

[GCG^{+}17] Jose A. Gonzalez, Lam A.

Cheah, Angel M. Gomez, Phil D. Green, James M. Gilbert, Stephen R. Ell, Roger K. Moore, and Ed Holdsworth. Direct speech reconstruction from articulatory sensor data by machine learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(12):2362–2374, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8114382/.

Gong:2021:PIA

[GCG21] Yuan Gong, Yu-An Chung, and James Glass. PSLA:

Improving audio tagging with pretraining, sampling, labeling, and aggregation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3292– 3306, 2021. CODEN ???? ISSN 2329-9290.

Giacobello:2014:SNE

[GCJ^{+}14] D. Giacobello, M. G. Christensen, T. L. Jensen, M. N. Murthi, S. H. Jensen, and M. Moonen. Stable 1-norm error minimization based linear predictors for speech modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(5):912–922, May 2014. CODEN ???? ISSN 2329-9290.

Gil-Cacho:2014:FDA

[GCvWMJ14] J. M. Gil-Cacho, T. van Waterschoot, M. Moonen, and S. H. Jensen. A frequency-domain adaptive filter (FDAF) prediction error method (PEM) framework for doubletalk-robust acoustic echo cancellation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2074– 2086, December 2014. CODEN ???? ISSN 2329-9290.

Giampiccolo:2021:WDM

[GdBBS21] Riccardo Giampiccolo, Mauro Giuseppe de Bari, Alberto Bernardini, and Augusto Sarti. Wave digital modeling and implementation of nonlinear audio circuits with nullors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3267–3279, 2021. CODEN ???? ISSN 2329-9290.

Guldenschuh:2014:DSP

[GdC14] M. Guldenschuh and R. de Callafon. Detection of secondary-path irregularities in active noise control headphones. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(7): 1148–1157, July 2014. CODEN ???? ISSN 2329-9290.

Gao:2019:MBC

[GDC19] J. Gao, J. Du, and E. Chen. Mixed-bandwidth cross-channel speech recognition via joint optimization of DNN-based bandwidth expansion and acoustic modeling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(3): 559–571, March 2019. CODEN ???? ISSN 2329-9290.

Gully:2018:DSU

[GDM18] Amelia J. Gully, Helena Daffern, and Damian T.

Murphy. Diphthong synthesis using the dynamic 3D digital waveguide mesh. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(2):243–255, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8114217/.

Galvez:2015:TDO

[GEC15] M. F. Simon Galvez, S. J. Elliott, and J. Cheer. Time domain optimization of filters used in a loudspeaker array for personal audio. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):1869– 1878, November 2015. CODEN ???? ISSN 2329-9290.

Gangeh:2014:MSD

[GFG^{+}14] M. J. Gangeh, P. Fewzee, A. Ghodsi, M. S. Kamel, and F. Karray. Multiview supervised dictionary learning in speech emotion recognition. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(6): 1056–1068, June 2014. CODEN ???? ISSN 2329-9290.

Gao:2022:WLP

[GFS22] Hao Gao, Xuelei Feng, and Yong Shen. Weighted loudspeaker placement method for sound field reproduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1263– 1276, 2022. CODEN ???? ISSN 2329-9290.

Gelly:2018:ORB

[GG18] G. Gelly and J. Gauvain. Optimization of RNN-based speech activity detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(3): 646–656, March 2018. CODEN ???? ISSN 2329-9290.

Gao:2023:PIF

[GGL23] Xiaoxue Gao, Chitralekha Gupta, and Haizhou Li. PoLyScriber: Integrated fine-tuning of extractor and lyrics transcriber for polyphonic music. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1968– 1981, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3275036.

Gu:2023:MSN

[GGZ23] Bin Gu, Wu Guo, and Jie Zhang. Memory storable network based feature aggregation for speaker representation learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):643– 655, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3231709.

Ghahabi:2017:DLB

[GH17] Omid Ghahabi and Javier Hernando. Deep learning backend for single and multisession ivector speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(4):807–817, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7847321/.

Ghorbani:2023:DEE

[GH23a] Shahram Ghorbani and John H. L. Hansen. Domain expansion for endto-end speech recognition: Applications for Accent/ dialect speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 762–774, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3233238.

Ghosh:2023:BCI

[GH23b] Ria Ghosh and John H. L. Hansen. Bilateral cochlear implant processing of coding strategies with CCi-MOBILE, an open-source research platform. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1839– 1850, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3267608.

Girin:2017:ECG

[GHAP17] Laurent Girin, Thomas Hueber, and Xavier Alameda-Pineda. Extending the cascaded Gaussian mixture regression framework for cross-speaker acousticarticulatory mapping. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (3):662–673, ???? 2017. CODEN ???? ISSN 2329-9290.

Guo:2023:CSF

[GHJ^{+}23] Peiming Guo, Shen Huang, Peijie Jiang, Yueheng Sun, Meishan Zhang, and Min Zhang. Curriculumstyle fine-grained adaption for unsupervised cross-lingual dependency transfer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):322– 332, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3224302.

Ge:2021:LFG

[GHS^{+}21] Jidong Ge, Yunyun Huang, Xiaoyu Shen, Chuanyi Li, and Wei Hu. Learning fine-grained fact-article correspondence in legal cases. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3694–3706, 2021. CODEN ???? ISSN 2329-9290.

Gerazov:2015:KPF

[GI15] B. Gerazov and Z. Ivanovski. Kernel power flow orientation coefficients for noiserobust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(2):407–419, February 2015. CODEN ???? ISSN 2329-9290.

Grozdic:2017:WSR

[GJ17] – Dor– de T. Grozdić and Slobodan T. Jovičić. Whispered speech recognition using deep denoising autoencoder and inverse filtering. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12): 2313–2322, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8114355/.

Grinewitschus:2022:HSA

[GJ22] Lukas Grinewitschus and Peter Jung. The harmonic shift algorithm for efficient multi-pitch detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):548–561, 2022. CODEN ???? ISSN 2329-9290.

Graja:2015:SFK

[GJB15] M. Graja, M. Jaoua, and L. H. Belguith. Statistical framework with knowledge base integration for robust speech understanding of the Tunisian dialect. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2311–2321, December 2015. CODEN ???? ISSN 2329-9290.

Gao:2023:MSL

[GJYW23] Shang Gao, Maoshen Jia, Dingding Yao, and Jing Wang. Multisource localization using optimized time-frequency representation and sparsity component analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3564–3578, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3316450.

Gkiokas:2016:TMP

[GKC16] Aggelos Gkiokas, Vassilis Katsouros, and George Carayannis. Towards multi-purpose spectral rhythm features: An application to dance style, meter and tempo estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):1885–1896, ???? 2016. CODEN ???? ISSN 2329-9290.

Gao:2014:SSF

[GLC^{+}14] Dehong Gao, Wenjie Li, Xiaoyan Cai, Renxian Zhang, and You Ouyang. Sequential summarization: a full view of Twitter trending topics. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): 293–302, February 2014. CODEN ???? ISSN 2329-9290.

Gu:2020:UUI

[GLL20] Jia-Chen Gu, Zhen-Hua Ling, and Quan Liu. Utterance-to-utterance interactive matching network for multi-turn response selection in retrievalbased chatbots. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 369–379, January 2020. CODEN ???? ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/abs/
                                        10.1109/TASLP.2019.2955290.

Gu:2021:DCU

[GLL^{+}21] Jia-Chen Gu, Tianda Li, Zhen-Hua Ling, Quan Liu, Zhiming Su, Yu-Ping Ruan, and Xiaodan Zhu. Deep contextualized utterance representations for response selection and dialogue analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2443–2455, 2021. CODEN ???? ISSN 2329-9290.

Gao:2016:SSA

[GLQ16] Min Gao, Jing Lu, and Xiaojun Qiu. A simplified subband ANC algorithm without secondary path modeling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(7): 1164–1174, July 2016. CODEN ???? ISSN 2329-9290.

Ge:2021:PMP

[GLQ21] Zhongshu Ge, Liang Li, and Tianshu Qu. Partially matching projection decoding method evaluation under different playback conditions. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1411–1423, 2021. CODEN ???? ISSN 2329-9290.

Galassi:2023:MTA

[GLT23] Andrea Galassi, Marco Lippi, and Paolo Torroni. Multi-task attentive residual networks for argument mining. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1877– 1892, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3275040.

Gupta:2020:ALE

[GLW20] Chitralekha Gupta, Haizhou Li, and Ye Wang. Automatic leaderboard: Evaluation of singing quality without a standard reference. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):13–26, January 2020. CODEN ???? ISSN 2329-9290.

                                        URL https://dl.acm.
                                        org/doi/abs/10.1109/
                                        TASLP.2019.2947737.

Gao:2022:SDM

[GLWQ22] Shan Gao, Jing Lin, Xihong Wu, and Tianshu Qu. Sparse DNN model for frequency expanding of higher order ambisonics encoding process. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1124–1135, 2022. CODEN ???? ISSN 2329-9290.

Gossling:2021:PAE

[GMD21] Nico Gößling, Daniel Marquardt, and Simon Doclo. Performance analysis of the extended binaural MVDR beamformer with partial noise estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):462–476, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3043674.

Gao:2024:LTE

[GMD24] Lijian Gao, Qirong Mao, and Ming Dong. On local temporal embedding for semi-supervised sound event detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 1687–1698, 2024. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2024.3369529.

Grijalva:2016:MLA

[GMFG16] Felipe Grijalva, Luiz Martini, Dinei Florencio, and Siome Goldenstein. A manifold learning approach for personalizing HRTFs from anthropometric features. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): 559–570, March 2016. CODEN ???? ISSN 2329-9290.

Ganapathy:2014:RFE

[GMH14] S. Ganapathy, S. H.

Mallidi, and H. Hermansky. Robust feature extraction using modulation filtering of autoregressive models. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(8): 1285–1295, August 2014. CODEN ???? ISSN 2329-9290.

Granell:2017:MCT

[GMH17] Emilio Granell and Carlos-D. Martínez-Hinarejos. Multimodal crowdsourcing for transcribing handwritten documents. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(2): 409–419, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Gajecki:2023:DLF

[GN23] Tom Gajecki and Waldo Nogueira. Deep latent fusion layers for binaural speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3127–3138, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3301223.

Grzeszick:2017:BFM

[GPF17] René Grzeszick, Axel Plinge, and Gernot A. Fink. Bag-of-features methods for acoustic event detection and classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(6):1242– 1252, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7933055/.

Guo:2019:LRL

[GQXZ19] Q. Guo, X. Qiu, X. Xue, and Z. Zhang. Low-rank and locality constrained self-attention for sequence modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2213– 2222, December 2019. CODEN ???? ISSN 2329-9290.

Ghalehjegh:2014:LRB

[GR14] S. H. Ghalehjegh and R. C. Rose. Linear regression based acoustic adaptation for the subspace Gaussian mixture model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (9):1391–1402, September 2014. CODEN ???? ISSN 2329-9290.

Grais:2017:TSS

[GRSP17] Emad M. Grais, Gerard Roma, Andrew J. R. Simpson, and Mark D. Plumbley. Two-stage single-channel audio source separation using deep neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(9): 1773–1783, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7951020/.

Gunnarsson:2021:BAM

[GS21] Viktor Gunnarsson and Mikael Sternad. Binaural auralization of microphone array room impulse responses using causal Wiener filtering. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2899–2914, 2021. CODEN ???? ISSN 2329-9290.

Geronazzo:2018:DWN

[GSA18] M. Geronazzo, S. Spagnol, and F. Avanzini. Do we need individual headrelated transfer functions for vertical localization? the case study of a spectral notch distance metric. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (7):1247–1260, July 2018. CODEN ???? ISSN 2329-9290.

Gotz:2023:CSM

[GSP23] Georg Götz, Sebastian J. Schlecht, and Ville Pulkki. Common-slope modeling of late reverberation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3945–3957, 2023. ISSN 2329-

                                        9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2023.3317572.

Gabrielli:2019:MSA

[GTS^{+}19] L. Gabrielli, S. Tomassetti, S. Squartini, C. Zinato, and S. Guaiana. A multi-stage algorithm for acoustic physical model parameters estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (8):1229–1240, August 2019. CODEN ???? ISSN 2329-9290.

Gelderblom:2024:PPO

[GTSM24] Femke B. Gelderblom, Tron Vedul Tronstad, Torbjørn Svendsen, and Tor Andre Myrvoll. On the predictive power of objective intelligibility metrics for the subjective performance of deep complex convolutional recurrent speech enhancement networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):215– 226, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3329378.

Gelderblom:2019:SEN

[GTV19] F. B. Gelderblom, T. V.

Tronstad, and E. M. Viggen. Subjective evaluation of a noisereduced training target for deep neural networkbased speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(3):583–594, March 2019. CODEN ???? ISSN 2329-9290.

Goli:2023:DLB

[GvdP23] Peyman Goli and Steven van de Par. Deep learningbased speech specific source localization by using binaural and monaural microphone arrays in hearing aids. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1652– 1666, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3268734.

Gannot:2017:CPM

[GVMGO17] Sharon Gannot, Emmanuel Vincent, Shmulik Markovich-Golan, and Alexey Ozerov. A consolidated perspective on multimicrophone speech enhancement and source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(4):692– 730, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

                                        ieeexplore.ieee.org/
                                        document/7805139/.

Gao:2023:BST

[GWCT23] Nan Gao, Yongjian Wang, Peng Chen, and Jijun Tang. Boosting short text classification by solving the OOV problem. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):4014– 4024, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3316422.

Geiger:2014:MEN

[GWG^{+}14] J. T. Geiger, F. Weninger, J. F. Gemmeke, M. Wollmer, B. Schuller, and G. Rigoll. Memory-enhanced neural networks and NMF for robust ASR. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(6):1037– 1046, June 2014. CODEN ???? ISSN 2329-9290.

Gong:2024:ALC

[GWL^{+}24] Xun Gong, Yu Wu, Jinyu Li, Shujie Liu, Rui Zhao, Xie Chen, and Yanmin Qian. Advanced longcontent speech recognition with factorized neural transducer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1803– 1815, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3350893.

Geng:2022:MLT

[GWLL22] Jianhua Geng, Sifan Wang, Qinglai Liu, and Xin Lou. Multi-level timefrequency bins selection for direction of arrival estimation using a single acoustic vector sensor. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1048–1060, 2022. CODEN ???? ISSN 2329-9290.

Gao:2023:PMB

[GWQ23] Shan Gao, Xihong Wu, and Tianshu Qu. A physical model-based selfsupervised learning method for signal enhancement under reverberant environment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2100– 2110, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3277246.

Guizzo:2023:LSE

[GWSC23] Eric Guizzo, Tillman Weyde, Simone Scardapane, and Danilo Comminiello. Learning speech emotion representations in the quaternion domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1200–1212, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3250840.

Guo:2023:MTM

[GXW^{+}23] Haohan Guo, Fenglong Xie, Xixin Wu, Frank K. Soong, and Helen Meng. MSMC-TTS: Multi-stage multi-codebook VQ-VAE based neural TTS. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1811–1824, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3272470.

Gasic:2014:GPP

[GY14] M. Gasic and S. Young.

Gaussian processes for POMDP-based dialogue manager optimization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (1):28–40, January 2014. CODEN ???? ISSN 2329-9290.

Guo:2021:PNS

[GYdL^{+}21] Gang Guo, Yi Yu, Rodrigo C. de Lamare, Zongsheng Zheng, Lu Lu, and Qiangming Cai. Proximal normalized subband adaptive filtering for acoustic echo cancellation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2174–2188, 2021. CODEN ???? ISSN 2329-

Guo:2023:LLP

[GYXY23] Junjun Guo, Junjie Ye, Yan Xiang, and Zhengtao Yu. Layer-level progressive transformer with modality difference awareness for multimodal neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3015– 3026, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3301210.

Gu:2023:DCF

[GZG23] Bin Gu, Jie Zhang, and Wu Guo. A dynamic convolution framework for session-independent speaker embedding learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3647–3658, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313431.

Guo:2021:AAE

[GZX^{+}21] Junliang Guo, Zhirui Zhang, Linli Xu, Boxing Chen, and Enhong Chen. Adaptive adapters: an efficient way to incorporate BERT into neural machine translation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??):

1740–1751, 2021. CODEN ???? ISSN 2329-9290.

Gu:2023:TUA

[GZZY23] Rongzhi Gu, Shi-Xiong Zhang, Yuexian Zou, and Dong Yu. Towards unified all-neural beamforming for time and frequency domain speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):849– 862, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3229261.

Hacihabiboglu:2014:TAO

[Hac14] Hüseyim Hacıhabiboğlu.

Theoretical analysis of open spherical microphone arrays for acoustic intensity measurements. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (2):465–476, February 2014. CODEN ???? ISSN 2329-9290.

Helmholz:2021:EAN

[HAGA21] Hannes Helmholz, David Lou Alon, Sebastià V. Amengual Garí, and Jens Ahrens. Effects of additive noise in binaural rendering of spherical microphone array signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3642–3653, 2021.

CODEN ???? ISSN 2329-9290.

Hajimolahoseini:2016:RET

[HAGSZ16] Habib Hajimolahoseini, Rassoul Amirfattahi, Saeed Gazor, and Hamid Soltanian-Zadeh. Robust estimation and tracking of pitch period using an efficient Bayesian filter. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(7): 1219–1229, July 2016. CODEN ???? ISSN 2329-9290.

Hu:2021:MSD

[HAS21] Yonggang Hu, Thushara D. Abhayapala, and Prasanga N. Samarasinghe. Multiple source direction of arrival estimations using relative sound pressure based MU-SIC. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):253–264, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3039569.

Hamilton:2017:FMD

[HB17] Brian Hamilton and Stefan Bilbao. FDTD methods for 3-D room acoustics simulation with highorder accuracy in space and time. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2112– 2124, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/8016655/.

He:2017:MUA

[HBB17] Qi He, Feng Bao, and Changchun Bao. Multiplicative update of autoregressive gains for codebookbased speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(3):457–468, ???? 2017. CODEN ???? ISSN 2329-9290.

Huang:2017:DFI

[HBC17] Gongping Huang, Jacob Benesty, and Jingdong Chen. On the design of frequency-invariant beampatterns with uniform circular microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):1140–1153, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7890381/.

Huang:2022:KPM

[HBCC22] Gongping Huang, Jacob Benesty, Israel Cohen, and Jingdong Chen. Kronecker product multichannel linear filtering for adaptive weighted prediction error-based speech dereverberation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1277–1289, 2022. CODEN ???? ISSN 2329-9290.

He:2016:UPB

[HBF^{+}16] Yanzhang He, P. Baumann, Hao Fang, B. Hutchinson, A. Jaech, M. Ostendorf, E. Fosler-Lussier, and J. Pierrehumbert. Using pronunciation-based morphological subword units to improve OOV handling in keyword search. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):79–92, January 2016. CODEN ???? ISSN 2329-9290.

Haubner:2024:EED

[HBK24] Thomas Haubner, Andreas Brendel, and Walter Kellermann. End-to-end deep learning-based adaptation control for linear acoustic echo cancellation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):227– 238, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3325923.

Huang:2014:FMS

[HBLC14] Gongping Huang, J. Benesty, Tao Long, and Jingdong Chen. A family of maximum SNR filters for noise reduction.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2034–2047, December 2014. CODEN ???? ISSN 2329-9290.

Hsu:2021:HSS

[HBT^{+}21] Wei-Ning Hsu, Benjamin Bolte, Yao-Hung Hubert Tsai, Kushal Lakhotia, Ruslan Salakhutdinov, and Abdelrahman Mohamed. HuBERT: Selfsupervised speech representation learning by masked prediction of hidden units. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3451– 3460, 2021. CODEN ???? ISSN 2329-9290.

Hu:2019:EFS

[HC19] J. Hu and N. Chen.

Enhanced feature summarizing for effective cover song identification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2113–2126, December 2019. CODEN ???? ISSN 2329-9290.

Huang:2018:IFI

[HCB18] G. Huang, J. Chen, and J. Benesty. Insights into frequency-invariant beamforming with concentric circular microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2305–2318, December 2018. CODEN ???? ISSN 2329-9290.

Huang:2020:DPD

[HCB20] Gongping Huang, Jingdong Chen, and Jacob Benesty. Design of planar differential microphone arrays with fractional orders. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):116–130, January 2020. CODEN ???? ISSN 2329-9290.

                                        URL https://dl.acm.
                                        org/doi/abs/10.1109/
                                        TASLP.2019.2949219.

He:2018:NRF

[HCBY18] H. He, J. Chen, J. Benesty, and T. Yang. Noise robust frequency-domain adaptive blind multichannel identification withℓ_{p}- norm constraint. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(9): 1608–1619, September 2018. CODEN ???? ISSN 2329-9290.

Hu:2023:CDG

[HCC^{+}23] Zhe Hu, Zhiwei Cao, Hou Pong Chan, Jiachen Liu, Xinyan Xiao, Jinsong Su, and Hua Wu. Controllable dialogue generation with disentangled multi-grained style specification and attribute consistency reward. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 188–199, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3221002.

Herzog:2023:AJA

[HCH23] Adrian Herzog, Srikanth Raj Chetupalli, and Emanuël A. P. Habets. AmbiSep: Joint ambisonicto-ambisonic speech separation and noise reduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3081– 3094, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3297954.

Hendriks:2015:ONE

[HCJT15] R. C. Hendriks, J. B.

Crespo, J. Jensen, and C. H. Taal. Optimal nearend speech intelligibility improvement incorporating additive noise and late reverberation under an approximation of the short-time SII. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(5): 851–862, May 2015. CO-DEN ???? ISSN 2329-9290.

Hsu:2023:ERG

[HCKW23] Jia-Hao Hsu, Jeremy Chang, Min-Hsueh Kuo, and Chung-Hsien Wu. Empathetic response generation based on plugand-play mechanism with empathy perturbation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):2032–2042, 2023. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2023.3277274.

Han:2024:SSL

[HCQ24] Bing Han, Zhengyang Chen, and Yanmin Qian. Self-supervised learning with cluster-aware-DINO for high-performance robust speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32 (??):529–541, 2024. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3331949.

Hu:2021:GCA

[HCY21a] De Hu, Zhe Chen, and Fuliang Yin. Geometry calibration for acoustic transceiver networks based on network Newton distributed optimization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1023–1032, 2021. CODEN ???? ISSN 2329-9290.

Hu:2021:PGC

[HCY21b] De Hu, Zhe Chen, and Fuliang Yin. Passive geometry calibration for microphone arrays based on distributed damped Newton optimization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):118–131, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3037532.

Hu:2024:WRC

[HCZC24] Yuchen Hu, Chen Chen, Qiushi Zhu, and Eng Siong Chng. Wav2code: Restore clean speech representations via codebook lookup for noiserobust ASR. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1145– 1156, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3332545.

Hadad:2016:BLB

[HDG16] Elior Hadad, Simon Doclo, and Sharon Gannot. The binaural LCMV beamformer and its performance analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): 543–558, March 2016. CODEN ???? ISSN 2329-

Hoang:2022:MSE

[HdHTJ22] Poul Hoang, Jan Mark de Haan, Zheng-Hua Tan, and Jesper Jensen. Multichannel speech enhancement with own voicebased interfering speech suppression for hearing assistive devices. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 706–720, 2022. CODEN ???? ISSN 2329-9290.

He:2023:AMM

[HDLL23] Mao-Kui He, Jun Du, Qing-Feng Liu, and Chin-Hui Lee. ANSD-MA-MSE: Adaptive neural speaker diarization using memory-aware multispeaker embedding. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1561–1573, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3265199.

Huang:2023:UDK

[HDW^{+}23] Tianyu Huang, Weisheng Dong, Fangfang Wu, Xin Li, and Guangming Shi. Uncertaintydriven knowledge distillation for language model compression. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2850– 2858, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3289303.

Helie:2014:SFO

[Hel14]  T. Helie.  Simulation

of fractional-order lowpass filters. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(11):1636– 1647, November 2014. CODEN ???? ISSN 2329-9290.

Hogg:2021:OSS

[HEMN21] Aidan O. T. Hogg, Christine Evers, Alastair H. Moore, and Patrick A. Naylor. Overlapping speaker segmentation using multiple hypothesis tracking of fundamental frequency. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1479– 1490, 2021. CODEN ???? ISSN 2329-9290.

Hoffmann:2015:TSA

[HF15] F.-M. Hoffmann and F. M. Fazi. Theoretical study of acoustic circular arrays with tangential pressure gradient sensors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1762–1774, November 2015. CODEN ???? ISSN 2329-9290.

Hamdan:2021:WOV

[HF21] Eric C. Hamdan and Filippo Maria Fazi. Weighted orthogonal vector rejection method for loudspeakerbased binaural audio reproduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1844– 1852, 2021. CODEN ???? ISSN 2329-9290.

Hollebon:2023:HOS

[HF23] Jacob Hollebon and Filippo Maria Fazi. Higherorder stereophony. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 2872–2885, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3297953.

He:2023:WNA

[HFL^{+}23] Ling He, Jia Fu, Yuanyuan Li, Xi Xiong, and Jing Zhang. WNSA-Net: an axial-attention-based network for schizophrenia detection using wideband and narrowband spectrograms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):721– 733, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3209941.

Horiguchi:2022:EDB

[HFW^{+}22] Shota Horiguchi, Yusuke Fujita, Shinji Watanabe, Yawen Xue, and Paola García. Encoderdecoder based attractors for end-to-end neural diarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1493– 1507, 2022. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3162080.

Hueber:2015:SAA

[HGAPB15] T. Hueber, L. Girin, X. Alameda-Pineda, and G. Bailly. Speakeradaptive acoustic-articulatory inversion using cascaded Gaussian mixture regression. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2246–2259, December 2015. CODEN ???? ISSN 2329-9290.

He:2015:PAE

[HGT15a] Jianjun He, Woon-Seng Gan, and Ee-Leng Tan. Primary-ambient extraction using ambient spectrum estimation for immersive spatial audio reproduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(9):1431– 1444, September 2015. CODEN ???? ISSN 2329-9290. See comments [LLS24].

He:2015:TSB

[HGT15b] Jianjun He, Woon-Seng Gan, and Ee-Leng Tan. Time-shifting based primaryambient extraction for spatial audio reproduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1576–1588, October 2015. CODEN ???? ISSN 2329-9290.

Hua:2015:TSE

[HGT15c] Guang Hua, J. Goh, and V. L. L. Thing. Timespread echo-based audio watermarking with optimized imperceptibility and robustness. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): 227–239, February 2015. CODEN ???? ISSN 2329-9290.

He:2023:DMA

[HGYN23] Qi He, Mingjie Gao, Ka Fai Cedric Yiu, and Sven Nordholm. Distributed microphone array localization problem via SDP-SOCP method. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):3579–3588, 2023. ISSN 2329-9290. URL

                                        https://dl.acm.org/doi/
                                        10.1109/TASLP.2023.3313437.

Hasan:2014:MLA

[HH14] T. Hasan and J. H. L.

Hansen. Maximum likelihood acoustic factor analysis models for robust speaker verification in noise. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):381–391, February 2014. CODEN ???? ISSN 2329-9290.

Hung:2016:RSR

[HHC16] Jeih-Weih Hung, Hsin-Ju Hsieh, and Berlin Chen. Robust speech recognition via enhancing the complex-valued acoustic spectrum in modulation domain. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(2): 236–251, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Huemmer:2018:EPN

[HHMK18] C. Huemmer, C. Hofmann, R. Maas, and W. Kellermann. Estimating parameters of nonlinear systems using the elitist particle filter based on evolutionary strategies. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):595–608, March 2018. CODEN ???? ISSN 2329-

Hono:2021:SDN

[HHO^{+}21] Yukiya Hono, Kei Hashimoto, Keiichiro Oura, Yoshihiko Nankaku, and Keiichi Tokuda. Sinsy: a deep neural networkbased singing voice synthesis system. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2803–2815, 2021. CODEN ???? ISSN 2329-9290.

Huang:2021:PTS

[HHW^{+}21] Wen-Chin Huang, Tomoki Hayashi, Yi-Chiao Wu, Hirokazu Kameoka, and Tomoki Toda. Pretraining techniques for sequenceto-sequence voice conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):745–755, 2021. CODEN ???? ISSN 2329-9290.

He:2021:ADA

[HHZ^{+}21] Fei He, Ling He, Jing Zhang, Yuanyuan Li, and Xi Xiong. Automatic detection of affective flattening in schizophrenia: Acoustic correlates to sound waves and auditory perception. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 3321–3334, 2021. CODEN ???? ISSN 2329-9290.

Higuchi:2017:OMB

[HIA^{+}17] Takuya Higuchi, Nobutaka Ito, Shoko Araki, Takuya Yoshioka, Marc Delcroix, and Tomohiro Nakatani. Online MVDR beamformer based on complex Gaussian mixture model with spatial prior for noise robust ASR. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(4):780–793, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7845594/.

Hansen:2019:EFF

[HJC19] M. W. Hansen, J. R.

Jensen, and M. G. Christensen. Estimation of fundamental frequencies in stereophonic music mixtures. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(2):296–310, February 2019. CODEN ???? ISSN 2329-9290.

Hauret:2023:CEE

[HJZB23] Julien Hauret, Thomas Joubaud, Véronique Zimpfer, and Éric Bavu. Configurable EBEN: Extreme bandwidth extension network to enhance body-conducted speech capture. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3499– 3512, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3313433.

Hamel:2024:FIA

[HK24] Emma Hamel and Nickvash Kani. Factors that influence automatic recognition of African-American vernacular English in machine-learning models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):509– 516, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3331139.

Huang:2015:JOM

[HKHJS15] Po-Sen Huang, Minje Kim, M. Hasegawa-Johnson, and P. Smaragdis. Joint optimization of masks and deep recurrent neural networks for monaural source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2136– 2147, December 2015. CODEN ???? ISSN 2329-9290.

Holighaus:2023:GBD

[HKHP23] Nicki Holighaus, Günther Koliander, Clara Hollomey, and Friedrich Pillichshammer. Gridbased decimation for wavelet transforms with stably invertible implementation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):789– 801, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3235197.

Hosoda:2023:CDP

[HKI23] Yuya Hosoda, Arata Kawamura, and Youji Iiguni. Complex-domain pitch estimation algorithm for narrowband speech signals. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 2067–2078, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3278488.

Hou:2024:CSE

[HKM^{+}24] Yuanbo Hou, Bo Kang, Andrew Mitchell, Wenwu Wang, Jian Kang, and Dick Botteldooren. Cooperative scene-event modelling for acoustic scene classification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 68–82, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3323135.

Hong:2021:SRR

[HKPR21] Joanna Hong, Minsu Kim, Se Jin Park, and Yong Man Ro. Speech reconstruction with reminiscent sound via visual voice memory. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 3654–3667, 2021. CODEN ???? ISSN 2329-9290.

Hawes:2014:SAD

[HL14] M. B. Hawes and Wei Liu. Sparse array design for wideband beamforming with reduced complexity in tapped delaylines. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):1236–1247, August 2014. CODEN ???? ISSN 2329-9290.

Hu:2015:SSV

[HL15] Ying Hu and Guizhong Liu. Separation of singing voice using nonnegative matrix partial co-factorization for singer identification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(4): 643–653, April 2015. CO-DEN ???? ISSN 2329-9290.

Hu:2018:ESF

[HL18] Y. Hu and Z. Ling. Extracting spectral features using deep autoencoders with binary distributed hidden units for statistical parametric speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(4):713–724, April 2018. CODEN ???? ISSN 2329-9290.

He:2024:LSE

[HLB^{+}24] Huang He, Hua Lu, Siqi Bao, Fan Wang, Hua Wu, Zheng-Yu Niu, and Haifeng Wang. Learning to select external knowledge with multi-scale negative sampling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 714–720, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3301222.

Heidel:2015:FCF

[HLL15] A. Heidel, Hsiang-Hung Lu, and Lin-Shan Lee. Finding complex features for guest language fragment recovery in resourcelimited code-mixed speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2148– 2161, December 2015. CODEN ???? ISSN 2329-9290.

Huang:2023:SSA

[HLZ^{+}23] Xinqi Huang, Yingsong Li, Yuriy Zakharov, Yongchun Miao, and Zhixiang Huang. Squared sine adaptive algorithm and its performance analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 3617–3628, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3313408.

Hui:2014:LSA

[HM14] Pui-Yu Hui and H. Meng. Latent semantic analysis for multimodal user input with speech and gestures. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):417–429, February 2014. CODEN ???? ISSN 2329-9290.

Hadad:2015:TAB

[HMDG15] E. Hadad, D. Marquardt, S. Doclo, and S. Gannot. Theoretical analysis of binaural transfer function MVDR beamformers with interference cue preservation constraints. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2449–2464, December 2015. CODEN ???? ISSN 2329-9290.

Hafezi:2017:AIV

[HMN17] Sina Hafezi, Alastair H. Moore, and Patrick A. Naylor. Augmented intensity vectors for direction of arrival estimation in the spherical harmonic domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(10):1956–1968, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8002621/.

He:2021:NNA

[HMO21] Weipeng He, Petr Motlicek, and Jean-Marc Odobez. Neural network adaptation and data augmentation for multi-speaker direction-of-arrival estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1303– 1317, 2021. CODEN ???? ISSN 2329-9290.

Hoffmann:2020:DEP

[HNF20] Falk-Martin Hoffmann, Philip Arthur Nelson, and Filippo Maria Fazi. DOA estimation performance with circular arrays in sound fields with finite rate of innovation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):171–184, January 2020. CODEN ???? ISSN 2329-9290.

URL https://dl.acm.
org/doi/abs/10.1109/
TASLP.2019.2950774.

Hasumi:2023:PIP

[HNT^{+}23] Takuya Hasumi, Tomohiko Nakamura, Norihiro Takamune, Hiroshi Saruwatari, Daichi Kitamura, Yu Takahashi, and Kazunobu Kondo. PoP-IDLMA: Product-of-prior independent deeply learned matrix analysis for multichannel music source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2680–2694, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3293044.

Hirao:2015:SDT

[HNY^{+}15] T. Hirao, M. Nishino, Y. Yoshida, J. Suzuki, N. Yasuda, and M. Nagata. Summarizing a document by trimming the discourse tree. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(11): 2081–2092, November 2015. CODEN ???? ISSN 2329-9290.

Hutchinson:2015:SPL

[HOF15] B. Hutchinson, M. Ostendorf, and M. Fazel. A sparse plus low-rank exponential language model for limited resource scenarios. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (3):494–504, March 2015. CODEN ???? ISSN 2329-9290.

Habibi:2015:KEC

[HPB15] M. Habibi and A. Popescu-Belis. Keyword extraction and clustering for document recommendation in conversations. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(4): 746–759, April 2015. CO-DEN ???? ISSN 2329-9290.

Hold:2024:CHO

[HPPM24] Christoph Hold, Ville Pulkki, Archontis Politis, and Leo McCormack. Compression of higher-order ambisonic signals using directional audio coding. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 651–665, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3328284.

Huang:2023:CBA

[HQTX23] Maochun Huang, Chunmei Qing, Junpeng Tan, and Xiangmin Xu. Contextbased adaptive multimodal fusion network for continuous framelevel sentiment prediction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3468–3477, 2023. ISSN 2329-

9290. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2023.3321971.

Haque:2021:GGA

[HRL^{+}21] Kazi Nazmul Haque, Rajib Rana, Jiajun Liu, John H. L. Hansen, Nicholas Cummins, Carlos Busso, and Björn W. Schuller. Guided generative adversarial neural network for representation learning and audio generation using fewer labelled audio data. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2575–2590, 2021. CODEN ???? ISSN 2329-9290.

Heo:2015:CBS

[HS15] Inseok Heo and W. A. Sethares. Classification based on speech rhythm via a temporal alignment of spoken sentences. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2209–2216, December 2015. CODEN ???? ISSN 2329-9290.

Su:2015:RDG

[hShWsL15] Pei hao Su, Chuan hsun Wu, and Lin shan Lee. A recursive dialogue game for personalized computer-aided pronunciation training. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(1):

127–141, January 2015. CODEN ???? ISSN 2329-9290.

Hatano:2017:CND

[HSK17] Yuta Hatano, Chuang Shi, and Yoshinobu Kajikawa. Compensation for nonlinear distortion of the frequency modulationbased parametric array loudspeaker. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(8):1709– 1717, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7930484/.

Hu:2023:DSS

[HSLB23] De Hu, Qintuya Si, Rui Liu, and Feilong Bao. Distributed sensor selection for speech enhancement with acoustic sensor networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 985–999, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3244525.

Hashimoto:2019:MMC

[HSM19] T. Hashimoto, D. Saito, and N. Minematsu. Manyto-many and completely parallel-data-free voice conversion based on eigenspace DNN. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(2):332–341, February 2019. CODEN ???? ISSN 2329-9290.

Heck:2018:DPM

[HSN18] M. Heck, S. Sakti, and S. Nakamura. Dirichlet process mixture of mixtures model for unsupervised subword modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):2027–2042, November 2018. CODEN ???? ISSN 2329-9290.

Hadian:2018:FSS

[HSPK18] H. Hadian, H. Sameti, D. Povey, and S. Khudanpur. Flat-start singlestage discriminatively trained HMM-based models for ASR. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):1949–1961, November 2018. CODEN ???? ISSN 2329-9290.

Hsu:2021:SER

[HSWC21] Jia-Hao Hsu, Ming-Hsiang Su, Chung-Hsien Wu, and Yi-Hsuan Chen. Speech emotion recognition considering nonverbal vocalization in affective conversations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1675– 1686, 2021. CODEN ???? ISSN 2329-9290.

Huang:2018:EHG

[HSZ^{+}18] J. Huang, Y. Sun, W. Zhang, H. Wang, and T. Liu. Entity highlight generation as statistical and neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):1860–1872, October 2018. CODEN ???? ISSN 2329-9290.

He:2014:LEB

[HTG14] Jianjun He, Ee-Leng Tan, and Woon-Seng Gan. Linear estimation based primary-ambient extraction for stereo audio signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):505–517, February 2014. CODEN ???? ISSN 2329-9290.

Han:2014:NNB

[HW14] Kun Han and Deliang Wang. Neural network based pitch tracking in very noisy speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2158–2168, December 2014. CODEN ???? ISSN 2329-9290.

Hu:2023:RDP

[HW23] Xinyu Hu and Xiaojun Wan. RST discourse parsing as text-to-text generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3278– 3289, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3306710.

Huang:2017:PSS

[HWC^{+}17] Yi-Chin Huang, Chung-Hsien Wu, Yan-You Chen, Ming-Ge Shie, and Jhing-Fa Wang. Personalized spontaneous speech synthesis using a small-sized unsegmented semispontaneous speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(5): 1048–1060, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7874173/.

Horiguchi:2023:OND

[HWG^{+}23] Shota Horiguchi, Shinji Watanabe, Paola García, Yuki Takashima, and Yohei Kawaguchi. Online neural diarization of unlimited numbers of speakers using global and local attractors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 706–720, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3233237.

Hon:2015:AFM

[HWRC15] Tsz-Kin Hon, Lin Wang, J. D. Reiss, and A. Cavallaro. Audio fingerprinting for multi-device selflocalization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1623– 1636, October 2015. CO-DEN ???? ISSN 2329-9290.

Hayashi:2017:DCL

[HWT^{+}17] Tomoki Hayashi, Shinji Watanabe, Tomoki Toda, Takaaki Hori, Jonathan Le Roux, and Kazuya Takeda. Duration-controlled LSTM for polyphonic sound event detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (11):2059–2070, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8010445/.

Han:2015:LSM

[HWW^{+}15] Kun Han, Yuxuan Wang, Deliang Wang, W. S. Woods, I. Merks, and Tao Zhang. Learning spectral mapping for speech dereverberation and denoising. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(6):982–992, June 2015. CODEN ???? ISSN 2329-

Huang:2016:IMP

[HWW16] Yi-Chin Huang, Chung-Hsien Wu, and Si-Ting Weng. Improving Mandarin prosody generation using alternative smoothing techniques. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(11): 1897–1907, ???? 2016. CODEN ???? ISSN 2329-9290.

Hong:2023:DRP

[HWW23a] Qian-Bei Hong, Chung-Hsien Wu, and Hsin-Min Wang. Decomposition and reorganization of phonetic information for speaker embedding learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1745– 1757, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3267833.

Hong:2023:GAI

[HWW23b] Qian-Bei Hong, Chung-Hsien Wu, and Hsin-Min Wang. Generalization ability improvement of speaker representation and antiinterference for speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):486– 499, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3221042.

Hu:2022:NAS

[HXC^{+}22] Shoukang Hu, Xurong Xie, Mingyu Cui, Jiajun Deng, Shansong Liu, Jianwei Yu, Mengzhe Geng, Xunying Liu, and Helen Meng. Neural architecture search for LF-MMI trained time delay neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1093–1107, 2022. CODEN ???? ISSN 2329-9290.

Hu:2021:BLL

[HXL^{+}21] Shoukang Hu, Xurong Xie, Shansong Liu, Jianwei Yu, Zi Ye, Mengzhe Geng, Xunying Liu, and Helen Meng. Bayesian learning of LF-MMI trained time delay neural networks for speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1514–1529, 2021. CODEN ???? ISSN 2329-9290.

Huang:2015:MFS

[HXWS15] Hao Huang, Haihua Xu, Xianhui Wang, and W. Silamu. Maximum F1score discriminative training criterion for automatic mispronunciation detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(4):787–797, April 2015. CODEN ???? ISSN 2329-9290.

Hirayama:2015:ASR

[HYI^{+}15] N. Hirayama, K. Yoshino, K. Itoyama, S. Mori, and H. G. Okuno. Automatic speech recognition for mixed dialect utterances by mixing dialect language models. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): 373–382, February 2015. CODEN ???? ISSN 2329-9290.

Hu:2023:DSR

[HZBW23] De De Hu, Huaiwen Zhang, Feilong Bao, and Rui Wang. Distributed sampling rate offset estimation over acoustic sensor networks based on asynchronous network Newton optimization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):301– 312, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3224256.

Hou:2022:MVS

[HZC^{+}22] Mixiao Hou, Zheng Zhang, Qi Cao, David Zhang, and Guangming Lu. Multiview speech emotion recognition via collective relation construction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):218–229, 2022. CO-DEN ???? ISSN 2329-9290.

Huang:2017:TSD

[HZF17] Qinghua Huang, Lin Zhang, and Yong Fang. Two-stage decoupled DOA estimation based on real spherical harmonics for spherical arrays. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(11): 2045–2058, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8003459/.

Huang:2018:TSS

[HZF18] Q. Huang, L. Zhang, and Y. Fang. Twostep spherical harmonics ESPRIT-type algorithms and performance analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1684–1697, September 2018. CODEN ???? ISSN 2329-9290.

Hu:2023:EPO

[HZL23] Guimin Hu, Yi Zhao, and Guangming Lu. Emotion prediction oriented method with multiple supervisions for emotioncause pair extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):1141–1152, 2023. ISSN 2329-9290. URL

                                        https://dl.acm.org/doi/
                                        10.1109/TASLP.2023.3250833.

Hou:2022:EAC

[HZW^{+}22] Wenxin Hou, Han Zhu, Yidong Wang, Jindong Wang, Tao Qin, Renjun Xu, and Takahiro Shinozaki. Exploiting adapters for cross-lingual low-resource speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):317– 329, 2022. CODEN ???? ISSN 2329-9290.

Itzhak:2021:DDK

[IBC21] Gal Itzhak, Jacob Benesty, and Israel Cohen. On the design of differential Kronecker product beamformers. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1397–1410, 2021. CODEN ???? ISSN 2329-9290.

Itakura:2018:BMA

[IBN^{+}18] K. Itakura, Y. Bando, E. Nakamura, K. Itoyama, K. Yoshii, and T. Kawahara. Bayesian multichannel audio source separation based on integrated source and spatial models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(4):831–846, April 2018. CODEN ???? ISSN 2329-9290.

Itturriet:2019:PRP

[IC19] F. P. Itturriet and M. H.

Costa. Perceptually relevant preservation of interaural time differences in binaural hearing aids. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (4):753–764, April 2019. CODEN ???? ISSN 2329-9290.

Ivry:2024:UCA

[ICB24] Amir Ivry, Israel Cohen, and Baruch Berdugo. A user-centric approach for deep residual-echo suppression in doubletalk. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1901–1914, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3374064.

Ito:2021:JDB

[IISN21] Nobutaka Ito, Rintaro Ikeshita, Hiroshi Sawada, and Tomohiro Nakatani. A joint diagonalization based efficient approach to underdetermined blind audio source separation using the multichannel Wiener filter. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1950–1965, 2021. CODEN ???? ISSN 2329-9290.

Ikemiya:2016:SVS

[IIY16] Yukara Ikemiya, Katsutoshi Itoyama, and Kazuyoshi Yoshii. Singing voice separation and vocal F0 estimation based on mutual combination of robust principal component analysis and subharmonic summation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(11): 2084–2095, ???? 2016. CODEN ???? ISSN 2329-9290.

Inaguma:2023:AKD

[IK23] Hirofumi Inaguma and Tatsuya Kawahara. Alignment knowledge distillation for online streaming attention-based speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1371– 1385, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2021.3133217.

Imoto:2017:SCS

[IO17] Keisuke Imoto and Nobutaka Ono. Spatial cepstrum as a spatial feature using a distributed microphone array for acoustic scene analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(6): 1335–1343, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7933053/.

Imoto:2019:ATM

[IO19] K. Imoto and N. Ono.

Acoustic topic model for scene analysis with intermittently missing observations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(2):367–382, February 2019. CODEN ???? ISSN 2329-9290.

Ibarrola:2019:SDS

[ISP19] F. J. Ibarrola, R. D.

Spies, and L. E. D. Persia. Switching divergences for spectral learning in blind speech dereverberation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(5):881–891, May 2019. CODEN ???? ISSN 2329-9290.

Islam:2015:SEB

[ISZA15] M. T. Islam, C. Shahnaz, Wei-Ping Zhu, and M. O. Ahmad. Speech enhancement based on student modeling of teager energy operated perceptual wavelet packet coefficients and a custom thresholding function. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):1800– 1811, November 2015. CODEN ???? ISSN 2329-9290.

Jensen:2016:NRO

[JBC16] Jesper Rindom Jensen, Jacob Benesty, and Mads Græsbøll Christensen. Noise reduction with optimal variable span linear filters. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (4):631–644, April 2016. CODEN ???? ISSN 2329-9290.

Jin:2023:DBG

[JBCH23] Jilu Jin, Jacob Benesty, Jingdong Chen, and Gongping Huang. Differential beamforming from a geometric perspective. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3042– 3054, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3301245.

Juvela:2019:GRW

[JBTA19] L. Juvela, B. Bollepalli, V. Tsiaras, and P. Alku. GlotNet — a raw waveform model for the glottal excitation in statistical parametric speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(6):1019– 1030, June 2019. CODEN ???? ISSN 2329-9290.

Jiao:2015:CWC

[JBTL15] Yishan Jiao, V. Berisha, Ming Tu, and J. Liss. Convex weighting criteria for speaking rate estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(9):1421–1430, September 2015. CODEN ???? ISSN 2329-9290.

Jiang:2024:CDA

[JBWZ24] Xiaotong Jiang, Ruirui Bai, Zhongqing Wang, and Guodong Zhou. Crossdomain aspect-based sentiment classification with tripartite graph modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1623–1635, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3365975.

Jo:2017:SHS

[JC17] Byeongho Jo and Jung-Woo Choi. Spherical harmonic smoothing for localizing coherent sound sources. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(10):1969– 1984, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

                                        ieeexplore.ieee.org/
                                        document/8008781/.

Jensen:2015:JST

[JCBJ15] J. R. Jensen, M. G. Christensen, J. Benesty, and S. H. Jensen. Joint spatiotemporal filtering methods for DOA and fundamental frequency estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):174–185, January 2015. CODEN ???? ISSN 2329-9290.

Jeong:2021:LCA

[JCYwH21] Myeongho Jeong, Seungtaek Choi, Jinyoung Yeo, and Seung won Hwang. Label and context augmentation for response selection at DSTC8. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2541–2550, 2021. CODEN ???? ISSN 2329-9290.

Janke:2017:ESD

[JD17] Matthias Janke and Lorenz Diener. EMG-to-speech: Direct generation of speech from facial electromyographic signals. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12): 2375–2385, ???? 2017.

CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8114359/.

Jin:2014:DOE

[JEP14] C. T. Jin, N. Epain, and A. Parthy. Design, optimization and evaluation of a dual-radius spherical microphone array. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): 193–204, January 2014. CODEN ???? ISSN 2329-9290.

Jalmby:2023:LRR

[JEvW23] Martin Jälmby, Filip Elvander, and Toon van Waterschoot. Low-rank room impulse response estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):957– 969, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3240650.

Jameel:2017:NRF

[JFG^{+}17] Abu Shafin Mohammad Mahdee Jameel, Shaikh Anowarul Fattah, Rajib Goswami, Wei-Ping Zhu, and M. Omair Ahmad. Noise robust formant frequency estimation method based on spectral model of repeated autocorrelation of speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (6):1357–1370, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7736088/.

Jati:2019:NPC

[JG19] A. Jati and P. Georgiou. Neural predictive coding using convolutional neural networks toward unsupervised learning of speaker characteristics. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(10):1577–1589, October 2019. CODEN ???? ISSN 2329-9290.

Jin:2024:PAD

[JGD^{+}24] Zengrui Jin, Mengzhe Geng, Jiajun Deng, Tianzi Wang, Shujie Hu, Guinan Li, and Xunying Liu. Personalized adversarial data augmentation for dysarthric and elderly speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32 (??):413–429, 2024. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3323888.

Jiao:2023:EMD

[JGHN23] Fangkai Jiao, Yangyang Guo, Minlie Huang, and Liqiang Nie. Enhanced multi-domain dialogue state tracker with secondorder slot interactions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):265–276, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2022.3221044.

Jin:2022:TTD

[JGK^{+}22] Di Jin, Shuyang Gao, Seokhwan Kim, Yang Liu, and Dilek Hakkani-Tür. Towards textual out-of-domain detection without in-domain labels. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1386– 1395, 2022. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3162081.

Jacome:2023:SEL

[JGM23] Karen Gissell Rosero Jacome, Felipe Leonel Grijalva, and Bruno Sanches Masiero. Sound events localization and detection using bio-inspired gammatone filters and temporal convolutional neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2314– 2324, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3284525.

Jorge:2022:LSS

[JGSC^{+}22] Javier Jorge, Adrià Giménez, Joan Albert Silvestre-Cerdà, Jorge Civera, Albert Sanchis, and Alfons Juan. Live streaming speech recognition using deep bidirectional LSTM acoustic models and interpolated language models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):148–161, 2022. CO-DEN ???? ISSN 2329-9290.

Jang:2018:CLN

[JHLK18] Y. Jang, J. Ham, B. Lee, and K. Kim. Crosslanguage neural dialog state tracker for large ontologies using hierarchical attention. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):2072– 2082, November 2018. CODEN ???? ISSN 2329-9290.

Jin:2021:SSL

[JHW^{+}21] Jilu Jin, Gongping Huang, Xuehan Wang, Jingdong Chen, Jacob Benesty, and Israel Cohen. Steering study of linear differential microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):158–170, January

                                        2021. URL https://dl.

acm.org/doi/10.1109/
TASLP.2020.3038566.

Ji:2016:PRS

[JJB16] An Ji, Michael T. Johnson, and Jeffrey J. Berry. Parallel reference speaker weighting for kinematicindependent acoustic-to-articulatory inversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (10):1865–1875, ???? 2016. CODEN ???? ISSN 2329-9290.

Jin:2015:TDM

[JK15] Wenyu Jin and W. B.

Kleijn. Theory and design of multizone soundfield reproduction using sparse methods. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2343– 2355, December 2015. CODEN ???? ISSN 2329-9290.

Jancovic:2019:BSR

[JK19] P. Jancovic and M. Köküer.

Bird species recognition using unsupervised modeling of individual vocalization elements. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(5): 932–947, May 2019. CO-DEN ???? ISSN 2329-

Jeong:2024:TLL

[JKC^{+}24] Myeonghun Jeong, Minchan Kim, Byoung Jin Choi, Jaesam Yoon, Won Jang, and Nam Soo Kim. Transfer learning for low-resource, multilingual, and zero-shot multi-speaker text-to-speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1519– 1530, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3364085.

Joy:2018:FSN

[JKU18] N. M. Joy, S. R. Kothinti, and S. Umesh. FM-LLR speaker normalization with i-Vector: In pseudo-FMLLR and distillation framework. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (4):797–805, April 2018. CODEN ???? ISSN 2329-9290.

Jeon:2021:TNT

[JLKK21] Kwang Myung Jeon, Geon Woo Lee, Nam Kyun Kim, and Hong Kook Kim. TAU-Net: Temporal activation U-net shared with nonnegative matrix factorization for speech enhancement in unseen noise environments. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3400– 3414, 2021. CODEN ???? ISSN 2329-9290.

Jia:2021:HRI

[JLP^{+}21] Bijue Jia, Jiancheng Lv, Xi Peng, Yao Chen, and Shenglan Yang. Hierarchical regulated iterative network for joint task of music detection and music relative loudness estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1–13, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3030484.

Jiang:2024:ERE

[JLZD24] Shu Jiang, Zuchao Li, Hai Zhao, and Weiping Ding. Entity-relation extraction as full shallow semantic dependency parsing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1088– 1099, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3350905.

Janod:2017:DBF

[JMD^{+}17] Killian Janod, Mohamed Morchid, Richard Dufour, Georges Linarès, and Renato De Mori. Denoised bottleneck features from deep autoencoders for telephone conversation analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(9): 1809–1820, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7954983/.

Jati:2021:TDW

[JNP^{+}21] Arindam Jati, Amrutha Nadarajan, Raghuveer Peri, Karel Mundnich, Tiantian Feng, Benjamin Girault, and Shrikanth Narayanan. Temporal dynamics of workplace acoustic scenes: Egocentric analysis and prediction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):756–769, 2021. CODEN ???? ISSN 2329-9290.

Jain:2014:EBM

[JP14] P. Jain and R. B. Pachori. Event-based method for instantaneous fundamental frequency estimation from voiced speech based on eigenvalue decomposition of the Hankel matrix. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (10):1467–1482, October 2014. CODEN ???? ISSN 2329-9290.

Jiang:2023:LDP

[JPX^{+}23] Xue Jiang, Xiulian Peng, Huaying Xue, Yuan Zhang, and Yan Lu. Latentdomain predictive neural speech coding. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 2111–2123, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3277693.

Jiang:2024:OGT

[JQL24] Congcong Jiang, Tieyun Qian, and Bing Liu. One general teacher for multi-data multi-task: a new knowledge distillation framework for discourse relation analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 239–249, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3328289.

Jokinen:2017:IET

[JRA17] Emma Jokinen, Ulpu Remes, and Paavo Alku. Intelligibility enhancement of telephone speech using Gaussian process regression for normalto-lombard spectral tilt conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(10):1985– 1996, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

                                        ieeexplore.ieee.org/
                                        document/8010472/.

Jiang:2022:ASE

[JSG^{+}22] Hui Jiang, Linfeng Song, Yubin Ge, Fandong Meng, Junfeng Yao, and Jinsong Su. An AST structure enhanced decoder for code generation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 468–476, 2022. CODEN ???? ISSN 2329-9290.

Jin:2018:LST

[JSMD18] Ma Jin, Yan Song, Ian McLoughlin, and Li-Rong Dai. LID-senones and their statistics for language identification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (1):171–183, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8080255/.

Jeffet:2016:TPE

[JSR16] Michael Jeffet, Noam R. Shabtai, and Boaz Rafaely. Theory and perceptual evaluation of the binaural reproduction and beamforming tradeoff in the generalized spherical array beamformer. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 708–718, April 2016. CO-DEN ???? ISSN 2329-9290.

Jao:2016:MMS

[JSYW16] Ping-Keng Jao, Li Su, Yi-Hsuan Yang, and Brendt Wohlberg. Monaural music source separation using convolutional sparse coding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):2158–2170, ???? 2016. CODEN ???? ISSN 2329-9290.

Jensen:2014:SIP

[JT14] J. Jensen and C. H. Taal.

Speech intelligibility prediction based on mutual information. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): 430–440, February 2014. CODEN ???? ISSN 2329-9290.

Jensen:2015:MMS

[JT15] J. Jensen and Zheng-Hua Tan. Minimum mean-square error estimation of mel-frequency cepstral features — a theoretically consistent approach. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):186– 197, January 2015. CO-DEN ???? ISSN 2329-9290.

Jensen:2016:API

[JT16] Jesper Jensen and Cees H. Taal. An algorithm for predicting the intelligibility of speech masked by modulated noise maskers. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(11): 2009–2022, ???? 2016. CODEN ???? ISSN 2329-9290.

Jarrett:2014:NRS

[JTHN14] D. P. Jarrett, M. Taseska, E. A. P. Habets, and P. A. Naylor. Noise reduction in the spherical harmonic domain using a tradeoff beamformer and narrowband DOA estimates. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (5):967–978, May 2014. CODEN ???? ISSN 2329-9290.

Jaureguiberry:2016:FMS

[JVR16]  Xabier  Jaureguiberry,

Emmanuel Vincent, and Gaël Richard. Fusion methods for speech enhancement and audio source separation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(7): 1266–1279, July 2016.

CODEN ???? ISSN 2329-9290.

Jukic:2015:MCL

[JvWGD15] A. Jukic, T. van Waterschoot, T. Gerkmann, and S. Doclo. Multichannel linear predictionbased speech dereverberation with sparse priors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (9):1509–1520, September 2015. CODEN ???? ISSN 2329-9290.

Jia:2021:MSD

[JWBR21] Maoshen Jia, Yuxuan Wu, Changchun Bao, and Christian Ritz. Multisource DOA estimation in reverberant environments by jointing detection and modeling of time-frequency points. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):379–392, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3042705.

Jiao:2022:EIE

[JWH^{+}22] Wenxiang Jiao, Xing Wang, Shilin He, Zhaopeng Tu, Irwin King, and Michael R. Lyu. Exploiting inactive examples for natural language generation with data rejuvenation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):931–943, 2022. CODEN ???? ISSN 2329-9290.

Jiang:2014:BCR

[JWLF14] Yi Jiang, Deliang Wang, Runsheng Liu, and Zhen-Ming Feng. Binaural classification for reverberant speech segregation using deep neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2112–2121, December 2014. CODEN ???? ISSN 2329-9290.

Jiang:2023:SEI

[JY23] Wenbin Jiang and Kai Yu. Speech enhancement with integration of neural homomorphic synthesis and spectral masking. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1758– 1770, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3271151.

Jia:2015:EMA

[JYB^{+}15] Maoshen Jia, Ziyu Yang, Changchun Bao, Xiguang Zheng, and C. Ritz. Encoding multiple audio objects using intra-object sparsity. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(6):1082– 1095, June 2015. CODEN ???? ISSN 2329-9290.

Jiang:2024:ESD

[JYC^{+}24] Xiaotong Jiang, Peiwen You, Chen Chen, Zhongqing Wang, and Guodong Zhou. Exploring scope detection for aspect-based sentiment analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 83–94, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3323136.

Jahromi:2019:ILH

[JZJØ19] M. Z. Jahromi, A. Zahedi, J. Jensen, and J. Østergaard. Information loss in the human auditory system. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(3):472–481, March 2019. CODEN ???? ISSN 2329-9290.

Kates:2016:HAA

[KA16] J. M. Kates and K. H.

Arehart. The Hearing-Aid Audio Quality Index (HAAQI). IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(2): 354–365, ???? 2016. CO-DEN ???? ISSN 2329-

Kanan:2024:IHC

[KAA^{+}24] Tarek Kanan, Amani AbedAlghafer, Shadi AlZu’bi, Bilal Hawashin, Ala Mughaid, Ghassan Kanaan, and M. M. Kamruzzaman. An intelligent health care system for detecting drug abuse in social media platforms based on low resource language. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 691–703, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3294699.

Karimian-Azari:2016:CEN

[KAJC16] Sam Karimian-Azari, Jesper Rindom Jensen, and Mads Græsbøll Christensen. Computationally efficient and noise robust DOA and pitch estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(9):1613–1625, ???? 2016. CODEN ???? ISSN 2329-9290.

Kamper:2023:WSD

[Kam23] Herman Kamper. Word segmentation on discovered phone units with dynamic programming and self-supervised scoring. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):684–694, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2022.3229264.

Kelly:2014:DAR

[KB14] I. J. Kelly and F. M.

Boland. Detecting arrivals in room impulse responses with dynamic time warping. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(7):1139–1147, July 2014. CODEN ???? ISSN 2329-9290.

Krawczyk-Becker:2016:FFI

[KBG16a] Martin Krawczyk-Becker and Timo Gerkmann. Fundamental frequency informed speech enhancement in a flexible statistical framework. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(5): 940–951, May 2016. CO-DEN ???? ISSN 2329-9290.

Krawczyk-Becker:2016:MBE

[KBG16b] Martin Krawczyk-Becker and Timo Gerkmann. On MMSE-based estimation of amplitude and complex speech spectral coefficients under phaseuncertainty. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2251– 2262, ???? 2016. CODEN ???? ISSN 2329-9290.

Krawczyk-Becker:2018:SEU

[KBG18] M. Krawczyk-Becker and T. Gerkmann. On speech enhancement under PSD uncertainty. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(6):1144– 1153, June 2018. CODEN ???? ISSN 2329-9290.

Krug:2023:AVL

[KBG^{+}23] Paul Konstantin Krug, Peter Birkholz, Branislav Gerazov, Daniel Rudolph van Niekerk, Anqi Xu, and Yi Xu. Artificial vocal learning guided by phoneme recognition and visual information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1734– 1744, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3264454.

Kounades-Bastian:2016:VEA

[KBGAP^{+}16] Dionyssos Kounades-Bastian, Laurent Girin, Xavier Alameda-Pineda, Sharon Gannot, and Radu Horaud. A variational EM algorithm for the separation of time-varying convolutive audio mixtures. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (8):1408–1423, ???? 2016.

CODEN ???? ISSN 2329-9290.

Koyama:2021:SAN

[KBI^{+}21] Shoichi Koyama, Jesper Brunnström, Hayato Ito, Natsuki Ueno, and Hiroshi Saruwatari. Spatial active noise control based on kernel interpolation of sound field. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 3052–3063, 2021. CODEN ???? ISSN 2329-9290.

Krishnan:2019:FAA

[KBT19] L. Krishnan, T. Betlehem, and P. D. Teal. Fast algorithms for acoustic impulse response shaping. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(2):392–403, February 2019. CODEN ???? ISSN 2329-9290.

Khalilian:2016:CLP

[KBV16] Hanieh Khalilian, Ivan V.

Bajić, and Rodney G. Vaughan. Comparison of loudspeaker placement methods for sound field reproduction. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(8): 1364–1379, ???? 2016. CODEN ???? ISSN 2329-

Khalilian:2017:SST

[KBV17] Hanieh Khalilian, Ivan V. Bajić, and Rodney G. Vaughan. A simulation study of a threedimensional sound field reproduction system for immersive communication. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(5):980–995, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7864381/.

Kawara:2021:PETa

[KCA21a] Yuki Kawara, Chenhui Chu, and Yuki Arase. Preordering encoding on transformer for translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):644–655, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3042001.

Kawara:2021:PETb

[KCA21b] Yuki Kawara, Chenhui Chu, and Yuki Arase. Preordering encoding on transformer for translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):644–655, 2021. CODEN ???? ISSN 2329-9290.

Koyama:2020:OSS

[KCD20] Shoichi Koyama, Gilles Chardon, and Laurent Daudet. Optimizing source and sensor placement for sound field control: an overview. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 696–714, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.2964958.

Kuo:2021:AAS

[KCL21] Chia-Chih Kuo, Kuan-Yu Chen, and Shang-Bao Luo. Audio-aware spoken multiple-choice question answering with pretrained language models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3170–3179, 2021. CODEN ???? ISSN 2329-9290.

Kim:2017:SIS

[KCMW17] Myungjong Kim, Beiming Cao, Ted Mau, and Jun Wang. Speakerindependent silent speech recognition from fleshpoint articulatory movements using an LSTM neural network. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12): 2323–2336, ???? 2017.

CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8114350/.

Kodrasi:2016:JDN

[KD16] Ina Kodrasi and Simon Doclo. Joint dereverberation and noise reduction based on acoustic multi-channel equalization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(4):680–693, April 2016. CODEN ???? ISSN 2329-9290.

Kodrasi:2017:SDP

[KD17] Ina Kodrasi and Simon Doclo. Signaldependent penalty functions for robust acoustic multi-channel equalization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(7):1512–1525, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7914695/.

Kodrasi:2018:AED

[KD18] I. Kodrasi and S. Doclo. Analysis of eigenvalue decomposition-based late reverberation power spectral density estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (6):1106–1118, June 2018.

CODEN ???? ISSN 2329-9290.

Kipnis:2021:GBC

[KD21] Dror Kipnis and Roee Diamant. Graph-based clustering of dolphin whistles. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2216–2227, 2021. CODEN ???? ISSN 2329-9290.

Kitic:2024:BIA

[KD24] Sr– dan Kitić and Jérôme Daniel. Blind identification of ambisonic reduced room impulse response. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):443– 458, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3332546.

Kuklasinski:2016:MLP

[KDJJ16] Adam Kuklasiński, Simon Doclo, Søren Holdt Jensen, and Jesper Jensen. Maximum likelihood PSD estimation for speech enhancement in reverberation and noise. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(9): 1599–1612, ???? 2016. CODEN ???? ISSN 2329-

Kuklasinski:2017:CML

[KDJJ17] Adam Kuklasiński, Simon Doclo, Søren Holdt Jensen, and Jesper Jensen. Correction to “Maximum Likelihood PSD Estimation for Speech Enhancement in Reverberation and Noise”. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(3): 687, ???? 2017. CODEN ???? ISSN 2329-9290. See [CG16].

Khanagha:2014:DGC

[KDY14] V. Khanagha, K. Daoudi, and H. M. Yahia. Detection of glottal closure instants based on the microcanonical multiscale formalism. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1941– 1950, December 2014. CODEN ???? ISSN 2329-9290.

Kirsch:2023:UFA

[KE23] Christoph Kirsch and Stephan D. Ewert. A universal filter approximation of edge diffraction for geometrical acoustics. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1636–1651, 2023. ISSN 2329-

9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2023.3264737.

Koutini:2021:RFR

[KEzW21] Khaled Koutini, Hamid Eghbal-zadeh, and Gerhard Widmer. Receptive field regularization techniques for audio classification and tagging with deep convolutional neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1987–2000, 2021. CODEN ???? ISSN 2329-9290.

Koyama:2014:WFR

[KFH^{+}14] S. Koyama, K. Furuya, Y. Hiwasaki, Y. Haneda, and Y. Suzuki. Wave field reconstruction filtering in cylindrical harmonic domain for withheight recording and reproduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(10):1546– 1557, October 2014. CO-DEN ???? ISSN 2329-9290.

Krawczyk:2014:SPR

[KG14] M. Krawczyk and T. Gerkmann. STFT phase reconstruction in voiced speech for an improved singlechannel speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1931–1940, December 2014. CODEN ???? ISSN 2329-9290.

Kroher:2016:ATF

[KG16] Nadine Kroher and Emilia Gómez. Automatic transcription of flamenco singing from polyphonic music recordings. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(5): 901–913, May 2016. CO-DEN ???? ISSN 2329-9290.

Korpusik:2017:SLU

[KG17] Mandy Korpusik and James Glass. Spoken language understanding for a nutrition dialogue system. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (7):1450–1461, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7902155/.

Korpusik:2019:DLD

[KG19] M. Korpusik and J. Glass. Deep learning for database mapping and asking clarification questions in dialogue systems. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(8): 1321–1334, August 2019. CODEN ???? ISSN 2329-9290.

Krause:2024:BSS

[KGBPM24] Daniel Aleksander Krause, Guillermo García-Barrios, Archontis Politis, and Annamaria Mesaros. Binaural sound source distance estimation and localization for a moving listener. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):996– 1011, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3346297.

Kumar:2019:DLV

[KGG19] A. Kumar, T. Guha, and P. K. Ghosh. Dirichlet latent variable model: a dynamic model based on Dirichlet prior for audio processing. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(5): 919–931, May 2019. CO-DEN ???? ISSN 2329-9290.

Kim:2021:OED

[KGG^{+}21] Seokhwan Kim, Michel Galley, Chulaka Gunasekara, Sungjin Lee, Adam Atkinson, Baolin Peng, Hannes Schulz, Jianfeng Gao, Jinchao Li, Mahmoud Adada, Minlie Huang, Luis Lastras, Jonathan K. Kummerfeld, Walter S. Lasecki, Chiori Hori, Anoop Cherian, Tim K. Marks, Abhinav Rastogi, Xiaoxue Zang, Srinivas Sunkara, and Raghav Gupta. Overview of the Eighth Dialog System Technology Challenge: DSTC8. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2529–2540, 2021. CODEN ???? ISSN 2329-9290.

Kortlang:2016:AMB

[KGH^{+}16] Steffen Kortlang, Giso Grimm, Volker Hohmann, Birger Kollmeier, and Stephan D. Ewert. Auditory model-based dynamic compression controlled by subband instantaneous frequency and speech presence probability estimates. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(10): 1759–1772, ???? 2016. CODEN ???? ISSN 2329-9290.

Kallasjoki:2014:EUI

[KGP14] H. Kallasjoki, J. F. Gemmeke, and K. J. Palomaki. Estimating uncertainty to improve exemplar-based feature enhancement for noise robust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):368–380, February 2014. CODEN ???? ISSN 2329-9290.

Kelly:2016:SAC

[KH16] Finnian Kelly and John H. L. Hansen. Score-aging calibration for speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2414– 2424, ???? 2016. CODEN ???? ISSN 2329-9290.

Kelly:2021:ACL

[KH21] Finnian Kelly and John H. L. Hansen. Analysis and calibration of Lombard effect and whisper for speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):927–942, 2021. CO-DEN ???? ISSN 2329-9290.

Kothapally:2024:MSD

[KH24] Vinay Kothapally and John H. L. Hansen. Monaural speech dereverberation using deformable convolutional networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1712– 1723, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3358720.

Krebs:2015:IMS

[KHCW15] F. Krebs, A. Holzapfel, A. T. Cemgil, and G. Widmer. Inferring metrical structure in music using particle filters. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(5): 817–827, May 2015. CO-DEN ???? ISSN 2329-9290.

Kourkounakis:2021:FEE

[KHE21] Tedd Kourkounakis, Amirhossein Hajavi, and Ali Etemad. FluentNet: End-to-end detection of stuttered speech disfluencies with deep learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2986–2999, 2021. CODEN ???? ISSN 2329-9290.

Koutrouvelis:2019:CAR

[KHHJ19a] A. I. Koutrouvelis, R. C. Hendriks, R. Heusdens, and J. Jensen. A convex approximation of the relaxed binaural beamforming optimization problem. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(2):321–331, February 2019. CODEN ???? ISSN 2329-9290.

Koutrouvelis:2019:RJE

[KHHJ19b] A. I. Koutrouvelis, R. C.

Hendriks, R. Heusdens, and J. Jensen. Robust joint estimation of multimicrophone signal model parameters. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(7):1136– 1150, July 2019. CODEN ???? ISSN 2329-9290.

Khademi:2017:IEB

[KHK17] Seyran Khademi, Richard C.

Hendriks, and W. Bastiaan Kleijn. Intelligibility enhancement based on mutual information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):1694–1708, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7946152/.

Kanervisto:2022:OTS

[KHKY22] Anssi Kanervisto, Ville Hautamäki, Tomi Kinnunen, and Junichi Yamagishi. Optimizing tandem speaker verification and anti-spoofing systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):477–488, 2022. CODEN ???? ISSN 2329-9290.

Kameoka:2021:MMV

[KHT^{+}21] Hirokazu Kameoka, Wen-Chin Huang, Kou Tanaka, Takuhiro Kaneko, Nobukatsu Hojo, and Tomoki Toda. Many-to-many voice transformer network. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 656–670, 2021. CODEN ???? ISSN 2329-9290.

Kameoka:2018:NMF

[KHTL18] H. Kameoka, T. Higuchi, M. Tanaka, and L. Li. Nonnegative matrix factorization with basis clustering using cepstral distance regularization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (6):1029–1040, June 2018. CODEN ???? ISSN 2329-9290.

Kabzinski:2021:CCF

[KJ21] Tobias Kabzinski and Peter Jax. A causalityconstrained frequencydomain least-squares filter design method for crosstalk cancellation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):2942–2956, 2021. CODEN ???? ISSN 2329-9290.

Kamper:2016:UWS

[KJG16] Herman Kamper, Aren Jansen, and Sharon Goldwater. Unsupervised word segmentation and lexicon discovery using acoustic word embeddings. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 669–679, April 2016. CO-DEN ???? ISSN 2329-9290.

Kim:2014:DAB

[KK14] Seon Man Kim and Hong Kook Kim. Directionof-arrival based SNR estimation for dual-microphone speech enhancement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2207–2217, December 2014. CODEN ???? ISSN 2329-9290.

Kato:2019:SRM

[KK19a] A. Kato and T. H. Kinnunen. Statistical regression models for noise robust F0 estimation using recurrent deep neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2336– 2349, December 2019. CODEN ???? ISSN 2329-9290.

Koriyama:2019:SPS

[KK19b] T. Koriyama and T. Kobayashi.

Statistical parametric speech synthesis using deep Gaussian processes. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(5): 948–959, May 2019. CO-DEN ???? ISSN 2329-9290.

Kita:2023:SSL

[KK23] Shunsuke Kita and Yoshinobu Kajikawa. Sound source localization inside a structure under semi-supervised conditions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1397– 1408, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3263776.

Koutrouvelis:2016:FMH

[KKGH16] A. I. Koutrouvelis, G. P. Kafentzis, N. D. Gaubitch, and R. Heusdens. A fast method for highresolution voiced/unvoiced detection and glottal closure/opening instant estimation of speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(2): 316–328, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Kim:2018:LSI

[KKGK18] Y. Kim, M. Kim, J. Goo, and H. Kim. Learning selfinformed feature contribution for deep learningbased acoustic modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):2204–2214, November 2018. CODEN ???? ISSN 2329-9290.

Kim:2023:MSM

[KKJK23] Sungjae Kim, Yewon Kim, Jewoo Jun, and Injung Kim. MuSE-SVS: Multi-singer emotional singing voice synthesizer that controls emotional intensity. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2751– 2764, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3294712.

Kim:2017:DIW

[KKJN17] Jung-Hee Kim, Jin Kim, Jae Hyeon Jeon, and Sang Won Nam. Delayless individual-weighting-factors sign subband adaptive filter with banddependent variable stepsizes. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(7):1526–1534, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7914653/.

Kim:2015:AIA

[KKK15] Myung Jong Kim, Younggwan Kim, and Hoirin Kim. Automatic intelligibility assessment of dysarthric speech using phonologically-structured sparse linear model. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(4): 694–704, April 2015. CO-DEN ???? ISSN 2329-9290.

Kameoka:2019:AVN

[KKTH19] H. Kameoka, T. Kaneko, K. Tanaka, and N. Hojo. ACVAE-VC: Non-parallel voice conversion with auxiliary classifier variational autoencoder. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(9):1432– 1443, September 2019. CODEN ???? ISSN 2329-9290.

Kim:2024:ASR

[KLCL24] Seung-Bin Kim, Sang-Hoon Lee, Ha-Yeong Choi, and Seong-Whan Lee. Audio superresolution with robust speech representation learning of masked autoencoder. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1012– 1022, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3349053.

Kanda:2017:MPB

[KLK17] Naoyuki Kanda, Xugang Lu, and Hisashi Kawai. Maximum-a-posteriori-based decoding for end-to-end acoustic models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(5):1023– 1034, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7870575/.

Kong:2021:HRP

[KLS^{+}21] Qiuqiang Kong, Bochen Li, Xuchen Song, Yuan Wan, and Yuxuan Wang. High-resolution piano transcription with pedals by regressing onset and offset times. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3707–3717, 2021. CODEN ???? ISSN 2329-9290.

Kan:2021:TLE

[KM21] Alan Kan and Qinglin Meng. The temporal limits encoder as a sound coding strategy for bilateral cochlear implants. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):265–273, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3039601.

Krause:2023:HCI

[KM23] Michael Krause and Meinard Müller. Hierarchical classification for instrument activity detection in orchestral music recordings. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2567–2578, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3291506.

Kheder:2018:UJM

[KMAB18] W. B. Kheder, D. Matrouf, M. Ajili, and J. Bonastre. A unified joint model to deal with nuisance variabilities in the i-vector space. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):633–645, March 2018. CODEN ???? ISSN 2329-9290.

Koldovsky:2015:SSS

[KMG15] Z. Koldovsky, J. Malek, and S. Gannot. Spatial source subtraction based on incomplete measurements of relative transfer function. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):1335– 1347, August 2015. CO-DEN ???? ISSN 2329-9290.

Koluguri:2017:SEU

[KMG17] Nithin Rao Koluguri, G. Nisha Meenakshi, and Prasanta Kumar Ghosh. Spectrogram enhancement using Multiple Window Savitzky–Golay (MWSG) filter for robust bird sound detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (6):1183–1192, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7933047/.

Kamper:2021:IAW

[KMG21] Herman Kamper, Yevgen Matusevych, and Sharon Goldwater. Improved acoustic word embeddings for zero-resource languages using multilingual transfer. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1107–1118, 2021. CODEN ???? ISSN 2329-9290.

Khan:2018:UVS

[KML18] F. U. Khan, B. P. Milner, and T. Le Cornu. Using visual speech information in masking methods for audio speaker separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):1742–1754, October 2018. CODEN ???? ISSN 2329-9290.

Katzberg:2018:CSF

[KMM^{+}18] F. Katzberg, R. Mazur, M. Maass, P. Koch, and A. Mertins. A compressed sensing framework for dynamic sound-field measurements. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):1962– 1975, November 2018. CODEN ???? ISSN 2329-9290.

Kavalekalam:2019:MBS

[KNBC19] M. S. Kavalekalam, J. K. Nielsen, J. B. Boldt, and M. G. Christensen. Model-based speech enhancement for intelligibility improvement in binaural hearing aids. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1): 99–113, January 2019. CODEN ???? ISSN 2329-9290.

Koizumi:2017:IAF

[KNH^{+}17] Yuma Koizumi, Kenta Niwa, Yusuke Hioka, Kazunori Kobayashi, and Hitoshi Ohmuro. Informative acoustic feature selection to maximize mutual information for collecting target sources. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):768–779, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7839197/.

Koizumi:2018:DBS

[KNH^{+}18] Y. Koizumi, K. Niwa, Y. Hioka, K. Kobayashi, and Y. Haneda. DNNbased source enhancement to increase objective sound quality assessment score. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(10):1780–1792, October 2018. CODEN ???? ISSN 2329-9290.

Kitamura:2016:DBS

[KOS^{+}16] Daichi Kitamura, Nobutaka Ono, Hiroshi Sawada, Hirokazu Kameoka, and Hiroshi Saruwatari. Determined blind source separation unifying independent vector analysis and nonnegative matrix factorization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(9):1626– 1641, ???? 2016. CODEN ???? ISSN 2329-9290.

Koudounas:2024:TCS

[KPA^{+}24] Alkis Koudounas, Eliana Pastor, Giuseppe Attanasio, Vittorio Mazzia, Manuel Giollo, Thomas Gueudre, Elisa Reale, Luca Cagliero, Sandro Cumani, Luca de Alfaro, Elena Baralis, and Daniele Amberti. Towards comprehensive subgroup performance analysis in speech models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1468–1480, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3363447.

Khoubrouy:2015:HDH

[KPH15] S. A. Khoubrouy, I. M. S. Panahi, and J. H. L. Hansen. Howling detection in hearing aids based on generalized Teager– Kaiser operator. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(1): 154–161, January 2015. CODEN ???? ISSN 2329-9290.

Kim:2016:PNC

[KS16] Chanwoo Kim and Richard M. Stern. Power-normalized cepstral coefficients (PNCC) for robust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(7):1315–1329, July 2016. CODEN ???? ISSN 2329-9290.

Kose:2021:MRS

[KS21] Öykü Deniz Köse and Murat Saraçlar. Multimodal representations for synchronized speech and realtime MRI video processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1912–1924, 2021. CODEN ???? ISSN 2329-9290.

Kim:2021:ESI

[KSG^{+}21] Seokhwan Kim, Hannes Schulz, Chulaka Gunasekara, Chiori Hori, Abhinav Rastogi, and Luis Fernando D. Haro. Editorial: Special issue on the eighth dialog system technology challenge. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2434–2436, 2021. CODEN ???? ISSN 2329-9290.

Krishna:2024:RLH

[KSG24] Varun Krishna, Tarun Sai, and Sriram Ganapathy. Representation learning with hidden unit clustering for low resource speech applications. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1036– 1047, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3337670.

Kaushik:2017:ASD

[KSH17] Lakshmish Kaushik, Abhijeet Sangwan, and John H. L. Hansen. Automatic sentiment detection in naturalistic audio. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):1668–1679, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7870612/.

Koutrouvelis:2018:LCR

[KSHH18] A. I. Koutrouvelis, T. W. Sherson, R. Heusdens, and R. C. Hendriks. A low-cost robust distributed linearly constrained beamformer for wireless acoustic sensor networks with arbitrary topology. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(8):1434– 1448, August 2018. CO-DEN ???? ISSN 2329-9290.

Kitamura:2015:MSS

[KSK^{+}15] D. Kitamura, H. Saruwatari, H. Kameoka, Yu. Takahashi, K. Kondo, and S. Nakamura. Multichannel signal separation combining directional clustering and nonnegative matrix factorization with spectrogram restoration. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (4):654–669, April 2015. CODEN ???? ISSN 2329-9290.

Kim:2021:RWN

[KSK21] Byoungjae Kim, Jungyun Seo, and Myoung-Wan Koo. Randomly wired network based on RoBERTa and dialog history attention for response selection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2437–2442, 2021. CODEN ???? ISSN 2329-9290.

Kamper:2019:SSR

[KSL19] H. Kamper, G. Shakhnarovich, and K. Livescu. Semantic speech retrieval with a visually grounded model of untranscribed speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (1):89–98, January 2019. CODEN ???? ISSN 2329-9290.

Koizumi:2019:UDA

[KSU^{+}19] Y. Koizumi, S. Saito, H. Uematsu, Y. Kawachi, and N. Harada. Unsupervised detection of anomalous sound based on deep learning and the Neyman– Pearson lemma. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1): 212–224, January 2019. CODEN ???? ISSN 2329-9290.

Kukanov:2020:MFM

[KTH^{+}20] Ivan Kukanov, Trung Ngo Trong, Ville Hautamäki, Sabato Marco Siniscalchi, Valerio Mario Salerno, and Kong Aik Lee. Maximal figure-of-merit framework to detect multilabel phonetic features for spoken language recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):682–695, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.2964953.

Kolbaek:2019:RBS

[KTJ19] M. Kolbæk, Z. Tan, and J. Jensen. On the relationship between short-time objective intelligibility and short-time spectralamplitude mean-square error for speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(2):283–295, February 2019. CODEN ???? ISSN 2329-9290.

Kataria:2024:TDS

[KVMV^{+}24] Saurabh Kataria, Jesús Villalba, Laureano Moro-Velázquez, Piotr Żelasko, and Najim Dehak. Timedomain speech superresolution with GAN based modeling for telephony speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1736– 1749, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3369536.

Karanasou:2017:VSN

[KWGW17] Penny Karanasou, Chunyang Wu, Mark Gales, and Philip C. Woodland. I-vectors and structured neural networks for rapid adaptation of acoustic models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):818–828, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7857716/.

Kong:2024:AES

[KWZ24] Jun Kong, Jin Wang, and Xuejie Zhang. Adaptive ensemble self-distillation with consistent gradients for fast inference of pretrained language models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):430– 442, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3331080.

Kong:2019:SED

[KXS^{+}19] Q. Kong, Y. Xu, I. Sobieraj, W. Wang, and M. D. Plumbley. Sound event detection and time frequency segmentation from weakly labelled data. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (4):777–787, April 2019. CODEN ???? ISSN 2329-9290.

Kameoka:2015:GMV

[KYI^{+}15] H. Kameoka, K. Yoshizato, T. Ishihara, K. Kadowaki, Y. Ohishi, and K. Kashino. Generative modeling of voice fundamental frequency contours. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(6):1042–1053, June 2015. CODEN ???? ISSN 2329-9290.

Karo:2024:CTD

[KYL24] Matan Karo, Arie Yeredor, and Itshak Lapidot. Compact timedomain representation for logical access spoofed audio. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):946– 958, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3341000.

Kolbaek:2017:MSS

[KYTJ17] Morten Kolbæk, Dong Yu, Zheng-Hua Tan, and Jesper Jensen. Multitalker speech separation with utterance-level permutation invariant training of deep recurrent neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(10): 1901–1913, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7979557/.

Ling:2018:WMG

[LAGD18] Z. Ling, Y. Ai, Y. Gu, and L. Dai. Waveform modeling and generation using hierarchical recurrent neural networks for speech bandwidth extension. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(5):883–894, May 2018. CODEN ???? ISSN 2329-9290.

Li:2016:SSA

[LAK16] Sheng Li, Yuya Akita, and Tatsuya Kawahara. Semi-supervised acoustic model training by discriminative data selection from multiple ASR systems’ hypotheses. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(9): 1524–1534, ???? 2016. CODEN ???? ISSN 2329-9290.

Lubeck:2023:SUS

[LAP23] Tim Lübeck, Johannes M.

Arend, and Christoph Pörschmann. Spatial upsampling of sparse spherical microphone array signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1163– 1174, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3250841.

Liu:2016:ABE

[LB16] Xin Liu and Changchun Bao. Audio bandwidth extension based on ensemble echo state networks with temporal evolution. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (3):594–607, March 2016. CODEN ???? ISSN 2329-9290.

Lotfian:2019:CLS

[LB19] R. Lotfian and C. Busso. Curriculum learning for speech emotion recognition from crowdsourced labels. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(4):815– 826, April 2019. CODEN ???? ISSN 2329-9290.

Lin:2023:SML

[LB23] Wei-Cheng Lin and Carlos Busso. Sequential modeling by leveraging non-uniform distribution of speech emotion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1087–1099, 2023. ISSN 2329-

                                        9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2023.3244527.

Leglaive:2016:MAS

[LBR16] Simon Leglaive, Roland Badeau, and Gaël Richard. Multichannel audio source separation with probabilistic reverberation priors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2453–2465, ???? 2016. CODEN ???? ISSN 2329-9290.

Leglaive:2018:SSM

[LBR18] S. Leglaive, R. Badeau, and G. Richard. Student’s t-source and mixing models for multichannel audio source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (6):1154–1168, June 2018. CODEN ???? ISSN 2329-9290.

Lee:2016:PLC

[LC16a] Bong-Ki Lee and Joon-Hyuk Chang. Packet loss concealment based on deep neural networks for digital speech transmission. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(2):378–387, ???? 2016. CODEN ???? ISSN 2329-

Liu:2016:ACM

[LC16b] Jinxin Liu and Xuefeng Chen. Adaptive compensation of misequalization in narrowband active noise equalizer systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2390–2399, ???? 2016. CODEN ???? ISSN 2329-9290.

Lee:2024:PAW

[LC24] Jae-Hong Lee and Joon-Hyuk Chang. Partitioning attention weight: Mitigating adverse effect of incorrect pseudo-labels for selfsupervised ASR. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 891–905, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3343615.

Loweimi:2023:PEA

[LCB^{+}23] Erfan Loweimi, Andrea Carmantini, Peter Bell, Steve Renals, and Zoran Cvetkovic. Phonetic error analysis beyond phone error rate. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3346– 3361, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313417.

Liu:2015:CRL

[LCC^{+}15] Shih-Hung Liu, Kuan-Yu Chen, B. Chen, Hsin-Min Wang, Hsu-Chun Yen, and Wen-Lian Hsu. Combining relevance language modeling and clarity measure for extractive speech summarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(6):957–969, June 2015. CODEN ???? ISSN 2329-9290.

Lu:2024:HTA

[LCG^{+}24] Ruiying Lu, Bo Chen, Dandan Guo, Dongsheng Wang, and Mingyuan Zhou. Hierarchical topicaware contextualized transformers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):841– 852, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3339344.

Lin:2021:PTD

[LCL^{+}21] Zehao Lin, Shaobo Cui, Guodun Li, Xiaoming Kang, Feng Ji, Fenglin Li, Zhongzhou Zhao, Haiqing Chen, and Yin Zhang. Predict-then-decide: a predictive approach for wait or answer task in dialogue systems. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 3012–3024, 2021. CODEN ???? ISSN 2329-9290.

Lehmann:2023:TMP

[LCL23a] Pedro Izquierdo Lehmann, Rodrigo F. Cádiz, and Carlos A. Sing Long. Towards maximizing a perceptual Sweet Spot for spatial sound with loudspeakers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3304– 3319, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3190729.

Liu:2023:VLS

[LCL^{+}23b] Hong Liu, Yucheng Cai, Zhenru Lin, Zhijian Ou, Yi Huang, and Junlan Feng. Variational latent-state GPT for semisupervised task-oriented dialog systems. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 970–984, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3240661.

LeLan:2018:AMC

[LCLM18] G. Le Lan, D. Charlet, A. Larcher, and S. Meignier. An adaptive method for cross-recording speaker diarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(10):1821– 1832, October 2018. CO-DEN ???? ISSN 2329-9290.

Luo:2018:SIS

[LCM18] Y. Luo, Z. Chen, and N. Mesgarani. Speakerindependent speech separation with deep attractor network. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(4): 787–796, April 2018. CO-DEN ???? ISSN 2329-9290.

Li:2022:DPM

[LCQ22] Chenda Li, Zhuo Chen, and Yanmin Qian. Dualpath modeling with memory embedding model for continuous speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1508– 1520, 2022. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3165712.

Liu:2023:DFN

[LCQ23] Bei Liu, Zhengyang Chen, and Yanmin Qian. Depthfirst neural architecture with attentive feature fusion for efficient speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1825– 1838, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3273417.

Lin:2024:MIF

[LCRY24] Changkai Lin, Hongju Cheng, Qiang Rao, and Yang Yang. M^{3}SA: Multimodal sentiment analysis based on multi-scale feature extraction and multitask learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1416– 1429, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3361374.

Li:2019:IMD

[LCSL19] W. Li, N. F. Chen, S. M. Siniscalchi, and C. Lee. Improving mispronunciation detection of Mandarin tones for non-native learners with soft-target tone labels and BLSTMbased deep tone models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2012–2024, December 2019. CODEN ???? ISSN 2329-9290.

Luo:2019:EVC

[LCTA19] Z. Luo, J. Chen, T. Takiguchi, and Y. Ariki. Emotional voice conversion using dual supervised adversarial networks with continuous wavelet transform F0 features. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(10): 1535–1548, October 2019. CODEN ???? ISSN 2329-9290.

Liu:2016:TEL

[LCW^{+}16] Xunying Liu, Xie Chen, Yongqiang Wang, Mark J. F. Gales, and Philip C. Woodland. Two efficient lattice rescoring methods using recurrent neural network language models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (8):1438–1449, ???? 2016. CODEN ???? ISSN 2329-9290.

Lee:2018:ISC

[LCW^{+}18] H. Lee, P. Chung, Y. Wu, T. Lin, and T. Wen. Interactive spoken content retrieval by deep reinforcement learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2447–2459, December 2018. CODEN ???? ISSN 2329-9290.

Liu:2021:MVC

[LCW^{+}21] Songxiang Liu, Yuewen Cao, Disong Wang, Xixin Wu, Xunying Liu, and Helen Meng. Anyto-many voice conversion with location-relative sequence-to-sequence modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1717–1728, 2021. CODEN ???? ISSN 2329-9290.

Li:2024:UDL

[LCW24] Linjian Li, Yi Cai, and Xin Wu. Unsupervised disentanglement learning model for exemplarguided paraphrase generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1885– 1900, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3374124.

Liao:2016:SAS

[LCWC16] I-Bin Liao, Chen-Yu Chiang, Yih-Ru Wang, and Sin-Horng Chen. Speaker adaptation of SR-HPM for speaking rate-controlled Mandarin TTS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):2046–2058, ???? 2016. CODEN ???? ISSN 2329-9290.

Lu:2023:ISE

[LCY^{+}23] Yen-Ju Lu, Chia-Yu Chang, Cheng Yu, Ching-Feng Liu, Jeih weih Hung, Shinji Watanabe, and Yu Tsao. Improving speech enhancement performance by leveraging contextual broad phonetic class information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2738–2750, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3288418.

Liu:2021:HRB

[LCYW21] Qing Liu, Lei Chen, Yuan Yuan, and Huarui Wu. History reuse and bag-of-words loss for long summary generation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2551–2560, 2021. CODEN ???? ISSN 2329-9290.

Li:2017:CPT

[LCZ^{+}17] Zhenghua Li, Jiayuan Chao, Min Zhang, Wenliang Chen, Meishan Zhang, and Guohong Fu. Coupled POS tagging on heterogeneous annotations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(3):557–571, ???? 2017. CODEN ???? ISSN 2329-9290.

Li:2019:SAE

[LCZ^{+}19] P. Li, C. Chen, W. Zheng, Y. Deng, F. Ye, and Z. Zheng. STD: an automatic evaluation metric for machine translation based on word embeddings. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(10):1497–1506, October 2019. CODEN ???? ISSN 2329-9290.

Lian:2023:PAP

[LCZ23] Zhicheng Lian, Haonan Cheng, and Jiawan Zhang. PQG-A2SA: Performance quantification guided audio-to-score alignment for orchestral music. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1982– 1992, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3277290.

Li:2016:ASF

[LD16] Bochen Li and Zhiyao Duan. An approach to score following for piano performances with the sustained effect. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(12): 2425–2438, ???? 2016. CODEN ???? ISSN 2329-9290.

Li:2023:AVE

[LDG^{+}23] Guinan Li, Jiajun Deng, Mengzhe Geng, Zengrui Jin, Tianzi Wang, Shujie Hu, Mingyu Cui, Helen Meng, and Xunying Liu. Audio-visual end-to-end multi-channel speech separation, dereverberation and recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2707– 2723, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3294705.

Li:2014:ONR

[LDGHU14] Jinyu Li, Li Deng, Yifan Gong, and R. Haeb-Umbach. An overview of noise-robust automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (4):745–777, April 2014. CODEN ???? ISSN 2329-9290.

Lin:2021:SEU

[LdLvWWS21] Ju Lin, Adriaan J.

de Lind van Wijngaarden, Kuang-Ching Wang, and Melissa C. Smith. Speech enhancement using multistage self-attentive temporal convolutional networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3440– 3450, 2021. CODEN ???? ISSN 2329-9290.

Lu:2019:AVD

[LDZ19] R. Lu, Z. Duan, and C. Zhang. Audio visual deep clustering for speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(11):1697– 1712, November 2019. CODEN ???? ISSN 2329-9290.

Lopez-Espejo:2021:NLF

[LETJ21] Iván López-Espejo, Zheng-Hua Tan, and Jesper Jensen. A novel loss function and training strategy for noise-robust keyword spotting. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2254–2266, 2021. CODEN ???? ISSN 2329-9290.

Lorente:2014:GIM

[LFdDG14] J. Lorente, M. Ferrer, M. de Diego, and A. Gonzalez. GPU implementation of multichannel adaptive algorithms for local active noise control. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (11):1624–1635, November 2014. CODEN ???? ISSN 2329-9290.

Li:2015:ISS

[LFH^{+}15] H. Li, M. Federico, X. He, H. Meng, and I. Trancoso. Introduction to the special section on continuous space and related methods in natural language processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (3):427–430, March 2015. CODEN ???? ISSN 2329-9290.

Lin:2023:CXC

[LFL^{+}23] Nankai Lin, Yingwen Fu, Xiaotian Lin, Dong Zhou, Aimin Yang, and Shengyi Jiang. CL-XABSA: Contrastive learning for crosslingual aspect-based sentiment analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 2935–2946, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3297964.

Lyu:2023:HEH

[LFM^{+}23] Boyang Lyu, Chunxiao Fan, Yue Ming, Panzi Zhao, and Nannan Hu. En-HACN: Enhancing hybrid architecture with fast attention and capsule network for end-to-end speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1050– 1062, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3245407.

Leem:2024:SAF

[LFO^{+}24] Seong-Gyun Leem, Daniel Fulford, Jukka-Pekka Onnela, David Gard, and Carlos Busso. Selective acoustic feature enhancement for speech emotion recognition with noisy speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 917–929, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3340603.

Liu:2019:BBF

[LFQL19] D. Liu, J. Fu, Q. Qu, and J. Lv. BFGAN: Backward and forward generative adversarial networks for lexically constrained sentence generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2350–2361, December 2019. CODEN ???? ISSN 2329-9290.

Liu:2017:TQE

[LFU^{+}17] Lemao Liu, Atsushi Fujita, Masao Utiyama, Andrew Finch, and Eiichiro Sumita. Translation quality estimation using only bilingual corpora. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(9): 1762–1772, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7949019/.

Liu:2023:SEC

[LFWZ23] Yinggang Liu, Hong Fu, Ying Wei, and Hanbing Zhang. Sound event classification based on frequency-energy feature representation and twostage data dimension reduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1290– 1304, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3260708.

Liu:2016:PAF

[LG16] Jianming Liu and Steven L.

Grant. Proportionate adaptive filtering for block-sparse system identification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(4):623– 630, April 2016. CODEN ???? ISSN 2329-9290.

Laufer:2019:BHM

[LG19] Y. Laufer and S. Gannot.

A Bayesian hierarchical model for speech enhancement with time-varying audio channel. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1): 225–239, January 2019. CODEN ???? ISSN 2329-9290.

Laufer:2020:SBM

[LG20] Yaron Laufer and Sharon Gannot. Scoring-based ML estimation and CRBs for reverberation, speech, and noise PSDs in a spatially homogeneous noise field. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):61–76, January 2020. CODEN ???? ISSN 2329-9290.

                                        URL https://dl.acm.
                                        org/doi/abs/10.1109/
                                        TASLP.2019.2948794.

Li:2018:MIN

[LGGH18] X. Li, S. Gannot, L. Girin, and R. Horaud. Multichannel identification and nonnegative equalization for dereverberation and noise reduction based on convolutive transfer function. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(10):1755–1768, October 2018. CODEN ???? ISSN 2329-9290.

Li:2019:MOD

[LGGH19a] X. Li, L. Girin, S. Gannot, and R. Horaud. Multichannel online dereverberation based on spectral magnitude inverse filtering. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(9):1365–1377, September 2019. CODEN ???? ISSN 2329-9290.

Li:2019:MSS

[LGGH19b] X. Li, L. Girin, S. Gannot, and R. Horaud. Multichannel speech separation and enhancement using the convolutive transfer function. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(3): 645–659, March 2019. CODEN ???? ISSN 2329-9290.

Liu:2021:RPC

[LGH^{+}21] Shansong Liu, Mengzhe Geng, Shoukang Hu, Xurong Xie, Mingyu Cui, Jianwei Yu, Xunying Liu, and Helen Meng. Recent progress in the CUHK dysarthric speech recognition system. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2267–2281, 2021. CODEN ???? ISSN 2329-9290.

Li:2017:MSL

[LGHG17] Xiaofei Li, Laurent Girin, Radu Horaud, and Sharon Gannot. Multiple-speaker localization based on direct-path features and likelihood maximization with spatial sparsity regularization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(10):1997– 2012, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/8010441/.

Liu:2022:RRO

[LGL^{+}22] Qianying Liu, Wenyu Guan, Sujian Li, Fei Cheng, Daisuke Kawahara, and Sadao Kurohashi. RODA: Reverse operation based data augmentation for solving math word problems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):1–11, 2022. CODEN ???? ISSN 2329-9290.

Lu:2024:TBO

[LGL^{+}24] Hua Lu, Zhen Guo, Chanjuan Li, Yunyi Yang, Huang He, and Siqi Bao. Towards building an open-domain dialogue system incorporated with Internet memes. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 721–726, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3288413.

Lu:2014:CLS

[LGR14] Liang Lu, A. Ghoshal, and S. Renals. Crosslingual subspace Gaussian mixture models for lowresource speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):17–27, January 2014. CODEN ???? ISSN 2329-9290.

Laufer-Goldshtein:2016:SSS

[LGTG16] Bracha Laufer-Goldshtein, Ronen Talmon, and Sharon Gannot. Semi-supervised sound source localization based on manifold regularization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(8):1393– 1407, ???? 2016. CODEN ???? ISSN 2329-9290.

Laufer-Goldshtein:2017:SSS

[LGTG17] Bracha Laufer-Goldshtein, Ronen Talmon, and Sharon Gannot. Semi-supervised source localization on multiple manifolds with distributed microphones. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (7):1477–1491, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7906609/.

Laufer-Goldshtein:2018:HAS

[LGTG18] B. Laufer-Goldshtein, R. Talmon, and S. Gannot. A hybrid approach for speaker tracking based on TDOA and datadriven models. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(4): 725–735, April 2018. CO-DEN ???? ISSN 2329-

Liu:2014:IBE

[LH14] Gang Liu and J. H. L. Hansen. An investigation into back-end advancements for speaker recognition in multi-session and noisy enrollment scenarios. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1978–1992, December 2014. CODEN ???? ISSN 2329-9290.

Li:2023:ALS

[LH23] Changheng Li and Richard C. Hendriks. Alternating least-squares-based microphone array parameter estimation for a single-source reverberant and noisy acoustic scenario. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3922– 3934, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3306713.

Li:2023:DTD

[LHC^{+}23] Tao Li, Chenxu Hu, Jian Cong, Xinfa Zhu, Jingbei Li, Qiao Tian, Yuping Wang, and Lei Xie. DiCLET-TTS: Diffusion model based cross-lingual emotion transfer for textto-speech — a study between English and Mandarin. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3418– 3430, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313413.

Leijon:2016:BAP

[LHD16] Arne Leijon, Gustav Eje Henter, and Martin Dahlquist. Bayesian analysis of phoneme confusion matrices. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): 469–482, March 2016. CODEN ???? ISSN 2329-9290.

Liu:2024:HMG

[LHG^{+}24] Tengfei Liu, Yongli Hu, Junbin Gao, Yanfeng Sun, and Baocai Yin. Hierarchical multi-granularity interaction graph convolutional network for long document classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1762– 1775, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3369530.

Luo:2023:DMD

[LHJ^{+}23] Xueqin Luo, Gongping Huang, Jilu Jin, Jingdong Chen, Jacob Benesty, Wen Zhang, Mengyao Zhu, and Chunjian Li. Design of maximum directivity beamformers with linear acoustic vector sensor arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1421– 1435, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3256093.

Liu:2022:SSS

[LHL^{+}22a] Yongkang Liu, Qingbao Huang, Jing Li, Linzhang Mo, Yi Cai, and Qing Li. SSAP: Storylines and sentiment aware pretrained model for story ending generation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 686–694, 2022. CODEN ???? ISSN 2329-9290.

Lu:2022:SFR

[LHL^{+}22b] Menglong Lu, Zhen Huang, Binyang Li, Yunxiang Zhao, Zheng Qin, and DongSheng Li. SIFTER: a framework for robust rumor detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 429–442, 2022. CODEN ???? ISSN 2329-9290.

Li:2023:MLS

[LHL^{+}23a] Shu’ang Li, Xuming Hu, Li Lin, Aiwei Liu, Lijie Wen, and Philip S. Yu. A multi-level supervised contrastive learning framework for lowresource natural language inference. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1771– 1783, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3270771.

Li:2023:NRR

[LHL^{+}23b] Yu Li, Bojie Hu, Jian Liu, Yufeng Chen, and Jinan Xu. A neighborhood re-ranking model with relation constraint for knowledge graph completion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):411– 425, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3225537.

Luo:2021:GCC

[LHM21] Yi Luo, Cong Han, and Nima Mesgarani. Group communication with context codec for lightweight source separation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1752–1761, 2021. CODEN ???? ISSN 2329-9290.

Liao:2021:EES

[LHW^{+}21] Xianwen Liao, Yongzhong Huang, Yongzhuang Wei, Chenhao Zhang, Fu Wang, and Yong Wang. Efficient estimate of sentence’s representation based on the difference semantics model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3384– 3399, 2021. CODEN ???? ISSN 2329-9290.

Liu:2023:UFA

[LHYJ23] Yuanzhi Liu, Min He, Qingqing Yang, and Gwanggil Jeon. An unsupervised framework with attention mechanism and embedding perturbed encoder for non-parallel text sentiment style transfer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2134–2144, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3278185.

Liu:2024:TSL

[LHZ^{+}24] Rui Liu, Yifan Hu, Haolin Zuo, Zhaojie Luo, Longbiao Wang, and Guanglai Gao. Text-to-speech for low-resource agglutinative language with morphology-aware language model pretraining. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1075– 1087, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3348762.

Li:2015:IEE

[Li15] H Li. Inaugural editorial: Embracing new opportunities for growth. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (1):5–6, January 2015. CODEN ???? ISSN 2329-9290.

Liberatore:2021:NNV

[Lib21] Christopher Liberatore.

Native-nonnative voice conversion by residual warping in a sparse, anchor-based representation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3040–3051, 2021. CODEN ???? ISSN 2329-9290.

Lin:2018:RRL

[Lin18] S. Lin. Reverberationrobust localization of speakers using distinct speech onsets and multichannel cross correlations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):2098–2111, November 2018. CODEN ???? ISSN 2329-9290.

Lin:2019:RPE

[Lin19] S. Lin. Robust pitch estimation and tracking for speakers based on subband encoding and the generalized labeled multi-Bernoulli filter. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(4): 827–841, April 2019. CO-DEN ???? ISSN 2329-9290.

Li:2024:LSV

[LJH^{+}24] Yanxiong Li, Zhongjie Jiang, Qisheng Huang, Wenchang Cao, and Jialong Li. Lightweight speaker verification using transformation module with feature partition and fusion. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 794–806, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3338533.

Li:2017:SQS

[LJK17] Songbin Li, Yizhen Jia, and C.-C. Jay Kuo. Steganalysis of QIM steganography in low-bit-rate speech signals. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(5): 1011–1022, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7867798/.

Liu:2023:TGT

[LJL^{+}23] Jiang Liu, Donghong Ji, Jingye Li, Dongdong Xie, Chong Teng, Liang Zhao, and Fei Li. TOE: a grid-tagging discontinuous NER model enhanced by embedding Tag/ word relations and more finegrained tags. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 177–187, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3221009.

Li:2023:MSS

[LJWC23] Lu Li, Maoshen Jia, Jing Wang, and Ruiyuan Cao. Multiple-speech-source DOA estimation based on single-source cluster detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 3667–3680, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3321213.

Lin:2016:UCL

[LJX^{+}16] Zheng Lin, Xiaolong Jin, Xueke Xu, Yuanzhuo Wang, Xueqi Cheng, Weiping Wang, and Dan Meng. An unsupervised cross-lingual topic model framework for sentiment classification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): 432–444, March 2016. CODEN ???? ISSN 2329-

Liu:2016:GBS

[LK16] Yuzong Liu and Katrin Kirchhoff. Graph-based semisupervised learning for acoustic modeling in automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):1946–1956, ???? 2016. CODEN ???? ISSN 2329-9290.

Lee:2019:JLA

[LK19] J. Lee and H. Kang. A joint learning algorithm for complex-valued T–F masks in deep learningbased single-channel speech enhancement systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (6):1098–1108, June 2019. CODEN ???? ISSN 2329-9290.

Lee:2021:GRC

[LKC^{+}21] Hyeonseung Lee, Woo Hyun Kang, Sung Jun Cheon, Hyeongju Kim, and Nam Soo Kim. Gated recurrent context: Softmax-free attention for online encoderdecoder speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):710–719, 2021. CODEN ???? ISSN 2329-9290.

Leutnant:2014:NOM

[LKHU14] V. Leutnant, A. Krueger, and R. Haeb-Umbach. A new observation model in the logarithmic mel power spectral domain for the automatic recognition of noisy reverberant speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (1):95–109, January 2014. CODEN ???? ISSN 2329-9290.

Li:2023:FIA

[LKM23] Li Li, Hirokazu Kameoka, and Shoji Makino. Fast-MVAE2: On improving and accelerating the fast variational autoencoderbased source separation algorithm for determined mixtures. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):96– 110, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3214763.

Laroche:2018:HPN

[LKPR18] C. Laroche, M. Kowalski, H. Papadopoulos, and G. Richard. Hybrid projective nonnegative matrix factorization with drum dictionaries for harmonic/percussive source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1499– 1511, September 2018. CODEN ???? ISSN 2329-9290.

Liu:2023:DTI

[LKR23] Zheng Liu, Xin Kang, and Fuji Ren. Dual-TBNet: Improving the robustness of speech features via dual-transformer-BiLSTM for speech emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2193– 2203, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3282092.

Lee:2022:NAF

[LLC22] Moa Lee, Junmo Lee, and Joon-Hyuk Chang. Nonautoregressive fully parallel deep convolutional neural speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1150–1159, 2022. CODEN ???? ISSN 2329-9290.

Liu:2023:PDF

[LLC23a] Chang Liu, Zhen-Hua Ling, and Ling-Hui Chen. Pronunciation dictionaryfree multilingual speech synthesis using learned phonetic representations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):3706–3716, 2023. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2023.3313424.

Liu:2023:LID

[LLC^{+}23b] Hanmeng Liu, Jian Liu, Leyang Cui, Zhiyang Teng, Nan Duan, Ming Zhou, and Yue Zhang. LogiQA 2.0 — an improved dataset for logical reasoning in natural language understanding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2947– 2962, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3293046.

Li:2024:JMC

[LLC^{+}24] Jingbei Li, Sipan Li, Ping Chen, Luwen Zhang, Yi Meng, Zhiyong Wu, Helen Meng, Qiao Tian, Yuping Wang, and Yuxuan Wang. Joint multiscale cross-lingual speaking style transfer with bidirectional attention mechanism for automatic dubbing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):517– 528, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3331813.

Liu:2022:ECO

[LLCT22] Shengcai Liu, Ning Lu, Cheng Chen, and Ke Tang. Efficient combinatorial optimization for wordlevel adversarial textual attack. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):98–111, 2022. CODEN ???? ISSN 2329-9290.

Liu:2023:HRM

[LLdCF23] Tongzheng Liu, Zhihua Lu, João Paulo J. da Costa, and Tai Fei. A hybrid reverberation model and its application to joint speech dereverberation and separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3000– 3014, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3301227.

Laufer:2020:MEC

[LLGG20] Yaron Laufer, Bracha Laufer-Goldshtein, and Sharon Gannot. ML estimation and CRBs for reverberation, speech, and noise PSDs in rankdeficient noise field. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 619–634, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2962689.

Li:2024:CLT

[LLHW24] Xiao Li, Ruirui Liu, Huichou Huang, and Qingyao Wu. Contrastive learning for target speaker extraction with attentionbased fusion. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 178–188, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3324550.

Li:2023:PTT

[LLJ^{+}23] Yachao Li, Junhui Li, Jing Jiang, Shimin Tao, Hao Yang, and Min Zhang. P-transformer: Towards better document-to-document neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3859– 3870, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313445.

Luo:2019:IAT

[LLL^{+}19] H. Luo, T. Li, B. Liu, B. Wang, and H. Unger. Improving aspect term extraction with bidirectional dependency tree representation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(7):1201– 1212, July 2019. CODEN ???? ISSN 2329-9290.

Lu:2022:EMS

[LLL^{+}22] Ziyao Lu, Xiang Li, Yang Liu, Chulun Zhou, Jianwei Cui, Bin Wang, Min Zhang, and Jinsong Su. Exploring multi-stage information interactions for multi-source neural machine translation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 562–570, 2022. CODEN ???? ISSN 2329-9290.

Luo:2023:DSI

[LLL^{+}23] Zhaojie Luo, Shoufeng Lin, Rui Liu, Jun Baba, Yuichiro Yoshikawa, and Hiroshi Ishiguro. Decoupling speaker-independent emotions for voice conversion via source-filter networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):11– 24, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3190715.

Liu:2024:CPE

[LLL24] Lei Liu, Li Liu, and Haizhou Li. Computation and parameter efficient multi-modal fusion transformer for cued speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1559– 1572, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3363446.

Liu:2019:AAV

[LLLL19] Y. Liu, T. Lee, T. Law, and K. Y. Lee. Acoustical assessment of voice disorder with continuous speech using ASR posterior features. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(6): 1047–1059, June 2019. CODEN ???? ISSN 2329-9290.

Lafay:2016:MMS

[LLR^{+}16] Grégoire Lafay, Mathieu Lagrange, Mathias Rossignol, Emmanouil Benetos, and Axel Roebel. A morphological model for simulating acoustic scenes and its application to sound event detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1854–1864, ???? 2016. CODEN ???? ISSN 2329-9290.

Lu:2024:CXP

[LLS24] Haisheng Lu, Jiangnan Liang, and Chuang Shi. Comments on “Primaryambient extraction using ambient spectrum estimation for immersive spatial audio reproduction”. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):200– 202, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3328286. See

[HGT15a].

Li:2021:AFN

[LLT^{+}21a] Xintong Li, Lemao Liu, Zhaopeng Tu, Guanlin Li, Shuming Shi, and Max Q.-H. Meng. Attending from foresight: a novel attention mechanism for neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2606–2616, 2021. CODEN ???? ISSN 2329-9290.

Lian:2021:CCT

[LLT21b] Zheng Lian, Bin Liu, and Jianhua Tao. CTNet: Conversational transformer network for emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):985– 1000, 2021. CODEN ???? ISSN 2329-9290.

Lee:2019:MCS

[LLW^{+}19] C. Lee, H. Lee, S. Wu, C. Liu, W. Fang, J. Hsu, and B. Tseng. Machine comprehension of spoken content: TOEFL listening test and spoken SQuAD. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (9):1469–1480, September 2019. CODEN ???? ISSN 2329-9290.

Liu:2021:AEG

[LLW21] Rui Liu, Zheng Lin, and Weiping Wang. Addressing extraction and generation separately: Keyphrase prediction with pre-trained language models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3180–3191, 2021. CODEN ???? ISSN 2329-9290.

Li:2023:CTR

[LLW23] Yuying Li, Yuchen Liu, and Donald S. Williamson. A composite T60 regression and classification approach for speech dereverberation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1013– 1023, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3245423.

Leong:2021:ETM

[LLWC21] Chongman Leong, Xuebo Liu, Derek F. Wong, and Lidia S. Chao. Exploiting translation model for parallel corpus mining. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2829–2839, 2021. CODEN ???? ISSN 2329-9290.

Liu:2016:RCM

[LLWJ16] Yang Liu, Sujian Li, Furu Wei, and Heng Ji. Relation classification via modeling augmented dependency paths. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(9): 1589–1598, ???? 2016. CODEN ???? ISSN 2329-9290.

Li:2024:UCM

[LLWQ24] Jiahong Li, Chenda Li, Yifei Wu, and Yanmin Qian. Unified cross-modal attention: Robust audiovisual speech recognition and beyond. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1941– 1953, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3375641.

Liu:2021:TSS

[LLyL21] Andy T. Liu, Shang-Wen Li, and Hung yi Lee. TERA: Self-supervised learning of transformer encoder representation for speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2351– 2366, 2021. CODEN ???? ISSN 2329-9290.

Li:2021:THB

[LLZ^{+}21a] Andong Li, Wenzhe Liu, Chengshi Zheng, Cunhang Fan, and Xiaodong Li. Two heads are better than one: a twostage complex spectral mapping approach for monaural speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1829–1843, 2021. CODEN ???? ISSN 2329-9290.

Li:2021:DSC

[LLZ^{+}21b] Guanlin Li, Lemao Liu, Conghui Zhu, Rui Wang, Tiejun Zhao, and Shuming Shi. Detecting source contextual barriers for understanding neural machine translation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 3158–3169, 2021. CODEN ???? ISSN 2329-9290.

Li:2021:BTV

[LLZ^{+}21c] Zekang Li, Zongjia Li, Jinchao Zhang, Yang Feng, and Jie Zhou. Bridging text and video: a universal multimodal transformer for audiovisual scene-aware dialog. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2476–2483, 2021. CODEN ???? ISSN 2329-9290.

Li:2023:DAN

[LLZ^{+}23a] Tianyou Li, Siyuan Lian, Sipei Zhao, Jing Lu, and Ian S. Burnett. Distributed active noise control based on an augmented diffusion FxLMS algorithm. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1449– 1463, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3261742.

Liu:2023:AEB

[LLZ23b] Zhidong Liu, Junhui Li, and Muhua Zhu. Alleviating exposure bias for neural machine translation via contextual augmentation and self distillation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2079– 2089, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3277245.

Lu:2023:HSD

[LLZ^{+}23c] Junyu Lu, Hongfei Lin, Xiaokun Zhang, Zhaoqing Li, Tongyue Zhang, Linlin Zong, Fenglong Ma, and Bo Xu. Hate speech detection via dual contrastive learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2787– 2795, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3294715.

Lyu:2023:RHF

[LLZ^{+}23d] Xinglin Lyu, Junhui Li, Min Zhang, Chenchen Ding, Hideki Tanaka, and Masao Utiyama. Refining history for futureaware neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):500– 512, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3226332.

Li:2015:SIP

[LM15] Na Li and Man-Wai Mak. SNR-invariant PLDA modeling in nonparametric subspace for robust speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1648–1659, October 2015. CODEN ???? ISSN 2329-9290.

LeCornu:2017:GIA

[LM17] Thomas Le Cornu and Ben Milner. Generating intelligible audio speech from visual speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(9): 1751–1761, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7949073/.

Luo:2019:CTS

[LM19] Y. Luo and N. Mesgarani. Conv-TasNet: Surpassing ideal time frequency magnitude masking for speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(8):1256–1266, August 2019. CODEN ???? ISSN 2329-9290.

Lin:2022:MRL

[LM22] Weiwei Lin and Man-Wai Mak. Mixture representation learning for deep speaker embedding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):968–978, 2022. CO-DEN ???? ISSN 2329-9290.

Lin:2023:MAM

[LM23a] Weiwei Lin and Man-Wai Mak. Modelagnostic meta-learning for fast text-dependent speaker embedding adaptation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1866– 1876, 2023. ISSN 2329-

                                        9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2023.3275029.

Lin:2023:RSV

[LM23b] Weiwei Lin and Man-Wai Mak. Robust speaker verification using deep weight space ensemble. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 802–812, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3233231.

Li:2017:DDM

[LMC17] Na Li, Man-Wai Mak, and Jen-Tzung Chien. DNNdriven mixture of PLDA for robust speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(6):1371– 1383, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7894162/.

Lin:2018:MVD

[LMC18] W. Lin, M. Mak, and J. Chien. Multisource Ivectors domain adaptation using maximum mean discrepancy based autoencoders. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2412– 2422, December 2018. CODEN ???? ISSN 2329-

Levin:2021:NFS

[LMGG21] Dovid Y. Levin, Shmulik Markovich-Golan, and Sharon Gannot. Nearfield superdirectivity: an analytical perspective. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):1661–1674, 2021. CODEN ???? ISSN 2329-9290.

Li:2023:JML

[LMH23] Changheng Li, Jorge Martinez, and Richard Christian Hendriks. Joint maximum likelihood estimation of microphone array parameters for a reverberant single source scenario. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):695– 705, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3231706.

Liski:2021:AGD

[LMV21] Juho Liski, Aki Mäkivirta, and Vesa Välimäki. Audibility of group-delay equalization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2189– 2201, 2021. CODEN ???? ISSN 2329-9290.

Liang:2024:ELR

[LMW^{+}24] Xiaobo Liang, Runze Mao, Lijun Wu, Juntao Li, Min Zhang, and Qing Li. Enhancing lowresource NLP by consistency training with data and model perturbations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):189– 199, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3325970.

Lee:2022:DCV

[LNNL22] Sang-Hoon Lee, Hyeong-Rae Noh, Woo-Jeoung Nam, and Seong-Whan Lee. Duration controllable voice conversion via phoneme-based information bottleneck. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1173–1183, 2022. CODEN ???? ISSN 2329-9290.

Luo:2022:EAE

[LP22] Yu Luo and Lina Pu. EC-ANC: Edge case-enhanced active noise cancellation for true wireless stereo earbuds. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1436– 1447, 2022. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3164211.

Lacouture-Parodi:2014:MNR

[LPHCB14] Y. Lacouture-Parodi, E. A. P. Habets, Jingdong Chen, and J. Benesty. Multichannel noise reduction in the Karhunen–Loève expansion domain. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(5): 923–936, May 2014. CO-DEN ???? ISSN 2329-9290.

Liu:2021:LIA

[LPI^{+}21] Yuanyuan Liu, Nelly Penttilä, Tiina Ihalainen, Juulia Lintula, Rachel Convey, and Okko Räsänen. Language-independent approach for automatic computation of vowel articulation features in dysarthric speech assessment. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2228–2243, 2021. CODEN ???? ISSN 2329-9290.

Li:2021:SDC

[LPL^{+}21] Jiangnan Li, Hongliang Pan, Zheng Lin, Peng Fu, and Weiping Wang. Sarcasm detection with commonsense knowledge. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):3192–3201, 2021. CODEN ???? ISSN 2329-9290.

Li:2022:RLB

[LPL^{+}22] Qian Li, Hao Peng, Jianxin Li, Jia Wu, Yuanxing Ning, Lihong Wang, Philip S. Yu, and Zheng Wang. Reinforcement learning-based dialogue guided event extraction to exploit argument relations. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 520–533, 2022. CODEN ???? ISSN 2329-9290.

Li:2023:AFN

[LPM^{+}23] Haozhou Li, Qinke Peng, Xu Mou, Ying Wang, Zeyuan Zeng, and Muhammad Fiaz Bashir. Abstractive financial news summarization via transformer-BiLSTM encoder and graph attention-based decoder. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3190– 3205, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3304473.

Le:2024:IFT

[LQW^{+}24] Yuquan Le, Zhe Quan, Jiawei Wang, Da Cao, and Kenli Li. R^{2}: a novel recall & ranking framework for legal judgment prediction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1609– 1622, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3365389.

Lu:2017:SFH

[LR17] Liang Lu and Steve Renals. Small-footprint highway deep neural networks for speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(7):1502–1511, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7913606/.

Lee:2021:PAF

[LRG21] Ching-Hua Lee, Bhaskar D. Rao, and Harinath Garudadri. Proportionate adaptive filtering algorithms derived using an iterative reweighting framework. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):171–186, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3038526.

Liu:2023:AAP

[LRP^{+}23] Yuanyuan Liu, Mittapalle Kiran Reddy, Nelly Penttila, Tiina Ihalainen, Paavo Alku, and Okko Rasanen. Automatic assessment of Parkinson&#x0027;s disease using speech representations of phonation and articulation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):242– 255, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3212829.

Lemercier:2023:SDB

[LRWG23] Jean-Marie Lemercier, Julius Richter, Simon Welker, and Timo Gerkmann. StoRM: a diffusionbased stochastic regeneration model for speech enhancement and dereverberation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2724– 2737, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3294692.

Li:2014:SMA

[LS14a] Bo Li and Khe Chai Sim.

A spectral masking approach to noise-robust speech recognition using deep neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (8):1296–1305, August 2014. CODEN ???? ISSN 2329-9290.

Liu:2014:TVW

[LS14b] Shilin Liu and Khe Chai Sim. Temporally varying weight regression:

A semi-parametric trajectory model for automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):151–160, January 2014. CODEN ???? ISSN 2329-9290.

Liu:2021:EMP

[LSB^{+}21] Rui Liu, Berrak Sisman, Feilong Bao, Jichen Yang, Guanglai Gao, and Haizhou Li. Exploiting morphological and phonological features to improve prosodic phrasing for Mongolian speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):274–285, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3040523.

Li:2019:LGO

[LSDM19] Z. Li, Y. Song, L. Dai, and I. McLoughlin. Listening and grouping: an online autoregressive approach for monaural speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(4):692– 703, April 2019. CODEN ???? ISSN 2329-9290.

Li:2023:ESR

[LSG^{+}23a] Jijie Li, Kai Shuang, Jinyu Guo, Zengyi Shi, and Hongman Wang. Enhancing semantic relation classification with shortest dependency path reasoning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1550– 1560, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3265205.

Liu:2023:DFR

[LSG^{+}23b] Yang Liu, Haoqin Sun, Wenbo Guan, Yuqi Xia, Yongwei Li, Masashi Unoki, and Zhen Zhao. A discriminative feature representation method based on cascaded attention network with adversarial strategy for speech emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1063– 1074, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3245401.

Luo:2024:DGF

[LSGH24] Zhengding Luo, Dongyuan Shi, Woon-Seng Gan, and Qirui Huang. Delayless generative fixedfilter active noise control based on deep learning and Bayesian filter. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1048–1060, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3337632.

Liu:2021:ETT

[LSGL21] Rui Liu, Berrak Sisman, Guanglai Gao, and Haizhou Li. Expressive TTS training with frame and style reconstruction loss. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1806–1818, 2021. CODEN ???? ISSN 2329-9290.

Lun:2014:NEM

[LSH14] D. P. K. Lun, Tak-Wai Shen, and K. C. Ho. A novel expectationmaximization framework for speech enhancement in non-stationary noise environments. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): 335–346, February 2014. CODEN ???? ISSN 2329-9290.

Li:2023:MTI

[LSH^{+}23] Xingfeng Li, Xiaohan Shi, Desheng Hu, Yongwei Li, Qingchen Zhang, Zhengxia Wang, Masashi Unoki, and Masato Akagi. Music theory-inspired acoustic representation for speech emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2534– 2547, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3289312.

Li:2024:SSA

[LSL24] Xian Li, Nian Shao, and Xiaofei Li. Selfsupervised audio teacherstudent transformer for both clip-level and framelevel tasks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1336– 1351, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3352248.

Liu:2024:GSV

[LSLK24] Xuechen Liu, Md Sahidullah, Kong Aik Lee, and Tomi Kinnunen. Generalizing speaker verification for spoof awareness in the embedding space. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1261– 1273, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3358056.

Lee:2021:FGS

[LSNC21] Taewoong Lee, Liming Shi, Jesper Kjær Nielsen, and Mads Græsbøll Christensen. Fast generation of sound zones using variable span trade-off filters in the DFT-domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):363–378, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3042701.

Lu:2021:CGM

[LSTK21] Xugang Lu, Peng Shen, Yu Tsao, and Hisashi Kawai. Coupling a generative model with a discriminative learning framework for speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3631–3641, 2021. CODEN ???? ISSN 2329-9290.

Lehner:2018:OLI

[LSW18] B. Lehner, J. Schlüter, and G. Widmer. Online, loudness-invariant vocal detection in mixed music signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(8):1369– 1380, August 2018. CO-DEN ???? ISSN 2329-9290.

Li:2023:UST

[LSWN23] Yi Li, Yang Sun, Wenwu Wang, and Syed Mohsen Naqvi. U-shaped transformer with frequencyband aware attention for speech enhancement.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):1511–1521, 2023. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2023.3265839.

Lubis:2019:PEE

[LSYN19] N. Lubis, S. Sakti, K. Yoshino, and S. Nakamura. Positive emotion elicitation in chatbased dialogue systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (4):866–877, April 2019. CODEN ???? ISSN 2329-9290.

Li:2023:MSD

[LSZ23] Shaokai Li, Peng Song, and Wenming Zheng. Multi-source discriminant subspace alignment for cross-domain speech emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2448– 2460, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3288415.

Li:2021:NRG

[LTE^{+}21] Yongwei Li, Jianhua Tao, Donna Erickson, Bin Liu, and Masato Akagi. F_{0}-noise-robust glottal source and vocal tract analysis based on ARX-LF model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3375– 3383, 2021. CODEN ???? ISSN 2329-9290.

Lee:2017:PRN

[LTWT17] Hung-Yi Lee, Bo-Hsiang Tseng, Tsung-Hsien Wen, and Yu Tsao. Personalizing recurrent-neural-network-based language model by social network. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (3):519–530, ???? 2017. CODEN ???? ISSN 2329-9290.

Liebich:2022:OEC

[LV22] Stefan Liebich and Peter Vary. Occlusion effect cancellation in headphones and hearing devices the sister of active noise cancellation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):35–48, 2022. CODEN ???? ISSN 2329-9290.

Liu:2019:DCD

[LW19] Y. Liu and D. Wang. Divide and conquer: A deep CASA approach to talkerindependent monaural speaker separation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2092–2102, December 2019. CODEN ???? ISSN 2329-9290.

Li:2018:LKA

[LWC^{+}18] Q. Li, D. F. Wong, L. S. Chao, M. Zhu, T. Xiao, J. Zhu, and M. Zhang. Linguistic knowledge-aware neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2341–2354, December 2018. CODEN ???? ISSN 2329-9290.

Li:2021:MAC

[LWCC21] Canguang Li, Guohua Wang, Jin Cao, and Yi Cai. A multi-agent communication based model for nested named entity recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2123– 2136, 2021. CODEN ???? ISSN 2329-9290.

Liu:2019:LAB

[LWCL19] X. Liu, D. F. Wong, L. S.

Chao, and Y. Liu. Latent attribute based hierarchical decoder for neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2103–2112, December 2019. CODEN ???? ISSN 2329-9290.

Lan:2019:KBQ

[LWJ19] Y. Lan, S. Wang, and J. Jiang. Knowledge base question answering with a matchingaggregation model and question-specific contextual relations. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(10): 1629–1638, October 2019. CODEN ???? ISSN 2329-9290.

Li:2022:PSE

[LWK^{+}22] Lantian Li, Dong Wang, Jiawen Kang, Renyu Wang, Jing Wu, Zhendong Gao, and Xiao Chen. A principle solution for enroll-test mismatch in speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):443–455, 2022. CODEN ???? ISSN 2329-9290.

Li:2019:RCK

[LWL^{+}19] L. Li, J. Wang, J. Li, Q. Ma, and J. Wei. Relation classification via keyword-attentive sentence mechanism and synthetic stimulation loss. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (9):1392–1404, September 2019. CODEN ???? ISSN 2329-9290.

Li:2021:DMS

[LWL^{+}21] Songbin Li, Jingang Wang, Peng Liu, Miao Wei, and Qiandong Yan. Detection of multiple steganography methods in compressed speech based on code element embedding, Bi-LSTM and CNN with attention mechanisms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1556– 1569, 2021. CODEN ???? ISSN 2329-9290.

Liang:2022:MTD

[LWL^{+}22] Xiaobo Liang, Lijun Wu, Juntao Li, Tao Qin, Min Zhang, and Tie-Yan Liu. Multi-teacher distillation with single model for neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):992–1002, 2022. CO-DEN ???? ISSN 2329-9290.

Lim:2024:AMD

[LWLL24] Jungwoo Lim, Taesun Whang, Dongyub Lee, and Heuiseok Lim. Adaptive multi-domain dialogue state tracking on spoken conversations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32 (??):727–732, 2024. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3302232.

Li:2024:SCS

[LWLS24] Songbin Li, Jingang Wang, Peng Liu, and Ke Shi. SANet: a compressed speech encoder and steganography algorithm independent steganalysis deep neural network. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 680–690, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3337667.

Li:2020:MSE

[LWM^{+}20] Ruizhi Li, Xiaofei Wang, Sri Harish Mallidi, Shinji Watanabe, Takaaki Hori, and Hynek Hermansky. Multi-stream endto-end speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):646–655, January 2020. CODEN ???? ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2019.2959721.

Liu:2023:ATS

[LWS^{+}23] Xuechen Liu, Xin Wang, Md Sahidullah, Jose Patino, Héctor Delgado, Tomi Kinnunen, Massimiliano Todisco, Junichi Yamagishi, Nicholas Evans, Andreas Nautsch, and Kong Aik Lee. ASVspoof 2021: Towards spoofed and deepfake speech detection in the wild. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2507– 2522, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3285283.

Li:2022:CSE

[LWX^{+}22] Tao Li, Xinsheng Wang, Qicong Xie, Zhichao Wang, and Lei Xie. Crossspeaker emotion disentangling and transfer for end-to-end speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1448– 1460, 2022. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3164181.

Lin:2023:TRR

[LWY^{+}23] Yi Lin, Qingyang Wang, Xincheng Yu, Zichen Zhang, Dongyue Guo, and Jizhe Zhou. Towards recognition for radio-echo speech in air traffic control: Dataset and a contrastive learning approach. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3249– 3262, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3307219.

Liu:2023:CCB

[LWYC23] Ben Liu, Jun Wang, Guanyuan Yu, and Shaolei Chen. CUPVC: a constraint-based unsupervised prosody transfer for improving telephone banking services. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 2695–2706, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3293042.

Lin:2023:CCI

[LWYD23] Yuqin Lin, Longbiao Wang, Yanbing Yang, and Jianwu Dang. CFDRN: a cognition-inspired feature decomposition and recombination network for dysarthric speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3824– 3836, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3319276.

Li:2014:FMM

[LWZ^{+}14] Weifeng Li, Longbiao Wang, Yicong Zhou, J. Dines, M. Magimai-Doss, H. Bourlard, and Qingmin Liao. Feature mapping of multiple beamformed sources for robust overlapping speech recognition using a microphone array. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2244–2255, December 2014. CODEN ???? ISSN 2329-9290.

Li:2021:RNN

[LWZG21] Hao Li, DeLiang Wang, Xueliang Zhang, and Guanglai Gao. Recurrent neural networks and acoustic features for frame-level signal-to-noise ratio estimation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2878–2887, 2021. CODEN ???? ISSN 2329-9290.

Li:2016:ISU

[LWZZ16] Lantian Li, Dong Wang, Chenhao Zhang, and Thomas Fang Zheng. Improving short utterance speaker recognition by modeling speech unit classes. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(6):1129– 1139, June 2016. CODEN ???? ISSN 2329-9290.

Li:2021:MTS

[LXK^{+}21] Mei Li, Lu Xiang, Xiaomian Kang, Yang Zhao, Yu Zhou, and Chengqing Zong. Medical term and status generation from Chinese clinical dialogue with multi-granularity transformer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3362– 3374, 2021. CODEN ???? ISSN 2329-9290.

Liu:2018:COU

[LXS^{+}18] Bingquan Liu, Zhen Xu, Chengjie Sun, Baoxun Wang, Xiaolong Wang, Derek F. Wong, and Min Zhang. Content-oriented user modeling for personalized response ranking in chatbots. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(1): 122–133, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8068225/.

Li:2021:MMO

[LY21] Haoyu Li and Junichi Yamagishi. Multi-metric optimization using generative adversarial networks for near-end speech intelligibility enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3000–3011, 2021. CODEN ???? ISSN 2329-9290.

Luo:2023:MSS

[LY23] Yi Luo and Jianwei Yu. Music source separation with bandsplit RNN. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1893– 1901, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3271145.

Loweimi:2023:MSA

[LYB^{+}23] Erfan Loweimi, Zhengjun Yue, Peter Bell, Steve Renals, and Zoran Cvetkovic. Multi-stream acoustic modelling using raw real and imaginary parts of the Fourier transform. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 876–890, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3237167.

Liu:2022:CDS

[LYCX22] Jian Liu, Mengshi Yu, Yufeng Chen, and Jinan Xu. Cross-domain slot filling as machine reading comprehension: a new perspective. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 673–685, 2022. CODEN ???? ISSN 2329-9290.

Lin:2022:SNC

[LYH^{+}22] Yu-Chen Lin, Cheng Yu, Yi-Te Hsu, Szu-Wei Fu, Yu Tsao, and Tei-Wei Kuo. SEOFP-NET: Compression and acceleration of deep neural networks for speech enhancement using sign-exponent-only floating-points. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1016–1031, 2022. CODEN ???? ISSN 2329-9290.

Lin:2021:DSM

[LYL21] Peiqin Lin, Meng Yang, and Jianhuang Lai. Deep selective memory network with selective attention and inter-aspect modeling for aspect level sentiment classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1093– 1106, 2021. CODEN ???? ISSN 2329-9290.

Li:2014:IBD

[LYN14] Zhibao Li, K. F. C. Yiu, and S. Nordholm. On the indoor beamformer design with reverberation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):1225–1235, August 2014. CODEN ???? ISSN 2329-9290.

Lei:2022:MMS

[LYWX22] Yi Lei, Shan Yang, Xinsheng Wang, and Lei Xie. MsEmoTTS: Multi-scale emotion transfer, prediction, and control for emotional speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):853–864, 2022. CO-DEN ???? ISSN 2329-9290.

Lei:2023:MCK

[LYY^{+}23] Lei Lei, Guoshun Yuan, Hongjiang Yu, Dewei Kong, and Yuefeng He. Multilingual customized keyword spotting using similar-pair contrastive learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2437– 2447, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3284523.

Li:2024:TSA

[LYY24] Changtao Li, Feiran Yang, and Jun Yang. A twostage approach to quality restoration of boneconducted speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 818–829, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3337988.

Li:2021:LCL

[LYZ21] Yuling Li, Kui Yu, and Yuhong Zhang. Learning cross-lingual mappings in imperfectly isomorphic embedding spaces. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2630–2642, 2021. CODEN ???? ISSN 2329-9290.

Li:2023:GUS

[LYZ^{+}23] Andong Li, Guochen Yu, Chengshi Zheng, Wenzhe Liu, and Xiaodong Li. A general unfolding speech enhancement method motivated by Taylor’s theorem. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3629–3646, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3313442.

Li:2016:JAI

[LZ16] Peifeng Li and Guodong Zhou. Joint argument inference in Chinese event extraction with argument consistency and event relevance. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 612–622, April 2016. CO-DEN ???? ISSN 2329-9290.

Lu:2023:TUM

[LZ23] Jinliang Lu and Jiajun Zhang. Towards unified multi-domain machine translation with mixture of domain experts. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3488–3498, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3316451.

Lian:2023:UTA

[LZAY23] Jiachen Lian, Chunlei Zhang, Gopala K. Anumanchipalli, and Dong Yu. Unsupervised TTS acoustic modeling for TTS with conditional disentangled sequential VAE. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2548–2557, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3290423.

Li:2014:JOC

[LZC^{+}14] Zhenghua Li, Min Zhang, Wanxiang Che, Ting Liu, and Wenliang Chen. Joint optimization for Chinese POS tagging and dependency parsing. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): 274–286, January 2014. CODEN ???? ISSN 2329-9290.

Lei:2023:MMS

[LZC^{+}23] Shun Lei, Yixuan Zhou, Liyang Chen, Zhiyong Wu, Xixin Wu, Shiyin Kang, and Helen Meng. MSStyleTTS: Multi-scale style modeling with hierarchical context information for expressive speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3290– 3303, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3301217.

Liu:2023:UIK

[LZD^{+}23] Juhua Liu, Qihuang Zhong, Liang Ding, Hua Jin, Bo Du, and Dacheng Tao. Unified instance and knowledge alignment pretraining for aspectbased sentiment analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2629–2642, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3290431.

Lim:2014:RMD

[LZHN14] F. Lim, Wancheng Zhang, E. A. P. Habets, and P. A. Naylor. Robust multichannel dereverberation using relaxed multichannel least squares. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (9):1379–1390, September 2014. CODEN ???? ISSN 2329-9290.

Liu:2023:HEE

[LZL^{+}23] Yaxin Liu, Yan Zhou, Ziming Li, Junlin Wang, Wei Zhou, and Songlin Hu. HIM: an endto-end hierarchical interaction model for aspect sentiment triplet extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2272– 2285, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3282379.

Lin:2023:TOD

[LZX^{+}23] Haitao Lin, Junnan Zhu, Lu Xiang, Feifei Zhai, Yu Zhou, Jiajun Zhang, and Chengqing Zong. Topic-oriented dialogue summarization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1797–1810, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3271118.

Li:2022:HIJ

[LZZP22] Zuchao Li, Junru Zhou, Hai Zhao, and Kevin Parnow. HPSG-inspired joint neural constituent and dependency parsing in O(n^{3}) time complexity. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):355–366, 2022. CO-DEN ???? ISSN 2329-9290.

Markovic:2016:EAS

[MAB^{+}16] Dejan Marković, Fabio Antonacci, Lucio Bianchi, Stefano Tubaro, and Augusto Sarti. Extraction of acoustic sources through the processing of sound field maps in the ray space. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2481–2494, ???? 2016. CODEN ???? ISSN 2329-9290.

Moritz:2015:AIA

[MAK15] N. Moritz, J. Anemuller, and B. Kollmeier. An auditory inspired amplitude modulation filter bank for robust feature extraction in automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1926–1937, November 2015. CODEN ???? ISSN 2329-9290.

Mirbagheri:2018:RFA

[MAL18] Majid Mirbagheri, Les Atlas, and Adrian K. C. Lee. Regression factor analysis with an application to continuous HRIR measurement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(2): 415–421, ???? 2018.

CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8169044/.

Mantena:2014:QES

[MAP14] G. Mantena, S. Achanta, and K. Prahallad. Queryby-example spoken term detection using frequency domain linear prediction and non-segmental dynamic time warping. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (5):946–955, May 2014. CODEN ???? ISSN 2329-9290.

Malik:2023:PLD

[MASC23] Arsalan Malik, Nipun Agarwal, Harshavardhan Settibhaktini, and Ananthakrishna Chintanpalli. Predicting leveldependent changes in concurrent vowel scores using the 2D-CNN models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2558–2566, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3288411.

Markovic:2015:MSI

[MAST15] Dejan Marković, Fabio Antonacci, Augusto Sarti, and Stefano Tubaro. Multiview soundfield imaging in the projective ray space. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(6):1054–1067, June 2015. CODEN ???? ISSN 2329-9290.

Momeni:2015:JDE

[MAT15] H. Momeni, H. R. Abutalebi, and A. Tadaion. Joint detection and estimation of speech spectral amplitude using noncontinuous gain functions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (8):1249–1258, August 2015. CODEN ???? ISSN 2329-9290.

May:2018:RSD

[May18]  Tobias May.  Robust

speech dereverberation with a neural networkbased post-filter that exploits multi-conditional training of binaural cues. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (2):406–414, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8080254/.

Ma:2021:MCA

[MAZ21] Fei Ma, Thushara D. Abhayapala, and Wen Zhang. Multiple circular arrays of vector sensors for real-time sound field analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):286–299, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3040533.

Manamperi:2022:DAS

[MAZS22] Wageesha Manamperi, Thushara D. Abhayapala, Jihui Zhang, and Prasanga N. Samarasinghe. Drone audition: Sound source localization using on-board microphones. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):508– 519, 2022. CODEN ???? ISSN 2329-9290.

Morchid:2015:CMR

[MBD^{+}15] M. Morchid, M. Bouallegue, R. Dufour, G. Linares, D. Matrouf, and R. De Mori. Compact multiview representation of documents based on the total variability space. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(8): 1295–1308, August 2015. CODEN ???? ISSN 2329-9290.

Magron:2018:MBS

[MBD18] P. Magron, R. Badeau, and B. David. Modelbased STFT phase recovery for audio source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(6):1095– 1105, June 2018. CODEN ???? ISSN 2329-9290.

Miaschi:2023:RSN

[MBDV23] Alessio Miaschi, Dominique Brunato, Felice Dell’Orletta, and Giulia Venturi. On robustness and sensitivity of a neural language model: a case study on Italian L1 learner errors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 426–438, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3226333.

Mowlaee:2017:NRM

[MBK17] Pejman Mowlaee, Martin Blass, and W. Bastiaan Kleijn. New results in modulation-domain single-channel speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2125– 2137, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

                                        ieeexplore.ieee.org/
                                        document/8022920/.

Marelli:2015:EAH

[MBM15] D. Marelli, R. Baumgartner, and P. Majdak. Efficient approximation of head-related transfer functions in subbands for accurate sound localization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(7):1130–1143, July 2015. CODEN ???? ISSN 2329-9290.

Molina:2014:DRP

[MBTB14] E. Molina, A. M. Barbancho, L. J. Tardon, and I. Barbancho. Dissonance reduction in polyphonic audio using harmonic reorganization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): 325–334, February 2014. CODEN ???? ISSN 2329-9290.

Ming:2017:SEB

[MC17] Ji Ming and Danny Crookes. Speech enhancement based on fullsentence correlation and clean speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (3):531–543, ???? 2017. CODEN ???? ISSN 2329-9290.

Mignot:2014:LFI

[MCD14] R. Mignot, G. Chardon, and L. Daudet. Low frequency interpolation of room impulse responses using compressed sensing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):205–216, January 2014. CODEN ???? ISSN 2329-9290.

Menzies:2021:RCM

[MCF21] Dylan Menzies, Philip Coleman, and Filippo Maria Fazi. A room compensation method by modification of reverberant audio objects. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):239–252, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3036781.

McLoughlin:2014:SAV

[McL14] I. V. McLoughlin. Superaudible voice activity detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(9):1424– 1433, September 2014. CODEN ???? ISSN 2329-9290.

Mao:2022:ESB

[MCL22] Shuiyang Mao, P. C. Ching, and Tan Lee. Enhancing segment-based speech emotion recognition by iterative selflearning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):123– 134, 2022. CODEN ???? ISSN 2329-9290.

Miramont:2021:VJE

[MCS21] Juan M. Miramont, Marcelo A.

Colominas, and Gastón Schlotthauer. Voice jitter estimation using high-order synchrosqueezing operators. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):527–536, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3045265.

Medina:2021:IND

[MCZ21] C. Medina, R. Coelho, and L. Zão. Impulsive noise detection for speech enhancement in HHT domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2244–2253, 2021. CODEN ???? ISSN 2329-9290.

Mohammadiha:2016:SDU

[MD16] N. Mohammadiha and S. Doclo. Speech dereverberation using nonnegative convolutive transfer function and spectrotemporal modeling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(2): 276–289, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Marquardt:2018:ICP

[MD18] D. Marquardt and S. Doclo. Interaural coherence preservation for binaural noise reduction using partial noise estimation and spectral postfiltering. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (7):1261–1274, July 2018. CODEN ???? ISSN 2329-9290.

Mimilakis:2020:EMF

[MDCS20] Stylianos Ioannis Mimilakis, Konstantinos Drossos, Estefanía Cano, and Gerald Schuller. Examining the mapping functions of denoising autoencoders in singing voice separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):266–278, January 2020. CODEN ???? ISSN 2329-9290.

URL https://dl.acm.
org/doi/abs/10.1109/
TASLP.2019.2952013.

Mesaros:2019:SED

[MDE^{+}19] A. Mesaros, A. Diment, B. Elizalde, T. Heittola, E. Vincent, B. Raj, and T. Virtanen. Sound event detection in the DCASE 2017 challenge. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(6): 992–1006, June 2019. CO-DEN ???? ISSN 2329-9290.

Morfi:2015:SAS

[MDM15] V. Morfi, G. Degottex, and A. Mouchtaris. Speech analysis and synthesis with a computationally efficient adaptive harmonic model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1950–1962, November 2015. CODEN ???? ISSN 2329-9290.

Mesnil:2015:URN

[MDY^{+}15] G. Mesnil, Y. Dauphin, Kaisheng Yao, Y. Bengio, Li Deng, D. Hakkani-Tur, Xiaodong He, L. Heck, G. Tur, Dong Yu, and G. Zweig. Using recurrent neural networks for slot filling in spoken language understanding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (3):530–539, March 2015. CODEN ???? ISSN 2329-9290.

Menzies:2017:DCC

[MF17] Dylan Menzies and Filippo Maria Fazi. Decoding and compression of channel and scene objects for spatial audio. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(11): 2138–2151, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8016390/.

Menzies:2018:CPM

[MF18] D. Menzies and F. M. Fazi. A complex panning method for nearfield imaging. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1539– 1548, September 2018. CODEN ???? ISSN 2329-9290.

McVicar:2015:ACA

[MFG15] M. McVicar, S. Fukayama, and M. Goto. AutoGuitarTab: Computer-aided composition of rhythm and lead guitar parts in the tablature space. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (7):1105–1117, July 2015. CODEN ???? ISSN 2329-9290.

Maeda:2021:SFR

[MFH21] Natsuko Maeda, Filippo Maria Fazi, and Falk-Martin Hoffmann. Sound field reproduction with a cylindrical loudspeaker array using first order wall reflections. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 3617–3630, 2021. CODEN ???? ISSN 2329-9290.

R:2018:PPS

[MG18] A. R. MV and P. K.

Ghosh. PSFM a probabilistic source filter model for noise robust glottal closure instant detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (9):1645–1657, September 2018. CODEN ???? ISSN 2329-9290.

R:2019:GIF

[MG19] A. R. MV and P. K.

Ghosh. Glottal inverse filtering using probabilistic weighted linear prediction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(1):114–124, January 2019. CODEN ???? ISSN 2329-9290.

Ma:2018:RBL

[MGB18] N. Ma, J. A. Gonzalez, and G. J. Brown. Robust binaural localization of a target sound source by combining spectral source models and deep neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):2122–2131, November 2018. CODEN ???? ISSN 2329-9290.

Menzies:2018:LFP

[MGF18] Dylan Menzies, Marcos F. Simón Gálvez, and Filippo Maria Fazi. A low-frequency panning method with compensation for head rotation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (2):304–317, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8115309/.

Markovich-Golan:2017:CLT

[MGGK17] Shmulik Markovich-Golan, Sharon Gannot, and Walter Kellermann. Combined LCMV-TRINICON beamforming for separating multiple speech sources in noisy and reverberant environments. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (2):320–332, ???? 2017. CODEN ???? ISSN 2329-9290.

Mu:2015:OAM

[MGT15] Hao Mu, Woon-Seng Gan, and Ee-Leng Tan. An objective analysis method for perceptual quality of a virtual Bass system. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (5):840–850, May 2015. CODEN ???? ISSN 2329-9290.

Mirsamadi:2016:GNT

[MH16] Seyedmahdad Mirsamadi and John H. L. Hansen. A generalized nonnegative tensor factorization approach for distant speech recognition with distributed microphones. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1721– 1731, ???? 2016. CODEN ???? ISSN 2329-9290.

Misra:2018:MLL

[MH18] A. Misra and J. H. L.

Hansen. Maximumlikelihood linear transformation for unsupervised domain adaptation in speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (9):1549–1558, September 2018. CODEN ???? ISSN 2329-9290.

Mesaros:2018:DCA

[MHB^{+}18] Annamaria Mesaros, Toni Heittola, Emmanouil Benetos, Peter Foster, Mathieu Lagrange, Tuomas Virtanen, and Mark D. Plumbley. Detection and classification of acoustic scenes and events: Outcome of the DCASE 2016 Challenge. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(2):379–393, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8123864/.

Martinez-Hinarejos:2015:UDA

[MHBT15] C.-D. Martinez-Hinarejos, J.-M. Benedi, and V. Tamarit. Unsegmented dialogue act annotation and decoding with N-gram transducers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):198–211, January 2015. CODEN ???? ISSN 2329-9290.

Marquardt:2015:ICP

[MHD15] D. Marquardt, V. Hohmann, and S. Doclo. Interaural coherence preservation in multi-channel Wiener filtering-based noise reduction for binaural hearing aids. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2162– 2176, December 2015. CODEN ???? ISSN 2329-9290.

Meltzner:2017:SSR

[MHD^{+}17] Geoffrey S. Meltzner, James T. Heaton, Yunbin Deng, Gianluca De Luca, Serge H. Roy, and Joshua C. Kline. Silent speech recognition as an alternative communication device for persons with laryngectomy. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(12):2386–2398, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8114357/.

Marquardt:2015:TAL

[MHGD15] D. Marquardt, E. Hadad, S. Gannot, and S. Doclo. Theoretical analysis of linearly constrained multi-channel Wiener filtering algorithms for combined noise reduction and binaural cue preservation in binaural hearing aids. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2384–2397, December 2015. CODEN ???? ISSN 2329-9290.

Meyer:2015:DDC

[MHPB15] T. Meyer, N. Hajlaoui, and A. Popescu-Belis. Disambiguating discourse connectives for statistical machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (7):1184–1197, July 2015. CODEN ???? ISSN 2329-9290.

Man:2024:WMD

[MHZ^{+}24] Zhibo Man, Zengcheng Huang, Yujie Zhang, Yu Li, Yuanmeng Chen, Yufeng Chen, and Jinan Xu. WDSRL: Multidomain neural machine translation with wordlevel domain-sensitive representation learning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 577–590, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3336515.

Mim:2021:CAB

[MIR^{+}21] Farjana Sultana Mim, Naoya Inoue, Paul Reisert, Hiroki Ouchi, and Kentaro Inui. Corruption is not all bad: Incorporating discourse structure into pre-training via corruption for essay scoring. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2202–2215, 2021. CODEN ???? ISSN 2329-9290.

Maezawa:2014:NBD

[MIYO14] A. Maezawa, K. Itoyama, K. Yoshii, and H. G. Okuno. Nonparametric Bayesian dereverberation of power spectrograms based on infinite-order autoregressive processes. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1918–1930, December 2014. CODEN ???? ISSN 2329-9290.

Mahe:2018:PCR

[MJ18] G. Mahé and M. Jaïdane.

Perceptually controlled reshaping of sound histograms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1671– 1683, September 2018. CODEN ???? ISSN 2329-9290.

Mu:2014:MCD

[MJJ^{+}14] Yongsheng Mu, Peifeng Ji, Wei Ji, Ming Wu, and Jun Yang. Modeling and compensation for the distortion of parametric loudspeakers using a onedimension Volterra filter. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2169–2181, December 2014. CODEN ???? ISSN 2329-9290.

Mamun:2015:PSI

[MJZ15] N. Mamun, W. A. Jassim, and M. S. A. Zilany. Prediction of speech intelligibility using a neurogram orthogonal polynomial measure (NOPM). IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (4):760–773, April 2015. CODEN ???? ISSN 2329-9290.

Mansikkaniemi:2015:AMB

[MK15a] A. Mansikkaniemi and M. Kurimo. Adaptation of morph-based speech recognition for foreign names and acronyms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (5):941–950, May 2015. CODEN ???? ISSN 2329-9290.

Mowlaee:2015:HPE

[MK15b] P. Mowlaee and J. Kulmer. Harmonic phase estimation in single-channel speech enhancement using phase decomposition and SNR information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (9):1521–1532, September 2015. CODEN ???? ISSN 2329-9290.

Mowlaee:2015:PES

[MK15c] P. Mowlaee and J. Kulmer. Phase estimation in single-channel speech enhancement: Limitspotential. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):1283– 1294, August 2015. CO-DEN ???? ISSN 2329-9290.

Movassagh:2016:SAC

[MK16] Mahmood Movassagh and Peter Kabal. Scalable audio coding using Trellisbased optimized joint entropy coding and quantization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2288– 2300, ???? 2016. CODEN ???? ISSN 2329-9290.

Markovic:2014:EAR

[MKA^{+}14] D. Markovic, K. Kowalczyk, F. Antonacci, C. Hofmann, A. Sarti, and W. Kellermann. Estimation of acoustic reflection coefficients through pseudospectrum matching. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):125–137, January 2014. CODEN ???? ISSN 2329-9290.

Moritz:2016:IOM

[MKA16] Niko Moritz, Birger Kollmeier, and Jörn Anemüller. Integration of optimized modulation filter sets into deep neural networks for automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2439–2452, ???? 2016. CODEN ???? ISSN 2329-9290.

Muckenhirn:2017:LTS

[MKMDM17] Hannah Muckenhirn, Pavel Korshunov, Mathew Magimai-Doss, and Sébastien Marcel. Long-term spectral statistics for voice presentation attack detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2098–2111, ????

2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8015145/.

Mathad:2023:CVT

[MLC^{+}23] Vikram C. Mathad, Julie M. Liss, Kathy Chapman, Nancy Scherer, and Visar Berisha. Consonantvowel transition models based on deep learning for objective evaluation of articulation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 86–95, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3209937.

Mao:2022:APT

[MLL^{+}22] Qianren Mao, Jianxin Li, Chenghua Lin, Congwen Chen, Hao Peng, Lihong Wang, and Philip S. Yu. Adaptive pre-training and collaborative fine-tuning: a win-win strategy to improve review analysis tasks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):622–634, 2022. CODEN ???? ISSN 2329-9290.

Mariotte:2024:CCA

[MLMT24] Théo Mariotte, Anthony Larcher, Silvio Montrésor, and Jean-Hugh Thomas. Channel-combination algorithms for robust distant voice activity and overlapped speech detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1859– 1872, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3369531.

Ma:2024:CSD

[MLP^{+}24] Penghui Ma, Jianfeng Li, Jingjing Pan, Xiaofei Zhang, and Roberto Gil-Pita. Coherent signal DOA estimation with coprime array: Exploiting signal subspace reconstructing strategy. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 497–508, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3334482.

Ma:2017:EDN

[MMB17] Ning Ma, Tobias May, and Guy J. Brown. Exploiting deep neural networks and head movements for robust binaural localization of multiple sources in reverberant environments. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12): 2444–2453, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8086216/.

Malviya:2021:HHD

[MMBT21] Shrikant Malviya, Rohit Mishra, Santosh Kumar Barnwal, and Uma Shanker Tiwary. HDRS: Hindi dialogue restaurant search corpus for dialogue state tracking in task-oriented environment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2517– 2528, 2021. CODEN ???? ISSN 2329-9290.

Martin-Morato:2018:AMT

[MMCF18] I. Martín-Morató, M. Cobos, and F. J. Ferri. Adaptive mid-term representations for robust audio event classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2381–2392, December 2018. CODEN ???? ISSN 2329-9290.

Mertins:2021:RIR

[MMK21] Alfred Mertins, Marco Maass, and Fabrice Katzberg. Room impulse response reshaping and crosstalk cancellation using convex optimization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):489–502, January

                                        2021. URL https://dl.

acm.org/doi/10.1109/
TASLP.2020.3044444.

Martin-Morato:2023:SLS

[MMM23] Irene Martín-Morató and Annamaria Mesaros. Strong labeling of sound events using crowdsourced weak labels and annotator competence estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):902–914, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2022.3233468.

Mezghani-Marrakchi:2014:NAS

[MMMDL^{+}14] I. Mezghani-Marrakchi, G. Mahe, S. Djaziri-Larbi, M. Jaidane, and M. Turki-Hadj Alouane. Nonlinear audio systems identification through audio input Gaussianization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): 41–53, January 2014. CO-DEN ???? ISSN 2329-9290.

Mingote:2022:ALF

[MMR^{+}22] Victoria Mingote, Antonio Miguel, Dayana Ribas, Alfonso Ortega, and Eduardo Lleida. aDCF loss function for deep metric learning in endto-end text-dependent speaker verification systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):772–784, 2022. CODEN ???? ISSN 2329-9290.

Maeno:2020:SHD

[MMS^{+}20] Yu Maeno, Yuki Mitsufuji, Prasanga N. Samarasinghe, Naoki Murata, and Thushara D. Abhayapala. Sphericalharmonic-domain feedforward active noise control using sparse decomposition of reference signals from distributed sensor arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):656– 670, January 2020. CO-DEN ???? ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2019.2962684.

Makishima:2019:IDL

[MMT^{+}19] N. Makishima, S. Mogami, N. Takamune, D. Kitamura, H. Sumino, S. Takamichi, H. Saruwatari, and N. Ono. Independent deeply learned matrix analysis for determined audio source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(10):1601– 1615, October 2019. CO-DEN ???? ISSN 2329-9290.

Mohsenipour:2023:PCL

[MMZ23] Reza Mohsenipour, Daniel Massicotte, and Wei-Ping Zhu. PI control of loudspeakers based on linear fractional order model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1153– 1162, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3250826.

Moller:2019:ITF

[MNFGO19] M. B. Møller, J. K.

Nielsen, E. Fernandez-Grande, and S. K. Olesen. On the influence of transfer function noise on sound zone control in a room. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(9):1405–1418, September 2019. CODEN ???? ISSN 2329-9290.

Moller:2020:MHF

[MØ20] Martin Bo Møller and Jan Østergaard. A moving horizon framework for sound zones. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 256–265, January 2020. CODEN ???? ISSN 2329-9290. URL https:

//dl.acm.org/doi/abs/
10.1109/TASLP.2019.2951995.

Matsubara:2023:HNF

[MOT^{+}23] Keisuke Matsubara, Takuma Okamoto, Ryoichi Takashima, Tetsuya Takiguchi, Tomoki Toda, and Hisashi Kawai. Harmonic-net: Fundamental frequency and speech rate controllable fast neural vocoder. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1902–1915, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3275032.

Marxer:2016:UIO

[MP16] Ricard Marxer and Hendrik Purwins. Unsupervised incremental online learning and prediction of musical audio signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(5):863–874, May 2016. CODEN ???? ISSN 2329-9290.

Mathad:2020:VOP

[MP20] Vikram C. Mathad and S. R. Mahadeva Prasanna. Vowel onset point based screening of misarticulated stops in cleft lip and palate speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 450–460, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2957887.

Mai:2015:REN

[MPAEBLB15] Van-Khanh Mai, D. Pastor, A. Aissa-El-Bey, and R. Le-Bidan. Robust estimation of non-stationary noise power spectrum for speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (4):670–682, April 2015. CODEN ???? ISSN 2329-9290.

Mak:2016:MPN

[MPC16] Man-Wai Mak, Xiaomin Pang, and Jen-Tzung Chien. Mixture of PLDA for noise robust I-vector speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):130–142, January 2016. CODEN ???? ISSN 2329-9290.

Miotello:2024:DPB

[MPC^{+}24] Federico Miotello, Mirco Pezzoli, Luca Comanducci, Fabio Antonacci, and Augusto Sarti. Deep prior-based audio inpainting using multi-resolution harmonic convolutional neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 113–123, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3324556.

Marafioti:2019:CEA

[MPHM19] A. Marafioti, N. Perraudin, N. Holighaus, and P. Majdak. A context encoder for audio inpainting. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2362–2372, December 2019. CODEN ???? ISSN 2329-9290.

Meyer:2023:VHR

[MPKS^{+}23] Julie Meyer, Sebastian Prepeliţă, Ali Khajeh-Saeed, Michael Smirnov, and Pablo Hoffmann. Verification on head-related transfer functions of a snowman model simulated using the finitedifference time-domain method. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2579– 2591, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3291529.

Morgenstern:2020:PTO

[MR20] Hai Morgenstern and Boaz Rafaely. Perceptuallytransparent online estimation of two-channel room transfer function for sound calibration. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 332–342, January 2020. CODEN ???? ISSN 2329-9290. URL https:

//dl.acm.org/doi/abs/
10.1109/TASLP.2019.2955286.

Masuda:2023:ISS

[MS23] Naotake Masuda and Daisuke Saito. Improving semi-supervised differentiable synthesizer sound matching for practical applications. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):863– 875, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3237161.

McFee:2018:APO

[MSB18] B. McFee, J. Salamon, and J. P. Bello. Adaptive pooling operators for weakly labeled sound event detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(11): 2180–2193, November 2018. CODEN ???? ISSN 2329-9290.

Mohammadi:2014:ESA

[MSD14] A. Mohammadi, S. S. Sarfjoo, and C. Demiroglu. Eigenvoice speaker adaptation with minimal data for statistical speech synthesis systems using a MAP approach and nearest-neighbors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2146–2157, December 2014. CODEN ???? ISSN 2329-9290.

Muller:2021:LDV

[MSEC21] Matthias Müller, Thilo Schulz, Tatiana Ermakova, and Philipp P. Caffier. Lyric or dramatic — vibrato analysis for voice type classification in professional opera singers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):943–955, 2021. CO-DEN ???? ISSN 2329-9290.

Matsuyama:2015:AEO

[MSFK15] Y. Matsuyama, A. Saito, S. Fujie, and T. Kobayashi. Automatic expressive opinion sentence generation for enjoyable conversational systems. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): 313–326, February 2015. CODEN ???? ISSN 2329-9290.

McVicar:2014:ACE

[MSRND14] M. McVicar, R. Santos-Rodriguez, Yizhao Ni, and Tijl De Bie. Automatic chord estimation from audio: a review of the state of the art. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):556–575, February 2014. CODEN ???? ISSN 2329-9290.

Maestre:2017:JMB

[MSS17] Esteban Maestre, Gary P.

Scavone, and Julius O. Smith. Joint modeling of bridge admittance and body radiativity for efficient synthesis of string instrument sound by digital waveguides. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(5): 1128–1139, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7888993/.

Ma:2022:EDS

[MSW^{+}22] Bing Ma, Haifeng Sun, Jingyu Wang, Qi Qi, and Jianxin Liao. Extractive dialogue summarization without annotation based on distantly supervised machine reading comprehension in customer service. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):87–97, 2022. CO-DEN ???? ISSN 2329-9290.

Meynard:2018:SAN

[MT18] A. Meynard and B. Torrésani.

Spectral analysis for nonstationary audio. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2371–2380, December 2018. CODEN ???? ISSN 2329-9290.

Molina:2015:SST

[MTBB15] E. Molina, L. J. Tardon, A. M. Barbancho, and I. Barbancho. SiPTH: Singing transcription based on hysteresis defined on the pitch-time curve. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(2):252–263, February 2015. CODEN ???? ISSN 2329-9290.

Mogami:2020:ILR

[MTK^{+}20] Shinichi Mogami, Norihiro Takamune, Daichi Kitamura, Hiroshi Saruwatari, Yu Takahashi, Kazunobu Kondo, and Nobutaka Ono. Independent lowrank matrix analysis based on time-variant sub-Gaussian source model for determined blind source separation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 503–518, January 2020. CODEN ???? ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2019.2959257.

Mitsufuji:2021:MBS

[MTKS21] Yuki Mitsufuji, Norihiro Takamune, Shoichi Koyama, and Hiroshi Saruwatari. Multichannel blind source separation based on evanescentregion-aware non-negative tensor factorization in spherical harmonic domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):607–617, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3045528.

Madmoni:2021:EPT

[MTNR21] Lior Madmoni, Shir Tibor, Israel Nelken, and Boaz Rafaely. The effect of partial timefrequency masking of the direct sound on the perception of reverberant speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2037– 2047, 2021. CODEN ???? ISSN 2329-9290.

Moriya:2019:ESB

[MTS^{+}19] T. Moriya, T. Tanaka, T. Shinozaki, S. Watanabe, and K. Duh. Evolutionstrategy-based automation of system development for high-performance speech recognition. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1): 77–88, January 2019. CO-DEN ???? ISSN 2329-

Michelsanti:2021:ODL

[MTZ^{+}21] Daniel Michelsanti, Zheng-Hua Tan, Shi-Xiong Zhang, Yong Xu, Meng Yu, Dong Yu, and Jesper Jensen. An overview of deep-learning-based audio-visual speech enhancement and separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1368–1396, 2021. CODEN ???? ISSN 2329-9290.

Mary:2022:VTS

[MUK22] Narla John Metilda Sagaya Mary, Srinivasan Umesh, and Sandesh Varadaraju Katta. S-vectors and TESA: Speaker embeddings and a speaker authenticator based on transformer encoder. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 404–413, 2022. CODEN ???? ISSN 2329-9290.

Mitsufuji:2020:MNN

[MUT^{+}20] Yuki Mitsufuji, Stefan Uhlich, Norihiro Takamune, Daichi Kitamura, Shoichi Koyama, and Hiroshi Saruwatari. Multichannel non-negative matrix factorization using banded spatial covariance matrices in wavenumber domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):49–60, January 2020. CODEN ???? ISSN 2329-9290.

URL https://dl.acm.
org/doi/abs/10.1109/
TASLP.2019.2948770.

Masiero:2014:FCD

[MV14] B. Masiero and M. Vorlander. A framework for the calculation of dynamic crosstalk cancellation filters. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(9):1345–1354, September 2014. CODEN ???? ISSN 2329-9290.

Magron:2019:CIP

[MV19] P. Magron and T. Virtanen. Complex ISNMF: a phase-aware model for monaural audio source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(1):20–31, January 2019. CODEN ???? ISSN 2329-9290.

Moliner:2023:BGB

[MV23] Eloi Moliner and Vesa Välimäki. BEHM-GAN: Bandwidth extension of historical music using generative adversarial networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):943– 956, 2023. ISSN 2329-

9290. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2022.3190726.

Mehta:2016:RBV

[MVH16] Daryush D. Mehta, Jarrad H. Van Stan, and Robert E. Hillman. Relationships between vocal function measures derived from an acoustic microphone and a subglottal neck-surface accelerometer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(4):659–668, April 2016. CODEN ???? ISSN 2329-9290.

Mrinalini:2022:SSB

[MVN22] K. Mrinalini, P. Vijayalakshmi, and T. Nagarajan. SBSim: a Sentence-BERT similarity-based evaluation metric for Indian language neural machine translation systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1396– 1406, 2022. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3161160.

Miao:2023:SAU

[MWC^{+}23] Xiaoxiao Miao, Xin Wang, Erica Cooper, Junichi Yamagishi, and Natalia Tomashenko. Speaker anonymization using orthogonal Householder neural network. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3681– 3695, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313429.

Ma:2022:RNF

[MWL22] Tong Ma, Ying Wei, and Xin Lou. Reconfigurable nonuniform filter bank for hearing aid systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):758–771, 2022. CO-DEN ???? ISSN 2329-9290.

Ma:2016:SWM

[MWT16] Xi Ma, Dong Wang, and Javier Tejedor. Similar word model for unfrequent word enhancement in speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1819–1830, ???? 2016. CODEN ???? ISSN 2329-9290.

Mei:2023:IBP

[MWT^{+}23] Jie Mei, Yufan Wang, Xinhui Tu, Ming Dong, and Tingting He. Incorporating BERT with probability-aware gate for spoken language understanding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):826– 834, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3237156.

Ma:2017:NSO

[MX17] Yaping Ma and Yegui Xiao. A new strategy for online secondarypath modeling of narrowband active noise control. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (2):420–434, ???? 2017. CODEN ???? ISSN 2329-9290.

Mai:2021:AMS

[MXH21] Sijie Mai, Songlong Xing, and Haifeng Hu. Analyzing multimodal sentiment via acoustic- and visual-LSTM with channelaware temporal convolution network. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1424–1437, 2021. CODEN ???? ISSN 2329-9290.

Moore:2019:NCM

[MXNB19] A. H. Moore, W. Xue, P. A. Naylor, and M. Brookes. Noise covariance matrix estimation for rotating microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (3):519–530, March 2019. CODEN ???? ISSN 2329-9290.

Morales:2023:GAE

[MYC^{+}23] Arturo Morales, Juan I.

Yuz, Juan P. Cortés, Javier G. Fontanet, and Matías Zañartu. Glottal airflow estimation using neck surface acceleration and low-order Kalman smoothing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2055– 2066, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3277269.

Masuyama:2024:CRD

[MYK^{+}24] Yoshiki Masuyama, Kouei Yamaoka, Yuma Kinoshita, Taishi Nakashima, and Nobutaka Ono. Causal and relaxed-distortionless response beamforming for online target source extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):310– 324, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3329377.

Masuyama:2024:EJO

[MYKO24] Yoshiki Masuyama, Kouei Yamaoka, Takao Kawamura, and Nobutaka Ono. Efficient joint optimization of sampling rate offsets using entire multichannel signal. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1816– 1828, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3369532.

Ma:2021:DSA

[MYL^{+}21] Qianli Ma, Jiangyue Yan, Zhenxi Lin, Liuhong Yu, and Zipeng Chen. Deformable self-attention for text classification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1570–1581, 2021. CODEN ???? ISSN 2329-9290.

Masuyama:2023:OPR

[MYNO23] Yoshiki Masuyama, Kohei Yatabe, Kento Nagatomo, and Yasuhiro Oikawa. Online phase reconstruction via DNN-Based phase differences estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):163– 176, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3221041.

Ma:2019:GLM

[MYT^{+}19] Q. Ma, L. Yu, S. Tian, E. Chen, and W. W. Y. Ng. Global-local mutual attention model for text classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2127– 2139, December 2019. CODEN ???? ISSN 2329-9290.

Ma:2020:ACO

[MZA20] Fei Ma, Wen Zhang, and Thushara Dheemantha Abhayapala. Active control of outgoing broadband noise fields in rooms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):529– 539, January 2020. CO-DEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2960716.

Miao:2024:EVE

[MZC^{+}24] Chenfeng Miao, Qingying Zhu, Minchuan Chen, Jun Ma, Shaojun Wang, and Jing Xiao. EfficientTTS 2: Variational end-to-end text-to-speech synthesis and voice conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1650– 1661, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3369528.

Munir:2021:ACS

[MZL21a] Kashif Munir, Hai Zhao, and Zuchao Li. Adaptive convolution for semantic role labeling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 782–791, 2021. CODEN ???? ISSN 2329-9290.

Munir:2021:LCA

[MZL21b] Kashif Munir, Hai Zhao, and Zuchao Li. Learning context-aware convolutional filters for implicit discourse relation classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2421– 2433, 2021. CODEN ???? ISSN 2329-9290.

Miao:2015:SAT

[MZM15] Yajie Miao, Hao Zhang, and F. Metze. Speaker adaptive training of deep neural network acoustic models using I-vectors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1938–1949, November 2015. CODEN ???? ISSN 2329-9290.

McLoughlin:2015:RSE

[MZX^{+}15] I. McLoughlin, Haomin Zhang, Zhipeng Xie, Yan Song, and Wei Xiao. Robust sound event classification using deep neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(3): 540–552, March 2015. CODEN ???? ISSN 2329-9290.

Ma:2023:ESA

[MZZ23] Xinbei Ma, Zhuosheng Zhang, and Hai Zhao. Enhanced speaker-aware multi-party multi-turn dialogue comprehension. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):2410–2423, 2023. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2023.3284516.

Nose:2017:SSB

[NAK^{+}17] Takashi Nose, Yusuke Arao, Takao Kobayashi, Komei Sugiura, and Yoshinori Shiga. Sentence selection based on extended entropy using phonetic and prosodic contexts for statistical parametric speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):1107–1116, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7888524/.

Nielsen:2023:ATN

[NCB23] Jesper Kjær Nielsen, Mads Græsbøll Christensen, and Jesper Bünsow Boldt. An analysis of traditional noise power spectral density estimators based on the Gaussian stochastic volatility model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2299– 2313, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3282107.

Niedzwiecki:2015:EID

[NCC15] M. Niedzwiecki, M. Ciolek, and K. Cisowski. Elimination of impulsive disturbances from stereo audio recordings using vector autoregressive modeling and variable-order Kalman filtering. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(6): 970–981, June 2015. CO-DEN ???? ISSN 2329-9290.

Niu:2023:QSQ

[NDS^{+}23] Shu-Tong Niu, Jun Du, Lei Sun, Yu Hu, and Chin-Hui Lee. QDM-SSD: Quality-aware dynamic masking for separationbased speaker diarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1037–1049, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3244513.

Nikunen:2018:SMS

[NDV18] Joonas Nikunen, Aleksandr Diment, and Tuomas Virtanen. Separation of moving sound sources using multichannel NMF and acoustic tracking. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (2):281–295, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8114273/.

Neo:2021:ENR

[NEN21] Vincent W. Neo, Christine Evers, and Patrick A. Naylor. Enhancement of noisy reverberant speech using polynomial matrix eigenvalue decomposition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3255–3266, 2021. CODEN ???? ISSN 2329-9290.

Neo:2023:SCU

[NEWN23] Vincent W. Neo, Christine Evers, Stephan Weiss, and Patrick A. Naylor. Signal compaction using polynomial EVD for spherical array processing with applications. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3537–3549, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313441.

Niwa:2016:OMA

[NHK16] Kenta Niwa, Yusuke Hioka, and Kazunori Kobayashi. Optimal microphone array observation for clear recording of distant sound sources. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1785– 1795, ???? 2016. CODEN ???? ISSN 2329-9290.

Nakatani:2022:SIV

[NIK^{+}22] Tomohiro Nakatani, Rintaro Ikeshita, Keisuke Kinoshita, Hiroshi Sawada, Naoyuki Kamo, and Shoko Araki. Switching independent vector analysis and its extension to blind and spatially guided convolutional beamforming algorithms. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1032–1047, 2022. CODEN ???? ISSN 2329-9290.

Norholm:2016:ENS

[NJC16a] Sidsel Marie Nørholm, Jesper Rindom Jensen, and Mads Græsbøll Christensen. Enhancement and noise statistics estimation for non-stationary voiced speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 645–658, April 2016. CO-DEN ???? ISSN 2329-9290.

Norholm:2016:IFF

[NJC16b] Sidsel Marie Nørholm, Jesper Rindom Jensen, and Mads Græsbøll Christensen. Instantaneous fundamental frequency estimation with optimal segmentation for nonstationary voiced speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(12): 2354–2367, ???? 2016. CODEN ???? ISSN 2329-9290.

Nakamura:2021:HTF

[NK21] Tomohiko Nakamura and Hirokazu Kameoka. Harmonictemporal factor decomposition for unsupervised monaural separation of harmonic sounds. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 68–82, January 2021.

URL https://dl.acm.
org/doi/10.1109/TASLP.
2020.3037487.

Nigro:2022:MSA

[NK22] Michael Nigro and Sridhar Krishnan. Multimodal system for audio scene source counting and analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1073–1082, 2022.

CODEN ???? ISSN 2329-9290.

Nakamura:2021:TDA

[NKS21] Tomohiko Nakamura, Shihori Kozuka, and Hiroshi Saruwatari. Timedomain audio source separation with neural networks based on multiresolution analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1687–1701, 2021. CODEN ???? ISSN 2329-9290.

Nie:2018:DLB

[NLL^{+}18] S. Nie, S. Liang, W. Liu, X. Zhang, and J. Tao. Deep learning based speech separation via NMF-style reconstructions. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(11): 2043–2055, November 2018. CODEN ???? ISSN 2329-9290.

Nugraha:2016:MAS

[NLV16] Aditya Arie Nugraha, Antoine Liutkus, and Emmanuel Vincent. Multichannel audio source separation with deep neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(9): 1652–1664, ???? 2016. CODEN ???? ISSN 2329-9290.

Ni:2024:DRD

[NLYK24] Shiwen Ni, Jiawen Li, Min Yang, and Hung-Yu Kao. DropAttack: a random dropped weight attack adversarial training for natural language understanding. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 364–373, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3330613.

Nelus:2021:PPA

[NM21] Alexandru Nelus and Rainer Martin. Privacypreserving audio classification using variational information feature extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2864– 2877, 2021. CODEN ???? ISSN 2329-9290.

Ngo:2019:PAS

[NNC19] G. H. Ngo, M. Nguyen, and N. F. Chen. Phonologyaugmented statistical framework for machine transliteration using limited linguistic resources. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1): 199–211, January 2019. CODEN ???? ISSN 2329-

Nguyen:2020:HCE

[NNC20] Minh Nguyen, Gia H. Ngo, and Nancy F. Chen. Hierarchical character embeddings: Learning phonological and semantic representations in languages of logographic origin using recursive neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 461–473, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2955246.

Nguyen:2021:DSC

[NNC21] Minh Nguyen, Gia H. Ngo, and Nancy F. Chen. Domain-shift conditioning using adaptable filtering via hierarchical embeddings for robust Chinese spell check. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2027–2036, 2021. CODEN ???? ISSN 2329-9290.

Nielsen:2015:PBP

[NNL15] J. B. B. Nielsen, J. Nielsen, and J. Larsen. Perceptionbased personalization of hearing aids using Gaussian processes and active learning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(1):

162–173, January 2015. CODEN ???? ISSN 2329-9290.

Nakamura:2016:RTA

[NNS16] T. Nakamura, E. Nakamura, and S. Sagayama. Real-time audio-to-score alignment of music performances containing errors and arbitrary repeats and skips. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(2):329–339, ???? 2016. CODEN ???? ISSN 2329-9290.

Nongpiur:2014:DMB

[Non14] R. C. Nongpiur. Design of minimax broadband beamformers that are robust to microphone gain, phase, and position errors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (6):1013–1022, June 2014. CODEN ???? ISSN 2329-9290.

Nose:2016:EIG

[Nos16] Takashi Nose. Efficient implementation of global variance compensation for parametric speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1694–1704, ???? 2016. CODEN ???? ISSN 2329-9290.

Nadiri:2014:LMS

[NR14a] O. Nadiri and B. Rafaely. Localization of multiple speakers under high reverberation using a spherical microphone array and the direct-path dominance test. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(10):1494–1505, October 2014. CODEN ???? ISSN 2329-9290.

Nanjundaswamy:2014:CLT

[NR14b] T. Nanjundaswamy and K. Rose. Cascaded long term prediction for enhanced compression of polyphonic audio signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (3):697–710, March 2014. CODEN ???? ISSN 2329-9290.

Ni:2022:FUN

[NR22] Xin Ni and Jia Ren. FC-U2-Net: a novel deep neural network for singing voice separation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 489–494, 2022. CODEN ???? ISSN 2329-9290.

Novotny:2014:AEA

[NRČR14] Michal Novotný, Jan Rusz, Roman Čmejla, and Evžen Růžička. Automatic evaluation of articulatory disorders in Parkinson’s disease. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(9):1366– 1378, September 2014. CODEN ???? ISSN 2329-9290.

Narendra:2021:DPD

[NSA21] N. P. Narendra, Björn Schuller, and Paavo Alku. The detection of Parkinson’s disease from speech using voice source information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1925– 1936, 2021. CODEN ???? ISSN 2329-9290.

Nishimura:2020:MSN

[NSNN20] Yuta Nishimura, Katsuhito Sudoh, Graham Neubig, and Satoshi Nakamura. Multi-source neural machine translation with missing data. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28 (??):569–580, January 2020. CODEN ???? ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2019.2959224.

Nakashika:2015:VCU

[NTA15] T. Nakashika, T. Takiguchi, and Y. Ariki. Voice conversion using RNN pretrained by recurrent temporal restricted Boltzmann machines. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(3): 580–587, March 2015. CODEN ???? ISSN 2329-9290.

Nakashika:2016:NPT

[NTM16] Toru Nakashika, Tetsuya Takiguchi, and Yasuhiro Minami. Non-parallel training in voice conversion using an adaptive restricted Boltzmann machine. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):2032–2045, ???? 2016. CODEN ???? ISSN 2329-9290.

Niizumi:2023:BAE

[NTO^{+}23] Daisuke Niizumi, Daiki Takeuchi, Yasunori Ohishi, Noboru Harada, and Kunio Kashino. BYOL for audio: Exploring pretrained general-purpose audio representations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):137–151, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2022.3221007.

Nakashika:2019:CVR

[NTY19] T. Nakashika, S. Takaki, and J. Yamagishi. Complexvalued restricted Boltzmann machine for speakerdependent speech parameterization from complex spectra. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(2): 244–254, February 2019. CODEN ???? ISSN 2329-9290.

Nikunen:2014:DAB

[NV14] J. Nikunen and T. Virtanen. Direction of arrival based spatial covariance model for blind sound source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (3):727–739, March 2014. CODEN ???? ISSN 2329-9290.

Niermann:2021:LEN

[NV21] Markus Niermann and Peter Vary. Listening enhancement in noisy environments: Solutions in time and frequency domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):699–709, 2021. CODEN ???? ISSN 2329-9290.

Narayanan:2014:ISS

[NW14] A. Narayanan and Deliang Wang. Investigation of speech separation as a front-end for noise robust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (4):826–835, April 2014. CODEN ???? ISSN 2329-9290.

Narayanan:2015:IRD

[NW15] A. Narayanan and Deliang Wang. Improving robustness of deep neural network acoustic models via speech separation and joint adaptive training. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (1):92–101, January 2015. CODEN ???? ISSN 2329-9290.

Nayem:2024:ABS

[NW24] Khandokar Md. Nayem and Donald S. Williamson. Attention-based speech enhancement using human quality perception modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):250– 260, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3328282.

Nagathil:2016:SCR

[NWM16] Anil Nagathil, Claus Weihs, and Rainer Martin. Spectral complexity reduction of music signals for mitigating effects of cochlear hearing loss. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (3):445–458, March 2016.

CODEN ???? ISSN 2329-9290.

Nguyen:2016:FAU

[NXCL16] Duc Hoang Ha Nguyen, Xiong Xiao, Eng Siong Chng, and Haizhou Li. Feature adaptation using linear spectro-temporal transform for robust speech recognition. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(6): 1006–1019, June 2016. CODEN ???? ISSN 2329-9290.

Natgunanathan:2017:PBM

[NXH^{+}17] Iynkaran Natgunanathan, Yong Xiang, Guang Hua, Gleb Beliakov, and John Yearwood. Patchworkbased multilayer audio watermarking. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(11): 2176–2187, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8025572/.

Nakashika:2021:GBM

[NY21] Toru Nakashika and Kohei Yatabe. Gamma Boltzmann machine for audio modeling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2591–2605, 2021. CODEN ???? ISSN 2329-9290.

Nakamura:2017:NVR

[NYD17] Eita Nakamura, Kazuyoshi Yoshii, and Simon Dixon. Note value recognition for piano transcription using Markov random fields. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (9):1846–1858, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7964741/.

Nakamura:2017:RTP

[NYS17] Eita Nakamura, Kazuyoshi Yoshii, and Shigeki Sagayama. Rhythm transcription of polyphonic piano music based on merged-output HMM for multiple voices. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):794–806, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7839215/.

Ni:2023:SPA

[NZL23] Jingen Ni, Ningning Zhang, and Haofen Li. Sparsity-promoting affine projection algorithm with periodically-updated gain matrix and its performance analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1993–2003, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3277289.

Olivieri:2024:AIC

[OBP^{+}24] Marco Olivieri, Amy Bastine, Mirco Pezzoli, Fabio Antonacci, Thushara Abhayapala, and Augusto Sarti. Acoustic imaging with circular microphone array: a new approach for sound field analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1750– 1761, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3369533.

Oglic:2021:LWB

[OCS21] Dino Oglic, Zoran Cvetkovic, and Peter Sollich. Learning waveform-based acoustic models using deep variational convolutional neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2850–2863, 2021. CODEN ???? ISSN 2329-9290.

Ottosen:2017:PVB

[OD17] Emil Solsbæk Ottosen and Monika Dörfler. A phase vocoder based on nonstationary Gabor frames. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2199– 2208, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/8031036/.

Ouali:2016:FAF

[ODG16] Chahid Ouali, Pierre Dumouchel, and Vishwa Gupta. Fast audio fingerprinting system using GPU and a clusteringbased technique. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(6): 1106–1118, June 2016. CODEN ???? ISSN 2329-9290.

Ochiai:2023:MBN

[ODNA23] Tsubasa Ochiai, Marc Delcroix, Tomohiro Nakatani, and Shoko Araki. Maskbased neural beamforming for moving speakers with self-attention-based tracking. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 835–848, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3237172.

Ouchi:2016:TBD

[ODSM16] Hiroki Ouchi, Kevin Duh, Hiroyuki Shindo, and Yuji Matsumoto. Transition-based dependency parsing exploiting supertags. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):2059– 2068, ???? 2016. CODEN ???? ISSN 2329-9290.

Olivieri:2017:GPS

[OFF^{+}17] Ferdinando Olivieri, Filippo Maria Fazi, Simone Fontana, Dylan Menzies, and Philip Arthur Nelson. Generation of private sound with a circular loudspeaker array and the weighted pressure matching method. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):1579–1591, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7918631/.

Osterwise:2014:DFD

[OG14] C. Osterwise and S. L.

Grant. On over-determined frequency domain BSS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (5):956–966, May 2014. CODEN ???? ISSN 2329-9290.

Ouyang:2018:SPP

[OGZ18] X. Ouyang, K. Gu, and P. Zhou. Spatial pyramid pooling mechanism in 3D convolutional network for sentence-level classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):2167– 2179, November 2018. CODEN ???? ISSN 2329-9290.

Ogawa:2016:ESR

[OHN16] Atsunori Ogawa, Takaaki Hori, and Atsushi Nakamura. Estimating speech recognition accuracy based on error type classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2400–2413, ???? 2016. CODEN ???? ISSN 2329-9290.

Otsuka:2014:BNM

[OISO14] T. Otsuka, K. Ishiguro, H. Sawada, and H. G. Okuno. Bayesian nonparametrics for microphone array processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):493–504, February 2014. CODEN ???? ISSN 2329-9290.

Otsuka:2014:MSS

[OIY^{+}14] T. Otsuka, K. Ishiguro, T. Yoshioka, H. Sawada, and H. G. Okuno. Multichannel sound source dereverberation and separation for arbitrary number of sources based on Bayesian nonparametrics. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2218–2232, December 2014. CODEN ???? ISSN 2329-9290.

Obin:2015:SMP

[OL15] N. Obin and P. Lanchantin. Symbolic modeling of prosody: From linguistics to statistics. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (3):588–599, March 2015. CODEN ???? ISSN 2329-9290.

Ozer:2024:SSP

[ÖM24] Yigitcan Özer and Meinard Müller. Source separation of piano concertos using musically motivated augmentation techniques. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1214– 1225, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3356980.

OHanlon:2016:NNG

[ONKP16] Ken O’Hanlon, Hidehisa Nagano, Nicolas Keriven, and Mark D. Plumbley. Non-negative group sparsity with subspace note modelling for polyphonic transcription. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): 530–542, March 2016.

CODEN ???? ISSN 2329-9290.

Omachi:2017:AMM

[OOK17] Motoi Omachi, Tetsuji Ogawa, and Tetsunori Kobayashi. Associative memory model-based linear filtering and its application to tandem connectionist blind source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(3):637– 650, ???? 2017. CODEN ???? ISSN 2329-9290.

Obin:2016:SSA

[OR16a] Nicolas Obin and Axel Roebel. Similarity search of acted voices for automatic voice casting. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (9):1642–1651, ???? 2016. CODEN ???? ISSN 2329-9290.

OLeary:2016:MAS

[OR16b] Seán O’Leary and Axel Röbel. A montage approach to sound texture synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(6):1094– 1105, June 2016. CODEN ???? ISSN 2329-9290.

OHanlon:2021:FSC

[OS21] Ken O Hanlon and Mark B. Sandler. Fifth-Net: Structured compact neural networks for automatic chord recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2671–2682, 2021. CODEN ???? ISSN 2329-9290.

OShaughnessy:2024:RMA

[O’S24] Douglas O’Shaughnessy.

Review of methods for automatic speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1776– 1789, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3346293.

Ong:2021:BSM

[OVN21] Jonah Ong, Ba Tuong Vo, and Sven Nordholm. Blind separation for multiple moving sources with labeled random finite sets. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2137–2151, 2021. CODEN ???? ISSN 2329-9290.

Ong:2022:AVB

[OVN^{+}22] Jonah Ong, Ba Tuong Vo, Sven Nordholm, Ba-Ngu Vo, Diluka Moratuwage, and Changbeom Shim. Audio-visual based online multi-source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1219–1234, 2022. CODEN ???? ISSN 2329-9290.

Okopal:2015:SAS

[OWA15] G. Okopal, S. Wisdom, and L. Atlas. Speech analysis with the strong uncorrelating transform. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1858–1868, November 2015. CODEN ???? ISSN 2329-9290.

Ozturk:2023:SRS

[OWW^{+}23] Muhammed Zahid Ozturk, Chenshu Wu, Beibei Wang, Min Wu, and K. J. Ray Liu. RadioSES: mmWave-based audioradio speech enhancement and separation system. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1333– 1347, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3250846.

Porschmann:2019:DES

[PAB19] C. Pörschmann, J. M. Arend, and F. Brinkmann. Directional equalization of sparse head-related transfer function sets for spatial upsampling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(6): 1060–1071, June 2019. CODEN ???? ISSN 2329-9290.

Palomo-Alonso:2023:FAU

[PACPJF^{+}23] Alberto Palomo-Alonso, David Casillas-Pérez, Silvia Jiménez-Fernández, Jose A. Portilla-Figueras, and Sancho Salcedo-Sanz. A flexible architecture using temporal, spatial and semantic correlationbased algorithms for story segmentation of broadcast news. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3055– 3069, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3301231.

Poirot:2023:PES

[PBA^{+}23] Samuel Poirot, Stefan Bilbao, Mitsuko Aramaki, Sølvi Ystad, and Richard Kronland-Martinet. A perceptually evaluated signal model: Collisions between a vibrating object and an obstacle. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2338– 2350, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3284515.

Pan:2016:DDP

[PBC16] Chao Pan, J. Benesty, and Jingdong Chen. Design of directivity patterns with a unique null of maximum multiplicity. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (2):226–235, ???? 2016. CODEN ???? ISSN 2329-9290.

Paleologu:2018:LSI

[PBC18] C. Paleologu, J. Benesty, and S. Ciochină. Linear system identification based on a Kronecker product decomposition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(10):1793–1808, October 2018. CODEN ???? ISSN 2329-9290.

Poignant:2015:USI

[PBQ15] J. Poignant, L. Besacier, and G. Quénot. Unsupervised speaker identification in TV broadcast based on written names. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (1):57–68, January 2015. CODEN ???? ISSN 2329-9290.

Prusa:2017:NMR

[PBS17] Zdeněk Průša, Peter Balazs, and Peter Lempel Søndergaard. A noniterative method for reconstruction of phase from STFT magnitude. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(5): 1154–1164, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7890450/.

Park:2019:SSM

[PC19] J. Park and J. Chang.

State-space microphone array nonlinear acoustic echo cancellation using multi-microphone nearend speech covariance. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (10):1520–1534, October 2019. CODEN ???? ISSN 2329-9290.

Park:2021:IEL

[PC21] Sooyeon Park and Jung-Woo Choi. Iterative echo labeling algorithm with convex hull expansion for room geometry estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1463–1478, 2021. CODEN ???? ISSN 2329-9290.

Peretz:2024:CEB

[PC24] Orel Peretz and Israel Cohen. Constant elevationbeamwidth beamforming with concentric ring arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1662– 1672, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3365390.

Pan:2014:PSM

[PCB14] Chao Pan, Jingdong Chen, and J. Benesty. Performance study of the MVDR beamformer as a function of the source incidence angle. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): 67–79, January 2014. CO-DEN ???? ISSN 2329-9290.

Pan:2015:TAD

[PCB15] Chao Pan, Jingdong Chen, and J. Benesty. Theoretical analysis of differential microphone array beamforming and an improved solution. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):2093–2105, November 2015. CODEN ???? ISSN 2329-9290.

Pan:2016:ROR

[PCB16a] Chao Pan, Jingdong Chen, and Jacob Benesty. Reduced-order robust superdirective beamforming with uniform linear microphone arrays.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (9):1548–1559, ???? 2016. CODEN ???? ISSN 2329-9290.

Parthasarathy:2016:UAD

[PCB16b] Srinivas Parthasarathy, Roddy Cowie, and Carlos Busso. Using agreement on direction of change to build rank-based emotion classifiers. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(11): 2108–2121, ???? 2016. CODEN ???? ISSN 2329-9290.

Pan:2019:DTB

[PCBS19] C. Pan, J. Chen, J. Benesty, and G. Shi. On the design of target beampatterns for differential microphone arrays. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(8): 1295–1307, August 2019. CODEN ???? ISSN 2329-9290.

Patel:2017:MPS

[PCG17] Vinal Patel, Jordan Cheer, and Nithin V. George. Modified phase-scheduled-command FxLMS algorithm for active sound profiling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(9):1799– 1808, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

                                        ieeexplore.ieee.org/
                                        document/7953648/.

Primavera:2014:OSI

[PCLP14] A. Primavera, S. Cecchi, Junfeng Li, and F. Piazza. Objective and subjective investigation on a novel method for digital reverberator parameters estimation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): 441–452, February 2014. CODEN ???? ISSN 2329-9290.

Perez-Carrillo:2015:IAV

[PCW15] A. Perez-Carrillo and M. M. Wanderley. Indirect acquisition of violin instrumental controls from audio signal with hidden Markov models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (5):932–940, May 2015. CODEN ???? ISSN 2329-9290.

Purushothaman:2024:SDF

[PDKG24] Anurenjan Purushothaman, Debottam Dutta, Rohit Kumar, and Sriram Ganapathy. Speech dereverberation with frequency domain autoregressive modeling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 29–38, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3317570.

Palangi:2016:DSE

[PDS^{+}16] Hamid Palangi, Li Deng, Yelong Shen, Jianfeng Gao, Xiaodong He, Jianshu Chen, Xinying Song, and Rabab Ward. Deep sentence embedding using long short-term memory networks: Analysis and application to information retrieval. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 694–707, April 2016. CO-DEN ???? ISSN 2329-9290.

Parekh:2020:WSR

[PEO^{+}20] Sanjeel Parekh, Slim Essid, Alexey Ozerov, Ngoc Q. K. Duong, Patrick Pérez, and Gaël Richard. Weakly supervised representation learning for audio-visual scene analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):416–428, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2957889.

Pei:2024:GFM

[PFLX24] Hao-Chen Pei, Hao Fang, Xin Luo, and Xin-Shun Xu. Gradformer: a framework for multiaspect multi-granularity pronunciation assessment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32 (??):554–563, 2024. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3335807.

Pepe:2022:DOP

[PGS^{+}22] Giovanni Pepe, Leonardo Gabrielli, Stefano Squartini, Carlo Tripodi, and Nicolò Strozzi. Deep optimization of parametric IIR filters for audio equalization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1136– 1149, 2022. CODEN ???? ISSN 2329-9290.

Pessentheiner:2016:LCM

[PHK16] Hannes Pessentheiner, Martin Hagmüller, and Gernot Kubin. Localization and characterization of multiple harmonic sources. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(8): 1348–1363, ???? 2016. CODEN ???? ISSN 2329-9290.

Phan:2016:LRN

[PHM^{+}16] Huy Phan, Lars Hertel, Marco Maass, Radoslaw Mazur, and Alfred Mertins. Learning representations for nonspeech audio events through their similarities to speech patterns. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(4):807–822, April 2016. CODEN ???? ISSN 2329-9290.

Phan:2017:IAS

[PHM^{+}17] Huy Phan, Lars Hertel, Marco Maass, Philipp Koch, Radoslaw Mazur, and Alfred Mertins. Improved audio scene classification based on labeltree embeddings and convolutional neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(6):1278– 1290, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7933052/.

Perraudin:2018:ILA

[PHMB18] N. Perraudin, N. Holighaus, P. Majdak, and P. Balazs. Inpainting of long audio segments with similarity graphs. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (6):1083–1094, June 2018. CODEN ???? ISSN 2329-9290.

Prabhavalkar:2024:EES

[PHS^{+}24] Rohit Prabhavalkar, Takaaki Hori, Tara N. Sainath, Ralf Schlüter, and Shinji Watanabe. End-to-end speech recognition: a survey. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):325– 351, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3328283.

Pedersen:2024:DDN

[PJTJ24] Mathias Bach Pedersen, Søren Holdt Jensen, Zheng-Hua Tan, and Jesper Jensen. Data-driven non-intrusive speech intelligibility prediction using speech presence probability. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):55– 67, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3321964.

Petkov:2015:SDR

[PK15] P. N. Petkov and W. B. Kleijn. Spectral dynamics recovery for enhanced speech intelligibility in noise. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(2):327–338, February 2015. CODEN ???? ISSN 2329-9290.

Panagakis:2014:MGC

[PKA14] Y. Panagakis, C. L.

Kotropoulos, and G. R. Arce. Music genre classification via joint sparse low-rank representation of audio features. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1905–1917, December 2014. CODEN ???? ISSN 2329-9290.

Pal:2021:MLL

[PKP^{+}21] Monisankha Pal, Manoj Kumar, Raghuveer Peri, Tae Jin Park, So Hyun Kim, Catherine Lord, Somer Bishop, and Shrikanth Narayanan. Meta-learning with latent space clustering in generative adversarial network for speaker diarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1204– 1219, 2021. CODEN ???? ISSN 2329-9290.

Paraskevopoulos:2024:SEU

[PKR^{+}24] Georgios Paraskevopoulos, Theodoros Kouzelis, Georgios Rouvalis, Athanasios Katsamanis, Vassilis Katsouros, and Alexandros Potamianos. Sampleefficient unsupervised domain adaptation of speech recognition systems: a case study for modern Greek. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):286– 299, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3328280.

Peng:2023:TSB

[PLQZ23] Wanli Peng, Sheng Li, Zhenxing Qian, and Xinpeng Zhang. Text steganalysis based on hierarchical supervised learning and dual attention mechanism. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3513– 3526, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3319975.

Proenca:2018:MDC

[PLT^{+}18] J. Proença, C. Lopes, M. Tjalve, A. Stolcke, S. Candeias, and F. Perdigão. Mispronunciation detection in children’s reading of sentences. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(7):1207– 1219, July 2018. CODEN ???? ISSN 2329-9290.

Pang:2017:BSL

[PLZL17] Cheng Pang, Hong Liu, Jie Zhang, and Xiaofei Li.

Binaural sound localization based on reverberation weighting and generalized parametric mapping. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(8):1618–1632, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7926345/.

Percybrooks:2015:NHB

[PM15] W. S. Percybrooks and E. Moore. A new HMMbased voice conversion methodology evaluated on monolingual and crosslingual conversion tasks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2298–2310, December 2015. CODEN ???? ISSN 2329-9290.

Prakash:2019:AIP

[PM19] J. J. Prakash and H. A.

Murthy. Analysis of interpausal units in Indian languages and its application to text-to-speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(10):1616–1628, October 2019. CODEN ???? ISSN 2329-9290.

Prakash:2023:ERL

[PM23] Anusha Prakash and Hema A. Murthy. Exploring the role of language families for building indic speech synthesisers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):734– 747, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3230453.

Politis:2021:OES

[PMA^{+}21] Archontis Politis, Annamaria Mesaros, Sharath Adavanne, Toni Heittola, and Tuomas Virtanen. Overview and evaluation of sound event localization and detection in DCASE 2019. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):684–698, 2021. CODEN ???? ISSN 2329-9290.

Payal:2019:ENP

[PMB^{+}19] S. S. Payal, V. J. Mathews, D. J. Button, A. Iyer, R. H. Lambert, J. Hutchings, and L. A. Azpicueta-Ruiz. Equalization of nonlinear propagation distortion in cylindrical waveguides. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(6):1072–1084, June 2019. CODEN ???? ISSN 2329-9290.

Parcollet:2020:RHS

[PMB^{+}20] Titouan Parcollet, Mohamed Morchid, Xavier Bost, Georges Linarès, and Renato De Mori. Real to H-space autoencoders for theme identification in telephone conversations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):198–210, January 2020. CODEN ???? ISSN 2329-9290.

URL https://dl.acm.
org/doi/abs/10.1109/
TASLP.2019.2950596.

Phan:2015:RRF

[PMMM15] H. Phan, M. Maas, R. Mazur, and A. Mertins. Random regression forests for acoustic event detection and classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (1):20–31, January 2015. CODEN ???? ISSN 2329-9290.

Pishdadian:2019:MRC

[PP19] F. Pishdadian and B. Pardo.

Multi-resolution common fate transform. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(2): 342–354, February 2019. CODEN ???? ISSN 2329-9290.

Parekh:2024:TIA

[PPM^{+}24] Jayneel Parekh, Sanjeel Parekh, Pavlo Mozharovskyi, Gaël Richard, and Florence d’Alché Buc. Tackling interpretability in audio classification networks with non-negative matrix factorization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1392– 1405, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3358049.

Pradhan:2017:IPD

[PPSG17] Somanath Pradhan, Vinal Patel, Dipen Somani, and Nithin V. George. An improved proportionate delayless multibandstructured subband adaptive feedback canceller for digital hearing aids. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):1633–1643, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7931717/.

Pal:2017:SMU

[PS17] Monisankha Pal and Goutam Saha. Spectral mapping using prior re-estimation of i-vectors and system fusion for voice conversion. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(11): 2071–2084, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8015184/.

Pak:2019:SLB

[PS19] J. Pak and J. W. Shin.

Sound localization based on phase difference enhancement using deep neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(8): 1335–1345, August 2019. CODEN ???? ISSN 2329-9290.

Parada:2016:SCN

[PSL^{+}16] Pablo Peso Parada, Dushyant Sharma, Jose Lainez, Daniel Barreda, Toon van Waterschoot, and Patrick A. Naylor. A single-channel non-intrusive C50 estimator correlated with speech recognition performance. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 719–732, April 2016. CO-DEN ???? ISSN 2329-9290.

Park:2017:ECD

[PSL17] Jeongsoo Park, Jaeyoung Shin, and Kyogu Lee. Exploiting continuity/discontinuity of basis vectors in spectrogram decomposition for harmonicpercussive sound separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(5):1061–1074, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7876808/.

Percival:2014:STE

[PT14] G. Percival and G. Tzanetakis. Streamlined tempo estimation based on autocorrelation and crosscorrelation with pulses. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1765–1776, December 2014. CODEN ???? ISSN 2329-9290.

Poletti:2021:STM

[PT21] Mark A. Poletti and Paul D. Teal. A superfast Toeplitz matrix inversion method for singleand multi-channel inverse filters and its application to room equalization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3144–3157, 2021. CODEN ???? ISSN 2329-9290.

Pandey:2019:NFC

[PW19] A. Pandey and D. Wang. A new framework for CNN-based speech enhancement in the time domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(7):1179–1188, July 2019. CODEN ???? ISSN 2329-9290.

Pandey:2021:DCS

[PW21] Ashutosh Pandey and DeLiang Wang. Dense CNN with self-attention for time-domain speech enhancement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1270–1279, 2021. CODEN ???? ISSN 2329-9290.

Pandey:2022:SAR

[PW22] Ashutosh Pandey and DeLiang Wang. Selfattending RNN for speech enhancement to improve cross-corpus generalization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1374–1385, 2022. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3161143.

Pandey:2023:ATN

[PW23] Ashutosh Pandey and DeLiang Wang. Attentive training: a new training framework for speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1360– 1370, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3260711.

Petermann:2023:TCF

[PWS^{+}23] Darius Petermann, Gordon Wichern, Aswin Shanmugam Subramanian, Zhong-Qiu Wang, and Jonathan Le Roux. Tackling the cocktail fork problem for separation and transcription of real-world soundtracks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2592– 2605, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3290428.

Pfeifenberger:2019:EBS

[PZP19] L. Pfeifenberger, M. Zöhrer, and F. Pernkopf. Eigenvectorbased speech mask estimation for multi-channel speech enhancement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2162–2172, December 2019. CODEN ???? ISSN 2329-9290.

Pan:2021:MTP

[PZW^{+}21] Zihan Pan, Malu Zhang, Jibin Wu, Jiadong Wang, and Haizhou Li. Multitone phase coding of interaural time difference for sound source localization with spiking neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??):

2656–2670, 2021. CODEN ???? ISSN 2329-9290.

Peng:2022:CSM

[PZZZ22] Yishu Peng, Sheng Zhang, Jiashu Zhang, and Wei Xing Zheng. Combined-sample multiband-structured subband filtering algorithms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):1083–1092, 2022. CODEN ???? ISSN 2329-9290.

Qian:2016:VDC

[QBTY16] Yanmin Qian, Mengxiao Bi, Tian Tan, and Kai Yu. Very deep convolutional neural networks for noise robust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2263–2276, ???? 2016. CODEN ???? ISSN 2329-9290.

Qian:2017:DFE

[QCDW17] Yanmin Qian, Nanxin Chen, Heinrich Dinkel, and Zhizheng Wu. Deep feature engineering for noise robust spoofing detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(10):1942– 1955, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7993036/.

Qin:2023:RMC

[QCL23] Xiaoyi Qin, Danwei Cai, and Ming Li. Robust multi-channel farfield speaker verification under different indomain data availability scenarios. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 71–85, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3212834.

Qin:2021:KWL

[QCN^{+}21] Libo Qin, Wanxiang Che, Minheng Ni, Yangming Li, and Ting Liu. Knowing where to leverage: Context-aware graph convolutional network with an adaptive fusion layer for contextual spoken language understanding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):1280–1289, 2021. CODEN ???? ISSN 2329-9290.

Qian:2021:AVD

[QCW21] Yanmin Qian, Zhengyang Chen, and Shuai Wang. Audio-visual deep neural network for robust person verification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??):

1079–1092, 2021. CODEN ???? ISSN 2329-9290.

Qi:2019:TDN

[QDSL19] J. Qi, J. Du, S. M. Siniscalchi, and C. Lee. A theory on deep neural network based vector-to-vector regression with an illustration of its expressive power in speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):1932– 1943, December 2019. CODEN ???? ISSN 2329-9290.

Quan:2024:SEL

[QL24] Changsheng Quan and Xiaofei Li. Spatial-Net: Extensively learning spatial information for multichannel joint speech separation, denoising and dereverberation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1310–1323, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3357036.

Qiang:2021:CLS

[QLL^{+}21] Jipeng Qiang, Xinyu Lu, Yun Li, Yunhao Yuan, and Xindong Wu. Chinese lexical simplification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1819–1828, 2021. CODEN ???? ISSN 2329-9290.

Qu:2024:DPR

[QLW^{+}24] Leyuan Qu, Taihao Li, Cornelius Weber, Theresa Pekarek-Rosin, Fuji Ren, and Stefan Wermter. Disentangling prosody representations with unsupervised speech reconstruction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):39– 54, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3320864.

Qiang:2021:LLS

[QLZ^{+}21] Jipeng Qiang, Yun Li, Yi Zhu, Yunhao Yuan, Yang Shi, and Xindong Wu. LSBert: Lexical simplification based on BERT. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3064– 3076, 2021. CODEN ???? ISSN 2329-9290.

Qian:2016:TPF

[QMS16] Xiaojun Qian, Helen Meng, and Frank Soong. A two-pass framework of mispronunciation detection and diagnosis for computer-aided pronunciation training. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(6): 1020–1028, June 2016. CODEN ???? ISSN 2329-9290.

Qian:2016:NNB

[QTY16] Yanmin Qian, Tian Tan, and Dong Yu. Neural network based multifactor aware joint training for robust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2231–2240, ???? 2016. CODEN ???? ISSN 2329-9290.

Quan:2024:MMA

[QVZY24] Zhibin Quan, Chi-Man Vong, Weili Zeng, and Wankou Yang. The Mor-PhEMe machine: an addressable neural memory for learning knowledgeregularized deep contextualized Chinese embedding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1673– 1686, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3364610.

Quan:2019:EFS

[QWL^{+}19] Z. Quan, Z. Wang, Y. Le, B. Yao, K. Li, and J. Yin. An efficient framework for sentence similarity modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(4):853–865, April 2019. CODEN ???? ISSN 2329-9290.

Qian:2023:AVC

[QWW^{+}23] Xinyuan Qian, Zhengdong Wang, Jiadong Wang, Guohui Guan, and Haizhou Li. Audiovisual cross-attention network for robotic speaker tracking. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):550– 562, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3226330.

Qin:2023:MPT

[QXW^{+}23] Libo Qin, Xiao Xu, Lehan Wang, Yue Zhang, and Wanxiang Che. Modularized pre-training for end-to-end task-oriented dialogue. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1601– 1610, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3244503.

Qi:2023:ELR

[QYCT23] Jun Qi, Chao-Han Huck Yang, Pin-Yu Chen, and Javier Tejedor. Exploiting low-rank tensortrain deep neural networks based on Riemannian gradient descent with illustrations of speech processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):633– 642, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3231714.

Qian:2022:ODU

[QZ22] Yanmin Qian and Zhikai Zhou. Optimizing data usage for low-resource speech recognition. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 394–403, 2022. CODEN ???? ISSN 2329-9290.

Qian:2021:JMN

[QZLH21] Tao Qian, Meishan Zhang, Yinxia Lou, and Daiwen Hua. A joint model for named entity recognition with sentence-level entity type attentions. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1438–1448, 2021. CODEN ???? ISSN 2329-9290.

Reddy:2023:EBS

[RA23] Mittapalle Kiran Reddy and Paavo Alku. Exemplarbased sparse representations for detection of Parkinson’s disease from speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1386– 1396, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3260709.

Ram:2018:SSM

[RAB18] D. Ram, A. Asaei, and H. Bourlard. Sparse subspace modeling for query by example spoken term detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(6):1130– 1143, June 2018. CODEN ???? ISSN 2329-9290.

Rakotomamonjy:2017:SRL

[Rak17] Alain Rakotomamonjy. Supervised representation learning for audio scene classification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(6): 1253–1265, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7933051/.

Rashwan:2015:DLF

[RARR15] M. A. A. Rashwan, A. A. Al Sallab, H. M. Raafat, and A. Rafea. Deep learning framework with confused sub-set resolution architecture for automatic Arabic diacritization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(3):505–516, March 2015. CODEN ???? ISSN 2329-9290.

Ruhland:2015:RGS

[RBBG15] M. Ruhland, J. Bitzer, M. Brandt, and S. Goetze. Reduction of Gaussian, supergaussian, and impulsive noise by interpolation of the binary mask residual. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1680–1691, October 2015. CODEN ???? ISSN 2329-9290.

Radmanesh:2016:LLO

[RBR16] Nasim Radmanesh, Ian S.

Burnett, and Bhaskar D. Rao. A lasso-LS optimization with a frequency variable dictionary in a multizone sound system. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (3):583–593, March 2016. CODEN ???? ISSN 2329-9290.

Rohmatillah:2023:HRL

[RC23] Mahdin Rohmatillah and Jen-Tzung Chien. Hierarchical reinforcement learning with guidance for multi-domain dialogue policy. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 748–761, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3235202.

Ruan:2018:SNE

[RCL18] Yu-Ping Ruan, Qian Chen, and Zhen-Hua Ling. A sequential neural encoder with latent structured description for modeling sentences. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(2): 231–242, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8106781/.

Rodriguez:2021:IWC

[RCR^{+}21] Demóstenes Z. Rodríguez, Dick Carrillo, Miguel A. Ramírez, Pedro H. J. Nardelli, and Sebastian Möller. Incorporating wireless communication parameters into the emodel algorithm. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 956–968, 2021. CODEN ???? ISSN 2329-9290.

Rafii:2014:CRB

[RDP14] Z. Rafii, Zhiyao Duan, and B. Pardo. Combining rhythm-based and pitch-based methods for background and melody separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1884– 1893, December 2014.

CODEN ???? ISSN 2329-9290.

Richter:2019:ICS

[RF19] J. Richter and J. Fels. On the influence of continuous subject rotation during high-resolution headrelated transfer function measurements. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(4): 730–741, April 2019. CO-DEN ???? ISSN 2329-9290.

Rakotomamonjy:2015:HGT

[RG15a] A. Rakotomamonjy and G. Gasso. Histogram of gradients of time– frequency representations for audio scene classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):142–153, January 2015. CODEN ???? ISSN 2329-9290.

Ranjan:2015:NLH

[RG15b] R. Ranjan and Woon-Seng Gan. Natural listening over headphones in augmented reality using adaptive filtering techniques. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):1988– 2002, November 2015. CODEN ???? ISSN 2329-

Rehr:2017:AAR

[RG17] Robert Rehr and Timo Gerkmann. An analysis of adaptive recursive smoothing with applications to noise PSD estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(2):397–408, ???? 2017. CODEN ???? ISSN 2329-9290.

Rehr:2018:ISG

[RG18] Robert Rehr and Timo Gerkmann. On the importance of super-Gaussian speech priors for machinelearning based speech enhancement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(2): 357–366, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8121999/.

Rehr:2021:SBF

[RG21] Robert Rehr and Timo Gerkmann. SNR-Based features and diverse training data for robust DNN-Based speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1937–1949, 2021. CODEN ???? ISSN 2329-9290.

Rouhe:2024:PCE

[RGK24] Aku Rouhe, Tamás Grósz, and Mikko Kurimo. Principled comparisons for end-to-end speech recognition: Attention vs hybrid at the 1000hour scale. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):623– 638, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3336517.

Ragni:2022:ICE

[RGR^{+}22] Anton Ragni, Mark J. F.

Gales, Oliver Rose, Katherine M. Knill, Alexandros Kastanos, Qiujia Li, and Preben M. Ness. Increasing context for estimating confidence scores in automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1319–1329, 2022. CODEN ???? ISSN 2329-9290.

Ranjan:2018:CLB

[RH18] Shivesh Ranjan and John H. L. Hansen. Curriculum learning based approaches for noise robust speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (1):197–210, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8080267/.

Rasumow:2016:RAS

[RHvdP^{+}16] E. Rasumow, M. Hansen, S. van de Par, D. Puschel, V. Mellert, S. Doclo, and M. Blau. Regularization approaches for synthesizing HRTF directivity patterns. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(2):215–225, ???? 2016. CODEN ???? ISSN 2329-9290.

Remaggi:2017:ARL

[RJCW17] Luca Remaggi, Philip J. B. Jackson, Philip Coleman, and Wenwu Wang. Acoustic reflector localization: Novel image source reversion and direct localization methods. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(2):296–309, ???? 2017. CODEN ???? ISSN 2329-9290.

Remaggi:2019:MCF

[RJW19] L. Remaggi, P. J. B. Jackson, and W. Wang. Modeling the comb filter effect and interaural coherence for binaural source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2263–2277, December 2019. CODEN ???? ISSN 2329-9290.

Rashobh:2014:MEK

[RKL14] R. S. Rashobh, A. W. H.

Khong, and Di Liu. Multichannel equalization in the KLT and frequency domains with application to speech dereverberation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (3):634–646, March 2014. CODEN ???? ISSN 2329-9290.

Reddy:2014:USD

[RKN14] V. V. Reddy, A. W. H.

Khong, and Boon Poh Ng. Unambiguous speech DOA estimation under spatial aliasing conditions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2133–2145, December 2014. CODEN ???? ISSN 2329-9290.

Liu:2022:LPR

[rLcHcC^{+}22] Da rong Liu, Po chun Hsu, Yi chen Chen, Sung feng Huang, Shun po Chuang, Da yi Wu, and Hung yi Lee. Learning phone recognition from unpaired audio and phone sequences based on generative adversarial network. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):230–243, 2022. CO-DEN ???? ISSN 2329-9290.

Rieser:2014:NLG

[RLK14] V. Rieser, O. Lemon, and S. Keizer. Natural language generation as incremental planning under uncertainty: Adaptive information presentation for statistical dialogue systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(5):979–994, May 2014. CODEN ???? ISSN 2329-9290.

Remes:2015:BCM

[RLPK15] U. Remes, A. Ramirez Lopez, K. Palomaki, and M. Kurimo. Bounded conditional mean imputation with observation uncertainties and acoustic model adaptation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(7): 1198–1208, July 2015. CODEN ???? ISSN 2329-9290.

Rafii:2018:OLA

[RLS^{+}18] Z. Rafii, A. Liutkus, F. Stöter, S. I. Mimilakis, D. FitzGerald, and B. Pardo. An overview of lead and accompaniment separation in music. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(8):1307–1335, August 2018. CODEN ???? ISSN 2329-9290.

Reindl:2014:MMI

[RMBK14] K. Reindl, S. Meier, H. Barfuss, and W. Kellermann. Minimum mutual information-based linearly constrained broadband signal extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (6):1096–1108, June 2014. CODEN ???? ISSN 2329-9290.

Rouvier:2015:ABV

[ROL^{+}15] M. Rouvier, S. Oger, G. Linares, D. Matrouf, B. Merialdo, and Y. Li. Audio-based video genre identification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(6): 1031–1041, June 2015. CODEN ???? ISSN 2329-9290.

r:2017:CRN

[rPH^{+}17] Emre Çakır, Giambattista Parascandolo, Toni Heittola, Heikki Huttunen, and Tuomas Virtanen. Convolutional recurrent neural networks for polyphonic sound event detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(6):1291–1303, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7933050/.

Raj:2023:SAT

[RPK23] Desh Raj, Daniel Povey, and Sanjeev Khudanpur. SURT 2.0: Advances in transducerbased multi-talker speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3800– 3813, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3318398.

Ren:2023:GVS

[RPL^{+}23] Yeqing Ren, Haipeng Peng, Lixiang Li, Xiaopeng Xue, Yang Lan, and Yixian Yang. Generalized voice spoofing detection via integral knowledge amalgamation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2461–2475, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3288416.

Raposo:2016:UGS

[RRdM16] Francisco Raposo, Ricardo Ribeiro, and David Martins de Matos. Using generic summarization to improve music information retrieval tasks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (6):1119–1128, June 2016. CODEN ???? ISSN 2329-9290.

Rahulamathavan:2019:PPI

[RSR^{+}19] Y. Rahulamathavan, K. R.

Sutharsini, I. G. Ray, R. Lu, and M. Rajarajan. Privacy-preserving iVector-based speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(3):496– 506, March 2019. CODEN ???? ISSN 2329-9290.

Raczynski:2014:GBM

[RV14] S. A. Raczyński and E. Vincent. Genre-based music language modeling with latent hierarchical Pitman-yor process allocation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(3): 672–681, March 2014. CODEN ???? ISSN 2329-9290.

Renkens:2017:WSL

[RV17] Vincent Renkens and Hugo Van hamme. Weakly supervised learning of hidden Markov models for spoken language acquisition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(2):285–295, ???? 2017.

CODEN ???? ISSN 2329-9290.

Ramo:2014:HPP

[RVB14] J. Ramo, V. Valimaki, and B. Bank. Highprecision parallel graphic equalizer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1894– 1904, December 2014. CODEN ???? ISSN 2329-9290.

Richard:2017:ISS

[RVB^{+}17] G. Richard, T. Virtanen, J. P. Bello, N. Ono, and H. Glotin. Introduction to the special section on sound scene and event analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(6): 1169–1171, ???? 2017. CODEN ???? ISSN 2329-9290.

Ruiz:2022:DCA

[RvWM22] Santiago Ruiz, Toon van Waterschoot, and Marc Moonen. Distributed combined acoustic echo cancellation and noise reduction in wireless acoustic sensor and actuator networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):534– 547, 2022. CODEN ???? ISSN 2329-9290.

Receveur:2016:TAS

[RWF16] Simon Receveur, Robin Weiß, and Tim Fingscheidt. Turbo automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (5):846–862, May 2016. CODEN ???? ISSN 2329-9290.

Richter:2023:SED

[RWL^{+}23] Julius Richter, Simon Welker, Jean-Marie Lemercier, Bunlong Lay, and Timo Gerkmann. Speech enhancement and dereverberation with diffusionbased generative models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2351–2364, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3285241.

Sarreshtedari:2015:WMD

[SAA15] S. Sarreshtedari, M. A.

Akhaee, and A. Abbasfar. A watermarking method for digital speech self-recovery. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):1917– 1925, November 2015. CODEN ???? ISSN 2329-

Saeidi:2016:FEU

[SAB16] R. Saeidi, P. Alku, and T. Backstrom. Feature extraction using powerlaw adjusted linear prediction with application to speaker recognition under severe vocal effort mismatch. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):42–53, January 2016. CODEN ???? ISSN 2329-9290.

Samarasinghe:2017:EDR

[SAC17] Prasanga N. Samarasinghe, Thushara D. Abhayapala, and Hanchi Chen. Estimating the direct-to-reverberant energy ratio using a spherical harmonics-based spatial correlation model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (2):310–319, ???? 2017. CODEN ???? ISSN 2329-9290.

Stafylakis:2016:TDS

[SAK16] Themos Stafylakis, Md. Jahangir Alam, and Patrick Kenny. Text-dependent speaker recognition with random digit strings. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (7):1194–1203, July 2016. CODEN ???? ISSN 2329-9290.

Samarasinghe:2014:WAL

[SAP14] P. Samarasinghe, T. Abhayapala, and M. Poletti. Wavefield analysis over large areas using distributed higher order microphones. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(3): 647–658, March 2014. CODEN ???? ISSN 2329-9290.

Samarasinghe:2015:EPR

[SAPB15] P. Samarasinghe, T. Abhayapala, M. Poletti, and T. Betlehem. An efficient parameterization of the room transfer function. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2217–2227, December 2015. CODEN ???? ISSN 2329-9290.

Sharma:2017:DSR

[SAS17] Pulkit Sharma, Vinayak Abrol, and Anil Kumar Sao. Deep-sparse-representation-based features for speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2162–2175, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8023884/.

Stankovic:2018:ARS

[SB18] Ljubiša Stanković and Miloš Brajović. Analysis of the reconstruction of sparse signals in the DCT domain applied to audio signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (7):1220–1235, July 2018. CODEN ???? ISSN 2329-9290.

Sigtia:2016:EEN

[SBD16] Siddharth Sigtia, Emmanouil Benetos, and Simon Dixon. An end-to-end neural network for polyphonic piano music transcription. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(5): 927–939, May 2016. CO-DEN ???? ISSN 2329-9290.

Stowell:2017:BSR

[SBG17] Dan Stowell, Emmanouil Benetos, and Lisa F. Gill. On-bird sound recordings: Automatic acoustic recognition of activities and contexts. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(6): 1193–1206, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7933044/.

Shimada:2019:USE

[SBM^{+}19] K. Shimada, Y. Bando, M. Mimura, K. Itoyama, K. Yoshii, and T. Kawahara. Unsupervised speech enhancement based on multichannel NMFinformed beamforming for noise-robust automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (5):960–971, May 2019. CODEN ???? ISSN 2329-9290.

Sekiguchi:2019:SSM

[SBN^{+}19] K. Sekiguchi, Y. Bando, A. A. Nugraha, K. Yoshii, and T. Kawahara. Semisupervised multichannel speech enhancement with a deep speech prior. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2197–2212, December 2019. CODEN ???? ISSN 2329-9290.

Stanciu:2024:DBW

[SBP^{+}24] Cristian-Lucian Stanciu, Jacob Benesty, Constantin Paleologu, Ruxandra-Liana Costea, Laura-Maria Dogariu, and Silviu Ciochină. Decompositionbased Wiener filter using the Kronecker product and conjugate gradient method. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 124–138, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3325136.

Sargent:2017:ESS

[SBV17] Gabriel Sargent, Frédéric Bimbot, and Emmanuel Vincent. Estimating the structural segmentation of popular music pieces under regularity constraints. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (2):344–358, ???? 2017. CODEN ???? ISSN 2329-9290.

Szurley:2016:BNC

[SBVM16] Joseph Szurley, Alexander Bertrand, Bas Van Dijk, and Marc Moonen. Binaural noise cue preservation in a binaural noise reduction system with a remote microphone signal. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(5):952–966, May 2016. CODEN ???? ISSN 2329-9290.

Sharma:2021:CEA

[SCB21] Rajib Sharma, Israel Cohen, and Baruch Berdugo. Controlling elevation and azimuth beamwidths with concentric circular microphone arrays. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1491–1502, 2021. CODEN ???? ISSN 2329-9290.

Shang:2023:MMN

[SCCM23] Xichen Shang, Chuxin Chen, Zipeng Chen, and Qianli Ma. Modularized mutuality network for emotion-cause pair extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):539– 549, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3228129.

Stoter:2019:CEN

[SCEH19] F. Stöter, S. Chakrabarty, B. Edler, and E. A. P. Habets. CountNet: Estimating the number of concurrent speakers using supervised learning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(2): 268–282, February 2019. CODEN ???? ISSN 2329-9290.

SongGong:2021:IMS

[SCW21] Kunkun SongGong, Huawei Chen, and Wenwu Wang. Indoor multi-speaker localization based on Bayesian nonparametrics in the circular harmonic domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1864–1880, 2021. CODEN ???? ISSN 2329-9290.

Shi:2022:USR

[SCY22] Xuan Shi, Erica Cooper, and Junichi Yamagishi. Use of speaker recognition approaches for learning and evaluating embedding representations of musical instrument sounds. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):367–377, 2022. CO-DEN ???? ISSN 2329-9290.

Schepker:2016:SPA

[SD16a] H. Schepker and S. Doclo. A semidefinite programming approach to minmax estimation of the common part of acoustic feedback paths in hearing aids. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(2):366–377, ???? 2016. CODEN ???? ISSN 2329-9290.

Schepker:2016:LSE

[SD16b] Henning Schepker and Simon Doclo. Least-squares estimation of the common pole-zero filter of acoustic feedback paths in hearing aids. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(8):1334–1347, ????

2016. CODEN ???? ISSN 2329-9290.

Sarmiento:2015:CFB

[SDDCC15] A. Sarmiento, I. Duran-Diaz, A. Cichocki, and S. Cruces. A contrast function based on generalized divergences for solving the permutation problem in convolved speech mixtures. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):1713– 1726, November 2015. CODEN ???? ISSN 2329-9290.

Salvati:2018:LCR

[SDF18] D. Salvati, C. Drioli, and G. L. Foresti. A low-complexity robust beamforming using diagonal unloading for acoustic source localization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):609–622, March 2018. CODEN ???? ISSN 2329-9290.

Sarfjoo:2017:UEN

[SDK17] Seyyed Saeed Sarfjoo, Cenk Demiroğlu, and Simon King. Using eigenvoices and nearestneighbors in HMM-based cross-lingual speaker adaptation with limited data. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):839–851, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7851067/.

Schmid:2014:VBI

[SEM^{+}14] D. Schmid, G. Enzner, S. Malik, D. Kolossa, and R. Martin. Variational Bayesian inference for multichannel dereverberation and noise reduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):1320–1335, August 2014. CODEN ???? ISSN 2329-9290.

Santos:2014:USC

[SF14] J. F. Santos and T. H. Falk. Updating the SRMR–CI metric for improved intelligibility prediction for cochlear implant users. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2197– 2206, December 2014. CODEN ???? ISSN 2329-9290.

Santos:2018:SDC

[SF18] J. F. Santos and T. H. Falk. Speech dereverberation with context-aware recurrent neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (7):1236–1246, July 2018.

CODEN ???? ISSN 2329-9290.

Schulze-Forster:2021:PLL

[SFDRB21] Kilian Schulze-Forster, Clement S. J. Doire, Gaël Richard, and Roland Badeau. Phoneme level lyrics alignment and textinformed singing voice separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2382– 2395, 2021. CODEN ???? ISSN 2329-9290.

Singh:2023:ETM

[SFEB23] Gopendra Vikram Singh, Mauajama Firdaus, Asif Ekbal, and Pushpak Bhattacharyya. EmoInt-Trans: a multimodal transformer for identifying emotions and intents in social conversations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):290– 300, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3224287.

Sheikh:2017:MSC

[SFIL17] Imran Sheikh, Dominique Fohr, Irina Illina, and Georges Linarès. Modelling semantic context of OOV words in large vocabulary continuous speech recognition. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(3): 598–610, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Shin:2015:PSE

[SFK15] Ho Seon Shin, T. Fingscheidt, and Hong-Goo Kang. A priori SNR estimation using air- and bone-conduction microphones. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):2015– 2025, November 2015. CODEN ???? ISSN 2329-9290.

Schulze-Forster:2023:UMS

[SFRK^{+}23] Kilian Schulze-Forster, Gaël Richard, Liam Kelley, Clement S. J. Doire, and Roland Badeau. Unsupervised music source separation using differentiable parametric source models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1276– 1289, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3252272.

Schlecht:2023:DFD

[SFV23] Sebastian J. Schlecht, Jon Fagerström, and Vesa Välimäki. Decorrelation in feedback delay networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3478– 3487, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313440.

Schwartz:2014:STU

[SG14] O. Schwartz and S. Gannot. Speaker tracking using recursive EM algorithms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):392– 402, February 2014. CO-DEN ???? ISSN 2329-9290.

Singh:2021:SSR

[SG21] Prachi Singh and Sriram Ganapathy. Selfsupervised representation learning with path integral clustering for speaker diarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1639– 1649, 2021. CODEN ???? ISSN 2329-9290.

Schroder:2015:STG

[SGA15] J. Schroder, S. Goetze, and J. Anemuller. Spectrotemporal Gabor filterbank features for acoustic event detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2198– 2208, December 2015. CODEN ???? ISSN 2329-

Stone:2022:ASV

[SGB22] Simon Stone, Yingming Gao, and Peter Birkholz. Articulatory synthesis of vocalized /r/ allophones in German. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):879– 889, 2022. CODEN ???? ISSN 2329-9290.

Song:2021:KGE

[SGF^{+}21] Wei Song, Jingjin Guo, Ruiji Fu, Ting Liu, and Lizhen Liu. A knowledge graph embedding approach for metaphor processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):406–420, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3040507.

Schwartz:2015:OSD

[SGH15a] B. Schwartz, S. Gannot, and E. A. P. Habets. Online speech dereverberation using Kalman filter and EM algorithm. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (2):394–406, February 2015. CODEN ???? ISSN 2329-9290.

Schwartz:2015:MMS

[SGH15b] O. Schwartz, S. Gannot, and E. A. P. Habets. Multi-microphone speech dereverberation and noise reduction using relative early transfer functions. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): 240–251, February 2015. CODEN ???? ISSN 2329-9290.

Schwartz:2016:EMA

[SGH16] Ofer Schwartz, Sharon Gannot, and Emanuël A. P. Habets. An expectation-maximization algorithm for multimicrophone speech dereverberation and noise reduction with coherence matrix estimation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(9): 1495–1510, ???? 2016. CODEN ???? ISSN 2329-9290.

Schwartz:2017:TMB

[SGH17a] Boaz Schwartz, Sharon Gannot, and Emanuël A. P. Habets. Two model-based EM algorithms for blind source separation in noisy environments. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2209– 2222, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

                                        ieeexplore.ieee.org/
                                        document/8007259/.

Schwartz:2017:CRB

[SGH17b] Ofer Schwartz, Sharon Gannot, and Emanuël A. P. Habets. Cramér– Rao bound analysis of reverberation level estimators for dereverberation and noise reduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):1680–1693, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7906601/.

Schwartz:2017:MLB

[SGH17c] Ofer Schwartz, Sharon Gannot, and Emanuël A. P. Habets. Multispeaker LCMV beamformer and postfilter for source separation and noise reduction. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(5): 940–951, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7842601/.

Shin:2020:LCU

[SgL20] Youhyun Shin and Sang goo Lee. Learning context using segment-level LSTM for neural sequence labeling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 105–115, January 2020. CODEN ???? ISSN 2329-9290. URL https:

//dl.acm.org/doi/abs/
10.1109/TASLP.2019.2948773.

Shi:2021:OOC

[SGL^{+}21] Dongyuan Shi, Woon-Seng Gan, Bhan Lam, Shulin Wen, and Xiaoyi Shen. Optimal outputconstrained active noise control based on inverse adaptive modeling leak factor estimate. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1256–1269, 2021. CODEN ???? ISSN 2329-9290.

Shi:2023:TLC

[SGL^{+}23] Dongyuan Shi, Woon-Seng Gan, Bhan Lam, Zhengding Luo, and Xiaoyi Shen. Transferable latent of CNN-based selective fixed-filter active noise control. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2910– 2921, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3261757.

Shi:2023:FDO

[SGLS23] Dongyuan Shi, Woon-Seng Gan, Bhan Lam, and Xiaoyi Shen. A frequency-domain outputconstrained active noise control algorithm based on an intuitive circulant convolutional penalty factor. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1318– 1332, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3260702.

Schasse:2015:TSF

[SGM^{+}15] A. Schasse, T. Gerkmann, R. Martin, W. Sorgel, T. Pilgrim, and H. Puder. Two-stage filter-bank system for improved singlechannel noise reduction in hearing aids. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): 383–393, February 2015. CODEN ???? ISSN 2329-9290.

Shao:2024:DIM

[SGY^{+}24] Qijie Shao, Pengcheng Guo, Jinghao Yan, Pengfei Hu, and Lei Xie. Decoupling and interacting multi-task learning network for joint speech and accent recognition. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 459–470, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3332542.

Sadjadi:2014:BSW

[SH14a] S. O. Sadjadi and J. H. L.

Hansen. Blind spectral weighting for robust speaker identification under reverberation mismatch. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(5): 937–945, May 2014. CO-DEN ???? ISSN 2329-9290.

Schuldt:2014:DRE

[SH14b] C. Schuldt and P. Handel. Decay rate estimators and their performance for blind reverberation time estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(8):1274–1284, August 2014. CODEN ???? ISSN 2329-9290.

Schlecht:2017:FDN

[SH17a] Sebastian J. Schlecht and Emanuël A. P. Habets. Feedback delay networks: Echo density and mixing time. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(2):374–383, ???? 2017. CODEN ???? ISSN 2329-9290.

Shokouhi:2017:TKE

[SH17b] Navid Shokouhi and John H. L. Hansen. Teager– Kaiser energy operators for overlapped speech detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(5):1035– 1047, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7872488/.

Schneider:2019:IDD

[SH19] M. Schneider and E. A. P. Habets. Iterative DFTdomain inverse filter optimization using a weighted least-squares criterion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):1957–1969, December 2019. CODEN ???? ISSN 2329-9290.

Sanchez-Hevia:2017:MLD

[SHAGPRZ17] Héctor A. Sánchez-Hevia, David Ayllón, Roberto Gil-Pita, and Manuel Rosa-Zurera. Maximum likelihood decision fusion for weapon classification in wireless acoustic sensor networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(6): 1172–1182, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7933058/.

Shankar:2023:DFB

[SHCV23] Ravi Shankar, Hsi-Wei Hsieh, Nicolas Charon, and Archana Venkataraman. A diffeomorphic flow-based variational framework for multi-speaker emotion conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):39– 53, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3209948.

Sarikaya:2014:ADB

[SHD14] R. Sarikaya, G. E. Hinton, and A. Deoras. Application of deep belief networks for natural language understanding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (4):778–784, April 2014. CODEN ???? ISSN 2329-9290.

Shum:2016:UAU

[SHDG16] Stephen H. Shum, David F.

Harwath, Najim Dehak, and James R. Glass. On the use of acoustic unit discovery for language recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(9):1665– 1676, ???? 2016. CODEN ???? ISSN 2329-9290.

Slizovskaia:2021:CSS

[SHG21] Olga Slizovskaia, Gloria Haro, and Emilia Gómez. Conditioned source separation for musical instrument performances. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2083– 2095, 2021. CODEN ???? ISSN 2329-9290.

Sari:2021:ANJ

[SHJT21] Leda Sari, Mark Hasegawa-Johnson, and Samuel Thomas. Auxiliary networks for joint speaker adaptation and speaker change detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):324–333, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3040626.

Sar:2021:CFA

[SHJY21] Leda Sar , Mark Hasegawa-Johnson, and Chang D. Yoo. Counterfactually fair automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3515–3525, 2021. CODEN ???? ISSN 2329-9290.

Schultz:2017:ISI

[SHKB17] T. Schultz, T. Hueber, D. J. Krusienski, and J. S. Brumberg. Introduction to the special issue on biosignal-based spoken communication. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12): 2254–2256, ???? 2017. CODEN ???? ISSN 2329-9290.

Su:2021:CLC

[SHL^{+}21] Yusheng Su, Xu Han, Yankai Lin, Zhengyan Zhang, Zhiyuan Liu, Peng Li, Jie Zhou, and Maosong Sun. CSS-LM: a contrastive framework for semi-supervised finetuning of pre-trained language models. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2930–2941, 2021. CODEN ???? ISSN 2329-9290.

Su:2023:RTD

[SHZ^{+}23] Xinxin Su, Zhen Huang, Yunxiang Zhao, Yifan Chen, Yong Dou, and Hengyue Pan. Recent trends in deep learning based textual emotion cause extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2765–2786, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3254166.

Sawada:2023:MFF

[SIKN23] Hiroshi Sawada, Rintaro Ikeshita, Keisuke Kinoshita, and Tomohiro Nakatani. Multi-frame full-rank spatial covariance analysis for underdetermined blind source separation and dereverberation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3589– 3602, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3313446.

Sakata:2021:VTL

[SIUW21] Tadashi Sakata, Naomitsu Ikeda, Yuichi Ueda, and Akira Watanabe. Vocal tract length estimation using accumulated means of formants and its effects on speakernormalization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1049–1064, 2021. CODEN ???? ISSN 2329-9290.

Schwarz:2015:CDP

[SK15] A. Schwarz and W. Kellermann. Coherent-to-diffuse power ratio estimation for dereverberation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(6): 1006–1018, June 2015. CODEN ???? ISSN 2329-9290.

Schneider:2016:MAE

[SK16a] Martin Schneider and Walter Kellermann. Multichannel acoustic echo cancellation in the wave domain with increased robustness to nonuniqueness. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(3):518–529, March 2016. CODEN ???? ISSN 2329-9290.

Sharma:2016:TSS

[SK16b] A. Sharma and S. Kaul.

Two-stage supervised learningbased method to detect screams and cries in urban environments. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(2): 290–299, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Saki:2017:RTU

[SK17] Fatemeh Saki and Nasser Kehtarnavaz. Real-time unsupervised classification of environmental noise signals. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(8): 1657–1667, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7937831/.

Surendran:2018:OPC

[SK18] S. Surendran and T. K.

Kumar. Oblique projection and cepstral subtraction in signal subspace speech enhancement for colored noise reduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2328–2340, December 2018. CODEN ???? ISSN 2329-9290.

Stafylakis:2016:SCF

[SKAK16] T. Stafylakis, P. Kenny, M. J. Alam, and M. Kockmann. Speaker and channel factors in textdependent speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):65–78, January 2016. CODEN ???? ISSN 2329-9290.

Souden:2014:LFI

[SKDN14] M. Souden, K. Kinoshita, M. Delcroix, and T. Nakatani. Location feature integration for clustering-based speech separation in distributed microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):354–367, February 2014. CODEN ???? ISSN 2329-9290.

Sugiura:2015:OCG

[SKH^{+}15a] R. Sugiura, Y. Kamamoto, N. Harada, H. Kameoka, and T. Moriya. Optimal coding of generalized-Gaussian-distributed frequency spectra for lowdelay audio Coder with powered all-pole spectrum estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):1309– 1321, August 2015. CO-DEN ???? ISSN 2329-9290.

Sugiura:2015:RWS

[SKH^{+}15b] R. Sugiura, Y. Kamamoto, N. Harada, H. Kameoka, and T. Moriya. Resolution warped spectral representation for low-delay and low-bit-rate audio Coder. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): 288–299, February 2015. CODEN ???? ISSN 2329-9290.

Shen:2014:DPR

[SKK14] Mo Shen, D. Kawahara, and S. Kurohashi. Dependency parse reranking with rich subtree features. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (7):1208–1218, July 2014. CODEN ???? ISSN 2329-9290.

Spille:2017:CBC

[SKM17] Constantin Spille, Birger Kollmeier, and Bernd T. Meyer. Combining binaural and cortical features for robust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(4):756–767, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7837638/.

Sugiura:2019:SCD

[SKM19] R. Sugiura, Y. Kamamoto, and T. Moriya. Shape control of discrete generalized Gaussian distributions for frequencydomain audio coding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2234–2248, December 2019. CODEN ???? ISSN 2329-9290.

Senoussaoui:2014:SCD

[SKSD14] M. Senoussaoui, P. Kenny, T. Stafylakis, and P. Dumouchel. A study of the cosine distance-based mean shift for telephone speech diarization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): 217–227, January 2014. CODEN ???? ISSN 2329-9290.

Supraja:2021:RPB

[SKT21] S. Supraja, Andy W. H. Khong, and S. Tatinati. Regularized phrase-based topic model for automatic question classification with domain-agnostic class labels. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3604– 3616, 2021. CODEN ???? ISSN 2329-9290.

Shen:2015:LCD

[SLC^{+}15] Qing Shen, Wei Liu, Wei Cui, Siliang Wu, Y. D. Zhang, and M. G. Amin. Low-complexity direction-of-arrival estimation based on wideband co-prime arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (9):1445–1456, September 2015. CODEN ???? ISSN 2329-9290.

Shah:2015:NAE

[SLGA15] P. Shah, I. Lewis, S. Grant, and S. Angrignon. Nonlinear acoustic echo cancellation using voltage and current feedback. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(10): 1589–1599, October 2015. CODEN ???? ISSN 2329-9290.

Lee:2015:SCR

[sLGyLaC15] Lin shan Lee, J. Glass, Hung yi Lee, and Chun an Chan. Spoken content retrieval — beyond cascading speech recognition with text retrieval. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (9):1389–1420, September 2015. CODEN ???? ISSN 2329-9290.

Sun:2015:SEU

[SLGZ15] Meng Sun, Yinan Li, J. F. Gemmeke, and Xiongwei Zhang. Speech enhancement under low SNR conditions via noise estimation using sparse and lowrank NMF with Kullback– Leibler divergence. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(7): 1233–1242, July 2015. CODEN ???? ISSN 2329-9290.

Steiner:2017:STM

[SLH17] Ingmar Steiner, Sébastien Le Maguer, and Alexander Hewer. Synthesis of tongue motion and acoustics from text using a multimodal articulatory database. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12): 2351–2361, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8114352/.

Sward:2018:GFF

[SLJ18] Johan Swärd, Hongbin Li, and Andreas Jakobsson. Off-grid fundamental frequency estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (2):296–303, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8115315/.

Sizov:2017:DOD

[SLK17] Aleksandr Sizov, Kong Aik Lee, and Tomi Kinnunen. Direct optimization of the detection cost for Ivector-based spoken language recognition. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(3): 588–597, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Shepstone:2016:TVM

[SLL^{+}16] Sven Ewan Shepstone, Kong Aik Lee, Haizhou Li, Zheng-Hua Tan, and Søren Holdt Jensen. Total variability modeling using source-specific priors. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (3):504–517, March 2016. CODEN ???? ISSN 2329-9290.

Souviraa-Labastie:2015:MCA

[SLOVB15] N. Souviraa-Labastie, A. Olivero, E. Vincent, and F. Bimbot. Multi-channel audio source separation using multiple deformed references. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):1775– 1787, November 2015. CODEN ???? ISSN 2329-9290.

Swietojanski:2016:LHU

[SLR16] Pawel Swietojanski, Jinyu Li, and Steve Renals. Learning hidden unit contributions for unsupervised acoustic model adaptation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(8):1450– 1463, ???? 2016. CODEN ???? ISSN 2329-9290.

Su:2015:ACC

[SLW15] Rongfeng Su, Xunying Liu, and Lan Wang. Automatic complexity control of generalized variable parameter HMMs for noise robust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):102–114, January 2015. CODEN ???? ISSN 2329-9290.

Su:2020:CDD

[SLWY20] Rongfeng Su, Xunying Liu, Lan Wang, and Jingzhou Yang. Crossdomain deep visual feature generation for Mandarin audio–visual speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):185– 197, January 2020. CO-DEN ???? ISSN 2329-9290. URL https:/

/dl.acm.org/doi/abs/
10.1109/TASLP.2019.2950602.

Su:2014:SMM

[SLY14] Li Su, Hsin-Ming Lin, and Yi-Hsuan Yang. Sparse modeling of magnitude and phase-derived spectra for playing technique classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2122– 2132, December 2014. CODEN ???? ISSN 2329-9290.

Shi:2021:GPS

[SLZ^{+}21] Liming Shi, Taewoong Lee, Lijun Zhang, Jesper Kjær Nielsen, and Mads Græsbøll Christensen. Generation of personal sound zones with physical meaningful constraints and conjugate gradient method. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 823–837, 2021. CODEN ???? ISSN 2329-9290.

Schasse:2014:ESS

[SM14] A. Schasse and R. Martin. Estimation of subband speech correlations for noise reduction via MVDR processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (9):1355–1365, September 2014. CODEN ???? ISSN 2329-9290.

Stahl:2018:PSS

[SM18] Johannes Stahl and Pejman Mowlaee. A pitchsynchronous simultaneous detection-estimation framework for speech enhancement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(2): 436–450, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8125741/.

Schroder:2017:CAA

[SMA^{+}17] Jens Schröder, Niko Moritz, Jörn Anemüller, Stefan Goetze, and Birger Kollmeier. Classifier architectures for acoustic scenes and events: Implications for DNNs, TDNNs, and perceptual features from DCASE 2016. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(6):1304–1314, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7933048/.

Stone:2018:CEP

[SMB18] S. Stone, M. Marxen, and P. Birkholz. Construction and evaluation of a parametric one-dimensional vocal tract model. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(8): 1381–1392, August 2018. CODEN ???? ISSN 2329-9290.

Speed:2014:MVT

[SMH14] M. Speed, D. Murphy, and D. Howard. Modeling the vocal tract transfer function using a 3D digital waveguide mesh. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): 453–464, February 2014. CODEN ???? ISSN 2329-9290.

Saeki:2024:TIG

[SML^{+}24] Takaaki Saeki, Soumi Maiti, Xinjian Li, Shinji Watanabe, Shinnosuke Takamichi, and Hiroshi Saruwatari. Text-inductive graphone-based language adaptation for low-resource speech synthesis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 1829–1844, 2024. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2024.3369537.

Spagnol:2021:ESN

[SMO^{+}21] Simone Spagnol, Riccardo Miccini, Marius George Onofrei, Runar Unnthorsson, and Stefania Serafin. Estimation of spectral notches from pinna meshes: Insights from a simple computational model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2683– 2695, 2021. CODEN ???? ISSN 2329-9290.

Stevens:2017:MSR

[SMSV17] Francis Stevens, Damian T. Murphy, Lauri Savioja, and Vesa Välimäki. Modeling sparsely reflecting outdoor acoustic scenes using the waveguide Web. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (8):1566–1578, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7918519/.

Sutojo:2022:SMM

[SMvdP22] Sarinah Sutojo, Tobias May, and Steven van de Par. Segmentation of multitalker mixtures based on local feature contrasts and auditory glimpses. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1249–1262, 2022. CODEN ???? ISSN 2329-9290.

Serizel:2014:LRA

[SMVW14] R. Serizel, M. Moonen, B. Van Dijk, and J. Wouters. Low-rank approximation based multichannel Wiener filter algorithms for noise reduction with application in cochlear implants. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(4): 785–799, April 2014. CO-DEN ???? ISSN 2329-9290.

Sato:2021:ASP

[SNK21] Ryotaro Sato, Kenta Niwa, and Kazunori Kobayashi. Ambisonic signal processing DNNs guaranteeing rotation, scale and time translation equivariance. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1449–1462, 2021. CODEN ???? ISSN 2329-9290.

Sundermeyer:2015:FRL

[SNS15] M. Sundermeyer, H. Ney, and R. Schluter. From feedforward to recurrent LSTM neural networks for language modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (3):517–529, March 2015. CODEN ???? ISSN 2329-9290.

Schepker:2019:NSB

[SNTD19] H. Schepker, S. E. Nordholm, L. T. T. Tran, and S. Doclo. Null-steering beamformer-based feedback cancellation for multi-microphone hearing aids with incoming signal preservation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(4): 679–691, April 2019. CO-DEN ???? ISSN 2329-9290.

Saito:2015:CBS

[SOF15] S. Saito, K. Oishi, and T. Furukawa. Convolutive blind source separation using an iterative least-squares algorithm for non-orthogonal approximate joint diagonalization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2434– 2448, December 2015. CODEN ???? ISSN 2329-9290.

Son:2021:DSA

[Son21] Phan Le Son. On the design of sparse arrays with frequency-invariant beam pattern. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):226–238, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3040033.

Strasser:2015:AFC

[SP15] F. Strasser and H. Puder. Adaptive feedback cancellation for realistic hearing aid applications. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2322–2333, December 2015. CODEN ???? ISSN 2329-9290.

Sailor:2016:NUA

[SP16] Hardik B. Sailor and Hemant A. Patil. Novel unsupervised auditory filterbank learning using convolutional RBM for speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (12):2341–2353, ???? 2016. CODEN ???? ISSN 2329-9290.

Sharma:2017:SMU

[SP17] Bidisha Sharma and S. R. Mahadeva Prasanna. Sonority measurement using system, source, and suprasegmental information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(3):505–518, ???? 2017. CODEN ???? ISSN 2329-9290.

Stefanakis:2017:CPC

[SPM17a] N. Stefanakis, D. Pavlidi, and A. Mouchtaris. Corrections to “Perpendicular Cross-Spectra Fusion for Sound Source Localization With a Planar Microphone Array” [Sep 17 1821–1835]. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(11): 2251, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

                                        ieeexplore.ieee.org/
                                        document/8082821/. See

[SPM17b].

Stefanakis:2017:PCS

[SPM17b] Nikolaos Stefanakis, Despoina Pavlidi, and Athanasios Mouchtaris. Perpendicular cross-spectra fusion for sound source localization with a planar microphone array. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(9): 1821–1835, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7955064/. See

corrections [SPM17a].

Shabtai:2014:GSA

[SR14] N. R. Shabtai and B. Rafaely. Generalized spherical array beamforming for binaural speech reproduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):238–247, January 2014. CODEN ???? ISSN 2329-9290.

Sahoo:2016:NMG

[SR16a] Subhasmita Sahoo and Aurobinda Routray. A novel method of glottal inverse filtering. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(7): 1230–1241, July 2016. CODEN ???? ISSN 2329-9290.

Swietojanski:2016:DPU

[SR16b] Pawel Swietojanski and Steve Renals. Differentiable pooling for unsupervised acoustic model adaptation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1773– 1784, ???? 2016. CODEN ???? ISSN 2329-9290.

Subakan:2023:ESA

[SRC^{+}23] Cem Subakan, Mirco Ravanelli, Samuele Cornell, François Grondin, and Mirko Bronzi. Exploring self-attention mechanisms for speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2169– 2180, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3282097.

Spa:2015:GIE

[SRH15] C. Spa, A. Rey, and E. Hernandez. A GPU implementation of an explicit compact FDTD algorithm with a digital impedance filter for room acoustics applications. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):1368–1380, August 2015. CODEN ???? ISSN 2329-9290.

Sun:2023:IST

[SRT23] Siqi Sun, Korin Richmond, and Hao Tang. Improving Seq2Seq TTS frontends with transcribed speech audio. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1940– 1952, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3273414.

Samarakoon:2016:FHL

[SS16] Lahiru Samarakoon and Khe Chai Sim. Factorized hidden layer adaptation for deep neural network based acoustic modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2241–2250, ???? 2016. CODEN ???? ISSN 2329-9290.

Shalyminov:2021:GGR

[SSAS21] Igor Shalyminov, Alessandro Sordoni, Adam Atkinson, and Hannes Schulz. GRTr: Generative-retrieval transformers for dataefficient dialogue domain adaptation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2484– 2492, 2021. CODEN ???? ISSN 2329-9290.

Serai:2022:HSR

[SSFL22] Prashant Serai, Vishal Sunder, and Eric Fosler-Lussier. Hallucination of speech recognition errors with sequence to sequence learning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 890–900, 2022. CODEN ???? ISSN 2329-9290.

Salehi:2018:LBR

[SSFP18] H. Salehi, D. Suelzle, P. Folkeard, and V. Parsa. Learning-based referencefree speech quality measures for hearing aid applications. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2277– 2288, December 2018. CODEN ???? ISSN 2329-9290.

Shen:2021:RRA

[SSGL21] Xingfa Shen, Xingkun Shao, Quanbo Ge, and Lili Liu. RARS: Recognition of audio recording source based on residual neural network. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):575–584, January

2021. URL https://dl.

acm.org/doi/10.1109/
TASLP.2020.3039597.

Sasaki:2021:SBC

[SSI21] Shota Sasaki, Jun Suzuki, and Kentaro Inui. Subwordbased compact reconstruction for open-vocabulary neural word embeddings. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):3551–3564, 2021. CODEN ???? ISSN 2329-9290.

Song:2017:ESE

[SSK17] Eunwoo Song, Frank K. Soong, and Hong-Goo Kang. Effective spectral and excitation modeling techniques for LSTM– RNN-based speech synthesis systems. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(11): 2152–2161, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8017571/.

Sigtia:2016:AES

[SSKP16] Siddharth Sigtia, Adam M. Stark, Sacha Krstulović, and Mark D. Plumbley. Automatic environmental sound recognition: Performance versus computational cost. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(11):

2096–2107, ???? 2016. CODEN ???? ISSN 2329-9290.

Sundar:2018:TBM

[SSS18] H. Sundar, T. V. Sreenivas, and C. S. Seelamantula. TDOA-Based multiple acoustic source localization without association ambiguity. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(11): 1976–1990, November 2018. CODEN ???? ISSN 2329-9290.

Shahrebabaki:2022:AAM

[SSSS22] Abdolreza Sabzi Shahrebabaki, Giampiero Salvi, Torbjørn Svendsen, and Sabato Marco Siniscalchi. Acoustic-to-articulatory mapping with joint optimization of deep speech enhancement and articulatory inversion models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):135–147, 2022. CO-DEN ???? ISSN 2329-9290.

Salvador:2018:BMF

[SSTS18] C. D. Salvador, S. Sakamoto, J. Treviño, and Y. Suzuki. Boundary matching filters for spherical microphone and loudspeaker arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):461–474, March 2018. CODEN ???? ISSN 2329-9290.

Sereshkeh:2017:ECC

[STBC17] Alborz Rezazadeh Sereshkeh, Robert Trott, Aurélien Bricout, and Tom Chau. EEG classification of covert speech using regularized neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (12):2292–2300, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8114360/.

Sahidullah:2018:RVL

[STH^{+}18] Md. Sahidullah, Dennis Alexander Lehmann Thomsen, Rosa Gonzalez Hautamäki, Tomi Kinnunen, Zheng-Hua Tan, Robert Parts, and Martti Pitkänen. Robust voice liveness detection and speaker verification using throat microphones. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (1):44–56, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8060607/.

Saito:2018:SPS

[STS18] Yuki Saito, Shinnosuke Takamichi, and Hiroshi Saruwatari. Statistical parametric speech synthesis incorporating generative adversarial networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (1):84–96, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8063435/.

Saito:2021:PSA

[STS21] Yuki Saito, Shinnosuke Takamichi, and Hiroshi Saruwatari. Perceptualsimilarity-aware deep speaker representation learning for multi-speaker generative modeling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1033–1048, 2021. CODEN ???? ISSN 2329-9290.

Sarkar:2019:TCL

[STT^{+}19] A. K. Sarkar, Z. Tan, H. Tang, S. Shon, and J. Glass. Time-contrastive learning based deep bottleneck features for textdependent speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(8):1267– 1279, August 2019. CO-DEN ???? ISSN 2329-9290.

Sugimoto:2022:LLC

[Sug22] Takehiro Sugimoto. Loudnesslevel-chasing algorithm for multiformat live audio production. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1290–1304, 2022. CODEN ???? ISSN 2329-9290.

Sharma:2024:TFS

[SUK24] Garima Sharma, Karthikeyan Umapathy, and Sridhar Krishnan. Timefrequency scattergrams for biomedical audio signal representation and classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):564– 576, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3332544.

enel:2018:SSI

[ŞUY^{+}18] Lütfi Kerem Şenel, İhsan Utlu, Veysel Yücesoy, Aykut Koç, and Tolga Çukur. Semantic structure and interpretability of word embeddings. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):1769–1779, October 2018. CODEN ???? ISSN 2329-9290.

Sahraeian:2017:CMS

[SV17] Reza Sahraeian and Dirk Van Compernolle. Crosslingual and multilingual speech recognition based on the speech manifold. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(12):2301–2312, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8114354/.

Sahraeian:2018:CET

[SV18] R. Sahraeian and D. Van Compernolle. Crossentropy training of DNN ensemble acoustic models for low-resource ASR. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):1991–2001, November 2018. CODEN ???? ISSN 2329-9290.

Sheaffer:2015:BRF

[SvWRK15] J. Sheaffer, M. van Walstijn, B. Rafaely, and K. Kowalczyk. Binaural reproduction of finite difference simulations using spherical array processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2125–2135, December 2015. CODEN ???? ISSN 2329-9290.

Sonnleitner:2016:RQB

[SW16] Reinhard Sonnleitner and Gerhard Widmer. Robust quad-based audio fingerprinting. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): 409–421, March 2016. CODEN ???? ISSN 2329-9290.

Sharma:2020:AES

[SW20] Bidisha Sharma and Ye Wang. Automatic evaluation of song intelligibility using singing adapted STOI and vocalspecific features. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 319–331, January 2020. CODEN ???? ISSN 2329-9290. URL https:

//dl.acm.org/doi/abs/
10.1109/TASLP.2019.2955253.

Su:2020:ABR

[SWC20] Ming-Hsiang Su, Chung-Hsien Wu, and Liang-Yu Chen. Attention-based response generation using parallel double Q-learning for dialog policy decision in a conversational system. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):131–143, January 2020. CODEN ???? ISSN 2329-9290.

URL https://dl.acm.
org/doi/abs/10.1109/
TASLP.2019.2949687.

Su:2021:PSD

[SWC^{+}21] Yixuan Su, Yan Wang, Deng Cai, Simon Baker, Anna Korhonen, and Nigel Collier. PROTOTYPE-TO-STYLE: Dialogue generation with styleaware editing on retrieval memory. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2152– 2161, 2021. CODEN ???? ISSN 2329-9290.

Sun:2019:TSM

[SWCN19] Y. Sun, W. Wang, J. Chambers, and S. M. Naqvi. Two-stage monaural source separation in reverberant room environments using deep neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1): 125–139, January 2019. CODEN ???? ISSN 2329-9290.

Sudholt:2023:PDN

[SWEV23] David Sudholt, Alec Wright, Cumhur Erkut, and Vesa Valimaki. Pruning deep neural network models of guitar distortion effects. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):256– 264, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3223257.

Schultz:2017:BBS

[SWH^{+}17] Tanja Schultz, Michael Wand, Thomas Hueber, Dean J. Krusienski, Christian Herff, and Jonathan S. Brumberg. Biosignal-based spoken communication: a survey. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(12):2257–2271, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8114358/.

Shao:2022:AGN

[SWH22] Zhihong Shao, Zhongqin Wu, and Minlie Huang. AdvExpander: Generating natural language adversarial examples by expanding text. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1184–1196, 2022. CODEN ???? ISSN 2329-9290.

Sainath:2017:MSP

[SWW^{+}17] Tara N. Sainath, Ron J. Weiss, Kevin W. Wilson, Bo Li, Arun Narayanan, Ehsan Variani, Michiel Bacchiani, Izhak Shafran, Andrew Senior, Kean Chin, Ananya Misra, and Chanwoo Kim. Multichannel signal processing with deep neural networks for automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):965–979, ???? 2017.

CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7859320/.

Su:2016:ETT

[SWZ16] Ming-Hsiang Su, Chung-Hsien Wu, and Yu-Ting Zheng. Exploiting turntaking temporal evolution for personality trait perception in dyadic conversations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(4):733– 744, April 2016. CODEN ???? ISSN 2329-9290.

Su:2015:CST

[SY15] Li Su and Yi-Hsuan Yang.

Combining spectral and temporal representations for multipitch estimation of polyphonic music. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1600–1612, October 2015. CODEN ???? ISSN 2329-9290.

Sun:2024:DBM

[SYG^{+}24] Linhui Sun, Shuo Yuan, Aifei Gong, Lei Ye, and Eng Siong Chng. Dual-branch modeling based on state-space model for speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1457– 1467, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3362691.

Sisman:2021:OVC

[SYKL21] Berrak Sisman, Junichi Yamagishi, Simon King, and Haizhou Li. An overview of voice conversion and its challenges: From statistical modeling to deep learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):132–157, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3038524.

Sisman:2019:GSR

[SZL19] B. Sisman, M. Zhang, and H. Li. Group sparse representation with WaveNet vocoder adaptation for spectrum and prosody conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(6):1085– 1097, June 2019. CODEN ???? ISSN 2329-9290.

Shifas:2022:EEN

[SZS22] Muhammed P. V. Shifas, Cătălin Zorilă, and Yannis Stylianou. End-to-end neural based modification of noisy speech for speech-in-noise intelligibility improvement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):162–173, 2022. CO-DEN ???? ISSN 2329-9290.

Sun:2016:UNE

[SZVZ16] Meng Sun, Xiongwei Zhang, H. Van Hamme, and T. F. Zheng. Unseen noise estimation using separable deep auto encoder for speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):93–104, January 2016. CODEN ???? ISSN 2329-9290.

Sun:2023:MBW

[SZW23] Guangzhi Sun, Chao Zhang, and Philip C. Woodland. Minimising biasing word errors for contextual ASR with the tree-constrained pointer generator. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):345– 354, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3224286.

Su:2018:HSA

[SZX^{+}18] J. Su, J. Zeng, D. Xiong, Y. Liu, M. Wang, and J. Xie. A hierarchyto-sequence attentional neural machine translation model. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(3):

623–632, March 2018. CODEN ???? ISSN 2329-9290.

Tan:2024:ASC

[TALP24] Yizhou Tan, Haojun Ai, Shengchen Li, and Mark D. Plumbley. Acoustic scene classification across cities and devices via feature disentanglement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1286– 1297, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3353578.

Tao:2018:GNN

[TB18] F. Tao and C. Busso. Gating neural network for large vocabulary audiovisual speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (7):1290–1302, July 2018. CODEN ???? ISSN 2329-9290.

Tu:2019:MOF

[TC19] Q. Tu and H. Chen. On mainlobe orientation of the first- and secondorder differential microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2025– 2040, December 2019. CODEN ???? ISSN 2329-9290.

Tu:2022:TLB

[TC22] Quansheng Tu and Huawei Chen. Theoretical lower bounds on the performance of the first-order differential microphone arrays with sensor imperfections. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):785– 801, 2022. CODEN ???? ISSN 2329-9290.

Tan:2019:GRN

[TCW19] K. Tan, J. Chen, and D. Wang. Gated residual networks with dilated convolutions for monaural speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(1):189–198, January 2019. CODEN ???? ISSN 2329-9290.

Tian:2015:DIU

[TCY15] Ye Tian, Zhe Chen, and Fuliang Yin. Distributed IMM-unscented Kalman filter for speaker tracking in microphone array networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1637– 1647, October 2015. CO-DEN ???? ISSN 2329-9290.

Tahon:2016:TSS

[TD16] M. Tahon and L. Devillers. Towards a small set of robust acoustic features for emotion recognition: Challenges. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(1): 16–28, January 2016. CO-DEN ???? ISSN 2329-9290.

Tammen:2023:PEP

[TD23] Marvin Tammen and Simon Doclo. Parameter estimation procedures for deep multi-frame MVDR filtering for singlemicrophone speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3237– 3248, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3306715.

Thulke:2024:TOD

[TDDN24] David Thulke, Nico Daheim, Christian Dugast, and Hermann Ney. Taskoriented document-grounded dialog systems by HLTPR@RWTH for DSTC9 and DSTC10. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32 (??):733–741, 2024. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3267832.

Tu:2019:SEB

[TDL19] Y. Tu, J. Du, and C. Lee. Speech enhancement based on teacher student deep learning using improved speech presence probability for noiserobust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2080–2091, December 2019. CODEN ???? ISSN 2329-9290.

Turan:2016:SFE

[TE16] M. A. Tugtekin Turan and E. Erzin. Source and filter estimation for throat-microphone speech enhancement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(2): 265–275, ???? 2016. CO-DEN ???? ISSN 2329-9290.

Thaleiser:2023:BPM

[TE23] Stefan Thaleiser and Gerald Enzner. Binauralprojection multichannel Wiener filter for cuepreserving binaural speech enhancement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 3730–3745, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3317569.

Tervo:2016:SSD

[Ter16]  Sakari Tervo.  Single

snapshot detection and estimation of reflections from room impulse responses in the spherical harmonic domain. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(12): 2466–2480, ???? 2016. CODEN ???? ISSN 2329-9290.

Torres:2016:NEM

[TG16] H. Torres and J. Gurlekian. Novel estimation method for the superpositional intonation model. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(1): 151–160, January 2016. CODEN ???? ISSN 2329-9290.

Tesch:2021:NSF

[TG21] Kristina Tesch and Timo Gerkmann. Nonlinear spatial filtering in multichannel speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1795–1805, 2021. CODEN ???? ISSN 2329-9290.

Tesch:2023:IDN

[TG23] Kristina Tesch and Timo Gerkmann. Insights into deep non-linear filters for improved multichannel speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):563–575, 2023. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2022.3221046.

Tesch:2024:MCS

[TG24] Kristina Tesch and Timo Gerkmann. Multi-channel speech separation using spatially selective deep non-linear filters. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 542–553, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3334101.

Torres:2021:FPD

[TGGE21] Humberto M. Torres, Mercedes Güemes, Jorge A. Gurlekian, and Diego A. Evin. F0 perturbation due to articulatory movements: Filtering, characterization and applications. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1977–1986, 2021. CODEN ???? ISSN 2329-9290.

Tong:2022:SIU

[TGZ22] Yu Tong, Jingzhi Guo, and Jizhe Zhou. Separation inference: a unified framework for word segmentation in East Asian languages. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1521– 1530, 2022. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3161142.

Taseska:2014:ISF

[TH14] M. Taseska and E. A. P. Habets. Informed spatial filtering for sound extraction using distributed microphone arrays. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(7): 1195–1207, July 2014. CODEN ???? ISSN 2329-9290.

Taseska:2016:SSF

[TH16] Maja Taseska and Emanuël A. P. Habets. Spotforming: Spatial filtering with distributed arrays for position-selective sound acquisition. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(7): 1291–1304, July 2016. CODEN ???? ISSN 2329-9290.

Taseska:2017:NNP

[TH17] Maja Taseska and Emanuël A. P. Habets. Nonstationary noise PSD matrix estimation for multichannel blind speech extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(11):2223–2236, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8030098/.

Taseska:2018:BSS

[TH18] M. Taseska and E. A. P.

Habets. Blind source separation of moving sources using sparsitybased source detection and tracking. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(3): 657–670, March 2018. CODEN ???? ISSN 2329-9290.

Tavakoli:2016:FSE

[TJCB16] Vincent Mohammad Tavakoli, Jesper Rindom Jensen, Mads Græsbøll Christensen, and Jacob Benesty. A framework for speech enhancement with ad hoc microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (6):1038–1051, June 2016. CODEN ???? ISSN 2329-9290.

Thuillier:2024:HIU

[TJV24] Etienne Thuillier, Craig T.

Jin, and Vesa Välimäki. HRTF interpolation using a spherical neural process meta-learner. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1790– 1802, 2024. ISSN 2329-

9290. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2024.3361377.

Togami:2014:SOA

[TK14] M. Togami and Y. Kawaguchi. Simultaneous optimization of acoustic echo reduction, speech dereverberation, and noise reduction against mutual interference. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(11):1612– 1623, November 2014. CODEN ???? ISSN 2329-9290.

Torcoli:2021:OMP

[TKH21] Matteo Torcoli, Thorsten Kastner, and Jürgen Herre. Objective measures of perceptual audio quality reviewed: an evaluation of their application domain dependence. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1530– 1541, 2021. CODEN ???? ISSN 2329-9290.

Tobing:2017:ACS

[TKT17] Patrick Lumban Tobing, Kazuhiro Kobayashi, and Tomoki Toda. Articulatory controllable speech modification based on statistical inversion and production mappings. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(12):

2337–2350, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8114384/.

Tuan:2019:ICS

[TL19] Y. Tuan and H. Lee. Improving conditional sequence generative adversarial networks by stepwise evaluation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(4): 788–798, April 2019. CO-DEN ???? ISSN 2329-9290.

Tao:2023:SST

[TLD^{+}23] Ruijie Tao, Kong Aik Lee, Rohan Kumar Das, Ville Hautamäki, and Haizhou Li. Self-supervised training of speaker encoder with multi-modal diverse positive pairs. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1706–1719, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3268568.

Tian:2017:EBA

[TLW^{+}17] Xiaohai Tian, Siu Wa Lee, Zhizheng Wu, Eng Siong Chng, and Haizhou Li. An exemplar-based approach to frequency warping for voice conversion. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(10): 1863–1876, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7968404/.

Tang:2017:CJT

[TLWV17] Zhiyuan Tang, Lantian Li, Dong Wang, and Ravichander Vipperla. Collaborative joint training with multitask recurrent model for speech and speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (3):493–504, ???? 2017. CODEN ???? ISSN 2329-9290.

Taghia:2014:OIM

[TM14] J. Taghia and R. Martin. Objective intelligibility measures based on mutual information for speech subjected to speech enhancement processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):6–16, January 2014. CODEN ???? ISSN 2329-9290.

Taghia:2016:FDA

[TM16] Jalal Taghia and Rainer Martin. A frequencydomain adaptive line enhancer with step-size control based on mutual information for harmonic noise reduction. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(6): 1140–1154, June 2016. CODEN ???? ISSN 2329-9290.

Trinh:2021:DCL

[TM21] Viet Anh Trinh and Michael Mandel. Directly comparing the listening strategies of humans and machines. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):312–323, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3040545.

Tu:2022:AFL

[TM22] Youzhi Tu and Man-Wai Mak. Aggregating frame-level information in the spectral domain with self-attention for speaker embedding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):944–957, 2022. CO-DEN ???? ISSN 2329-9290.

Tsao:2014:MBO

[TMH^{+}14] Yu Tsao, S. Matsuda, C. Hori, H. Kashioka, and Chin-Hui Lee. A MAP-based online estimation approach to ensemble speaker and speaking environment modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(2):403–416, February 2014. CODEN ???? ISSN 2329-9290.

Trowitzsch:2017:RDE

[TMKO17] Ivo Trowitzsch, Johannes Mohr, Youssef Kashef, and Klaus Obermayer. Robust detection of environmental sounds in binaural auditory scenes. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (6):1344–1356, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7933059/.

Tan:2018:DBS

[TMM18] Z. Tan, M. Mak, and B. K. Mak. DNN-based score calibration with multitask learning for noise robust speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (4):700–712, April 2018. CODEN ???? ISSN 2329-9290.

Tan:2018:DSV

[TMMZ18] Z. Tan, M. Mak, B. K. Mak, and Y. Zhu. Denoised Senone I-vectors for robust speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(4):820– 830, April 2018. CODEN ???? ISSN 2329-9290.

Tran:2018:TMH

[TNS^{+}18] L. T. T. Tran, S. E.

Nordholm, H. Schepker, H. H. Dam, and S. Doclo. Two-microphone hearing aids using prediction error method for adaptive feedback control. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(5): 909–923, May 2018. CO-DEN ???? ISSN 2329-9290.

Tachibana:2014:HPS

[TOKS14] H. Tachibana, N. Ono, H. Kameoka, and S. Sagayama. Harmonic/percussive sound separation based on anisotropic smoothness of spectrograms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2059– 2073, December 2014. CODEN ???? ISSN 2329-9290.

Terenzi:2022:CFE

[TON^{+}22] Alessandro Terenzi, Nicola Ortolani, Inês Nolasco, Emmanouil Benetos, and Stefania Cecchi. Comparison of feature extraction methods for sound-based classification of honey bee activity. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):112– 122, 2022. CODEN ???? ISSN 2329-9290.

Tachibana:2014:SVE

[TOS14] H. Tachibana, N. Ono, and S. Sagayama. Singing voice enhancement in monaural music signals based on two-stage harmonic/percussive sound separation on multiple resolution spectrograms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):228– 237, January 2014. CO-DEN ???? ISSN 2329-9290.

Tervo:2015:DAE

[TP15] S. Tervo and A. Politis. Direction of arrival estimation of reflections from room impulse responses using a spherical microphone array. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(10):1539–1551, October 2015. CODEN ???? ISSN 2329-9290.

Torbati:2016:DHD

[TP16] A. H. Harati Nejad Torbati and J. Picone. A doubly hierarchical Dirichlet process hidden Markov model with a non-ergodic structure. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(1):174– 184, January 2016. CO-DEN ???? ISSN 2329-9290.

Tiwari:2019:SEU

[TP19] N. Tiwari and P. C.

Pandey. Speech enhancement using noise estimation with dynamic quantile tracking. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2301– 2312, December 2019. CODEN ???? ISSN 2329-9290.

Turchet:2022:MER

[TP22] Luca Turchet and Johan Pauwels. Music emotion recognition: Intention of composers-performers versus perception of musicians, non-musicians, and listening machines. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):305–316, 2022. CO-DEN ???? ISSN 2329-9290.

Tripathi:2023:DDN

[TP23] Achyut Mani Tripathi and Om Jee Pandey. Divide and distill: New outlooks on knowledge distillation for environmental sound classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1100– 1113, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3244507.

Tan:2018:AVD

[TQH^{+}18] T. Tan, Y. Qian, H. Hu, Y. Zhou, W. Ding, and K. Yu. Adaptive very deep convolutional residual network for noise robust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(8):1393–1405, August 2018. CODEN ???? ISSN 2329-9290.

Tian:2023:SSS

[TQM^{+}23] Hui Tian, Yiqin Qiu, Wojciech Mazurczyk, Haizhou Li, and Zhenxing Qian. STFF-SM: Steganalysis model based on spatial and temporal feature fusion for speech streams. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):277– 289, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3224295.

Tang:2015:JSC

[TQW^{+}15] Duyu Tang, Bing Qin, Furu Wei, Li Dong, Ting Liu, and Ming Zhou. A joint segmentation and classification framework for sentence level sentiment classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1750–1761, November 2015. CODEN ???? ISSN 2329-9290.

Tan:2016:CAT

[TQY16] Tian Tan, Yanmin Qian, and Kai Yu. Cluster adaptive training for deep neural network based acoustic model. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(3): 459–468, March 2016. CODEN ???? ISSN 2329-9290.

Tomar:2014:FDM

[TR14a] V. S. Tomar and R. C.

Rose. A family of discriminative manifold learning algorithms and their application to speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(1):161– 171, January 2014. CO-DEN ???? ISSN 2329-9290.

Tourbabin:2014:TFO

[TR14b] V. Tourbabin and B. Rafaely.

Theoretical framework for the optimization of microphone array configuration for humanoid robot audition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1803–1814, December 2014. CODEN ???? ISSN 2329-9290.

Tourbabin:2015:DAE

[TR15] V. Tourbabin and B. Rafaely. Direction of arrival estimation using microphone array processing for moving humanoid robots. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):2046–2058, November 2015. CODEN ???? ISSN 2329-9290.

Traa:2014:MSS

[TS14] J. Traa and P. Smaragdis. Multichannel source separation and tracking with RANSAC and directional statistics. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2233– 2243, December 2014. CODEN ???? ISSN 2329-9290.

Tsai:2016:REM

[TS16] T. J. Tsai and Andreas Stolcke. Robust and efficient multiple alignment of unsynchronized meeting recordings. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(5): 833–845, May 2016. CO-DEN ???? ISSN 2329-9290.

Trowitzsch:2020:JSE

[TSKO20] Ivo Trowitzsch, Christopher Schymura, Dorothea Kolossa, and Klaus Obermayer. Joining sound event detection and localization through spatial segregation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 487–502, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2958408.

Takamichi:2016:PMM

[TTB^{+}16] Shinnosuke Takamichi, Tomoki Toda, Alan W. Black, Graham Neubig, Sakriani Sakti, and Satoshi Nakamura. Postfilters to modify the modulation spectrum for statistical parametric speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(4):755– 767, April 2016. CODEN ???? ISSN 2329-9290.

Thiergart:2014:IPS

[TTH14] O. Thiergart, M. Taseska, and E. A. P. Habets. An informed parametric spatial filter based on instantaneous direction-of-arrival estimates. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2182–2196, December 2014. CODEN ???? ISSN 2329-9290.

Tran:2015:NUE

[TVJ15] D. T. Tran, E. Vincent, and D. Jouvet. Nonparametric uncertainty estimation and propagation for noise robust ASR. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1835–1846, November 2015. CODEN ???? ISSN 2329-9290.

Tan:2020:LCS

[TW20] Ke Tan and DeLiang Wang. Learning complex spectral mapping with gated convolutional recurrent networks for monaural speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):380–390, January 2020. CODEN ???? ISSN 2329-9290.

                                        URL https://dl.acm.
                                        org/doi/abs/10.1109/
                                        TASLP.2019.2955276.

Tan:2021:TMC

[TW21] Ke Tan and DeLiang Wang. Towards model compression for deep learning based speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1785– 1794, 2021. CODEN ???? ISSN 2329-9290.

Tang:2018:PTN

[TWC^{+}18] Zhiyuan Tang, Dong Wang, Yixiang Chen, Lantian Li, and Andrew Abel. Phonetic temporal neural model for language identification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (1):134–144, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8070977/.

Thien:2023:IFP

[TWIN23] Nguyen Binh Thien, Yukoh Wakabayashi, Kenta Iwai, and Takanobu Nishiura. Inter-frequency phase difference for phase reconstruction using deep neural networks and maximum likelihood. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1667–1680, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3268577.

Traa:2016:RSL

[TWSS16] Johannes Traa, David Wingate, Noah D. Stein, and Paris Smaragdis. Robust source localization and enhancement with a probabilistic steered response power model.

IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (3):493–503, March 2016. CODEN ???? ISSN 2329-9290.

Tan:2022:NSF

[TWW22] Ke Tan, Zhong-Qiu Wang, and DeLiang Wang. Neural spectrospatial filtering. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):605–621, 2022. CO-DEN ???? ISSN 2329-9290.

Tan:2018:CAA

[TWZ^{+}18] C. Tan, F. Wei, Q. Zhou, N. Yang, B. Du, W. Lv, and M. Zhou. Contextaware answer sentence selection with hierarchical gated recurrent neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(3): 540–549, March 2018. CODEN ???? ISSN 2329-9290.

Tao:2023:NUA

[TXH^{+}23] Qian Tao, Zhihao Xiong, Bocheng Han, Xiaoyang Fan, and Lusi Li. A novel unsupervised approach for cross-lingual word alignment in low isomorphic embedding spaces. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3027– 3041, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3301208.

Tian:2023:ILF

[TYW^{+}23] Jinchuan Tian, Jianwei Yu, Chao Weng, Yuexian Zou, and Dong Yu. Integrating lattice-free MMI into end-to-end speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):25– 38, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3198555.

Tan:2022:DMB

[TYZ^{+}22] Zhixing Tan, Zeyuan Yang, Meng Zhang, Qun Liu, Maosong Sun, and Yang Liu. Dynamic multi-branch layers for on-device neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):958– 967, 2022. CODEN ???? ISSN 2329-9290.

Talagala:2014:EMC

[TZA14] D. S. Talagala, Wen Zhang, and T. D. Abhayapala. Efficient multichannel adaptive room compensation for spatial soundfield reproduction using a modal decomposition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(10):1522–1532, October 2014. CODEN ???? ISSN 2329-9290.

Tong:2023:QEB

[TZL^{+}23] Chuxuan Tong, Xi Zheng, Jianhua Li, Xingjun Ma, Longxiang Gao, and Yong Xiang. Query-efficient black-Box adversarial attacks on automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3981– 3992, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3304476.

Tang:2021:MGS

[TZS^{+}21] Jian Tang, Jie Zhang, Yan Song, Ian McLoughlin, and Li-Rong Dai. Multigranularity sequence alignment mapping for encoderdecoder based end-to-end ASR. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2816–2828, 2021. CODEN ???? ISSN 2329-9290.

Tan:2021:DLB

[TZW21] Ke Tan, Xueliang Zhang, and DeLiang Wang. Deep learning based real-time speech enhancement for dual-microphone mobile phones. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1853– 1863, 2021. CODEN ???? ISSN 2329-9290.

Tong:2015:RTF

[TZZ^{+}15a] Renjie Tong, Yingyue Zhou, Long Zhang, Guangzhao Bao, and Zhongfu Ye. A robust time-frequency decomposition model for suppression of mixed Gaussian-impulse noise in audio signals. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(1): 69–79, January 2015. CO-DEN ???? ISSN 2329-9290.

Tu:2015:EDF

[TZZ15b] Mei Tu, Yu Zhou, and Chengqing Zong. Exploring diverse features for statistical machine translation model pruning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1847–1857, November 2015. CODEN ???? ISSN 2329-9290.

Ueno:2019:TDS

[UKS19] N. Ueno, S. Koyama, and H. Saruwatari. Threedimensional sound field reproduction based on weighted mode-matching method. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):1852– 1867, December 2019. CODEN ???? ISSN 2329-9290.

Ueda:2024:BSR

[UNI^{+}24] Tetsuya Ueda, Tomohiro Nakatani, Rintaro Ikeshita, Keisuke Kinoshita, Shoko Araki, and Shoji Makino. Blind and spatially-regularized online joint optimization of source separation, dereverberation, and noise reduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1157– 1172, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3351353.

Varanasi:2019:NFA

[VAH19] V. Varanasi, A. Agarwal, and R. M. Hegde. Near-field acoustic source localization using spherical harmonic features. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2054–2066, December 2019. CODEN ???? ISSN 2329-9290.

Vikram:2019:DNV

[VAP19] C. M. Vikram, N. Adiga, and S. R. M. Prasanna. Detection of nasalized voiced stops in cleft palate speech using epochsynchronous features. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(7): 1189–1200, July 2019. CODEN ???? ISSN 2329-9290.

Valentini-Botinhao:2018:SEN

[VBY18] C. Valentini-Botinhao and J. Yamagishi. Speech enhancement of noisy and reverberant speech for text-to-speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(8): 1420–1433, August 2018. CODEN ???? ISSN 2329-9290.

Vairetti:2017:SAP

[VDC^{+}17] Giacomo Vairetti, Enzo De Sena, Michael Catrysse, Søren Holdt Jensen, Marc Moonen, and Toon van Waterschoot. A scalable algorithm for physically motivated and sparse approximation of room impulse responses with orthonormal basis functions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (7):1547–1561, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7918506/.

Varzandeh:2024:SAB

[VDH24] Reza Varzandeh, Simon Doclo, and Volker Hohmann. Speech-aware binaural DOA estimation utilizing periodicity and spatial features in convolutional neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1198– 1213, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3356987.

Vilkamo:2015:PRS

[VDM15] J. Vilkamo and S. Delikaris-Manias. Perceptual reproduction of spatial sound using loudspeaker-signal-domain parametrization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (10):1660–1669, October 2015. CODEN ???? ISSN 2329-9290.

Vered:2024:PAD

[VE24] Yoav Vered and Stephen Elliott. A parallel analog and digital adaptive feedforward controller for active noise control. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1100– 1108, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3350882.

Valero:2019:LCM

[VH19] M. Luis Valero and E. A. P. Habets. Lowcomplexity multi-microphone acoustic echo control in the short-time Fourier transform domain. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(3): 595–609, March 2019. CODEN ???? ISSN 2329-9290.

Vargas:2021:ITC

[VHBS21] Elizabeth Vargas, James R.

Hopgood, Keith Brown, and Kartic Subr. On improved training of CNN for acoustic source localisation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):720– 732, 2021. CODEN ???? ISSN 2329-9290.

vanHengel:2014:CST

[vHK14] P. W. J. van Hengel and J. D. Krijnders. A comparison of spectrotemporal representations of audio signals. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): 303–313, February 2014. CODEN ???? ISSN 2329-9290.

VanKuyk:2018:EII

[VKH18] S. Van Kuyk, W. B.

Kleijn, and R. C. Hendriks. An evaluation of intrusive instrumental intelligibility metrics. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):2153–2166, November 2018. CODEN ???? ISSN 2329-9290.

Vijayan:2015:APS

[VM15] K. Vijayan and K. S. R. Murty. Analysis of phase spectrum of speech signals using allpass modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2371–2383, December 2015. CODEN ???? ISSN 2329-9290.

Vaizman:2014:CBA

[VML14] Y. Vaizman, B. McFee, and G. Lanckriet. Codebookbased audio feature representation for music information retrieval. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(10): 1483–1493, October 2014. CODEN ???? ISSN 2329-9290.

vanMourik:2014:EHO

[vMM14] J. van Mourik and D. Murphy. Explicit higher-order FDTD schemes for 3D room acoustic simulation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2003–2011, December 2014. CODEN ???? ISSN 2329-9290.

Villalba:2016:BNM

[VMOL16] Jesús Villalba, Antonio Miguel, Alfonso Ortega, and Eduardo Lleida. Bayesian networks to model the variability of speaker verification scores in adverse environments. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2327– 2340, ???? 2016. CODEN ???? ISSN 2329-9290.

vonNeumann:2023:SLC

[vNKB^{+}23] Thilo von Neumann, Keisuke Kinoshita, Christoph Boeddeker, Marc Delcroix, and Reinhold Haeb-Umbach. Segmentless continuous speech separation of meetings: Training and evaluation criteria. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):576– 589, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3228629.

Vikram:2017:EET

[VP17] C. M. Vikram and S. R. Mahadeva Prasanna. Epoch extraction from telephone quality speech using single pole filter. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(3): 624–636, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Valimaki:2021:LRS

[VP21] Vesa Välimäki and Karolina Prawda. Late-reverberation synthesis using interleaved velvet-noise sequences. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1149– 1160, 2021. CODEN ???? ISSN 2329-9290.

Vlimki:2019:NCG

[VR19] V. Välimäki and J. Rämö. Neurally controlled graphic equalizer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2140– 2149, December 2019. CODEN ???? ISSN 2329-9290.

Vaz:2018:ADU

[VRN18] C. Vaz, V. Ramanarayanan, and S. Narayanan. Acoustic denoising using dictionary learning with spectral and temporal regularization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(5):967– 980, May 2018. CODEN ???? ISSN 2329-9290.

Venkitaraman:2014:BSP

[VS14] A. Venkitaraman and C. S. Seelamantula. Binaural signal processing motivated generalized analytic signal construction and AM–FM demodulation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(6):1023–1036, June 2014. CODEN ???? ISSN 2329-9290.

Varela:2024:ERS

[VSK^{+}24] Alejandro Santorum Varela, Svetlana Stoyanchev, Simon Keizer, Rama Doddipatla, and Kate Knill. Entity resolution in situated dialog with unimodal and multimodal transformers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):704– 713, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3304468.

Vukovic:2021:CLE

[VSL21] Maria Vukovic, Melissa Stolar, and Margaret Lech. Cognitive load estimation from speech commands to simulated aircraft. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1011–1022, 2021. CODEN ???? ISSN 2329-9290.

VanSegbroeck:2015:RLI

[VTN15] M. Van Segbroeck, R. Travadi, and S. S. Narayanan. Rapid language identification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(7):1118–1129, July 2015. CODEN ???? ISSN 2329-9290.

Venturini:2014:SFF

[VZC14] A. Venturini, L. Zao, and R. Coelho. On speech features fusion, α-integration Gaussian modeling and multi-style training for noise robust speaker classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1951– 1964, December 2014. CODEN ???? ISSN 2329-9290.

Wu:2015:JPD

[WAJL15] Yuntao Wu, L. Amir, J. R. Jensen, and Guisheng Liao. Joint pitch and DOA estimation using the ESPRIT method. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(1): 32–45, January 2015. CO-DEN ???? ISSN 2329-9290.

Wachowski:2014:DCN

[WAS14] N. Wachowski and M. R. Azimi-Sadjadi. Detection and classification of nonstationary transient signals using sparse approximations and Bayesian networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1750– 1764, December 2014. CODEN ???? ISSN 2329-9290.

Winter:2016:AMD

[WAS16] Fiete Winter, Jens Ahrens, and Sascha Spors. On analytic methods for 2.5-D local sound field synthesis using circular distributions of secondary sources. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (5):914–926, May 2016. CODEN ???? ISSN 2329-9290.

Wang:2018:MBS

[WB18] Y. Wang and M. Brookes.

Model-based speech enhancement in the modulation domain. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(3): 580–594, March 2018. CODEN ???? ISSN 2329-9290.

Wang:2021:BCM

[WBC^{+}21] Xuehan Wang, Jacob Benesty, Jingdong Chen, Gongping Huang, and Israel Cohen. Beamforming with cube microphone arrays via Kronecker product decompositions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1774– 1784, 2021. CODEN ???? ISSN 2329-9290.

Wang:2022:AST

[WBLC22] Changhong Wang, Emmanouil Benetos, Vincent Lostanlen, and Elaine Chew. Adaptive scattering transforms for playing technique recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1407– 1421, 2022. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3156785.

Weisz:2018:SED

[WBSG18] G. Weisz, P. Budzianowski, P. Su, and M. Gašić. Sample efficient deep reinforcement learning for dialogue systems with large action spaces. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(11): 2083–2097, November 2018. CODEN ???? ISSN 2329-9290.

Wu:2016:DFF

[WC16a] Xiaoguang Wu and Huawei Chen. Directivity factors of the first-order steerable differential array with microphone mismatches: Deterministic and worst-case analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (2):300–315, ???? 2016. CODEN ???? ISSN 2329-9290.

Wu:2016:GAA

[WC16b] Yi-Chan Wu and Homer H.

Chen. Generation of affective accompaniment in accordance with emotion flow. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2277–2287, ???? 2016. CODEN ???? ISSN 2329-9290.

Wang:2018:SSS

[WC18a] D. Wang and J. Chen.

Supervised speech separation based on deep learning: an overview. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):1702–1726, October 2018. CODEN ???? ISSN 2329-9290.

Wang:2018:PDB

[WC18b] L. Wang and A. Cavallaro.

Pseudo-determined blind source separation for adhoc microphone networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (5):981–994, May 2018. CODEN ???? ISSN 2329-9290.

Wang:2015:SEV

[WCC^{+}15] Jia-Ching Wang, Yu-Hao Chin, Bo-Wei Chen, Chang-Hong Lin, and Chung-Hsien Wu. Speech emotion verification using emotion variance modeling and discriminant scale-frequency maps. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(10): 1552–1562, October 2015. CODEN ???? ISSN 2329-9290.

Wang:2021:MIN

[WCC^{+}21] Xianghui Wang, Jie Chen, Xiaoyi Chen, Jing Guo, and Qian Xiang. Multichannel iterative noise reduction filters in the short-time-fourier-transform domain based on Kronecker product decomposition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2725–2740, 2021. CODEN ???? ISSN 2329-9290.

Wang:2023:TGI

[WCC^{+}23] Zhong-Qiu Wang, Samuele Cornell, Shukjae Choi, Younglo Lee, Byeong-Yeol Kim, and Shinji Watanabe. TF-GridNet: Integrating full- and subband modeling for speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3221– 3236, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3304482.

Wang:2019:RHD

[WCCB19] X. Wang, I. Cohen, J. Chen, and J. Benesty. On robust and high directive beamforming with small-spacing microphone arrays for scattered sources. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(4): 842–852, April 2019. CO-DEN ???? ISSN 2329-9290.

Wu:2021:EBE

[WCL^{+}21a] Xixin Wu, Yuewen Cao, Hui Lu, Songxiang Liu, Shiyin Kang, Zhiyong Wu, Xunying Liu, and Helen Meng. Exemplar-based emotive speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):874–886, 2021. CO-DEN ???? ISSN 2329-9290.

Wu:2021:SER

[WCL^{+}21b] Xixin Wu, Yuewen Cao, Hui Lu, Songxiang Liu, Disong Wang, Zhiyong Wu, Xunying Liu, and Helen Meng. Speech emotion recognition using sequential capsule networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3280– 3291, 2021. CODEN ???? ISSN 2329-9290.

Wu:2023:GNL

[WCL^{+}23] Xin Wu, Yi Cai, Zetao Lian, Ho fung Leung, and Tao Wang. Generating natural language from logic expressions with structural representation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1499– 1510, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3263784.

Wang:2021:SSU

[WCW^{+}21] Peidong Wang, Zhuo Chen, DeLiang Wang, Jinyu Li, and Yifan Gong. Speaker separation using speaker inventories and estimated speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):537–546, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3045556.

Wang:2015:NHD

[WCY15] Lin Wang, Zhe Chen, and Fuliang Yin. A novel hierarchical decomposition vector quantization method for high-order LPC parameters. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(1): 212–221, January 2015.

CODEN ???? ISSN 2329-9290.

Wang:2019:DBT

[WCY19] R. Wang, Z. Chen, and F. Yin. DOA-based threedimensional node geometry calibration in acoustic sensor networks and its Cramér–Rao bound and sensitivity analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (9):1455–1468, September 2019. CODEN ???? ISSN 2329-9290.

Wang:2021:ISG

[WCZ21] Yile Wang, Leyang Cui, and Yue Zhang. Improving skip-gram embeddings using BERT. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1318–1328, 2021. CODEN ???? ISSN 2329-9290.

Wang:2023:EEM

[WCZR23] Mou Wang, Junqi Chen, Xiao-Lei Zhang, and Susanto Rahardja. End-to-end multi-modal speech recognition on an air and bone conducted speech corpus. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):513– 524, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3224305.

Wang:2016:CMB

[WD16] Lin Wang and Simon Doclo. Correlation maximizationbased sampling rate offset estimation for distributed microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(3):571–582, March 2016. CODEN ???? ISSN 2329-9290.

Wu:2016:AST

[WDD^{+}16] Zhizheng Wu, Phillip L. De Leon, Cenk Demiroglu, Ali Khodabakhsh, Simon King, Zhen-Hua Ling, Daisuke Saito, Bryan Stewart, Tomoki Toda, Mirjam Wester, and Junichi Yamagishi. Anti-spoofing for textindependent speaker verification: an initial database, comparison of countermeasures, and human performance. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(4): 768–783, April 2016. CO-DEN ???? ISSN 2329-9290.

Wang:2017:GMD

[WDDL17] Yannan Wang, Jun Du, Li-Rong Dai, and Chin-Hui Lee. A gender mixture detection approach to unsupervised singlechannel speech separation based on deep neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(7): 1535–1546, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7917323/.

Wang:2018:MLE

[WDDL18] Q. Wang, J. Du, L. Dai, and C. Lee. A multiobjective learning and ensembling approach to highperformance speech enhancement with compact neural network architectures. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(7):1185–1197, July 2018. CODEN ???? ISSN 2329-9290.

Wu:2018:RAD

[WDS^{+}18] C. Wu, C. Dittmar, C. Southall, R. Vogl, G. Widmer, J. Hockman, M. Müller, and A. Lerch. A review of automatic drum transcription. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1457–1483, September 2018. CODEN ???? ISSN 2329-9290.

Wang:2022:AAR

[WDW^{+}22] Chengyu Wang, Suyang Dai, Yipeng Wang, Fei Yang, Minghui Qiu, Kehan Chen, Wei Zhou, and Jun Huang. ARoBERT: an ASR robust pretrained language model for spoken language understanding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1207– 1218, 2022. CODEN ???? ISSN 2329-9290.

Wang:2023:FSD

[WDW^{+}23] Qing Wang, Jun Du, Hua-Xin Wu, Jia Pan, Feng Ma, and Chin-Hui Lee. A four-stage data augmentation approach to ResNet-Conformer based acoustic modeling for sound event localization and detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1251– 1264, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3256088.

Wang:2016:REJ

[WED16] Siying Wang, Sebastian Ewert, and Simon Dixon. Robust and efficient joint alignment of multiple musical performances. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):2132– 2145, ???? 2016. CODEN ???? ISSN 2329-9290.

Wang:2017:IME

[WED17] Siying Wang, Sebastian Ewert, and Simon Dixon. Identifying missing and extra notes in piano recordings using scoreinformed dictionary learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(10):1877–1889, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7971931/.

Wakabayashi:2018:SCS

[WFN^{+}18] Y. Wakabayashi, T. Fukumori, M. Nakayama, T. Nishiura, and Y. Yamashita. Single-channel speech enhancement with phase reconstruction based on phase distortion averaging. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(9):1559–1569, September 2018. CODEN ???? ISSN 2329-9290.

Wu:2021:HOP

[WFR^{+}21] Shengqiong Wu, Hao Fei, Yafeng Ren, Bobo Li, Fei Li, and Donghong Ji. High-order pairwise aspect and opinion terms extraction with edge-enhanced syntactic graph convolution. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2396–2406, 2021. CODEN ???? ISSN 2329-9290.

Wang:2022:NND

[WFY^{+}22] Tao Wang, Ruibo Fu, Jiangyan Yi, Jianhua Tao, and Zhengqi Wen. NeuralDPS: Neural deterministic plus stochastic model with multiband excitation for noisecontrollable waveform generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):865– 878, 2022. CODEN ???? ISSN 2329-9290.

Wang:2015:NPS

[WGD15] Lin Wang, T. Gerkmann, and S. Doclo. Noise power spectral density estimation using MaxNSR blocking matrix. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(9):1493– 1508, September 2015. CODEN ???? ISSN 2329-9290.

Wang:2021:MVM

[WGLZ21] Jianyu Wang, Shanzheng Guan, Shupei Liu, and Xiao-Lei Zhang. Minimumvolume multichannel nonnegative matrix factorization for blind audio source separation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 3089–3103, 2021. CODEN ???? ISSN 2329-9290.

Wu:2018:IIR

[WGR^{+}18] Chunyang Wu, Mark J. F. Gales, Anton Ragni, Penny Karanasou, and Khe Chai Sim. Improving interpretability and regularization in deep learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(2):256–265, ???? 2018. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8114209/.

Wang:2022:MSD

[WH22] Zhenyu Wang and John H. L. Hansen. Multisource domain adaptation for text-independent forensic speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):60–75, 2022. CO-DEN ???? ISSN 2329-9290.

Westhausen:2021:RSL

[WHB^{+}21] Nils L. Westhausen, Rainer Huber, Hannah Baumgartner, Ragini Sinha, Jan Rennies, and Bernd T. Meyer. Reduction of subjective listening effort for TV broadcast signals with recurrent neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??):

3541–3550, 2021. CODEN ???? ISSN 2329-9290.

Wu:2014:SSS

[WHLG14] Chung-Hsien Wu, Yi-Chin Huang, Chung-Han Lee, and Jun-Cheng Guo. Synthesis of spontaneous speech with syllable contraction using state-based context-dependent voice transformation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(3): 585–595, March 2014. CODEN ???? ISSN 2329-9290.

Wu:2021:QPP

[WHO^{+}21] Yi-Chiao Wu, Tomoki Hayashi, Takuma Okamoto, Hisashi Kawai, and Tomoki Toda. Quasi-periodic parallel WaveGAN: a nonautoregressive raw waveform generative model with pitch-dependent dilated convolution neural network. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):792– 806, 2021. CODEN ???? ISSN 2329-9290.

Wang:2019:DNE

[WHQY19] S. Wang, Z. Huang, Y. Qian, and K. Yu. Discriminative neural embedding learning for shortduration text-independent speaker verification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(11): 1686–1696, November 2019. CODEN ???? ISSN 2329-9290.

Wang:2016:IAS

[WHRC16] Lin Wang, Tsz-Kin Hon, Joshua D. Reiss, and Andrea Cavallaro. An iterative approach to source counting and localization using two distant microphones. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(6):1079– 1093, June 2016. CODEN ???? ISSN 2329-9290.

Wu:2021:QPW

[WHT^{+}21] Yi-Chiao Wu, Tomoki Hayashi, Patrick Lumban Tobing, Kazuhiro Kobayashi, and Tomoki Toda. Quasi-periodic WaveNet: an autoregressive raw waveform generative model with pitch-dependent dilated convolution neural network. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1134–1148, 2021. CODEN ???? ISSN 2329-9290.

Widmark:2018:CIA

[Wid18] S. Widmark. Causal IIR audio precompensator filters subject to quadratic constraints. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(10):1897– 1912, October 2018. CO-DEN ???? ISSN 2329-9290.

Widmark:2019:CMO

[Wid19]  S. Widmark.  Causal

MSE-optimal filters for personal audio subject to constrained contrast. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (5):972–987, May 2019. CODEN ???? ISSN 2329-9290.

Wang:2015:IQS

[WJ15a] Chung-Che Wang and J.-S. R. Jang. Improving query-by-singing/ humming by combining melody and lyric information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(4):798–806, April 2015. CODEN ???? ISSN 2329-9290.

Weng:2015:DTU

[WJ15b] Chao Weng and B.- H. F. Juang. Discriminative training using nonuniform criteria for keyword spotting on spontaneous speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): 300–312, February 2015. CODEN ???? ISSN 2329-9290.

Wang:2023:TST

[WJ23] Jun-You Wang and Jyh-Shing Roger Jang. Training a singing transcription model using connectionist temporal classification loss and crossentropy loss. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 383–396, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3224297.

Wu:2016:ITM

[WK16] Zhizheng Wu and Simon King. Improving trajectory modelling for DNNbased speech synthesis by using stacked bottleneck features and minimum generation error training. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (7):1255–1265, July 2016. CODEN ???? ISSN 2329-9290.

Wilkinghoff:2024:WDA

[WK24a] Kevin Wilkinghoff and Frank Kurth. Why do angular margin losses work well for semi-supervised anomalous sound detection? IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):608– 622, 2024. ISSN 2329-

9290. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2023.3337153.

Wu:2024:SAS

[WK24b] Jin Chu Wu and Raghu N. Kacker. Statistical analysis for speaker recognition evaluation with data dependence and three score distributions. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 1–14, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3313447.

Wu:2019:AAC

[WL19] Y. Wu and W. Li. Automatic audio chord recognition with MIDI-trained deep feature and BLSTM-CRF sequence decoding model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(2):355– 366, February 2019. CO-DEN ???? ISSN 2329-9290.

Wu:2023:PFH

[WLF23] Guowei Wu, Shipei Liu, and Xiaoya Fan. The power of fragmentation: a hierarchical transformer model for structural segmentation in symbolic music generation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1409–1420, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3263797.

Wang:2015:ASM

[WLL^{+}15a] Haipeng Wang, Tan Lee, Cheung-Chi Leung, Bin Ma, and Haizhou Li. Acoustic segment modeling with spectral clustering methods. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(2): 264–277, February 2015. CODEN ???? ISSN 2329-9290.

Wu:2015:PFR

[WLL15b] Ji Wu, Miao Li, and Chin-Hui Lee. A probabilistic framework for representing dialog systems and entropy-based dialog management through dynamic stochastic state evolution. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):2026– 2035, November 2015. CODEN ???? ISSN 2329-9290.

Wang:2016:CSB

[WLL^{+}16] Jia-Ching Wang, Yuan-Shan Lee, Chang-Hong Lin, Shu-Fan Wang, Chih-Hao Shih, and Chung-Hsien Wu. Compressive sensing-based speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):2122– 2131, ???? 2016. CODEN ???? ISSN 2329-9290.

Wu:2022:IAR

[WLL^{+}22] Haibin Wu, Xu Li, Andy T. Liu, Zhiyong Wu, Helen Meng, and Hung-Yi Lee. Improving the adversarial robustness for speaker verification by self-supervised learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):202–217, 2022. CO-DEN ???? ISSN 2329-9290.

Wu:2023:HSM

[WLL^{+}23] Xixin Wu, Hui Lu, Kun Li, Zhiyong Wu, Xunying Liu, and Helen Meng. Hiformer: Sequence modeling networks with hierarchical attention mechanisms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3993– 4003, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3313428.

Wang:2017:EAC

[WLLZ17] Zhongqing Wang, Sophia Yat Mei Lee, Shoushan Li, and Guodong Zhou. Emotion analysis in codeswitching text with joint factor graph model. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(3): 469–480, ???? 2017. CO-DEN ???? ISSN 2329-9290.

Wang:2018:SSWa

[WLT^{+}18] S. Wang, P. Lin, Y. Tsao, J. Hung, and B. Su. Suppression by selecting wavelets for feature compression in distributed speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(3):564–579, March 2018. CODEN ???? ISSN 2329-9290.

Wang:2024:DCT

[WLT24] Rui Wang, Li Li, and Tomoki Toda. Dualchannel target speaker extraction based on conditional variational autoencoder and directional information. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1968– 1979, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3376154.

Wang:2021:MIE

[WLZH21] Run-Ze Wang, Zhen-Hua Ling, Jing-Bo Zhou, and Yu Hu. A multipleintegration encoder for multi-turn text-to-SQL semantic parsing. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1503–1513, 2021. CODEN ???? ISSN 2329-9290.

Wang:2014:TTS

[WNW14] Yuxuan Wang, A. Narayanan, and Deliang Wang. On training targets for supervised speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1849–1858, December 2014. CODEN ???? ISSN 2329-9290.

Wang:2021:ABP

[WPY^{+}21] Weiqing Wang, Jin Pan, Hua Yi, Zhanmei Song, and Ming Li. Audiobased piano performance evaluation for beginners with convolutional neural network and attention mechanism. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1119–1133, 2021. CODEN ???? ISSN 2329-9290.

Wang:2024:NJT

[WQ24a] Shouhui Wang and Biao Qin. A novel joint training model for knowledge base question answering. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):666– 679, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3336526.

Wang:2024:UCL

[WQ24b] Wei Wang and Yanmin Qian. Universal crosslingual data generation for low resource ASR. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 973–983, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3345150.

Wang:2021:GIS

[WQZ^{+}21] Xinsheng Wang, Tingting Qiao, Jihua Zhu, Alan Hanjalic, and Odette Scharenborg. Generating images from spoken descriptions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):850– 865, 2021. CODEN ???? ISSN 2329-9290.

Wang:2016:DSS

[WRC16] Lin Wang, Joshua D.

Reiss, and Andrea Cavallaro. Over-determined source separation and localization using distributed microphones. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (9):1573–1588, ???? 2016. CODEN ???? ISSN 2329-9290.

Wood:2017:BSS

[WRDP17] Sean U. N. Wood, Jean Rouat, Stéphane Dupont, and Gueorgui Pironkov. Blind speech separation and enhancement with GCC-NMF. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(4):745– 755, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7828077/.

Wu:2018:MDE

[WRK18] K. Wu, V. G. Reju, and A. W. H. Khong. Multisource DOA estimation in a reverberant environment using a single acoustic vector sensor. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):1848–1859, October 2018. CODEN ???? ISSN 2329-9290.

Wu:2017:SIB

[WRKG17] Kai Wu, Vaninirappuputhenpurayil Gopalan Reju, Andy W. H. Khong, and Shu Ting Goh. Swarm intelligence based particle filter for alternating talker localization and tracking using microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (6):1384–1397, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7898417/.

Wang:2014:RBC

[WS14] Guangsen Wang and Khe Chai Sim. Regressionbased context-dependent modeling of deep neural networks for speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(11):1660– 1669, November 2014. CODEN ???? ISSN 2329-9290.

Winter:2019:GMP

[WSFS19] F. Winter, F. Schultz, G. Firtha, and S. Spors. A geometric model for prediction of spatial aliasing in 2.5D sound field synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(6):1031– 1046, June 2019. CODEN ???? ISSN 2329-9290.

Wu:2021:STS

[WSG21] Jingyi Wu, Lin Shang, and Xiaoying Gao. Sentiment time series calibration for event detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2407–2420, 2021. CODEN ???? ISSN 2329-9290.

Wu:2015:CSE

[WSH15] Chung-Hsien Wu, Han-Ping Shen, and Chun-Shan Hsu. Code-switching event detection by using a latent language space model and the delta-Bayesian information criterion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(11):1892– 1903, November 2015. CODEN ???? ISSN 2329-9290.

Wang:2023:DMP

[WSJ^{+}23] Ante Wang, Linfeng Song, Lifeng Jin, Junfeng Yao, Haitao Mi, Chen Lin, Jinsong Su, and Dong Yu. D^{2}PSG: Multiparty dialogue discourse parsing as sequence generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):4004– 4013, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313415.

Wang:2015:SDU

[WsL15] Yow-Bang Wang and Lin shan Lee. Supervised detection and unsupervised discovery of pronunciation error patterns for computer-assisted language learning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(3): 564–579, March 2015. CODEN ???? ISSN 2329-9290.

Wang:2020:NSD

[WSL^{+}20] Zhongqing Wang, Qingying Sun, Shoushan Li, Qiaoming Zhu, and Guodong Zhou. Neural stance detection with hierarchical linguistic representations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):635–645, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.2963954.

Wood:2019:BCB

[WSM19] S. U. N. Wood, J. K. W.

Stahl, and P. Mowlaee. Binaural codebook-based speech enhancement with atomic speech presence probability. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2150– 2161, December 2019. CODEN ???? ISSN 2329-9290.

Weisman:2021:RAR

[WST^{+}21] Ran Weisman, Tom Shlomo, Vladimir Tourbabin, Paul Calamia, and Boaz Rafaely. Robustness of acoustic rake filters in minimum variance beamforming. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3668–3678, 2021.

CODEN ???? ISSN 2329-9290.

Wu:2014:CEP

[WSY14] Chung-Hsien Wu, Han-Ping Shen, and Yan-Ting Yang. Chinese–English phone set construction for code-switching ASR using acoustic and DNNextracted articulatory features. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(4):858–862, April 2014. CODEN ???? ISSN 2329-9290.

Wu:2021:TPB

[WSZN21] Bin Wu, Sakriani Sakti, Jinsong Zhang, and Satoshi Nakamura. Tackling perception bias in unsupervised phoneme discovery using DPGMM-RNN hybrid model and functional load. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):348–362, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3042016.

Wu:2022:MUE

[WSZN22] Bin Wu, Sakriani Sakti, Jinsong Zhang, and Satoshi Nakamura. Modeling unsupervised empirical adaptation by DPGMM and DPGMM-RNN hybrid model to extract perceptual features for lowresource ASR. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 901–916, 2022. CODEN ???? ISSN 2329-9290.

Weng:2014:LSR

[WTHJ14] Chao Weng, D. L. Thomson, P. Haffner, and B.- H. F. Juang. Latent semantic rational kernels for topic spotting on conversational speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1738–1749, December 2014. CODEN ???? ISSN 2329-9290.

Wu:2019:BEP

[WTQ^{+}19] L. Wu, X. Tan, T. Qin, J. Lai, and T. Liu. Beyond error propagation: Language branching also affects the accuracy of sequence generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):1868–1879, December 2019. CODEN ???? ISSN 2329-9290.

Wang:2020:BGB

[WTW20] Peidong Wang, Ke Tan, and De Liang Wang. Bridging the gap between monaural speech enhancement and recognition with distortionindependent acoustic modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):39–48, January 2020. CODEN ???? ISSN 2329-9290.

                                        URL https://dl.acm.
                                        org/doi/abs/10.1109/
                                        TASLP.2019.2946789.

Wang:2018:ANF

[WTY18] X. Wang, S. Takaki, and J. Yamagishi. Autoregressive neural F0 model for statistical parametric speech synthesis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(8): 1406–1419, August 2018. CODEN ???? ISSN 2329-9290.

Wang:2020:NSF

[WTY20a] Xin Wang, Shinji Takaki, and Junichi Yamagishi. Neural source-filter waveform models for statistical parametric speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):402– 415, January 2020. CO-DEN ???? ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2019.2956145.

Wang:2020:VQV

[WTY^{+}20b] Xin Wang, Shinji Takaki, Junichi Yamagishi, Simon King, and Keiichi Tokuda. A vector quantized variational autoencoder (VQ-VAE) autoregressive neural F_{0} model for statistical parametric speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):157–170, January 2020. CODEN ???? ISSN 2329-9290.

URL https://dl.acm.
org/doi/abs/10.1109/
TASLP.2019.2950099.

Wang:2023:MVB

[WTY^{+}23] Xianke Wang, Bowen Tian, Weiming Yang, Wei Xu, and Wenqing Cheng. MusicYOLO: a visionbased framework for automatic singing transcription. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):229– 241, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3221005.

Wang:2018:ISM

[WTZ18] X. Wang, Z. Tu, and M. Zhang. Incorporating statistical machine translation word knowledge into neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2255– 2266, December 2018. CODEN ???? ISSN 2329-9290.

Wang:2018:SSWb

[WUF^{+}18] R. Wang, M. Utiyama, A. Finch, L. Liu, K. Chen, and E. Sumita. Sentence selection and weighting for neural machine translation domain adaptation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (10):1727–1741, October 2018. CODEN ???? ISSN 2329-9290.

Wu:2014:EBS

[WVCL14] Zhizheng Wu, T. Virtanen, Eng Siong Chng, and Haizhou Li. Exemplarbased sparse representation with residual compensation for voice conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(10):1506– 1521, October 2014. CO-DEN ???? ISSN 2329-9290.

Wang:2021:SSD

[WvdHZ^{+}21] Xinsheng Wang, Justin van der Hout, Jihua Zhu, Mark Hasegawa-Johnson, and Odette Scharenborg. Synthesizing spoken descriptions of images. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3242–3254, 2021. CODEN ???? ISSN 2329-9290.

Wei:2015:DLC

[WW15] Ying Wei and Yinfeng Wang. Design of low complexity adjustable filter bank for personalized hearing aid solutions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (5):923–931, May 2015. CODEN ???? ISSN 2329-9290.

Wang:2016:JTF

[WW16] Zhong-Qiu Wang and DeLiang Wang. A joint training framework for robust automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(4):796– 806, April 2016. CODEN ???? ISSN 2329-9290.

Williamson:2017:TFM

[WW17] Donald S. Williamson and DeLiang Wang. Timefrequency masking in the complex domain for speech dereverberation and denoising. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(7): 1492–1501, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7906509/.

Wang:2019:CSS

[WW19] Z. Wang and D. Wang.

Combining spectral and spatial features for deep learning based blind speaker separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(2):457– 468, February 2019. CO-DEN ???? ISSN 2329-9290.

Wang:2021:TRS

[WW21] Heming Wang and DeLiang Wang. Towards robust speech super-resolution. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):2058–2066, 2021. CODEN ???? ISSN 2329-9290.

Wang:2022:NCA

[WW22] Heming Wang and DeLiang Wang. Neural cascade architecture with tripledomain loss for speech enhancement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 734–743, 2022. CODEN ???? ISSN 2329-9290.

Wang:2014:ISC

[WWD14] Qi Wang, W. L. Woo, and S. S. Dlay. Informed single-channel speech separation using HMM– GMM user-generated exemplar source. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2087–2100, December 2014. CODEN ???? ISSN 2329-9290.

Wang:2023:DAD

[WWG^{+}23] Yixin Wang, Wei Wei, Xiangming Gu, Xiaohong Guan, and Ye Wang. Disentangled adversarial domain adaptation for phonation mode detection in singing and speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3746– 3759, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3317568.

Wang:2024:MLI

[WWG^{+}24] Jiapu Wang, Boyue Wang, Junbin Gao, Simin Hu, Yongli Hu, and Baocai Yin. Multi-level interaction based knowledge graph completion. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 386–396, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3331121.

Winter:2018:CLW

[WWH^{+}18] F. Winter, H. Wierstorf, C. Hold, F. Krüger, A. Raake, and S. Spors. Colouration in local wave field synthesis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(10): 1913–1924, October 2018.

CODEN ???? ISSN 2329-9290.

Wang:2021:CPM

[WWL21] Zhong-Qiu Wang, Gordon Wichern, and Jonathan Le Roux. Convolutive prediction for monaural speech dereverberation and noisy-reverberant speaker separation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 3476–3490, 2021. CODEN ???? ISSN 2329-9290.

Wei:2014:EPF

[WWLL14] Wen-Li Wei, Chung-Hsien Wu, Jen-Chun Lin, and Han Li. Exploiting psychological factors for interaction style recognition in spoken conversation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (3):659–671, March 2014. CODEN ???? ISSN 2329-9290.

Wang:2018:SER

[WWS^{+}18] C. Wang, J. Wang, A. Santoso, C. Chiang, and C. Wu. Sound event recognition using auditory-receptive-field binary pattern and hierarchical-diving deep belief network. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(8): 1336–1351, August 2018.

CODEN ???? ISSN 2329-9290.

Williamson:2016:CRM

[WWW16] Donald S. Williamson, Yuxuan Wang, and DeLiang Wang. Complex ratio masking for monaural speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (3):483–492, March 2016. CODEN ???? ISSN 2329-9290.

Wang:2021:MMC

[WWW21] Zhong-Qiu Wang, Peidong Wang, and DeLiang Wang. Multi-microphone complex spectral mapping for utterance-wise and continuous speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2001–2014, 2021. CODEN ???? ISSN 2329-9290.

Wang:2023:SDN

[WWWR23] Zhong-Qiu Wang, Gordon Wichern, Shinji Watanabe, and Jonathan Le Roux. STFT-Domain neural speech enhancement with very low algorithmic latency. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 397–410, 2023. ISSN 2329-

9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2022.3224285.

Wang:2023:MVH

[WWX^{+}23] Zhichao Wang, Xinsheng Wang, Qicong Xie, Tao Li, Lei Xie, Qiao Tian, and Yuping Wang. MSM-VC: High-fidelity source style transfer for nonparallel voice conversion by multi-scale style modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3883– 3895, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313414.

Wang:2024:USR

[WWX^{+}24] Siyuan Wang, Zhongyu Wei, Jiarong Xu, Taishan Li, and Zhihao Fan. Unifying structure reasoning and language pretraining for complex reasoning tasks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1586– 1595, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3325973.

Wu:2021:EFB

[WXL^{+}21] Wenyi Wu, Yegui Xiao, Jianhui Lin, Liying Ma, and Khashayar Khorasani. An efficient filter bank structure for adaptive notch filtering and applications. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 3226–3241, 2021. CODEN ???? ISSN 2329-9290.

Wu:2024:SAD

[WXS24] Han Wu, Kun Xu, and Linqi Song. Structureaware dialogue modeling methods for conversational semantic role labeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):742– 752, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3331576.

Wang:2019:SSN

[WXZ^{+}19] Y. Wang, Y. Xia, L. Zhao, J. Bian, T. Qin, E. Chen, and T. Liu. Semisupervised neural machine translation via marginal distribution estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (10):1564–1576, October 2019. CODEN ???? ISSN 2329-9290.

Wu:2023:MFS

[WY23] Shih-Lun Wu and Yi-Hsuan Yang. MuseMorphose: Full-song and finegrained piano music style transfer with one transformer VAE. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1953– 1967, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3270726.

Wang:2015:SBT

[WYH^{+}15] Xun Wang, Y. Yoshida, T. Hirao, M. Nagata, and K. Sudoh. Summarization based on taskoriented discourse parsing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):1358–1367, August 2015. CODEN ???? ISSN 2329-9290.

Wang:2017:RHF

[WYH17] Dongmei Wang, Chengzhu Yu, and John H. L. Hansen. Robust harmonic features for classificationbased pitch estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):952–964, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7851049/.

Wen:2020:PUA

[WYHL20] Hao-Xiang Wen, Sen-Quan Yang, Yuan-Quan Hong, and Huan Luo. A partial update adaptive algorithm for sparse system identification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 240–255, January 2020. CODEN ???? ISSN 2329-9290. URL https:

//dl.acm.org/doi/abs/
10.1109/TASLP.2019.2949928.

Wang:2023:MSG

[WYL^{+}23] Jie Wang, Yan Yang, Keyu Liu, Zhiping Zhu, and Xiaorong Liu. M3S: Scene graph driven multigranularity multi-task learning for multi-modal NER. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):111–120, 2023. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2022.3221017.

Wang:2016:CBW

[WYLZ16] Jin Wang, Liang-Chih Yu, K. Robert Lai, and Xuejie Zhang. Community-based weighted graph model for valence-arousal prediction of affective words. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(11): 1957–1968, ???? 2016. CODEN ???? ISSN 2329-9290.

Wang:2020:TSR

[WYLZ20] Jin Wang, Liang-Chih Yu, K. Robert Lai, and Xuejie Zhang. Treestructured regional CNN-LSTM model for dimensional sentiment analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):581–591, January 2020. CODEN ???? ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2019.2959251.

Wakabayashi:2023:SFI

[WYO23] Yukoh Wakabayashi, Kouei Yamaoka, and Nobutaka Ono. Sound field interpolation for rotationinvariant multichannel array signal processing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2286–2298, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3282098.

Weng:2015:DNN

[WYSD15] Chao Weng, Dong Yu, M. L. Seltzer, and J. Droppo. Deep neural networks for single-channel multitalker speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (10):1670–1679, October 2015. CODEN ???? ISSN 2329-9290.

Wang:2022:CTF

[WYY22] Taihui Wang, Feiran Yang, and Jun Yang. Convolutive transfer functionbased multichannel nonnegative matrix factorization for overdetermined blind source separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):802–815, 2022. CO-DEN ???? ISSN 2329-9290.

Wang:2024:MLP

[WYY24] Taihui Wang, Feiran Yang, and Jun Yang. Multichannel linear predictionbased speech dereverberation considering sparse and low-rank priors. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 1724–1735, 2024. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2024.3369535.

Wang:2023:TRA

[WYZ^{+}23] Qing Wang, Jixun Yao, Li Zhang, Pengcheng Guo, and Lei Xie. Timbrereserved adversarial attack in speaker identification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3848– 3858, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3306714.

Wu:2021:DGB

[WZ21] Linzhi Wu and Meishan Zhang. Deep graphbased character-level Chinese dependency parsing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1329–1339, 2021. CODEN ???? ISSN 2329-9290.

Wang:2024:LCS

[WZ24] Yile Wang and Yue Zhang. Lost in context? On the sense-wise variance of contextualized word embeddings. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 639–650, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3337643.

Wong:2023:MIR

[WZC23] Jeremy H. M. Wong, Huayun Zhang, and Nancy F. Chen. Modelling interrater uncertainty in spoken language assessment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):2886–2898, 2023. ISSN 2329-9290. URL

                                        https://dl.acm.org/doi/
                                        10.1109/TASLP.2023.3297958.

Wang:2023:HAM

[WZG^{+}23] Tianrui Wang, Weibin Zhu, Yingying Gao, Shilei Zhang, and Junlan Feng. Harmonic attention for monaural speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2424– 2436, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3284522.

Wu:2022:AMW

[WZH22] Qinzhuo Wu, Qi Zhang, and Xuanjing Huang. Automatic math word problem generation with topicexpression co-attention mechanism and reinforcement learning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 1061–1072, 2022. CODEN ???? ISSN 2329-9290.

Wang:2015:BCS

[WZL^{+}15] Rui Wang, Hai Zhao, Bao-Liang Lu, M. Utiyama, and E. Sumita. Bilingual continuous-space language model growing for statistical machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(7):1209– 1220, July 2015. CODEN ???? ISSN 2329-9290.

Wang:2024:GSL

[WZLL24] Yile Wang, Yue Zhang, Peng Li, and Yang Liu. Gradual syntactic label replacement for language model pretraining. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):486– 496, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3331096.

Wei:2021:GSL

[WZM21] Penghui Wei, Jiahao Zhao, and Wenji Mao. A graph-to-sequence learning framework for summarizing opinionated texts. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):1650–1660, 2021. CODEN ???? ISSN 2329-9290.

Wu:2022:DHT

[WZMH22] Junshuang Wu, Richong Zhang, Yongyi Mao, and Jinpeng Huai. Dealing with hierarchical types and label noise in fine-grained entity typing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):1305–1318, 2022. CODEN ???? ISSN 2329-9290.

Wu:2014:TCS

[WZS14] Dalei Wu, Wei-Ping Zhu, and M. N. S. Swamy. The theory of compressive sensing matching pursuit considering time-domain noise with application to speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (3):682–696, March 2014. CODEN ???? ISSN 2329-9290.

Wang:2019:RSL

[WZW19] Z. Wang, X. Zhang, and D. Wang. Robust speaker localization guided by deep learning-based timefrequency masking. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1): 178–188, January 2019. CODEN ???? ISSN 2329-9290.

Wu:2018:DDN

[WZZ^{+}18] S. Wu, D. Zhang, Z. Zhang, N. Yang, M. Li, and M. Zhou. Dependencyto-dependency neural machine translation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(11): 2132–2141, November 2018. CODEN ???? ISSN 2329-9290.

Wang:2024:DCC

[WZZ^{+}24] Xu Wang, Hainan Zhang, Shuai Zhao, Hongshen Chen, Zhuoye Ding, Zhiguo Wan, Bo Cheng, and Yanyan Lan. Debiasing counterfactual context with causal inference for multi-turn dialogue reasoning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1125– 1132, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3343608.

Xue:2014:FAD

[XAHJ^{+}14] Shaofei Xue, O. Abdel-Hamid, Hui Jiang, Lirong Dai, and Qingfeng Liu. Fast adaptation of deep neural network based on discriminant codes for speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):1713–1725, December 2014. CODEN ???? ISSN 2329-9290.

Xiong:2024:JDE

[XBZ^{+}24] Wenmeng Xiong, Changchun Bao, Jing Zhou, Maoshen Jia, and José Picheral. Joint DOA estimation and dereverberation based on multi-channel linear prediction filtering and azimuth sparsity. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 1481–1493, 2024. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2024.3363441.

Xu:2024:MCG

[XCHX24] Jiaming Xu, Jian Cui, Yunzhe Hao, and Bo Xu. Multi-cue guided semisupervised learning toward target speaker separation in real environments. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):151– 163, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3323856.

Xiao:2023:GHT

[XCZ^{+}23] Yingying Xiao, Shanmou Chen, Qiangqiang Zhang, Dongyuan Lin, Minglin Shen, Junhui Qian, and Shiyuan Wang. Generalized hyperbolic tangent based random Fourier conjugate gradient filter for nonlinear active noise control. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 619–632, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3230545.

Xu:2017:TDF

[XDC^{+}17] Xinzhou Xu, Jun Deng, Nicholas Cummins, Zixing Zhang, Chen Wu, Li Zhao, and Björn Schuller. A two-dimensional framework of multiple kernel subspace learning for recognizing emotion in speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(7):1436– 1449, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

                                        ieeexplore.ieee.org/
                                        document/7902189/.

Xu:2015:RAS

[XDDL15] Yong Xu, Jun Du, Li-Rong Dai, and Chin-Hui Lee. A regression approach to speech enhancement based on deep neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(1):7– 19, January 2015. CO-DEN ???? ISSN 2329-9290.

Xiong:2017:THP

[XDH^{+}17] Wayne Xiong, Jasha Droppo, Xuedong Huang, Frank Seide, Michael L. Seltzer, Andreas Stolcke, Dong Yu, and Geoffrey Zweig. Toward human parity in conversational speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(12):2410–2423, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8049322/.

Xie:2023:EPQ

[XFH^{+}23] Jiayuan Xie, Wenhao Fang, Qingbao Huang, Yi Cai, and Tao Wang. Enhancing paraphrase question generation with prior knowledge. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1464–1475, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2022.3221045.

Xiong:2018:EAI

[XGKM18] F. Xiong, S. Goetze, B. Kollmeier, and B. T. Meyer. Exploring auditoryinspired acoustic features for room acoustic parameter estimation from monaural speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(10): 1809–1820, October 2018. CODEN ???? ISSN 2329-9290.

Xiong:2019:JER

[XGKM19] F. Xiong, S. Goetze, B. Kollmeier, and B. T. Meyer. Joint estimation of reverberation time and early-to-late reverberation ratio from singlechannel speech signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (2):255–267, February 2019. CODEN ???? ISSN 2329-9290.

Xia:2023:ADB

[XH23] Wei Xia and John H. L.

Hansen. Attention and DCT based global context modeling for textindependent speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2668– 2679, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3284521.

Xiang:2024:TSD

[XHRC24] Yang Xiang, Jesper Lisby Højvang, Morten Højfeldt Rasmussen, and Mads Græsbøll Christensen. A twostage deep representation learning-based speech enhancement method using variational autoencoder and adversarial training. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):164– 177, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3321975.

Xu:2018:CDS

[XHS^{+}18] J. Xu, H. He, X. Sun, X. Ren, and S. Li. Crossdomain and semisupervised named entity recognition in Chinese social media: a unified model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (11):2142–2152, November 2018. CODEN ???? ISSN 2329-9290.

Xu:2017:UFL

[XHW^{+}17] Yong Xu, Qiang Huang, Wenwu Wang, Peter Foster, Siddharth Sigtia, Philip J. B. Jackson, and Mark D. Plumbley. Unsupervised feature learning based on deep models for environmental audio tagging. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(6):1230–1241, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7933054/.

Xu:2023:MWR

[XLG^{+}23] Yingrui Xu, Hao Liu, Jingguo Ge, Xiaodan Zhang, Jingyuan Hu, Yulei Wu, Honglei Lv, Hongbin Shi, and Wei Zhou. Mining weak relations between reviews for opinion spam detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):152– 162, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3221008.

Xie:2019:SEC

[XLL^{+}19] Y. Xie, R. Liang, Z. Liang, C. Huang, C. Zou, and B. Schuller. Speech emotion classification using attention-based LSTM. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (11):1675–1685, November 2019. CODEN ???? ISSN 2329-9290.

Xie:2021:BLD

[XLLW21] Xurong Xie, Xunying Liu, Tan Lee, and Lan Wang. Bayesian learning for deep neural network adaptation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2096–2110, 2021. CODEN ???? ISSN 2329-9290.

Xu:2018:GVE

[XLLY18] L. Xu, K. A. Lee, H. Li, and Z. Yang. Generalizing I-vector estimation for rapid speaker recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(4):749–759, April 2018. CODEN ???? ISSN 2329-9290.

Xie:2023:CAF

[XLZ^{+}23] Binhong Xie, Yu Li, Hongyan Zhao, Lihu Pan, and Enhui Wang. A crossattention fusion based graph convolution autoencoder for open relation extraction. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 476–485, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3226680.

Xue:2018:MDM

[XMBN18] W. Xue, A. H. Moore, M. Brookes, and P. A. Naylor. Modulationdomain multichannel Kalman filtering for speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(10):1833– 1847, October 2018. CO-DEN ???? ISSN 2329-9290.

Xue:2021:SEB

[XMBN21] Wei Xue, Alastair H.

Moore, Mike Brookes, and Patrick A. Naylor. Speech enhancement based on modulationdomain parametric multichannel Kalman filtering. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):393–405, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3040850.

Xiang:2014:PBA

[XNG^{+}14] Yong Xiang, I. Natgunanathan, Song Guo, Wanlei Zhou, and S. Nahavandi. Patchworkbased audio watermarking method robust to desynchronization attacks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (9):1413–1423, September 2014. CODEN ???? ISSN 2329-9290.

Xiang:2018:SSA

[XNP^{+}18] Y. Xiang, I. Natgunanathan, D. Peng, G. Hua, and B. Liu. Spread spectrum audio watermarking using multiple orthogonal PN sequences and variable embedding strengths and polarities. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):529–539, March 2018. CODEN ???? ISSN 2329-9290.

Xiang:2015:SSB

[XNRG15] Yong Xiang, I. Natgunanathan, Yue Rong, and Song Guo. Spread spectrum-based high embedding capacity watermarking method for audio signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2228– 2237, December 2015. CODEN ???? ISSN 2329-9290.

Xu:2016:SDA

[XO16] Haotian Xu and Zhijian Ou. Scalable discovery of audio fingerprint motifs in broadcast streams with determinantal point process based motif clustering. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(5):978–989, May 2016. CODEN ???? ISSN 2329-9290.

Xie:2022:DDG

[XPC^{+}22] Jiayuan Xie, Ningxin Peng, Yi Cai, Tao Wang, and Qingbao Huang. Diverse distractor generation for constructing highquality multiple choice questions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):280– 291, 2022. CODEN ???? ISSN 2329-9290.

Xu:2021:TSV

[XRWL21] Chenglin Xu, Wei Rao, Jibin Wu, and Haizhou Li. Target speaker verification with selective auditory attention for single and multi-talker speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2696– 2709, 2021. CODEN ???? ISSN 2329-9290.

Xu:2023:VCM

[XSC^{+}23] Xuexin Xu, Liang Shi, Xunquan Chen, Pingyuan Lin, Jie Lian, Jinhui Chen, Zhihong Zhang, and Edwin R. Hancock. Any-to-any voice conversion with multilayer speaker adaptation and content supervision. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3431– 3445, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3306716.

Xu:2019:DWM

[XSL^{+}19] Z. Xu, C. Sun, Y. Long, B. Liu, B. Wang, M. Wang, M. Zhang, and X. Wang. Dynamic working memory for context-aware response generation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(9): 1419–1431, September 2019. CODEN ???? ISSN 2329-9290.

Xie:2023:NNV

[XT23] Chao Xie and Tomoki Toda. Noisy-to-noisy voice conversion under variations of noisy condition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3871– 3882, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3313426.

Xie:2021:ZSA

[XV21] Huang Xie and Tuomas Virtanen. Zero-shot audio classification via semantic embeddings. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??):

1233–1242, 2021. CODEN ???? ISSN 2329-9290.

Xu:2021:CSR

[XWS^{+}21] Kun Xu, Han Wu, Linfeng Song, Haisong Zhang, Linqi Song, and Dong Yu. Conversational semantic role labeling. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2465–2475, 2021. CODEN ???? ISSN 2329-9290.

Xiao:2018:RNS

[XWWW18] K. Xiao, S. Wang, M. Wan, and L. Wu. Radiated noise suppression for electrolarynx speech based on multiband time-domain amplitude modulation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (9):1585–1593, September 2018. CODEN ???? ISSN 2329-9290.

Xiao:2019:RME

[XWWW19] K. Xiao, S. Wang, M. Wan, and L. Wu. Reconstruction of Mandarin electrolaryngeal fricatives with hybrid noise source. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (2):383–391, February 2019. CODEN ???? ISSN 2329-9290.

Xiao:2023:TEP

[XWY^{+}23] Rong Xiao, Yu Wan, Baosong Yang, Haibo Zhang, Huajin Tang, Derek F. Wong, and Boxing Chen. Towards energy-preserving natural language understanding with spiking neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):439– 447, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3221011.

Xiao:2016:LAA

[XWZ16] Tong Xiao, Derek F. Wong, and Jingbo Zhu. A loss-augmented approach to training syntactic machine translation systems. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 24(11): 2069–2083, ???? 2016. CODEN ???? ISSN 2329-9290.

Xiao:2023:TAP

[XXS^{+}23] Lin Xiao, Pengyu Xu, Mingyang Song, Huafeng Liu, Liping Jing, and Xiangliang Zhang. Triple alliance prototype orthotist network for long-tailed multi-label text classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2616– 2628, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3265860.

Xu:2024:BSQ

[XXWY24] Xuenan Xu, Zeyu Xie, Mengyue Wu, and Kai Yu. Beyond the status quo: a contemporary survey of advances and challenges in audio captioning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 95–112, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3321968.

Xu:2014:UAA

[XYH14] Jian Xu, Zhi-Jie Yan, and Qiang Huo. An unsupervised adaptation approach to leveraging feedback loop data by using i-vector for data clustering and selection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (11):1581–1589, November 2014. CODEN ???? ISSN 2329-9290.

Xu:2021:MPL

[XYH^{+}21] Junhao Xu, Jianwei Yu, Shoukang Hu, Xunying Liu, and Helen Meng. Mixed precision low-bit quantization of neural network language models for speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3679–3693, 2021. CODEN ???? ISSN 2329-9290.

Xu:2023:DFB

[XYY^{+}23] Longting Xu, Jichen Yang, Chang Huai You, Xinyuan Qian, and Daiyu Huang. Device features based on linear transformation with parallel training data for replay speech detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1574– 1586, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3267610.

Xu:2023:CSQ

[XZF23] Ziyi Xu, Ziyue Zhao, and Tim Fingscheidt. Coded speech quality measurement by a non-intrusive PESQ-DNN. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3404– 3417, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3317574.

Xu:2023:ICI

[XZFN23] Lin Xu, Qixian Zhou, Jinlan Fu, and See-Kiong Ng. CET2: Modelling topic transitions for coherent and engaging knowledge-grounded conversations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3527– 3536, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3313418.

Xie:2021:HNP

[XZL^{+}21] Zhiwen Xie, Runjie Zhu, Jin Liu, Guangyou Zhou, and Jimmy Xiangji Huang. Hierarchical neighbor propagation with bidirectional graph attention network for relation prediction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):1762–1773, 2021. CODEN ???? ISSN 2329-9290.

Xie:2023:TTA

[XZL^{+}23] Zhiwen Xie, Runjie Zhu, Jin Liu, Guangyou Zhou, and Jimmy Xiangji Huang. TARGAT: a time-aware relational graph attention model for temporal knowledge graph embedding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2246– 2258, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3282101.

Xue:2023:DLF

[XZSX23] Zhengshan Xue, Xiaolei Zhang, Tingxun Shi, and Deyi Xiong. DetTrans: a lightweight framework to detect and translate noisy inputs simultaneously. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3696– 3705, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3284513.

Xiong:2015:TBC

[XZW15] Deyi Xiong, Min Zhang, and Xing Wang. Topicbased coherence modeling for statistical machine translation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(3): 483–493, March 2015. CODEN ???? ISSN 2329-9290.

Yang:2022:ADL

[Yan22]  Feiran Yang.  Anal-

ysis of deficient-length partitioned-block frequencydomain adaptive filters. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):456–467, 2022. CO-DEN ???? ISSN 2329-9290.

Yella:2014:OSD

[YB14] S. H. Yella and H. Bourlard.

Overlapping speech detection using long-term conversational features for speaker diarization in meeting room conversations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1688–1700, December 2014. CODEN ???? ISSN 2329-9290.

Yang:2014:BNR

[YBC^{+}14] Na Yang, He Ba, Weiyang Cai, I. Demirkol, and W. Heinzelman. BaNa: a noise resilient fundamental frequency detection algorithm for speech and music. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1833– 1848, December 2014. CODEN ???? ISSN 2329-9290.

Yang:2021:NCD

[YBHC21] Wenxing Yang, Jacob Benesty, Gongping Huang, and Jingdong Chen. A new class of differential beamformers. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):594–606, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3045561.

Yang:2022:VWV

[YC22] Joon-Young Yang and Joon-Hyuk Chang. VACE-WPE: Virtual acoustic channel expansion based on neural networks for weighted prediction errorbased speech dereverberation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):174–189, 2022. CODEN ???? ISSN 2329-9290.

Yoshino:2024:OTD

[YCC^{+}24] Koichiro Yoshino, Yun-Nung Chen, Paul Crook, Satwik Kottur, Jinchao Li, Behnam Hedayatnia, Seungwhan Moon, Zhengcong Fei, Zekang Li, Jinchao Zhang, Yang Feng, Jie Zhou, Seokhwan Kim, Yang Liu, Di Jin, Alexandros Papangelis, Karthik Gopalakrishnan, Dilek Hakkani-Tur, Babak Damavandi, Alborz Geramifard, Chiori Hori, Ankit Shah, Chen Zhang, Haizhou Li, João Sedoc, Luis F. D’Haro, Rafael Banchs, and Alexander Rudnicky. Overview of the tenth dialog system technology challenge: DSTC10. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):765– 778, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3293030.

Yu:2022:NAA

[YCL22] Fu-Hao Yu, Kuan-Yu Chen, and Ke-Han Lu. Non-autoregressive ASR modeling using pre-trained language models for Chinese speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):1474–1482, 2022. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2022.3166400.

Yang:2017:IWR

[YCLS17] Liner Yang, Xinxiong Chen, Zhiyuan Liu, and Maosong Sun. Improving word representations with document labels. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):863–870, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7833113/.

Yang:2021:KSU

[YCM^{+}21] Runyan Yang, Gaofeng Cheng, Haoran Miao, Ta Li, Pengyuan Zhang, and Yonghong Yan. Keyword search using attentionbased end-to-end ASR and frame-synchronous phoneme alignments. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??):

3202–3215, 2021. CODEN ???? ISSN 2329-9290.

Yusuf:2023:EEO

[YČS23] Bolaji Yusuf, Jan Černocký, and Murat Saraçlar. Endto-end open vocabulary keyword search with multilingual neural representations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3070– 3080, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3301239.

Yuan:2021:EMR

[YDW^{+}21] Weitao Yuan, Bofei Dong, Shengbei Wang, Masashi Unoki, and Wenwu Wang. Evolving multi-resolution pooling CNN for monaural singing voice separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):807–822, 2021. CODEN ???? ISSN 2329-9290.

Yang:2019:EOS

[YDZ19] J. Yang, R. K. Das, and N. Zhou. Extraction of octave spectra information for spoofing attack detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2373– 2384, December 2019. CODEN ???? ISSN 2329-9290.

Yang:2017:SCA

[YEY17] Feiran Yang, Gerald Enzner, and Jun Yang. Statistical convergence analysis for optimal control of DFT-domain adaptive echo canceler. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(5): 1095–1106, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7858641/.

Yadav:2024:JDB

[YG24] Shekhar Kumar Yadav and Nithin V. George. Joint dereverberation and beamforming with blind estimation of the shape parameter of the desired source prior. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 779–793, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3335000.

Yusuf:2019:LRK

[YGS19] B. Yusuf, B. Gundogdu, and M. Saraclar. Low resource keyword search with synthesized crosslingual exemplars. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(7): 1126–1135, July 2019.

CODEN ???? ISSN 2329-9290.

Yilmaz:2014:NRE

[YGV14] E. Yilmaz, J. F. Gemmeke, and H. Van hamme. Noise robust exemplar matching using sparse representations of speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (8):1306–1319, August 2014. CODEN ???? ISSN 2329-9290.

Yu:2017:ALB

[YH17] Chengzhu Yu and John H. L. Hansen. Active learning based constrained clustering for speaker diarization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(11): 2188–2198, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/8030331/.

Yousefi:2021:BBH

[YH21] Midia Yousefi and John H. L. Hansen. Blockbased high performance CNN architectures for frame-level overlapping speech detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 28–40, January 2021.

                                        URL https://dl.acm.

org/doi/10.1109/TASLP.
2020.3036237.

Yang:2019:ESE

[YHC19] G. Yang, H. He, and Q. Chen. Emotionsemantic-enhanced neural network. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(3): 531–543, March 2019. CODEN ???? ISSN 2329-9290.

Yu:2020:EBN

[YHC^{+}20] Yi Yu, Hongsen He, Badong Chen, Jianghui Li, Youwen Zhang, and Lu Lu. M-estimate based normalized subband adaptive filter algorithm: Performance analysis and improvements. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28 (??):225–239, January 2020. CODEN ???? ISSN 2329-9290. URL https:

//dl.acm.org/doi/abs/
10.1109/TASLP.2019.2950597.

Yuan:2023:RGR

[YHL^{+}23] Lin Yuan, Guoheng Huang, Fenghuan Li, Xiaochen Yuan, Chi-Man Pun, and Guo Zhong. RBA-GCN: Relational bilevel aggregation graph convolutional network for emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2325– 2337, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3284509.

Yoshimura:2017:SOM

[YHO^{+}17] Takenori Yoshimura, Kei Hashimoto, Keiichiro Oura, Yoshihiko Nankaku, and Keiichi Tokuda. Simultaneous optimization of multiple tree-based factor analyzed HMM for speech synthesis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(9): 1836–1845, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7961246/.

Yoshimura:2018:MCB

[YHO^{+}18] T. Yoshimura, K. Hashimoto, K. Oura, Y. Nankaku, and K. Tokuda. Melcepstrum-based quantization noise shaping applied to neural-network-based speech waveform synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(7):1177–1184, July 2018. CODEN ???? ISSN 2329-9290.

Yu:2019:GIA

[YJX19] J. Yu, J. Jiang, and R. Xia. Global inference for aspect and opinion terms co-extraction based on multi-task neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1): 168–177, January 2019. CODEN ???? ISSN 2329-9290.

Yu:2020:ESA

[YJX20] Jianfei Yu, Jing Jiang, and Rui Xia. Entitysensitive attention and fusion network for entitylevel multimodal sentiment classification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 429–439, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2957872.

Yang:2014:OSD

[YK14] Jae-Mo Yang and Hong-Goo Kang. Online speech dereverberation algorithm based on adaptive multichannel linear prediction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (3):608–619, March 2014. CODEN ???? ISSN 2329-9290.

Yang:2017:CTF

[YK17] Wenjun Yang and Sridhar Krishnan. Combining temporal features by local binary pattern for acoustic scene classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(6):1315–1321, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7933046/.

Yatabe:2021:DBB

[YK21a] Kohei Yatabe and Daichi Kitamura. Determined BSS based on timefrequency masking and its application to harmonic vector analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1609–1625, 2021. CODEN ???? ISSN 2329-9290.

Yu:2021:RAP

[YK21b] Wangyang Yu and W. Bastiaan Kleijn. Room acoustical parameter estimation from room impulse responses using deep neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):436–447, January

                                        2021. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.3043115.

Yoon:2023:OTL

[YKL^{+}23] Ji Won Yoon, Hyung Yong Kim, Hyeonseung Lee, Sunghwan Ahn, and Nam Soo Kim. Oracle teacher: Leveraging target information for better knowledge distillation of CTC models. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2974– 2987, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3297955.

Yee:2018:NRP

[YKPMP18] Dianna Yee, Homayoun Kamkar-Parsi, Rainer Martin, and Henning Puder. A noise reduction postfilter for binaurally linked singlemicrophone hearing aids utilizing a nearby external microphone. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(1): 5–18, ???? 2018. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7987050/.

Yan:2024:DOM

[YLC^{+}24] Jingsong Yan, Piji Li, Haibin Chen, Junhao Zheng, and Qianli Ma. Does the order matter? A random generative way to learn label hierarchy for hierarchical text classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):276– 285, 2024. ISSN 2329-

9290. URL https://dl.

                                        acm.org/doi/10.1109/
                                        TASLP.2023.3329374.

Yang:2015:PEU

[YLD15] Guang Yang, R. F. Lyon, and E. M. Drakakis. Psychophysical evaluation of an ultra-low power, analog biomimetic cochlear implant processor filterbank architecture with across channels AGC. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2465–2473, December 2015. CODEN ???? ISSN 2329-9290.

Yang:2022:TGD

[YLFL22] Sen Yang, Yang Liu, Dawei Feng, and Dongsheng Li. Text generation from data with dynamic planning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 26–34, 2022. CODEN ???? ISSN 2329-9290.

Yang:2022:CJM

[YLGG22] Jingxuan Yang, Si Li, Sheng Gao, and Jun Guo. CorefDPR: a joint model for coreference resolution and dropped pronoun recovery in Chinese conversations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):571– 581, 2022. CODEN ???? ISSN 2329-9290.

Yen:2023:RNA

[YLH23] Benjamin Yen, Yameizhen Li, and Yusuke Hioka. Rotor noise-aware noise covariance matrix estimation for unmanned aerial vehicle audition. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 2491–2506, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3288410.

Yang:2016:SAS

[YLJ16] Cheng-Yen Yang, Chih-Wei Liu, and Shyh-Jye Jou. A systematic ANSI S1.11 filter bank specification relaxation and its efficient multirate architecture for hearing-aid systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(8):1380–1392, ???? 2016. CODEN ???? ISSN 2329-9290.

Yoon:2021:TTF

[YLK^{+}21] Ji Won Yoon, Hyeonseung Lee, Hyung Yong Kim, Won Ik Cho, and Nam Soo Kim. TutorNet: Towards flexible knowledge distillation for endto-end speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1626–1638, 2021.

CODEN ???? ISSN 2329-9290.

Yeung:2015:SSM

[YLL15] Yu Ting Yeung, Tan Lee, and Cheung-Chi Leung. Supervised singlemicrophone multi-talker speech separation with conditional random fields. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2334–2342, December 2015. CODEN ???? ISSN 2329-9290.

Yang:2021:LDD

[YLL21] Bing Yang, Hong Liu, and Xiaofei Li. Learning deep direct-path relative transfer function for binaural sound source localization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3491–3503, 2021. CODEN ???? ISSN 2329-9290.

Yu:2019:STP

[YLLC19] J. Yu, Q. Ling, C. Luo, and C. W. Chen. Synthesizing 3D Trump: Predicting and visualizing the relationship between text, speech, and articulatory movements. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2223– 2233, December 2019. CODEN ???? ISSN 2329-9290.

Yang:2019:MSS

[YLPL19] B. Yang, H. Liu, C. Pang, and X. Li. Multiple sound source counting and localization based on TF-wise spatial spectrum clustering. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(8):1241–1255, August 2019. CODEN ???? ISSN 2329-9290.

Yao:2024:IST

[YLQZ24] Jiadi Yao, Hong Luo, Jun Qi, and Xiao-Lei Zhang. Interpretable spectrum transformation attacks to speaker recognition systems. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1531– 1545, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3364100.

Lee:2014:ISR

[yLsL14] Hung yi Lee and Lin shan Lee. Improved semantic retrieval of spoken content by document/query expansion with random walk over acoustic similarity graphs. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): 80–94, January 2014. CO-DEN ???? ISSN 2329-

Lee:2014:SKO

[yLSY^{+}14] Hung yi Lee, Sz-Rung Shiang, Ching-Feng Yeh, Yun-Nung Chen, Yu Huang, Sheng-Yi Kong, and Lin shan Lee. Spoken knowledge organization by semantic structuring and a prototype course lecture system for personalized learning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(5): 883–898, May 2014. CO-DEN ???? ISSN 2329-9290.

Yoo:2015:FBR

[YLY15] In-Chul Yoo, Hyeontaek Lim, and Dongsuk Yook. Formant-based robust voice activity detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(12):2238–2245, December 2015. CODEN ???? ISSN 2329-9290.

Yang:2017:TIF

[YM17] Xuefeng Yang and Kezhi Mao. Task independent fine tuning for word embeddings. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(4): 885–894, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7797241/.

Yu:2019:ASI

[YMM19] J. Yu, K. Markov, and T. Matsui. Articulatory and spectrum information fusion based on deep recurrent neural networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (4):742–752, April 2019. CODEN ???? ISSN 2329-9290.

Yong:2014:EBM

[YND14] Pei Chee Yong, S. Nordholm, and Hai Huyen Dam. Effective binaural multi-channel processing algorithm for improved environmental presence. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (12):2012–2024, December 2014. CODEN ???? ISSN 2329-9290.

Yamaoka:2021:TFW

[YOM21] Kouei Yamaoka, Nobutaka Ono, and Shoji Makino. Time-frequency-bin-wise linear combination of beamformers for distortionless signal enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3461– 3475, 2021. CODEN ???? ISSN 2329-9290.

Yao:2022:SRS

[YPC^{+}22] Zengwei Yao, Wenjie Pei, Fanglin Chen, Guangming Lu, and David Zhang. Stepwise-refining speech separation network via fine-grained encoding in high-order latent domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):378–393, 2022. CO-DEN ???? ISSN 2329-9290.

Yang:2023:FHF

[YPH23] Runxuan Yang, Yuyang Peng, and Xiaolin Hu. A fast high-fidelity sourcefilter vocoder with lightweight neural modules. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 3362–3373, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3321191.

Yin:2021:RQG

[YPL21] Kai-Li Yin, Yi-Fei Pu, and Lu Lu. Robust Qgradient subband adaptive filter for nonlinear active noise control. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2741–2752, 2021. CODEN ???? ISSN 2329-9290.

Yu:2015:CMB

[YSCZ15] Kai Yu, Kai Sun, Lu Chen, and Su Zhu. Constrained Markov Bayesian polynomial for efficient dialogue state tracking. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2177–2188, December 2015. CODEN ???? ISSN 2329-9290.

Yeh:2015:IFR

[YsL15] Ching-Feng Yeh and Lin shan Lee. An improved framework for recognizing highly imbalanced bilingual code-switched lectures with cross-language acoustic modeling and frame-level language identification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(7):1144– 1159, July 2015. CODEN ???? ISSN 2329-9290.

Yi:2023:AMT

[YTF^{+}23] Jiangyan Yi, Jianhua Tao, Ruibo Fu, Tao Wang, Chu Yuan Zhang, and Chenglong Wang. Adversarial multi-task learning for Mandarin prosodic boundary prediction with multi-modal embeddings. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2963– 2973, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3301235.

Yi:2019:LAT

[YTWB19] J. Yi, J. Tao, Z. Wen, and Y. Bai. Languageadversarial transfer learning for low-resource speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(3):621– 630, March 2019. CODEN ???? ISSN 2329-9290.

Yuksel:2021:SCD

[YUK21] Arda Yüksel, Berke Uğurlu, and Aykut Koç. Semantic change detection with Gaussian word embeddings. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3349–3361, 2021. CODEN ???? ISSN 2329-9290.

Yan:2014:SLS

[YW14] Su Yan and Xiaojun Wan. SRRank: Leveraging semantic roles for extractive multi-document summarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):2048– 2058, December 2014. CODEN ???? ISSN 2329-9290.

Yang:2021:MMP

[YWDQ21] Jichen Yang, Hongji Wang, Rohan Kumar Das, and Yanmin Qian. Modified magnitude-phase spectrum information for spoofing detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 1065–1078, 2021. CODEN ???? ISSN 2329-9290.

Yu:2018:RWE

[YWLZ18] L. Yu, J. Wang, K. R. Lai, and X. Zhang. Refining word embeddings using intensity scores for sentiment analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(3): 671–681, March 2018. CODEN ???? ISSN 2329-9290.

Yoneyama:2023:HFP

[YWT23] Reo Yoneyama, Yi-Chiao Wu, and Tomoki Toda. High-fidelity and pitchcontrollable neural vocoder based on unified sourcefilter networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 3717–3729, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3313410.

Yuan:2023:UDU

[YWW^{+}23] Weitao Yuan, Shengbei Wang, Jianming Wang, Masashi Unoki, and Wenwu Wang. Unsupervised deep unfolded representation learning for singing voice separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3206– 3220, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3302237.

Yu:2022:DDR

[YXX^{+}22] Weijie Yu, Chen Xu, Jun Xu, Liang Pang, and Ji-Rong Wen. Distribution distance regularized sequence representation for text matching in asymmetrical domains. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):721–733, 2022. CO-DEN ???? ISSN 2329-9290.

Yang:2023:DDD

[YYW^{+}23] Dongchao Yang, Jianwei Yu, Helin Wang, Wen Wang, Chao Weng, Yuexian Zou, and Dong Yu. Diffsound: Discrete diffusion model for text-to-sound generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1720– 1733, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3268730.

Yang:2023:GGF

[YYY^{+}23] Jian Yang, Yuwei Yin, Liqun Yang, Shuming Ma, Haoyang Huang, Dongdong Zhang, Furu Wei, and Zhoujun Li.

GTrans: Grouping and fusing transformer layers for neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1489– 1498, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3221040.

Yao:2024:KLD

[YYZY24] Yu Yao, Peng Yang, Guangzhen Zhao, and Guoshun Yin. KGAgent: Learning a deep reinforced agent for keyphrase generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1928– 1940, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3375630.

Yang:2021:MAS

[YZCW21] Jielong Yang, Xionghu Zhong, Weiguang Chen, and Wenwu Wang. Multiple acoustic source localization in microphone array networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29 (??):334–347, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3039390.

Yang:2018:JPT

[YZL^{+}18] L. Yang, M. Zhang, Y. Liu, M. Sun, N. Yu, and G. Fu. Joint POS tagging and dependence parsing with transition-based neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(8): 1352–1358, August 2018. CODEN ???? ISSN 2329-9290.

Ying:2017:WDS

[YZLY17] Dongwen Ying, Ruohua Zhou, Junfeng Li, and Yonghong Yan. Windowdominant signal subspace methods for multiple short-term speech source localization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 25(4): 731–744, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7736102/.

Yu:2018:RST

[YZW^{+}18] K. Yu, Z. Zhao, X. Wu, H. Lin, and X. Liu. Rich short text conversation using semantic-key-controlled sequence generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(8):1359–1368, August 2018. CODEN ???? ISSN 2329-9290.

Yu:2021:AVM

[YZW^{+}21] Jianwei Yu, Shi-Xiong Zhang, Bo Wu, Shansong Liu, Shoukang Hu, Mengzhe Geng, Xunying Liu, Helen Meng, and Dong Yu. Audio-visual multi-channel integration and recognition of overlapped speech. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 2067–2082, 2021. CODEN ???? ISSN 2329-9290.

Zhang:2014:TDS

[ZA14] Wen Zhang and T. D.

Abhayapala. Three dimensional sound field reproduction using multiple circular loudspeaker arrays: Functional analysis guided approach. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(7): 1184–1194, July 2014. CODEN ???? ISSN 2329-9290.

Zheng:2024:IUT

[ZAL24] Rui-Chen Zheng, Yang Ai, and Zhen-Hua Ling. Incorporating ultrasound tongue images for audiovisual speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1430–1444, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3361376.

Zhang:2018:ANC

[ZAZ^{+}18] J. Zhang, T. D. Abhayapala, W. Zhang, P. N. Samarasinghe, and S. Jiang. Active noise control over space: a wave domain approach. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (4):774–786, April 2018. CODEN ???? ISSN 2329-9290.

Zitouni:2014:APC

[ZB14] I. Zitouni and Y. Benajiba. Aligned-parallel-corpora based semi-supervised learning for Arabic mention detection. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(2): 314–324, February 2014. CODEN ???? ISSN 2329-9290.

Zhao:2014:DRD

[ZBC14] Liheng Zhao, J. Benesty, and Jingdong Chen. Design of robust differential microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (10):1455–1466, October 2014. CODEN ???? ISSN 2329-9290.

Zhao:2021:DBB

[ZBCH21] Xudong Zhao, Jacob Benesty, Jingdong Chen, and Gongping Huang.

Differential beamforming from the beampattern factorization perspective. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):632–643, January

2021. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.3046082.

Zhou:2024:OAN

[ZBW^{+}24] Yongwei Zhou, Junwei Bao, Youzheng Wu, Xiaodong He, and Tiejun Zhao. Operation-augmented numerical reasoning for question answering. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 15–28, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3316448.

Zhang:2021:SRM

[ZCDH21] Jie Zhang, Huawei Chen, Li-Rong Dai, and Richard Christian Hendriks. A study on reference microphone selection for multi-microphone speech enhancement. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 29(??): 671–683, 2021. CODEN ???? ISSN 2329-9290.

Zhang:2015:CLA

[ZCDS15] Zixing Zhang, E. Coutinho, Jun Deng, and B. Schuller. Cooperative learning and its application to emotion recognition from speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(1):115– 126, January 2015. CO-DEN ???? ISSN 2329-9290.

Zao:2014:SEE

[ZCF14] L. Zão, R. Coelho, and P. Flandrin. Speech enhancement with EMD and Hurst-based mode selection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(5):899–911, May 2014. CODEN ???? ISSN 2329-9290.

Zhang:2018:MSS

[ZCHH18] J. Zhang, S. P. Chepuri, R. C. Hendriks, and R. Heusdens. Microphone subset selection for MVDR beamformer based noise reduction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (3):550–563, March 2018. CODEN ???? ISSN 2329-9290.

Zhao:2024:FFD

[ZCHW24] Yaru Zhao, Bo Cheng, Yakun Huang, and Zhiguo Wan. FluGCF: a fluent dialogue generation model with coherent concept entity flow. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 853–867, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3340610.

Zhu:2019:RPA

[ZCQ^{+}19] Q. Zhu, P. Coleman, X. Qiu, M. Wu, J. Yang, and I. Burnett. Robust personal audio geometry optimization in the SVD-based modal domain. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(3):610–620, March 2019. CODEN ???? ISSN 2329-9290.

Zhu:2024:BCD

[ZCW^{+}24] Han Zhu, Gaofeng Cheng, Jindong Wang, Wenxin Hou, Pengyuan Zhang, and Yonghong Yan. Boosting cross-domain speech recognition with selfsupervision. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):471– 485, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3301230.

Zhang:2016:DMA

[ZCY16] Qiaoling Zhang, Zhe Chen, and Fuliang Yin. Distributed marginalized auxiliary particle filter for speaker tracking in distributed microphone networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(11):1921– 1934, ???? 2016. CODEN ???? ISSN 2329-9290.

Zhu:2022:ICN

[ZCY^{+}22] Peng Zhu, Dawei Cheng, Fangzhou Yang, Yifeng Luo, Dingjiang Huang, Weining Qian, and Aoying Zhou. Improving Chinese named entity recognition by large-scale syntactic dependency graph. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):979–991, 2022. CO-DEN ???? ISSN 2329-9290.

Zhang:2018:MEP

[ZCYZ18] W. Zhang, Z. Chen, F. Yin, and Q. Zhang. Melody extraction from polyphonic music using particle filter and dynamic programming. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(9): 1620–1632, September 2018. CODEN ???? ISSN 2329-9290.

Zhang:2014:BCC

[ZDC14] Min Zhang, Xiangyu Duan, and Wenliang Chen. Bayesian constituent context model for grammar induction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22 (2):531–541, February 2014. CODEN ???? ISSN 2329-9290.

Zhang:2021:SSR

[ZDD21] Jie Zhang, Jun Du, and Li-Rong Dai. Sensor selection for relative acoustic transfer function steered linearlyconstrained beamformers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1220–1232, 2021. CODEN ???? ISSN 2329-9290.

Zapata:2014:MFB

[ZDG14] J. R. Zapata, M. E. P.

Davies, and E. Gomez. Multi-feature beat tracking. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(4):816–825, April 2014. CODEN ???? ISSN 2329-9290.

Zhao:2021:CFA

[ZDGO21] Guanlong Zhao, Shaojin Ding, and Ricardo Gutierrez-Osuna. Converting foreign accent speech without a reference. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2367–2381, 2021. CODEN ???? ISSN 2329-

Zheng:2018:SAM

[ZDLK18] C. Zheng, A. Deleforge, X. Li, and W. Kellermann. Statistical analysis of the multichannel Wiener filter using a bivariate normal distribution for sample covariance matrices. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(5): 951–966, May 2018. CO-DEN ???? ISSN 2329-9290.

Zhou:2021:IFA

[ZDZ^{+}21] Hengshun Zhou, Jun Du, Yuanyuan Zhang, Qing Wang, Qing-Feng Liu, and Chin-Hui Lee. Information fusion in attention networks using adaptive and multi-level factorized bilinear pooling for audiovisual emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2617–2629, 2021. CODEN ???? ISSN 2329-9290.

Zhang:2023:PPE

[ZDZ^{+}23] Chen Zhang, Luis Fernando D’Haro, Qiquan Zhang, Thomas Friedrichs, and Haizhou Li. PoE: a panel of experts for generalized automatic dialogue assessment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1234– 1250, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3250825.

Zohourian:2018:BSL

[ZEM18] M. Zohourian, G. Enzner, and R. Martin. Binaural speaker localization integrated into an adaptive beamformer for hearing aids. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(3):515–528, March 2018. CODEN ???? ISSN 2329-9290.

Zhang:2014:DTS

[ZF14] Weibin Zhang and P. Fung.

Discriminatively trained sparse inverse covariance matrices for speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(5):873– 882, May 2014. CODEN ???? ISSN 2329-9290.

Zezario:2023:DLB

[ZFC^{+}23] Ryandhimas E. Zezario, Szu-Wei Fu, Fei Chen, Chiou-Shann Fuh, Hsin-Min Wang, and Yu Tsao. Deep learning-based nonintrusive multi-objective speech assessment model with cross-domain features. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):54–70, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2022.3205757.

Zhang:2016:FFE

[ZG16] Geliang Zhang and Simon Godsill. Fundamental frequency estimation in speech signals with variable rate particle filters. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(5):890–900, May 2016. CODEN ???? ISSN 2329-9290.

Zeghidour:2021:WEE

[ZG21] Neil Zeghidour and David Grangier. Wavesplit: End-to-end speech separation by speaker clustering. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2840–2849, 2021. CODEN ???? ISSN 2329-9290.

Zhu:2023:APL

[ZGC^{+}23] Han Zhu, Dongji Gao, Gaofeng Cheng, Daniel Povey, Pengyuan Zhang, and Yonghong Yan. Alternative pseudo-labeling for semi-supervised automatic speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3320– 3330, 2023. ISSN 2329-

                                        9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2023.3306709.

Zhang:2020:MIH

[ZGL^{+}20] Mengfan Zhang, Zhongshu Ge, Tiejun Liu, Xihong Wu, and Tianshu Qu. Modeling of individual HRTFs based on spatial principal component analysis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 785–797, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2020.2967539.

Zhang:2023:CSN

[ZGL^{+}23] Hongsheng Zhang, Jizhang Gan, Ting Liu, Kui Huang, and Hong Yang. Coefficients-switched normalized least-mean- squares adaption in echo canceler of sparse-echo-path. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):1191–1199, 2023. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2023.3250824.

Zhao:2019:UPP

[ZGO19] G. Zhao and R. Gutierrez-Osuna. Using phonetic posteriorgram based frame pairing for segmental accent conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (10):1649–1660, October 2019. CODEN ???? ISSN 2329-9290.

Zhao:2023:RSA

[ZGZ23] Haiquan Zhao, Yuan Gao, and Yingying Zhu. Robust subband adaptive filter algorithms-based mixture correntropy and application to acoustic echo cancellation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1223– 1233, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3250845.

Zeng:2014:DDS

[ZH14] Yuan Zeng and R. C. Hendriks. Distributed delay and sum beamformer for speech enhancement via randomized gossip. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(1): 260–273, January 2014. CODEN ???? ISSN 2329-9290.

Zhong:2015:TFM

[ZH15] Xionghu Zhong and J. R. Hopgood. A Time– Frequency masking based random finite set particle filtering method for multiple acoustic source detection and tracking. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2356–2370, December 2015. CODEN ???? ISSN 2329-9290.

Zhang:2018:LDR

[ZH18] Q. Zhang and J. H. L.

Hansen. Language/dialect recognition based on unsupervised deep learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (5):873–882, May 2018. CODEN ???? ISSN 2329-9290.

Zakeri:2019:AIH

[ZH19] V. Zakeri and A. J. Hodgson. Automatic identification of hard and soft bone tissues by analyzing drilling sounds. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(2): 404–414, February 2019. CODEN ???? ISSN 2329-9290.

Zhang:2018:SNF

[ZHB^{+}18] W. Zhang, C. Hofmann, M. Buerger, T. D. Abhayapala, and W. Kellermann. Spatial noise-field control with online secondary path modeling: a wave-domain approach. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (12):2355–2370, December 2018. CODEN ???? ISSN 2329-9290.

Zhao:2021:DSB

[ZHCB21] Xudong Zhao, Gongping Huang, Jingdong Chen, and Jacob Benesty. On the design of 3D steerable beamformers with uniform concentric circular microphone arrays. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2764–2778, 2021. CODEN ???? ISSN 2329-9290.

Zhao:2023:DDM

[ZHCB23] Xudong Zhao, Gongping Huang, Jingdong Chen, and Jacob Benesty. Design of 2D and 3D differential microphone arrays with a multistage framework. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2016– 2031, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3278182.

Zheng:2023:BCE

[Zhe23] Xiaoqing Zheng. Building conventional “Experts” with a dialogue logic programming language. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1784– 1796, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3271119.

Zhang:2018:RDS

[ZHH18] J. Zhang, R. Heusdens, and R. C. Hendriks. Ratedistributed spatial filtering based noise reduction in wireless acoustic sensor networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(11):2015– 2026, November 2018. CODEN ???? ISSN 2329-9290.

Zhang:2019:RAT

[ZHH19] J. Zhang, R. Heusdens, and R. C. Hendriks. Relative acoustic transfer function estimation in wireless acoustic sensor networks. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(10):1507– 1519, October 2019. CO-DEN ???? ISSN 2329-9290.

Zhou:2019:SES

[ZHZ19] M. Zhou, M. Huang, and X. Zhu. Story ending selection by finding hints from pairwise candidate endings. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(4): 719–729, April 2019. CO-DEN ???? ISSN 2329-

Zivanovic:2015:HBC

[Ziv15]  M. Zivanović.  Har-

monic bandwidth companding for separation of overlapping harmonics in pitched signals. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 23(5): 898–908, May 2015. CO-DEN ???? ISSN 2329-9290.

Zhou:2015:SCB

[ZJD^{+}15] Pan Zhou, Hui Jiang, Li-Rong Dai, Yu Hu, and Qing-Feng Liu. Stateclustering based multiple deep neural networks modeling approach for speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (4):631–642, April 2015. CODEN ???? ISSN 2329-9290.

Zhou:2021:RED

[ZJL21] Mengjia Zhou, Donghong Ji, and Fei Li. Relation extraction in dialogues: a deep learning model based on the generality and specialty of dialogue text. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2015–2026, 2021. CODEN ???? ISSN 2329-9290.

Zhang:2015:EDF

[ZKA15] Mengqiu Zhang, R. A.

Kennedy, and T. D. Abhayapala. Empirical determination of frequency representation in spherical harmonics-based HRTF functional modeling. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(2):351–360, February 2015. CODEN ???? ISSN 2329-9290.

Zhao:2018:SSS

[ZKDS18] Y. Zhao, M. Kuruvilla-Dugdale, and M. Song. Structured sparse spectral transforms and structural measures for voice conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26(12):2267– 2276, December 2018. CODEN ???? ISSN 2329-9290.

Zhang:2018:TIS

[ZKH18] C. Zhang, K. Koishida, and J. H. L. Hansen. Textindependent speaker verification based on triplet convolutional neural network embeddings. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 26(9): 1633–1644, September 2018. CODEN ???? ISSN 2329-9290.

Zhao:2022:WRA

[ZKKC22] Yuting Zhao, Mamoru Komachi, Tomoyuki Kajiwara, and Chenhui Chu. Word-region alignmentguided multimodal neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):244–259, 2022. CO-DEN ???? ISSN 2329-9290.

Zhang:2021:QAB

[ZL21a] Jie Zhang and Changheng Li. Quantization-aware binaural MWF based noise reduction incorporating external wireless devices. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3118– 3131, 2021. CODEN ???? ISSN 2329-9290.

Zhang:2021:EPW

[ZL21b] Ya-Jie Zhang and Zhen-Hua Ling. Extracting and predicting word-level style variations for speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1582– 1593, 2021. CODEN ???? ISSN 2329-9290.

Zhou:2024:CCI

[ZLC^{+}24] Jie Zhou, Yuanbiao Lin, Qin Chen, Qi Zhang, Xuanjing Huang, and Liang He. CausalABSC: Causal inference for aspect debiasing in aspectbased sentiment classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):830– 840, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3340606.

Zhang:2020:NPS

[ZLD20] Jing-Xuan Zhang, Zhen-Hua Ling, and Li-Rong Dai. Non-parallel sequenceto-sequence voice conversion with disentangled linguistic and speaker representations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):540– 552, January 2020. CO-DEN ???? ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2019.2960721.

Zhou:2021:USS

[ZLD21] Xiao Zhou, Zhen-Hua Ling, and Li-Rong Dai. UnitNet: a sequence-to-sequence acoustic model for concatenative speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2643– 2655, 2021. CODEN ???? ISSN 2329-9290.

Zhao:2019:CNN

[ZLF19] Z. Zhao, H. Liu, and T. Fingscheidt. Convolutional neural networks to enhance coded speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (4):663–678, April 2019. CODEN ???? ISSN 2329-9290.

Zhou:2022:MCI

[ZLG^{+}22] Ying Zhou, Xuefeng Liang, Yu Gu, Yifei Yin, and Longshan Yao. Multi-classifier interactive learning for ambiguous speech emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):695–705, 2022. CODEN ???? ISSN 2329-9290.

Zhang:2021:DSH

[ZLHL21] Chen Zhang, Grandee Lee, Luis Fernando D. Haro, and Haizhou Li. D-Score: Holistic dialogue evaluation without reference. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2502–2516, 2021. CODEN ???? ISSN 2329-9290.

Zhang:2017:NNS

[ZLJ^{+}17] Shiliang Zhang, Cong Liu, Hui Jiang, Si Wei, Lirong Dai, and Yu Hu. Nonrecurrent neural structure for long-term dependence. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (4):871–884, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/7859324/.

Zheng:2022:UCE

[ZLJ22] Kengtao Zheng, Nankai Lin, and Shengyi Jiang. Unsupervised character embedding correction and candidate word denoising. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):76–86, 2022. CODEN ???? ISSN 2329-9290.

Zhang:2023:RVQ

[ZLJB23] Jianwei Zhang, Julie Liss, Suren Jayasuriya, and Visar Berisha. Robust vocal quality feature embeddings for dysphonic voice detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1348– 1359, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3261753.

Zhang:2019:SSA

[ZLL^{+}19] J. Zhang, Z. Ling, L. Liu, Y. Jiang, and L. Dai. Sequence-to-sequence acoustic modeling for voice conversion. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(3):631– 644, March 2019. CODEN ???? ISSN 2329-9290.

Zhang:2023:ISSa

[ZLL^{+}23] Yuxiang Zhang, Zhuo Li, Jingze Lu, Hua Hua, Wenchao Wang, and Pengyuan Zhang. The impact of silence on speech anti-spoofing. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 3374–3389, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3306711.

Zhou:2024:CMI

[ZLL^{+}24a] Dong Zhou, Fang Lei, Lin Li, Yongmei Zhou, and Aimin Yang. Cross-modal interaction via reinforcement feedback for audiolyrics retrieval. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 1248–1260, 2024. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2024.3358048.

Zhu:2024:MME

[ZLL^{+}24b] Xinfa Zhu, Yi Lei, Tao Li, Yongmao Zhang, Hongbin Zhou, Heng Lu, and Lei Xie. METTS: Multilingual emotional textto-speech by cross-speaker and cross-lingual emotion transfer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1506– 1518, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3363444.

Zhang:2016:LRF

[ZLLS16] Meng Zhang, Yang Liu, Huanbo Luan, and Maosong Sun. Listwise ranking functions for statistical machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (8):1464–1472, ???? 2016. CODEN ???? ISSN 2329-9290.

Zhang:2023:NMC

[ZLLZ23] Chenggang Zhang, Jinjiang Liu, Hao Li, and Xueliang Zhang. Neural multi-channel and multimicrophone acoustic echo cancellation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2181– 2192, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3282103.

Zeghidour:2022:SEE

[ZLO^{+}22] Neil Zeghidour, Alejandro Luebs, Ahmed Omran, Jan Skoglund, and Marco Tagliasacchi. SoundStream: an end-to-end neural audio codec. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):495–507, 2022. CO-DEN ???? ISSN 2329-9290.

Zhang:2021:NMT

[ZLS^{+}21] Jiacheng Zhang, Huanbo Luan, Maosong Sun, Feifei Zhai, Jingfang Xu, and Yang Liu. Neural machine translation with explicit phrase alignment. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1001–1010, 2021. CODEN ???? ISSN 2329-9290.

Zhu:2023:NVT

[ZLSC23] Wenzhao Zhu, Lei Luo, Jinwei Sun, and Mads Græsbøll Christensen. A new virtual tracking subalgorithm based hybrid active control system for narrowband noise with impulsive interference. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):915– 926, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3240645.

Zhang:2022:AKL

[ZLW^{+}22] Yi Zhang, Lei Li, Yunfang Wu, Qi Su, and Xu Sun. Alleviating the knowledge-language inconsistency: a study for deep commonsense knowledge. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):594–604, 2022. CODEN ???? ISSN 2329-9290.

Zhao:2024:JQJ

[ZLW^{+}24] Yun Zhao, Dexi Liu, Changxuan Wan, Xiping Liu, Jian yun Nie, and Jiaming Liu. JMS-QA: a joint hierarchical architecture for mental health question answering. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):352– 363, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3329295.

Zhang:2024:ERF

[ZLX^{+}24] Qiangqiang Zhang, Dongyuan Lin, Yingying Xiao, Yunfei Zheng, and Shiyuan Wang. Error reused filtered-X least mean square algorithm for active noise control. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 397–412, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3330077.

Zhang:2021:MTD

[ZLZ21] Zhuosheng Zhang, Junlong Li, and Hai Zhao. Multi-turn dialogue reading comprehension with pivot turns and knowledge. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1161–1173, 2021. CODEN ???? ISSN 2329-9290.

Zhang:2024:QDR

[ZLZ^{+}24] Geng Zhang, Jin Liu, Guangyou Zhou, Kunsong Zhao, Zhiwen Xie, and Bo Huang. Questiondirected reasoning with relation-aware graph attention network for complex question answering over knowledge graph. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32 (??):1915–1927, 2024. ISSN 2329-9290. URL

                                        https://dl.acm.org/doi/
                                        10.1109/TASLP.2024.3375631.

Zhang:2021:FFS

[ZLZW21] Yun Zhang, Yongguo Liu, Jiajing Zhu, and Xindong Wu. FSPRM: a feature subsequence based probability representation model for Chinese word embedding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1702– 1716, 2021. CODEN ???? ISSN 2329-9290.

Zhao:2017:TAD

[ZM17] Rui Zhao and Kezhi Mao. Topic-aware deep compositional models for sentence classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (2):248–260, ???? 2017. CODEN ???? ISSN 2329-9290.

Zohourian:2020:BDR

[ZM20] Mehdi Zohourian and Rainer Martin. Binaural direct-to-reverberant energy ratio and speaker distance estimation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 92–104, January 2020. CODEN ???? ISSN 2329-9290. URL https:

//dl.acm.org/doi/abs/
10.1109/TASLP.2019.2948730.

Zalkow:2021:CBL

[ZM21] Frank Zalkow and Meinard Müller. CTC-Based learning of chroma features for score audio music retrieval. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2957– 2971, 2021. CODEN ???? ISSN 2329-9290.

Zhu:2023:BSA

[ZM23] Yingke Zhu and Brian Mak. Bayesian selfattentive speaker embeddings for text-independent speaker verification. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1000–1012, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3244502.

Zhang:2024:CQE

[ZMC^{+}24] Ying Zhang, Fandong Meng, Yufeng Chen, Jinan Xu, and Jie Zhou. Complex question enhanced transfer learning for zero-shot joint information extraction. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 261–275, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3304481.

Zhang:2021:RAC

[ZMX^{+}21] Licheng Zhang, Zhendong Mao, Benfeng Xu, Quan Wang, and Yongdong Zhang. Review and arrange: Curriculum learning for natural language understanding. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3307–3320, 2021.

CODEN ???? ISSN 2329-9290.

Zhao:2024:DTR

[ZMZ24] Jiahao Zhao, Wenji Mao, and Daniel Dajun Zeng. Disentangled text representation learning with information-theoretic perspective for adversarial robustness. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1237– 1247, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3358052.

Zakharov:2014:SWR

[ZN14] Y. Zakharov and V. H.

Nascimento. Slidingwindow RLS low-cost implementation of proportionate affine projection algorithms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(12):1815– 1824, December 2014. CODEN ???? ISSN 2329-9290.

Zahoransky:2015:TSS

[ZP15] D. Zahoransky and I. Polasek. Text search of surnames in some Slavic and other morphologically rich languages using rule based phonetic algorithms. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(3):553– 563, March 2015. CODEN ???? ISSN 2329-9290.

Zhang:2019:SSC

[ZPD19] Y. Zhang, B. Pardo, and Z. Duan. Siamese style convolutional neural networks for sound search by vocal imitation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(2): 429–441, February 2019. CODEN ???? ISSN 2329-9290.

Zahedi:2021:MPB

[ZPØ^{+}21] Adel Zahedi, Michael Syskind Pedersen, Jan Østergaard, Thomas Ulrich Christiansen, Lars Bramsløw, and Jesper Jensen. Minimum processing beamforming. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2710– 2724, 2021. CODEN ???? ISSN 2329-9290.

Zohrer:2015:RLS

[ZPP15] M. Zohrer, R. Peharz, and F. Pernkopf. Representation learning for singlechannel source separation and bandwidth extension. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (12):2398–2409, December 2015. CODEN ???? ISSN 2329-9290.

Zhang:2023:LLA

[ZPW23] Hao Zhang, Ashutosh Pandey, and De Liang Wang. Low-latency active noise control using attentive recurrent network. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1114– 1123, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3244528.

Zhu:2024:BCI

[ZQF^{+}24] Tiantian Zhu, Yang Qin, Ming Feng, Qingcai Chen, Baotian Hu, and Yang Xiang. BioPRO: Contextinfused prompt learning for biomedical entity linking. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):374– 385, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3331149.

Zhang:2022:EFR

[ZQL22] Mi Zhang, Tieyun Qian, and Bing Liu. Exploit feature and relation hierarchy for relation extraction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):917–930, 2022. CODEN ???? ISSN 2329-

Zhang:2023:TFA

[ZQN^{+}23] Qiquan Zhang, Xinyuan Qian, Zhaoheng Ni, Aaron Nicolson, Eliathamby Ambikairajah, and Haizhou Li. A time-frequency attention module for neural speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):462–475, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2022.3225649.

Zhang:2023:ITR

[ZQZ^{+}23] Guangyan Zhang, Ying Qin, Wenjie Zhang, Jialun Wu, Mei Li, Yutao Gai, Feijun Jiang, and Tan Lee. iEmoTTS: Toward robust cross-speaker emotion transfer and control for speech synthesis based on disentanglement between prosody and timbre. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1693– 1705, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3268571.

Zhao:2019:SLD

[ZQZH19] L. Zhao, X. Qiu, Q. Zhang, and X. Huang. Sequence labeling with deep gated dual path CNN. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (12):2326–2335, December 2019. CODEN ???? ISSN 2329-9290.

Zhang:2020:TSD

[ZS20] Weijian Zhang and Peng Song. Transfer sparse discriminant subspace learning for cross-corpus speech emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):307–318, January 2020. CODEN ???? ISSN 2329-9290.

URL https://dl.acm.
org/doi/abs/10.1109/
TASLP.2019.2955252.

Zeinali:2017:HBP

[ZSB17] Hossein Zeinali, Hossein Sameti, and Lukáš Burget. HMM-based phrase-independent ivector extractor for textdependent speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(7):1421– 1435, ???? 2017. CO-DEN ???? ISSN 2329-9290. URL http://

ieeexplore.ieee.org/
document/7902120/.

Zhang:2023:IST

[ZSC^{+}23a] Hao Zhang, Nianwen Si, Yaqi Chen, Wenlin Zhang, Xukui Yang, Dan Qu, and Wei-Qiang Zhang.

Improving speech translation by cross-modal multi-grained contrastive learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1075– 1086, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3244521.

Zhang:2023:CBS

[ZSC^{+}23b] Junqing Zhang, Liming Shi, Mads Græsbøll Christensen, Wen Zhang, Lijun Zhang, and Jingdong Chen. CGMMbased sound zone generation using robust pressure matching with ATF perturbation constraints. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):3331–3345, 2023. ISSN 2329-9290. URL

https://dl.acm.org/doi/
10.1109/TASLP.2023.3306712.

Zhang:2020:PTP

[ZSH20] Liwen Zhang, Ziqiang Shi, and Jiqing Han. Pyramidal temporal pooling with discriminative mapping for audio classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 28(??):770–784, January 2020. CODEN ???? ISSN 2329-

9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2020.2966868.

Zorila:2016:NFF

[ZSIA16] Tudor-Cătălin Zorilă, Yannis Stylianou, Tatsuma Ishihara, and Masami Akamine. Near and far field speech-in-noise intelligibility improvements based on a time–frequency energy reallocation approach. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1808– 1818, ???? 2016. CODEN ???? ISSN 2329-9290.

Zhen:2022:SEN

[ZSL^{+}22] Kai Zhen, Jongmo Sung, Mi Suk Lee, Seungkwon Beack, and Minje Kim. Scalable and efficient neural speech coding: a hybrid design. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):12–25, 2022. CODEN ???? ISSN 2329-9290.

Zorrilla:2023:AEA

[ZTC23] Asier López Zorrilla, María Inés Torres, and Heriberto Cuayáhuitl. Audio embedding-aware dialogue policy learning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31 (??):525–538, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2022.3225658.

Zhang:2023:EES

[ZTDD23a] Jie Zhang, Rui Tao, Jun Du, and Li-Rong Dai. Energy-efficient sparsitydriven speech enhancement in wireless acoustic sensor networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 215–228, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2022.3221013.

Zhang:2023:SSS

[ZTDD23b] Jie Zhang, Rui Tao, Jun Du, and Li-Rong Dai. SDW-SWF: Speech distortion weighted singlechannel Wiener filter for noise reduction. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 3176–3189, 2023. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2023.3304474.

Zhou:2021:LAS

[ZTL21] Yi Zhou, Xiaohai Tian, and Haizhou Li. Language agnostic speaker embedding for cross-lingual personalized speech generation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3427–3439, 2021.

CODEN ???? ISSN 2329-9290.

Zhu:2023:CDF

[ZTM^{+}23] Yi Zhu, Abhishek Tiwari, João Monteiro, Shruti Kshirsagar, and Tiago Henrique Falk. COVID-19 detection via fusion of modulation spectrum and linear prediction speech features. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 1536–1549, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3265603.

Zhang:2019:DTC

[ZTQB19] G. Zhang, J. Tao, X. Qiu, and I. Burnett. Decentralized two-channel active noise control for single frequency by shaping matrix eigenvalues. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (1):44–52, January 2019. CODEN ???? ISSN 2329-9290.

Zheng:2019:FBD

[ZTWY19] Y. Zheng, J. Tao, Z. Wen, and J. Yi. Forward backward decoding sequence for regularizing end-to-end TTS. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2067– 2079, December 2019.

CODEN ???? ISSN 2329-9290.

Zhang:2016:BCI

[ZW16a] Xiao-Lei Zhang and DeLiang Wang. Boosting contextual information for deep neural network based voice activity detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (2):252–264, ???? 2016. CODEN ???? ISSN 2329-9290.

Zhang:2016:DEL

[ZW16b] Xiao-Lei Zhang and DeLiang Wang. A deep ensemble learning method for monaural speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(5):967–977, May 2016. CODEN ???? ISSN 2329-9290.

Zhang:2017:DLB

[ZW17] Xueliang Zhang and DeLiang Wang. Deep learning based binaural speech separation in reverberant environments. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25 (5):1075–1084, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

                                        /ieeexplore.ieee.org/
                                        document/7886357/.

Zhang:2019:CLF

[ZW19] T. Zhang and J. Wu.

Constrained learned feature extraction for acoustic scene classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27 (8):1216–1228, August 2019. CODEN ???? ISSN 2329-9290.

Zhang:2023:PDC

[ZWC^{+}23] Lin Zhang, Xin Wang, Erica Cooper, Nicholas Evans, and Junichi Yamagishi. The Partial-Spoof database and countermeasures for the detection of short fake speech segments embedded in an utterance. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 813–825, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3233236.

Zeng:2015:GBL

[ZWCT15] Xiaodong Zeng, D. F.

Wong, L. S. Chao, and I. Trancoso. Graph-based lexicon regularization for PCFG with latent annotations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(3):441–450, March 2015. CODEN ???? ISSN 2329-9290.

Zhao:2023:LCC

[ZWHD23] Fei Zhao, Zhen Wu, Liang He, and Xin-Yu Dai. Label-correction capsule network for hierarchical text classification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2158– 2168, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3282099.

Zhang:2023:OQC

[ZWL^{+}23a] Cuilian Zhang, Derek F. Wong, Eddy S. K. Lei, Runzhe Zhan, and Lidia S. Chao. Obscurity-quantified curriculum learning for machine translation evaluation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2259– 2271, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3282105.

Zhang:2023:MGD

[ZWL^{+}23b] Hanyi Zhang, Longbiao Wang, Kong Aik Lee, Meng Liu, Jianwu Dang, and Helen Meng. Metageneralization for domaininvariant speaker verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1024– 1036, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3244518.

Zhao:2023:TTB

[ZWS^{+}23] Zhengdao Zhao, Yuhua Wang, Guang Shen, Yuezhu Xu, and Jiayuan Zhang. TDFNet: Transformer-based deepscale fusion network for multimodal emotion recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3771– 3782, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3316458.

Zhou:2023:OCL

[ZWTL23] Yi Zhou, Zhizheng Wu, Xiaohai Tian, and Haizhou Li. Optimization of cross-lingual voice conversion with linguistics losses to reduce foreign accents. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1916– 1926, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3271107.

Zhao:2014:RSI

[ZWW14] Xiaojia Zhao, Yuxuan Wang, and Deliang Wang. Robust speaker identification in noisy and reverberant conditions. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(4): 836–845, April 2014. CO-DEN ???? ISSN 2329-9290.

Zhao:2015:CSI

[ZWW15] Xiaojia Zhao, Yuxuan Wang, and Deliang Wang. Cochannel speaker identification in anechoic and reverberant conditions. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (11):1727–1736, November 2015. CODEN ???? ISSN 2329-9290.

Zhao:2019:TSD

[ZWW19] Y. Zhao, Z. Wang, and D. Wang. Two-stage deep learning for noisyreverberant speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(1):53–62, January 2019. CODEN ???? ISSN 2329-9290.

Zhang:2023:IFT

[ZWW23a] Yixuan Zhang, Heming Wang, and DeLiang Wang. F 0 estimation and voicing detection with cascade architecture in noisy speech. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3760– 3770, 2023. ISSN 2329-

                                        9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2023.3313427.

Zheng:2023:EIC

[ZWW^{+}23b] Qinghua Zheng, Yuefei Wu, Guangtao Wang, Yanping Chen, Wei Wu, Zai Zhang, Bin Shi, and Bo Dong. Exploring interactive and contrastive relations for nested named entity recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2899– 2909, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3293047.

Zhou:2015:COT

[ZWX15] Xinjie Zhou, Xiaojun Wan, and Jianguo Xiao. CLOpinionMiner: Opinion target extraction in a cross-language scenario. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23 (4):619–630, April 2015. CODEN ???? ISSN 2329-9290.

Zhang:2021:PPE

[ZWZ^{+}21] Lu Zhang, Mingjiang Wang, Qiquan Zhang, Xinsheng Wang, and Ming Liu. PhaseDCN: a phase-enhanced dualpath dilated convolutional network for singlechannel speech enhancement. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2561–2574, 2021. CODEN ???? ISSN 2329-9290.

Zhou:2016:LMT

[ZXH^{+}16] Guangyou Zhou, Zhiwen Xie, Tingting He, Jun Zhao, and Xiaohua Tony Hu. Learning the multilingual translation representations for question retrieval in community question answering via non-negative matrix factorization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(7):1305– 1314, July 2016. CODEN ???? ISSN 2329-9290.

Zong:2021:NLE

[ZXN^{+}21] Tianrui Zong, Yong Xiang, Iynkaran Natgunanathan, Longxiang Gao, Guang Hua, and Wanlei Zhou. Non-linear-echo based anti-collusion mechanism for audio signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):969–984, 2021. CO-DEN ???? ISSN 2329-9290.

Zhang:2017:CAR

[ZXSD17] Biao Zhang, Deyi Xiong, Jinsong Su, and Hong Duan. A context-aware recurrent encoder for neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 25(12):2424–2432, ???? 2017. CODEN ???? ISSN 2329-9290. URL http:/

/ieeexplore.ieee.org/
document/8031316/.

Zhang:2019:FAK

[ZXSL19] B. Zhang, D. Xiong, J. Su, and J. Luo. Future-aware knowledge distillation for neural machine translation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(12):2278–2287, December 2019. CODEN ???? ISSN 2329-9290.

Zhang:2021:MCM

[ZXY^{+}21] Zhuohuang Zhang, Yong Xu, Meng Yu, Shi-Xiong Zhang, Lianwu Chen, Donald S. Williamson, and Dong Yu. Multichannel multi-frame ADL-MVDR for target speech separation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3526– 3540, 2021. CODEN ???? ISSN 2329-9290.

Zhang:2023:LDR

[ZXZZ23] Hanlei Zhang, Hua Xu, Shaojie Zhao, and Qianrui Zhou. Learning discriminative representations and decision boundaries for open intent detection. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1611– 1623, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3265203.

Zhang:2021:CIE

[ZYD^{+}21] Ningyu Zhang, Hongbin Ye, Shumin Deng, Chuanqi Tan, Mosha Chen, Songfang Huang, Fei Huang, and Huajun Chen. Contrastive information extraction with generative transformer. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3077– 3088, 2021. CODEN ???? ISSN 2329-9290.

Zhang:2023:AAC

[ZYD^{+}23] Yiming Zhang, Hong Yu, Ruoyi Du, Zheng-Hua Tan, Wenwu Wang, Zhanyu Ma, and Yuan Dong. ACTUAL: Audio captioning with caption feature space regularization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2643– 2657, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3293015.

Zhang:2018:SEN

[ZYF18] M. Zhang, N. Yu, and G. Fu. A simple and effective neural model for joint word segmentation and POS tagging. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 26 (9):1528–1538, September 2018. CODEN ???? ISSN 2329-9290.

Zhao:2014:MPA

[ZYG^{+}14] Haiquan Zhao, Yi Yu, Shibin Gao, Xiangping Zeng, and Zhengyou He. Memory proportionate APA with individual activation factors for acoustic echo cancellation. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 22(6): 1047–1055, June 2014. CODEN ???? ISSN 2329-9290.

Zeng:2021:DAM

[ZYL^{+}21] Jiali Zeng, Yongjing Yin, Yang Liu, Yubin Ge, and Jinsong Su. Domain adaptive meta-learning for dialogue state tracking. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2493–2501, 2021. CODEN ???? ISSN 2329-9290.

Zhou:2020:JSS

[ZYW^{+}20] Qingyu Zhou, Nan Yang, Furu Wei, Shaohan Huang, Ming Zhou, and Tiejun Zhao. A joint sentence scoring and selection framework for neural extractive document summarization. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 28(??): 671–681, January 2020. CODEN ???? ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2020.2964427.

Zhao:2024:DDN

[ZYW^{+}24] Yangyang Zhao, Kai Yin, Zhenyu Wang, Mehdi Dastani, and Shihan Wang. Decomposed deep Q-network for coherent task-oriented dialogue policy learning. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 32(??): 1380–1391, 2024. ISSN 2329-9290. URL https:

                                        //dl.acm.org/doi/10.
                                        1109/TASLP.2024.3357038.

Zhang:2022:WAK

[ZYZU22] Zhuosheng Zhang, Haojie Yu, Hai Zhao, and Masao Utiyama. Which apple keeps which doctor away? Colorful word representations with visual oracles. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):49–59, 2022. CODEN ???? ISSN 2329-9290.

Zheng:2019:PAS

[ZZ19] N. Zheng and X. Zhang. Phase-aware speech enhancement based on deep neural networks. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 27(1): 63–76, January 2019. CO-DEN ???? ISSN 2329-9290.

Zhang:2022:PUB

[ZZ22] Jie Zhang and Guanghui Zhang. A parametric unconstrained beamformer based binaural noise reduction for assistive hearing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):292–304, 2022. CODEN ???? ISSN 2329-9290.

Zhao:2023:MGE

[ZZD23] Yilin Zhao, Hai Zhao, and Sufeng Duan. Multigrained evidence inference for multi-choice reading comprehension. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 3896–3907, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3313885.

Zhu:2021:KEF

[ZZGD21] Biru Zhu, Xingyao Zhang, Ming Gu, and Yangdong Deng. Knowledge enhanced fact checking and verification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):3132– 3143, 2021. CODEN ???? ISSN 2329-9290.

Zhang:2016:RVS

[ZZH16] Sheng Zhang, Jiashu Zhang, and Hongyu Han. Robust variable step-size decorrelation normalized least-mean-square algorithm and its application to acoustic echo cancellation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(12):2368–2376, ???? 2016. CODEN ???? ISSN 2329-9290.

Zhou:2021:GRG

[ZZH21] Yi Zhou, Xiaoqing Zheng, and Xuanjing Huang. Generating responses with a given syntactic pattern in Chinese dialogues. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):2888–2898, 2021. CODEN ???? ISSN 2329-9290.

Zeng:2023:ESR

[ZZH23] Jiandian Zeng, Jiantao Zhou, and Caishi Huang. Exploring semantic relations for social media sentiment analysis. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):2382–2394, 2023. ISSN 2329-

                                        9290. URL https://dl.

acm.org/doi/10.1109/
TASLP.2023.3285238.

Zeng:2023:DLR

[ZZJ^{+}23] Daojian Zeng, Chao Zhao, Chao Jiang, Jianling Zhu, and Jianhua Dai. Document-level relation extraction with context guided mention integration and inter-pair reasoning. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3659– 3666, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3316454.

Zhang:2019:ESS

[ZZL^{+}19] Z. Zhang, H. Zhao, K. Ling, J. Li, Z. Li, S. He, and G. Fu. Effective subword segmentation for text comprehension. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(11):1664–1674, November 2019. CODEN ???? ISSN 2329-9290.

Zhong:2023:SSS

[ZZL^{+}23] Jiaxin Zhong, Tao Zhuang, Mengtong Li, Ray Kirby, Mahmoud Karimi, Jing Lu, and Dong Zhang. Sidelobe suppression for a steerable parametric source using the sparse random array technique. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3152– 3161, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3304491.

Zhang:2019:ASR

[ZZLZ19] J. Zhang, Y. Zhao, H. Li, and C. Zong. Attention with sparsity regularization for neural machine translation and summarization. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(3):507– 518, March 2019. CODEN ???? ISSN 2329-9290.

Zhang:2016:PAU

[ZZN^{+}16] Xueliang Zhang, Hui Zhang, Shuai Nie, Guanglai Gao, and Wenju Liu. A pairwise algorithm using the deep stacking network for speech separation and pitch estimation. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24 (6):1066–1078, June 2016. CODEN ???? ISSN 2329-9290.

Zhang:2023:PMP

[ZZS^{+}23] Ya-Jie Zhang, Chao Zhang, Wei Song, Zhengchen Zhang, Youzheng Wu, and Xiaodong He. Prosody modelling with pre-trained cross-utterance representations for improved speech synthesis. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 31(??): 2812–2823, 2023. ISSN 2329-9290. URL https:

//dl.acm.org/doi/10.
1109/TASLP.2023.3278184.

Zhao:2021:DAR

[ZZX^{+}21] Juan Zhao, Tianrui Zong, Yong Xiang, Longxiang Gao, Wanlei Zhou, and Gleb Beliakov. Desynchronization attacks resilient watermarking method based on frequency singular value coefficient modification. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):2282– 2295, 2021. CODEN ???? ISSN 2329-9290.

Zhao:2023:SSB

[ZZX^{+}23a] Juan Zhao, Tianrui Zong, Yong Xiang, Longxiang Gao, Guang Hua, Keshav Sood, and Yushu Zhang. SSVS-SSVD based desynchronization attacks resilient watermarking method for stereo signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):448– 461, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2022.3225668.

Zhou:2023:DND

[ZZX^{+}23b] Aolong Zhou, Wen Zhang, Guojun Xu, Xiaoyong Li, Kefeng Deng, and Junqiang Song. DBSA-Net: Dual branch selfattention network for underwater acoustic signal denoising. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1851– 1865, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3275030.

Zhang:2022:HLH

[ZZXH22] Yi Zhang, Guangyou Zhou, Zhiwen Xie, and Jimmy Xiangji Huang. HGEN: Learning hierarchical heterogeneous graph encoding for math word problem solving. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):816–828, 2022. CO-DEN ???? ISSN 2329-9290.

Zhang:2019:TIN

[ZZY19] Y. Zhang, P. Zhang, and Y. Yan. Tailoring an interpretable neural language model. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(7):1164– 1178, July 2019. CODEN ???? ISSN 2329-9290.

Zhang:2024:ISSb

[ZZY^{+}24] Puning Zhang, Rongjian Zhao, Boran Yang, Yuexian Li, and Zhigang Yang. Integrated syntactic and semantic tree for targeted sentiment classification using dual-channel graph convolutional network. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1109–1124, 2024. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2024.3350877.

Zhang:2016:ACL

[ZZZ16] Jiajun Zhang, Yu Zhou, and Chengqing Zong. Abstractive cross-language summarization via translation model enhanced predicate argument structure fusing. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 24(10):1842– 1853, ???? 2016. CODEN ???? ISSN 2329-9290.

Zhang:2021:SAN

[ZZZ^{+}21] Junqing Zhang, Wen Zhang, Jihui Aimee Zhang, Thushara Dheemantha Abhayapala, and Lijun Zhang. Spatial active noise control in rooms using higher order sources. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29 (??):3577–3591, 2021. CODEN ???? ISSN 2329-

Zhang:2022:SAM

[ZZZ22a] Zhuosheng Zhang, Yiqing Zhang, and Hai Zhao. Syntax-aware multi-spans generation for reading comprehension. IEEE/ ACM Transactions on Audio, Speech, and Language Processing, 30(??): 260–268, 2022. CODEN ???? ISSN 2329-9290.

Zhao:2022:RKN

[ZZZ22b] Yilin Zhao, Zhuosheng Zhang, and Hai Zhao. Reference knowledgeable network for machine reading comprehension. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30 (??):1461–1473, 2022. ISSN 2329-9290. URL

                                        https://dl.acm.org/doi/
                                        10.1109/TASLP.2022.3164219.

Zhou:2024:ATS

[ZZZ^{+}24] Xuehao Zhou, Mingyang Zhang, Yi Zhou, Zhizheng Wu, and Haizhou Li. Accented text-to-speech synthesis with limited data. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 32(??):1699–1711, 2024. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2024.3363414.

Zhu:2023:JSE

[ZZZD23] Qiu-Shi Zhu, Jie Zhang, Zi-Qiang Zhang, and Li-Rong Dai. A joint speech enhancement and self-supervised representation learning framework for noise-robust speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):1927– 1939, 2023. ISSN 2329-

9290. URL https://dl.
acm.org/doi/10.1109/
TASLP.2023.3275033.

Zhang:2021:TLS

[ZZZL21] Mingyang Zhang, Yi Zhou, Li Zhao, and Haizhou Li. Transfer learning from speech synthesis to voice conversion with non-parallel training data. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 29(??):1290–1302, 2021. CODEN ???? ISSN 2329-9290.

Zhu:2022:DRC

[ZZZL22] Pengfei Zhu, Zhuosheng Zhang, Hai Zhao, and Xiaoguang Li. DUMA: Reading comprehension with transposition thinking. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 30(??):269–279, 2022. CODEN ???? ISSN 2329-9290.

Zhou:2023:LPC

[ZZZW23] Kun Zhou, Yuanhang Zhou, Wayne Xin Zhao, and Ji-Rong Wen. Learning to perturb for contrastive learning of unsupervised sentence representations. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 31(??):3935– 3944, 2023. ISSN 2329-

                                        9290. URL https://dl.
                                        acm.org/doi/10.1109/
                                        TASLP.2023.3304485.